A Complete Bibliography of Publications in the ICES Journal of Marine Science (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.06

Title word cross-reference

< [BGG^{+}06]. ^{18} [VH08]. _{2} [HMMB^{+}08]. a [GAP^{+}00, GHC09, VM09]. ≈ [SQN08]. δ [VH08]. log_{10} [MMC03].

-related [HMMB^{+}08].

/ICES [RDF^{+}03].

10-year [íJCMR07, VM09]. 120-kHz [GJH^{+}09]. 17th [LLD^{+}05]. 18th [LLD^{+}05]. 1940s [EKPT07]. 1956/1957 [CDD^{+}07]. 1976/1977 [Lit06]. 1979 [DBDA^{+}02]. 1980s [AMD^{+}05, GF01, Tan00]. 1989 [Hel00]. 1990s [ADO02, BR04, Tan00]. 1995 [MLNC01]. 1995/1996 [NRR^{+}09]. 1998 [AJ00].

2-m [RKE06]. 2-year [BVB^{+}07]. 20 [BGG^{+}06]. 200-kHz [TSK03]. 2000 [BI08, PFK^{+}09]. 2000s [PHDC^{+}09]. 2001 [Chr02]. 2005/2006 [CDD^{+}07].

2006 [DPW07]. 2007 [MPD^{+}08, NM09]. 2008 [DKMO09]. 20^{o} [ODCN09]. 20th [RE00]. 21st [GWvM07]. 224 [HHAB09]. 23^{o} [GGV^{+}04]. 2D [LHKGS00, SB04]. 2J3KL [AG00].

300-kHz [GJH^{+}09]. 34^{o} [MVMH04]. 36^{o} [SQN08]. 38-kHz [JFCH05]. 3D [MLM02, MLMC02].

41^{o} [MVMH04]. 4D [BFZ05].

5.8S [KHS^{+}08].

66 [HHAB09, WBK09a].

9-year [SL04].

abalone [MMCD08, SMH09]. abandoned [LGH^{+}09]. ABC [YFL05]. abiotic [HHAB09, ODRN05, WBK^{+}09b, WBK09a]. absence [GC02, GHFA09, HFWB05]. absolute [RT03]. absorption [CDB05, DCM03, Rho08]. absorptivity [DLS01]. Abundance [DGC00a, KA01, AP09, AHS08, ANNG01, AO08, BRP02, BJ00b, BJ00a, BMJ08, BCL03, BBSK09, CR04, CRvCB08, Cor07, DNP03, DCN^{+}04, DLR02, DHKV01, Gaa00, GDL04, GDH02, GAM^{+}06, Gud04, HOP09, HH01, HHO08, HBST02, Kan07, Kas09, LPA^{+}00, LAO^{+}07, LME02, MG07, Mam06, MGTS00, MFB^{+}08, MKB01, NEJH05, NIF^{+}09, ÓMP^{+}04, PSO^{+}04, PLP^{+}07, Pet01, PGM01, PCS^{+}04, RAB^{+}07, RW01, Rob05, RR02a, RD01, RLdAW06, SHT^{+}09, SW06c, SB03, Tan00, TS05, UPK^{+}08, UP00, VLBB08, WHG07, WPB^{+}03, WMS^{+}03, WRF09, WvdMF06, YFL05]. abundant [AE02]. acanthias [EK08, MF07, TM09]. accommodated [MPJ07]. according [NH09]. account [CLR^{+}05, FGP07, FGP09]. Accounting [BS03, CD07, FGFP08, UASN07]. accumulated [SRS^{+}07]. accumulation [URMS04]. Accuracy [HCEM06, ES02, MMF09]. accurate [GKOV05, KFM02]. Acetylcholinesterase [PN06]. achievable [ES09]. achieved [PKH^{+}08]. achieving [Arn00, SPS00a]. acid [Cor00a, CSdQB06, JGN04]. acidification [FSFO08]. Acipenser [FSDB09]. ACME [Ric00b]. Acoustic [AGC02, AVK^{+}08, BJ00b, BJ00a, CMC^{+}06a, CMN^{+}07, CMM03, CKS03, DSJ03, EGB02, FWW06, FSQ^{+}03, GOK05, GKO07, HSA05, JO02, KRYL09, KHEJ09, MANT07, OGD09, OMTS03, Orl03, ZKP03, AAV^{+}04, BRC09, BGW03, BML^{+}05, BPWS09, BH07, CBDB02, CWC00, CWC^{+}03, DNP03, DHWW08, DBC03, ES02, ES09, EHG06, EZ03, Erm09, ETB07, FLK^{+}09, FGR04, FR09, GH04, GR05, GO03a, GO03b, GHD^{+}09, GJH^{+}09, HSR01, HS01, HOP09, HMD^{+}08, HRM04, HMHI09, HMAN03, HHO08, HSA^{+}09, HK00, JH01b, JGST09, JG07, Jør03, JB00, KMI^{+}05, Kas09, KO02, KDO^{+}08, LBF01, LKK^{+}09, LW04, MvdKN05, MMC03, MSS^{+}05, MR09, MLMC02, MCL03, Miy03, MCP03, NTJ04, NHK09, OR01, ORA02, O’D04, OM05, Orł00, Orl01, Orl05, Ost09, PB05b, Pet03, PMB^{+}03a, PGMB09]. acoustic [PR01, PR03, Rho08, RR02a, RT03, RPSSW09, RKM09, STA^{+}09, SMB09, SS09, Sim03, SGC^{+}09, SSI07, SC00, SRJ03, Tje02, War01b, WSWS03, WB02, WPJ09, WSP03, WRF09, WWWB03, YM08, ZPK05, dR01]. acoustic- [Sim03]. acoustic-abundance [DNP03, HHO08]. acoustic-frequency [Miy03]. acoustic-optical [RKM09, STA^{+}09]. acoustic-scattering [FGR04, NHK09, WSWS03]. acoustic-survey [BGW03, HOP09]. acoustic-trawl [MSS^{+}05]. acoustical [CD03]. acoustically [BH08]. Acoustics [DKMO09, MGM03, CSH00, FSB^{+}03, HDG^{+}09, Kos09, LDCH^{+}09, MFD02]. Across [Jag02, GWF01, MPD^{+}08, MM05, Mui03, PGD09, PPH09, SRS^{+}09]. Across-channel [Jag02]. action [KHN03]. active [FR09, LKL08]. active-acoustic [FR09]. activities [CMO^{+}06, EMA^{+}07, Fru02, RS03]. Activity [JDA^{+}06, BR02, Ber04, BS02, CRTS04, DH09, FS02, FLH06, GHFA09, LBL06, OMBP06, PN06, TPT^{+}09, UA04]. aculeatus [PVLP04]. acutorostrata [LHHJ^{+}09, LHR02]. Ad [Vin01]. ADAPT [BR08b]. ADAPT-VPA [BR08b]. adaptability [Gef09]. adaptation [Cor00b, OK05]. adaptive [BBS09a, CS02, Fik00, HOP09, MSS^{+}05]. ADCP [FGR04]. added [TL05]. additional [RDB09]. Additive [DLR02, BR02, MNY^{+}09]. address [GMGN06, GFP09]. Addressing [SGAC00]. adequacy [Kal01, Vec00]. Adige [SAPP04]. adjacent [HHH00, KMV^{+}07, MWF^{+}05, SKR^{+}06, WPB^{+}03]. Adriatic [MDM03, FS02, FLP^{+}02, FGLT02, MLG^{+}09, Mis02, MFA07, PAC02, PRF^{+}00, SAPP04]. adult [BPM^{+}09, GWF01, GKO07, MFB^{+}08, NTSM07, Sec00b, WC01]. advance [Mil02]. Advances [Gef09, GJR04]. advantage [MLOT09]. Advection [IWP00, BP08, GH00, GWF01, MGH08, MLNC01, SCLG00]. adverse [BHN06]. advice [BdP07, CLR^{+}05, CGV03, DRDC06, HNK07, Hoy07, ÓMP^{+}04, PR07, PR04, Ree03, RP07, Ric09, TRM07, VRP04]. Aegean [KCCM03]. aeglefinus [ATM02, BDO^{+}04, GML06, ÖG04, PKRT06, RW01, TSH^{+}06]. Aequipecten [VBSB07]. affect [BO05, ZPRJ02]. affected [Bra05, PFLFR08, SWG06]. affecting [STAN02, STJ^{+}07, WSP03]. affinis [GKFM09, Rob08]. Africa [BKN^{+}07, DBS06, ECC06, ECC08, HM04a, HL09, HR00, MLS07, OSB06, OR09, PS09, PRB^{+}07, Rob05, SIT^{+}05, WLS07a, dBMS09]. African [Mur00b, CLK^{+}09, CvdLHF08, CUUD07, DB04, ERBP09, LBF01, LDCH^{+}09]. aft [KHM09]. after [Agn08, Dek04, FBD^{+}08, NRR^{+}09, NC06, Pen07, PBLFR06, PSFY07, SBD^{+}09, UBP^{+}09, WBC^{+}08]. Afterthought [Daa05]. against [BP08, Cip09, SW06b]. Age [CAAJ07, DWC03, EDG03, FM04, KHO06, KK06a, LCC07, LHJJ^{+}01, Pel02, PS03, SRGC04, SNV^{+}09, Sim07b, SGS^{+}05, WK01, YYY^{+}02, AK04, BA03, BdP07, BWC00, BD03, BR08b, CA00, CDR05, Coo04, Cor01, CMP07, FL06, FPKH03, GML06, HLL^{+}08, HDG02, HBST02, HL09, KNKT06, KCR07, KTRG06, KGT01, LDQ08, MB01, MC09, MN02, MLOT09, MV09, NH09, PGG05, PKP07, PJ08, Pie08, Pie02, Ree03, SPFF^{+}08, SP03b, WPM^{+}09, WS06, dBMS09, dPGPB06]. age-0 [BWC00, WS06]. age-at-maturity [FPKH03]. age-at-recruitment [CDR05]. age-reading [Ree03]. age-specific [Coo04, HBST02]. age-structured [MLOT09, dBMS09]. Aggregata [GGP07]. aggregate [Des00]. aggregated [AS02a, GC05b]. aggregating [DPN^{+}09, JB00, D00]. aggregation [FR09, PMM^{+}09, WWWB03]. aggregations [BH07, CMM03, DBC03, NCM^{+}03, ZO03]. agitation [GBBG06]. agreed [dPGPB06]. agreement [RPB^{+}08]. agricultural [SP05]. Agulhas [Rob05]. agulhensis [HR00]. aid [SGMN^{+}06]. airbladder [YTS^{+}06]. airborne [CCB^{+}06, CD09]. al [HHAB09, WBK09a]. alalunga [GA05, LMVdZ^{+}07, PVH^{+}05, SA05]. Alan [AJ00]. alarms [CBDB02]. Alaska [APD09, Bro02b, DBDA^{+}02, HHMM01, HIL00, KK06a, LBNS00, Lit06, LDNS08, MM05, PSO^{+}04, SZ07, SNB^{+}02, WS06, WPF00, YM08]. Alaskan [ZK00]. albacore [GA05, LMVdZ^{+}07, PVH^{+}05, SA05]. albomaculata [RLdAW06]. algae [RMKT01]. algal [Law08, MMKR^{+}00]. Algiers [SDG^{+}08]. algorithm [GRMR07, MV09]. algorithms [CTW09]. alizarin [LTA00]. alizarin-red [LTA00]. allele [KM05]. alletteratus [BKN^{+}07]. allis [AAV^{+}04, AVJ^{+}06]. allocation [KF08, PM06, RDHP00, VEP^{+}09]. allocations [PB08b]. allochthonous [RRT00]. allometric [HSdLP06]. allow [BRP02]. allowable [GV02, dVA07]. allowing [RDB09]. allozyme [TJAS04]. Allozymic [STMM06]. ALOHA [SL04]. along [CAGV05, Cas07, CLK^{+}09, De 04, Des00, KWZ00, KTRG06, LDCH^{+}09, LSGD02, Mam06, MLS07, NGNB^{+}04, Pet03, PMB^{+}03b, SLvdB^{+}09, SLMCRM05, SSA08, TCC08, WHA08, YBF^{+}03]. Alosa [AAV^{+}04]. alpinus [BF02]. altered [CUUD07]. alternative [Aks05, Aks06, HK06, Kat05, MNCU09]. alternatively [Her04]. alterniflora [TDE08]. Alveolata [HSM09]. always [DHWW08]. America [RPSSW09, TCC08]. American [FRC03, BHMD05, GWSV08, Mor04, ON09, PCS^{+}07b, PCS^{+}07a, TCC08]. americanus [IWP00, JRCS08, RNS08, TST^{+}09]. Ammodytes [GAM^{+}06, HKD^{+}04, JWM03, LPA^{+}00, TM02]. AMOEBAs [CGV03]. among [CMDN02, Cip09, JDN01, JKSO06, PSSD08, RvMBV00, RDHP00, Roc00a, Sil03, YMF02]. Amphipoda [DBBM01]. amplitude [MS02]. anadromous [CMO^{+}06, GSN^{+}03, Gro06]. anaemia [Cip09]. analyse [BIdL^{+}08]. Analyses [HFWB05, BPM^{+}09, EDG03, FMF02, GPP09, HAvH06, MKR06]. Analysis [FT05, GDL04, GA05, Joy02, SIT^{+}05, AK04, ADC^{+}08, BHMS02, BMDBM09, BH07, BBK08, BR08b, CF05, CG07, Cam08, CA00, CMJ09, DC05a, DGMM02, DLR02, EIS05, Esm06, EN02, FB02b, FMB01, GWG06, Gud04, Har07, Hea00a, ISS^{+}07, JAC00, JMLG05, KHE^{+}09, LJM00, MLG^{+}09, Mar08, MASA06, MML^{+}00, MH01, MHD02, MLM05, ÓMP^{+}04, PR01, PH03, RPT02, RD03, SLvdB^{+}09, SH07, SBL07, Ste01a, Str05, Syr00, TSK07, TLMO08, VSC06, ZCH06]. analytic [Mye01]. Analytical [Zit01, EKPT07, RPR02]. anatomical [Aló08, HSA^{+}09]. ancestry [OCWV06]. Anchoa [BW08, SCLG00]. anchoveta [SGC^{+}09]. anchovy [BCT05, Ber00, Ber04, BDÑ04, BBGA05, BW08, CD03, DM07a, GCS^{+}04, GBBG06, GMKS06, IFUR08, KCL^{+}09, LDCH^{+}09, Miy03, NKOK00, SGMV^{+}08, STA^{+}09, SCLG00, SHS01, SPG^{+}04, ZPI^{+}09, ZWD08]. ancillary [LBF01]. Anderson [WBK09a]. Andrew [WBK09a]. angel [CLM07]. angle [BNF^{+}07, Jaf06, Jaf08, KMI^{+}05, KNS^{+}04, LN08, MW03, STA^{+}09]. Anglerfish [DSV^{+}08, RAB^{+}07, ADC^{+}08, DAH^{+}08, His01, LDQ08, MPD^{+}08, MSB04, VLBB08, WWGG02]. anglers [Aló08, BBB06b]. Anguilla [AWW^{+}07, BMC^{+}07, BD07, JHKZ09, KKS^{+}07, MYAT09, Sim07b]. angular [CLM07]. animal [ASC01, BB09]. Anisakinae [MMM00]. Anisakis [HP01, PH03, PN06]. anisotropic [SB04]. Annual [Kol06, RNWS08, Arm01, BS03, Bri02, KPK^{+}06, LFD^{+}09, MSB04, Ska07, SA03, vDBF^{+}09]. annularis [SJGRRRE02]. annulus [WWGG02]. anomalies [BO05, HR09, dPVV04]. Antarctic [CdlMA^{+}00, AF06, CW05, CdlMA^{+}00, CRC^{+}09, DC03b, DC04b, HTA09, HHKL04, LD03a, RCBM05, RCLD08, UR01, WB02]. Anthropogenic [SGMMGB09, Bjö02, WBV09]. anti [FB02b]. anti-grazing [FB02b]. Antibody [Cip09]. any [Le 09]. Aphanopus [FBMR^{+}03, MN02, MD01, QGdS04]. Apicomplexa [GGP07]. appear [Gre08]. appearance [KMT08]. Applicability [STM^{+}08, SLvdB^{+}09]. Application [ADC^{+}08, CBHM07, HS07, MNY^{+}09, ÓMP^{+}04, RO05, RPE^{+}03, dBMS09, AK04, AS02b, AWW^{+}07, Arm01, Ber00, BMC^{+}07, BBPW07, Bri02, BFZ05, BBK08, BFMJ03, EHG06, EIS05, Fle05, GBC^{+}05, HO01, HC09, HDG02, MSB04, MYAT09, MCP03, MLOT09, NGNB^{+}04, OMTS03, RW01, RGG^{+}04, RRY08, SH07, STG06, SLMCRM05, WPR^{+}07]. applications [TZ03]. Applied [HHMN01, CFMdP07, HSR01, KKC04, Vin01]. apply [Ric09]. Applying [Aks06, WCMK05]. apportionment [FB03]. appraisal [íJCMR07]. Approach [DKMO09, But07, CLL^{+}09, CFMdP07, CMGS05, DDGR07, DC01, EHG06, Eig09, FBF09, GC05a, GNC08, GFKM07, GR06, HS07, HO01, IFUR08, JR06, JR07, JRN06, JHC09, KTM^{+}05, LAB^{+}05, LBN09, LO05, MFD02, MMV^{+}08, NPPO06, PJR08, PCRW04, Ric09, Roc00a, RD01, SRN00, Tri00, WPF00, YNX^{+}05, YSF09]. approaches [GMGN06, KGRW07, Mye01, WS02a]. approaching [PHG04]. approximation [DC03a, DC03b, DC04a, DC04b]. April [Chr02]. Aquaculture [HW06, Ben01, CMO^{+}06, Den08, HBC01, JR01, LPH^{+}08, SMEK01, Zit01]. Aquatic [Jen02b, MGM03, KYG03]. Arabia [YBF^{+}03]. Arabian [GAFA06, HAvH06, HL07, KA01]. arbitrarily [OMTS03]. Archipelago [AFGR09, DHKV01, SML01, RLH01]. Arctic [ANNG01, HAN02, BF02, BFK^{+}07, Buc00, HDG02, Joh02, KMNP01, KB07, LHJJT04, ODRN05, PK09, PLJ01, SRN00, SP03b, YM00, Ynd01, Ynd06]. Arctica [KCR07]. Arcto [SBB^{+}05, SN08, VSÅF05]. Arcto-Norwegian [SBB^{+}05, SN08, VSÅF05]. Arctocephalus [MKR06]. arcuatus [DSG05]. Area [AHS08, CHB09, LPSL09, MPJ07, Ric00b, SQN08, BM01a, BM01b, CW09b, CTF02, DDR^{+}03, DDM^{+}05, DBL07, FSS00, FSQ^{+}03, FHDM00, GAM^{+}06, HJB^{+}08, Hen04, HHJK06, HAvH06, JSR06, LPM^{+}09, MAAN09, NFM^{+}02, PRvB00, PRF^{+}00, RAB^{+}07, SAPP04, SP07b, SNB^{+}02, UMSA09, VC02, VLJM^{+}07, WHG07, CBBL09, MMB09]. area-closure [GAM^{+}06]. Areas [AFM^{+}09, CW09a, ERBP09, Hal01, HKBK09, HPR09, KPD^{+}09, SEOR09, TSK00, VBF09, YSF09, AGH^{+}09, BCD^{+}02, BF02, CHB09, DTC01, GWG06, HJ03, HMDS09, HHMM01, HD00, HHH00, HSS07, Jen09, KS08, KTH^{+}00, KF08, KMJH01, LD03b, MBC^{+}09, MSGC^{+}09, RK04, SKS^{+}00, SJ08, SGY08, SSA08, SIT^{+}05, SJM03, SGAC00, SVRF08, SFM01, TCS^{+}09, WDRP09, WLK02]. Argentina [Alo01, LME05, SRM00]. Argentine [CMM03, SdlRdA06a, SdlRdA06b]. Argyrosomus [PS09, PMM^{+}09]. array [HMD^{+}08]. arrow [TJAS04]. articles [NM09]. Artificial [CTM09, Jen02a, Jen02b, LMU^{+}02, SGS02, AS02a, AS02b, BCD^{+}02, BFSC02, CSR^{+}02, CWYM^{+}02, CJM^{+}02, DGMM02, DDGJ02, FS02, FLP^{+}02, FB02b, GDL04, GC02, GAZ02, GZND02, HEGH02, HRB02, HS09, KWL^{+}02, KTS02, LSM07, Mil02, Mor02, PSC02, PAC02, RRTdA02, RRTP02, SJGRRRE02, SMG02, SW02, Sea02, wScY02, SFKC02, SP02, TB02, WS02a, WS02b, WLK02, ZNGF02]. artisanal [CBBL09, GWSV08, RSC^{+}09]. ascendance [MGH08]. ascidian [Den08]. ash [KTS02]. Asia [TYH04, ZCH06]. asiaeorientalis [XZW05]. Asian [CBDS08]. aspect [BW08, BRC09, BNF^{+}07, PHO09]. Aspects [ECC08, ERGT07, HNK07, SB03]. assemblage [De 04, FGLT02, Mor02, PPW^{+}09, SJGRRRE02, SMG02]. assemblages [BdMAL00, CLFS02, CSR^{+}02, FJSJBS^{+}08, GWF01, GH04, GAZ02, GSdFB01, HS09, JI05, KCCM03, LVHU00, MM03b, MM03a, PCDM08, PAC02, RE00, SPGT00, SF09, SCWD08, VCC07, YFL05]. assess [ARMM09, MSF^{+}06, MM05, RTB^{+}05, TCM^{+}08, dBMS09]. assessed [AGY^{+}05]. Assessing [BGAM00, CG07, FN02, JI05, LZS09, Pow00, RR02a, HNLR04, NPPO06, RL07, Vec00]. Assessment [BVB^{+}07, BFSC02, Col02, EHL07, GBT02, Mol00, PFF01, SDG^{+}08, VC02, AAV^{+}04, AFHJ04, ARMM09, ADC^{+}08, Bai09, BHH^{+}08, BNBR05, BdP07, BMC^{+}07, BKR09, Boo00, Cam08, CDB09, CJM^{+}02, CKS03, CMM01, CA02, DM07a, Dek00b, DCPvK07, EMA^{+}07, EKPT07, FS02, Fle05, GHBR08, GAM^{+}06, GRMR07, HO01, HT05, HMQ^{+}08, HOGH07, JAC00, KMNP01, Kou00, KDP09, LAB^{+}05, MMV^{+}08, MLM02, MKB01, MFA07, MLOT09, NGNB^{+}04, PCM09, RS04, Ree03, RUA07, RBGJ08, SGM09, Sim03, Sim09, SFØ07, SP07b, SKC^{+}00, Ste01b, TH08a, TH08b, TL05, TB02, TSK00, WCMK05, WWR^{+}08, WB05, WvdMF06]. assessments [BBPW07, FN00, HIL00, MPN^{+}08, RPR02]. assignment [GSS08]. associated [AFM^{+}09, BBÁMC06, CSR^{+}02, FJSJBS^{+}08, GP00, HTA09, HK06, JB00, MF07, O’D04, TCC08, ZFFT01]. associating [PMB^{+}03a]. association [HBS^{+}06]. Associations [TST^{+}09, SBT^{+}09]. assumptions [SMP09]. assurance [WMS^{+}03]. Asterias [BM01a]. At-sea [BHM^{+}04]. Atlantic [Ber04, CMC^{+}06b, GP00, Hen04, MVMH04, RMAO^{+}03, SRGC04, Sar09, SRS^{+}09, AK04, ABB^{+}08, AMJ^{+}06, AMGV06, BPM^{+}09, BMV05, BGL08, BCAN^{+}06, Bar05, BBM^{+}02, BJN^{+}06, BR02, Ben01, Ber00, BFK^{+}07, BD04, Bra07, BD02, BD03, BK07, Buc00, CH09, CA00, CMC^{+}06b, CW06, CCB^{+}06, Cas07, CPR06, Cha04, CLR^{+}05, CWC^{+}03, Cip09, CBHM07, Cla00, CCC02, CSC^{+}04, CSdQB06, DC05a, DCN^{+}04, Dem01, DLR02, DSV^{+}08, Dri05, DP03a, DP03b, DR08b, ERP01, EK08, Erm09, FCM09, FSDB09, FBMR^{+}03, FLH06, FMB01, FPKH03, FHDM00, FRC03, FCM05, FMH^{+}09, FR09, GWG06, GD05, GR01, GR02, GR05, GKOV05, GSS08, GA05, GOA^{+}09, GOK05, GKO07, GW04, GHD^{+}09, GP00, Gro06, GJH^{+}09, HMK^{+}07, HS07, HJ03, HW06, Han06, HHB^{+}00, HSS^{+}05, HSCN06, HHHH06, Hea00a, Hea00b, HBG^{+}04]. Atlantic [Hea05c, HM04b, HHH00, HAG^{+}08, HHC^{+}09, JH01a, JGN04, JNF^{+}09, JLR^{+}08, JDN01, JV05, JJ06, JKSO06, JSMK06, Joy02, KNKT06, KM05, KPS^{+}05, KCR07, Kol06, KTRG06, LMC^{+}01, LP00, LDQ08, LGH^{+}09, LMVdZ^{+}07, LD05, LME02, MG07, MCB09, MASA06, Mar07, MPN^{+}08, MFB^{+}08, MGvH06, MW03, MSS^{+}05, MR05, MSP09, MB05, MG02, MLOT09, MCI03, MSGC^{+}09, Mur00b, NHKJK09, NTJ04, NHK09, OMBP06, OCWV06, OV04, OED^{+}04, OSLO06, OGL06, OdSBS09, OAJ06, OFN02b, OFN02c, OL07, PSSD08, PGD09, PPK^{+}06, PAA06, PS03, PB00, PPH09, PCS^{+}04, PPC^{+}03, D00, RF01, RML06, RR06, Rob08, RGG^{+}04, RPK^{+}03, Ros05b, Ros09, RHBR04, RLdAW06, SJKN^{+}04, SA05, SNM05, Sar09, SCCM06, Sec02, SSKE06, Sil03, SSC^{+}06, SWG06, SW06b, SW06c, SHAH09, SMK08, SA03, SRS^{+}07, STMM06, SRMB07]. Atlantic [Ste02, SMP09, SPWHR04, SGS^{+}05, Str05, SVRF08, SPS00b, SFM01, SGMN^{+}06, TM00, TAHK06, TCSW06, TES^{+}05, TCC08, UKR05, WBC^{+}08, WBC^{+}06, WPB^{+}03, WGMM08, WPJ09, Wei05, WPM^{+}09, WC01, WBD^{+}06, YW05, ZPRJ02, vdVBMR00]. atlanticus [CB07, KH03b]. Atlantoraja [OV05]. Atmospheric [DAd02]. Atoll [AGY^{+}05]. attached [RKM09]. attempt [HMAN03]. attitudes [DOBT02]. attractants [SGS02]. attracted [RKKM06]. attraction [OSWL02]. attributable [Rho08, RSC^{+}09]. attributes [PPW^{+}09]. August [AJ00]. auratus [EB04, WM01]. Aurelia [HMHI09, KHS^{+}08, LPH^{+}08]. aurita [HMHI09, LPH^{+}08]. austral [MVMH04]. Australasian [Bun01]. australasicus [TCTC09]. Australes [AFGR09]. Australia [HHAB09, WBK09a, Bun01, Cor00a, GKFM09, HOGH07, JP03, JMWJ08, LHHF03, MS00, PPHB00, RUCG07, TJAS04, VEP^{+}09, WBK^{+}09b]. Australian [BBB06b, EH04, MM01, Ray07, SJM03, TA05, War01a]. australis [DM06, HBD05, JP03, MS04, MLS07, PM06, SRM00]. Automated [HW08, PMB^{+}03a]. automatic [MLMC02, TES^{+}05, TGS09]. Autonomous [FSB^{+}03, GEM01, PHG04]. autosub [GEM01]. autumn [IA04]. availability [GG04, dLMACC00, PF08b, YW05]. Average [GHC09, SCWD08, ASB05, Mac09, TNF09]. Avilés [GQCÁMI03]. avoidance [GCS^{+}04, HHO08]. Azerbaijan [Mam06]. Azores [AFM^{+}09, PGM01]. Azov [SB00a].

B. [ABB^{+}08]. baby [LD03b]. Back [CRB08, EKPT07, LP00]. back-calculation [LP00]. background [DH07, Vuo02]. backscatter [BB09, DHWW08, FGR04, GKO07, GJH^{+}09, LW04, MCL03, WWWB03, YM08, ZPK05]. backscattering [GO03b, GOK05, GC05b, KFM02, LN08]. Bacteria [KNS^{+}06]. bad [SJKN^{+}04]. bag [LS04]. Bahia [RSNB^{+}08]. Bahrain [YBF^{+}03]. bairdi [NTSM07]. bait [VHF^{+}04]. baitboat [RMAO^{+}03]. baited [Col02]. Balaenoptera [LHHJ^{+}09, LHR02]. balance [AKLL07, SAMS02]. Balancing [BFZ05]. Balearic [GM06, MM03b, MM03a, dPVV04, dPM08]. Ballast [BHH^{+}08, DB08]. ballast-water [DB08]. Balsfjord [UR01]. Baltic [DHKV01, ET07, KSD01, PVLP04, AMD^{+}05, AKLL07, Bac08, CA00, CCA04, CBHM07, DH04, EKPT07, FG09, FL06, Fjä05, FWW06, GS03, GFH04, HMK^{+}07, HCE^{+}03, HSA^{+}01, HSPM05, HKBK09, HP01, JHL05, KASA07, Kas09, Kol06, KM00, KMH^{+}05, KPD^{+}09, MNHL01, MRV^{+}08, MKFK05, NK00, NWH02, OKG^{+}09, OK05, Orł00, Orl01, Orl05, PCM09, Pel02, PB05b, PLP^{+}07, PMB^{+}03b, PH03, PN06, Rad03, RKP03, RS04, RLH01, RPK^{+}03, SSKE06, STJ^{+}07, UP00, VGBH09, Vin01, VSS07, Vuo02, WU03, WLS07b, WJTH00, WW01]. ban [FBD^{+}08]. bandwidth [CD03, CDB05, DC03b, DC04b, LN08]. Bank [MSGC^{+}09, OL07, BMDBM09, BQHG00, CGS09, DGC00a, DGC00b, DWC03, FSS00, GL00, GHBR08, Hol03, Kan07, MLNC01, MS07, MBM02, Rob05, RMM05, TCS^{+}09, TC01, VLBB08]. Banquereau [GGM^{+}05]. Barbara [CLFS02]. barbatus [FBD^{+}08, MLG^{+}09]. barbless [APGD08]. Barents [CRW^{+}01, BVD01, DBBM01, Dol02, GBT02, GDH02, HP07, Hel00, HM05, ISHB07, Joh02, JGST09, KMJH01, LLD^{+}05, LHR02, LT06, OGR^{+}07, OBNU02, OUNB02, OL00, PPTS09, SFØ07, SA03, Tan00, Ter02, Tje02, UP02, YS02, Ynd03]. barnacle [BLMB06]. barndoor [FMF02]. barotrauma [NC06]. barred [HAvH06]. barrens [PJ08]. Barriers [MHF^{+}09]. bartramii [CZC07]. based [BPM^{+}05, BBS09a, BDJ^{+}05, BKR09, BVB^{+}07, BFMJ03, CDQL06, CFN03, CTLN09, DDGR07, DH08, DBDA^{+}02, FGR04, FPS06, GAA^{+}04, GHFA09, HHMN01, HJBG04, JD05, JYW09, KS08, Kas09, Kat05, KMNP01, Kos09, KPD^{+}09, LP00, LKK^{+}09, LAO^{+}07, MMV^{+}08, MKR06, MH01, MHD02, MTJ^{+}07, MYAT09, MM05, NB08, NJ04, OSK^{+}05, ODCN09, OGR^{+}07, PG08, PCM01, PF08a, PCS^{+}07b, PCS^{+}07a, PMB^{+}03a, PKH^{+}08, PSC02, PRD^{+}06, RMDB05, RBD^{+}07, RUN07, RGG^{+}04, RR05b, RR09, RMB^{+}09, SRJ^{+}05, SFH^{+}07, SLMCRM05, Str05, TRM07, UASN07, Ung07, WYM09, WPF00, YSO^{+}03, YSF09]. basin [Cad00, GP00, KRYL09, MLG^{+}09]. basin-scale [GP00]. basis [GOS07, TCP05]. Basque [BLMB06]. bass [Gro06, PPL^{+}07, PKP07, RRC03, Sec00a, Sec00b]. batch [GCM09]. Bathymetric [BOC^{+}08, Wei05]. bathymetry [BM02, SKH02]. Bathyraja [ABB^{+}08, RLdAW06, SdlRdA06a, SdlRdA06b, SB01]. Bay [AO08, Ber04, Bun01, BFSC02, LSH^{+}09, Mor02, MNY^{+}09, PFF01, RR06, SFKC02, WBD^{+}06, BPD^{+}03, BW08, GFH04, SCLG00, SP03a, SHSKR01, dCA03, Ber00, BDTW06, CRvCB08, CPR06, EHG06, IFUR08, KMV^{+}07, KIDY09, LMM^{+}08, MAB^{+}07, ML08, PB05c, PVH^{+}05, RMAO^{+}03, SG00, Sec00a, Sec00b, SZ07, WWR^{+}08, WPR^{+}07, YM08, dPGPB06]. Bayesian [BKR09, BGW03, FBF09, HO01, HSR01, HK06, IFUR08, JRN06, JHC09, ÓMP^{+}04, PPC^{+}03]. bays [BI08]. BC [SHT^{+}09]. be [HSdLP06, JAC00, MPJ07, PKH^{+}08]. beach [NT02, NW02]. beach-spawning [NT02]. beam [AF06, BPWS09, GF00, GJH^{+}09, HPB09, HHT08, HF08a, HSA05, HMPC04, KRM05, MM07, MS09, MCL03, Pie00, RKE06, RvMBV00, RDHP00, RBGJ08, wScY02, TSK03, TK01, YCCH07]. beam-trawl [GF00]. Bear [CRW^{+}01]. bearing [War01b]. beat [HPB09]. bed [HSA05]. beds [GC02, GZND02, MCRF06, PJ08]. before [CMO^{+}06, SW06b, WBC^{+}08]. behavior [DSJ03, Dor01, ES03, dR01]. Behaviour [FLH04, GHI^{+}04, HA03, PHG04, WB04, AAV^{+}04, Aka02, ADDH04, BBC^{+}04, BDÑ04, Bjö02, BD07, CD07, GCS^{+}04, GG04, GLR06, GJR04, Hor03, HAG^{+}08, KK06b, LD03a, NW02, NC06, OMBP06, OGL06, ÖG04, PMM^{+}09, SB06, Sec00b, SDWQ09, SAN^{+}05, TMI^{+}04, TF04, TK01, UASN07, VHF^{+}04, WGM04, WM04, WLS07b, WWHB04, ZPK05, ZMM^{+}07]. Behavioural [Orl01, EB04, GG04, Rye04]. behind [RP07]. being [DCPvK07, GOK05, Mor04]. bellianus [PGM01]. belt [JP03]. beluga [AFHJ04]. Benefits [JvD07, BSS07, JYW09, PBH02]. Benguela [BCT05, CUUD07, FDD^{+}05, MJA^{+}05, SCJ00, UC05, YSF09]. Benthic [LPSL09, SFKC02, BvS00, BMDBM09, CEV00, De 04, DDR^{+}03, DBL07, Eno01, FHHH00, GF00, HH01, HHJK06, IWP00, JR01, KPD^{+}07, LVHU00, MM07, MML^{+}00, Mis02, Pie00, PR03, PFF01, PRF^{+}00, RS03, Tri00]. Benthos [Ele00, BFM00, GF00, MMD00]. Berg [PMB00]. bergii [MPG^{+}09]. Bering [SBC^{+}00, BWC00, HFWB05, HSR01, HM05, JMLG05, LJM00, MMD00, MS09, MM05, RO05, SDWQ09, Wal07, WFIM00, YTS^{+}06]. berried [AKJ07]. Bertalanffy [HC09]. best [SEOR09]. better [PPW^{+}09, SS07]. between [APGC04, BPM^{+}09, BGL08, BR02, BDÑ04, BBBF02, BLMB06, Buc00, BBK08, CAWD09, CA00, CCHV05, CDDM05, ES09, ELR01, FLH06, FHDM00, Fur02, GDL04, GZS^{+}09, GWF01, GWSV08, GHBR08, Gro06, HW06, HR01, Hoy07, HJBG04, JSMK06, KASA07, LLC^{+}08, LD03a, Lóp06, MvdKN05, MS09, Mor04, MNY^{+}09, NW02, OSK^{+}05, OBD^{+}05, PTTS00, Pet01, RM01, RRTdA02, RPB^{+}08, RASS09, SAAFCA07, SKC^{+}00, SG05, SAMS02, SJM03, TM02, TES^{+}05, UA04, WMS^{+}03, WFIM00, dPM08]. bi [NPPO06]. bi-dimensional [NPPO06]. Bias [aFADN08, BdP07, Bet04, SHdLP04, SGS^{+}05, dPGPB06]. Biased [KDP09]. Bicol [NC08]. bigeye [LCRS08]. Bight [ŠCBD09]. bio [HDG^{+}09, LO05, NM08, SP05, SRS^{+}07, UASN07]. bio-accumulated [SRS^{+}07]. bio-economic [LO05, SP05, UASN07]. bio-invasions [NM08]. bio-optics [HDG^{+}09]. bioassays [SLvdB^{+}09]. Biochemical [KCBC00, PFLFR08]. biodeposits [CFR^{+}01]. Biodiversity [Gre08, FGP09, HHAB09, Vec00, WBK^{+}09b, WBK09a]. Bioeconomic [MMB09, RD03]. biofouling [DBS06]. biogeochemical [LHKGS00, LC09b]. biogeography [ZPI^{+}09]. bioindicator [SDG^{+}08]. bioinvasions [PB08a].

Biological [DF00, FGBS00, KNO00, LDML08, MG02, PSFY07, WWR^{+}08, BO08, BBPW07, Bro02a, Des00, FGR04, GJL08, HH03, Mol00, PKH^{+}08, PPC^{+}03, RBD^{+}07, Rho08, TVH08, TH08b, WSWS03, WvdMF06]. biological-effect [TVH08]. Biology [ABB^{+}08, GAA^{+}03, PGM01, Bag04, Box06, Dun01, ECC06, ECC08, Gef09, GAFA06, Hen04, HL09, HR00, Mam06, MNGB07, NAK^{+}08, OV05, RNS08, RLdAW06, TAC00, VGF03]. biomarkers [SLvdB^{+}09]. Biomass [RUN07, YFL05, Arm01, BLMB06, BRC09, Bra05, BGG^{+}06, Bri02, CRC^{+}09, DDGR07, DCM03, DPN^{+}09, ERP01, FS02, FGP07, HHJK06, IFUR08, MMCD08, NGNB^{+}04, PF06, RCLD08, RUA07, RDB09, RFT02, SL04, SS09, SSI07, SLMCRM05, SBD^{+}09, Tje02, WMS^{+}03, War01a, WYMF08, Ynd03, vDBF^{+}09]. biophysical [HM05, MHD02, SS00]. biota [FN00, KTS02, NRS09]. biotic [ODRN05]. biotopes [BMDBM09]. bioturbation [PR01]. biovalue [RD03]. Bird [AFM^{+}09]. birds [MHF^{+}09, Tas00]. birthdates [AK04]. Biscay [Ber00, Ber04, RMAO^{+}03, CRvCB08, CPR06, IFUR08, KMV^{+}07, MAB^{+}07, PB05c, PVH^{+}05, SG00, SSU^{+}09, WPR^{+}07, dPGPB06]. bivalve [HHSM03]. bivalves [SL01]. Bivalvia [FM04, JGM^{+}08]. Bjørnøya [CRW^{+}01]. black [FBMR^{+}03, MN02, MD01, QGdS04, SDÖ09, SB00a]. Blackwater [Fox01, ROB04]. bleaching [CVL^{+}09]. blind [SMB09, TJG^{+}09]. blocks [WS02b]. blood [UBP^{+}09]. bloom [HP05]. blooms [KHS^{+}08]. blue [BK07, DM06, HBD05, JG07, PS06, RMM05, TCTC09, UBP^{+}09, WGMM08]. bluefin [CMC^{+}06b, GOA^{+}09, RF01, RMAO^{+}03, Sec02]. bluefish [CAAJ07, MC00]. bluemouth [SNV^{+}09, SGMN^{+}06]. Board [Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i]. Bodden [ODCN09]. bodies [BO08]. Body [KTRG06, BGL08, FL06, KNS^{+}04, OMTS03, RNK05, RAKS06, RBGJ08]. body-weight [RNK05]. Bomb [MC09]. Bonamia [CDDM05]. bonapartii [OV04]. Bonn [SA05]. Bonnaterre [Joy02]. boreal [GH04, KMT08, LT00]. borealis [Har07, HK06, KKC04, MASA06, Wie05]. Boreogadus [CMHN05]. Born [DC03a, DC03b, DC04a, DC04b]. Bothnia [JSMK06]. bottlenose [Lóp06]. Bottom [AE02, Cot01, OR09, ADDH04, BS03, BBS^{+}09b, CBDB02, CEV00, DWDD03, Eig09, FLP^{+}02, GR05, GAA^{+}04, HFWB05, HMAN03, HNLR04, KWBR08, MSB04, MMD00, PvHG09, RMDB05, SPGT00, SMB09, SL01, SW06a, SPS00b, TST^{+}09, ZWW^{+}03]. bottom-mounted [ADDH04]. bottom-set [CBDB02]. bottom-trawl [BS03, GR05, GAA^{+}04, HFWB05, RMDB05, SPS00b]. bottom-trawlers [MSB04]. bottom-up [Eig09, SL01]. bound [BO08]. boundaries [GWF01]. boundary [OMTS03]. boundary-element [OMTS03]. Bouvier [Rob08]. Bowdich [GAA^{+}03]. box [PRvB00, KF08]. brachydactyla [CF06]. Brachyura [PGM01]. brachyurops [ABB^{+}08]. brachyurus [LME05]. brackish [HCV03, NWH02]. Brazil [DNLSM08, FFL06, GAZ02, LO05, MNGB07, OV05, RSNB^{+}08, dCA03]. Brazilian [MNGB07, ZNGF02]. Break [MFB^{+}08, GQCÁMI03]. bream [Her04, KH03a, LSM07]. bred [ERGT07]. breeding [Bun01, CRW^{+}01, PCRW04, UC05, dLMS06]. Breidafjordur [JTE^{+}07]. Brendan [WBK09a]. Brevoortia [BW08]. bridging [OSK^{+}05]. bridles [Som04]. brief [SBC^{+}07]. Bristol [SZ07, YM08]. British [SML01, His01, MS02, RE00, SKC^{+}00]. Brittany [AVJ^{+}06, EHG06, MSB04]. Broad [CDB05, DC03b, DC04b, LN08]. Broad-bandwidth [CDB05, LN08]. broadband [SRJ03]. broadband-acoustic [SRJ03]. broadly [Ric09]. Brodsky [SRM00]. broodstock [DJRO06]. Brooke [WBK09a]. Broughton [SML01]. brown [BFK^{+}07, CRIP08, GSN^{+}03, Ano06j]. Brunswick [Ben01, CW06, OCWV06, JR01, WC01]. bubbles [Ost09]. budegassa [LDM08, MSB04]. Buffered [KKC04]. Buoyancy [SSA08]. Burrow [SP03a, CAWD09]. Bushehr [NAK^{+}08]. by-catch [CBDB02]. by-products [WS02b]. Bycatch [GLR06, MANT07, SJM03, Tal07, TM09, WLS07a].

C [Ber00, ODCN09]. Côte [BKN^{+}07]. Cabot [SFM01]. Cádiz [dHET04]. Cadmium [NCC^{+}07]. caesium [SRS^{+}07]. cage [AS02b, But01, GOA^{+}09, JO02, KTH^{+}00]. caged [CBS^{+}06]. cages [CH06, GLKPCP01, TAHK06, WSC^{+}06]. calamary [MS04, TA05]. calanoid [FR04, HLCG04]. Calanoida [LD05, PM04]. Calanus [Buc00, CH00, Cor00b, CM00, DK00, DGC00a, DGC00b, EH00, Fik00, Gaa00, GH00, GAP^{+}00, GA00, GP00, HHB^{+}00, Har00, Hea00a, Hea00b, HBG^{+}04, Hel00, HR00, HRHC00, Iri00, JLR^{+}08, Kaa00, MGH08, MCM00, Nie00, OED^{+}04, PTTS00, PB00, PPH09, PS06, SRM00, SOMT00, SBB^{+}05, TM00, TH05, URMS04, Uye00]. calculation [LP00]. Caledonia [CW09b]. calibrated [MCL03, MJA^{+}05]. Calibration [KWBR08, OMA09, GMM^{+}08]. California [ŠCBD09, Bot01, CLFS02, DDGJ02, Fru02, GZND02, Hel02, RNWS08, SHS01, SP02]. Caligus [UPK^{+}08]. callarias [EKPT07, KPD^{+}09, Rad03]. calls [PGD09, ZPI^{+}09]. Camargue [BMC^{+}07]. camera [MCP03]. cameras [Col02]. Campbell [HBD05]. campechanus [SBG06, WSFH02]. Can [GCC^{+}09, GC02, JAC00, MPJ07, PKH^{+}08, BJN^{+}06, FSP05, HSdLP06, MPG^{+}09, RO02, SMI07, Ska07]. Canada [BJN^{+}06, Ben01, CW06, CRB08, Dem01, Dup05, GGM^{+}05, LMM^{+}08, MMM00, MC09, MS02, OCWV06, She07, SKC^{+}00]. Canadian [Cla00, NSP06, SRS^{+}07, TH08a, TH08b]. canary [Ros05a, CFRM08, CHB09, HEGH02]. Cancer [BC07, PGM01, Ste08, Ung07, UMSA09, WvdMF06]. candidate [NJ04]. cannibalism [NK00, UP00, WFIM00]. Canyon [GQCÁMI03]. Cap [SRGC04, SF09]. capacity [ARMM09, CUUD07, PCM01, UKR05]. Cape [CDD^{+}07, GLR06, MKR06, MNCU09, MC00, SR03]. Capelin [CRW^{+}01, íJR02, NW02, RO02, Ros05a, Vil02, ADO02, Bro02b, Bro02a, CFL00, CMDN02, CF02, CMHN05, DAd02, Dol02, DBDA^{+}02, EN02, FRK02, GBT02, GDH02, GV02, GW09, HP07, HJBG04, JO02, Jør03, Mow02, NT02, Nau02, OR01, ORA02, OBNU02, OUNB02, OBD^{+}05, PPTS09, Ter02, Tje02, UP02, Vel02, VC02, YS02]. capensis [CDD^{+}07, GLR06, LC06]. captive [ERGT07, GJH^{+}09, NHKJK09, OAJ06]. captive-bred [ERGT07]. captivity [MF07]. capture [Aló08, BDO^{+}04, MF07, TS05, UBP^{+}09, WHG07]. captured [MDM03, VH08]. capturing [WPR^{+}07]. carapace [Har07]. carbo [FBMR^{+}03, MN02, MD01, QGdS04]. Carbon [GGV^{+}04, BCL03, BAO04, LBN09, UR01]. Carcharhinus [LME05, MSW07]. Carcharias [DBS06, LME02]. Carcinus [YG08]. care [Gre08]. career [NM09]. Carlo [LN03]. carry [OM05]. carrying [CUUD07, UKR05]. cascade [MMKKJ08]. case [AFM^{+}09, BCD^{+}02, BGG^{+}06, BR08b, CCB^{+}06, CRIP08, CMM03, CRTS04, DMvD07, GJL08, GLR06, GFKM07, HMK^{+}07, HF08a, Hol03, HIL00, MCB09, MMB09, Mis02, MNY^{+}09, NM08, PCM01, PKH^{+}08, PVH^{+}05, RRT00, RRC03, SW02, VBF09, VLJM^{+}07, WPB^{+}03, YM08, dHET04]. Caspian [Mam06, AP07, DM07a, GJL08, Kar06]. Caspian/Ural [GJL08]. Castellammare [BCD^{+}02, FBD^{+}08]. Catalan [SPGT00]. catch [AE02, APGD08, AK04, BVDS08, BR08b, CBDB02, CR04, CLR^{+}05, GDL04, GF01, GA05, HT05, HBD05, HB07, HA03, KS08, LLD^{+}05, LDML08, LZS09, LHJS02, Mac09, Mar08, MSF^{+}06, MRV^{+}08, MWF^{+}05, ÓMP^{+}04, PKP07, PCS^{+}07b, PCS^{+}07a, PPB03, PMN01, Pie02, PB08b, RSC^{+}09, RMAO^{+}03, SH07, She05, SS09, SW06c, SP03b, TS05, VM07, WvdMF06, YMF02]. catch-at-age [AK04, BR08b, PKP07, Pie02]. catch-at-length [VM07]. catchability [BSA09, BS03, EW07, FGP07, FGFP08, FGP09, GG04, JRN06, ORA02, RAB^{+}07, SPS00b, WGM04]. catches [BTR06, CMK09, GZS^{+}09, Gas02, GV02, Hor08, LHHJ^{+}09, LDM08, LCRS08, MS01a, NEJH05, SSI07, TES^{+}05, dVA07]. catching [RKE06]. catchment [Cad00]. categorical [PGG05]. catsharks [ECC06]. caudal [SPFF^{+}08]. caught [AK04, CMC^{+}06b, FSDB09, aFADN08, Kol06]. Caulerpa [RRT00]. cause [PCRW04]. caused [AS02b, BvS00, Dek04, HSA^{+}01, KNO00, NKOK00, NM08, RDB09]. Causes [ON09, DM07a, WJTH00]. CCAMLR [CdlMA^{+}00]. Celtic [BDJ^{+}05, BD02, BBK08, MAB^{+}07, PPB03, RPT02]. Censored [HT05]. censuses [RRTP02]. central [AMD^{+}05, AJR00, Bro03, CFN03, DGK^{+}09, FLP^{+}02, FBD^{+}08, GCS^{+}04, HA03, MKFK05, MFA07, OBNU02, OUNB02, KSD01, UP00]. central-west [CFN03]. Centriscidae [KTM^{+}05]. Centropages [DM04, LD05]. Centroscymnus [VGF03]. centuries [LLD^{+}05]. century [GWvM07, RE00, SBL07]. cephalopod [BQHG00, KCBC00]. Cephalopoda [MLS07, SKS^{+}00, TJAS04]. cephalopods [JMWJ08]. Cepphus [LPA^{+}00]. Cerastoderma [KPD^{+}07]. Cercopagis [GFH04]. cetacean [LHJJT04, MANT07]. cetaceans [KMV^{+}07]. Chaceon [Rob08, Tal07, WBC^{+}08]. Chaetognatha [DCN^{+}04]. chaetognaths [NC08]. Chagos [HR09]. Chain [LN03]. chair [Ric00b]. chalcogramma [BWC00, HH04, Hor03, KK06a, SBC^{+}00, Som04, WFIM00]. Challenges [AGH^{+}09, HW06, MV07, CSC^{+}04, KTT06, KRYL09, PK07, Pet04]. Chamcook [OCWV06]. Chamelea [MDM03]. Champlain [RPE^{+}03]. Change [PR07, APD09, BR02, BMJ08, CSVGTP09, Cor00b, Dri05, FQS01, GLDD00, GG09, HSdLP06, HB09, HBW^{+}09, KPO05, OK05, PB00, PB05c, RTDJ09, RKP03, Ric08, RPE^{+}09, Ros05a, Ros05b, SWG06, SE09, SN08, VPC^{+}09, VBF09, WYMF08]. Changes [BBÁMC06, DGPR05, FBD^{+}08, Hea05a, Hea05b, Mow02, RSC^{+}09, RE00, SKS^{+}00, SPGT00, WJTH00, Wie05, BdMAL00, BMV05, Bai09, BGAM00, BF04, Bra07, CCHV05, CSH00, Cur00, DLC03, FB02a, FHHH00, FMK07, GHC09, HR01, Kol06, LT06, LBN09, MNHL01, Miy03, MB05, NH09, PF06, PF08b, PPW^{+}09, RMDB05, Rob05, RTB^{+}05, SF09, wScY02, SS09, TF02, VSS07, WPM^{+}09, WS06, dPM08]. changing [CMDN02, CRB08, HM04b, MSP09, PHDC^{+}09, SKS^{+}00, VPC^{+}09]. Channel [CLFS02, HS06, Jag02, dPVJM04, CDR05, Des00, DWDD03, GLDD00, HBS^{+}06, RPR02, VCC07, WPB^{+}03]. channels [LCC08]. char [BF02]. Characteristics [JLR^{+}08, BFMJ03, DJRO06, DAAD09, ES09, HMHI09, HM05, KHO06, KMV^{+}07, LCRS08, MSB04, Miy03, MS04, MG02, MRT01, NCM^{+}03, Orl05, OLB01, PMB00, SMK08, WWHB04, WWWB03, dPVJM04]. Characterization [BH08, SML01, CMN^{+}07, CMM03, CD03, MCI03, Pet03]. Characterizing [KH03b]. characters [Tur04, Ung07]. charr [BFK^{+}07]. Cheilodactylidae [BES^{+}01]. Chela [RNK05, HSdLP06]. Chela-height [RNK05]. chemical [BGG^{+}08, HBC01, KI04, KTS02, WS02b]. chemistry [SGMN^{+}06, Zit01]. chemometry [JGN04]. Chesapeake [Sec00a, Sec00b]. chick [MNCU09]. chierchiae [LD05]. Chile [BBC^{+}04, CR04, EH00, GCS^{+}04, GG08, GWSV08, GGV^{+}04, HA03, LC06, LCC08, SQN08, WGLJM04]. Chilean [Peñ08]. chilensis [Alo01, CDDM05, EH00]. chimaeras [SBDW00]. China [LLC^{+}08, LC09b, YYY^{+}02, YCCH07, Aka02, HLL^{+}08]. Chinook [RHD09]. Chionoecetes [NTSM07]. Chirp [LKK^{+}09]. Chlamys [JTE^{+}07]. chlorinated [SVRF08]. chlorophyll [GAP^{+}00, GHC09, VM09]. choice [HMPC04, KPD^{+}07]. choices [KBDC^{+}08]. Chokka [Rob05, OSB06, OR09]. chondrichthyans [SBDW00, WHA08]. Chondrichthyes [ECC06, ECC08, MG07]. chromatography [NNT01]. chromophoric [URMS04]. chronic [MMD00]. chum [aFADN08, MMS01]. ciliata [BBB06b]. Ciona [TH08a, TH08b]. circulation [HHB^{+}00, HSA^{+}01, LTI09, SR03]. circumference [BM09]. cirripede [MWS04]. clam [AGY^{+}05, AFGR09, DSG05, Gas02, GLS^{+}03, GAYR06, GGM^{+}05, MCRF06, MDM03, PDRG04]. class [Cot01, CMP07, EN02, GOS07, HSS^{+}09, OKG^{+}09, OL00, RT03, SBB^{+}05, WS06]. classes [LHHF03]. Classification [Fer09, AGC02, AVK^{+}08, BPWS09, BMDBM09, BH08, DCRB09, EGB02, FSQ^{+}03, HNLR04, KCD^{+}03]. Classifying [FLK^{+}09, HHAB09, WBK^{+}09b, WBK09a]. Clausocalanus [PM04]. clavata [SPFF^{+}08]. cleaning [HF08b]. clear [OGD09]. Climate [Lit06, MKFK05, APD09, BDJ^{+}05, BMJ08, Bra05, CBM09, CSVGTP09, CH05, DiUVH08, DHKV01, Dri05, FHDM00, FCM05, GP00, HLCG04, Hea05a, Hea05b, HB09, HBW^{+}09, KPO05, KMH^{+}05, LLD^{+}05, MMKKJ08, OK05, PB05c, Ric08, RPE^{+}09, Ros05a, Ros05b, RCL05, SGM09, SBB^{+}05, SE09, SN08, VPC^{+}09, WFIM00, WYMF08, Ynd06, ZK00, dPM08]. Climatic [HEGL05, SA03, dPVV04]. climatological [SR03]. cline [CLK^{+}09]. clockworks [GC07]. Closed [AHS08, DBL07, FSS00, HD00, LD03b, PRvB00, SJ08, SGY08, SJM03, TCS^{+}09]. closure [DDR^{+}03, GAM^{+}06, LC09a]. closures [HHJK06, JSR06, LPM^{+}09]. Clupea [BRP02, BD02, BD03, BBK08, BBSK09, CA00, CCA04, CMHN05, CBHM07, DEMD00, EDG03, Fox01, GFH04, HMDS09, HP01, HK00, HSS^{+}09, Joh02, JHL05, LMC^{+}01, LHR02, NTJ04, OL07, PHG04, PMB00, PHO09, PVLP04, PB05b, PH03, PN06, RLH01, STAN02, SHT^{+}09, SDCR07, TK03, Vuo02, WPM^{+}09]. clupeid [HM04a, KM00, PPMH04]. clupeids [FG09]. clupeoid [Orl05]. Clupeonella [Mam06]. cluster [CS02, HSR01]. clustered [HM04a]. clusters [HS01, Pet01, Pet03, SP03b]. Clyde [TAC00, BM01a, BM01b, MAAN09]. Cnidaria [WBV09]. CO [HMMB^{+}08]. coal [KTS02]. Coast [GAA^{+}03, BPM^{+}09, BJN^{+}06, BRE^{+}08, BML^{+}05, CAGV05, CEH03, CH06, CSR^{+}02, CLK^{+}09, De 04, Des00, DBS06, DGK^{+}09, ECC06, HM04a, JSR06, KTRG06, LDCH^{+}09, LSGD02, Mam06, MC09, MLS07, MCI03, Nau02, NC08, OSB06, OB05, PZTE05, RSNB^{+}08, SNM05, She07, SP07b, SSA08, Sve03, TCC08, WHA08, ZNGF02, Cas07, SGM09]. coastal [AP07, AGC02, BdMAL00, Bai09, BA03, BO08, BBBF02, BAO04, BF04, BRHG^{+}06, CCB^{+}06, CTF02, CRB08, CSVGTP09, DJRO06, EGO^{+}07, FJSJBS^{+}08, FTDVC^{+}08, FSQ^{+}03, GMKS06, HAG^{+}08, HSS07, KCL^{+}09, KTH^{+}00, KHS^{+}08, KWZ00, LDNS08, LPH^{+}08, LFW03, MML^{+}00, MSR03, MNMG^{+}05, NAK^{+}08, NNT01, PCD05, RTDJ09, RE00, Ros03, RRY08, SLMCRM05, SGMMGB09, SSA08, TCM^{+}08, Vel02, VBF09, VLJM^{+}07, WPB^{+}03, WW01, YBF^{+}03]. coastal-shelf [FSQ^{+}03]. coastal-zone [RRY08]. coasts [RTB^{+}05]. Cobscook [WBD^{+}06]. cockles [KPD^{+}07, MPJ07]. cocktails [BG07]. Cod [BA03, Bra05, MGvH06, OLB01, Ste01a, UP00, AG00, Arm01, AGA^{+}04, BGL08, BBM^{+}02, BSO01, Bjö02, BD04, BÓ06, BDS01, Bra07, BMLH07, BDTW06, BR08b, CH06, CMHN05, CWC^{+}03, CCC02, CSdQB06, DJRO06, DMvD07, Dri05, DR08b, DLC03, EKPT07, ET07, Erm09, EGO^{+}07, FL06, FMB01, FGBS00, FR09, GWG06, GTOJA06, GF01, GMGN06, GW04, GJH^{+}09, HP07, Ham06, HRM04, HDG02, HSPM05, HKBK09, HOF04, HWF08, HOD06, HLS00, HJBG04, ISHB07, Joh02, JDN01, JCM06, JMC07, JV05, JKSO06, KNKT06, KM05, KPO05, KCR06, KTRG06, KMNP01, KNS^{+}06, KM00, KMH^{+}05, KPD^{+}09, KMJH01, KB07, LGR08, MNHL01, MGTS00, MGS00, MW03, MSS^{+}05, MR05, MJB08, MB05, MG02, NK00, NC06, OMBP06, OSLO06, ODRN05, OBD^{+}05, ON09, OAJ06, OFN02b, OFN02c]. cod [OL00, PSSD08, PPK^{+}06, PK09, PF08b, PLJ01, Rad03, RL05, RP07, RML06, RR02b, RR06, RO02, Ros03, Ros09, RHBR04, She05, Sin01, SMK08, SBB^{+}05, Som04, SP03b, SRMB07, SG05, SSA08, SPWHR04, SN08, STJ^{+}07, SB03, SPS00b, SFM01, TAHK06, VGBH09, VSÅF05, VSC06, WSC^{+}06, WJTH00, YM00, Ynd01, YW05, vdKRS^{+}07, MC00]. code [Ynd03]. codend [BM09, BMU09, OH07, ÖFR^{+}06, ÖTTM07, RBGJ08, ZFFT01]. codends [GM06]. coefficients [JMLG05]. coelolepis [VGF03]. coexist [BJN^{+}06]. Coexistence [GPRD08]. coherence [Ard08]. coherent [BBPW07]. Coho [RHD09]. cohort [CZC07, Hor08]. cohorts [BHØ^{+}04, HRHC00]. COI [KHS^{+}08]. Coilia [HLL^{+}08]. coindetii [AJR00]. cold [MSGC^{+}09, NRR^{+}09, NFM^{+}02, SSJL02, Tan00]. cold-front [NFM^{+}02]. cold-water [MSGC^{+}09]. Collapse [JTE^{+}07, DM07a, She05, SP07a, SRMB07]. collect [CLL^{+}09]. collected [DJRO06, Peñ08]. collection [KDO^{+}08]. collections [DBDA^{+}02]. collide [RL07]. Cololabis [STA^{+}09]. colonization [HEGH02]. colourations [GW04]. columba [LPA^{+}00]. Columbia [MS02, SKC^{+}00, CBDS08, PP08, SML01]. column [SAAFCA07]. Comacchio [Mis02]. comber [Aló08]. combination [BMDBM09]. Combined [GWG06, BBBF02, BVB^{+}07, Erm09, HF08a, VSÅF05]. Combining [KHE^{+}09, PGMB09, RTB^{+}05]. Comment [HHAB09, WBK09a]. Commercial [MLMC02, Bag04, BNBR05, BLRC05, Bun01, DLR02, Eig09, GS03, GGP07, KNS^{+}06, MMCD08, MS01a, MB06, OM05, PPB03, She05, SSI07, SPD00, SBT^{+}09, SP03b, WM01]. commerson [HAvH06]. Commission [PD07]. commitment [HMK^{+}07]. common [DNLSM08, FAL^{+}08, GCS^{+}04, JYW09, KMI^{+}05, LHHJ^{+}09, MPD^{+}08, RNWS08]. communicating [Deg05, PPKM07]. communication [vDM07]. communities [AS02a, BGAM00, Bia00, CHB09, CJS02, CBS^{+}06, DDR^{+}03, FHHH00, GRE06, GLDD00, GAYR06, GGM^{+}05, GPRD08, HHJK06, MSF^{+}06, Mis02, MM05, NRR^{+}09, PRF^{+}00, RK04, RTB^{+}05, SDRK00, SJM03, SBD^{+}09, TD00, Tri00, VCC07, WCP08, YSF09, Zwa00]. Community [TCM^{+}08, AFP^{+}09, ASB05, Bla01, BDJ^{+}05, BBÁMC06, Cal02, CTF02, CFN03, Cor00a, DGPR05, FLP^{+}02, FB02a, GL00, GCC^{+}09, GCS09, GR06, HSM09, JSR06, JD05, Kou00, LLC^{+}08, Lit06, MCRF06, MML^{+}00, NJ04, PGJ^{+}05, PJ05, PvHG09, PR03, PRD^{+}06, PB05c, RS03, RBBB00, Ric00a, SMI07, SQN08, SAPP04, SW06a, Sve03, TPRR04, TMG^{+}08, TC01, UR01, dBP02]. comparable [BCD^{+}02]. Comparative [Hea00a, HBG^{+}04, MLM05, VHI^{+}04, DNP03, FMB01, HS09, LCC09, Mar08, OED^{+}04, Roc00a]. compare [Lit06]. compared [DLC03]. Comparing [HH04, JFCH05, MJA^{+}05, PPB03, GML06]. Comparison [AS02a, ATM02, CCB^{+}06, CTW09, HP04, HHC^{+}09, HM05, JKSO06, NGNB^{+}04, Sim07a, WSC^{+}06, YFL05, AHS08, AFGR09, BIdL^{+}08, BML^{+}05, CDM03, CF06, DHWW08, FGR04, GLS^{+}03, GAYR06, HLCG04, LSH^{+}09, MUK^{+}02, MCL03, ORA02, RS03, RL07, Sec02, SMEK01, SJM03]. comparisons [OLS00]. compensate [GC02]. compensation [BP07, ZO03]. competition [ATH^{+}07, BCT05, KASA07]. Competitive [RvMBV00, PK07, RDHP00]. Complementary [DKMO09, GLR06]. completed [HSCN06]. complex [BBS09a, BBBF02, Gef09, OL07, PJ08, TA05]. complexes [JHC09]. complexity [BNF^{+}07, CSR^{+}02, HS09, SGS02]. compliance [STJ^{+}07]. component [Boo00, LDM08]. components [JNF^{+}09, JMC07, KYG03, YMF02]. Composition [MWS04, OBNU02, BKN^{+}07, BB09, Bia00, BGW05, CAWD09, Cas07, DBL07, FGFP08, aFADN08, HSR01, HS01, HLL^{+}08, KI04, KTS02, KGT01, MGTS00, O’D03, PPK^{+}06, RSC^{+}09, SB01, SCWD08, SP03b, URMS04, WvdMF06]. Compositional [SH07]. compositions [CSdQB06]. Comprehensive [AP09, FN00]. compression [GO03a, GO03b]. compressive [KTS02]. compromise [KF08]. computational [TSK07]. computationally [MHH06]. computed [LKK^{+}09]. Concarneau [EHG06]. concentration [FHJS09, IB00, PF06, URMS04, VM09]. concentrations [BGL08, Ped05, SRS^{+}07, UR01]. Concepción [SQN08]. concept [Roc00b]. conceptual [HNK07]. conclusions [VPC^{+}09]. concrete [KTS02]. condition [BPM^{+}09, Ber00, Ber04, BWC00, CF02, DLC03, Jør03, KWZ00, LdSSG02, MR05, MAAN09, Mor04, OB05, OLB01, RAKS06, Sim07b, SCLK01, WPM^{+}09, YM00]. conditions [AE02, DEMD00, DM04, FHJS09, FRC03, HM04b, JKSO06, LLC^{+}08, LSGD02, MVMH04, MSP09, OBD^{+}05, OAJ06, PS06, SKS^{+}00, vdMBD00]. Conference [Ano01h, Jen02b, PB08a, NM09]. Conflict [Cla00]. conflicts [HMQ^{+}08, TT08]. congeners [HLCG04]. Congruence [TES^{+}05]. conical [WW07]. Connecticut [AMJ^{+}06]. connection [HR01, PTTS00]. connections [SCHR07]. Connectivity [SKC09, SMP09]. Consecutive [WPJ09, JMWJ08]. Conseil [RDF^{+}03]. Consequences [BdP07, RS04, BFM00, Ber04, BFK^{+}07, DM07b, DDR^{+}03, ERBP09, FGP09, HP05, KPD^{+}07, LO05, ON09, TSK03]. Conservation [CdlMA^{+}00, ERGT07, HTSB04, MPJ07, BMC^{+}07, BFZ05, GM07, GWvM07, Gro06, HS07, HPR09, KGRW07, MBC^{+}09, Pen07, XZW05, YCCH07]. conservationist [Aga00]. conserve [CdlMA^{+}00, FGP09]. consideration [CDD^{+}07]. Considerations [Cha04, CSC^{+}04, DDGJ02, FB07, HHMN01]. considered [BF04]. Considering [PPW^{+}09]. Consistency [Mui03, ZWW^{+}03]. consistent [MFD02]. constraint [KWL^{+}02]. Constructing [BBPW07, HHMN01]. Consultation [CWYM^{+}02, SW02]. consumer [DMDE04]. Consumption [OLS00, OL07, BANGC02, BCAN^{+}06, BS02, Bun01, GKFM09, Hea07, HA03, MKR06, RNWS08]. consumptive [Fru02]. containing [OSLO06]. contaminant [BG07, FN00, FN02]. Contaminants [TVH08, SVRF08]. contaminated [SLvdB^{+}09]. contamination [BG07]. content [BVD01, TM02]. contents [ATM02, Joy02]. context [Bai09, FTDVC^{+}08, KCD^{+}03, PB00]. Continental [Kos00, BSA09, BKN^{+}07, BvS00, CMN^{+}07, CMM03, GKOV05, GSdFB01, JI05, LBF01, MM03b, MM03a, MS00, PB05c, SdlRdA06a, SdlRdA06b, SNV^{+}09, TLMO08, WCP08]. Continuous [PF06, TAHK06, PCS^{+}07b, PCS^{+}07a]. Contrasting [Ber04]. contrasts [CW05]. contributed [CvdLHF08]. Contribution [JMC07, LSM07, BDTW06, JGM^{+}08, RFM^{+}02, TMB08]. control [CFN03, CGS09, CCC02, KM00, Lun01, MC07, NK00, PGMB09, RL07, RD07, SL01, SZ07, TR09, ZCH06]. controlled [FHJS09, JO02]. controlling [KS08]. controls [JSMK06, LPM^{+}09]. controversy [OSWL02]. Conveners [HW06, RC07]. Convention [CdlMA^{+}00, Joh08]. conventional [BHR^{+}05, DHWW08, LCC09]. conversion [BSO01]. conveyor [JP03]. Cook [O’D04]. cooperative [JvD07]. Copepod [Mil08, BF04, FR04, GNC08, LLC^{+}08, Tan00, URMS04, WHP08]. Copepoda [FSDB09, LD05, PM04, SBB^{+}05]. copepodite [Hel00]. copepodites [CM00, FR04]. copepods [CBDS08, DM04, KMT08, OBNU02, PSO^{+}04, PGG05, SC00, UR01]. coping [Gef09]. copper [LME05]. Coquimbo [GWSV08]. Coral [MMKR^{+}00, AS02a, ASB05, CW09b, CVL^{+}09, LPM^{+}09, MSGC^{+}09, PPW^{+}09]. coral-algal [MMKR^{+}00]. coral-reef [ASB05]. cordgrass [TDE08]. cormorants [ATH^{+}07]. corrected [Mac09]. Correcting [HHO08, RW01, WHG07]. correction [aFADN08]. correlated [MMS01]. correlation [KYG03, Mui03]. Coryphaena [DNLSM08]. Cost [WHP01, GKOV05, MGvH06]. Cost-effective [WHP01, GKOV05]. count [HCEM06]. counter [NGNB^{+}04, TES^{+}05]. Counting [ORA02, dR01]. Country [BLMB06]. coupled [KIDY09, LBN09, RR02a]. Covariability [OL00]. covariance [MRV^{+}08]. covariate [FN02]. covariate-dependent [FN02]. cover [RBGJ08]. coverage [ANNG01]. cpue [CCHV05, CMP07]. cpue-at-age [CMP07]. crab [Col02, CF06, HSdLP06, Rob08, Ste08, Tal07, Ung07, UMSA09, WBC^{+}08, WW07, YG08]. crabs [BC07, JYW09, MS09, NTSM07, RNK05, SZ07, SHdLP04, UBP^{+}09, ZK00]. Crangon [CRIP08, CDB05]. Crassostrea [HKI01, HPBK04]. creating [NM09]. creep [MAC^{+}07]. crested [MPG^{+}09]. crimson [ZCH06]. cristata [CGN^{+}04]. criteria [Lin05, NB08, PCM09, RL07, RR05b]. critical [CBM09, KRM05]. Critique [Syr00]. Cros [CBBL09]. Cross [MFB^{+}08, Fra06, Pel02]. Cross-front [MFB^{+}08]. cross-validations [Fra06]. cruises [ODCN09]. crustacean [Eno01, GM06, HSM09]. crustaceans [BMP^{+}08, BM01b, RBGJ08]. CUFES [PGMB09]. cultivation [BBB06a]. culture [BHN06, But01, CFR^{+}01, LSH^{+}09, RS06b]. Cultured [JJ06, Agn08, OGL06, WC01]. cumulative [SW06a]. cupressina [WBV09]. Current [BWK07, DGK^{+}09, RAR^{+}07, RNWS08, SG00, SHS01, AVK^{+}08, GNC08, IA04, Kat05, WBC^{+}06, ZKP03, ZPRJ02, Bot01, EH00, GG08, GGV^{+}04, KWZ00, NKOK00, RFM^{+}02, VLJM^{+}07, WGLJM04]. curve [SH07]. curves [Cot01, CMP07, FSDC09, JI05, YFL05]. Cushing [RRC03]. cuttlefish [CDR05, WPB^{+}03]. Cyanea [HMHI09]. cycle [ÅD07, AKJ07, CLM07, CRTS04, DCN^{+}04, DSG05, FBMR^{+}03, GAP^{+}00, HBG^{+}04, Orl01, Orl05, Rob05, SA03, Ynd06, dLMS06]. cycles [MR05, MAAN09, Ynd03]. cycling [GGV^{+}04]. cyclophora [OV05]. cygnus [dLMS06]. Cynoscion [Kup04, LO05]. Cystophora [CGN^{+}04].

D [WBK09a]. dab [NWH02]. dactylopterus [SNV^{+}09, SGMN^{+}06]. Daily [SPG^{+}04, DDM^{+}05, GKFM09, ODCN09]. Dakhla [FB07]. damage [HU04, LS04, Lun01, MDM03, NM08]. damaged [Gas02, GF00]. dangerous [GGP07]. Danish [MNHL01]. dark [dVA07]. Data [HMQ^{+}08, MM02, AGY^{+}05, Ard08, BABB08, BR08a, BM02, BNBR05, BPT09, BIdL^{+}08, BHR^{+}05, Boo00, BGW03, BPWS09, BVD01, BH07, CDB09, CLL^{+}09, CTW09, CMK09, DLR02, DLS01, DCPvK07, DAH^{+}08, EIS05, FB03, FGD02, GPP09, HS01, HT05, HJB^{+}08, HC09, HRM04, HCEM06, HWF08, Iri00, JHC09, KMV^{+}07, KO02, KDO^{+}08, KHE^{+}09, LLD^{+}05, LEP04, LFD^{+}09, LKK^{+}09, LBN09, Mac09, MSF^{+}06, MMF09, MHD02, MLMC02, MTJ^{+}07, Mol00, dLMACC00, MB06, NGNB^{+}04, ORVP09, PGG05, PCM09, PGMB09, PQRG07, PR03, PF06, PH05, PPC^{+}03, RD07, SGM09, SH07, She05, SLN02b, Ste01a, Syr00, TM02, TES^{+}05, TT08, TS05, UE01, WM04, WPB^{+}03, WSWS03, WWR^{+}08, ZPK05, ZCR09, vdKRS^{+}07]. data-diagnostic [LBN09]. data-limited [Ard08]. data-poor [Mac09]. data-processing [CTW09]. data-rich [PPC^{+}03]. data-selection [PCM09]. data-storage [vdKRS^{+}07]. datasets [HHC^{+}09]. date [MGS00]. dating [MC09]. Daugava [PMB^{+}03b]. Day [PMN01, AE02, OGD09]. daylight [RW01]. days [GHFA09]. dead [MSS^{+}05, MR09, TJG^{+}09]. debate [Bai02]. debris [CLFS02]. Decadal [LBN09, dPM08, LTI09, ZK00]. decades [CF02, NCC^{+}07]. decapod [BM01b, Mil08]. Decapoda [LCC08, PGM01]. decipiens [MMM00]. decision [Cla00, DMvD07, Dor01, Lin05]. decision-making [DMvD07, Dor01]. decisions [KPD^{+}07]. decline [Bra07, Dek04, SB03]. declines [FMK07]. decommissioned [JLS02]. decomposing [LAO^{+}07]. Decreasing [CA00]. Deep [LHJJT04, SB00b, BSA09, DNP03, DCM03, ERBP09, HHAB09, HL09, Kos00, LGH^{+}09, MMC03, Rob08, Sar09, STG06, SB01, Str05, Tal07, WBC^{+}08, WBK^{+}09b, WBK09a]. Deep-ocean [LHJJT04]. deep-sea [HL09, Kos00, MMC03, STG06, Str05, WBC^{+}08]. deep-towed [DNP03]. deep-water [BSA09, DCM03, ERBP09, HHAB09, LGH^{+}09, Rob08, Tal07, WBK^{+}09b, WBK09a]. deepwater [GM06, KH03b, MN02, TLM04]. defining [DB08]. Definition [KRM05, SA03]. Definitions [Mur00a, MFD02]. deformity [KTRG06]. degrees [BdMAL00]. Delaware [BFSC02, SFKC02]. Delay [SP07a]. delayed [Dav07]. delays [TAHK06]. Delile [FTDVC^{+}08]. Delineating [SFM01]. deliver [SMI07]. Delphinapterus [AFHJ04]. Delphinid [CRvCB08]. Delta [MCRF06, Syr00]. Demersal [MAMO02, MM03b, MM03a, Bia00, Bla01, BZRO06, CCHV05, CTF02, CH05, FGP07, FGFP08, GJR04, GHFA09, HFWB05, íJCMR07, JSR06, JR06, JI05, KHM09, LdSSG02, MUK^{+}02, Mar08, MR09, MHV09, MS00, dLMACC00, NW02, RBD^{+}07, RDD06, RR06, RE00, Sve03, SBD^{+}09, WMÖ06, Zwa00]. demersus [CUUD07]. Demographic [MPG^{+}09, DP03a, DP03b, MRT01]. Demographics [DOBT02, Sec00b]. demography [Hea00a, MSH07]. demonstrating [GFKM07]. Denmark [ATH^{+}07, RFM^{+}02, UA04]. dense [ZO03]. densely [GC05b]. densities [EGO^{+}07, PO09]. Density [ELR01, RR06, BB09, BÓ06, BMJ08, BGW03, CAWD09, CW05, Col02, DLS01, EW07, GJH^{+}09, HK00, HSS07, JH01b, MCRF06, MR09, Mye01, ORA02, PCD05, PSC02, RRC03, SP03a, SAMS02, TLM04, Wie05, vdMBD00]. density-dependence [PCD05]. Density-dependent [ELR01, RR06, BMJ08, RRC03, Wie05]. denticles [SPFF^{+}08]. dependence [CVG08, JHKZ09, Mye01, PCD05, vdMBD00]. dependencies [BRP02]. dependent [BS03, BMJ08, ELR01, FN02, Fur02, HOF04, JPO09, Kup04, LW04, OFN02b, OFN02c, RR06, RRC03, WHG07, Wie05, ZPK05, ZWD08]. Depensation [RHBR04]. depleted [BMJ08]. Depletion [Mac09, BDD06, MPG^{+}09, RUA07, RPR02, WHG07]. Depletion-corrected [Mac09]. deployed [BFSC02]. deployment [KWL^{+}02, WS02a, WLK02]. deposited [CFMdP07]. depredation [MCB09, RSC^{+}09]. deprivation [DLC03]. Depth [DGC00b, SRS^{+}09, ZWD08, Aló08, BS03, MM03b, MM03a, PMN01, SAAFCA07]. Depth-dependent [ZWD08, BS03]. derive [PvHG09]. derived [BW08, MJA^{+}05, NGNB^{+}04, PKP07, PK09, PR03, PH05, RM01, SRN00, SHdLP04, WM04]. dermal [SPFF^{+}08]. describing [HLL^{+}08]. Description [SPFF^{+}08, TMB08]. Descriptors [FTDVC^{+}08, DNLSM08, LBF01]. Design [DDGJ02, SPS00a, BG04, CW09b, GFKM07, HOP09, HKBK09, LCC09, MSS^{+}05, MSI07, NCM^{+}03, SGS02, SGC^{+}09]. Designing [SMI07, CMJ09, RPB07]. designs [WYM09, dVA07]. despite [BHMD05]. detail [Mil08]. details [Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i]. detect [BMJ08, FSP05, NJ04]. detected [BH08]. Detecting [VM07, BH07]. Detection [CM00, NNT01, OCWV06, BM02, BBS^{+}09b, BPWS09, CTW09, HSA05, MYAT09, RO05]. detections [BAB^{+}04]. deter [TM09]. determinant [SG00]. determinants [DGK^{+}09]. determinate [CF06]. Determination [GOA^{+}09, dHET04, DWC03, FR04, Iri00, MC09, Pel02, SBT^{+}09, SGS^{+}05, SVRF08]. determine [Bri02, HSdLP06, JAC00]. determined [BBM^{+}02, BHMS02, MASA06, OKG^{+}09, VSC06]. Determining [HL07, O’D03, DM06, Vec00]. deterministic [Cor01]. deterrent [GHD^{+}09]. Developing [DB04, PD07, RBGJ08, SEOR09, TCP05, HCV03, RUCG07, TVH08]. Development [Den08, HBD05, LHKGS00, MMV^{+}08, RPK^{+}03, RRY08, AS02a, AWW^{+}07, CBS^{+}06, CA02, CVG08, Eig09, Gaa00, GNC08, HM04b, KMG^{+}07, KTT06, NHKJK09, NB08, NIF^{+}09, OR09, OSLO06, PRvB00, PMB^{+}03b, SW02, VGF03]. developmental [WHP08]. developments [BBS^{+}09b]. Deviation [KM05]. deviations [AE02]. device [CRIP08, DPN^{+}09, EHL07, GHD^{+}09, HB07, JB00]. devices [FWW06, D00]. Diadromous [HW06]. diagnostic [LBN09, RTB^{+}05]. diagnostics [CF05]. diagrams [Kat05]. diamond [GM06]. diamond- [GM06]. diapause [Fik00]. Diatom [SCLK01]. Dicentrarchus [PPL^{+}07, PKP07]. Dichelesthiidae [FSDB09]. Dichelesthium [FSDB09]. Did [She05, SRMB07]. Didemnum [AO08, Den08]. DIDSON [HW08, HCEM06]. didymozoid [PAA06]. Diel [ADDH04, BBC^{+}04, DK00, GR05, GPP09, HK00, JG07, Orł00, Sab04, SBC^{+}00, SS09, TPT^{+}09, ZMM^{+}07, BSA09, BS03, FS02, NCM^{+}03, Orl01, Orl05, PMM^{+}09, VSS07]. diel-vertical [NCM^{+}03]. Diet [ATH^{+}07, BCAN^{+}07, CSdQB06, RS06a, EH04, FR04, GKFM09, HAN02, HOGH07, LLHK07, LSM07, MAB^{+}07, ODRN05, SJGRRRE02, TSH^{+}06, URMS04]. Diet-induced [CSdQB06]. dietary [BGW05, CSdQB06]. Diets [PS06, FJSJBS^{+}08, KNS^{+}06, LPA^{+}00, OSLO06]. difference [KFM02]. Differences [BFK^{+}07, OAJ06, Buc00, CLM07, CSdQB06, GML06, HSS07, NH09, Vel02]. Different [LPM^{+}09, TPRR04, DM04, HLCG04, JMC07, JKSO06, KA01, KTRG06, KNS^{+}06, LO05, OUNB02, OBD^{+}05, RK04, RBGJ08, SGM09, SSA08]. Differential [vDEM^{+}00, ASC01]. differentiation [JDN01, RMM05, TCTC09]. Differing [GSN^{+}03, SL01]. Digestion [CMHN05]. digestive [OMBP06]. digital [BI08, GPWG04]. dimensional [CRC^{+}09, GCS^{+}04, LC09b, MHD02, Mol00, NIF^{+}09, NPPO06, SS00, SKH02, TNF09, TCP05, VGBH09]. dimensions [CW09a]. diminished [SP07a]. dimorphism [Bro03, CLM07]. dioxide [LBN09]. Diplodus [LSM07, SJGRRRE02]. Dipturus [Alo01, FMF02, LCC07]. Direct [KNS^{+}04, EMA^{+}07, FGR04, GLS^{+}03, PvHG09]. direction [HHT08]. directions [AVK^{+}08, JD05]. Directive [FTDVC^{+}08]. directivity [CD07, CRC^{+}09]. directly [PFLFR08]. Dirichlet [HBST02]. discard [AP09, BZRO06, BvKvH^{+}08, BS02, BMU09, DSV^{+}08, DCPvK07, Tal07]. discarded [BM01a, BM01b, MBPW06]. Discarding [HU04, UBP^{+}09]. discards [GS03, GFKM07, GF00, MNCU09, RPT02, TMG^{+}08]. discharge [JWBP07]. discourse [DMvD07]. discovered [SQN08]. Discrete [SB04]. discriminant [EDG03]. discriminating [BBK08, DNLSM08]. discrimination [AOSD09, BML^{+}05, GH04, GKOV05, LW04, LN08, MCAS04, PR01, WSP03]. discriminatory [FB03]. Diseases [BBB06a, Ste08]. Disentangling [TS05]. disequilibrium [KK06a]. disguised [RR09]. Dispersal [LCC08, TCS^{+}09, Aco02, BBMS01, FSDC09, LC09a]. Dispersant [CSW06]. dispersion [DH09, HSA^{+}01]. Displacement [MS02, GFP09]. display [CGV03]. displaying [MC07]. disposal [CTF02, WWR^{+}08]. dissolved [URMS04]. distance [HP04, Kal01]. distant [BF02]. distinct [GWF01, JHL05, VCC07]. distinctness [SCWD08]. distorted [DC03a, DC03b, DC04a, DC04b]. distorted-wave [DC03a, DC03b, DC04a, DC04b]. Distribution [ANNG01, AJNM07, DBBM01, Hel00, KYG03, KMV^{+}07, MM05, SRM00, SB00b, VLBB08, AFM^{+}09, ADDH04, BRE^{+}08, BREB09, BA03, BDÑ04, BBGA05, BR00, BR04, BHM^{+}04, BGW03, Bro02b, BES^{+}01, CCHV05, CRvCB08, CvdLHF08, DGC00a, DGC00b, EW07, FM04, GL00, GA00, GQCÁMI03, HM04a, Hea00b, HD00, HHH00, Jag02, JWM03, JGST09, Kal01, KMHS04, KA01, LBL06, LC06, LG08, LDCH^{+}09, LND05, MVMH04, MM03b, MM03a, Mow02, MWF^{+}05, MWS04, NH09, NCM^{+}03, NTSM07, NC08, NGNB^{+}04, OKRK04, OBNU02, ÓGS09, PTTS00, PCS^{+}07b, PCS^{+}07a, PM04, PP08, PAC02, PS06, RS03, RM01, RL08, RAKS06, RR06, Ros05a, RSNB^{+}08, Sab04, SAM09, SB01, SLMCRM05, SRS^{+}07, SSA08, SNB^{+}02, Syr00, TH08a, TCC08, UPK^{+}08, UP00, VSÅF05, VSS07, WGLJM04, Wei05, WHA08, WS06, ZPRJ02]. distribution [ZMM^{+}07, dPVJM04]. distributional [Ros05b]. Distributions [JMWJ08, BH08, HDG02, JH01b, LD05, MGS00, MMM00, MS09, MHV09, MNY^{+}09, PR03, SAMS02, Wal07, YM08]. disturbance [DDR^{+}03, Kou00, WWR^{+}08]. disturbances [KNO00, LVHU00, NKOK00]. Diurnal [Miy03, SMG02, CT07]. divergence [PSSD08, YMF02]. divers [DOBT02, LPSL09]. Diversity [Cal02, BR08a, Bia00, ERGT07, FGFP08, GLDD00, GCS09, GAA^{+}03, HMDS09, KHN03, LKK^{+}09, MVM^{+}08, XZW05]. Diving [Aka02, LD03a]. Division [BNBR05, GCM09]. d’Ivoire [BKN^{+}07]. DNA [BABB08, Ber00, CFRM08, GKOV05, HAvH06, KTM^{+}05, QGdS04, TYH04, WW01]. DNA/C [Ber00]. Do [BDJ^{+}05, BHR^{+}05, BO05, RR05b, Som04, TM09, TLM04, Agn08, BvKvH^{+}08, DHWW08, DCCS09, Gre08, JHL05, Lit06, SCLG00]. documentation [Hen04]. Does [HSdLP06, OH00, Roc00b, ZPRJ02, CH00, KMHS04, Uye00, VLJM^{+}07]. dogfish [TM09]. Dollard [Jag02]. dolphin [HSS^{+}05, RSC^{+}09]. dolphinfish [DNLSM08]. dolphins [Lóp06]. domestic [BBMS01]. dominance [JI05]. dominant [RNWS08]. Doñana [SGMMGB09]. doors [Som04]. Doppler [TZ03, ZKP03, ZPK05, ZCR09]. dormancy [JLR^{+}08]. dorsal [GO03b]. Dory [Dun01, YYY^{+}02]. dotted [RLdAW06]. double [BMU09, KTM^{+}05, dLMS06]. Dover [ERGT07]. down [CHB09, SL01]. downstream [JWBP07, WJB07]. Dr [Ano06j]. draco [Bag04]. Drake [CT07]. dredge [Gas02, GCC^{+}09]. dredged [WWR^{+}08]. dredged-material [WWR^{+}08]. dredges [GLS^{+}03]. Dredging [KPD^{+}07, BBR08, BLRC05, Des00, GRE06, GGM^{+}05, HSM00, HHSM03, KCCM03, MDM03]. Drift [BMLH07, DAd02, HSA^{+}01, HSPM05, JV05, RW01]. drifting [GLKPCP01]. Drilling [PAC02]. driven [HSA^{+}01]. driver [SRS^{+}09]. drove [GZS^{+}09]. Dual [HW08, HCEM06, MCP03, RR07]. dual-frequency [HCEM06, MCP03]. due [MSB04, SKS^{+}00, SPGT00]. Dungeness [BC07]. duorarum [ELR01]. during [AG00, BR04, BCL03, CF02, CSH00, CMO^{+}06, DGC00a, DGC00b, DLC03, Erm09, GGP07, GHI^{+}04, HA03, IA04, JWBP07, JNF^{+}09, KA01, MBPW06, MGH08, MVMH04, MAMO02, MW03, MSR03, OUNB02, OL07, PSO^{+}04, PMN01, Rob05, RE00, Ros09, Sab04, SBL07, SM02, Tan00, URMS04, WJB07]. dusky [PS09]. Dutch [BvS00, HF08a, RDHP00]. DYMONIS [LHKGS00]. dynamic [ADC^{+}08, DC01, IFUR08, KMJH01, LKL08, MHH06, MM01, MC07]. Dynamics [DF00, Ele00, SJ08, TM00, UA04, VHI^{+}04, Aco02, BBA03, BSMB03, Bla01, DP03a, DP03b, EKPT07, EH00, FMF02, FGBS00, FQS01, GJL08, GF01, GG04, HP07, HP05, HWF08, HK06, JP03, JHC09, Kar06, KV06, LNLS09, LMVdZ^{+}07, LN03, LT06, MS07, MMKKJ08, MC00, Nau02, Niw07, NFM^{+}02, Orł00, SPGT00, SSU^{+}09, SK07, SHS01, SMEK01, Ter02, TLMO08, VEP^{+}09, YSF09, YS02, vdMBD00].

Early [DBDA^{+}02, Bar05, BGAM00, Ber00, Ber04, BR00, CWC^{+}03, CBS^{+}06, CM00, GAYR06, GF01, GAW^{+}08, Hea07, His01, MHD02, NM09, OFN02b, OFN02c, OL00, PKRT06, Rob05, YMF02]. early-running [YMF02]. early-stage [CWC^{+}03]. earth [TM09, Ynd01]. East [VC02, Vil02, DLR02, DBS06, JSR06, She07, LLC^{+}08, PSSD08, RFM^{+}02, TYH04, YYY^{+}02, YCCH07, ZCH06]. Eastern [AGY^{+}05, RMAO^{+}03, Bar05, BFZ05, CMC^{+}06b, CCHV05, Cor00b, DWDD03, GGM^{+}05, GF01, GKFM09, HKBK09, JMLG05, KV06, KPD^{+}09, LMM^{+}08, LCRS08, LD05, LJM00, MMM00, MMD00, MS09, MC09, PJ08, PB05c, D00, RF01, SDWQ09, SB01, SPD00, SB03, VGBH09, Wal07, YBF^{+}03, Des00, EKPT07, Rad03, VCC07, LP00]. Ebrie [GAA^{+}03]. echo [Aks05, Aks06, BAB^{+}04, BNF^{+}07, DCRB09, DLS01, FB03, GPP09, GC05b, Kor00, NIF^{+}09, TK01, dR01]. echo-envelope [FB03]. echo-integrating [Aks05]. echo-integration [Aks06, NIF^{+}09]. echoes [Bet04, KHO06, MCAS04, SRJ03, TGS09]. echograms [KO03]. echoic [CD03]. echolocation [TPT^{+}09]. echosounder [BBS^{+}09b, CDM03, CD07, DH07, DR08a, DH09, DGO^{+}09, HSA05, Knu09, KHE^{+}09, MS09, Peñ08, wScY02, TMB08]. echosounders [BPT09, CCB^{+}06, GJH^{+}09, HPB09, JFCH05]. echotraces [Fer09]. ECOHAM1 [Mol00]. Ecological [DGO^{+}09, HE08, SLMCRM05, AMD^{+}05, Ard08, FSP05, GFP09, HHSM03, HFMD06, HR01, Mil02, dPM08]. Ecology [MGM03, AFP^{+}09, BGW05, CPR06, CRB08, HBG^{+}04, HR00, JJ06, LME02, MG07, MLS07, PVH^{+}05, D00, SDÖ09, SdlRdA06a, SdlRdA06b, SB00b]. Economic [Arn00, LT06, NM08, CA02, KBDC^{+}08, LO05, MLLK09, PKH^{+}08, RG07, SP05, UASN07]. economically [MB06]. Economics [Pas06, Kin02]. Ecopath [PCW00]. Ecosim [PCW00]. Ecospace [PCW00]. Ecosystem [Ano01h, Daa05, DKMO09, FPS06, GC05a, GS00, Hol00a, KGRW07, PH05, RPB^{+}08, TD00, VHI^{+}04, Arn00, BPM^{+}05, BBS09a, BJN^{+}06, BCL03, BAO04, BFZ05, CMDN02, CTLN09, CdlMA^{+}00, CUUD07, CC05, CMGS05, EH00, FSFO08, FDD^{+}05, GBC^{+}05, GSSO00, GM07, GWvM07, GFKM07, GR06, HCE^{+}03, HB09, HPR09, JR07, KYG03, KIDY09, Kos00, Kos09, Lin05, LAB^{+}05, Mis02, MMKKJ08, Mur00a, NJ04, NTSM07, OSK^{+}05, OGR^{+}07, Orl03, OLS00, PCW00, PST^{+}07, PJR08, PBH02, D00, RKP03, RCBM05, Ric09, Rob05, RR05b, RO05, Ros05a, SPS00a, SCJ00, SRJ^{+}05, SFH^{+}07, SGAC00, TL05, TMG^{+}08, TCP05, TLMO08, UC05, Uye00, Vil02, WCMK05, WPF00, YNX^{+}05]. Ecosystem-based [FPS06, BPM^{+}05, BBS09a, CTLN09, Kos09, NJ04, OSK^{+}05, OGR^{+}07, RR05b, SFH^{+}07, WPF00]. Ecosystem-sensitive [KGRW07]. Ecosystems [FLH04, WB04, Aga00, Bai09, Bea05, Bla00, BRHG^{+}06, CRB08, Cur00, HDG^{+}09, Hol00b, LT00, MJA^{+}05, STM^{+}08, SBDW00, TVH08]. eddies [KNO00, NKOK00]. Eddy [PZTE05]. edible [KPD^{+}07, Ste08, Ung07, UMSA09]. Editing [Daa03]. Editorial [Daa01a, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i]. Editors [Fra00]. edule [KPD^{+}07]. edulis [CFR^{+}01]. Edwards [Rob08]. edwardsii [GG09, LHHF03, MM01]. eel [ÅD07, AWW^{+}07, BWK07, BB07, BMC^{+}07, BVB^{+}07, Dek00a, Dek00b, Dek04, FMK07, JHKZ09, Jel07, KKS^{+}07, MV07, Sim07b, WLS07b, WJB07]. eelgrass [BRE^{+}08]. Eels [BG07, BGG^{+}08, BD07, DPW07, JWBP07]. Eems [Jag02]. EFA [CSdQB06]. EFAs [CSdQB06]. Effect [AE02, ASC01, BB07, Ber00, HD00, KKS^{+}07, MH01, MB05, Nie00, PPK^{+}06, PR03, SJGRRRE02, SPS00b, YW05, Bla01, BDS01, GMKS06, GG04, GO03a, GO03b, GFKM07, HFWB05, Iri00, JSR06, KBDC^{+}08, KB07, MLNC01, MR09, ODRN05, Orl01, PR01, RKP03, RPE^{+}09, RW01, SAAFCA07, Sar09, SDWQ09, TVH08, TLMO08, VCC07, VSÅF05, WCP08, YSF09, ZO03, Sec00b]. Effective [KFM02, GKOV05, HMAN03, RDD06, WHG07, WHP01]. Effectiveness [KS08, CSW06, DB08, GHD^{+}09, TSK00]. Effects [Aga00, APGD08, Bjö02, BÓ06, Bla00, CSR^{+}02, DH09, DBL07, EW07, Eno01, FLP^{+}02, FHJS09, GRE06, GS00, GF00, GZND02, HH01, Hol00a, HLSW01, HSS^{+}09, JH01b, KNKT06, LCRS08, LVHU00, LdSSG02, MMKKJ08, MS00, Mor02, PRvB00, RDHP00, SB00a, URMS04, VPC^{+}09, Vor00, War01a, AP07, ACD^{+}03, BTR06, BHN06, BB09, BMJ08, BRC09, BLRC05, Cad00, CG07, CRB08, CSVGTP09, Cor00a, CSdQB06, DGPR05, DRRS01, DCD00, DC03a, DC04a, FSP05, GTOJA06, GL00, GR05, GLKPCP01, GFP09, HEGL05, HC09, HH03, HRM04, HHJK06, HB09, HFMD06, HS09, IA04, Jag02, Jør03, LBL06, LC09a, LSH^{+}09, MML^{+}00, MMD00, MAAN09, PMN01, PPHB00, PF08b, PRF^{+}00, PDRG04, PPW^{+}09, PKRT06, RS03, RG07, RBGJ08, SKC09, SCJ00, SRJ^{+}05, SW06a, SBB^{+}05, SGMMGB09, SBDW00, SGAC00]. effects [TD00, TSK03, TS05, TDE08, UR01, Vec00, WBV09, Wie05, WHP08, YFL05, Zit01, Zwa00]. efficacy [CRIP08, WDRP09]. efficiencies [BDD06]. efficiency [BBR08, Esm06, Her04, LKL08, RKE06, TS05, WHG07, dHET04]. Efficient [WYM09]. Effort [íJCMR07, KF08, MWF^{+}05, APGD08, CR04, DCPvK07, EW07, GDL04, GCC^{+}09, GA05, GHFA09, GFP09, HBD05, HHJK06, HF08a, KS08, LZS09, LPM^{+}09, MAC^{+}07, MSF^{+}06, MRV^{+}08, MTJ^{+}07, PPB03, RS03, RDHP00, RDD06, RMAO^{+}03, VEP^{+}09, lPV01]. egg [Arm01, Bar05, BIdL^{+}08, DDM^{+}05, EGO^{+}07, GA00, MGTS00, OLB01, PCS^{+}07b, PCS^{+}07a, SPG^{+}04, vDBF^{+}09]. eggs [Agn08, CH00, CWC^{+}03, CVG08, GAW^{+}08, MYAT09, MV09, OR09, PCS^{+}07b, PCS^{+}07a, PÁMGV05, SDCR07, SSA08]. eiders [RLF01]. eight [PM04]. Eilat [AS02a]. Ekoreef [CA02]. elasmobrachs [MM03a]. elasmobranch [MF07]. Elasmobranchii [OV04, OV05]. elasmobranchs [MM03b]. electrical [YCCH07]. electronic [BHR^{+}05, vdKRS^{+}07]. electropositive [TM09]. ELEFAN [JAC00]. Element [Cas07, OMTS03, TL05]. elements [BGL08]. elephant [BHMS02, BHM^{+}04, AHS08]. elevated [EGO^{+}07]. Elminius [MWS04]. elongated [SC00]. elongatus [UPK^{+}08]. embiotocids [PSC02]. embryonic [CVG08, VGF03]. Emerald [FSS00]. Emerald/Western [FSS00]. emerging [FCM05]. Emperor [GHBR08]. emphasis [vdVBMR00]. Empirical [BPM^{+}09, MKB01, BDÑ04, DC03a, DC04a, MR09]. Empirically [PF08b]. encaged [GR01]. enclosed [Cad00, CJS02, dLMACC00, SP03a]. encounter [DHWW08, KMV^{+}07]. encrasicolus [Ber00, ZPI^{+}09]. End [CGS09, BTR06, KHM09]. End-to-end [CGS09]. endangered [AGY^{+}05, BDTW06, Pow00, XZW05]. endurance [BDO^{+}04]. energetics [BBBF02]. Energy [DSG05, DLC03]. enflata [GG08]. England [BBS09a, BWK07, BO05, Dun01, MB01, MWS04, PKP07, Tal07, TDE08]. English [CDR05, Des00, DWDD03, GLDD00, HBS^{+}06, HMPC04, RPR02, VCC07, WPB^{+}03]. Engraulis [Ber00, BDÑ04, BBGA05, KCL^{+}09, Miy03, STA^{+}09, ZPI^{+}09, ZWD08]. enhance [HRB02]. Enhancement [LPH^{+}08, AP07, Agn08, PS09, SFKC02]. enhancing [MHV09]. enough [BVDS08]. enriched [KNS^{+}06, PPK^{+}06]. enrichment [AS02b, BR08a, SAPP04, WHP01]. Ensis [DSG05]. ENSO [PSO^{+}04]. entered [BJN^{+}06]. entomelas [SKC^{+}00, SKH02]. Entrainment [DLT^{+}00]. entropy [BGW03]. Enumeration [GPWG04]. envelope [FB03]. envelopes [BNF^{+}07]. environment [ANNG01, BFM00, BJN^{+}06, CG07, CFL00, CSC^{+}04, DPN^{+}09, Dup05, Fra06, FN02, GF01, GAA^{+}03, HPBK04, MLM05, MHV09, PMB^{+}08, Pas06, PK07, PHDC^{+}09, PCD05, Ric00b, SHSKR01, SCLK01, UC05, WDRP09]. Environmental [CJM^{+}02, CA02, DCD00, Ele00, GH07, HSS^{+}05, HBC01, HLSW01, Joh08, LMVdZ^{+}07, RHH^{+}08, AFM^{+}09, Bac08, Bai09, BPD^{+}03, BG07, BO05, But01, CMM01, DRRS01, DM04, FGBS00, GAZ02, GA05, GLKPCP01, Jen09, JWM03, LDCH^{+}09, LSGD02, MMV^{+}08, MS01a, MSP09, MNY^{+}09, OL00, Ped05, PLP^{+}07, SKS^{+}00, SAMS02, WPB^{+}03, WHP08, ZPRJ02]. environments [CRTS04]. enzyme [OMBP06]. epibenthic [Cal02, RK04]. epibenthos [HD00]. epibiota [CJS02]. epifauna [CEV00, HFWB05, NRR^{+}09, RKE06, RK00]. epizootic [CDDM05]. equilibrium [Aco02, KM05]. Erie [RPE^{+}03, JRN06]. erinacea [FMF02]. Erratum [Ano01a, Ano01b, MGS01, OFN02a, PPHB01]. Error [BM02, Cor07, JRN06]. error-in-variable [JRN06]. errors [Hor08, Ree03]. ERSEM [Mol00, PST^{+}07]. Erxleben [CGN^{+}04, FPKH03]. erythrogramma [PJ08]. erythropterus [ZCH06]. Escape [ISHB07, Tal07, CW06, WW07]. escaped [HJ03, JH01a, OCWV06, RS06a, SW06c, SHAH09, WBC^{+}06, WC01]. escapees [GSS08, SWG06]. escapement [AWW^{+}07, BB07]. escaping [Rye04]. esmarkii [LNLS09, PK09, SLN02a, SLN02b]. essential [CSdQB06, Hel02]. establish [MMCD08]. Establishing [SCHR07]. establishment [HRB02]. ester [MGH08]. estimate [Arm01, DH07, EN02, HOP09, MMCD08, RFT02, Tje02, Wal07]. Estimated [PGB03, GTOJA06, SSKE06, UE01, VM07]. Estimates [GCM09, BPT09, BRC09, Bri02, CRvCB08, DSV^{+}08, ETB07, ERGT07, HHO08, Hor08, Jag02, JH01b, KCR07, KB07, LAO^{+}07, MR09, NTJ04, OGD09, PGMB09, PvHG09, RAB^{+}07, RT03, SS09, SHdLP04, TLM04, WHG07, WMS^{+}03, WRF09]. Estimating [BTR06, CMK09, CMP07, GOS07, HPB09, HC09, HDG02, LAO^{+}07, MMF09, MTJ^{+}07, O’D04, PCS^{+}04, RKE06, SK04, TM02, CD06, CF06, ES02, FLK^{+}09, FBF09, FGP07, HS01, HLL^{+}08, HMAN03, HBST02, LHHF03, Mac09, MS01b, NIF^{+}09, SP03b, VSC06]. Estimation [AK04, Coo04, Dem01, DLS01, DDM^{+}05, Har07, HSCN06, SKC^{+}00, UKR05, ZO03, AP09, BdP07, BJ00b, BJ00a, CRC^{+}09, DNP03, DCM03, EZ03, Fjä05, HSR01, KCBC00, MKB01, MM01, MSS^{+}05, PCS^{+}07b, PCS^{+}07a, PS03, RW01, RUN07, SPFF^{+}08, SSI07, SBP07, dPGPB06]. estimations [RPE^{+}03]. estimator [ASB05, VM07]. estuaries [Kup04, ZAJ01]. estuarine [Bla00, CMC^{+}06a, Cor00a, HCV03, HAG^{+}08, RUCG07, RBGJ08, RD01, RD03]. Estuary [HLL^{+}08, Ber04, BBB06b, But01, FR04, HKI01, Jag02, LMC^{+}01, STM^{+}08, NCC^{+}07, ROB04]. Ethmalosa [GAA^{+}03]. EU [DMvD07, FTDVC^{+}08]. Eulalia [Mor02]. Euphausia [AF06, DSJ03, HTA09, HHKL04, RCLD08]. euphausiid [CSH00, DSJ03, KHE^{+}09]. euphausiids [SC00]. Europe [CCB^{+}06, FMH^{+}09, GZS^{+}09, Jen02a, dLMACC00, PPL^{+}07, vDM07]. European [AKJ07, Agn08, AWW^{+}07, BWK07, BOC^{+}08, BdP07, BMC^{+}07, CBS^{+}06, Dek00a, Dek00b, FMK07, GAA^{+}04, JHKZ09, KBW09, LFD^{+}09, LD05, MV07, MAC^{+}07, NM08, OCWV06, Pen07, PJR08, PS03, PF06, SGMV^{+}08, Sim07b, SEOR09, SPG^{+}04, Ste08, SS07, SGMN^{+}06, Sym07, WPR^{+}07, YG08, ZPI^{+}09, dPBB^{+}03, dPGPB06]. Euthynnus [BKN^{+}07, GKFM09]. eutrophication [BdMAL00, RTDJ09, RRY08]. evacuation [RR02a, TM02]. evaluate [BHR^{+}05, OSWL02, SMI07, SRJ^{+}05, UASN07]. Evaluating [CRIP08, DC01, PJR08, RP07, Ric00a, SZ07, TSK07, TLMO08, WS02a, WRF09, HH03, MSH07, PCW00, KDCH^{+}09]. Evaluation [ASB05, CFL00, FB02a, Kat05, KPS^{+}05, MDM03, Rad03, RT03, RRC03, Sim09, TZ03, APD09, Bet04, BML^{+}05, CEV00, DB08, FSS00, GD05, GAM^{+}06, HBD05, KPK^{+}05, KPK^{+}06, KMG^{+}07, Knu09, KPD^{+}09, NIF^{+}09, Pie00, RR09, Sea02]. event [BR08a, PSO^{+}04]. events [GZND02, MPG^{+}09, War01a]. Evidence [BGG^{+}06, CLK^{+}09, CHB09, HM04a, HWF08, SWG06, WGMM08, dPGPB06, BK07, Buc00, DGPR05, IB00, Iri00, PSSD08, RPB^{+}08, Rye04, Ste02, SMP09, TA05, TB02]. evidenced [PZTE05]. Evolution [FGLT02, GD05, FB02b, Law00, NPPO06, WPR^{+}07]. evolutionarily [Fik00]. evolving [VBF09]. Ex [KH03a, KCL^{+}09, XZW05, BW08]. examination [BRP02, MMD00]. examine [KMM07]. example [BNBR05, Cla00, KPD^{+}09, MKR^{+}09, O’D03]. examples [HM04b, Kas09, KHS^{+}08, KHE^{+}09, LLD^{+}05, RPE^{+}03]. Exceptional [BJN^{+}06]. Excess [Rho08]. exchange [NNT01]. excluder [EHL07]. excluding [GHD^{+}09]. exclusion [SBD^{+}09]. excretions [ASC01]. exercise [KNKT06, SKC^{+}00]. exercises [LGH^{+}09]. existence [JNF^{+}09]. exitiosa [CDDM05]. expanded [Ona03]. expansion [KHS^{+}08]. expatriated [KMT08]. expected [TMB08]. experience [ETB07]. experiment [CBDB02, JO02, KIDY09, dPBB^{+}03]. Experimental [BD07, CVG08, Orl05, SOB^{+}07, ZNGF02, BG04, De 04, DTC01, FHJS09, HHSM03, HFMD06, Iri00, PRF^{+}00]. experimentally [MCRF06]. Experiments [MCAS04, BIdL^{+}08, BHR^{+}05, BW08, Dav07, EZ03, HU04, HRB02, RKE06, SK04, UE01, URMS04]. expert [TH08b, UKR05]. explain [YM08]. Explaining [SCLG00, Kup04]. explicit [Aco02, HHMN01, MHD02, RR05b, RD01]. explicitly [MLOT09]. Exploitation [D00, ASB05, BFZ05, Dem01, Dun01, FMF02, GAYR06, HMK^{+}07, Rad03, dBP02]. Exploited [DF00, FLH04, WB04, AGA^{+}04, LdSSG02, MF07, Niw07, NPPO06, PRD^{+}06, RTB^{+}05]. exploiting [Her04, RvMBV00]. exploration [LSH^{+}09, dLMACC00]. explore [SLvdB^{+}09]. explored [BBBF02]. Exploring [PG08, YFL05, YSF09, HNK07]. exponential [BMM03]. export [LC09b]. exposure [BI08, BLMB06]. expression [SMK08]. extending [EKPT07]. Extension [KNO00]. extensive [LPH^{+}08]. external [Pie02, RW01]. extinction [RL07]. Extracting [ZCR09, BKR09]. Extraction [MCL03, BLRC05, Des00, SPGT00, vDEM^{+}00]. extreme [MML09]. extremes [RMM05]. Exxon [TT08].

faber [Dun01, YYY^{+}02]. face [BP08, dVA07]. faced [MSI07]. facilitate [MHD02]. facing [AGH^{+}09, ADDH04, PK07, Pet04, SAPP04]. factor [ADC^{+}08, DLC03]. Factors [AG00, AO08, STAN02, STJ^{+}07, WSP03, AFM^{+}09, BO05, CF02, CDD^{+}07, FGBS00, HH03, MS01a, MNY^{+}09, ODRN05, TCC08, TDE08]. factory [Dor01]. faecal [UR01]. failing [FGP09]. failure [PHDC^{+}09, PCRW04, Sim07a]. Falkland [ABB^{+}08, RUA07]. fall [YCCH07]. false [HHAB09, WBK^{+}09b, WBK09a]. families [KTRG06]. family [MC09, YSO^{+}03, YTS^{+}06]. Fangatau [AGY^{+}05]. fanskate [OV04]. Farfantepenaeus [ELR01, PCD05]. farm [GSS08, HH01, SWG06, SHSKR01]. farmed [BJN^{+}06, CW06, DRRS01, FLH06, HJ03, Han06, JH01a, JKSO06, KNKT06, OCWV06, SW06c, SHAH09, SMK08, WBC^{+}06, WSC^{+}06, WBD^{+}06]. farming [AS02b, BDTW06, BRHG^{+}06, FLH06, HFMD06, KTH^{+}00, LBL06, MML^{+}00, MSIL09]. farms [BF02, CMM01, FJSJBS^{+}08, Han06, MHF^{+}09, RLF01]. Faroe [Gaa00, GH00, HJ03, HEGL05, íJCMR07, MGH08, SG05]. Faroese [íJR02, PSSD08]. fasciatus [SRGC04]. fast [GG09]. fatty [CSdQB06, JGN04]. fauna [DBL07, Eno01, FSFO08, HH01, KPD^{+}07, Pie00, RRTP02, TK03]. faunal [HS09]. Feasibility [LTA00, TM09, WW07]. features [Bro02a, FAL^{+}08, LDML08, LD03a]. featuring [GRMR07]. Fecundity [Agn08, EK08, ÓT06, Bri02, DRDC06, GCM09, SRMB07, Ter02, TAC00]. fed [KNS^{+}06]. feed [BSO01, KNS^{+}06]. feedback [KS08, MMKKJ08]. Feeding [CCA04, Har00, HHHH06, JH01a, JR06, MG07, MJB08, BKN^{+}07, Bjö02, BGW05, BWC00, CPR06, Cor00b, DEMD00, GMM^{+}08, GG08, GHI^{+}04, IA04, JRCS08, KWZ00, MGvH06, MH01, OKRK04, OGL06, OUNB02, Ped05, PVLP04, PVH^{+}05, D00, RS06a, RAKS06, SB00b, SPK05, SVRF08, TM02]. fees [PS09]. Female [HL09, BHM^{+}04, Cos09, FPKH03, MGH08, NTSM07, dLMS06]. females [Agn08]. feral [Cip09]. ferruginea [DWC03]. fertilization [RHBR04]. few [Pay04, Sin09]. Fidelity [SPK05, BBMS01, FSDC09, WSFH02]. field [Ber00, BD07, CBDB02, Iri00, LSH^{+}09, MBM02, TM02]. fields [Kal01, KYG03]. Fifth [PB08a, CM00]. fimbriata [GAA^{+}03]. fin [LG08]. Finding [PCM09]. Fine [LD03a, GLDD00]. fine-sand [GLDD00]. Fine-scale [LD03a]. finfish [But01]. fingerprinting [SVRF08]. fingerprints [BGG^{+}08]. Finland [DHKV01, PVLP04, PLP^{+}07]. finless [Aka02, XZW05]. finmarchicus [DGC00b, Buc00, CH00, Cor00b, CM00, DK00, DGC00a, Fik00, Gaa00, GH00, GAP^{+}00, GA00, GP00, HHB^{+}00, Hea00a, Hea00b, HBG^{+}04, Hel00, HRHC00, JLR^{+}08, Kaa00, MGH08, MCM00, Nie00, OED^{+}04, PTTS00, PB00, PPH09, PS06, SOMT00, SBB^{+}05, TM00, TH05, URMS04]. finned [AJR00, DCD00]. First [OB05, AG00, BD03, IWP00, KWZ00, PG08, VPC^{+}09, VH08, WWGG02, YW05]. first-feeding [KWZ00]. first-principles-based [PG08]. first-time [YW05]. Firth [BML^{+}05, TAC00]. Fish [FLH04, GAZ02, HW06, KSD01, Lun01, MML^{+}00, PSC02, RKKM06, WB04, WGM04, ZPK05, AS02a, APGD08, AFP^{+}09, AS02b, ASB05, ADDH04, AOSD09, BEB^{+}09, BHN06, BS03, BPT09, BJ00c, BDÑ04, Bet04, Bia00, Bla01, BDJ^{+}05, BRC09, BGG^{+}06, BBB06a, BPWS09, BH08, BFMJ03, BNF^{+}07, CTM09, Cal02, CCHV05, CSR^{+}02, CW09b, CTF02, CMJ09, CDM03, CGS09, CKS03, CD07, DGPR05, DRRS01, Dav07, DHWW08, DLS01, DH09, DCM03, DBC03, DPN^{+}09, Dup05, ES02, ES03, ES09, EW07, EZ03, EJR01, FS02, FGLT02, FLK^{+}09, Fer09, FJSJBS^{+}08, FLH06, FGP07, FGFP08, GD05, GL00, GH04, GCS09, GPRD08, GPP09, GLR06, GOK05, GJR04, GR06, Gre08, Gro06, Gud04, GLDB04a, GLDB04b, GAA^{+}03, HFWB05, HSR01, HS01, HCV03, HW08, Han06, HOP09, HM04a, HH03, Hea05a]. fish [Hea05b, Hea07, HMD^{+}08, HH01, Hel02, HHT08, HHMN01, HEGH02, HD00, HBW^{+}09, HCEM06, HSS07, Jaf06, Jaf08, JSR06, JR06, JHC09, JJ06, JLS02, JB00, JI05, Kaa00, KFM02, KTH^{+}00, KH03b, KRYL09, KHEJ09, KHM09, LLD^{+}05, LBF01, LS04, LSGD02, LN03, LD03b, LT06, LdSSG02, LW04, MCAS04, MLNC01, MPG^{+}09, MKR06, MHD02, MR09, MS07, MKFK05, MHV09, Mor02, Mor04, MB06, MNCU09, MNY^{+}09, MV09, NJ04, NIF^{+}09, Niw07, OM05, OGD09, OK05, OKRK04, Orł00, Orl01, Ost09, OLS00, OL07, PLP^{+}07, PCS^{+}07b, PCS^{+}07a, PGJ^{+}05, Pet01, PGMB09, PR04, PJ05, PvHG09, PRD^{+}06, PPMH04, PB05c, PPW^{+}09, D00, RRTP02, RPE^{+}09, Roc00a, RTB^{+}05, RE00, Ros05b, RUCG07, RBGJ08, RCL05, RKM09, Rye04, Sab04, SB06, SJGRRRE02, SMG02, SAAFCA07, SMB09, SLvdB^{+}09, SDCR07]. fish [SF09, SB00a, STG06, SP05, SB04, SSJL02, SLMCRM05, SBL07, SHSKR01, SCLK01, SP07b, SRJ03, SAMS02, Ste01b, SJM03, Sve03, SBD^{+}09, TMI^{+}04, TNF09, TF04, TES^{+}05, TK01, TPRR04, TLM04, TMG^{+}08, TC01, VHF^{+}04, VCC07, Vin01, VSC06, WMS^{+}03, WCP08, WDRP09, WMÖ06, WWHB04, YTS^{+}06, YFL05, YSF09, YM08, ZCR09, ZO03, Zwa00, CLFS02, Rot00]. fish-abundance [WMS^{+}03]. fish-aggregating [DPN^{+}09]. fish-count [HCEM06]. fish-farm [HH01]. Fish-farming [MML^{+}00]. fish-processing [OM05]. fish-school [BPT09, Fer09, NIF^{+}09]. fish-species [CTM09]. fish-stock [PGMB09]. fish-tag [ES09]. fish-tagging [STG06]. fished [WHG07]. Fisher [RG07, BDÑ04, DAH^{+}08, HMK^{+}07, KPD^{+}07, RPB^{+}08]. Fisheries [DKMO09, Ele00, GAFA06, KCL04, LT00, MGM03, RC07, Sym07, Aga00, AKLL07, AFP08, Arn00, ARMM09, BvS00, BDD06, BBBF02, BD04, Boo00, BZRO06, Bun01, Cad00, Cha04, Cla00, CGV03, CdlMA^{+}00, CH05, CC05, CMGS05, DM07a, DM07b, DMvD07, DSV^{+}08, DAH^{+}08, DBL07, FB07, FSB^{+}03, Fjä05, Fle05, FFL06, FPS06, Fur02, GC05a, GS03, GSSO00, GAA^{+}04, GH07, GAM^{+}06, GRMR07, HMK^{+}07, HT05, HCE^{+}03, HNK07, Hea05c, Hoy07, HLS00, HMPC04, íJCMR07, JR07, JRN06, JvD07, JI05, KGRW07, Kat05, KM02, Kos00, LLD^{+}05, LGH^{+}09, LS04, LZS09, LT06, LO05, MCB09, MFD02, MKR^{+}09, MNHL01, MUK^{+}02, MLM02, MSW07, MPG^{+}09, MR05, MS00, MB06, MCP03, MLLK09, ÓMP^{+}04, Pas06, PR07, PCW00, PKP07, PFK^{+}09, Pet04, PST^{+}07, Pet03, PMB^{+}03a, PSFY07, PCS^{+}04, RBD^{+}07, RBBB00]. fisheries [RR05a, RR07, RL07, Ric09, Ric00b, RDD06, RR05b, RSC^{+}09, SBC^{+}07, SFH^{+}07, SJ08, SS07, TZ03, TMB08, TCP05, UA04, UP02, Vor00, WBC^{+}06, WYM09, WGLJM04, WLK02, WPF00, YNX^{+}05, Ye00, dHET04, lPV01]. fisheries-acoustic [Pet03]. fishers [CWYM^{+}02, Her04, vDM07]. Fishery [BS02, GC07, GM07, SE09, APGD08, APD09, ATH^{+}07, BPM^{+}05, Bag04, BSS07, BBS09a, BQHG00, BFM00, BM01a, BM01b, BDÑ04, BSMB03, CRIP08, Cla00, CvdLHF08, DB04, DAAD09, EHL07, ET07, Eig09, Esm06, FLK^{+}09, FJK^{+}07, FWW06, GDL04, Gas02, Gat00, GWSV08, GF00, GM06, HM08, HB07, HTA09, HF08a, HPR09, Hol03, HMQ^{+}08, ISHB07, JGM^{+}08, JTE^{+}07, JMC07, JSMK06, Kos09, KBDC^{+}08, LCRS08, LCC09, LFW03, MKB01, MNMG^{+}05, MFA07, dLMACC00, MNCU09, MRT01, O’D03, ÓMP^{+}04, OH07, OLS00, PPL^{+}07, Pen07, PJR08, PCS^{+}04, D00, Ric00a, RGG^{+}04, RMAO^{+}03, RD07, RUCG07, RPSSW09, RD03, RK00, SPS00a, SMH09, SMI07, SCHR07, SP07a, SKR^{+}06, SP05, SJM03, STJ^{+}07, Tal07, TCS^{+}09, TR09, TB02, VBSB07, VEP^{+}09, WBV09, WHG07, WPB^{+}03, WM01, YCCH07]. fishery-acoustic [FLK^{+}09]. fishery-dependent [WHG07]. fishery-independent [LCC09, MFA07, RD07, RUCG07]. fishes [MMC03, MH01, RRTdA02, RR02a]. Fishing [Dor01, GL00, GS00, GWvM07, GFKM07, HM04b, Hol00a, Law00, PCRW04, ARMM09, BHMD05, BDD06, Bia00, Bla00, Bla01, BDJ^{+}05, BDO^{+}04, CBBL09, CG07, CR04, CEV00, DGPR05, Dav07, DCPvK07, DDR^{+}03, FHHH00, Fru02, FSP05, GCC^{+}09, GM07, GHI^{+}04, GJR04, GHFA09, GFP09, Hea05a, Hea05b, HHJK06, HMAN03, Hol00b, Hor08, HMPC04, JYW09, KGRW07, KDP09, LHHF03, LC09a, Lun01, MSB04, MM07, MNHL01, MUK^{+}02, MAC^{+}07, MSH07, MMKR^{+}00, MLMC02, MTJ^{+}07, Peñ08, PJ05, PQRG07, PvHG09, PMD^{+}00, RRT00, RDHP00, RDD06, SCJ00, SDWQ09, SRJ^{+}05, SW06a, SPD00, SBT^{+}09, SKC^{+}00, SRM08, SBDW00, SGAC00, TSK07, Tas00, TD00, Vec00, YFL05, Zwa00, lPV01]. Fitting [LEP04, Cot01]. five [OED^{+}04]. fixation [GAW^{+}08]. fixed [ES03, HMD^{+}08, PÁMGV05]. fjord [Bjö02, GSS08, HSS07, NTSM07, SHAH09, SOMT00]. fjordic [CRTS04]. fjords [BA03, BFK^{+}07, LC06, LGR08]. flatfish [KPS^{+}05, NWH02, RvMBV00, WWHB04, vdVBMR00, vdVBMR00]. flats [KPD^{+}07]. flavescens [GWSV08, HA03, JRN06]. flawed [Le 09]. fleet [DAAD09, GF01, HM04b, RBD^{+}07, RDHP00, RPT02, UASN07, VEP^{+}09]. fleet-based [RBD^{+}07, UASN07]. fleets [MAC^{+}07, Mar08]. Flemish [SRGC04]. flesus [Jag02, KBW09, NWH02]. flexibility [UA04]. flexible [AP09, LKL08, MBC^{+}09]. flexuosus [CDB05]. Florida [ASB05, ELR01, Kup04]. flounder [CS05, DWC03, Jag02, KASA07, KNS^{+}04, KBW09, MFIO04, NWH02, WM04]. flounders [NCC^{+}07]. flow [BP08, JWBP07, JV05, MBM02]. FLR [KMG^{+}07]. Fluctuations [VBSB07, RF01, SP03a]. fluid [SC00]. fluid-like [SC00]. fluorescent [URMS04]. flux [HKI01]. fluxes [CJS02]. fly [KTS02]. flying [CZC07]. foes [Law08]. following [AMGV06, GGM^{+}05, GAW^{+}08, WSC^{+}06]. Food [Alo01, BKN^{+}07, BANGC02, CPR06, DEMD00, GTOJA06, Kin02, OUNB02, PVH^{+}05, BD04, CRW^{+}01, CDD^{+}07, DLC03, HCE^{+}03, HPBK04, Iri00, MAMO02, SB00a, URMS04, YW05]. food-deprivation [DLC03]. foodweb [BCL03, CGS09, Dol02, GGV^{+}04, Hea05a, Hea05b]. footrope [MS01b]. forage [TLMO08]. foraging [BHMS02, OKRK04, WHP08]. forcing [DiUVH08, DAd02, GP00, Tri00]. forecast [MLM05]. Forecasting [MML09, PBH02, PCS^{+}04, SHT^{+}09]. forecasts [VBSB07]. forest [TD00]. forever [Daa03]. Foreword [Daa01b]. form [GH07, JHL05]. formaldehyde [GAW^{+}08, PÁMGV05]. formaldehyde-fixed [PÁMGV05]. formalization [SB04]. formation [BR04, HHKL04, PZTE05, SBB^{+}05, WWGG02]. formula [Mac09]. forum [GM07]. forward [GM07, MHH06]. fouling [CBS^{+}06, HF08b]. four [CW09b, CMM03, DJRO06, KNS^{+}06, OAJ06, Roc00a, SF09]. Foveaux [CDDM05]. fractal [Dek00a]. fraction [GCM09]. fractionated [IA04]. fractionation [HOF04]. fragile [Kos00, SYR^{+}08]. fragmented [CW09b]. Framework [FTDVC^{+}08, AKLL07, BBR08, DDGR07, GD05, HOHS05, HNK07, HBW^{+}09, HK06, KMG^{+}07, LAB^{+}05, OSWL02, RR05a, TVH08, TCP05]. France [BMV05, EHG06, AVJ^{+}06, RTB^{+}05, lPV01]. free [Bjö02, CGN^{+}04, NNT01]. free-ranging [Bjö02, CGN^{+}04]. freezer [BvKvH^{+}08]. freezer-trawlers [BvKvH^{+}08]. French [AGY^{+}05, AFGR09, SF09, CSR^{+}02, Des00, DP03a, DP03b, GAYR06, HM04b, Mar08, RPT02, WPB^{+}03]. frequencies [GLDB04a, GLDB04b, KM05]. frequency [AK04, BB09, BPT09, CH09, CSH00, CDSC05, FLH06, HW08, HPB09, HDG^{+}09, HCEM06, JAC00, JPO09, KO02, KO03, LEP04, LW04, Miy03, MCP03, PK09, RML06, UE01, WSWS03]. frequency-dependent [LW04]. fresh [BD07, Jel07, PK07]. fresh-water [Jel07]. freshwater [Aka02, CW06, GLDB04a, GLDB04b]. friderici [DCN^{+}04]. front [CMC^{+}06a, GQCÁMI03, LND05, MFB^{+}08, NFM^{+}02, UR01]. frontal [BWC00, KNO00, NKOK00, SBC^{+}00]. frontier [dHET04]. fronts [ORVP09]. Fuel [STW^{+}08]. full [ÅD07]. fullness [GMM^{+}08]. fulmars [CRW^{+}01]. Fulmarus [CRW^{+}01]. function [BNF^{+}07, Hea05a, Hea05b, NT02, Orl03, PM06, vdKRS^{+}07]. Functional [GCS09, Bai09, LFD^{+}09, SCWD08]. functionally [GWF01]. functioning [Bla00]. functions [AP09, BP07, TM02]. fund [PS09]. Fundy [LMM^{+}08, BDTW06, ML08, PFF01]. furcatum [GC02]. furnieri [dCA03]. Further [PR01, PGD09, ZPI^{+}09]. Furtive [Law08]. future [AVK^{+}08, CRB08, Dri05, FSB^{+}03, HBW^{+}09, Jen02a, KTT06, WBC^{+}06]. Fuzzy [BSA09, MvdKN05]. fyllae [SB01].

G. [ISS^{+}07]. Gadoid [KTT06, BHN06, BBB06a, GAW^{+}08, RS06b]. gadoids [JGST09, KMHS04, LN08]. Gadus [Arm01, AGA^{+}04, BGL08, BBM^{+}02, BSO01, BD04, BÓ06, BDS01, BDTW06, CCC02, CSdQB06, DJRO06, DMvD07, Dri05, EKPT07, ET07, Erm09, FL06, FR09, GW04, Ham06, HOF04, Joh02, JDN01, JCM06, JKSO06, KNKT06, KPO05, KTRG06, KNS^{+}06, KPD^{+}09, KMJH01, MGTS00, MGvH06, MW03, MSS^{+}05, MR05, MSR03, MJB08, MB05, MG02, OMBP06, OSLO06, OFN02b, OFN02c, PPK^{+}06, PK09, Rad03, RL05, RML06, RR06, RHBR04, Sin01, SMK08, SBB^{+}05, Som04, SSA08, SPWHR04, SB03, SPS00b, SFM01, TAHK06, UP00, VGBH09, VSC06, WSC^{+}06, YM00, YW05]. gahi [RUA07]. Galatheidae [LCC08]. Galicia [BCL03, SCCM06]. gallina [MDM03]. galloprovincialis [PBLFR06]. GAM [MNY^{+}09]. Gamma [HBST02]. Gamma/Dirichlet [HBST02]. gammarus [AKJ07, Agn08, LFD^{+}09]. GAMs [Pie02]. gannet [MNCU09]. gannets [Bun01, CDD^{+}07]. gap [OSK^{+}05]. gas [FGLT02, Hel02, TPT^{+}09, War01b]. gas-bearing [War01b]. Gasterosteus [PVLP04]. gastric [CGN^{+}04, TM02]. Gastropoda [SDÖ09]. Gauging [PMD^{+}00]. gauntlet [Dup05]. Gaztelugatxe [BLMB06]. gear [DAAD09, aFADN08, GFKM07, JR07, LCRS08, Lun01, RRT00, TSK07, TM09]. gears [BDO^{+}04, GA05, RAR^{+}07]. gelatinous [CMM03, SB00a]. gene [CMO^{+}06]. General [DLR02, HS01]. generalist [VLJM^{+}07]. generalization [Ric09]. generalizations [Mye01]. Generalized [BP07, BR02, MNY^{+}09, PH03]. generate [Bro03, GHFA09]. generated [KO03]. Genetic [AMJ^{+}06, BHN06, CFRM08, CMC^{+}06b, DJRO06, GSS08, HFMD06, JDN01, PSSD08, PÁMGV05, RMM05, SBG06, SCCM06, TA05, BBM^{+}02, Buc00, CDQL06, CLK^{+}09, DC05a, ERGT07, FCM09, HTSB04, HAvH06, KHN03, MV07, MASA06, PGD09, SWG06, SRS^{+}09, TYH04, UMSA09, ZPI^{+}09, ZCH06]. genetically [JHL05]. genetics [Box06]. genome [GMGN06]. genotypes [JKSO06]. genus [Har00, HOGH07]. geo [BHMS02]. geo-location [BHMS02]. Geographic [SRN00, SSU^{+}09, Str05, CLM07, Ped05, SSC^{+}06, BBBF02]. Geographical [TF02]. geometric [aFADN08]. geometry [CR04, Dek00a, OH00, TSK07]. geophysical [TST^{+}09]. Georges [OL07, CGS09, DGC00a, DGC00b, GL00, Hol03, Kan07, MLNC01, MS07, TCS^{+}09, TC01]. Georgia [SHT^{+}09, Col02]. Geostatistical [AHS08, Wal07, WRF09]. geostatistics [RUN07]. German [PFK^{+}09]. Germany [Sim07b, JHKZ09, WBV09]. Gerres [GAFA06]. ghost [TSK07]. Gialova [Kou00]. giant [AFGR09, DDGJ02, GAYR06]. Gibraltar [Her04]. gigas [HKI01, HPBK04]. gill [PGG05]. gill-infesting [PGG05]. Gillnet [MSW07, aFADN08, HLS00, LGH^{+}09, LHJS02, LO05, TSK07]. gillnets [CBDB02, Lóp06, MANT07, UBP^{+}09]. Girella [ISS^{+}07]. Gironde [Ber04]. GIS [BBBF02, BM02, BMDBM09, FGD02, ZPRJ02]. given [RD07]. glacial [NTSM07]. glacialis [CRW^{+}01]. gladius [CPR06]. glass [BB07, BD07]. Global [RTDJ09, SYR^{+}08, VHI^{+}04, GMGN06, HB09, KHS^{+}08, MMKR^{+}00, NM09, PBM^{+}04, Sea02, VM09]. Gmelin [BLMB06, SBT^{+}09]. go [JWBP07]. goals [Mil02]. goby [KASA07]. Going [BP08]. golden [Str05]. gonad [CF06]. gonads [Jør03]. good [Le 09, RO02]. goose [BLMB06]. gouldi [TJAS04]. governance [GH07]. government [Ray07]. gradients [PCDM08]. grading [Kin02]. Grand [DWC03]. grappling [BBS09a]. gravel [BLRC05]. grazing [FB02b, LDNS08, TDE08, URMS04]. Great [RPSSW09, ATH^{+}07, BSMB03, HS06]. greater [Bag04]. green [NAK^{+}08, YG08, YBF^{+}03]. Greenland [RFM^{+}02, VC02, Vil02, ANNG01, AFHJ04, FRK02, HSCN06, HAN02, JRM^{+}03, KKC04, LHHJ^{+}09, LHJJ^{+}01, RL05, SPWHR04, Wie05, WB05]. gregaria [LCC08]. Greifswalder [ODCN09]. grey [BMM03, FT05, JSMK06]. grid [HB07, MFIO04]. gridded [HHC^{+}09]. griseocauda [ABB^{+}08]. groenlandicus [CGN^{+}04, FPKH03, HSCN06, PGB03, SFØ07, Ste02]. ground [BML^{+}05, EGO^{+}07, JYW09, RR02b, SPD00, WSP03]. ground-discrimination [BML^{+}05]. groundfish [BO05, FGP07, FGP09, GSdFB01, Hol03, LBNS00, MS09, MM05, She07, WPF00]. grounds [DRSD09, FHHH00, HKBK09, MJB08, OR09, RMM05, SW06a, SPK05]. groups [AJR00, JCM06, SGMV^{+}08, WSC^{+}06]. grow [DDGJ02]. growing [TSH^{+}06]. Growth [AKJ07, AJR00, BHØ^{+}04, GDH02, HSS07, JRCS08, KKF^{+}06, LDQ08, Mar07, MMS01, OMBP06, OSLO06, PB05a, PJ08, PCD05, PBLFR06, PFLFR08, TSH^{+}06, UE01, ZAJ01, AFP^{+}09, ACD^{+}03, AGA^{+}04, ASC01, BPD^{+}03, BSO01, Bjö02, BÓ06, BMM03, Bra07, Bro03, BD03, BK07, CAAJ07, CDR05, CF06, CRTS04, DWC03, FM04, FHJS09, FPKH03, FHDM00, FCM05, FQS01, GPRD08, GG09, HP07, Har00, HS06, HC09, HLL^{+}08, JAC00, JKSO06, KNKT06, KCR07, KCBC00, LNLS09, LP00, LC06, LEP04, LCC07, Mil08, MKFK05, MNCU09, OH00, ODCN09, OFN02b, OFN02c, OL00, PPK^{+}06, PM06, PS03, PF08b, RS04, RL05, RMB^{+}09, SRGC04, SA05, SNV^{+}09, SPFF^{+}08, Sim07b, SW06b, SCLK01, SGS^{+}05, TAHK06, VLBB08, VSÅF05, Wie05, YBF^{+}03, YYY^{+}02, dPGPB06]. grypus [JSMK06]. guatucupa [LO05]. guggenheim [CLM07]. guide [HB09]. guidelines [GC05a]. guild [GL00]. guillemot [LPA^{+}00]. Guinea [WYMF08]. Gulf [APD09, BCD^{+}02, BW08, BGG^{+}06, Bro02b, BR08b, CDM03, CCC02, DBDA^{+}02, Dup05, DR08b, Esm06, FBD^{+}08, GAFA06, HB07, HHMM01, HIL00, HAvH06, HL07, JSMK06, KK06a, KCCM03, Lit06, LDNS08, MM05, NAK^{+}08, OL07, PVLP04, PLP^{+}07, PGJ^{+}05, SBG06, SW06a, Sin01, SGY08, SPS00b, SFM01, TM09, WCP08, WYMF08, dHET04]. gull [VLJM^{+}07]. Gullmarsfjord [ETB07]. Gullmarsfjorden [LVHU00]. Gunnerus [HRHC00, MGH08]. Günther [MG07]. gurnard [FT05]. gut [KNS^{+}06].

Habitat [AFP08, HCV03, HBS^{+}06, MB06, WCP08, WSFH02, BMV05, BI08, BMJ08, BML^{+}05, CLFS02, CSR^{+}02, CW09b, CTLN09, HSS^{+}05, Hel02, HL07, HAG^{+}08, HS09, JDA^{+}06, KCD^{+}03, KMV^{+}07, LdSSG02, SW06a, SBT^{+}09, SP07b, VLJM^{+}07]. habitats [AGH^{+}09, AGC02, BML^{+}05, CMN^{+}07, CMJ09, CKS03, DK00, EHG06, EGB02, FSQ^{+}03, HSM00, NRS09, SOMT00, SN08, VCC07, Jen02b]. habits [Alo01, BKN^{+}07, JH01a, JR06]. haddock [ATM02, BNBR05, BDO^{+}04, CH06, FMB01, FSS00, GML06, HLS00, ISHB07, OL00, ÖG04, ÖFR^{+}06, PKRT06, RW01, TSH^{+}06]. hairtail [Zha06]. hake [BOC^{+}08, BdP07, CSH00, CFMdP07, DC05a, ERBP09, GAA^{+}04, GLR06, MAB^{+}07, PÁMGV05, PS03, SG00, SHS01, SGMN^{+}06, WPR^{+}07, dPBB^{+}03, dPGPB06]. half [SBL07]. halibut [ANNG01, HAN02, JRM^{+}03]. Halichoerus [JSMK06]. Hampshire [CH06]. hangs [LD03b]. Hansen [DSJ03]. haplotype [TCTC09]. harassment [FWW06]. Harbor [SP02]. harbour [ATH^{+}07, BBBF02, CBDB02, LHJJ^{+}01, Lun01, TPT^{+}09]. harbours [MMV^{+}08, MVM^{+}08]. hardly [KDP09]. Hardy [KM05]. harengus [HP01, PH03, RLH01, BRP02, BD02, BD03, BBK08, BBSK09, CA00, CCA04, CMHN05, CBHM07, DEMD00, EDG03, Fox01, GFH04, HK00, HSS^{+}09, Joh02, JHL05, LMC^{+}01, LHR02, NTJ04, OL07, PHG04, PHO09, PVLP04, PB05b, PN06, STAN02, SDCR07, TK03, Vuo02, WPM^{+}09]. Harmful [RMKT01]. harp [CGN^{+}04, FPKH03, HSCN06, PGB03, SFØ07, Ste02]. Harris [WBK09a]. harsh [íJR02]. Harvest [MC07, RD07, Aco02, AFHJ04, HF08a, JSMK06, PCM01, Ray07, TR09]. harvest-based [PCM01]. harvestable [MMCD08]. harvested [HTSB04, LPM^{+}09, MM01]. harvesting [HTA09, MPJ07, PDRG04, WBC^{+}08, dBMS09]. hatch [MGS00]. hatcheries [CW06]. hatchery [Kol06, MSM^{+}06, RS06a, SSKE06]. hatchery-reared [SSKE06]. hatching [CH00, HSS^{+}09, PM06]. Hatteras [MC00]. Hatton [MSGC^{+}09]. hauls [PMB^{+}03a]. haumela [Zha06]. Havel [Sim07b]. Hawaii [SL04]. health [EJR01, GR06, HE08, Ste01b, TCM^{+}08, UC05]. Heap [WBK09a]. heart [JGN04]. heavily [AGA^{+}04]. heavy [BO08, SDG^{+}08]. Hebrides [RMM05]. height [HMAN03, RNK05, RASS09]. helgolandicus [SOMT00]. Helicolenus [SNV^{+}09, SGMN^{+}06]. Heliocidaris [PJ08]. help [RR05b]. Hematodinium [HSM09, SNA01]. Henle [OV04]. herding [Som04, WWHB04]. hermit [RNK05]. Herring [DCCS09, DMDE04, Sin09, TT08, BBS09a, BRP02, BHØ^{+}04, BD02, BD03, BBK08, BBSK09, CA00, CCA04, CMHN05, CBHM07, Cla00, Cor00b, DEMD00, DH04, EDG03, FSDC09, FHJS09, Fox01, Gef09, GFH04, GO03a, GO03b, GOS07, HMDS09, HHO08, HP01, HBST02, HMQ^{+}08, HK00, HSS08, HSS07, HSS^{+}09, JNF^{+}09, Joh02, JHL05, Kas09, KDCH^{+}09, LMC^{+}01, LHR02, LAO^{+}07, MKR^{+}09, MS07, MSP09, MRV^{+}08, NTJ04, NFM^{+}02, ODCN09, OKG^{+}09, Ona03, ÓT06, ÓGS09, OL00, OL07, PHG04, PO09, PMB00, PCM09, PHDC^{+}09, PPTS09, PHO09, Pel02, PVLP04, PB05b, PH03, PN06, PS06, RKP03, RS04, RLH01, RM01, ROB04, RDB09, RT03, RMB^{+}09, STAN02, SHT^{+}09, SKC09, SDCR07, Sim03, Sim07a, SK07, Sim09, SAN^{+}05, SMP09, TL05, TR09, TK03, Vuo02, WPJ09, WPM^{+}09]. herring [ZKP03]. heterogeneity [CMC^{+}06b, GR05, VM07]. heuristics [Ard08]. hexapterus [LPA^{+}00]. hidden [Fjä05, SMB09]. Hierarchical [JHC09, Dor01, ÓMP^{+}04, PPC^{+}03]. High [HDG^{+}09, CKS03, HR01, Kin02, LHJJT04, MCRF06, MTJ^{+}07, MS02, PF08a, PST^{+}07]. High-frequency [HDG^{+}09]. high-grading [Kin02]. high-relief [CKS03]. high-resolution [PF08a]. highlight [MPG^{+}09]. highly [HOP09]. Hippoglossoides [BHMD05, MMM00, Mor04, PCS^{+}07b, PCS^{+}07a, ANNG01, HAN02, JRM^{+}03]. hippurus [DNLSM08]. hispida [KKF^{+}06, LLHK07]. Histioteuthis [HL09]. Histological [VSC06]. Historical [SFØ07, DBDA^{+}02, LLD^{+}05, PSSD08, Sin09, TT08]. histories [BPD^{+}03, BK07]. history [AMJ^{+}06, BABB08, Bar05, Bro02b, DC05a, DBDA^{+}02, HLCG04, His01, Kaa00, MHD02, ON09, PM06, PKRT06, Roc00a, WBV09]. hoc [Vin01]. hoki [O’D03, O’D04]. Hokkaido [MSM^{+}06]. holobenthic [LZS09]. Homarus [AKJ07, Agn08, IWP00, LFD^{+}09, TST^{+}09]. Home [EGO^{+}07, JDA^{+}06]. home-range [JDA^{+}06]. homogeneity [HHAB09, WBK^{+}09b, WBK09a]. Hong [CWYM^{+}02, KWL^{+}02, WLK02]. hooded [CGN^{+}04]. hook [APGD08, TM09]. hooking [APGD08, Aló08]. hooks [APGD08, WM01]. Hoplostethus [CB07, KH03b]. horizontal [FJK^{+}07, LHKGS00, PP08]. Horse [GMM^{+}08, Cos09, DRDC06, DDM^{+}05, GCM09, LMVdZ^{+}07, Mur00b, NHKJK09, RD07, RFT02, SdFBG01, Tur04, WK01]. Host [PAA06]. hosts [PN06]. hot [Ric08]. hull [DNP03, HF08b]. hull-mounted [DNP03]. Human [CW09a, DiUVH08, EMA^{+}07, BO08, PBH02, dBP02]. human-made [PBH02]. Humboldt [EH00, GG08, GGV^{+}04, MJA^{+}05, VLJM^{+}07, WGLJM04]. humpback [RSNB^{+}08]. hurricane [TB02]. husbandry [TSH^{+}06]. hybrids [KBW09]. hydraulic [GGM^{+}05, HHSM03, MDM03]. hydroacoustic [BFMJ03, CS02, FS02, FB03, HMQ^{+}08, NCM^{+}03, SSJL02, WMS^{+}03]. hydrodynamic [HKBK09]. Hydrographic [SG00, LLC^{+}08, LBNS00, MVMH04, dPM08]. Hydrographical [dPVJM04]. hydrography [BBÁMC06, Gaa00, GAP^{+}00, PTTS00, WGLJM04]. Hydrozoa [WBV09]. hyperborea [BREB09, SB01]. hypotheses [ZPRJ02]. hypothesis [Bla01, HSA^{+}01, IB00]. hypoxia [PSHL09].

Iberia [ORVP09]. Iberian [ADC^{+}08, BIdL^{+}08, BBÁMC06, CFRM08, CAGV05, Cas07, LMVdZ^{+}07, Mur00b, NGNB^{+}04, PS03, SSU^{+}09]. Ibiza [Mor02]. ice [Pie08]. Iceland [AJNM07, BGAM00, CMC^{+}06b, GRE06, GAP^{+}00, GA00, JSR06, JTE^{+}07, VC02, Vil02]. Icelandic [Bjö02, BMLH07, Buc00, GV02, GOS07, GHC09, JR06, JDN01, JCM06, JMC07, JV05, MGTS00, MGS00, ÓT06, ÓGS09, PSSD08, RS03, SPK05]. ICES [DKMO09, HHAB09, WBK09a, BNBR05, Daa03, DPW07, DDM^{+}05, GCM09, HNK07, HLSW01, KPK^{+}05, KPK^{+}06, MPD^{+}08, Mol00, Ric00b, Sin09, SBL07, Pay04, PB05a, PK07, RDF^{+}03]. ichthyolarvae [KA01]. Ichthyoplankton [GQCÁMI03, GPZ^{+}05, LND05, SP02]. Icy [PSO^{+}04]. ideas [Sin09]. identical [OAJ06]. IDentification [HW08, ISS^{+}07, WWWB03, CTM09, CBHM07, Cos09, FCM09, Fer09, GWG06, GPWG04, HCEM06, KBW09, KHEJ09, LBF01, MBC^{+}09, MYAT09, MCP03, Mur00b, PÁMGV05, Pet03, PMB^{+}03a, TGS09, Tur04]. identifies [GSS08]. identify [KFM02, MLM05, SMH09, VBF09, WB02]. Identifying [HKBK09, PBM^{+}04, GAW^{+}08, GR06, KTM^{+}05, dPM08]. Idiosepius [SAM09]. if [KDP09]. ignorance [RR09]. ignoring [DM07b]. II [TAC00]. IJsselmeer [Dek04]. illecebrosus [DCD00, Hen04]. illegal [CMK09]. Illex [AJR00, DCD00, Hen04]. illicia [WWGG02]. illustrated [HDG02]. image [Har07]. imagery [SBP07]. images [PMB^{+}03a, TST^{+}09]. imaging [GPWG04, HCEM06, KHE^{+}09]. immature [BSO01, HDG02]. Immediate [GGM^{+}05, PDRG04]. immersion [LTA00]. immune [CMO^{+}06, EJR01]. Impact [Bia00, De 04, DDR^{+}03, KTH^{+}00, MAC^{+}07, SPD00, WYMF08, BRP02, BHN06, BDTW06, Cam08, CSW06, CJM^{+}02, CA02, Des00, HBC01, HSS08, JI05, KMH^{+}05, LFW03, OK05, PCW00, Pie00, PvHG09, PFF01, PMD^{+}00, PB05c, RLF01, RK00, SDRK00, SBT^{+}09, TSK07, TCP05]. Impacts [CFR^{+}01, Ele00, FSFO08, MRT01, RBBB00, BRHG^{+}06, Cad00, CEV00, CMM01, GGM^{+}05, HSM00, HHSM03, Hea05a, Hea05b, Hol00b, HNLR04, JR01, KCCM03, LT06, LAB^{+}05, LPSL09, MM07, MHF^{+}09, MSH07, OL07, PGJ^{+}05, Ric00a, SBG06, SW06a, SBP07, Tas00, WCMK05]. impairment [Dav07, Rye04]. implementation [CdlMA^{+}00, GC05a, SFH^{+}07]. implemented [CDBS08, PRB^{+}07]. implementing [LKL08, TL05]. Implications [KPO05, NKOK00, PPHB00, BD04, BGW05, BP07, CW09b, CH05, Eig09, FGP07, Fur02, GPRD08, GPP09, HHB^{+}00, HHAB09, HBG^{+}04, Kos00, LHHF03, Lun01, MSW07, MMKR^{+}00, MNMG^{+}05, Mur00b, NCM^{+}03, NTSM07, ÓT06, PPKM07, PKP07, RS03, Ree03, RD01, SGMV^{+}08, SK07, SSI07, SCLK01, SRM08, SBDW00, SPWHR04, WJTH00, WBK^{+}09b, WBK09a, WvdMF06, YCCH07]. implicit [KPK^{+}05]. Importance [HHMM01, Vec00, Fra06, HH01, MNCU09, ODRN05]. Important [AFM^{+}09, CRW^{+}01, MB06, OLS00]. imprints [PSSD08]. improve [DRDC06, GWG06, JGST09, PCM09, Pie02]. Improved [CD06, ES03, BSS07, BBS^{+}09b, CDSC05, ZWW^{+}03]. improvement [Hol00b]. improvements [BMP^{+}08, KGRW07]. Improving [DAH^{+}08, ETB07, SBP07, vDM07, LKL08]. in-stream [BMV05]. in-water [HF08b]. incidence [LN08, WBC^{+}06]. incidences [GO03b]. incidental [WM01]. inclination [FB02a]. include [HHMN01]. included [DCPvK07]. Including [HIL00, MHH06, Fra06, KB07, Sea02]. Incorporating [Boo00, MM01, Cor01, GSSO00, HMQ^{+}08, MLOT09, SGM09]. increase [STW^{+}08]. increased [CSR^{+}02, SP07a, Ste02]. increases [TAHK06]. increasing [SL04]. Increment [CEH03, HC09]. incremental [SB04]. incubation [BIdL^{+}08]. incursion [Cam08]. independent [AHS08, LCC09, MFA07, RD07, RUCG07]. independently [BNBR05]. Index [Ano00a, Ano01c, TCM^{+}08, Ber00, BFZ05, GMM^{+}08, LKK^{+}09, RDB09, STM^{+}08, TCP05, YM00]. Indexing [UC05]. India [CMN^{+}07]. Indian [LD03a]. indicated [CFRM08]. indicates [DC05b, VH08]. Indications [HSS08, KASA07]. indicator [BBGA05, GBC^{+}05, LAB^{+}05, MMV^{+}08, MVM^{+}08, RDD06, RTB^{+}05, SIT^{+}05, TVH08, TRM07]. indicator-based [MMV^{+}08, TRM07]. Indicators [Bac08, Daa05, Deg05, GR06, OSK^{+}05, BPM^{+}05, BBR08, Bea05, BO08, BDJ^{+}05, BF04, CC05, DDGR07, DH08, FB07, FDD^{+}05, FSP05, GHFA09, HMMB^{+}08, HOHS05, HE08, JD05, Joh08, Lin05, MPG^{+}09, MJA^{+}05, MM05, NB08, NJ04, PMB^{+}08, PJ05, PQRG07, PJR08, PH05, RMDB05, RHH^{+}08, RCBM05, RR05a, RR07, RR05b, SRJ^{+}05, SP05, SLMCRM05, VBF09, WWR^{+}08, YNX^{+}05]. Indices [WPR^{+}07, BR08a, Cor07, CDBS08, GAA^{+}04, Gud04, HBD05, Kas09, KCBC00, MUK^{+}02, MKB01, Pie02, RD01, RRY08, SHT^{+}09, Sim03]. indigenous [KASA07, LMM^{+}08, RL08, TDE08]. Indirect [SP05, DGPR05]. individual [FM04, GOK05, HHMN01, HJBG04, MSF^{+}06, MH01, MHD02, PMM^{+}09, RMB^{+}09, SRMB07, TK01, TPRR04, VSC06, YSF09]. individual-based [HHMN01, HJBG04, MH01, MHD02, RMB^{+}09, YSF09]. individually [PK09]. induced [Bet04, CSdQB06, LS04, SAAFCA07]. industry [HBC01, JR01, SKC^{+}00]. inert [GTOJA06]. inertiograms [BR00]. infauna [DGMM02]. infaunal [DTC01, MCRF06]. infecting [PAA06]. Infection [PH03, GSN^{+}03, HP01, SNA01]. infectious [Cip09]. infer [GCC^{+}09]. Inference [CWC00, WSWS03, BGW03]. inferred [AJNM07, BABB08, Ber00, KHS^{+}08, LBN09, Miy03, NCM^{+}03, NSP06, PF06, SKS^{+}00, SYR^{+}08, TYH04]. Inferring [SRS^{+}07, SRJ03, BGW05]. infestation [BF02, BFK^{+}07, DRRS01, FSDB09, JHKZ09]. infesting [PGG05]. inflicted [Fjä05]. Influence [Aló08, BES^{+}01, CZC07, DGMM02, EH00, Gas02, GLDD00, GF01, GG04, HKD^{+}04, MS01a, PP08, ZWW^{+}03, Bar05, BO08, BDJ^{+}05, BF04, BD03, CF05, CH09, CMJ09, CWC^{+}03, CBS^{+}06, CDD^{+}07, DHKV01, DPN^{+}09, FB02a, HHT08, HPBK04, Hor03, KMI^{+}05, Kas09, LMC^{+}01, LHHJ^{+}09, LLD^{+}05, LMVdZ^{+}07, NK00, NFM^{+}02, PSHL09, SRMB07, UP00, WMÖ06, Ynd01, Ynd06, lPV01]. influenced [VBF09]. Influences [DMvD07, KIDY09, Bot01, MSH07, PLJ01, WPB^{+}03, YM00]. influencing [AO08]. Information [BBBF02, ARMM09, BVDS08, CH06, DAH^{+}08, Dor01, GBBG06, GLR06, LBF01, MKR^{+}09, Pie02, PPC^{+}03, RPB^{+}08]. informing [TH08a]. ingestion [GTOJA06]. inhibited [KKC04]. initial [Ard08, Hen04]. initiative [CWYM^{+}02]. initiatives [Jel07]. Injury [BC07, APGD08]. Inlet [PFF01]. inlets [BI08]. innate [EJR01]. innovations [PRB^{+}07]. inorganic [PMB^{+}03b]. Input [Gat00, BVD01]. inshore [BDTW06, EGO^{+}07, KMHS04, Sve03]. insights [vdKRS^{+}07, CFMdP07]. inspecting [FGD02]. instability [KM05]. installations [TPT^{+}09]. installed [Peñ08]. instantaneous [KCBC00]. Institutional [dVA07, Sym07]. instruments [Arn00]. Insurance [MLLK09]. intact [CH00]. integer [CDB05]. Integrated [DM07a, BKR09, EHG06, MSS^{+}05, SA05, TVH08]. Integrating [Hol03, MMCD08, PPTS09, Ric00b, Aks05, GM07, GWvM07]. integration [Aks06, Bet04, Kor00, NIF^{+}09]. integrity [SMP09]. intensity [DH09, GMM^{+}08]. intensive [MHH06]. Inter [Sec02, MJA^{+}05]. inter-calibrated [MJA^{+}05]. Inter-laboratory [Sec02]. interaction [ATH^{+}07, Cur00, FWW06, JJ06, KSD01, NW02, PAA06, SBC^{+}00]. Interactions [BDÑ04, BBBF02, DF00, GWSV08, Gro06, HW06, JSMK06, Lóp06, AFP08, BGW05, DBBM01, FB07, Fur02, GG04, HP07, HCE^{+}03, HJBG04, NRS09, PPL^{+}07, RvMBV00, RDHP00, SLN02b]. interactive [TRM07]. Interannual [BMV05, CT07, PSO^{+}04, SBR07, WS06, dLMS06, CCC02, HSS07, LC09b, MM01, MML09, Ped05, SA03, UP00, YM00]. Intercalibration [Cot01, MN02, aFADN08]. Interdecadal [PGJ^{+}05, PB00]. interest [AKLL07]. interface [BD07, Hoy07]. interference [GC05b]. intermediate [OUNB02, Sar09]. internal [WSWS03]. International [ATM02, AE02, Cot01, Jen02b, PB08a, PD07, VPC^{+}09]. interpolated [SKH02]. interpolation [HMD^{+}08, PR03]. interpretation [PR03]. Interpreting [MSF^{+}06]. interrelations [UP02]. interspecific [ATH^{+}07]. intertidal [DGK^{+}09, KPD^{+}07, dBP02]. intestinalis [TH08a, TH08b]. intra [BCT05]. intra-species [BCT05]. introduced [MPN^{+}08]. Introduction [CV02, DiUVH08, DLM^{+}05, Ele00, Fog01, HPR09, KTT06, MGM03, MM02, NRS09, PB08a, WB04, DPW07, DKMO09, HW06, MPD^{+}08, NM09, RC07, Jen02b]. intrusion [MLNC01]. invaded [RL08]. invaders [Law08]. invalidates [VSC06]. invasion [HHB^{+}00]. invasions [BP08, CBDS08, NM08]. invasive [DB08, SDÖ09]. invertebrate [Bot01, HHJK06, IWP00, PCDM08, WCP08, YM08]. invertebrates [AFGR09, BS03, Cal02]. investigate [BPD^{+}03, Bro03]. Investigating [BR02, CAWD09, CDB09, DRDC06, ERBP09, wScY02]. Investigation [MM07, CDQL06, PGD09, SW06a, TT08, TJAS04, ZPI^{+}09]. investment [YW05]. ion [NNT01]. Ionian [Kou00]. Iranian [AP07, Esm06]. Ireland [VLBB08, WBC^{+}06]. Irish [Arm01, AGA^{+}04, BFM00, BZRO06, Bri02, BD02, BBK08, BBSK09, CH09, CMO^{+}06, KCR06, LHKGS00, LND05, ÓMP^{+}04, RDB09, SCHR07, SRS^{+}07]. Irminger [Ped05, SKR^{+}06]. Island [AFGR09, BMM03, CRW^{+}01, HBD05, MNMG^{+}05, MSM^{+}06, AO08]. islandica [JTE^{+}07, KCR07]. Islands [ABB^{+}08, GM06, LBNS00, MM03b, MM03a, MMB09, SBC^{+}00, SNB^{+}02, TSK00, WS06, AFGR09, GAYR06, BDD06, CFRM08, CHB09, HJ03, HEGH02, íJCMR07, OB05, RCLD08, RUA07]. Isle [BSMB03, VBSB07]. isolation [MLG^{+}09]. isotopes [BCL03, BAO04, HOF04]. Israel [Ost09]. issue [Gre08]. issues [Fle05, PBM^{+}04]. Istiophorus [HL07]. Italy [De 04, MCRF06, MFA07, Lóp06, Mis02]. ITQs [Kin02]. ITS-5.8S [KHS^{+}08]. itself [BVDS08]. IUCN [RL07]. Ivory [GAA^{+}03]. IXa [GCM09].

J [WBK09a]. jack [BBC^{+}04, HA03, NH09, Peñ08]. jackass [BES^{+}01]. Jakarta [KIDY09]. Jan [VC02, Vil02]. Janeiro [GAZ02, ZNGF02]. Japan [HKI01, MSM^{+}06, MNY^{+}09]. Japanese [KMI^{+}05, KCL^{+}09, KNS^{+}04, KKS^{+}07, KWZ00, Miy03, NH09, SAM09, STA^{+}09]. japonica [KKS^{+}07, MYAT09]. japonicus [KCL^{+}09, Mar07, Miy03, NH09, PS09, PMM^{+}09, STA^{+}09, TYH04, ZWD08]. Jasus [GG09, LHHF03, MM01]. Jeffreys [DSG05]. jelly [KHS^{+}08]. jellyfish [BAB^{+}04, HMHI09, LPH^{+}08]. John [Dun01, YYY^{+}02]. joint [BBPW07, DB04, SKC^{+}00]. Joseph [Ano06j]. Journal [Daa03, HHAB09, Pay04, PB05a, PK07, RDF^{+}03, WBK09a, RDF^{+}03]. joy [Daa03]. Juan [TSK00]. July [Hel00]. Just [JWBP07, Le 09]. Juvenile [ADO02, ANNG01, ACD^{+}03, BOC^{+}08, BÓ06, BBSK09, CW06, CDR05, EHL07, FL06, FSS00, GTOJA06, GA05, HOF04, HSS07, Joh02, JGST09, KMHS04, KNS^{+}06, KMJH01, LTA00, LBNS00, LN08, MBPW06, MGS00, MFIO04, NTSM07, NEJH05, OR01, OFN02b, OFN02c, PVH^{+}05, RW01, RR06, SBC^{+}00, WSFH02, ZAJ01, dCA03]. juveniles [ELR01, HP07, Ham06, VSÅF05, WSC^{+}06].

Kalloni [KCCM03]. Kamchatka [Nau02, OB05]. Kaohsiung [CTF02]. Katsuwonus [AK04]. Kattegat [Bag04, CCHV05, LNLS09, SB03, Ung07, UMSA09, VSC06]. kelp [BREB09, DDGJ02, GZND02, TD00, VLJM^{+}07]. kelts [HAG^{+}08]. kept [OAJ06]. Kerguelen [BDD06]. Kernel [BPWS09]. keta [aFADN08, MMS01]. key [KMH^{+}05, VBF09]. keys [GML06]. kHz [AF06, GJH^{+}09, JFCH05, TSK03]. kilka [DM07a, Mam06]. kinase [Ber04]. kind [BFSC02]. king [SZ07, SP02]. Kingdom [PCM01, WBC^{+}06]. Kinneret [Ost09]. kisutch [RHD09]. knowledge [Bac08, Deg05, HKBK09, SBC^{+}07, TH08b, UKR05]. kob [PS09]. Kola [OGL06]. Kong [CWYM^{+}02, KWL^{+}02, WLK02]. Korean [KHS^{+}08]. kriging [RW01]. krill [AF06, CDB05, CD06, CRC^{+}09, CRTS04, DC03a, DC03b, DC04a, DC04b, DC05b, ETB07, HTA09, HHKL04, KK06b, MNY^{+}09, OKRK04, RCBM05, RCLD08, SS09, WB02]. Kristian [Chr02]. Krøyer [BF02, BFK^{+}07, GSN^{+}03, UPK^{+}08]. Kuril [OB05]. Kuroshio [KNO00, KWZ00, KMT08, NKOK00]. kuruma [TYH04]. Kuwait [YBF^{+}03]. KwaZulu [dBMS09]. KwaZulu-Natal [dBMS09].

L. [AAV^{+}04, Arm01, AGA^{+}04, AMJ^{+}06, AMGV06, BRP02, Ben01, Ber00, BSMB03, BF02, BSO01, BD04, BÓ06, BDO^{+}04, Bri02, Bro03, BD02, BD03, CFR^{+}01, CLR^{+}05, CMHN05, CCC02, CSC^{+}04, CSdQB06, Dem01, DDM^{+}05, DP03a, DP03b, FTDVC^{+}08, FLH06, Fox01, FHDM00, GKOV05, GSN^{+}03, HJ03, Han06, HU04, HHHH06, HOF04, HHH00, HAG^{+}08, HK00, JH01a, Jag02, Joh02, JDN01, JCM06, JHL05, KNKT06, KCBC00, KPD^{+}09, KMJH01, LMC^{+}01, LHJJ^{+}01, MGTS00, MS01a, MR05, MG02, MS02, ÓMP^{+}04, OCWV06, OSLO06, OGL06, PAA06, PHG04, PVLP04, PB05b, PR01, PCS^{+}04, Rad03, RHBR04, STAN02, SCCM06, SWG06, SW06b, SHAH09, SRS^{+}07, SB03, TAHK06, TM02, TK03, TAC00, UPK^{+}08, UP00, Vuo02, WBC^{+}06, WSC^{+}06]. Laboratory [Rye04, DLC03, Sec02, Zit01]. Labrador [ADO02, BBM^{+}02, GW04, MG02, RML06]. labrax [PPL^{+}07, PKP07]. Lacépède [GAFA06]. Lack [UMSA09, GOK05, Som04]. lactating [MKR06]. laevis [FMF02]. Lagoon [GAA^{+}03, Kou00, PDRG04, LPH^{+}08, MCRF06, PCD05, RD01, STM^{+}08, SLMCRM05]. lagoonal [Mis02]. lagoons [BdMAL00, BMC^{+}07, SGMMGB09]. Lagrangian [HHKL04]. Lake [Dek04, JRN06, Ost09, SP05]. lakes [RPE^{+}03, Sim07b, WMS^{+}03, RPE^{+}03, RPSSW09]. lalandii [MNGB07]. Laminaria [BREB09]. Lamna [Joy02]. land [CRB08]. landing [Rob08]. landings [GAA^{+}04, GHFA09, KGT01, LLS00, dLMACC00, PR07]. landscape [CRB08]. landscapes [SRM08]. lanternfish [YTS^{+}06]. lanternfishes [YSO^{+}03]. lanthanide [GTOJA06]. large [BPD^{+}03, BI08, CSR^{+}02, DDR^{+}03, HR09, JRCS08, KWL^{+}02, KHE^{+}09, LTA00, LdSSG02, LFW03, MMF09]. large-mesh [LFW03]. large-scale [DDR^{+}03, HR09, KWL^{+}02, LdSSG02]. largest [HSdLP06]. larvae [Ber00, Ber04, BMLH07, CWC^{+}03, DLT^{+}00, FHJS09, HP07, Ham06, HPBK04, Jag02, JV05, KWZ00, LCC08, MLNC01, MGvH06, Mil08, MYAT09, Miy03, MWS04, ODCN09, PMB00, PPTS09, PF08b, PH03, PN06, RRC03, Sab04, SNM05, SAAFCA07, TCS^{+}09, VSÅF05, VSS07]. Larval [BK07, LC06, SP02, BPD^{+}03, Bri02, BES^{+}01, CSdQB06, DAd02, DBDA^{+}02, Fox01, GWF01, HHMN01, HSA^{+}01, HSPM05, HP01, Jag02, KKS^{+}07, LC09a, MMM00, MH01, OMBP06, OFN02b, OFN02c, PPK^{+}06, SCLG00, SCLK01]. last [Pie08]. late [AMD^{+}05]. Lateral [PHO09]. Lateral-aspect [PHO09]. latitudinal [HSS07, OUNB02]. L’Aulne [AVJ^{+}06]. Laurentian [RPSSW09]. Lawrence [Dup05, DR08b, CCC02, HSA05, LMC^{+}01, Sin01, SPS00b, SFM01]. layer [BBA03, BB09, CT07, SR03]. layers [CD09, HU04]. learned [KCR06]. learning [ÖG04]. Lectins [EJR01]. legal [CMK09]. legislation [CRIP08]. lemon [HBS^{+}06]. Length [MMC03, AK04, ASB05, BPM^{+}09, BS03, BA03, BJ00c, CLL^{+}09, FB03, GML06, Har07, JAC00, KB07, LEP04, LHJJ^{+}01, STA^{+}09, SSC^{+}06, SB01, UE01, VM07, WPM^{+}09, WWHB04]. length- [BS03]. length-at-age [WPM^{+}09]. leonina [ISS^{+}07]. Lepechin [CMHN05]. Lepeophtheirus [BF02, BFK^{+}07, GSN^{+}03]. Lepeoptheirus [UPK^{+}08]. Lepidorhombus [LP00]. Lérez [SCCM06]. lesser [HKD^{+}04]. lessoniana [TA05]. lessons [BHN06, KCR06, NM08, PD07, SH06]. Letang [PFF01]. Letter [Fra00]. leucas [AFHJ04]. leucopsarus [YTS^{+}06]. Leucoraja [FMF02]. level [RDHP00]. levels [Cha04, GR06, HHMN01, MML09, RT03, SLMCRM05]. Lévy [BBGA05]. LFA [AK04]. lice [BF02, Box06, DRRS01, SW06b]. licence [PS09]. licensing [SW02]. Lidar [TCSW06, Bro02a, CCB^{+}06, CDM03, CTW09, CD09]. Life [Bro02b, CRTS04, GAP^{+}00, HLCG04, Kaa00, PM06, vdKRS^{+}07, ÅD07, AG00, Bar05, DCN^{+}04, DBDA^{+}02, FAL^{+}08, HBG^{+}04, Hea07, His01, LPL03, MHD02, ON09, PKRT06, Rob05, Roc00a, VH08]. life-cycle [ÅD07, HBG^{+}04]. Life-history [HLCG04, ON09, Roc00a]. light [DH09, FHJS09, TAHK06]. Ligurian [FB02b, RRT00]. like [OdSBS09, SC00]. likelihood [MS01b, RUN07]. likelihood-based [RUN07]. Limanda [DWC03, NWH02]. Limfjord [ATH^{+}07]. limit [Fru02, OH00]. limitation [SAMS02]. limitations [GNC08]. Limited [DTC01, Ard08]. limiting [FPS06]. Limits [HNK07, FMF02, SAM09]. line [BGW03]. line-transect [BGW03]. linear [GDL04, HC09, MRV^{+}08, PH03, TSK03]. linefishery [LPM^{+}09]. linefishing [PS09]. lines [Pet03]. link [AFP08, CAWD09, OL00]. Linkage [FHDM00, ELR01]. linkages [DM07b, GC07, LD03a, Sin09]. linked [Rob05]. Linking [DCCS09]. links [AFP^{+}09, GWF01, dPM08]. Linnaeus [Dun01, GAA^{+}04, His01, PS03, UE01, YYY^{+}02]. Linnhe [SW02]. lintea [SB01]. lion [GWSV08, HA03, BGG^{+}06]. lipid [MNCU09, OMBP06]. lipid-rich [MNCU09]. List [Ano00b, Ano00c, Ano00d, Ano01d, Ano01e, Ano01f, Ano01g, Ano02a, Ano02b, Ano02c, Ano03a, Ano04, Ano07a, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d]. Lithodidae [Col02]. little [BKN^{+}07, FMF02]. live [BMP^{+}08, KNS^{+}06]. lived [GAFA06, MSH07, PM06]. liver [KNKT06, YM00]. Liverpool [WWR^{+}08]. lividus [SDG^{+}08]. Living [CdlMA^{+}00]. Loano [FB02b, RRTP02]. Loano-Ligurian [FB02b]. lobster [Agn08, CEH03, GG09, HU04, LHHF03, LFD^{+}09, MM01, MAAN09, MFA07, SJ08, STMM06, TAC00, UE01, dLMS06]. lobsters [AKJ07, TST^{+}09]. Local [CF05, HB09, AKJ07, Bai09, BRP02, BDD06, CWYM^{+}02, KM05, RvMBV00, ROB04]. localities [OED^{+}04].

Localization [PMM^{+}09]. localized [DBC03]. locate [SEOR09]. location [BHMS02, ES03, HHMM01, HMPC04]. locations [Hea00a, KTRG06]. Loch [SW02]. loci [KM05]. Lofoten [Hel00, MJB08]. logbook [BVDS08]. logging [MLMC02]. Loliginid [DLR02, JP03, RPR02]. Loligo [MB01, OSB06, OR09, RUA07, Rob05]. Long [BFM00, CTF02, CFN03, ERP01, FHHH00, Knu09, PF06, PPH09, RL05, RF01, BHMD05, BGAM00, CH09, DC05a, Des00, HSM00, MSH07, NEJH05, Pen07, WYMF08, Ynd03]. Long- [BFM00]. long-lived [MSH07]. long-tailed [DC05a]. Long-term [CTF02, CFN03, ERP01, FHHH00, Knu09, PF06, PPH09, RL05, RF01, BHMD05, BGAM00, CH09, Des00, HSM00, NEJH05, Pen07, WYMF08, Ynd03]. longer [SP07a]. Longevity [Sec00a, BHØ^{+}04]. longirostris [GAFA06]. longline [BJ00b, BJ00a, HLS00, MCB09, WM01]. looking [BFMJ03]. loops [MMKKJ08]. Lophius [FAL^{+}08, His01, JRCS08, LDQ08, LDML08, LDM08, MSB04, MD01, RNS08, VLBB08, WWGG02]. Lorn [BML^{+}05]. Loss [AMGV06, UBP^{+}09]. losses [Fjä05]. Lost [LGH^{+}09]. louse [BFK^{+}07, GSN^{+}03]. Louvenga [OGL06]. low [BHMD05, Bra05, CH09, HU04, MCRF06, OLB01]. low-frequency [CH09]. Lowe [FBMR^{+}03, MN02]. Lower [BMV05]. lucioperca [AP07]. Lumpers [KDCH^{+}09]. lunar [Ynd06]. Lutjanus [GHBR08, SBG06, WSFH02, ZCH06].

m [BGG^{+}06, RKE06, SYR^{+}08]. M. [GLR06]. M74 [Vuo02]. mackerel [Bar05, BR02, BBC^{+}04, BR00, CTW09, Cos09, DRDC06, DDM^{+}05, GMM^{+}08, GHI^{+}04, GCM09, GOK05, GKO07, GKFM09, HAvH06, HA03, LMVdZ^{+}07, Mar07, MLOT09, Mur00b, NH09, NHKJK09, NHK09, PAA06, Peñ08, PS06, RD07, RFT02, SdFBG01, SSI07, TCSW06, Tur04, TCTC09, WK01]. macleayi [MBPW06]. macloviana [SdlRdA06a, SdlRdA06b]. macroalgal [FB02b, PJ08]. Macrobenthic [DWDD03, Kou00, Cor00a, EHG06, GRE06, GCC^{+}09, PAC02]. macrobenthos [BLRC05, MS00]. macrocephalus [EH04, Som04]. macrocrustaceans [HCV03]. Macrocystis [DDGJ02]. macrofauna [SFKC02]. macrofaunal [GLS^{+}03, GGM^{+}05, SW06a]. macrophyte [BdMAL00]. macrophytobenthic [FB02a]. macroplankton [ODRN05]. macropterus [BES^{+}01]. Macrorhamphosus [KTM^{+}05]. Macroscopic [Cos09]. macroscopically [VSC06]. macrostructures [CFMdP07]. macrozoobenthic [SAPP04]. macrozoobenthos [BdMAL00, vDEM^{+}00]. macrozooplankton [SYR^{+}08]. Macruronus [DC05a, O’D03, O’D04]. mactroides [FM04]. made [PBH02, WS02b, vdVBMR00]. Madsen [Chr02]. maenas [YG08]. maerl [HSM00, HHSM03, KMHS04]. Magaguadavic [OCWV06, WC01]. magellanicus [DC05a, HS06, HC09, SBT^{+}09]. magister [BC07]. Maine [OL07, BR08b, HB07, PGJ^{+}05, SW06a, SGY08, TM09, WBD^{+}06]. mainland [VGF03]. Maja [CF06]. major [CvdLHF08, DMDE04, OBNU02, KH03a]. Majorca [MNMG^{+}05]. make [Roc00b, SS07]. Making [PRB^{+}07, DMvD07, DAH^{+}08, Dor01, HOHS05, Le 09]. male [YW05]. males [HSdLP06]. maliger [SH07]. Mallorca [dPVJM04]. Mallotus [ADO02, CFL00, CMDN02, CF02, CRW^{+}01, CMHN05, DAd02, Dol02, DBDA^{+}02, FRK02, GBT02, GDH02, GW09, Mow02, NT02, NW02, Nau02, OUNB02, Ros05a, Tje02, Vel02, Vil02]. mammal [BBPW07]. mammals [DGO^{+}09, OL07]. Man [BSMB03, VBSB07]. manage [HB09]. managed [dBMS09]. Management [AWW^{+}07, DB08, Fur02, HMK^{+}07, HW06, HF08b, RC07, ROB04, SK07, Ska07, APD09, AKLL07, Arn00, BPM^{+}05, BBS09a, BB07, BdP07, BD04, BP07, But07, CBBL09, Cam08, Cla00, CTLN09, CH05, CC05, DDGR07, DB04, DRDC06, DMvD07, DOBT02, Eig09, Fle05, FPS06, FQS01, GWF01, GSSO00, GFKM07, GR06, GAM^{+}06, HS07, HM08, HNK07, HTA09, HPR09, Hol03, Hoy07, HK06, Jel07, Jen09, JYW09, JvD07, KS08, Kat05, KPK^{+}05, KPS^{+}05, KPO05, KPK^{+}06, KMG^{+}07, KDCH^{+}09, KM02, KMM07, KF08, Kos09, KBDC^{+}08, LO05, MV07, MKR^{+}09, MSW07, MPG^{+}09, MNMG^{+}05, MS00, MRT01, NJ04, OSK^{+}05, OGR^{+}07, PPKM07, PPL^{+}07, PKP07, PST^{+}07, PR04, PJR08, PKH^{+}08, PRB^{+}07, PSFY07, PCS^{+}04, PD07, RPB07, RS04, RBD^{+}07, Ree03, RP07, RR05a, RR07, RL07, Ric00b, RR05b, RR09]. management [RPK^{+}03, RG07, RD01, SPS00a, SGMV^{+}08, SMH09, SH06, SMI07, SBC^{+}07, She07, SCHR07, SP07a, Sim07a, SFH^{+}07, SRM08, Ste01b, SE09, SS07, Sym07, TR09, TCP05, TB02, UASN07, VRP04, Ynd01]. managers [HB09, OH07, Pet04, vDM07]. Managing [CdlMA^{+}00, CSC^{+}04, KHN03, LC09a, RCL05, WLS07a, Cha04]. Manila [MCRF06]. many [Agn08]. Mapping [BI08, BML^{+}05, FGFP08, CTLN09, KWL^{+}02, KCD^{+}03, MSGC^{+}09, NRS09, SA03]. maps [BML^{+}05, TST^{+}09]. Mar [STM^{+}08]. March [DBDA^{+}02]. margaritacea [dBMS09]. margin [BSA09]. mariculture [CBS^{+}06, GMGN06, GLKPCP01, HKI01, KTT06, PFF01, RMKT01, SB06, WHP01, HLSW01]. marina [BRE^{+}08]. Marine [Ano01h, BLMB06, Cad00, CW09a, CHB09, DF00, Daa03, ERBP09, Ele00, Hal01, HHAB09, HKBK09, HPR09, KPD^{+}09, LPSL09, NM09, Pay04, PB05a, PK07, RDF^{+}03, SEOR09, VBF09, WBK09a, YSF09, AGH^{+}09, Aco02, Aga00, AFM^{+}09, AGC02, ASC01, BBR08, BJN^{+}06, BLRC05, BBPW07, BRHG^{+}06, BP08, Cal08, CG07, CW09b, CBM09, CTLN09, CdlMA^{+}00, Dem01, Des00, DGO^{+}09, DB08, EB04, FSFO08, FSQ^{+}03, FN00, FN02, GCS09, Gre08, HJB^{+}08, HHMN01, HB09, Hol00b, Jen09, KS08, KHN03, KCD^{+}03, KF08, KTS02, MBC^{+}09, MHD02, MJA^{+}05, MLLK09, NRS09, NM08, OK05, OSLO06, Orl03, OL07, PMB^{+}08, Pas06, PB08a, PBM^{+}04, PRD^{+}06, Pow00, RAKS06, RR02a, RSC^{+}09, Ros05a, SL01, SRS^{+}07, SCLK01, SRM08, SBDW00, SGAC00, Tas00, TVH08, TPRR04, Vec00, VLJM^{+}07, WDRP09, WLK02, vDEM^{+}00]. marine [vdMBD00, CBBL09, MMB09, TSK00]. marinus [GAM^{+}06, HKD^{+}04, JWM03, PGD09, SRGC04, SGS^{+}05, Str05, TM02]. marisalbi [PMB00]. Maritimes [NSP06]. mark [AJNM07, BVB^{+}07, NSP06, SSKE06]. mark-recapture [AJNM07, BVB^{+}07, NSP06, SSKE06]. marker [QGdS04, SMH09]. markers [CDQL06, GTOJA06, GKOV05, PSSD08]. market [Bro03, CMK09]. markings [LTA00]. Markov [LN03]. marks [O’D03]. Marmara [CAAJ07]. marsh [HCV03]. marshes [TDE08]. Marxan [SEOR09]. Maryland [RRC03]. Mass [CVL^{+}09, BR04, BF04, MBM02, War01a]. Material [CWC^{+}03, CWC00, WWR^{+}08]. maternal [OLB01]. Mathematical [TMI^{+}04, GJL08]. mathematics [RR09]. mating [RHBR04]. matrix [FMF02]. matter [HH01, SML01, URMS04]. matters [SJKN^{+}04]. Maturation [BHMD05, AJR00, BHØ^{+}04, BWG^{+}07, EDG03, HDG02, JRM^{+}03, OSB06, PM06, SAM09, TAHK06, VGBH09, WPM^{+}09]. mature [Jør03, MJB08, Mor04, SPK05, SP02, dLMS06]. Maturity [LNLS09, AGA^{+}04, BA03, Bro03, CF06, Cos09, FPKH03, HSdLP06, KTRG06, LFD^{+}09, MB01, OV04, PS03, SSC^{+}06, SHdLP04, TAC00, Ung07, VSC06]. Mauritanian [EIS05, JI05]. maxima [AGY^{+}05, GAYR06]. Maximum [MS01b, BGW03, Mye01]. maximus [BSMB03, ERGT07, LTA00, LPL03]. May [Chr02, MLNC01]. Mayen [VC02, Vil02]. McEachran [MG07]. ME70 [BBS^{+}09b, TMB08]. meadows [FTDVC^{+}08, SJGRRRE02]. meals [TM02]. Mean [ODCN09, KFM02, PF06, TM02, WPM^{+}09]. means [Deg05]. measure [SCWD08]. measured [AF06, CSH00, HH04, JO02, MSIL09, XZW05, YTS^{+}06, ZWD08]. Measurement [Kor00, RKM09, GPWG04, KNS^{+}04, TSK03, Zha06]. Measurements [CW05, GJH^{+}09, NHK09, BW08, CWC00, CDM03, CD03, CB07, DC03b, DC04b, EDG03, Erm09, FL06, GLDB04a, GLDB04b, GW09, HK00, JB00, KCL^{+}09, KRM05, KH03b, MR09, Orł00, Orl01, PHO09, Peñ08, STA^{+}09, TCSW06, War01b]. measures [BB07, BMC^{+}07, CRIP08, HMK^{+}07, HRM04, Hol03, PCM01, RS03, RPK^{+}03, RG07, SS07]. Measuring [FSDC09, Fra06, PO09, CDSC05, Hol00b, VCC07]. meat [RASS09]. mechanical [PDRG04]. mechanics [SH06]. mechanisms [Hus04, MLLK09, WW07]. Medes [MMB09]. mediate [MLLK09, TDE08]. mediated [TC01]. medicines [DRRS01]. Mediterranean [GM06, KCCM03, LLS00, MM03b, MM03a, MMB09, MNMG^{+}05, Mor02, SPGT00, SF09, dPVJM04, dPVV04, dPM08, AGH^{+}09, APGD08, AJR00, BSS07, BS02, CBBL09, CSR^{+}02, CFMdP07, FJSJBS^{+}08, FBD^{+}08, GAA^{+}04, GOA^{+}09, JDA^{+}06, KTH^{+}00, KV06, KTS02, LG08, LdSSG02, Lóp06, LBN09, LPSL09, MLG^{+}09, MBC^{+}09, MAMO02, RMDB05, RF01, RSC^{+}09, Sab04, SDRK00, Sec02, Sil03, SSC^{+}06, SPD00, STMM06, SGMN^{+}06, TMG^{+}08, Tur04, vDEM^{+}00]. mediterraneus [Tur04]. medium [ACD^{+}03]. medium-term [ACD^{+}03]. meeting [Joh08, RHH^{+}08]. Megafauna [SQN08]. megafaunal [BvS00]. megalops [BGW05]. Meganyctiphanes [CDB05, CRTS04, ETB07, KK06b, OKRK04]. Megaptera [RSNB^{+}08]. megazoobenthos [PPHB00]. Megrim [LP00, PÁMGV05, PPB03]. meiofauna [DGMM02, MVM^{+}08]. meiofaunal [MML^{+}00]. Melanogrammus [ATM02, BDO^{+}04, GML06, ÖG04, PKRT06, RW01, TSH^{+}06]. membras [HP01, PH03, RLH01]. Memoriam [Ano06j]. menhaden [BW08]. Menor [STM^{+}08]. mentella [DNP03, JGN04, Ped05, SRN00, SRGC04, SKR^{+}06, SRS^{+}09, SGS^{+}05, Str05]. Mercator [CDBS08]. Merceneria [BABB08]. mercury [NCC^{+}07]. meristic [Tur04]. Merlangius [TM02]. merlangus [TM02]. Merluccius [BOC^{+}08, BdP07, CAGV05, ERBP09, GAA^{+}04, MAB^{+}07, PS03, SG00, SGMN^{+}06, WPR^{+}07, dPBB^{+}03]. mesh [BM09, aFADN08, Gas02, GM06, LFW03, MSW07, ÖG04, RBGJ08, STJ^{+}07, ZFFT01]. meso [KNO00, NKOK00]. meso-scale [KNO00, NKOK00]. mesocosm [LBL06]. mesocosms [HRHC00]. Mesodesma [FM04]. mesopelagic [KI04, OGD09, TK01, YSO^{+}03]. mesoplankton [NGNB^{+}04]. mesoscale [BES^{+}01, SG00]. Mesozooplankton [IA04, LDNS08, BBÁMC06, BAO04, MVMH04, PP08, SL04]. messenger [íJR02]. meta [Mye01, SBL07]. meta-analysis [SBL07]. meta-analytic [Mye01]. Metabolic [MGvH06]. Metabolism [KI04, IA04]. metals [BO08, SDG^{+}08, TM09]. Metapenaeus [MBPW06]. metapopulation [JNF^{+}09, SKC09]. metapopulations [KDCH^{+}09]. Methane [SQN08, Ost09]. method [Aks05, Aks06, AF06, Arm01, BPT09, Bri02, CBHM07, CDSC05, DCRB09, Den08, ES02, FCM09, GML06, HH03, HLL^{+}08, KYG03, LKL08, LN08, MSH07, MFA07, NIF^{+}09, OMTS03, Pet03, RO05, SOB^{+}07, SSKE06, dPGPB06, vDBF^{+}09]. Methodological [BBS^{+}09b]. methodology [Fle05, GBT02, dPBB^{+}03]. Methods [Erm09, BHH^{+}08, BPWS09, CF06, GWG06, GAM^{+}06, GOS07, HK06, LCC09, MPJ07, MS01b, OMA09, RPR02, SGM09, SGS^{+}05]. métiers [Mar08]. Metrics [Daa05, NJ04, Ric00a, TMG^{+}08]. Meuse [WJB07]. Mexican [VH08]. Mexico [CDM03, PCD05, SBG06, SLMCRM05, WCP08]. MHC [CMO^{+}06]. MIBA [MMV^{+}08]. microcalorimetry [MGvH06]. microconstituents [Sec02]. Micromesistius [BK07, DM06, HBD05, JG07, RMM05, WGMM08]. micronekton [BBA03, KRYL09]. Micropogonias [dCA03]. Microsatellite [MLG^{+}09, TCTC09, ZPI^{+}09, CFRM08, CDQL06, KM05]. microsatellites [XZW05]. microscopic [Cos09]. microstructure [BD02, CBHM07, PMB00, WWGG02]. Microzooplankton [LC04, Cal08]. mid [EKPT07, OR09]. mid-1940s [EKPT07]. mid-shelf [OR09]. midnight [DK00]. midshelf [PTTS00]. midwater [PSHL09]. migrating [HAG^{+}08, JYW09, MHF^{+}09]. Migration [Han06, AAV^{+}04, AVJ^{+}06, AJNM07, BEB^{+}09, BVB^{+}07, DK00, GR05, GA00, GHI^{+}04, GPP09, HJ03, JWBP07, KKS^{+}07, ÓGS09, PSHL09, Ros09, SBC^{+}00, SK04, Ste01a, SPWHR04, VSS07, WLS07b, WJB07]. Migrations [PPL^{+}07, CCC02, Cor00b, DEMD00]. migratory [Bjö02, BD07, LPM^{+}09, WDRP09]. Milne [Rob08]. mimics [GC02]. mineralization [HH01]. minimizing [BHN06]. Minimum [Rob08]. minke [LHHJ^{+}09, LHR02, TF02, TL05]. miranda [HL09]. mischmetal [TM09]. missing [AFP08]. mistakes [Le 09]. mitchilli [BW08, SCLG00]. Mitigating [GLKPCP01, CDB09, LFW03]. Mitigation [LS04, BFSC02].

Mitochondrial [HAvH06, WW01, BABB08, KHS^{+}08, QGdS04, TYH04, TCTC09, ZCH06]. Mixed [KCL04, APGD08, Cha04, CBHM07, Cla00, CSC^{+}04, DAAD09, HC09, HMPC04, JMC07, KMT08, KBDC^{+}08, MRV^{+}08, O’D03, PCS^{+}04, RBD^{+}07, RDD06, UR01, VRP04]. mixed-effects [HC09]. mixed-species [APGD08, DAAD09, O’D03, VRP04]. mixed-stock [Cha04, Cla00, PCS^{+}04]. mixing [LND05, RDB09, Sab04, VH08]. mixotrophy [HP05]. Mixture [FB03]. mobile [LC09a, RAR^{+}07]. mobility [LC09a]. mobilization [MGH08]. model [ÅD07, BBR08, CFL00, DC03b, DC04b, DC05b, Dor01, FGR04, GJL08, HP07, HO01, HC09, HCE^{+}03, HSPM05, HBST02, HMQ^{+}08, HK06, IFUR08, KIDY09, KPD^{+}09, Kup04, LHHF03, LHKGS00, LSH^{+}09, LC09b, LBN09, MKR06, MH01, MHD02, MMB09, MRV^{+}08, MNY^{+}09, OH07, ORVP09, PG08, PKP07, PCS^{+}07b, PCS^{+}07a, PST^{+}07, PRD^{+}06, Ric09, RGG^{+}04, RDB09, RMB^{+}09, RRC03, SKC09, SB04, SS00, TMI^{+}04, TNF09, Tje02, TRM07]. model-based [KPD^{+}09, TRM07]. modelled [Bar05, HH04, KMJH01]. Modelling [BEB^{+}09, BZRO06, BDS01, CRC^{+}09, GNC08, GO03a, GO03b, GHFA09, HP07, HHB^{+}00, HF08a, HP01, HMPC04, Kar06, KMM07, LFD^{+}09, LN03, MLNC01, PF08a, PvHG09, PRD^{+}06, RKP03, SCJ00, TF04, ZFFT01, BRP02, BI08, BRE^{+}08, BREB09, BBBF02, DM06, FSDC09, GFP09, HLL^{+}08, HFMD06, HKBK09, HHKL04, HPBK04, HBW^{+}09, HJBG04, JHC09, JGST09, LKL08, LLS00, MRV^{+}08, MB06, ÓMP^{+}04, PF08b, PCRW04, PPC^{+}03, RMB^{+}09, SC00, WHP08, YSF09]. models [AWW^{+}07, BR02, BI08, BGG^{+}08, BIdL^{+}08, Boo00, Cor01, DC01, DLR02, FGR04, FB03, HJB^{+}08, HSS^{+}05, HHMN01, Hol00b, HMQ^{+}08, LEP04, MHH06, Miy03, MJA^{+}05, MLOT09, PPKM07, PH03, RUA07, SP05, TS05, ZO03, dBMS09, dHET04]. moderate [OBNU02]. modes [EB04]. modestus [MWS04]. modification [LS04, Rho08]. Modified [WM01]. Molecular [KBW09]. mollusc [BCD^{+}02]. Mollusca [JGM^{+}08]. molt [CEH03]. moment [HP07]. Mondego [STM^{+}08]. monetary [Sim07a]. monitor [SW06c, SHSKR01]. Monitoring [Bea05, JR01, LPA^{+}00, PPH09, Ros03, WBC^{+}06, BMJ08, But01, CTF02, DRRS01, EHG06, HMMB^{+}08, MTJ^{+}07, NCC^{+}07, NJ04, NEJH05, RCBM05, ŠCBD09, WHP01, ZWW^{+}03, Zit01]. monkfish [HM08, JRCS08, LDM08, MD01, RNS08, RGG^{+}04, SGY08, MPD^{+}08]. Monkfish/ [MPD^{+}08]. monsoon [KA01]. Monte [LN03]. monthly [DPN^{+}09]. moon [KHS^{+}08]. moored [DPN^{+}09]. moorings [GLKPCP01]. morhua [EKPT07, KPD^{+}09, Rad03, Arm01, AGA^{+}04, BGL08, BBM^{+}02, BSO01, BD04, BÓ06, BDS01, BDTW06, CCC02, CSdQB06, DJRO06, DMvD07, Dri05, ET07, Erm09, FL06, FR09, GW04, Ham06, HOF04, Joh02, JDN01, JCM06, JKSO06, KNKT06, KPO05, KTRG06, KNS^{+}06, KMJH01, MGTS00, MGvH06, MW03, MSS^{+}05, MR05, MSR03, MJB08, MB05, MG02, OMBP06, OSLO06, OFN02b, OFN02c, PPK^{+}06, PK09, RL05, RML06, RR06, RHBR04, Sin01, SMK08, SBB^{+}05, SSA08, SPWHR04, SB03, SPS00b, SFM01, TAHK06, UP00, VGBH09, VSC06, WSC^{+}06, YM00, YW05]. Morone [Gro06, RRC03]. Morphological [Mur00b, MPG^{+}09]. morphology [YSO^{+}03]. Morphometric [CS05, CF06, Sil03, HSdLP06, SRN00, SMH09, Tur04, Ung07]. morphometry [PF08a]. Mortality [BvS00, BM01a, BBB06b, UBP^{+}09, Aló08, AFP^{+}09, BHMD05, BMU09, Coo04, Dav07, DC01, FQS01, GLS^{+}03, GPRD08, HP07, HU04, HIL00, Hor08, ISHB07, JWBP07, KCR07, LHHF03, LC09a, MBPW06, MF07, MAC^{+}07, MSH07, MMF09, MPG^{+}09, MS04, MLOT09, OED^{+}04, PPTS09, PCD05, PvHG09, PPH09, PMD^{+}00, RDD06, Sin01, SLN02a, TLMO08, War01a, WM01, WJB07]. Morus [Bun01, CDD^{+}07]. morwong [BES^{+}01]. moult [CRTS04, dLMS06]. moulting [SNA01]. mounted [ADDH04, DNP03, DH09, DGO^{+}09]. mouth [PMB^{+}03b]. mouths [SAPP04]. move [CBDS08]. Movement [CW09b, AKJ07, MHF^{+}09, ZKP03]. Movements [WBD^{+}06, HRM04, LMC^{+}01, SHAH09]. MPA [SF09]. MPAs [CMJ09, FGP09, GFP09, Le 09, LC09a, MWF^{+}05]. MRI [PF08a]. MSVPA [BVD01]. MSW [YMF02]. MSY [PKH^{+}08, SZ07, WCMK05]. MSY-based [PKH^{+}08]. mtDNA [DC05a, XZW05]. much [SMB09]. mud [JYW09, SDRK00, SW06a]. mud-bottom [SW06a]. mudclouds [Som04]. mudflat [dBP02]. Müller [CMHN05, GBT02, NT02, OV04, Tje02]. mullet [FBD^{+}08, MLG^{+}09]. Mullidae [MLG^{+}09]. mulloway [PMM^{+}09]. Mullus [FBD^{+}08, MLG^{+}09]. Multi [BMC^{+}07, KCL04, CSH00, DAAD09, HJB^{+}08, Jaf06, KPK^{+}06, KO02, MM07, MKB01, MHD02, MCL03, wScY02]. multi- [MCL03]. multi-angle [Jaf06]. multi-annual [KPK^{+}06]. multi-area [HJB^{+}08]. multi-beam [MM07, wScY02]. multi-dimensional [MHD02]. multi-frequency [CSH00, KO02]. multi-gear [DAAD09]. Multi-objective [BMC^{+}07]. multi-species [MKB01]. Multi-Stock [KCL04]. multi-zone [MKB01]. multibeam [BPT09, BBS^{+}09b, BPWS09, CD07, CRC^{+}09, GBBG06, GJH^{+}09, KHE^{+}09, OMA09, TMB08, TGS09]. multibeam-echosounder [CD07]. multibeam-sonar [KHE^{+}09]. multidecadal [LTI09, SN08]. multidisciplinary [PDRG04]. multifleet [CGV03]. Multifrequency [AOSD09, BPT09, GPP09, KDO^{+}08, KHE^{+}09, LDCH^{+}09]. multifrequency-echosounder [KHE^{+}09]. multinomial [BIdL^{+}08]. Multiple [BBM^{+}02, MPN^{+}08, BKR09, Bot01, CDSC05, CA02, EB04, HHMN01, Jaf08, KPS^{+}05, KDP09, MKB01, MSI07, MM05, PPW^{+}09, RL08, WSWS03]. multiple-angle [Jaf08]. multiple-frequency [CDSC05, WSWS03]. multiple-species [PPW^{+}09]. multiplicative [EN02]. multipolluted [GAA^{+}03]. multishape [DNLSM08]. multispecies [BSS07, CGV03, DB04, GAA^{+}04, GRMR07, HJB^{+}08, Hol03, Hol00b, JMLG05, LPM^{+}09, LJM00, RD03, SLN02b, TLMO08, Vin01, lPV01]. multistep [MMV^{+}08]. multivariate [BMDBM09, EIS05, FFL06, JR06]. Munida [LCC08]. Muricidae [SDÖ09]. murphyi [BBC^{+}04, PF08a, Peñ08]. murre [RNWS08]. muscle [PAA06]. muscle-infecting [PAA06]. mussel [CFR^{+}01, RLF01]. mussels [NCC^{+}07, PFLFR08]. Mustelus [FCM09]. Myctophidae [YSO^{+}03, YTS^{+}06]. mykiss [BBMS01, RHD09, RS06a]. mysids [AOSD09, CDB05]. mystus [HLL^{+}08]. Mytilus [CFR^{+}01, PBLFR06].

NAFO [AG00]. Namibian [GLR06]. Narragansett [AO08]. narrow [HAvH06]. narrow-barred [HAvH06]. narrowband [BNF^{+}07]. nasus [Joy02]. Natal [dBMS09]. National [CBBL09, SGMMGB09]. nations [BBR08]. native [KASA07, PJ08, RL08]. Natura [BI08, PFK^{+}09]. Natural [Sin01, VH08, AFP^{+}09, BCD^{+}02, BCL03, Coo04, ERGT07, GPRD08, GC02, GZND02, HS09, LHHF03, PCRW04, SGMMGB09, SLN02a, WM04]. nature [JJ06]. nauplii [FR04, DGC00b]. near [BF02, CT07, GO03b, LBNS00, SBC^{+}00, WS06]. near-normal [GO03b]. near-surface [CT07]. nearby [CHB09]. nearshore [BBA03, Bla00, EHG06, GW04]. nebulosus [Kup04]. need [Fru02, SB06]. needs [Joh08, RHH^{+}08]. negative [HSS08]. negatives [But07]. Negotiation [AKLL07]. neighbourhood [SDG^{+}08]. neighbouring [BCD^{+}02, CMJ09]. nekton [PSHL09]. Nemadactylus [BES^{+}01]. Nematoda [MMM00]. Nemopilema [HMHI09]. Neocalanus [LDNS08]. Neomysis [CDB05]. neon [CZC07]. Neophocaena [Aka02, XZW05]. Nephrops [BFM00, BM01a, BM01b, BD04, Bri02, CAWD09, CDB09, CEH03, HU04, MS01a, MAAN09, MFA07, PR01, SP03a, STMM06, SNA01, TAC00, UE01]. neritic [GKFM09]. net [AS02b, FGR04, FWW06, GPWG04, NGNB^{+}04, SML01, TMI^{+}04]. net-cage [AS02b]. net-pen [SML01]. netpens [TSH^{+}06]. nets [FB02b, LGH^{+}09]. netting [SOB^{+}07]. network [AGH^{+}09, But01, EHG06, EDG03, HHAB09, WBK^{+}09b, WBK09a]. networks [CTM09, CW09b, GDL04]. neural [CTM09, EDG03, GDL04]. neutral [PSSD08, Pel02]. Newfoundland [AGC02, ADO02, BHMD05, BPD^{+}03, BBM^{+}02, CF02, Dem01, GW04, MSR03, Mow02, NW02, RML06, RR02b, RR06, Ros03, WPM^{+}09]. next [Pay04, VRP04]. nick [Sec00b]. night [OGD09, PMN01]. Niña} [CZC07, EH00, OdSBS09]. Niña-like [OdSBS09]. Niño [CZC07, OdSBS09, EH00, GZND02, KKS^{+}07]. Niño}- [OdSBS09]. Niño/La [CZC07]. nitrogen [BCL03, BAO04]. no [EB04, MBC^{+}09, PBH02]. no-take [EB04, MBC^{+}09, PBH02]. nodal [Ynd06]. noise [Bet04, DH07, DHWW08, GEM01, KRM05, Kor00, PCM09]. noise- [Bet04]. noise-reduced [DHWW08]. noisy [DCPvK07]. nomurai [HMHI09]. Non [LMM^{+}08, SLN02a, TSK03, CLR^{+}05, Cor07, EIS05, GDL04, KASA07, KPD^{+}07, LKL08, MS01b, PSSD08, PvHG09, PMD^{+}00, RL08, SMEK01, TDE08]. non-active [LKL08]. non-aquaculture [SMEK01]. Non-indigenous [LMM^{+}08, KASA07, RL08, TDE08]. Non-linear [TSK03, GDL04]. non-neutral [PSSD08]. non-parametric [MS01b]. Non-predation [SLN02a]. non-random [Cor07]. non-stationarity [CLR^{+}05]. non-stationary [EIS05]. non-target [KPD^{+}07, PvHG09, PMD^{+}00]. nordenskioeldii [SMEK01]. Nordic [BR04, SBR07, TH05]. Nordmann [UPK^{+}08]. normal [GO03b]. Normandy [BMV05]. normani [MG07]. norms [BHMD05, HDG02, VGBH09]. North [FJK^{+}07, HMPC04, De 04, DLR02, GAZ02, HJ03, LD05, LdSSG02, LME05, MS00, RSNB^{+}08, ZNGF02, ATM02, AE02, AMD^{+}05, Bai02, BCAN^{+}06, Bar05, BRP02, BA03, BvS00, BKR09, BWG^{+}07, Bra07, Bro03, Bro02a, Cal02, CH09, CMC^{+}06b, CCHV05, CWC^{+}03, CFN03, CBHM07, CH05, Cor00b, Cot01, DGPR05, DMvD07, DWDD03, DCPvK07, DBL07, FT05, FMB01, FGFP08, FHHH00, FHDM00, FRC03, Fur02, GLDD00, GPRD08, GA05, GP00, GR06, Gre08, GFP09, GPZ^{+}05, GF00, GW09, HSCN06, HBG^{+}04, Hea05a, Hea05b, Hea07, HE08, HF08a, HM04b, HR01, HWF08, HOD06, HHC^{+}09, JWM03, JGN04, JV05, JLS02, KI04, KPO05, KF08, LNLS09, LP00, LMVdZ^{+}07, LHJS02, MvdKN05, MKR^{+}09, MUK^{+}02, Mol00, MMS01, Mur00b, NRR^{+}09, NGNB^{+}04, OED^{+}04, ON09, ÖFR^{+}06, PGD09, PRvB00, PHDC^{+}09, PZTE05]. North [PMN01, Pie00, Pie02, PR04, PvHG09, PKH^{+}08, PB00, PH05, RBD^{+}07, RP07, RK04, RNK05, RvMBV00, RW01, Ros05b, RFT02, RPSSW09, RK00, SA05, Sar09, SDCR07, Sim03, Sim07a, Sim09, SB00b, SB01, SS00, SSJL02, SRS^{+}07, SLN02a, SLN02b, SRS^{+}09, SE09, Str05, TM00, TPT^{+}09, TCTC09, Vin01, ZPRJ02, vDEM^{+}00, vDBF^{+}09, vdVBMR00]. north-east [DLR02]. north-eastern [LD05, LP00]. north-western [LdSSG02, MS00]. Northeast [Ber00, DSV^{+}08, MSGC^{+}09, KB07, PK09, ANNG01, BK07, CPR06, ERP01, EK08, FCM09, FPKH03, GD05, HHB^{+}00, HSS^{+}05, HHHH06, Hea00a, Hea00b, Hea05c, HDG02, HAN02, JH01a, JNF^{+}09, Joh02, KPS^{+}05, KMNP01, LDQ08, LGH^{+}09, MASA06, ODRN05, PLJ01, PCS^{+}04, Rob08, RGG^{+}04, SRN00, SP03b, STMM06, SGMN^{+}06, TCSW06, TLMO08, WGMM08, YM00, Ynd01]. northeastern [DNLSM08, FFL06, HM08, KCCM03, OLS00, PPHB00, Sil03, SSC^{+}06, SB00b, SB01, Tan00]. Northern [CRTS04, AG00, BABB08, BdP07, BFK^{+}07, BLMB06, Bro02b, CR04, CRW^{+}01, Dup05, DR08b, EH00, ETB07, FGLT02, GWSV08, GGV^{+}04, HS06, HBG^{+}04, Hen04, His01, HS09, JHKZ09, KKC04, KWBR08, KK06b, NC08, ÖFR^{+}06, PB05b, PLP^{+}07, PAC02, PRF^{+}00, RS04, RLH01, RAKS06, RPK^{+}03, RO02, Ros03, SBG06, SAM09, She05, SSKE06, SAPP04, SOMT00, SM02, TA05, VEP^{+}09, WGLJM04, WCP08, Wie05, WHA08, DHKV01, Mis02, MDM03]. Northumberland [BF04]. Northwest [NGNB^{+}04, SRGC04, BSS07, BRP02, CBBL09, CSVGTP09, DBDA^{+}02, Joy02, MMB09, Ste02, SIT^{+}05, CLR^{+}05, CLK^{+}09, GAA^{+}04, GP00, Hen04, JLR^{+}08, KH03a, KCL^{+}09, LLS00, MCB09, MPN^{+}08, MB05, PPL^{+}07, PF06, PPH09, SRMB07, SMP09, Uye00, WBC^{+}08, WPJ09, YG08]. northwestern [BBÁMC06, BS02, CZC07, KA01, LLHK07, LG08, SDRK00, lPV01, Sab04, TF02]. Norton [HFWB05]. norvegica [CDB05, CRTS04, ETB07, KK06b, OKRK04]. norvegicus [Bri02, CAWD09, CDB09, CEH03, HU04, MS01a, MAAN09, MFA07, PR01, SP03a, STMM06, SNA01, TAC00, UE01]. Norway [AKJ07, Agn08, BRE^{+}08, BA03, CEH03, FLH06, HU04, HSS08, KKF^{+}06, LLHK07, LNLS09, MAAN09, MFA07, OAJ06, PZTE05, PK09, SJ08, SLN02a, SLN02b, STMM06, TAC00, UE01, UR01, WHA08]. Norwegian [SB00b, BANGC02, BHØ^{+}04, BMDBM09, CTW09, DJRO06, DEMD00, DMDE04, EDG03, FGBS00, GSS08, GHI^{+}04, Han06, HM04a, HHH00, HBST02, HSS07, HSS^{+}09, Kaa00, KTRG06, LSGD02, LAO^{+}07, LGR08, NFM^{+}02, OGR^{+}07, PTTS00, PS06, RT03, STAN02, SHAH09, SB01, SOMT00, SBB^{+}05, SSA08, SN08, Tan00, TL05, TR09, TK03, VSÅF05, WvdMF06]. note [Cor01, Cor07]. Notes [SB01]. Nototodarus [TJAS04]. Nova [BMM03, TST^{+}09]. novaeangliae [RSNB^{+}08]. novaezelandiae [O’D03, O’D04]. Novel [GKOV05, HP07, LHHF03]. November [AJ00]. nozakii [HMHI09]. NSW [Cor00a]. nuclear [KHS^{+}08]. number [DLS01, SFM01]. numbers [BCAN^{+}06, CDD^{+}07, MM01]. numerical [BB09, SNM05, Tri00]. Nursery [DRSD09, FRC03, HHMM01, HKBK09, HSS07, SNB^{+}02]. nutation [Ynd01]. nutrient [dLMACC00, SS00]. nutrients [PMB^{+}03b]. Nutrition [Ham06]. Nutritional [KWZ00, Ber00, Ber04, Bjö02, OLB01]. NW [BCD^{+}02, BCL03, IA04, ORVP09].

O [VH08]. O. [RHD09]. obesus [LCRS08]. objective [BMC^{+}07, PCM09]. Objectives [HE08, TR09, Arn00, GSSO00, GFP09, KPS^{+}05, MSI07, PKH^{+}08, Ric00b, SPS00a]. oblivious [Gre08]. oblongum [FSDB09]. Observation [BG04, CKS03, GLDD00]. Observations [FBMR^{+}03, HEGH02, MBM02, WM04, DSJ03, DBS06, KRYL09, LSH^{+}09, NCM^{+}03, ORVP09, SS00, WPJ09, ZKP03]. observed [ADDH04, CD09, DLC03, GBBG06, GHI^{+}04]. obtained [HSPM05]. occupation [BCT05, BBC^{+}04]. Occurrence [RSNB^{+}08, LME05, SAM09]. Ocean [FRC03, SL04, BHM^{+}04, CDBS08, FSFO08, FHDM00, HMMB^{+}08, KCR07, KRYL09, LHJJT04, Wei05, AFGR09, CMC^{+}06b, CZC07, CDBS08, DC05b, GW09, Hen04, KI04, LD03a, LCRS08, LD05, MVMH04, SRS^{+}09, Uye00, WWWB03]. Oceanic [FMK07, HJ03, HSS^{+}05, KWZ00]. oceanica [FTDVC^{+}08]. Oceanographic [PCDM08, BES^{+}01, LD03a, OdSBS09, PH05]. oceanography [PSFY07, ZK00]. Oceans [VHI^{+}04, LC04, VPC^{+}09, VPC^{+}09]. ocellata [FMF02]. October [DBDA^{+}02]. octopiana [GGP07]. octopus [LZS09, CFRM08, FB07, GGP07, KV06]. off [AKJ07, Agn08, Alo01, ADO02, APGC04, BBM^{+}02, BBC^{+}04, BBÁMC06, BML^{+}05, CH06, CPR06, Cor00a, DBS06, DLT^{+}00, DNLSM08, ECC06, FFL06, GMM^{+}08, GG08, GWSV08, GGV^{+}04, GAM^{+}06, GM06, HM04a, HA03, HR00, IA04, JSR06, Lóp06, Mar07, MM03b, MM03a, MC09, MNMG^{+}05, Mow02, MCI03, Nau02, NC08, OV05, OSB06, OB05, ORVP09, PVH^{+}05, RL05, RML06, RTB^{+}05, SRM00, SdFBG01, SNM05, She07, SKC^{+}00, SIT^{+}05, SM02, TST^{+}09, ZMM^{+}07]. offal [GS03]. officinalis [CDR05, KCBC00, WPB^{+}03]. offshore [BMDBM09, CLFS02, DOBT02, EMA^{+}07, HM04b, PCDM08, PFK^{+}09, Sve03, TPT^{+}09, WMÖ06]. offspring [PFLFR08]. Ofunato [HKI01]. ogives [BZRO06, Bro03, VSC06]. Oil [MCM00, CSW06, Hel02, JLS02, LHJS02, PCDM08, PBLFR06, PFLFR08, TT08]. oilfield [CLFS02]. Oir [BMV05]. Okhotsk [MSIL09]. old [Le 09]. oligotrophic [PM04]. olivaceus [KNS^{+}04]. Oman [HAvH06]. Ommastrephes [CZC07]. Ommastrephidae [SKS^{+}00, TJAS04]. on-growing [TSH^{+}06]. Oncorhynchus [BBMS01, aFADN08, MMS01, RHD09, RS06a]. one [PO09, SK04, Zit01]. ongoing [RTB^{+}05]. ongrowing [GGP07]. onset [TAC00, WBC^{+}08]. onshore [MM01]. onto [GH00, OSLO06]. Ontogenetic [MAB^{+}07, GA00, HAN02]. ontogeny [Hor03, MSR03]. Oocyte [NHKJK09]. open [KMG^{+}07, SJM03, PMB^{+}03b]. open-source [KMG^{+}07]. opening [SOB^{+}07]. operating [RPSSW09]. operational [BBR08, CDBS08, HMMB^{+}08, HNK07, KO02, PRB^{+}07, SPS00a, TGS09, TCP05]. opercularis [VBSB07]. Ophiura [BM01a]. opportunities [PBM^{+}04, Sea02]. opportunity [KMV^{+}07]. optical [NGNB^{+}04, RKM09, STA^{+}09]. optics [HDG^{+}09]. Optimal [BSO01, KMJH01, PCS^{+}07b, PCS^{+}07a, BHMS02, RR07, Ros03]. optimisation [KMJH01]. optimization [GRMR07, KYG03, LCC09]. optimized [HMD^{+}08]. Optimizing [MSI07, SGC^{+}09]. options [DB08, HF08b, HK06, ROB04, SK07]. orange [CB07, DRSD09, KH03b, SCHR07]. orca [MS02]. Orcinus [MS02]. orders [Roc00a]. Orectolobus [HOGH07]. Oregon [DGK^{+}09, PP08]. Organic [SAPP04, AS02b, BR08a, HH01, PMB^{+}03b, SVRF08, URMS04, WHP01]. Organism [Cam08]. organisms [Gas02]. organization [JvD07]. organs [Hus04]. orientation [BRC09, SRJ03, ZPK05]. orientation-dependent [ZPK05]. Origin [HJ03, BQHG00, BBSK09, FHJS09, GKOV05, GSS08, MGS00]. Oscillating [MSP09]. Oscillation [GP00, Sar09]. oscillations [SN08]. OSPAR [Ard08, HE08, Joh08]. ostracods [KI04]. Ostrea [CDDM05]. Otaria [GWSV08, HA03]. other [BBR08, Bro02a, CDD^{+}07, Gro06, HW06, LDNS08]. Otolith [BBK08, JCM06, PMB00, SGMN^{+}06, BD02, BK07, CBHM07, CFMdP07, FL06, GWG06, LTA00, LP00, OFN02b, OFN02c, Pel02, Sec02, SGS^{+}05, Str05, WWGG02]. otoliths [Cas07, CMHN05, CFMdP07, DNLSM08, HOF04, Pel02, VH08, WK01, WWGG02]. otter [MF07, MS00, RS03, SDRK00, SPD00, TMG^{+}08]. otter-trawl [MF07, TMG^{+}08]. Our [VPC^{+}09, MHV09]. out-of-kind [BFSC02]. outcomes [LGH^{+}09]. outflow [Ber04]. outliers [VM07]. outplanted [HRB02]. Ovarian [RLH01, BGL08, JRM^{+}03]. over- [PB08b]. over-reliance [She05]. over-wintering [HBG^{+}04]. overall [BGW05]. overcapacity [ET07, STW^{+}08]. overexploitation [WDRP09]. overfished [JGM^{+}08]. overfishing [MMKKJ08, Mur00a]. overflow [RFM^{+}02]. overlap [Bun01, HSPM05]. overlapping [KDP09]. Overview [Ele00, GSSO00, HF08b, KCD^{+}03, RHH^{+}08]. overwintering [HK00]. oxbow [Aka02]. Oxygen [CJS02, HOF04]. oxyrinchus [FSDB09, FSDB09]. Oyashio [KMT08]. oyster [HKI01, LSH^{+}09, dBMS09]. oyster-culture [LSH^{+}09].

P. [MG07]. Pacific [GW09, KI04, NC08, Uye00, AFGR09, BJN^{+}06, Bro02b, Bro02a, BRHG^{+}06, CZC07, CBM09, FSDC09, HMDS09, HMQ^{+}08, KH03a, KCL^{+}09, LCRS08, LCC07, MMS01, MSM^{+}06, NC06, ON09, PH05, STA^{+}09, SHT^{+}09, Som04, SE09, TF02, TCTC09, VH08, YG08]. pacifica [DSJ03]. pacificus [KMI^{+}05, SKS^{+}00]. Pagophilus [CGN^{+}04, FPKH03, HSCN06, PGB03, SFØ07, Ste02]. Pagrus [EB04, KH03a, WM01]. Paguridae [RNK05]. Paguro [PAC02]. pagurus [Ste08, Ung07, UMSA09, WvdMF06]. painted [Aló08]. paints [Hal01]. Pallas [AFHJ04]. pallasi [PMB00, HMDS09, SHT^{+}09]. pan [CBS^{+}06]. pan-European [CBS^{+}06]. Pandalus [FQS01, Har07, HK06, KKC04, MASA06, Wie05]. panel [FJK^{+}07, ZFFT01]. panels [FB02a, SOB^{+}07]. Panulirus [dLMS06]. papers [DPW07, MPD^{+}08]. Paracentrotus [SDG^{+}08]. paradigm [Kar06]. paradoxus [ERBP09, GLR06, SAM09]. Paralichthys [KNS^{+}04]. parameter [BH07, MHH06, MKB01]. parameterization [CD06]. parameters [CMP07, DP03a, DP03b, FPKH03, HC09, KKF^{+}06, LDCH^{+}09, LHJJ^{+}01, Mui03, PRD^{+}06, SZ07, WSWS03]. parametric [MS01b]. parasite [JHKZ09, PAA06, TK03]. parasites [PN06]. parent [PLJ01]. Parental [PKRT06, FHJS09]. Park [SGMMGB09, CBBL09]. Part [MGM03]. Partial [RDD06]. partially [PRvB00]. particle [GTOJA06]. particular [BVD01, MWS04]. particulate [SML01]. parts [NHK09]. Passage [CT07]. Passive [FR09]. Passive- [FR09]. past [CF02, Ric00b, Rot00, SBL07]. Patagonia [Alo01, LME05]. Patagonian [LCC08, MFB^{+}08, SRM00, SdlRdA06a, SdlRdA06b]. patagonica [MFB^{+}08]. patch [KPD^{+}07]. patch-choice [KPD^{+}07]. patches [JGM^{+}08, RvMBV00]. patchiness [BMJ08, MSR03]. Patchy [Kal01]. pathogen [GGP07]. pathogens [SHSKR01]. patronus [BW08]. pattern [FBD^{+}08, MSR03, Mil08, NH09, SRM00, SDWQ09, SB04]. Patterns [Tan00, AVJ^{+}06, ADDH04, BSA09, BPD^{+}03, BRP02, BBMS01, CW09b, CRB08, Cur00, DK00, DLR02, DSV^{+}08, DDR^{+}03, FB07, FJSJBS^{+}08, GSdFB01, HBG^{+}04, HSPM05, JDA^{+}06, JLR^{+}08, MVMH04, MWF^{+}05, OED^{+}04, OSB06, OKRK04, PCDM08, PSO^{+}04, PVLP04, PPH09, SRM08, Ste01a, SNB^{+}02, VSS07, WPB^{+}03, WPR^{+}07, ZK00]. paucity [RD07]. pay [PS09]. Pb [KK06a]. Pb-210 [KK06a]. PCB [Vuo02]. PCDD [Vuo02]. PCDF [Vuo02]. PCR [ISS^{+}07, MYAT09, QGdS04]. PCR-RFLP [ISS^{+}07]. pealeii [MB01]. Pecten [BSMB03, LPL03]. pectinid [LPL03]. Pelagic [PLP^{+}07, ADDH04, BEB^{+}09, Bea05, BCL03, BvKvH^{+}08, BGG^{+}06, DNP03, DB04, DH09, GGV^{+}04, KA01, LBF01, LPL03, MCB09, Mar08, MGS00, MLM02, MKFK05, dLMACC00, MNY^{+}09, OLS00, Pet01, SKR^{+}06, VSÅF05]. pelagic-demersal [dLMACC00]. pelagics [Cur00]. pelagicus [UBP^{+}09]. pelamis [AK04]. pellet [UR01]. pen [SML01]. penaeid [Ye00]. Penaeus [NAK^{+}08, TYH04, YBF^{+}03]. Penalties [PB08b]. penetration [ÖG04]. pengoi [GFH04]. penguins [CUUD07]. Peninsula [OGL06, BBÁMC06, CFRM08, CAGV05, Cas07]. Perca [JRN06]. perceived [MB05]. perceptions [RPB^{+}08]. perch [JRN06]. Perciformes [MLG^{+}09, PSC02, dCA03]. Performance [PR04, BR08a, BÓ06, CSdQB06, DR08a, HTA09, Hus04, Knu09, KBDC^{+}08, MPG^{+}09, WSC^{+}06, WS02a, WS02b, WSP03, ZWW^{+}03]. performances [ES09]. period [PMM^{+}09, Sab04]. periods [KA01, Sim07a]. Persian [NAK^{+}08, Esm06]. persist [BP08, YG08]. Persistence [Mis02, SKC09]. perspective [Aga00, BPM^{+}05, DBDA^{+}02, Her04, Jen02a, KM02, Mur00a, PSFY07, RBD^{+}07, RS06b, Sym07, WJB07, Zit01]. Perspectives [Ano01h, NEJH05]. Peru [APGC04, BDÑ04]. Peruvian [BBGA05, SGC^{+}09]. perverse [DM07b]. pesticide [MD01]. Peter [WBK09a]. phantom [PO09]. phase [DC03a, DC04a, MMKR^{+}00]. phenology [GPZ^{+}05]. phenotypic [Law00]. philippinarum [MCRF06]. Philippines [NC08]. Phillip [Bun01]. Phoca [BBBF02, LLHK07, Lun01]. phocaenoides [XZW05, Aka02]. Phocoena [LHJJ^{+}01, TPT^{+}09]. Photographic [CEV00]. photographs [CLL^{+}09]. photoperiod [KNKT06]. Phyllospadix [HRB02]. Physeter [EH04]. Physical [Bot01, DF00, Des00, PCM01, TDE08, ANNG01, EMA^{+}07, KIDY09, LC09b, MAAN09, Mol00, Orl01, WSWS03, WFIM00, WS02b, ZK00]. physical-biological [Mol00]. physicochemical [Tri00]. Physiological [Hus04, MF07, SAMS02]. physiology [Hor03]. Phytoplankton [TCM^{+}08, WU03, ASC01, CH00, ERP01, Gaa00, ML08, PMB^{+}03b, SG05, VM09]. picarel [ÖTTM07]. picture [Hal01]. pictures [TPRR04]. pigeon [LPA^{+}00]. piked [BGW05]. pikeperch [AP07]. pilchardus [Cas07, CLK^{+}09, MCI03, Sil03, ZMM^{+}07]. pilot [dPBB^{+}03]. pingers [CBDB02]. pink [BJN^{+}06, ELR01, HB07]. pinpointing [BG07]. piscatorius [His01, LDQ08, LDML08, MSB04, MD01, VLBB08, WWGG02]. Pisces [Alo01, KTM^{+}05, LC06, OV04, SB01]. place [NM09]. Placentia [RR06]. Placopecten [HS06, HC09, SBT^{+}09]. Plaice [KF08, Arm01, BHMD05, BKR09, DCPvK07, HBS^{+}06, HF08a, KBW09, Mor04, NWH02, PRvB00, PCS^{+}07b, PCS^{+}07a, Pie02, PKH^{+}08, SPK05, vDBF^{+}09]. plan [KCR06, OGR^{+}07]. planktivorous [Kaa00]. Plankton [Pie08, PF06, Bea05, BDS01, CMM03, GQCÁMI03, HP05, HEGL05, KFM02, NGNB^{+}04, OBD^{+}05, PÁMGV05, RBBB00, SB00a]. Planning [SW02, MBC^{+}09, SRM08]. Plans [AWW^{+}07, KDCH^{+}09]. plastic [GC02]. plasticity [DCCS09, Gef09]. Plata [CMC^{+}06a]. platessa [Arm01, BKR09, KBW09, NWH02, PKH^{+}08, SPK05]. platessoides [BHMD05, MMM00, Mor04, PCS^{+}07b, PCS^{+}07a]. platform [CA02, FGLT02, JLS02, KMV^{+}07, SSJL02]. platform-of-opportunity [KMV^{+}07]. platforms [FSB^{+}03, Hel02, LHJS02, PCDM08, PAC02]. Platichthys [Jag02, KBW09]. platypterus [HL07]. play [KMHS04]. playback [RLF01]. Pleuronectes [Arm01, BKR09, KBW09, NWH02, PKH^{+}08, SPK05]. Pleuronectidae [MMM00]. plumbeus [MSW07]. plume [ORVP09, PP08]. point [GLKPCP01, KKC04]. points [HS07, HP04, JD05, LKL08, MRT01, PR04, PKH^{+}08, PPC^{+}03, RL07, SH06]. Polar [UR01, CMHN05]. Poleward [SG00, IA04]. policies [KMM07, WCMK05]. Policy [SBC^{+}07, Bac08, Pen07, Ray07, Ste01b]. politicians [Gre08]. Pollachius [AJNM07, CDQL06, NCM^{+}03, NSP06, PK09]. pollack [CDQL06]. Pollicipes [BLMB06]. pollock [APD09, BWC00, DW06, HH04, HHMM01, HIL00, Hor03, HSA^{+}09, KHO06, KK06a, LBNS00, NCM^{+}03, SBC^{+}00, SDWQ09, Som04, SNB^{+}02, Wal07, WFIM00, WS06, NSP06]. Pollution [BGG^{+}08, MVM^{+}08, SDG^{+}08]. polyamine [NNT01]. Polychaeta [Vor00]. polychaetes [vdMBD00]. Polychlorobiphenyl [MD01]. Polynesia [AGY^{+}05, AFGR09, GAYR06]. POM [PST^{+}07]. Pomatomus [CAAJ07, LO05, MC00]. poor [Mac09]. Popp [Chr02]. Population [BSMB03, Buc00, CAGV05, DC05a, DM04, Hen04, RNS08, SMP09, TM00, TAC00, TYH04, ZCH06, vdMBD00, AVJ^{+}06, Aco02, AGY^{+}05, AMJ^{+}06, AMGV06, BMV05, BABB08, BMM03, BVB^{+}07, CF05, CDQL06, CMJ09, CSH00, Cor01, DCN^{+}04, DH08, DP03a, DP03b, EH00, FMF02, FQS01, GAYR06, HP07, JP03, JHC09, JMLG05, Kar06, KM05, KV06, KKF^{+}06, LNLS09, LT06, LJM00, MHH06, MMCD08, MPG^{+}09, MG02, MM05, QGdS04, RF01, SRM00, SGMV^{+}08, ŠCBD09, SFØ07, SMEK01, SRMB07, SNA01, Ter02, TLM04, TS05, TLMO08, WBC^{+}08, WGMM08, WJB07, XZW05]. population-based [MM05]. Populations [DF00, AFP^{+}09, BHN06, BvS00, BF02, BFK^{+}07, BMJ08, Bot01, BDTW06, BD02, Buc00, CDB09, CRTS04, DiUVH08, DJRO06, ERGT07, FLH06, FGBS00, GSN^{+}03, GP00, HJB^{+}08, HOP09, HTSB04, LLD^{+}05, LPH^{+}08, Niw07, PSSD08, PGG05, PCD05, PBM^{+}04, RPE^{+}09, RMB^{+}09, RMM05, SCCM06, SSKE06, Sil03, SWG06, SGY08, STMM06, TPRR04, WB05, dBMS09]. porbeagle [Joy02]. Porcupine [RMM05, MBM02, VLBB08]. porpoise [CBDB02, XZW05]. porpoises [Aka02, LHJJ^{+}01, TPT^{+}09]. port [GHFA09, Bun01, CBBL09]. Port-Cros [CBBL09]. Portugal [STM^{+}08, BMP^{+}08, CEH03, FSQ^{+}03, GMM^{+}08, GSdFB01, Mar07, SR03, SdFBG01, SM02, ZMM^{+}07]. Portuguese [CSVGTP09, DAAD09, Gas02, SNM05, SNV^{+}09, VGF03]. Portunus [UBP^{+}09]. posed [HF08b]. Posidonia [FTDVC^{+}08, SJGRRRE02]. position [ES02, HHT08]. positions [HMD^{+}08]. positives [But07]. possibilities [BHN06]. Possible [Pet04, WCMK05, BBR08, BD04, Cor00b, CSdQB06, DM07a, LN08, PTTS00, WJTH00]. post [BLRC05, Cam08, DH07, Des00, FHDM00, FRC03, FCM05, HHHH06, HHH00, MSR03, Miy03, TB02, WS02a]. post-deployment [WS02a]. post-dredging [BLRC05, Des00]. post-hurricane [TB02]. post-incursion [Cam08]. post-larvae [Miy03]. post-processing [DH07]. post-settled [MSR03]. post-smolt [FHDM00, FRC03, FCM05]. post-smolts [HHHH06, HHH00]. Potential [BSS07, CSVGTP09, ET07, GH04, PQRG07, SHdLP04, Bun01, KRYL09, LBL06, Law08, MB05, Nie00, ÓT06, PB05a, PGJ^{+}05, PJ05, PJR08, PF08b, ŠCBD09, SRS^{+}09, TH08a, TB02]. Potomac [RRC03]. pout [HSS08, LNLS09, PK09, SLN02a, SLN02b]. poutassou [BK07, JG07, RMM05, WGMM08]. Power [BMJ08, MSB04, MNHL01, MUK^{+}02, NJ04, RDHP00]. practical [CdlMA^{+}00, MYAT09, SFH^{+}07]. practice [AVK^{+}08, GFKM07, OSK^{+}05, SEOR09]. Praunus [CDB05]. prawn [NAK^{+}08, SJM03, TYH04, VEP^{+}09, YBF^{+}03]. prawns [MBPW06]. Pre [TB02, BD03, JWM03, LMC^{+}01, NCM^{+}03, ÓMP^{+}04, PCS^{+}04, SHS01, WS02a]. Pre- [TB02, WS02a]. pre-fishery [ÓMP^{+}04, PCS^{+}04]. pre-recruit [SHS01]. pre-recruitment [BD03]. pre-settled [JWM03]. pre-spawning [LMC^{+}01, NCM^{+}03]. precautionary [GD05, HS07, HNK07, ÓMP^{+}04, PR04, Ric09]. preceding [NEJH05]. precipitate [She05]. Precisely [DCPvK07]. precision [HCEM06, MMF09, SGS^{+}05, Wal07]. Predation [GFH04, LHR02, TC01, CFN03, Dup05, GBBG06, HP07, HEGH02, HIL00, Joh02, LHJJT04, LBNS00, MLOT09, PLJ01, SLN02a, TL05, TLMO08, Wie05, vdMBD00]. predation-based [CFN03]. Predation-mediated [TC01]. predator [BBC^{+}04, BGW05, DBBM01]. predator-prey [DBBM01]. predators [CHB09, KM00]. predatory [OLS00]. predict [RL08]. predicted [EDG03, Hor08]. Predicting [FL06, GV02, HHJK06, TH08a, Dav07, FMF02, HSS^{+}05]. Prediction [BMDBM09, LG08, MHV09, WS06]. Predictions [RL08, BHN06, BSMB03, FGR04]. predictive [BREB09, OH07]. predictor [Fra06, PGG05]. predictors [BFMJ03]. Preface [FLH04, Hol00a, Hol02]. preference [FT05, Lun01]. preferences [BOC^{+}08, CCA04, DOBT02, WSFH02]. Preliminary [CH06, DBS06, AS02b, Buc00, JGM^{+}08, dLMACC00, SNM05, dPBB^{+}03]. preparation [SGS^{+}05]. Prerecruit [DR08b]. prescriptive [TRM07]. PRESEMO [OH07]. presence [CH00, LCRS08, Ost09]. presented [DPW07, MPD^{+}08]. prespawning [Nau02]. Pressure [Daa05, EMA^{+}07, HSA^{+}09, PQRG07, SRM08, Wie05, vdKRS^{+}07]. pressures [Ric09]. Prestige [PBLFR06, PFLFR08]. Prevalence [HSM09, UPK^{+}08, SNA01]. Preventing [WDRP09]. previously [DTC01]. Prey [Bun01, SDCR07, BCAN^{+}06, BBC^{+}04, BGW05, DBBM01, FT05, FHJS09, HSPM05, MPG^{+}09, MNCU09, OBNU02, PF08b, RNWS08, TF02]. Pribilof [LBNS00, SBC^{+}00, SNB^{+}02, WS06]. price [SP05, STW^{+}08]. primary [LC09b, MSIL09, SBR07, SVRF08]. Prince [Bro02b, HMQ^{+}08]. principles [BHH^{+}08, PG08]. prior [BBPW07]. prioritization [KWL^{+}02]. prioritize [Fle05]. prioritizing [AGH^{+}09]. probabilities [BMLH07]. probability [BRE^{+}08, HSPM05, Mor04, NB08, RHBR04, SP07a]. probability-based [NB08]. probes [GAW^{+}08]. problem [CF05]. problems [GMGN06, MPD^{+}08]. procedure [But07, DB04, HS01, KS08, KPK^{+}05, RPSSW09]. procedures [PMB^{+}03a, PRB^{+}07, PD07, RPB07]. proceedings [DKMO09]. process [HB09, KKC04]. processed [MS09]. processes [BHH^{+}08, BES^{+}01, FSFO08, FGBS00, KIDY09, KMH^{+}05, Mil02, NTSM07]. processing [CTW09, DH07, Dor01, KO02, OM05]. Procrustean [Dek00b]. produce [Agn08]. produced [BML^{+}05, GF00, HBC01]. production [AFP^{+}09, AWW^{+}07, Arm01, Bar05, BDS01, Bri02, CH06, CGS09, DDM^{+}05, Fox01, GA00, GQCÁMI03, HSCN06, HHJK06, KNO00, LC04, LC09b, MGTS00, MC07, MSIL09, MWS04, OSWL02, PSC02, PGB03, SBR07, SPG^{+}04, SCLK01, SG05, Ste02, dBMS09, vDBF^{+}09]. productivity [CLR^{+}05, HP05, HEGL05, MB05, SKC09, SFKC02, SP02, WGLJM04]. products [WS02b]. profile [JGN04]. profiler [SYR^{+}08, ZKP03, ZPK05, ZCR09]. profiles [Mar08]. profound [HSM00]. Programme [PPH09, NEJH05, ZNGF02, Fru02]. programmes [Gro06, MSM^{+}06, WBC^{+}06]. Progress [DM06, RMB^{+}09, vdVBMR00, CRC^{+}09, WBC^{+}06]. projection [RUA07]. projections [MHH06]. Projects [ATM02]. prolonged [BDO^{+}04]. properties [BHM^{+}04, BP07, CGN^{+}04, CWC00, CWC^{+}03, FR09, MANT07, Ric09]. proportion [Gas02]. proportions [FLK^{+}09, Kol06]. proposal [RBD^{+}07]. Proposals [KDO^{+}08]. prosper [Uye00]. Protected [CBBL09, CW09a, CHB09, ERBP09, Hal01, HKBK09, HPR09, KPD^{+}09, LPSL09, MMB09, SEOR09, TSK00, VBF09, YSF09, AGH^{+}09, CW09b, Jen09, KS08, KF08, MBC^{+}09, SGAC00, WDRP09, WLK02]. protecting [Pow00]. protection [AGH^{+}09, MSGC^{+}09, PPW^{+}09, SF09]. protein [JKSO06, OSLO06]. Protista [GGP07]. proto [HP07]. proto-moment [HP07]. protocol [AE02, VBF09]. provide [HMMB^{+}08, TPRR04, TLM04]. provides [PSSD08]. providing [HNK07, PR04]. Provision [CLR^{+}05, ÓMP^{+}04]. proxies [DRDC06, VSC06]. proximate [CCC02]. proxy [dLMACC00]. Psammobatis [MG07]. Pseudoterranova [MMM00]. puberty [KNKT06]. publication [Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i]. Publishing [PK07]. Puerto [ZAJ01]. punctata [ISS^{+}07]. pup [HSCN06, PGB03, Ste02]. purse [BBGA05, GDL04, HA03, LCRS08, LO05, D00, SSI07, SM02]. purse-seine [GDL04, LCRS08, LO05, D00, SSI07]. purse-seiners [BBGA05]. purse-seining [HA03, SM02]. Pusa [KKF^{+}06]. pusillus [MKR06, MKR06]. pygmy [SAM09]. pyrifera [DDGJ02]. pyruvate [Ber04].

Qatar [YBF^{+}03]. QTC [EGB02]. quahog [BABB08, KCR07]. qualitative [FB02b, Fle05, TH08b]. qualities [HOHS05]. Quality [BD04, HE08, WMS^{+}03, BÓ06, DAH^{+}08, MMV^{+}08, NB08, PGMB09, Sim09]. Quantification [MSB04, Ost09, SSJL02]. quantified [Mol00]. quantify [HRM04, HK06]. Quantifying [TJG^{+}09, RCBM05, SOB^{+}07]. Quantitative [CC05, HOGH07, SP07b, CFMdP07, FB02b, OSWL02, Pie00, SBP07, SA03]. quantum [SH06]. quarry [WS02b]. queen [VBSB07]. quinquedens [Tal07, WBC^{+}08]. Quoddy [SMEK01]. quota [MMCD08].

R. [SB01]. Ra [KK06a]. Ra-226 [KK06a]. racks [LSH^{+}09]. radiata [SB00b]. radiated [BPD^{+}03, GEM01]. radiation [SBB^{+}05]. Radiotracer [FR04]. rafts [LPH^{+}08]. raggedtooth [DBS06]. rainbow [RS06a, RPE^{+}03, SW06c]. rainbow-smelt [RPE^{+}03]. Raitt [TM02]. Raivavae [AFGR09]. Raja [SPFF^{+}08, SB00b, SB01]. Rajidae [Alo01, MG07, MC09, OV04, OV05, SB01]. Rajiformes [ECC08]. Rajoidei [ECC08]. ranched [SW06b]. Random [Niw07, Cor07, HP04, KKC04, TS05]. Random-walk [Niw07]. randomly [CFMdP07]. Range [BABB08, EGO^{+}07, FM04, JDA^{+}06, ODCN09, Ped05, RMM05]. Range-wide [BABB08]. ranging [Bjö02, CGN^{+}04]. ranked [WYM09]. Rapana [SDÖ09]. RAPD [MASA06]. Raphaël [SF09]. Rapid [RMDB05, GKOV05]. Rapid-response [RMDB05]. Rapido [PRF^{+}00]. rare [TM09]. rare-earth [TM09]. rate [Bjö02, BS02, Coo04, GKFM09, JKSO06, Mye01, She05]. rates [AE02, BVDS08, Dav07, Dem01, GF01, GA00, IWP00, KMV^{+}07, LHJS02, MCB09, PMN01, RR02a, SK04]. ratio [CM00, DH07, Iri00, KRM05, LHHJ^{+}09, dLMACC00, VM07]. ration [GKFM09]. rays [SBDW00]. razor [DSG05]. rDNA [KHS^{+}08]. Reaction [MFIO04, BHMD05, HDG02, VGBH09]. readers [SGS^{+}05]. reading [Ree03]. readings [MN02]. real [MYAT09, PO09]. real-time [MYAT09]. realism [BBPW07]. realistic [GNC08]. really [DCCS09]. reanalysis [vDBF^{+}09]. reared [SJKN^{+}04, SB06, SSKE06]. Rearing [HRHC00]. reasoning [WvdMF06]. REBENT [EHG06]. rebuild [RO02]. Rebuilding [JGM^{+}08, GRMR07, Gro06, SP07a, SGY08]. recapture [AJNM07, BVB^{+}07, NSP06, SSKE06]. receivers [HMD^{+}08]. recently [SQN08]. recommendations [BCAN^{+}07, SC00, WBC^{+}06]. Reconciling [DC03a, DC04a, KGRW07]. reconstructed [HDG02]. Reconstruction [BPD^{+}03, AP09, BGW03, Kal01]. record [SMEK01]. recorded [DHWW08]. Recorder [PF06]. recordings [HMAN03, JG07]. recover [ÅD07]. recovering [MS07]. recovery [Bra07, FSDC09, GGM^{+}05, HOD06, KDCH^{+}09, KCR06, LHHF03, MML^{+}00, MMF09, SK04]. recreational [APGD08, Aló08, BBB06b, LPSL09, MNMG^{+}05, PS09]. recruit [BP07, Kat05, KB07, RKP03, Roc00b, SMK08, SHS01, UP00]. recruit- [SMK08]. Recruitment [DF00, FMB01, GAA^{+}04, MC00, PHDC^{+}09, SHT^{+}09, ZK00, vdVBMR00, AO08, BMV05, Bot01, BDS01, Bra05, BD03, CFL00, CDR05, Cor01, DAd02, DCD00, DGK^{+}09, DR08b, ELR01, FBD^{+}08, Fox01, Fra06, FMK07, FMH^{+}09, HSS08, IWP00, KMM07, KM00, KMH^{+}05, Kup04, MFB^{+}08, MM01, MLM05, Mye01, NKOK00, NK00, NTSM07, ÓMP^{+}04, PGG05, PGJ^{+}05, PCRW04, PLJ01, RL05, RDB09, Rot00, STAN02, SG00, SdFBG01, Sec00b, Sve03, SNB^{+}02, TC01, WFIM00, WJTH00, Wie05, Ye00]. recruits [BO05, JP03]. Recuperation [NC06]. Red [AS02a, FBD^{+}08, GHBR08, KH03a, LTA00, MLG^{+}09, Pel02, Rob08, SBG06, SZ07, Tal07, WBC^{+}08, WSFH02]. redfish [DNP03, DLT^{+}00, GR01, GR02, GR05, Ped05, SRGC04, STG06, SKR^{+}06, SGS^{+}05, Str05]. reduce [AS02b, BMU09, CH00, CBDB02, Den08, FWW06, RLF01, TM09, WM01]. reduced [DHWW08]. Reducing [HB07, MBPW06, BPT09, Ska07, TJG^{+}09]. reef [ASB05, BFSC02, CSR^{+}02, CW09b, CHB09, CJM^{+}02, CJS02, DDGJ02, DOBT02, FS02, FLP^{+}02, FB02b, FFL06, GAZ02, LPM^{+}09, MMKR^{+}00, Mil02, Mor02, PSC02, PPW^{+}09, RRTdA02, RRTP02, SJGRRRE02, SMG02, SW02, Sea02, SGS02, SFKC02, SP02, WS02a, WS02b, ZNGF02]. Reefs [Fru02, Jen02b, AS02a, AS02b, BCD^{+}02, Bai02, CWYM^{+}02, CA02, DGMM02, GZND02, HEGH02, HS09, Jen02a, KWL^{+}02, KTS02, LSM07, LMU^{+}02, PBH02, PAC02, wScY02, SSJL02, TB02, Vor00, WLK02]. Referees [Ano01d, Ano01g, Ano02d, Ano03b, Ano05a, Ano05b, Ano06k, Ano06l, Ano06m, Ano07b, Ano07c, Ano00b, Ano00c, Ano00d, Ano01e, Ano01f, Ano02a, Ano02b, Ano02c, Ano03a, Ano04, Ano07a, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c]. Reference [JD05, SH06, BVD01, Cha04, CSdQB06, GR06, HS07, MMC03, MRT01, MWS04, PR04, PKH^{+}08, PPC^{+}03, RL07, SDCR07, TCP05, WS06]. refined [MKR06]. refining [DB04]. reflections [Ric00b]. reflex [Dav07]. reform [Sym07]. refuges [Aco02]. refugia [TB02]. Regan [OV05]. regime [AMD^{+}05, FHJS09, HFWB05, Lit06, MMKKJ08, RO05]. regimes [HLCG04, MC07, OdSBS09]. Region [SMEK01, BDTW06, CAAJ07, GL00, KMT08, MSIL09, PPH09, SR03, SBC^{+}00, SPFF^{+}08, SSU^{+}09, WYMF08, WW01, ZCH06]. Regional [BWG^{+}07, Hea05c, TCC08, AMGV06, Bac08, ERP01, GML06, GFP09, HWF08, PLP^{+}07, PFF01]. regional-scale [GFP09]. regions [BWC00, HHB^{+}00, OAJ06]. regular [HP04]. Regulating [BBS09a, AG00, CRTS04, FGBS00]. regulation [íJCMR07, Sve03]. regulations [HF08a, STJ^{+}07]. regulatory [BBR08, Joh08, RHH^{+}08]. Reinhardtius [ANNG01, HAN02, JRM^{+}03]. relate [MANT07]. Related [Jen02b, BJ00c, CCHV05, FGFP08, GQCÁMI03, GKFM09, HMMB^{+}08, LBN09, NEJH05, SN08, Ye00]. relating [CBM09]. relation [ANNG01, BGL08, BHMS02, CR04, DSG05, DCM03, Fox01, Gaa00, GAP^{+}00, GDH02, GA05, GJR04, HSPM05, HSS07, JWBP07, JWM03, JV05, Kaa00, LND05, MR05, MAAN09, NWH02, OMBP06, ODCN09, OLB01, PLP^{+}07, Pet01, PS06, RLH01, SAN^{+}05, Ste01a, WGLJM04, ZK00]. Relationship [SG05, ES09, MvdKN05, MVMH04, Mor04, Ona03, OBD^{+}05, RASS09, SNA01, SNB^{+}02, dPGPB06]. Relationships [BLMB06, FLH06, RM01, SAAFCA07, BMV05, BBC^{+}04, CMDN02, Fra06, GDL04, KMM07, LLC^{+}08, LDCH^{+}09, MS09, MLM05, MHV09, MNY^{+}09, RNK05, RRTdA02, SQN08, WFIM00]. Relative [HLS00, BHM^{+}04, CG07, KO03, LBNS00, PK09, vdMBD00]. release [SK04, SW06b]. release/one [SK04]. released [Aló08, BBMS01, BBB06b, Han06, WBD^{+}06, Wie05]. relevance [EJR01]. relevant [WGM04]. Reliability [LGR08, RGG^{+}04]. reliance [She05]. relief [CKS03]. remarkable [AGY^{+}05]. Remarks [WBK09a]. Remote [Bro02a, AGY^{+}05]. remotely [PH05]. removal [Kor00, LSH^{+}09, RKE06]. remove [DH07]. rent [ET07]. reorganization [Lit06]. repeat [ETB07, SMK08]. repeat-spawning [SMK08]. repeatability [WMS^{+}03]. repeated [ODCN09]. Repetitive [BDD06]. report [OB05]. reporting [DBS06]. representative [AGH^{+}09, GOK05]. Reproduction [CLM07, GG08, LME05, Har00, JRCS08, Kup04, LCC07, SBB^{+}05, Vel02]. Reproductive [ECC06, LME02, MNGB07, NAK^{+}08, NWH02, OV05, RLdAW06, SDÖ09, VGF03, AAV^{+}04, AVJ^{+}06, AKJ07, Bro03, CAAJ07, CLM07, CRTS04, DSG05, ECC08, FBMR^{+}03, HL09, Kup04, LHJJ^{+}01, MSP09, MS04, MB05, Mye01, Nie00, ÓT06, Ung07, YW05]. requiem [MMC03]. requirements [Hea05c, PQRG07]. research [CBM09, DHWW08, HDG^{+}09, JvD07, MV07, MLM02, MMCD08, PO09, PPB03, Sea02, SSI07, TMB08, WYM09, WS02a]. Reserve [BLMB06, EB04, HHAB09, RSC^{+}09, WBK^{+}09b, WBK09a]. reserves [DLC03]. Residence [JLS02, HL07]. resident [MG02]. residues [MD01]. resilience [Sec00a]. resistance [SOB^{+}07]. resistivity [CWC00]. resolution [Cla00, HNLR04, MTJ^{+}07, PF08a, PST^{+}07]. resolved [CMJ09, HTA09]. Resolving [RPE^{+}09]. resonance [DLS01, GPP09]. resource [ERBP09, GJL08, HB09, MNMG^{+}05, PM06, RMDB05, RAR^{+}07, STW^{+}08]. Resources [CdlMA^{+}00, DWDD03, SB00a]. respecting [BBPW07]. respiration [PG08]. respond [DW06]. Response [PJ05, PPMH04, BR08a, BB09, BPT09, CMO^{+}06, DTC01, Dri05, GBBG06, GP00, JPO09, KO03, MSP09, OdSBS09, PK09, RMDB05, RCBM05, RPK^{+}03, Som04, Tri00, Wei05, vDEM^{+}00]. responses [HBW^{+}09, HSA^{+}09, LPM^{+}09, MPG^{+}09, PR07, PFLFR08, Ros05b, RG07, SE09]. restocking [PS09, SB06]. Restoration [WLK02, AMJ^{+}06, Des00]. restored [SCCM06]. restrained [Dav07]. restrictions [HHJK06]. resulting [WHP01]. Results [EZ03, TM00, AS02b, BF04, BLRC05, DM06, Erm09, HSCN06, HFMD06, NCC^{+}07, PPKM07, PRD^{+}06, ZNGF02, dPBB^{+}03]. retention [HSA^{+}01, TH05]. retrieval [LGH^{+}09]. Retrospective [GHBR08, SVRF08, CF05, SLvdB^{+}09]. return [BPM^{+}09, PR01, RR02b]. Returns [WC01, MS09]. reveal [MNY^{+}09]. revealed [DC05a, JGN04, ZCH06]. reveals [EB04, GML06, MLG^{+}09]. revenue [Kat05]. revenue-per-recruit [Kat05]. reverberation [CD03]. reversals [DGK^{+}09]. Review [GJR04, SC00, BCAN^{+}07, Box06, BBB06a, BRHG^{+}06, CBM09, FRK02, GC05a, HR00, JJ06, LDQ08, LPL03, MSM^{+}06, PPMH04, Rye04, Ste08, SMP09, UP02]. reviewers [Ano09d]. Revisiting [PLJ01]. rewards [PB08b]. reynaudii [OSB06, OR09, Rob05]. RFLP [DC05a, ISS^{+}07, QGdS04]. Rhine [BVB^{+}07]. Rhizoprionodon [MNGB07]. Rhode [AO08]. Rican [ZAJ01]. rich [MNCU09, PPC^{+}03]. Richmond [Cor00a]. richness [HHJK06, LSGD02]. ridge [HR09]. right [DCPvK07]. Rigs [Fru02, SSJL02, Bai02, CA02]. Rigs-to-Reefs [Fru02, SSJL02, Bai02, CA02]. ring [Tal07]. ringed [KKF^{+}06, LLHK07]. ringens [BDÑ04, BBGA05]. rings [LP00]. Rio [GAZ02, ZNGF02, CMC^{+}06a]. Rise [HBD05, YCCH07]. Risk [Her04, AFHJ04, BHH^{+}08, BO08, CG07, Cam08, Fle05, GRMR07, HTA09, HK06, MS04, RL07, Ste01b, TH08a, TH08b]. risks [BFK^{+}07, HF08b, MLLK09]. Risso [RMM05]. Ritter [DCN^{+}04]. River [Aka02, AMJ^{+}06, Cor00a, MCRF06, OCWV06, OGL06, ORVP09, PP08, RRC03, WC01, AVJ^{+}06, BFMJ03, CMO^{+}06, DP03a, DP03b, HAG^{+}08, JWBP07, AWW^{+}07, BMV05, BVB^{+}07, CBDS08, PMB^{+}03b, Sim07b, WJB07]. rivers [BGG^{+}08, GHD^{+}09, RPK^{+}03, SCCM06, TES^{+}05, UKR05, WBC^{+}06, WW01, YMF02]. robust [Fik00, PD07]. robustly [FSP05]. rock [AS02a, GG09, LHHF03, MM01, dLMS06]. rock-aggregated [AS02a]. Rockall [HR01, MD01]. Rockfish [ŠCBD09, KH03a, LC06, SKC^{+}00, SKH02]. rocky [BCD^{+}02, CHB09]. rocky-reef [CHB09]. Rod [YMF02, TES^{+}05]. Role [ÖG04, Ste01b, VHI^{+}04, BCT05, BR00, Bra07, BDO^{+}04, CVL^{+}09, CHB09, CTLN09, Dol02, FCM05, FQS01, HP05, JR07, Jen09, KMHS04, KRM05, Kos09, Ped05, PCD05, PMB^{+}03b, RRT00, RR07, She07, SGY08, SS07, WWHB04, vdMBD00]. roles [Cal08]. ROPME [HAvH06]. rotational [dBMS09]. rotifer [OMBP06]. rotifers [PPK^{+}06]. roughskin [LCC07]. roughy [CB07, DRSD09, KH03b, SCHR07]. round [KASA07]. roundfish [KPK^{+}05, KPK^{+}06]. Route [JWBP07]. routes [BVB^{+}07]. routine [EDG03]. Roux [SF09]. RoxAnn [HNLR04, RM01, BML^{+}05]. RoxAnn^{T M} [PR01, PR03, WSP03]. RPR [Kat05]. rubens [BM01a]. rudis [MG07]. rule [SZ07, STJ^{+}07]. rules [Cla00, MC07, RL07, RD07, TR09]. run [Cor00a]. run-off [Cor00a]. Running [Dup05, YMF02]. Russia [OGL06, PGB03]. Russian [CMK09].

S [GGV^{+}04, LTA00, MVMH04, SQN08]. S. [SRGC04, SGS^{+}05, Str05]. Sabellaria [Vor00]. Sable [BMM03]. sablefish [OB05, SGM09]. sagax [CR04, GWF01, War01a]. Sagitta [DCN^{+}04, GG08]. sagittal [DNLSM08]. Saharan [BQHG00]. saida [CMHN05]. sailfish [HL07]. sailing [HP04]. Saint [SF09]. saira [STA^{+}09]. saithe [AJNM07, FT05, ISHB07, PK09]. Sakhalin [Vel02]. salar [AMJ^{+}06, AMGV06, BMV05, Ben01, CLR^{+}05, Cip09, CSC^{+}04, Dem01, DP03a, DP03b, FLH06, FHDM00, GKOV05, GSS08, Gro06, HJ03, Han06, HHHH06, HHH00, HAG^{+}08, JH01a, JSMK06, Kol06, ÓMP^{+}04, OCWV06, OGL06, PCS^{+}04, SCCM06, SWG06, SW06b, SHAH09, SRS^{+}07, SVRF08, TES^{+}05, Vuo02, WBC^{+}06]. Salinity [JHKZ09, CSW06, CVL^{+}09, HU04, MMS01, NWH02, Sar09, SLvdB^{+}09].

Salmo [AMJ^{+}06, AMGV06, BMV05, Ben01, BF02, CLR^{+}05, Cip09, CMO^{+}06, CSC^{+}04, Dem01, DP03a, DP03b, FLH06, FHDM00, GKOV05, GSN^{+}03, GSS08, Gro06, HJ03, Han06, HHHH06, HHH00, HAG^{+}08, JH01a, JSMK06, Kol06, ÓMP^{+}04, OCWV06, OGL06, PCS^{+}04, RAKS06, SCCM06, SWG06, SW06b, SHAH09, SRS^{+}07, SVRF08, TES^{+}05, UPK^{+}08, Vuo02, WBC^{+}06]. Salmon [BF02, HW06, AMJ^{+}06, AMGV06, BPM^{+}09, BMV05, BJN^{+}06, Ben01, BFK^{+}07, BDTW06, CW06, Cha04, CLR^{+}05, CBM09, Cip09, CMK09, CMM01, CSC^{+}04, Dem01, DP03a, DP03b, FLH06, Fjä05, FWW06, FHDM00, FRC03, FCM05, FMH^{+}09, aFADN08, GKOV05, GSN^{+}03, GSS08, GHD^{+}09, Gro06, HMK^{+}07, HJ03, Han06, HHHH06, HBC01, HTSB04, HFMD06, HHH00, HAG^{+}08, JH01a, JR01, JJ06, JSMK06, Kol06, LS04, LFW03, MMS01, MSM^{+}06, NEJH05, ÓMP^{+}04, OCWV06, OGL06, PFF01, PCS^{+}04, PPC^{+}03, RPK^{+}03, SJKN^{+}04, SCCM06, SSKE06, SWG06, SW06b, SW06c, SHAH09, SRS^{+}07, SVRF08, TES^{+}05, UKR05, Vuo02, WBC^{+}06, WC01, WBD^{+}06, WHP01, YMF02]. salmon-trap [FWW06]. salmonid [BRHG^{+}06, SML01]. salmonids [BHN06, RHD09]. salmonis [BF02, BFK^{+}07, GSN^{+}03, UPK^{+}08]. salpa [JDA^{+}06]. salt [BD07]. saltatrix [CAAJ07, LO05, MC00]. saltmarsh [TDE08]. Salvelinus [BF02]. same [JNF^{+}09]. sample [CLL^{+}09, SP03b]. sampler [PCS^{+}07b, PCS^{+}07a]. samples [FGR04, GPWG04, NGNB^{+}04, PÁMGV05, SHdLP04]. Sampling [ATM02, PMB^{+}03a, Bro03, CS02, ETB07, HP04, Kal01, KKC04, LFD^{+}09, LCC09, MSI07, NB08, PZTE05, RAR^{+}07, RKE06, RUCG07, RBGJ08, VM07, Wal07, WYM09]. San [TSK00]. sand [BLRC05, BBB06b, GLDD00, LME02, MG07, SPGT00, vDEM^{+}00]. sandbank [NPPO06]. sandbar [MSW07]. sandeel [Fur02, HKD^{+}04, JWM03, JPO09, KGT01, LPA^{+}00, TM02]. sandeel-dependent [Fur02]. sandeels [BWG^{+}07, Coo04, GAM^{+}06, PCRW04]. Sander [AP07]. sandy [SPGT00]. sandy-bottom [SPGT00]. Santa [CLFS02, Mor02]. Sarcophyton [CVL^{+}09]. Sardina [Cas07, CLK^{+}09, MCI03, Sil03, ZMM^{+}07]. Sardine [SdFBG01, SM02, BCT05, BIdL^{+}08, CR04, Cas07, CvdLHF08, CD03, GCS^{+}04, GMKS06, KWZ00, MCI03, SNM05, Sil03, SSC^{+}06, SSU^{+}09, SHS01, VH08, War01a, ZMM^{+}07]. Sardinia [Lóp06]. Sardinops [CR04, GWF01, War01a]. Sargasso [FMK07]. Sargassum [GC02]. sargus [LSM07]. Sarpa [JDA^{+}06]. Satellite [MSIL09, MTJ^{+}07, ORVP09, PH05]. satellite-based [MTJ^{+}07]. Satellite-measured [MSIL09]. Saudi [YBF^{+}03]. saury [STA^{+}09]. Saville [AJ00]. saxatilis [Gro06, RRC03]. scabbard [MD01]. scabbardfish [FBMR^{+}03, MN02, QGdS04]. scale [BHMS02, DDR^{+}03, EDG03, GP00, GFP09, HR09, JRM^{+}03, KMHS04, KWL^{+}02, KNO00, KRYL09, LD03a, LND05, LSH^{+}09, LdSSG02, NKOK00, RD03, SB04]. scalefish [MS00]. scales [Bot01, GLDD00, RL08, SL01, VEP^{+}09]. Scallop [HSM00, AHS08, BSMB03, GRE06, HC09, HSA05, JTE^{+}07, KCCM03, MFB^{+}08, MSIL09, RASS09, TCS^{+}09]. scallop-farming [MSIL09]. scallops [HS06, SBT^{+}09, VBSB07]. scanner [PF08a]. scat [MKR06]. scattered [Jaf06, Jaf08]. scatterers [OdSBS09]. scattering [BBA03, BB09, CT07, CWC^{+}03, CD09, CDB05, CD07, CRC^{+}09, FGR04, Miy03, NHK09, OMTS03, RHD09, SC00, WSWS03]. scattering-layer [BB09]. scavengers [GF00]. scavenging [BS02]. Scenario [PST^{+}07, LD03b]. scenarios [Kat05, LO05]. schlegeli [KH03a]. School [DH08, BRP02, BPT09, Fer09, HSR01, HS01, JNF^{+}09, KHO06, MBPW06, Mui03, NIF^{+}09, NFM^{+}02, RM01, TGS09, WPJ09]. School-based [DH08]. Schooling [SDWQ09, KHEJ09, TMI^{+}04, ZMM^{+}07]. schools [BPWS09, CR04, CDM03, FLK^{+}09, GCS^{+}04, GBBG06, GHI^{+}04, GKO07, HM04a, JPO09, KHE^{+}09, LBF01, MCI03, Pet01, Pet03]. Sciaenidae [dCA03]. Science [Daa03, HW06, HHAB09, WBK09a, FPS06, NM09, RP07, She07, Pay04, PB05a, PK07, RDF^{+}03]. Scientific [SFH^{+}07, Bac08, Hoy07, JFCH05, Knu09, MHD02, OMA09, PR07, Peñ08, RPB^{+}08, RUCG07, SMI07]. scientific-echosounder [Knu09]. Scientists [Ano01h, Gre08, NM09, Pet04]. scolopax [KTM^{+}05]. Scomber [Bar05, DDM^{+}05, Mar07, PAA06, TCSW06, TCTC09]. Scomberomorus [HAvH06]. scombrus [Bar05, DDM^{+}05, Mar07, PAA06, TCSW06]. Scophthalmus [ERGT07, LTA00]. Scotia [BMM03, TST^{+}09]. Scotian [BFZ05, CSH00, DLT^{+}00, GF01, OSK^{+}05, Zwa00]. Scotland [BM01a, BM01b, BML^{+}05, GHD^{+}09, GAM^{+}06, HWF08, SK07, TAC00]. Scottish [BPM^{+}09, CAWD09, CDB09, DAH^{+}08, FJK^{+}07, HSM09, LDML08, LDM08, SNA01, TES^{+}05, YMF02]. scrap [CJM^{+}02]. screening [Fra06]. scriba [Aló08]. Scyliorhinidae [ECC06]. SDWBA [CD06]. Sea [De 04, DHKV01, KCCM03, MFA07, PGB03, RLH01, SA03, WBV09, dPVV04, dPM08, AHS08, BR00, BF02, Box06, BHMS02, BHM^{+}04, CH09, DRRS01, GWSV08, GG09, HS06, HC09, Hea05c, Her04, HAvH06, HL09, HA03, HHC^{+}09, KH03a, Kos00, LHHJ^{+}09, LSM07, MMC03, MML09, MMS01, PPL^{+}07, PKP07, PJ08, RAKS06, Ros05a, RASS09, STG06, SW06b, SW06c, SBT^{+}09, SDG^{+}08, Str05, TAHK06, TCS^{+}09, UPK^{+}08, WBC^{+}08, WSC^{+}06, WW01, AP07, AS02a, ATM02, AE02, AMD^{+}05, APGD08, Arm01, AGA^{+}04, Bac08, Bai02, BFM00, BRP02, BvS00, BM01a, BM01b, BDJ^{+}05, BKR09, BWG^{+}07, BS02, Bri02, BWC00, Bro03, BVD01, BBK08, BBSK09, CBBL09, Cal02, CA00, CCA04, CCHV05, CRW^{+}01, CTW09, CFN03, CBHM07, CH05, Cor00b, Cot01, DGPR05, DBBM01, DM07a, DMvD07, DWDD03, DCPvK07]. Sea [DH04, Dol02, DMDE04, DBL07, ET07, FS02, FLP^{+}02, FGLT02, FB02b, FJK^{+}07, FBD^{+}08, Fjä05, FT05, FGFP08, FHHH00, FHDM00, FMK07, Fur02, GS03, GLDD00, GPRD08, GBT02, GDH02, GHI^{+}04, GFH04, GR06, Gre08, GFP09, GPZ^{+}05, GF00, HMK^{+}07, HP07, HFWB05, HSR01, HCE^{+}03, HSCN06, HM04a, Hea05a, Hea05b, Hea07, Hel00, HE08, HSA^{+}01, HF08a, HR01, HHH00, HWF08, HAvH06, HOD06, HMPC04, ISHB07, JDA^{+}06, JWM03, Joh02, JGST09, JLS02, JHL05, JMLG05, Kaa00, KASA07, Kas09, KPO05, KCR06, KA01, KF08, Kol06, KSD01, Kou00, KMJH01, LLD^{+}05, LNLS09, LLC^{+}08, LG08, LHKGS00, LND05, LHR02, LT06, LC09b, LJM00, LLS00, LHJS02, LBN09, MvdKN05, MLG^{+}09, MAB^{+}07, Mam06, MKR^{+}09, MNHL01, MUK^{+}02, MMD00, MS09, MAAN09, Mol00, MKFK05]. Sea [MDM03, MM05, MSIL09, NRR^{+}09, NWH02, NFM^{+}02, OK05, OGR^{+}07, OBNU02, OUNB02, OL00, ÖFR^{+}06, PRvB00, PMB00, PHDC^{+}09, Ped05, PPTS09, PVLP04, PB05b, PLP^{+}07, PPB03, PMN01, Pie00, Pie02, PR04, PvHG09, PKH^{+}08, PAC02, PRF^{+}00, PS06, RBD^{+}07, RP07, RK04, RNK05, RRT00, RvMBV00, RW01, RPT02, RO05, RDB09, RPK^{+}03, RFT02, RK00, SDÖ09, SPGT00, SBC^{+}00, SDCR07, SDWQ09, SB00a, SKR^{+}06, SSKE06, Sim03, Sim07a, Sim09, SAPP04, SFØ07, SB00b, SB01, SS00, SSJL02, SLN02a, SLN02b, Tan00, Ter02, Tje02, TPT^{+}09, UP02, Vin01, Wal07, WU03, WFIM00, YTS^{+}06, YS02, Ynd03, YYY^{+}02, YCCH07, Zha06, vDEM^{+}00, vDBF^{+}09]. sea-change [GG09]. sea-lice [DRRS01]. sea-ranched [SW06b]. sea-surface [MMS01]. seabass [ACD^{+}03]. Seabed [MSGC^{+}09, AVK^{+}08, BML^{+}05, DH09, DGO^{+}09, EMA^{+}07, EGB02, FSQ^{+}03, KTH^{+}00, KCD^{+}03, LKK^{+}09, MCAS04, MM07, PR01, RM01, TJG^{+}09]. seabed-mapping [KCD^{+}03]. seabed-mounted [DH09, DGO^{+}09]. Seabird [BCAN^{+}06, WGLJM04, AFM^{+}09, PCRW04, RNWS08, VLJM^{+}07]. Seabirds [PMB^{+}08, BANGC02, BCAN^{+}07, CMDN02, Fur02, GS03, MAMO02, OL07, UC05]. SEAFACTS [DKMO09]. seafloor [CLFS02, CMN^{+}07]. seal [BHMS02, CMHN05, Fjä05, FWW06, HSCN06, JSMK06, LS04, Lun01, LFW03]. seal-induced [LS04]. seal-inflicted [Fjä05]. seals [ATH^{+}07, BBBF02, BMM03, BHM^{+}04, CGN^{+}04, FPKH03, Fur02, GHD^{+}09, HS07, KKF^{+}06, LLHK07, LD03a, MKR06, PGB03, SFØ07, Ste02, TCC08]. sealworm [MMM00]. search [Fik00, VHF^{+}04]. seas [Cad00, dLMACC00, PF08b, SBR07, BR04, BD02, HM05, TH05]. seascape [PGD09]. season [CAAJ07, Cor00b, PM06, SNA01]. Seasonal [BdMAL00, FJSJBS^{+}08, Gaa00, KV06, KMT08, LLC^{+}08, LC09b, MB01, MR05, NH09, ÖFR^{+}06, ÖTTM07, RK04, RAKS06, SAM09, TMG^{+}08, AJR00, BAO04, CCC02, FB02a, GC02, GHC09, HAN02, LFD^{+}09, LME05, MML09, MAAN09, MS04, MSIL09, PM04, PPH09, RRY08, TF02, YM00]. seasonality [HSM09]. seatrout [Kup04]. seawater [DCM03, NNT01, Rho08, SDG^{+}08]. seaweeds [MPN^{+}08]. SeaWiFS [VM09]. sebae [GHBR08]. Sebastes [DNP03, DLT^{+}00, GR01, GR02, JGN04, KH03a, LC06, PGD09, Ped05, SRN00, SRGC04, SH07, SKR^{+}06, SKC^{+}00, SKH02, SRS^{+}09, SGS^{+}05, Str05]. Sebastidae [LC06]. secondary [MWS04, SMK08]. sections [Pel02]. sedentary [LPM^{+}09]. Sediment [BO08, NRS09, CAWD09, LVHU00, SLvdB^{+}09]. Sediment-bound [BO08]. Sedimentation [HKI01]. sediments [CFR^{+}01, DWDD03, HH01, MML^{+}00]. seeded [MCRF06]. seedlings [HRB02]. seen [HU04]. Seep [SQN08]. Segmentation [DAAD09]. segregation [LHHJ^{+}09]. seine [GDL04, LCRS08, LO05, D00, SSI07, NCC^{+}07]. seiners [BBGA05]. seining [HA03, SM02]. seismic [HKD^{+}04, LKK^{+}09]. selected [Bai09, NM09].

Selecting [FGP09, KWL^{+}02, MSGC^{+}09, RR05a, RR05b]. selection [BH07, CRIP08, JWBP07, KMJH01, Law00, MAC^{+}07, ÖFR^{+}06, PCM09, SDCR07]. selective [BPD^{+}03, RS06a, SCLG00]. selectivities [ZFFT01]. Selectivity [GM06, AP09, BSS07, BM09, GFKM07, HLS00, MSW07, MS01b, OH07, ÖTTM07, Tal07]. self [LFD^{+}09]. self-sampling [LFD^{+}09]. semi [Cad00, MR09, dLMACC00, Orl03, SP03a, SSJL02]. semi-cold [SSJL02]. semi-demersal [MR09]. semi-enclosed [Cad00, dLMACC00, SP03a]. semi-tomography [Orl03]. semisulcatus [NAK^{+}08, YBF^{+}03]. Sendai [MNY^{+}09]. sense [Roc00b]. sensed [PH05]. sensing [AGY^{+}05, Bro02a]. sensitive [KGRW07]. Sensitivity [BH07, Hor08, LTI09, HR09, MH01]. sensory [Hus04]. Sentinels [SHSKR01]. separable [PKP07]. separated [PK09]. Separating [FJK^{+}07]. separation [KHM09, SGMN^{+}06]. separators [MFIO04]. Sepia [CDR05, KCBC00, MLS07, WPB^{+}03]. Sepiidae [MLS07]. Sepioteuthis [JP03, MS04, PM06, TA05]. September [DPW07, MPD^{+}08]. sequence [BABB08, MYAT09]. sequences [KHS^{+}08, TYH04]. sequential [CF05, KKC04, RUA07, RO05]. series [CMP07, EIS05, FN00, FN02, aFADN08, Gud04, HHC^{+}09, KDP09, LLS00, SL04, VM09, WWR^{+}08, dPM08]. Serranus [Aló08]. serrator [Bun01]. Sertularia [WBV09]. sessile [KTS02, PPHB00]. Session [DPW07, MPD^{+}08]. Seston [UR01]. set [CBDB02, Fjä05, TMI^{+}04]. set-net [TMI^{+}04]. set-trap [Fjä05]. sets [WYM09]. Setting [PPC^{+}03]. settled [JWM03, MSR03]. settlement [IWP00, TCS^{+}09]. Seven [BF04]. Seventh [Jen02b]. several [PO09]. Severn [AWW^{+}07]. sex [CM00, FQS01, Iri00, LHHJ^{+}09, SNA01]. Sexual [LHHJ^{+}09, Ung07, Bro03, CLM07, KTRG06, SAM09, SMK08, TAHK06, TAC00]. Seychelles [GHBR08, HR09]. shad [AAV^{+}04, AVJ^{+}06]. shadowing [ZO03]. shallow [BI08, BGG^{+}06, DWDD03, KMHS04, SPGT00, SP07b]. shanny [BPD^{+}03]. shape [BBK08, GWG06, JCM06, Str05, WWWB03]. shaped [OMTS03]. Shark [MCB09, CLM07, Joy02, LME02, LME05, MSW07, MSH07, MNGB07]. sharks [DBS06, FCM09, HOGH07, SBDW00]. sharpnose [MNGB07]. shedding [DBS06]. Shelf [JV05, MFB^{+}08, NC08, NGNB^{+}04, BKN^{+}07, BvS00, CRvCB08, CMN^{+}07, CMM03, FSQ^{+}03, GSdFB01, GQCÁMI03, HHB^{+}00, Hea05c, JI05, LBF01, LD05, MM03b, MM03a, MS00, OR09, OLS00, PP08, PPHB00, PF06, PF08b, PB05c, SdlRdA06a, SdlRdA06b, TLMO08, Uye00, WCP08, BFZ05, CSH00, DLT^{+}00, Gaa00, GH00, GF01, HEGL05, MGH08, OSK^{+}05, SRM00, SG05, Zwa00]. Shell [HS06, MDM03, RASS09]. shellfish [HBW^{+}09, LBL06]. Shepherd [RRC03]. Shetland [RCLD08, Coo04, PCRW04]. shift [CvdLHF08, HFWB05, RO05, Wei05]. shifts [AMD^{+}05, Lit06, MMKR^{+}00, MMKKJ08, SN08, ZK00]. ships [DHWW08]. shirasu [Miy03]. shoal [SKC^{+}00, SKH02]. shoaling [GR05]. shooting [HKD^{+}04]. Short [RUA07, AJR00, BFM00, DCD00, Des00, EIS05, GAFA06, KPD^{+}07, MBPW06, PM06, RGG^{+}04]. short- [Des00]. short-finned [AJR00, DCD00]. short-lived [GAFA06, PM06]. Short-term [RUA07, BFM00, KPD^{+}07, MBPW06]. shortest [HP04]. shortfin [Hen04, MG07]. show [BK07, SCLG00]. shrimp [CRIP08, CDB05, EHL07, ELR01, Har07, HB07, HK06, KKC04, KWBR08, PCD05, SBG06, SW06a, Vor00, WCP08, Wie05, Ye00, YCCH07]. shrimps [HB07]. Sicily [BCD^{+}02, AJR00, SBD^{+}09]. Side [BW08, BRC09, BFMJ03, BNF^{+}07, dVA07]. Side-aspect [BW08, BRC09, BNF^{+}07]. side-looking [BFMJ03]. sidescan [BML^{+}05, HNLR04, MM07, SBP07, YM08]. sidescan-sonar [SBP07]. signal [BFMJ03, DH07, ES09, KRM05, PCM09]. signal-to-noise [DH07, KRM05]. signals [BKR09]. significant [GML06]. Silent [DHWW08]. Sillago [BBB06b]. Silver [WLS07b, WJB07, BVB^{+}07, CSH00, JWBP07]. similar [TPRR04]. similarity [LKK^{+}09, Mil08]. simple [FCM09, Gef09, GML06, Mac09, MMKKJ08, PRD^{+}06, SMH09]. simplex [HP01, PH03, PN06]. Simulated [Dav07, BR08a, CDB09, CMHN05, GDL04, MMF09, UBP^{+}09]. Simulating [BRC09, VHF^{+}04, WWHB04]. Simulation [RR09, HU04, HHMM01, KBDC^{+}08, LKL08, PPKM07, PBH02, Ree03, SHAH09, UASN07, VEP^{+}09]. Simulation-based [RR09]. simulations [BEB^{+}09, DCPvK07, Mol00, Wal07, WRF09]. Simultaneous [GLDB04a, GLDB04b]. since [Bra07, FGBS00, ML08]. Single [BAB^{+}04, Bet04, BPWS09, Cha04, GML06, GLKPCP01, GRMR07, Hol00b, HAvH06, HSA05, JO02, MS09, MMF09, MCL03, VRP04, WCMK05]. single- [Cha04]. single-beam [BPWS09, HSA05, MS09, MCL03]. single-fish [Bet04]. single-point [GLKPCP01]. single-species [Hol00b, VRP04, WCMK05]. Single-target [BAB^{+}04, JO02]. sinicus [Uye00]. siphonophores [SAAFCA07, War01b]. Site [BBMS01, BLRC05, FSDC09, GG09, WWR^{+}08, WSFH02]. sites [KWL^{+}02, PFK^{+}09, SMEK01]. situ [AGY^{+}05, BW08, CH00, CDSC05, CB07, DH04, EZ03, Erm09, GR02, GW09, HHC^{+}09, JH01b, JB00, KH03a, KCL^{+}09, KK06b, NTJ04, OR01, OR09, PHO09, Peñ08, RPE^{+}03, STG06, TK01, War01b, XZW05, Zha06, ZWD08]. situation [Ard08]. situations [Mac09, PPC^{+}03]. six [Sim07b]. Sixty [RDF^{+}03]. Size [JPO09, OV04, SSKE06, SSI07, APGD08, AGY^{+}05, BGL08, BCT05, BRP02, BB09, Bia00, BDJ^{+}05, BVB^{+}07, Bro02b, BFMJ03, CF06, EN02, FL06, FT05, FGFP08, aFADN08, GTOJA06, Gas02, GKFM09, HB07, HH01, HDG02, Hor08, IA04, JDA^{+}06, Jaf06, Jaf08, JD05, JWM03, Jør03, KS08, KNKT06, LHHF03, LFD^{+}09, LC09a, MB01, MGTS00, NGNB^{+}04, OUNB02, PGG05, PM06, PF06, PRD^{+}06, RS06a, Rob08, SJKN^{+}04, SRJ^{+}05, SP03a, SHdLP04, SIT^{+}05, STJ^{+}07, TAC00, WSC^{+}06, Wie05, ZFFT01, dLMS06, lPV01]. size-at-age [PGG05]. size-based [BDJ^{+}05, JD05, PRD^{+}06, SRJ^{+}05]. Size-dependent [JPO09]. size-fractionated [IA04]. size-related [FGFP08, GKFM09]. size-selective [RS06a]. size-sorting [HB07]. sizes [Aco02, aFADN08, GAYR06, RBGJ08]. Skagerrak [CCHV05, FGBS00, LNLS09, LSGD02, SB00b, Sve03, SB03, UE01, Ung07, UMSA09]. skate [FMF02, LCC07, MG07, RLdAW06, SdlRdA06a, SdlRdA06b]. skates [ABB^{+}08, ECC08, MG07, MC09]. sketch [Sin09]. skewed [HOP09]. skipjack [AK04]. skipped [RML06]. slag [CTF02]. sledge [HNLR04]. slime [DPW07]. slipping [SM02]. slope [Col02, Kos00, MLNC01, MM03b, MM03a, SNV^{+}09, SB01]. slope-water [MLNC01]. Small [Cur00, KMHS04, LND05, AVJ^{+}06, AF06, Dup05, GPP09, HB07, MANT07, D00, RD03]. Small-scale [KMHS04, LND05, RD03]. smallnose [OV04]. smallthorn [MG07]. smaris [ÖTTM07]. smelt [RPE^{+}03]. Smith [Ros03]. smolt [FHDM00, FRC03, FCM05, UKR05]. smolts [HHHH06, HHH00, SJKN^{+}04]. smoothers [FN00]. smoothhound [FCM09]. smoothing [BVD01]. snail [TDE08]. Snake [CBDS08]. snapper [EB04, GHBR08, SBG06, WM01, WSFH02, ZCH06]. snapshot [PGD09]. snow [WW07]. social [GG04]. socialis [Nau02, Vel02]. socio [CA02, RG07]. socio-economic [CA02, RG07]. sockeye [CMK09]. soft [CVL^{+}09, DWDD03, FLP^{+}02, GC07, LVHU00, MMD00]. soft-bottom [DWDD03, FLP^{+}02, MMD00]. soft-sediment [LVHU00]. software [UASN07]. sole [ACD^{+}03, Arm01, Bro03, ERGT07, HBS^{+}06, HF08a, LBL06, Pie02, PKH^{+}08]. Solea [Arm01, Bro03, ERGT07, LBL06, PKH^{+}08]. solida [JGM^{+}08]. Solutions [HW06, HB09, LTA00, MPD^{+}08, Pet04]. Some [SBB^{+}05, AE02, GPP09, KPK^{+}05, KCR06, MUK^{+}02, Mar08, MNMG^{+}05, OLS00, Pet04, But07]. SONar [HW08, BML^{+}05, CRC^{+}09, GBBG06, HCEM06, HNLR04, KHE^{+}09, MM07, MCL03, MCP03, SBP07, TNF09, TZ03, TGS09, YM08]. sonars [GHI^{+}04]. Sonic [BDTW06]. sonically [WBD^{+}06]. sorting [HB07]. Sound [HMQ^{+}08, RHD09, Ros03, BBA03, CW05, CDB05, CRC^{+}09, DCM03, GC05b, Jaf06, Jaf08, MS02, OdSBS09, TSK03, Bro02b, HFWB05]. Sound-scattering [RHD09, BBA03]. sound-speed [CW05]. sounder [DLS01]. sounds [ŠCBD09]. source [CRW^{+}01, KMG^{+}07, MAMO02]. Sources [BGW05, KGT01, OSLO06, Sve03]. South [EHG06, HS06, HOGH07, RUCG07, dBMS09, Bro03, CMC^{+}06b, CEH03, DP03a, GAP^{+}00, MLS07, CvdLHF08, Col02, DCN^{+}04, DB04, DBS06, ERBP09, GWSV08, HR00, LBF01, LDCH^{+}09, LC09b, MSB04, MM01, MLS07, OSB06, OR09, PS09, PRB^{+}07, RCLD08, Rob05, TCC08, WLS07a, War01a]. South-Brittany [MSB04]. south-west [DP03a]. south-western [Bro03]. Southampton [MWS04]. southeast [Bun01, BRHG^{+}06, GAM^{+}06]. southeastern [KTS02, LCC07, MNGB07, OdSBS09, PSO^{+}04, SDÖ09, dCA03]. southern [BCT05, BHM^{+}04, CCB^{+}06, CCA04, CCC02, CDDM05, DWDD03, DBL07, DM06, ECC06, ECC08, ELR01, EH04, FDD^{+}05, GZS^{+}09, GCS^{+}04, GLDD00, GCM09, GG09, GF00, GZND02, HBD05, HS06, Hel02, HL09, HA03, LLC^{+}08, LC06, LHHF03, LD03a, LCC08, LO05, MB01, MS04, MG02, NRR^{+}09, OV05, Pie00, PN06, RK04, RR02b, RK00, SRM00, SG00, SPS00b, WW01, YSF09, DC05b, DDGJ02, SCJ00, Sin01, ŠCBD09, WWWB03]. southernmost [FM04]. Southwest [ABB^{+}08, SR03, DP03b, GA00, HM04a, JR01, MG07, SGMMGB09, AK04, RLdAW06]. southwestern [AKJ07, Agn08, JHL05, LSH^{+}09, LME02, TST^{+}09, MVMH04]. sp.1 [KHS^{+}08]. space [BCT05, BMP^{+}08, BRP02, BBC^{+}04, SGS02]. spacing [Gas02, PR03]. Spain [BCL03, BLMB06, STM^{+}08, dHET04, FTDVC^{+}08, IA04, SCCM06, SGMMGB09]. Spanish [DSV^{+}08, HAvH06, Mar07, MCI03]. Sparidae [WM01, JDA^{+}06]. sparse [PPC^{+}03]. sparse-data [PPC^{+}03]. Spartina [TDE08]. Spatial [AFM^{+}09, BRE^{+}08, BREB09, BBA03, BWC00, CR04, FB07, GSdFB01, HMDS09, HHH00, HAN02, LG08, MVMH04, MGTS00, MGS00, MMM00, MM02, MSR03, MS04, MHV09, NTSM07, NC08, PVLP04, RS03, SHS01, SN08, SB03, SNB^{+}02, Vel02, WPB^{+}03, WvdMF06, YBF^{+}03, BPM^{+}05, BDÑ04, BBGA05, BR00, Boo00, BHMS02, BH08, CCHV05, CRvCB08, ES03, FGFP08, GL00, GLDD00, GMKS06, GG04, HM04a, HSPM05, Hol03, HNLR04, HJBG04, JH01b, Joh02, KKC04, LBL06, LDCH^{+}09, LVHU00, LPA^{+}00, MAB^{+}07, ÓGS09, PPTS09, PBH02, PPH09, RD03, SMH09, SL01, SRM08, UMSA09, Wal07, WPR^{+}07, ZPRJ02]. Spatialized [FDD^{+}05, BPM^{+}05]. Spatially [Aco02, CMJ09, EHG06, HHMN01, HTA09, KDP09, MHD02, RD01]. spatially-explicit [RD01]. Spatio [BRP02, DLR02, LHJS02, OSB06, SRM08, Bar05, BR02, GMM^{+}08, Kup04, LND05]. Spatio-temporal [BRP02, DLR02, LHJS02, OSB06, SRM08, Bar05, BR02, GMM^{+}08, Kup04, LND05]. spatiotemporal [KM02]. spawned [TCS^{+}09]. spawner [BO05, Cha04, JGM^{+}08]. spawners [BMV05]. Spawning [SGMV^{+}08, Sec00b, Arm01, BR02, BHØ^{+}04, Bri02, BD03, DEMD00, DDGR07, EDG03, EGO^{+}07, FBD^{+}08, FSDC09, FR09, GWG06, GCM09, GOA^{+}09, GOS07, HMDS09, Hen04, HHMM01, HBST02, HSS07, HSS^{+}09, JP03, JNF^{+}09, JCM06, JMC07, JHL05, Kat05, LMC^{+}01, LAO^{+}07, MGTS00, MJB08, MS04, MRT01, NT02, NW02, NCM^{+}03, O’D04, OR09, OB05, ÓT06, ÓGS09, OAJ06, PMM^{+}09, PMB00, PCM09, Ped05, RKP03, RLH01, RML06, Rob05, RR02a, RR02b, Roc00b, ROB04, RT03, RMM05, STAN02, SKS^{+}00, SAN^{+}05, SMK08, SPK05, SN08, TL05, TR09, TK03, War01a, WPM^{+}09, WJTH00, Ye00, YW05, vDBF^{+}09]. spawning-per-recruit [Kat05]. spawning-site [FSDC09]. spawning-stock [DDGR07, FBD^{+}08, vDBF^{+}09]. Special [MPJ07, MMC03, SDCR07, vdVBMR00]. Species [HW06, LBF01, LSGD02, LW04, MYAT09, AE02, APGD08, ASC01, BSA09, BCT05, BREB09, BAO04, BF04, BFMJ03, CTM09, CCHV05, CW09b, CF06, DAAD09, DB08, FLK^{+}09, Fer09, FJK^{+}07, FB03, FGFP08, GAW^{+}08, GAFA06, Gro06, HSdLP06, HHJK06, HMHI09, Hol00b, JHC09, JYW09, KHEJ09, LMM^{+}08, LC09a, LPM^{+}09, LN08, ML08, MKB01, MHV09, MB06, NWH02, O’D03, PM04, PMB^{+}03a, PvHG09, PMD^{+}00, Pow00, PPW^{+}09, RRT00, RL08, RMM05, TF02, TA05, TLMO08, TDE08, VLJM^{+}07, VRP04, WCMK05, vdVBMR00]. species- [FGFP08]. species-based [JYW09]. species-specific [GAW^{+}08, JHC09]. specific [Coo04, DK00, GAW^{+}08, HBST02, JHC09, SOMT00]. spectra [LN08, NHK09, NGNB^{+}04, RHD09, SIT^{+}05, SBD^{+}09]. spectral [DCRB09]. spectrum [GBC^{+}05]. speed [AE02, CW05, KNS^{+}04]. speeds [BDO^{+}04, HMD^{+}08]. sperm [EH04]. Spheniscus [CUUD07]. spheroid [TNF09]. Spicare [ÖTTM07]. spider [CF06]. spill [PBLFR06, PFLFR08, TT08]. spills [CSW06]. spinal [KTRG06]. spined [PVLP04]. spinicauda [SB01]. spinulosa [Vor00]. spiny [TM09]. Spisula [JGM^{+}08]. Spitsbergen [LLHK07]. split [AF06, GJH^{+}09, HPB09, KRM05, TK01]. split-beam [AF06, GJH^{+}09, HPB09, KRM05, TK01]. splitters [KDCH^{+}09]. SPMs [GLKPCP01]. sponge [KCCM03]. sport [DOBT02]. spotted [Kup04]. spp [CVL^{+}09, DBBM01, DLT^{+}00, FCM09, GR01, GR02, Iri00, LDNS08, Mam06, PCD05]. SPR [Kat05]. sprat [CCA04, DH04, GFH04, Kas09, PVLP04, VSS07, Vuo02]. Sprattus [CCA04, GFH04, PVLP04, Vuo02]. spread [BP08, Den08]. spreading [MS04, RRT00]. Spring [JHL05, LLHK07, SR03, BHØ^{+}04, BCL03, CZC07, DEMD00, DLC03, EDG03, HHB^{+}00, HBST02, HSS07, HSS^{+}09, LAO^{+}07, MGH08, MVMH04, PMB00, PCM09, PMB^{+}03b, ROB04, RT03, STAN02, Tan00, TL05, TR09, TK03, WPM^{+}09, dPVJM04]. Spring-spawning [JHL05, BHØ^{+}04, DEMD00, EDG03, HBST02, HSS07, HSS^{+}09, LAO^{+}07, PMB00, PCM09, ROB04, RT03, STAN02, TL05, TR09, WPM^{+}09]. springtime [Ber04]. spurdog [BGW05, EK08]. Squalus [BGW05, EK08, MF07, TM09]. Square [BM09, GM06, ZFFT01]. Square-mesh [BM09, GM06]. Squatina [CLM07]. squid [AJR00, CZC07, DCD00, DLR02, Hen04, HL09, JAC00, JP03, KMI^{+}05, OH00, OSB06, OR09, OLS00, PGG05, PM06, Rob05, SAM09, TJAS04]. squids [RPR02]. St [CCC02, Dup05, DR08b, LMC^{+}01, SR03, Sin01, SPS00b, SFM01]. St. [HSA05]. stability [JDN01, JCM06, JMLG05, KPS^{+}05, PR07, SKC09]. stable [BCL03, BAO04, HOF04]. staff [SKC^{+}00]. Stage [SOMT00, CWC^{+}03, CM00, DK00, GAW^{+}08, IFUR08, IWP00, Ric09, RDB09, WHP08]. Stage-specific [SOMT00, DK00]. stages [Bar05, BR00, Cos09, GAYR06, Hea07, Hel00, SHS01]. stained [Pel02]. stakeholder [SW02, SCHR07, VBF09]. stakeholder-influenced [VBF09]. stakeholders [PPKM07]. standard [DDM^{+}05, HS01, RPSSW09]. Standardization [RMAO^{+}03]. Standardizing [BVDS08]. Star [DBC03]. starvation [Nie00]. state [CDBS08, GD05, RMDB05, RSNB^{+}08]. static [RAR^{+}07]. Station [SL04, BF04]. stationarity [CLR^{+}05]. stationary [ADDH04, EIS05, FS02]. Statistical [BR08b, MS09, CA00, CFMdP07, DCRB09, HJB^{+}08, IB00, MHH06, MLM05, PH03, ZPRJ02]. statistical-spectral [DCRB09]. statistics [RGG^{+}04, SW06c]. Status [Jel07, ASB05, BWK07, BGG^{+}08, Bjö02, DH08, Fru02, HL07, LZS09, MF07, MSF^{+}06, MMKR^{+}00, MM05, PKP07, PPW^{+}09, RAR^{+}07, RS06b, SBL07, VSC06, XZW05]. steel [CTF02]. steel-slag [CTF02]. steelhead [BBMS01, RHD09]. Steeper [SBD^{+}09]. Stenobrachius [YTS^{+}06]. step [VRP04]. sterile [Ben01]. Sterna [MPG^{+}09]. stewardship [GH07]. stickleback [PVLP04]. stiffness [MANT07]. stochastic [Cor01, DC03b, DC04b, LN03, MSH07, RUA07, WDRP09, dHET04]. Stock [JGN04, KCL04, Mye01, RPR02, SPWHR04, Tur04, AP07, AKJ07, Agn08, AGA^{+}04, ARMM09, BCT05, BBM^{+}02, BNBR05, BdP07, Boo00, Bra05, BBPW07, BGW03, Bri02, BP07, BK07, BES^{+}01, BBSK09, Cha04, CBHM07, Cla00, CMP07, CSC^{+}04, DDGR07, Dek00a, Dek00b, Dek04, DCPvK07, EKPT07, FB07, FBD^{+}08, Fox01, GML06, GCM09, GHBR08, GAM^{+}06, GRMR07, HO01, HT05, HIL00, HAvH06, HK06, JMC07, KMM07, KMNP01, KDP09, LZS09, LN03, MKB01, MS07, MFA07, MLOT09, Mur00b, NSP06, ÓMP^{+}04, PS09, PRvB00, PGMB09, PLJ01, PCS^{+}04, Rad03, Ree03, RUA07, ROB04, RD07, RO02, SP07a, SZ07, Sim07a, SK07, SP03a, SKC^{+}00, SGMN^{+}06, TJAS04, Wie05, Ye00, YMF02, vDBF^{+}09]. stock-abundance [MKB01]. stock-recruit [BP07]. stock-recruitment [Fox01]. stocking [AMGV06, BÓ06]. stocks [ADC^{+}08, BWK07, Bra07, BBK08, CBHM07, Dem01, DNLSM08, ERGT07, GD05, Jel07, KPK^{+}05, KPS^{+}05, KPK^{+}06, MSF^{+}06, MSH07, MC07, MB05, Mui03, ON09, PKP07, Pet01, PR04, PPC^{+}03, Rob08, RDB09, RPK^{+}03, Rot00, RCL05, SP05, SBL07, SFM01, Vin01, WDRP09, HW06]. stomach [ATM02, BVD01, CMHN05, GMM^{+}08, Joy02, TM02]. stone [Col02]. storage [DSG05, HRM04, MCM00, Ste01a, WM04, vdKRS^{+}07]. story [CBBL09]. strains [KTRG06]. Strait [O’D04, PSO^{+}04, AJR00, CDDM05, Her04, RFM^{+}02, SHT^{+}09, SFM01]. Strategies [BCT05, RC07, APD09, CRTS04, FAL^{+}08, Fik00, HLCG04, KPK^{+}06, KMG^{+}07, MSP09, PST^{+}07, RR07, Roc00a, SPS00a, SH06, SMI07, Ska07, UASN07]. strategy [Bro03, But01, DB08, GFKM07, GRMR07, Her04, HM04b, Ray07, RL08, RR09, RUCG07, TB02, VEP^{+}09, lPV01]. stratified [MSI07, Ost09, SSKE06, SOMT00]. Stream [OCWV06, BMV05, SCLG00]. streams [CW06, LMC^{+}01]. Strength [GR01, AF06, BJ00c, BW08, CD06, CB07, DC03b, DC04b, DC05b, DH04, DM06, EZ03, Erm09, FBF09, FG09, Fra06, GR02, GOS07, GW09, HH03, HH04, HHT08, Hor03, HSS^{+}09, JO02, Jør03, JB00, KFM02, KH03a, KMI^{+}05, KCL^{+}09, Kas09, KK06b, KH03b, KTS02, MMC03, MW03, MLM05, OR01, OKG^{+}09, Ona03, OL00, PHO09, PB05b, PF08a, Peñ08, Ros09, RPE^{+}03, RKM09, STA^{+}09, TNF09, TCSW06, TSK03, WS06, YSO^{+}03, YTS^{+}06, Zha06, ZWD08]. strengths [CDSC05, DC03a, DC04a, War01b]. stress [CVL^{+}09]. striking [AKLL07]. Striostrea [dBMS09]. striped [Gro06, RRC03, Sec00a, Sec00b]. strong [Ber00]. strongly [Bra05]. structural [Cur00]. structure [AMGV06, BABB08, BBM^{+}02, Bla00, BDJ^{+}05, BAO04, BK07, Bro02b, BES^{+}01, CFRM08, Cal02, CR04, CAGV05, CDQL06, Cor07, DC05a, DCN^{+}04, DM04, FBD^{+}08, GL00, GCC^{+}09, GCS^{+}04, GLDD00, GMKS06, Hea05a, Hea05b, JD05, JGN04, Kou00, KB07, LBNS00, LVHU00, MM07, MASA06, MML^{+}00, MMKKJ08, NSP06, Orl03, PJ08, PB05c, RBBB00, Ric00a, RD03, SRM00, SGMV^{+}08, SQN08, SAPP04, SW06a, SB04, SRS^{+}09, SPWHR04, TJAS04, TYH04, ZPI^{+}09, ZCH06, dBP02]. structured [Cor01, MLOT09, dBMS09]. structures [GAYR06, HRB02, LKL08, MRV^{+}08, SG00]. structuring [CHB09, VCC07, WGMM08]. studied [LDCH^{+}09, TK01]. Studies [VHI^{+}04, AJNM07, BCAN^{+}07, CMM03, DH09, DGO^{+}09, ES09, FSQ^{+}03, GR02, GFKM07, HSPM05, HHKL04, Iri00, LDQ08, NSP06, NM08, Orl03, QGdS04, dR01]. Study [TM00, AFM^{+}09, BCD^{+}02, BDÑ04, BAO04, BVB^{+}07, BD07, CMC^{+}06a, CCB^{+}06, CRIP08, CBS^{+}06, CVG08, ES09, FFL06, GJL08, GLR06, GJR04, HMK^{+}07, HHSM03, HHMM01, HPBK04, HIL00, KBDC^{+}08, LCC09, LC09b, MCB09, MMB09, Mis02, MNY^{+}09, OED^{+}04, PPTS09, PKH^{+}08, PDRG04, PVH^{+}05, Ree03, RRC03, SA05, SNM05, SW02, SHAH09, TM09, WPB^{+}03, WHP08, YM08]. Studying [HJBG04, ES03]. sturgeon [FSDB09, GJL08, Kar06]. stylifera [DM04, LD05]. sub [VCC07]. sub-communities [VCC07]. subadult [dCA03]. subbifurcata [BPD^{+}03]. sublittoral [CHB09, CBS^{+}06]. submerged [CH06]. subpolar [Sar09]. subpopulation [VH08]. subsampling [WYM09]. subsequent [BD03, NEJH05]. subsidies [STW^{+}08]. substitution [Gat00]. substrate [FB02a, Mor02, RM01]. subtidal [Cor00a]. success [BFSC02, CBBL09, CH00, KM00, NT02, NK00, NWH02, Sim07a, WFIM00]. successful [HHMM01]. successfully [BJN^{+}06]. succession [BAO04]. successive [AJR00, PPW^{+}09]. sufficient [LZS09]. suggest [HAvH06, KHS^{+}08]. suggests [BJN^{+}06, VSC06]. suitability [JMLG05, MB06]. suite [CDBS08, RR05a]. sulphate [Cor00a]. summer [BGAM00, GHI^{+}04, GOS07, HL07, JP03, MAMO02, ÓT06, ÓGS09, PMB^{+}03b, SRM00, SR03, Tan00]. summer-spawning [GOS07, JP03, ÓT06, ÓGS09]. summers [JMWJ08]. sun [DK00]. superba [AF06, HTA09, HHKL04, RCLD08]. supply [HPBK04, OUNB02]. support [JNF^{+}09, NJ04, SFH^{+}07]. supporting [But01, JR07, NB08]. surface [BR00, BHMS02, BHM^{+}04, CH09, CT07, GA05, HHC^{+}09, LBN09, MMS01, SMB09, TJG^{+}09]. surfclams [Wei05]. surfgrass [HRB02]. surficial [CFR^{+}01]. surrogates [HHAB09, WBK^{+}09b, WBK09a]. surrounding [CFR^{+}01, DGMM02, FLP^{+}02, SML01]. Survey [AE02, RD01, ANNG01, BNBR05, BS03, BKR09, BGW03, But01, Cor07, DDM^{+}05, EIS05, FGD02, FGP07, FGFP08, FGP09, GOS07, HSR01, HS01, HOP09, HMAN03, HWF08, KKC04, KMNP01, LAO^{+}07, MSS^{+}05, NCM^{+}03, PPB03, PGMB09, RAB^{+}07, RGG^{+}04, Ros03, Sim03, SAN^{+}05, Som04, SLN02b, SPS00b, Syr00, ZWW^{+}03]. survey-based [BKR09, KMNP01, LAO^{+}07]. surveying [EHG06]. Surveys [Cot01, AHS08, AFGR09, BMJ08, BKR09, CS02, CD07, DBC03, Erm09, FLK^{+}09, GR05, HFWB05, HSCN06, Kas09, LCC09, LGR08, MvdKN05, MMCD08, NJ04, O’D04, OM05, Pet03, PMB^{+}03a, PMN01, PR03, RMDB05, RAR^{+}07, RUN07, RR02a, RUCG07, RPE^{+}03, RPSSW09, SGC^{+}09, TNF09, TPRR04, VC02, WMS^{+}03, WRF09]. survivability [Tal07]. Survival [BM01b, HSPM05, HRB02, HAG^{+}08, SJKN^{+}04, SNM05, SW06b, AG00, Bar05, BPD^{+}03, BJN^{+}06, DR08b, FHJS09, FHDM00, GG09, Han06, HPBK04, OSLO06, PPK^{+}06, PPH09, RRC03, TSH^{+}06, WBD^{+}06]. Surviving [GG09]. susceptibility [GSN^{+}03]. suspended [CFR^{+}01, SML01]. sustainability [AFHJ04, CRB08, JGM^{+}08, KPS^{+}05, KBDC^{+}08, Pen07, STW^{+}08]. Sustainable [RBD^{+}07, DDGR07, HMK^{+}07, Mac09]. Sustained [BMM03, BDO^{+}04]. Sv [GLDB04a, GLDB04b]. Svalbard [KKF^{+}06]. SW [DHKV01, DC05a, MFB^{+}08, OV04]. swarm [HHKL04]. Sweden [ETB07, LVHU00]. Swedish [ET07, SP07b, Sve03]. swept [RAB^{+}07]. swept-area [RAB^{+}07]. swimbladder [BJ00c, DM06, GPP09, GO03a, GO03b, GOK05, Jør03, NTJ04, PF08a, YSO^{+}03, vdKRS^{+}07]. Swimbladders [HSA^{+}09, Jaf06, Jaf08]. swimmer [UBP^{+}09]. Swimming [BDO^{+}04, DSJ03, HHT08, Hus04, KNS^{+}04, LBL06, TK01, WWHB04]. swordfish [CPR06]. symbols [MFD02]. symmetricus [PF08a, Peñ08, HA03]. sympatric [BF02, BFK^{+}07, JYW09, MG07]. Symposia [vdVBMR00]. Symposium [DKMO09, GSSO00, HLSW01, VPC^{+}09]. Sympterygia [OV04]. synchronies [PBM^{+}04]. synchronized [GJH^{+}09]. Synchronous [AMD^{+}05]. Syndinea [HSM09]. syndrome [Vuo02]. synoptic [GD05, PZTE05]. synthesis [DKMO09, FAL^{+}08, His01]. Synthetic [KO03]. System [BBBF02, YS02, BBS09a, BML^{+}05, CBDS08, EGB02, EJR01, GG08, GPWG04, HMMB^{+}08, HCEM06, HNLR04, KO02, LSH^{+}09, MTJ^{+}07, MYAT09, PCM01, PR01, RNWS08, RHBR04, RKM09, STA^{+}09, STG06, Sim07b, TZ03, TGS09, WSP03]. system-scale [LSH^{+}09]. Systematic [MBC^{+}09]. systems [Cal08, Cur00, DNP03, DM07b, GC07, Hol03, MLMC02, OMA09, WMS^{+}03].

T [WBK09a]. TACs [PR07]. tactics [Roc00a]. tag [DBS06, ES09, FSDC09, LHHF03, MMF09, SK04, Ste01a]. tag-recovery [FSDC09, LHHF03, MMF09]. tag-signal [ES09]. tagged [DBS06, LDQ08, WBD^{+}06]. Tagging [ACD^{+}03, BHR^{+}05, HJB^{+}08, STG06, UE01, dPBB^{+}03]. tags [BHR^{+}05, ES02, HRM04, WM04, vdKRS^{+}07]. tail [HPB09]. tail-beat [HPB09]. tailbeat [KNS^{+}04]. tailed [DC05a]. Taiwan [CTF02, LSH^{+}09, LPH^{+}08]. take [EB04, MBC^{+}09, PBH02]. Taking [FGP07, CLR^{+}05, VRP04]. tank [AF06, CD03]. tanks [TSH^{+}06]. Tanner [NTSM07]. Tapes [MCRF06]. Tapong [LSH^{+}09]. TaqMan [GAW^{+}08]. Tara [WBK09a]. Target [GR01, STA^{+}09, YSO^{+}03, YTS^{+}06, AE02, AF06, BJ00c, BW08, BAB^{+}04, CDSC05, CD06, CB07, DC03a, DC03b, DC04a, DC04b, DC05b, DH04, DM06, EZ03, Erm09, FBF09, FG09, GR02, GW09, HH03, HH04, HHT08, Hor03, JH01b, JO02, Jør03, JB00, KH03a, KMI^{+}05, KCL^{+}09, Kas09, KK06b, KH03b, KPD^{+}07, MMC03, MW03, OR01, Ona03, PHO09, PB05b, PF08a, Peñ08, PvHG09, PMD^{+}00, Ros09, RPE^{+}03, RKM09, TNF09, TCSW06, TSK03, TK01, War01b, Zha06, ZWD08]. Target-strength [STA^{+}09, BJ00c, BW08, DC05b, Erm09, HH04, KCL^{+}09, KH03b, Ona03, PHO09, Peñ08, RPE^{+}03, TCSW06, TSK03, Zha06]. Targeted [DRRS01]. targeting [LO05, MSB04]. targets [GC05b, PK09, WWWB03]. tarpon [ZAJ01]. Tasmania [JP03, LHHF03, PJ08]. taurus [DBS06, LME02]. taxa [MM05]. taxifolia [RRT00]. Taxonomic [BAO04, Vec00]. Technical [Esm06, CRIP08, SS07, dHET04]. technique [DH07, FS02, HDG02]. techniques [BIdL^{+}08, BG04, ES03, EIS05, KWL^{+}02, TVH08, WB02, ZPRJ02]. technological [BMP^{+}08, Eig09, GJR04, MAC^{+}07]. Technologies [DKMO09, KCD^{+}03]. technologists [JR07]. technology [Bro02a, GWvM07, JGST09, KGRW07, Kar06, TJG^{+}09, YCCH07]. telemetric [RR02a]. telemetry [ACD^{+}03, BVB^{+}07]. teleost [Roc00a]. Teleostei [JDA^{+}06, PSC02]. television [CDB09, MFA07]. tell [BHR^{+}05]. TEMAS [UASN07]. Temora [DM04, LD05]. temperate [EB04, FR04, HS09, MWF^{+}05, PSC02]. Temperature [HOF04, OFN02b, OFN02c, BR00, BSO01, CH09, CVL^{+}09, CVG08, DR08a, HHC^{+}09, HSS07, KMJH01, Kup04, LHHJ^{+}09, ODCN09, OR09, OBD^{+}05, OLB01, PF08b, PLJ01, PS06, Sar09, Ste01a, SPS00b, VSÅF05, Wei05, Wie05, YW05]. Temperature-dependent [HOF04, Kup04]. temperatures [BHMS02]. Temporal [BBSK09, CF02, CMO^{+}06, GKFM09, Joh02, NRR^{+}09, SSC^{+}06, Tri00, VM09, WPM^{+}09, BWK07, Bar05, BR02, BRP02, BWC00, DLR02, ES03, GMM^{+}08, GLDD00, GG04, HHH00, JDN01, JCM06, KM05, Kup04, LG08, LND05, LVHU00, LPA^{+}00, LHJS02, MMM00, MRV^{+}08, MWS04, NJ04, OSB06, PPTS09, SWG06, SRM08, SB03, TMG^{+}08, WPB^{+}03]. Ten [RRTP02, SMEK01]. Ten-year [SMEK01]. term [ACD^{+}03, BFM00, BHMD05, BGAM00, CH09, CTF02, CFN03, Des00, ERP01, FHHH00, HSM00, Knu09, KPD^{+}07, MBPW06, NEJH05, Pen07, PF06, PPH09, RL05, RF01, RUA07, WYMF08, Ynd03]. terminal [Hor08]. Terminos [SLMCRM05]. tern [MPG^{+}09]. terrain [BI08, wScY02]. Testing [GHD^{+}09, HSA^{+}01, JMLG05, LBL06, NJ04, SGM09, BR08a, DB08, MMF09, PST^{+}07, ZPRJ02]. tests [Ard08]. Texas [DOBT02]. Thailand [CVL^{+}09]. Thalassiosira [SMEK01]. Thames [ROB04]. their [AG00, BBR08, BDO^{+}04, CWC^{+}03, EJR01, FLK^{+}09, GAYR06, GG04, HSPM05, HNLR04, HSS07, Joh08, JJ06, PMM^{+}09, PQRG07, D00, RG07, RD01, SAM09, SPFF^{+}08, SB00a, Sin09, ŠCBD09, SRM08, SNB^{+}02, Tri00, WPR^{+}07]. them [RO02, SS07]. Theme [DPW07, MPD^{+}08]. Themisto [DBBM01]. theoretical [DC03a, DC04a, HHMN01, MR09, Miy03]. theory [CMGS05, GBC^{+}05, OSK^{+}05]. Theragra [BWC00, HH04, Hor03, KK06a, SBC^{+}00, WFIM00]. there [DPW07, HR01, Ste02]. Theregra [Som04]. thermal [FRC03]. thermocline [HR09, SAAFCA07]. thermohaline [LTI09]. thiamine [Vuo02]. Thin [CD09]. thirty [Rot00]. Thought [Kin02]. thousand [Hal01]. Three [Ard08, CDBS08, GCS^{+}04, SKH02, BdMAL00, CRC^{+}09, GLS^{+}03, GAYR06, GFKM07, GLDB04a, GLDB04b, HSdLP06, HMHI09, KTH^{+}00, LC09b, MUK^{+}02, MMB09, Mol00, MB05, NIF^{+}09, NWH02, PVLP04, Ric09, SS00, TNF09, VGBH09]. Three-dimensional [GCS^{+}04, SKH02, CRC^{+}09, LC09b, Mol00, NIF^{+}09, SS00, TNF09, VGBH09]. three-spined [PVLP04]. three-stage [Ric09]. three-zone [MMB09]. threshold [Bet04]. threshold-induced [Bet04]. throughout [PMM^{+}09]. Thunnus [GA05, GOA^{+}09, LMVdZ^{+}07, LCRS08, PVH^{+}05, RMAO^{+}03, SA05]. thynnus [GOA^{+}09, RMAO^{+}03]. tidal [LMC^{+}01, LND05, SCLG00]. tidal-mixing [LND05]. tidal-stream [SCLG00]. tiger [LME02, NAK^{+}08, YBF^{+}03]. Tilt [MW03, KMI^{+}05, STA^{+}09]. tilt-angle [STA^{+}09]. Time [BFMJ03, Gud04, JRM^{+}03, LLS00, MNHL01, SL04, AE02, BRP02, EIS05, FN00, FN02, GOA^{+}09, HHC^{+}09, HSS^{+}09, KHN03, MYAT09, OM05, OAJ06, RLH01, Sec00b, TM02, VM09, WWR^{+}08, WPR^{+}07, YW05, dPM08]. Time-based [BFMJ03]. Time-series [Gud04, SL04, EIS05, FN02, HHC^{+}09, VM09, WWR^{+}08, dPM08]. times [íJR02, MMF09]. times-at-large [MMF09]. Timing [CCC02, BDS01, Cor00b, Fik00, HHMM01, WJTH00]. Tips [RPB07]. tissue [JGN04, SRS^{+}07]. Todarodes [KMI^{+}05, SKS^{+}00]. tomography [Orl03]. tool [BR08a, FGD02, HPR09, MBC^{+}09, OH07]. tools [MHD02, MLM05, PCW00, RUCG07, SMI07, SFH^{+}07]. tooth [Gas02]. toothed [KMV^{+}07]. top [CHB09, SL01]. top-down [SL01]. topography [GMKS06]. torreyi [HRB02]. total [DC03b, DC04b, GV02, LAO^{+}07, dVA07]. towed [BDO^{+}04, DNP03]. trace [BGL08, LLD^{+}05]. Trachinus [Bag04]. Trachurus [BBC^{+}04, CVG08, DRDC06, DDM^{+}05, GMM^{+}08, GCM09, HA03, LMVdZ^{+}07, Mur00b, NH09, NHKJK09, PF08a, Peñ08, RD07, Tur04, WK01]. trachyderma [LCC07]. Tracing [BD02]. track [PR03]. tracking [AAV^{+}04, BDTW06, EB04, HW08, JO02, MW03, TK01]. trade [BMP^{+}08, CMK09]. traditional [BIdL^{+}08, Hol03, MPJ07]. traits [PKRT06]. trajectories [BBGA05, GGM^{+}05]. Trans [TM00]. Trans-Atlantic [TM00]. transducer [ADDH04, DR08a]. transect [BGW03, Pet03, PMB^{+}03b]. transects [SLvdB^{+}09, TLM04]. transfer [PPC^{+}03, WSC^{+}06]. transfers [Den08]. transient [PCD05]. transitional [STM^{+}08]. Translating [Lin05]. translocated [GG09]. transmit [GBBG06]. transmitters [ACD^{+}03]. transparency [HNK07]. transparent [HOHS05]. transplanted [RR02b, WC01]. transport [GOA^{+}09, HHMM01, HR01, Jag02, KKS^{+}07, MF07, SCLG00, SPWHR04, VSÅF05, WFIM00]. transportation [JGST09]. trap [Fjä05, FWW06, Gat00, LFW03, SHdLP04]. trapnet [LS04]. trapped [BC07]. traps [Eno01, WW07]. Trash [CLFS02, EHL07]. Trawl [AE02, Cot01, LD03b, BSS07, BFM00, BTR06, BS03, BvS00, Cor07, CMP07, DSV^{+}08, EHL07, ET07, Erm09, EIS05, FJK^{+}07, FBD^{+}08, FGD02, GR05, GAA^{+}04, GF00, GM06, HFWB05, HMAN03, HF08a, HLS00, ISHB07, KKC04, KHM09, LGR08, MvdKN05, MF07, MS01a, MSS^{+}05, MTJ^{+}07, MS01b, ORA02, ÖFR^{+}06, ÖTTM07, PMB^{+}03a, PMN01, RS03, RMDB05, RAB^{+}07, RKE06, RDHP00, RBGJ08, RK00, Sim03, SPD00, SLN02b, STJ^{+}07, SPS00b, Syr00, Tje02, TMG^{+}08, WLS07a, WPB^{+}03, WWHB04, ZWW^{+}03, ZFFT01]. trawl-acoustic [Erm09, Tje02]. trawl-survey [Sim03]. trawler [DLR02, RPT02, SBD^{+}09]. trawlers [BvKvH^{+}08, Dor01, HMPC04, MSB04, RvMBV00]. trawling [De 04, DW06, DTC01, HNLR04, LVHU00, MBPW06, MAMO02, MMD00, MW03, MAAN09, MS00, Pie00, PvHG09, PPHB00, PRF^{+}00, RS03, SBG06, SDRK00, SW06a, SPD00, SBP07, SJM03, WCP08, YCCH07]. trawlnet [RKM09]. trawls [FGP07, FGFP08, FGP09, GJR04, HW08, KWBR08, Rye04, SSI07, Som04]. treat [DRRS01]. treated [SW06b]. trees [Fer09]. trend [SL04]. Trends [CCHV05, CDR05, CDD^{+}07, FPKH03, APGC04, BWK07, BHMD05, BMJ08, BBSK09, CH09, DCPvK07, EN02, Fox01, Kan07, KDP09, NJ04, RTB^{+}05, RPB^{+}08, Sea02, TPRR04, TMG^{+}08, WU03, YMF02]. Trichiurus [Zha06]. tricks [RPB07]. Tridacna [AGY^{+}05, GAYR06]. trip [RDHP00, RDD06]. triploid [Ben01, BBMS01]. Trisopterus [LNLS09, PK09, SLN02a, SLN02b]. TRIX [STM^{+}08]. Trophic [CMDN02, RRTdA02, SdlRdA06a, SdlRdA06b, YM00, BAO04, Cal08, GL00, GBC^{+}05, GLDD00, Hea05c, HHMN01, HM05, HJBG04, MMKKJ08, STM^{+}08, SQN08, SLMCRM05, UP02]. Trophodynamic [KM00, NK00, YNX^{+}05]. tropical [GAFA06, JMWJ08, PPHB00, D00, RD01, SIT^{+}05, SJM03, dCA03, dBP02]. Trough [HR01, MD01]. trout [BF02, BFK^{+}07, BBMS01, GSN^{+}03, RS06a, RAKS06, SW06c, UPK^{+}08, WW01]. true [TLM04]. truncatus [Lóp06]. Trunk [AHS08]. truth [BHR^{+}05]. trutta [BF02, CMO^{+}06, GSN^{+}03, RAKS06, UPK^{+}08]. TS [GLDB04a, GLDB04b]. TSDV [FGD02]. tshawytscha [RHD09]. Tuamotu [AGY^{+}05]. Tuna [BJ00c, AK04, ARMM09, BJ00b, BJ00a, CMC^{+}06b, CLL^{+}09, DH08, DPN^{+}09, GDL04, GZS^{+}09, GOA^{+}09, GKFM09, Her04, JB00, LCRS08, RF01, RMAO^{+}03, Sec02]. tunas [D00]. tunicate [LMM^{+}08, TH08a, TH08b]. tuning [KDP09, Vin01]. tunny [BKN^{+}07]. turbot [ERGT07, LTA00]. turbulence [MH01, SAAFCA07]. Turkey [CAAJ07, LMU^{+}02]. Tursiops [Lóp06]. Tuscany [De 04]. Two [JNF^{+}09, VGBH09, AHS08, APD09, AFGR09, AJR00, CF02, EIS05, GOS07, HLCG04, Han06, HEGH02, IFUR08, JMWJ08, JFCH05, KTRG06, LHHF03, MG07, MSH07, NCC^{+}07, NM08, OMA09, RDB09, STM^{+}08, SK04, Sim07a, TCP05, VEP^{+}09, WSC^{+}06, vdMBD00]. Two- [VGBH09]. two-dimensional [TCP05]. two-stage [IFUR08, RDB09]. type [TST^{+}09]. types [GLS^{+}03]. typicus [DM04]. tyre [CJM^{+}02]. Tyrrhenian [De 04].

UK [BBR08, CRIP08, EMA^{+}07, PF08b, SRM08, WWR^{+}08]. Ulla [SCCM06]. Ultrasonic [EB04]. ultrasound [PPMH04]. ultraviolet [SBB^{+}05]. Ulvaria [BPD^{+}03]. uncertain [SK07]. uncertainties [CDB09]. uncertainty [BPT09, HO01, HBST02, Kas09, KH03b, LAO^{+}07, MHH06, O’D04, PD07, RCL05, Tje02, WRF09]. underages [PB08b]. underestimation [dPGPB06]. undergo [HSdLP06]. underrepresented [AGH^{+}09]. understand [CRB08, DCCS09, PPW^{+}09, SB06]. understanding [BDTW06, GG04, GLR06, MHV09]. Underwater [MFA07, CDB09, FSB^{+}03, GEM01, LKL08, PHG04, RLF01, STG06, SBP07, SYR^{+}08]. underwater-video [SBP07]. underway [PCS^{+}07b, PCS^{+}07a]. unexploited [MMCD08]. Unifying [Sea02]. Unintended [DM07b]. Union [Pen07]. unit [APGD08, GDL04, GA05, HBD05, KS08, Kal01, LZS09, MSF^{+}06, MRV^{+}08, PCM01, PPB03, RMAO^{+}03]. United [PCM01, WBC^{+}06]. units [CSR^{+}02, PPL^{+}07, SMH09]. univariate [FFL06]. unsampled [BGG^{+}06]. unstable [SK07]. unsupervised [CFMdP07]. untrawled [DTC01]. updated [RPR02]. upper [KA01, SR03, SYR^{+}08]. Upstream [AVJ^{+}06, AAV^{+}04]. upward [ADDH04]. upward-facing [ADDH04]. upwelling [BCL03, BAO04, CUUD07, Cur00, MJA^{+}05, SdFBG01, VLJM^{+}07, WYMF08]. Ural [GJL08]. urchin [SDG^{+}08]. urchins [PJ08]. USA [PP08, AHS08, AMJ^{+}06, BFSC02, CH06, CBDS08, Gro06, HM08, HS06, OLS00, RRC03, SFKC02, TSK00, WBD^{+}06]. Use [BR08a, Ben01, HRM04, NB08, RLF01, TCM^{+}08, dCA03, BPM^{+}05, BMJ08, Bro03, CRB08, CS02, DRDC06, DAH^{+}08, DNLSM08, GAW^{+}08, HJB^{+}08, HOP09, HL07, HNLR04, KFM02, KTS02, LLD^{+}05, MMC03, MVM^{+}08, PKH^{+}08, RUA07, SPFF^{+}08, ŠCBD09, SW06c, SDG^{+}08, Syr00, TM09, TB02, VLJM^{+}07, WCP08]. used [DRRS01, HSdLP06, JAC00, KPK^{+}05, RAR^{+}07, WB02]. useful [dLMACC00, RDD06, SIT^{+}05]. Using [ARMM09, BNBR05, BIdL^{+}08, BMU09, CGV03, FN00, GMGN06, GFP09, Jaf06, Jaf08, JGST09, MHD02, Mil02, MV09, OM05, Pie02, SMH09, SRJ^{+}05, YM08, AP09, AAV^{+}04, AK04, AGY^{+}05, BM02, BR02, Bea05, BPT09, BR00, Bri02, BD02, Bro02a, BML^{+}05, BMDBM09, CDB09, CBDB02, Cha04, CRB08, Cla00, Col02, CMP07, CRC^{+}09, Dav07, DLR02, DGO^{+}09, DNLSM08, DB08, EGB02, EDG03, ERGT07, FS02, FJK^{+}07, FSDC09, FMF02, aFADN08, GJL08, GLS^{+}03, GPWG04, GLDB04a, GLDB04b, HP07, HS01, HW08, HPB09, HC09, HKBK09, HMAN03, HJBG04, HK06, JRN06, JI05, KK06a, KDP09, LTA00, LBF01, LN03, LAB^{+}05, LW04, MM07, MGvH06, MMF09, MH01, Mye01, NNT01, PS09, PKP07, PF08a, PST^{+}07, PBH02, PH03, Ric00a, Ric09, RR02a, RRC03, RKM09, STA^{+}09, SHT^{+}09]. using [wScY02, SEOR09, SS00, SHSKR01, SGAC00, SVRF08, SFM01, TSK07, TVH08, TH08b, TJG^{+}09, TS05, Tur04, TLMO08, UKR05, VM07, WPB^{+}03, WSWS03, WWR^{+}08, WS02b, WRF09, WWWB03, WWGG02, YFL05, ZPRJ02, vDBF^{+}09, BHR^{+}05]. utility [HE08, Joh08, RHH^{+}08]. Utilization [GS03, DSG05, JDA^{+}06].

vaguely [DCPvK07]. Valdez [TT08]. Valencia [FTDVC^{+}08]. Valenciennes [SDÖ09]. Validated [KCR07]. Validating [PPKM07].

Validation [DC03b, DC04b, HLL^{+}08, SS00, dR01, CBHM07, DWC03, FGR04, KK06a, Mol00, WK01]. validations [Fra06]. Valli [Mis02]. value [HCV03, MKR^{+}09, NM08, Sim07a]. values [VH08]. Variability [AGA^{+}04, DP03a, DP03b, RASS09, TH05, AFM^{+}09, Bai09, Bar05, BWC00, CH09, CF02, CT07, CCC02, DC03a, DC04a, DHKV01, DPN^{+}09, ERP01, GP00, HH04, Hea05c, HWF08, KMH^{+}05, LMVdZ^{+}07, LND05, LPA^{+}00, LC09b, LdSSG02, LTI09, Mar07, MML09, MCM00, Mye01, NRR^{+}09, PGJ^{+}05, RL05, RCBM05, RK04, STAN02, SSC^{+}06, SSU^{+}09, SBR07, SA03, UP00, VH08, VM09, WFIM00, WHP08, ZPI^{+}09, dPVV04]. variable [BGG^{+}06, FB03, JRN06]. variables [GAZ02, GA05, HMMB^{+}08, JWM03, LBN09]. variance [PMB^{+}03a, RDB09]. Variant [GW04]. Variation [GAM^{+}06, JRN06, LBNS00, ÓGS09, RML06, SL01, BBM^{+}02, BS03, BA03, BWG^{+}07, BGW05, BHM^{+}04, Buc00, CFRM08, CS05, CMO^{+}06, FGFP08, GKFM09, HTSB04, HAN02, HK00, Joh02, KGT01, LFD^{+}09, MAB^{+}07, MGTS00, MGS00, MM01, MS04, MRT01, Mur00b, ON09, ÓT06, ÖFR^{+}06, ÖTTM07, PLP^{+}07, PCRW04, RS04, SRN00, SCCM06, Sil03, Ska07, STMM06, Str05, UMSA09, WW01, WvdMF06, YM00, YBF^{+}03, ZMM^{+}07, dLMS06]. Variations [DR08a, GH00, LD05, RCLD08, Ros09, GMM^{+}08, JG07, Kup04, LHJS02, MSB04, MFB^{+}08, MSIL09, Ped05, SMG02, Tan00, Tri00]. various [CRTS04, OSLO06]. vary [KDP09]. varying [BdMAL00]. vase [TH08a, TH08b]. vegetation [FB02b]. vehicle [GEM01, PHG04]. vehicles [FSB^{+}03]. velocity [ZCR09]. Venice [PDRG04]. venosa [SDÖ09]. venting [Aló08]. verifiable [CLL^{+}09]. Verification [SLN02b, WB02, WWGG02, Orl05, RKM09]. versus [BAO04, Cad00, DC03a, DC04a, HSA^{+}01, PCM01]. vertebral [SFM01]. Vertical [JWM03, OKRK04, PM04, SHAH09, SAMS02, VSS07, BBC^{+}04, DK00, GR05, HRM04, HL07, KHM09, Mow02, NH09, NCM^{+}03, PSHL09, PCS^{+}07b, PCS^{+}07a, PP08, Ros09, Sab04, SBC^{+}00, SSA08, ZMM^{+}07]. vessel [DHWW08, GHI^{+}04, HHO08, HF08b, MSB04, MTJ^{+}07, PCM01, Peñ08, SAN^{+}05, SSI07, UA04, lPV01]. vessels [CLL^{+}09, DW06, MLMC02, OM05, PO09, RDHP00, RKKM06]. vexillum [Den08]. via [BBS09a, BNBR05]. Viability [CH00, CMGS05, DDGR07, OLB01]. viable [JGM^{+}08, MGTS00]. vicinity [JLS02, LHJS02, SSJL02]. Victoria [SP05]. Video [CMM01, AHS08, Bro02a, Col02, EZ03, HNLR04, MM07, SBP07, SYR^{+}08]. video-acoustic [EZ03]. video-sledge [HNLR04]. Vietnamese [EHL07]. view [Ber04]. Viewpoint [Hoy07]. VIEW^{T M} [EGB02]. villosus [Nau02, Vel02, ADO02, CFL00, CMDN02, CF02, CRW^{+}01, CMHN05, DAd02, Dol02, DBDA^{+}02, FRK02, GBT02, GDH02, GW09, Mow02, NT02, NW02, OUNB02, Ros05a, Tje02, Vil02]. Vincent [SR03]. virens [AJNM07, NCM^{+}03, NSP06, PK09]. virtual [JMLG05, LJM00, TLMO08]. virus [Cip09]. viruses [Law08]. viscosity [Rho08]. vision [NM09]. visual [RKM09, TLM04]. Visualization [KM02, MM02, BM02, MLM02, MHD02, MLMC02, SKH02]. visualizing [KO02]. vitulina [BBBF02, Lun01]. void [SGS02]. Volume [Ano00a, Ano01c, APGC04, BB09, BJ00c, KFM02, NTJ04]. volumes [RDF^{+}03]. volumetric [NPPO06]. VPA [BR08b, Vin01]. vs [BR08b, Cos09, KPS^{+}05, LDNS08, LN08, MR09, PCRW04, RNK05]. vulgaris [OSB06, OR09, Rob05, CFRM08, FB07, GGP07, KV06].

Wadden [WBV09]. waist [BBS09a, Cur00]. Walbaum [BBMS01, JRM^{+}03, MCI03, RS06a]. Wales [HOGH07, RUCG07, BWK07, Dun01, PKP07]. walk [Niw07]. Walleye [DW06, APD09, BWC00, HH04, HHMM01, HIL00, Hor03, HSA^{+}09, KHO06, KK06a, LBNS00, SBC^{+}00, SDWQ09, Som04, Wal07, WFIM00, WS06]. walrus [WB05]. warm [OBNU02, OUNB02, Tan00]. warmer [Wei05]. warming [CH05]. Washington [TSK00, PP08]. wasp [BBS09a, Cur00]. wasp-waist [BBS09a, Cur00]. waste [MAMO02]. wastes [HBC01]. watches [GC07]. Water [BR04, FTDVC^{+}08, MWS04, SA03, BSA09, BHH^{+}08, BO08, BÓ06, BF04, BGG^{+}06, BD07, DR08a, DCM03, DB08, ERBP09, HHAB09, HR01, HF08b, Jel07, JV05, LGH^{+}09, MLNC01, MSGC^{+}09, NB08, NWH02, Ric08, Rob08, SAAFCA07, SPS00b, Tal07, WBK^{+}09b, WBK09a, ZCR09]. Water-mass [BR04, BF04]. water-quality [NB08]. waters [AP07, ABB^{+}08, BKN^{+}07, BANGC02, BA03, BBBF02, Buc00, CAWD09, CH09, CCB^{+}06, CTF02, CW05, DCN^{+}04, DOBT02, Dun01, EMA^{+}07, EH04, FTDVC^{+}08, FRK02, GW04, GKFM09, GHC09, His01, JMWJ08, íJR02, JR06, JDN01, JMC07, KMHS04, KHS^{+}08, KA01, KWZ00, KKC04, KMV^{+}07, LDML08, LDM08, LMVdZ^{+}07, LD05, MB01, MGTS00, MGS00, NW02, NAK^{+}08, PFK^{+}09, PM04, PJR08, PS03, RTDJ09, RS03, RE00, Sar09, SKR^{+}06, SEOR09, SPG^{+}04, SPWHR04, TCM^{+}08, TH08a, TH08b, TA05, VH08, Vel02, VGF03, WPB^{+}03, War01a, WPM^{+}09, WvdMF06, YBF^{+}03]. wave [BI08, BLMB06, DC03a, DC03b, DC04a, DC04b, WSWS03]. wave-exposure [BI08]. Waves [GBBG06, JHL05]. Wax [MGH08]. Wax-ester [MGH08]. way [GBBG06, GM07, Le 09, LFW03]. weakening [She07]. weaned [OSLO06]. web [HCE^{+}03]. weekly [ODCN09]. weever [Bag04]. weight [BPM^{+}09, CA00, KTRG06, MR05, RLH01, RNK05, RASS09]. weight-at-age [CA00]. weighted [EN02]. Weighting [BVD01, Sim03, MKB01]. Weinberg [KM05]. West [BKN^{+}07, JTE^{+}07, SGM09, BRE^{+}08, BML^{+}05, CFN03, DP03a, ECC06, GRE06, SK07, SP07b, VLBB08, AFHJ04, HWF08, KKC04, LHJJ^{+}01, Nau02, SPWHR04, Wie05]. Western [RD07, SPGT00, dPVJM04, dPVV04, APGD08, Bro03, CMN^{+}07, CZC07, CBHM07, DRDC06, FSQ^{+}03, GOA^{+}09, GFH04, GM06, KI04, KCR07, LD05, LdSSG02, MM03b, MM03a, MSP09, MNMG^{+}05, MS00, Mor02, OKG^{+}09, PCM09, Sil03, SSC^{+}06, SHAH09, TCTC09, dLMS06, dPM08, FSS00, JMWJ08, MAMO02, vDEM^{+}00]. whale [LG08, TF02]. whales [EH04, LHHJ^{+}09, LHR02, NFM^{+}02, RSNB^{+}08, TL05]. Whaling [PD07]. whelk [SDÖ09]. Which [FSP05]. whiffiagonis [LP00]. while [BBPW07]. White [LLD^{+}05, LDQ08, LSM07, RLdAW06, WWGG02, PMB00, PGB03]. white-dotted [RLdAW06]. whitefish [FJK^{+}07, LS04]. whiteweed [WBV09]. whiting [BK07, BBB06b, DM06, FT05, HBD05, JG07, PS06, RMM05, TM02, WGMM08, ZPRJ02]. whole [NHK09, Pel02]. Wide [CD03, BABB08]. Wide-bandwidth [CD03]. widow [SKC^{+}00, SKH02]. Wild [HW06, Agn08, BHN06, BBMS01, BDTW06, DLC03, FSDB09, FJSJBS^{+}08, FLH06, HJ03, HFMD06, JH01a, JJ06, Kol06, OGL06, RPK^{+}03, SJKN^{+}04, SSKE06, SWG06, SMK08, UPK^{+}08]. wild-caught [FSDB09]. Will [YG08, ÅD07]. William [Bro02b, HMQ^{+}08]. Williams [HHAB09]. willingness [PS09]. win [LD03b]. wind [HR09, HSA^{+}01, MHF^{+}09, MM01, SAAFCA07]. wind-driven [HSA^{+}01]. wind-induced [SAAFCA07]. windpower [WMÖ06]. winds [Ber00]. Winter [DLC03, GA00, Hea00b, CZC07, FMF02, NRR^{+}09, Sab04, vdMBD00]. winter-mixing [Sab04]. wintering [HBG^{+}04]. within [BBR08, BFK^{+}07, CBBL09, EGO^{+}07, Fra06, GD05, GH04, GML06, GSSO00, GOA^{+}09, HMPC04, JNF^{+}09, Orl01, PMM^{+}09, PFF01, SF09, VH08, WMS^{+}03]. without [JHC09, RO02, YTS^{+}06]. wobbegong [HOGH07]. words [Hal01]. work [Ska07, SS07]. World [VHI^{+}04, FAL^{+}08, MPD^{+}08, RL08, VPC^{+}09]. Worldwide [RS06b]. wounds [DBS06]. wreck [PAC02]. wrong [DCPvK07].

Xiphias [CPR06].

Yangtze [Aka02, HLL^{+}08, XZW05]. year [BVB^{+}07, Cot01, CMP07, EN02, GOS07, Hor08, HSS^{+}09, íJCMR07, Kup04, OKG^{+}09, OL00, Pie08, PB00, RT03, SHT^{+}09, SL04, SBB^{+}05, SMEK01, VH08, VM09, WS06, GLDB04a, GLDB04b]. year-class [Cot01, CMP07, EN02, GOS07, HSS^{+}09, OKG^{+}09, OL00, WS06]. yearly [MM01, Tje02]. years [AG00, Ber04, BLRC05, KTRG06, Mui03, OBNU02, OUNB02, ODRN05, OL07, Pay04, PSFY07, RRTP02, Rot00, SJKN^{+}04, SF09, Tan00]. yellow [JRN06, Zha06]. yellowtail [CS05, DWC03, WM04]. yield [DC01, Kat05, KPS^{+}05, KB07, SP07a, Ska07]. yield-mortality [DC01]. yield-per-recruit [Kat05, KB07]. yields [Mac09]. Young [Ano01h, Gud04, JMWJ08, KCBC00, Kup04, SHT^{+}09, VH08, Zha06, GLDB04a, GLDB04b]. young-of-the-year [Kup04, SHT^{+}09, GLDB04a, GLDB04b]. YOY [GLDB04a, GLDB04b]. YPR [Kat05].

Záhony [DCN^{+}04]. Zealand [CG07, CDDM05, DRSD09, HBD05, Jel07, O’D03, O’D04, RMKT01, WM01]. Zeus [Dun01, YYY^{+}02]. Zone [PPH09, MKB01, MR09, MMB09, OUNB02, RRY08, TJG^{+}09]. zones [CSVGTP09, KYG03, MSS^{+}05, SMB09]. zoning [BPM^{+}05]. zoogeography [SYR^{+}08]. Zooplankton [APGC04, Kan07, LDCH^{+}09, VHI^{+}04, dPVV04, BGAM00, Ber04, CWC00, CW05, CFN03, DiUVH08, DHKV01, GDH02, GPWG04, Hea07, KSD01, LND05, MKFK05, MMKKJ08, PG08, PBM^{+}04, PGJ^{+}05, PF06, Ric08, SBC^{+}00, SC00, SNB^{+}02, WYMF08, dPVJM04, dPM08, dR01]. ZOOSCAN [GPWG04]. Zostera [BRE^{+}08]. Zygochlamys [MFB^{+}08].

References

Acolas:2004:AUM

[AAV^{+}04] M. L. Acolas, M. L. Bégout Anras, V. Véron, H. Jourdan, M. R. Sabatié, and J. L. Baglinière. An assessment of the upstream migration and reproductive behaviour of allis shad (Alosa alosa L.) using acoustic tracking. ICES Journal of Marine Science, 61 (8):1291–1304, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/8/1291/630120.

Arkhipkin:2008:BSB

[ABB^{+}08] Alexander I. Arkhipkin, Nina Baumgartner, Paul Brickle, Vladimir V. Laptikhovsky, Joost H. W. Pompert, and Zhanna N. Shcherbich. Biology of the skates Bathyraja brachyurops and B. griseocauda in waters around the Falkland Islands, Southwest Atlantic. ICES Journal of Marine Science, 65(4):560–570, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/4/560/635895.

Anras:2003:TJS

[ACD^{+}03] M. L. Bégout Anras, D. Covés, G. Dutto, P. Laffargue, and F. Lagardère. Tagging juvenile seabass and sole with telemetry transmitters: medium-term effects on growth. ICES Journal of Marine Science, 60(6):1328–1334, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/6/1328/653814.

Acosta:2002:SED

[Aco02] Charles A. Acosta. Spatially explicit dispersal dynamics and equilibrium population sizes in marine harvest refuges. ICES Journal of Marine Science, 59(3):458–468, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/3/458/610816.

Aastrom:2007:WWE

[ÅD07] Mårten Åström and Willem Dekker. When will the eel recover? A full life-cycle model. ICES Journal of Marine Science, 64(7): 1491–1498, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/7/1491/728670.

Azevedo:2008:ADF

[ADC^{+}08] Manuela Azevedo, Rafael Duarte, Fátima Cardador, Pedro Sousa, Celso Fariña, Paz Sampedro, Jorge Landa, and Gersom Costas. Application of dynamic factor analysis in the assessment of Iberian anglerfish stocks. ICES Journal of Marine Science, 65(7):1362–1369, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/7/1362/643955.

Axenrot:2004:DPP

[ADDH04] Thomas Axenrot, Tomas Didrikas, Charlotte Danielsson, and Sture Hansson. Diel patterns in pelagic fish behaviour and distribution observed from a stationary, bottom-mounted, and upward-facing transducer. ICES Journal of Marine Science, 61 (7):1100–1104, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/7/1100/878657.

Anderson:2002:JCM

[ADO02] John T. Anderson, Edgar L. Dalley, and Richard L. O’Driscoll. Juvenile capelin (Mallotus villosus) off Newfoundland and Labrador in the 1990s. ICES Journal of Marine Science, 59 (5):917–928, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/5/917/675207.

Adlerstein:2002:EDT

[AE02] Sara Adlerstein and Siegfried Ehrich. Effect of deviations from target speed and of time of day on catch rates of some abundant species under North Sea international bottom trawl survey protocol conditions. ICES Journal of Marine Science, 59 (3):594–603, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/3/594/610843.

Amakasu:2006:TSA

[AF06] Kazuo Amakasu and Masahiko Furusawa. The target strength of Antarctic krill (Euphausia superba) measured by the split-beam method in a small tank at 70 kHz. ICES Journal of Marine Science, 63(1):36–45, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/63/1/36/624934.

Fukuwaka:2008:BSC

[aFADN08] Masa aki Fukuwaka, Tomonori Azumaya, Nancy D. Davis, and Toru Nagasawa. Bias in size composition of chum salmon (Oncorhynchus keta) caught by a gillnet with a geometric series of mesh sizes, and its correction using gear intercalibration. ICES Journal of Marine Science, 65(6):930–936, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/6/930/603502.

Andrefouet:2009:CTS

[AFGR09] Serge Andréfouët, Kim Friedman, Antoine Gilbert, and Georges Remoissenet. A comparison of two surveys of invertebrates at Pacific Ocean islands: the giant clam at Raivavae Island, Australes Archipelago, French Polynesia. ICES Journal of Marine Science, 66(9):1825–1836, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/9/1825/721848.

Alvarez-Flores:2004:RAS

[AFHJ04] Carlos M. Alvarez-Flores and Mads Peter Heide-Jørgensen. A risk assessment of the sustainability of the harvest of beluga (Delphinapterus leucas (Pallas 1776)) in West Greenland. ICES Journal of Marine Science, 61(2):274–286, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/2/274/621806.

Amorim:2009:SVS

[AFM^{+}09] Patrícia Amorim, Miguel Figueiredo, Miguel Machete, Telmo Morato, Ana Martins, and Ricardo Serrão Santos. Spatial variability of seabird distribution associated with environmental factors: a case study of marine important bird areas in the Azores. ICES Journal of Marine Science, 66(1):29–40, January 2009.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/1/29/632961.

Armstrong:2008:HFI

[AFP08] Claire W. Armstrong and Jannike Falk-Petersen. Habitat– fisheries interactions: a missing link? ICES Journal of Marine Science, 65(6):817–821, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/6/817/604979.

Andersen:2009:HCE

[AFP^{+}09] K. H. Andersen, K. D. Farnsworth, M. Pedersen, H. Gislason, and J. E. Beyer. How community ecology links natural mortality, growth, and production of fish populations. ICES Journal of Marine Science, 66(9):1978–1984, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/9/1978/723268.

Anderson:2000:FRS

[AG00] John T. Anderson and Robert S. Gregory. Factors regulating survival of northern cod (NAFO 2J3KL) during their first 3 years of life. ICES Journal of Marine Science, 57(2):349–359, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/2/349/620448.

Agardy:2000:EFM

[Aga00] Tundi Agardy. Effects of fisheries on marine ecosystems: a conservationist’s perspective. ICES Journal of Marine Science, 57 (3):761–765, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/761/636008.

Armstrong:2004:VMG

[AGA^{+}04] M. J. Armstrong, H. D. Gerritsen, M. Allen, W. J. McCurdy, and J. A. D. Peel. Variability in maturity and growth in a heavily exploited stock: cod (Gadus morhua L.) in the Irish Sea. ICES Journal of Marine Science, 61(1):98–112, ???? 2004. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/61/1/98/
702771.

Anderson:2002:ACM

[AGC02] John T. Anderson, Robert S. Gregory, and William T. Collins. Acoustic classification of marine habitats in coastal Newfoundland. ICES Journal of Marine Science, 59(1):156–167, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/1/156/649992.

Abdulla:2009:CFN

[AGH^{+}09] Ameer Abdulla, Marina Gomei, David Hyrenbach, Giuseppe Notarbartolo di Sciara, and Tundi Agardy. Challenges facing a network of representative marine protected areas in the Mediterranean: prioritizing the protection of underrepresented habitats. ICES Journal of Marine Science, 66(1):22–28, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/1/22/632501.

Agnalt:2008:FEL

[Agn08] Ann-Lisbeth Agnalt. Fecundity of the European lobster (Homarus gammarus) off southwestern Norway after stock enhancement: do cultured females produce as many eggs as wild females? ICES Journal of Marine Science, 65(2):164–170, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/2/164/733653.

Andrefouet:2005:RPS

[AGY^{+}05] Serge Andréfouët, Antoine Gilbert, Laurent Yan, Georges Remoissenet, Claude Payri, and Yannick Chancerelle. The remarkable population size of the endangered clam Tridacna maxima assessed in Fangatau Atoll (Eastern Tuamotu, French Polynesia) using in situ and remote sensing data. ICES Journal of Marine Science, 62(6):1037–1048, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/6/1037/615844.

Adams:2008:GCT

[AHS08] Charles F. Adams, Bradley P. Harris, and Kevin D. E. Stokesbury. Geostatistical comparison of two independent video surveys of sea scallop abundance in the Elephant Trunk Closed Area, USA. ICES Journal of Marine Science, 65(6):995– 1003, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/6/995/599934.

Adams:2000:ASA

[AJ00] James A. Adams and Jakob Jakobsson. Alan Saville 14 August 1923–1 November 1998. ICES Journal of Marine Science, 57(4): 1296–1297, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/4/1296/647328.

Armannsson:2007:DMS

[AJNM07] Hlynur Armannsson, Sigurdur Th. Jonsson, John D. Neilson, and Gudrun Marteinsdottir. Distribution and migration of saithe (Pollachius virens) around Iceland inferred from markrecapture studies. ICES Journal of Marine Science, 64(5): 1006–1016, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/5/1006/642916.

Arkhipkin:2000:GMT

[AJR00] A. Arkhipkin, P. Jereb, and S. Ragonese. Growth and maturation in two successive seasonal groups of the short-finned squid, Illex coindetii from the Strait of Sicily (central Mediterranean). ICES Journal of Marine Science, 57(1):31–41, February 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/1/31/641142.

Andrade:2004:EBC

[AK04] Humber A. Andrade and Paul G. Kinas. Estimation of birthdates and catch-at-age using length frequency analysis (LFA), with application for skipjack tuna (Katsuwonus pelamis) caught in the Southwest Atlantic. ICES Journal of Marine Science, 61 (5):798–811, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/5/798/866259.

Akamatsu:2002:DBF

[Aka02] T. Akamatsu. Diving behaviour of freshwater finless porpoises (Neophocaena phocaenoides) in an oxbow of the Yangtze River, China. ICES Journal of Marine Science, 59(2):438–443, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/2/438/619669.

Agnalt:2007:GRC

[AKJ07] Ann-Lisbeth Agnalt, Tore S. Kristiansen, and Knut E. Jørstad. Growth, reproductive cycle, and movement of berried European lobsters (Homarus gammarus) in a local stock off southwestern Norway. ICES Journal of Marine Science, 64(2):288–297, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/2/288/2182156.

Aps:2007:NFB

[AKLL07] Robert Aps, Laurence T. Kell, Hans Lassen, and Innar Liiv. Negotiation framework for Baltic fisheries management: striking the balance of interest. ICES Journal of Marine Science, 64 (4):858–861, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/4/858/638828.

Aksland:2005:AEI

[Aks05] Magnar Aksland. An alternative echo-integrating method. ICES Journal of Marine Science, 62(2):226–235, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/2/226/603539.

Aksland:2006:AAM

[Aks06] Magnar Aksland. Applying an alternative method of echointegration. ICES Journal of Marine Science, 63(8):1438–1452, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/8/1438/713017.

Alonso:2001:FHD

[Alo01] M. Koen Alonso. Food habits of Dipturus chilensis (Pisces: Rajidae) off Patagonia, Argentina. ICES Journal of Marine Science, 58(1):288–297, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/58/1/288/603530.

Alos:2008:IAH

[Aló08] Josep Alós. Influence of anatomical hooking depth, capture depth, and venting on mortality of painted comber (Serranus scriba) released by recreational anglers. ICES Journal of Marine Science, 65(9):1620–1625, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/9/1620/630844.

Alheit:2005:SER

[AMD^{+}05] J. Alheit, C. Möllmann, J. Dutz, G. Kornilovs, P. Loewe, V. Mohrholz, and N. Wasmund. Synchronous ecological regime shifts in the central Baltic and the North Sea in the late 1980s. ICES Journal of Marine Science, 62(7):1205–1215, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/7/1205/654639.

Ayllon:2006:LRP

[AMGV06] Fernando Ayllon, Jose L. Martinez, and Eva Garcia-Vazquez. Loss of regional population structure in Atlantic salmon, Salmo salar L., following stocking. ICES Journal of Marine Science, 63 (7):1269–1273, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/7/1269/756978.

Ayllon:2006:GHP

[AMJ^{+}06] Fernando Ayllon, Jose L. Martinez, Francis Juanes, Stephen Gephard, and Eva Garcia-Vazquez. Genetic history of the population of Atlantic salmon, Salmo salar L., under restoration in the Connecticut River, USA. ICES Journal of Marine Science, 63(7):1286–1289, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/7/1286/757404.

Albert:2001:DAJ

[ANNG01] Ole Thomas Albert, Einar M. Nilssen, Kjell H. Nedreaas, and Agnes C. Gundersen. Distribution and abundance of juvenile Northeast Arctic Greenland halibut (Reinhardtius hippoglossoides) in relation to survey coverage and the physical environment. ICES Journal of Marine Science, 58(5):1053–1062, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/5/1053/730824.

Anonymous:2000:IV

[Ano00a] Anonymous. Index to volume 57. ICES Journal of Marine Science, 57(6):1875, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/6/1875/768704.

Anonymous:2000:LRa

[Ano00b] Anonymous. List of referees. ICES Journal of Marine Science, 57(3):792, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/792/636013.

Anonymous:2000:LRb

[Ano00c] Anonymous. List of referees. ICES Journal of Marine Science, 57(5):1524, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/5/1524/660974.

Anonymous:2000:LRc

[Ano00d] Anonymous. List of referees. ICES Journal of Marine Science, 57(6):1875, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/6/1875/768708.

Anonymous:2001:Ea

[Ano01a] Anonymous. Erratum. ICES Journal of Marine Science, 58 (6):1340, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1340/641590.

Anonymous:2001:Eb

[Ano01b] Anonymous. Erratum. ICES Journal of Marine Science, 58 (6):1341, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1341/641593.

Anonymous:2001:IV

[Ano01c] Anonymous. Index to volume 58. ICES Journal of Marine Science, 58(6):1343, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/58/6/1343/641596.

Anonymous:2001:LRa

[Ano01d] Anonymous. List of referees. ICES Journal of Marine Science, 58(2):530, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/2/530/617867.

Anonymous:2001:LRb

[Ano01e] Anonymous. List of referees. ICES Journal of Marine Science, 58(4):934, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/4/934/630234.

Anonymous:2001:LRc

[Ano01f] Anonymous. List of referees. ICES Journal of Marine Science, 58(5):1114, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/5/1114/730852.

Anonymous:2001:LRd

[Ano01g] Anonymous. List of referees. ICES Journal of Marine Science, 58(6):1342–1343, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1342/641595.

Anonymous:2001:YSC

[Ano01h] Anonymous. The Young Scientists Conference on Marine Ecosystem Perspectives. ICES Journal of Marine Science, 58 (3):687–688, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/3/687/810124.

Anonymous:2002:LRa

[Ano02a] Anonymous. List of referees. ICES Journal of Marine Science, 59(1):235, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/1/235/650021.

Anonymous:2002:LRb

[Ano02b] Anonymous. List of referees. ICES Journal of Marine Science, 59(6):1337–1338, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/6/1337/608203.

Anonymous:2002:LRc

[Ano02c] Anonymous. List of referees. ICES Journal of Marine Science, 59(S1):S63, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S63/618006.

Anonymous:2002:R

[Ano02d] Anonymous. Referees. ICES Journal of Marine Science, 59 (5):1131, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/5/1131/675185.

Anonymous:2003:LR

[Ano03a] Anonymous. List of referees. ICES Journal of Marine Science, 60(6):1399–1400, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/6/1399/654964.

Anonymous:2003:R

[Ano03b] Anonymous. Referees. ICES Journal of Marine Science, 60 (3):707, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/3/707/661432.

Anonymous:2004:LR

[Ano04] Anonymous. List of referees. ICES Journal of Marine Science, 61(7):1242, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/7/1242/883509.

Anonymous:2005:Ra

[Ano05a] Anonymous. Referees. ICES Journal of Marine Science, 62 (3):614, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/3/614/667179.

Anonymous:2005:Rb

[Ano05b] Anonymous. Referees. ICES Journal of Marine Science, 62 (7):1542, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/7/1542/661793.

Anonymous:2006:EBPa

[Ano06a] Anonymous. Editorial Board & publication details. ICES Journal of Marine Science, 63(1):176, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/1/176/624594.

Anonymous:2006:EBPb

[Ano06b] Anonymous. Editorial Board & publication details. ICES Journal of Marine Science, 63(2):400, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/2/400/642575.

Anonymous:2006:EBPc

[Ano06c] Anonymous. Editorial Board & publication details. ICES Journal of Marine Science, 63(3):571, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/3/571/723966.

Anonymous:2006:EBPd

[Ano06d] Anonymous. Editorial Board & publication details. ICES Journal of Marine Science, 63(4):774, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/4/774/694972.

Anonymous:2006:EBPe

[Ano06e] Anonymous. Editorial Board & publication details. ICES Journal of Marine Science, 63(5):959, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/5/959/664655.

Anonymous:2006:EBPf

[Ano06f] Anonymous. Editorial Board & publication details. ICES Journal of Marine Science, 63(6):1158, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/6/1158/617924.

Anonymous:2006:EBPg

[Ano06g] Anonymous. Editorial Board & publication details. ICES Journal of Marine Science, 63(7):1372, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/7/1372/762250.

Anonymous:2006:EBPh

[Ano06h] Anonymous. Editorial Board & publication details. ICES Journal of Marine Science, 63(8):1566, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/8/1566/715997.

Anonymous:2006:EBPi

[Ano06i] Anonymous. Editorial Board & publication details. ICES Journal of Marine Science, 63(9):1767, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/9/1767/701760.

Anonymous:2006:MDJ

[Ano06j] Anonymous. In memoriam — Dr Joseph A. Brown. ICES Journal of Marine Science, 63(2):192–193, ???? 2006. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/2/192/636972.

Anonymous:2006:Ra

[Ano06k] Anonymous. Referees. ICES Journal of Marine Science, 63 (2):400, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/2/400/642579.

Anonymous:2006:Rb

[Ano06l] Anonymous. Referees. ICES Journal of Marine Science, 63 (7):1372, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/7/1372/762265.

Anonymous:2006:Rc

[Ano06m] Anonymous. Referees. ICES Journal of Marine Science, 63 (9):1766–1767, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/9/1766/701759.

Anonymous:2007:LR

[Ano07a] Anonymous. List of referees. ICES Journal of Marine Science, 64(8):1616, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/8/1616/614004.

Anonymous:2007:Ra

[Ano07b] Anonymous. Referees. ICES Journal of Marine Science, 64(4): 862, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/4/862/642430.

Anonymous:2007:Rb

[Ano07c] Anonymous. Referees. ICES Journal of Marine Science, 64(9): 1845–1847, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/9/1845/789202.

Anonymous:2008:LRa

[Ano08a] Anonymous. List of referees. ICES Journal of Marine Science, 65(3):495, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/3/495/788886.

Anonymous:2008:LRb

[Ano08b] Anonymous. List of referees. ICES Journal of Marine Science, 65(5):816, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/5/816/715153.

Anonymous:2008:LRc

[Ano08c] Anonymous. List of referees. ICES Journal of Marine Science, 65(8):1540, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/8/1540/714051.

Anonymous:2008:LRd

[Ano08d] Anonymous. List of referees. ICES Journal of Marine Science, 65(9):1773–1775, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/9/1773/633227.

Anonymous:2009:LRa

[Ano09a] Anonymous. List of referees. ICES Journal of Marine Science, 66(1):225, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/1/225/635227.

Anonymous:2009:LRb

[Ano09b] Anonymous. List of referees. ICES Journal of Marine Science, 66(6):1433, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1433/697791.

Anonymous:2009:LRc

[Ano09c] Anonymous. List of referees. ICES Journal of Marine Science, 66(7):1647, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/7/1647/658983.

Anonymous:2009:LRd

[Ano09d] Anonymous. List of reviewers. ICES Journal of Marine Science, 66(8):1823, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/8/1823/675136.

Auker:2008:FIR

[AO08] Linda A. Auker and Candace A. Oviatt. Factors influencing the recruitment and abundance of Didemnum in Narragansett Bay, Rhode Island. ICES Journal of Marine Science, 65(5):765–769, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/5/765/712427.

Axenrot:2009:MDF

[AOSD09] Thomas Axenrot, Martin Ogonowski, Alfred Sandström, and Tomas Didrikas. Multifrequency discrimination of fish and mysids. ICES Journal of Marine Science, 66(6):1106–1110, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1106/697303.

Abdolmalaki:2007:ESE

[AP07] Shahram Abdolmalaki and Iwona Psuty. The effects of stock enhancement of pikeperch (Sander lucioperca) in Iranian coastal waters of the Caspian Sea. ICES Journal of Marine Science, 64(5):973–980, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/5/973/640995.

Aarts:2009:CDR

[AP09] G. Aarts and J. J. Poos. Comprehensive discard reconstruction and abundance estimation using flexible selectivity functions. ICES Journal of Marine Science, 66(4):763–771, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/4/763/603549.

Amar:2009:ETM

[APD09] Z. Teresa A’mar, André E. Punt, and Martin W. Dorn. The evaluation of two management strategies for the Gulf of Alaska walleye pollock fishery under climate change. ICES Journal of Marine Science, 66(7):1614–1632, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/7/1614/656485.

Ayon:2004:ZVT

[APGC04] Patricia Ayón, Sara Purca, and Renato Guevara-Carrasco. Zooplankton volume trends off Peru between 1964 and 2001. ICES Journal of Marine Science, 61(4):478–484, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/4/478/602728.

Alos:2008:EHS

[APGD08] Josep Alós, Miquel Palmer, Antoni Maria Grau, and Salud Deudero. Effects of hook size and barbless hooks on hooking injury, catch per unit effort, and fish size in a mixed-species recreational fishery in the western Mediterranean Sea. ICES Journal of Marine Science, 65(6):899–905, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/6/899/602201.

Ardron:2008:TIO

[Ard08] Jeff A. Ardron. Three initial OSPAR tests of ecological coherence: heuristics in a data-limited situation. ICES Journal of Marine Science, 65(8):1527–1533, November 2008. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/8/1527/712765.

Armstrong:2001:AAE

[Arm01] M. Armstrong. An application of the annual egg production method to estimate the spawning biomass of cod (Gadus morhua L.), plaice (Pleuronectes platessa L.) and sole (Solea solea L.) in the Irish Sea. ICES Journal of Marine Science, 58(1):183–203, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/1/183/603516.

Arrizabalaga:2009:USA

[ARMM09] Haritz Arrizabalaga, Victor R. Restrepo, Mark N. Maunder, and Jacek Majkowski. Using stock assessment information to assess fishing capacity of tuna fisheries. ICES Journal of Marine Science, 66(9):1959–1966, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/9/1959/724556.

Arnason:2000:EIA

[Arn00] Ragnar Arnason. Economic instruments for achieving ecosystem objectives in fisheries management. ICES Journal of Marine Science, 57(3):742–751, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/3/742/636005.

Abelson:2002:CDC

[AS02a] Avigdor Abelson and Yehiam Shlesinger. Comparison of the development of coral and fish communities on rock-aggregated artificial reefs in Eilat, Red Sea. ICES Journal of Marine Science, 59(S1):S122–S126, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/59/suppl/S122/617902.

Angel:2002:AAR

[AS02b] Dror L. Angel and Ehud Spanier. An application of artificial reefs to reduce organic enrichment caused by net-cage fish farming: preliminary results. ICES Journal of Marine Science, 59 (S1):S324–S329, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S324/617984.

Ault:2005:EAL

[ASB05] Jerald S. Ault, Steven G. Smith, and James A. Bohnsack. Evaluation of average length as an estimator of exploitation status for the Florida coral-reef fish community. ICES Journal of Marine Science, 62(3):417–423, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/3/417/661641.

Arzul:2001:EMA

[ASC01] Geneviève Arzul, Miriam Seguel, and Alejandro Clément. Effect of marine animal excretions on differential growth of phytoplankton species. ICES Journal of Marine Science, 58(2): 386–390, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/2/386/617829.

Andersen:2007:DHS

[ATH^{+}07] Signe M. Andersen, Jonas Teilmann, Pernille B. Harders, Else H. Hansen, and Dorthe Hjøllund. Diet of harbour seals and great cormorants in Limfjord, Denmark: interspecific competition and interaction with fishery. ICES Journal of Marine Science, 64(6): 1235–1245, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/6/1235/615338.

Adlerstein:2002:CSC

[ATM02] Sara A. Adlerstein, Axel Temming, and Nils Mergardt. Comparison of stomach contents of haddock (Melanogrammus aeglefinus) from the 1981 and 1991 North Sea international stomach sampling projects. ICES Journal of Marine Science, 59 (3):497–515, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/3/497/610820.

Acolas:2006:UMR

[AVJ^{+}06] M. L. Acolas, V. Véron, H. Jourdan, M. L. Bégout, M. R. Sabatié, and J. L. Baglinière. Upstream migration and reproductive patterns of a population of allis shad in a small river (L’Aulne, Brittany, France). ICES Journal of Marine Science, 63(3):476–484, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/3/476/721164.

Anderson:2008:ASC

[AVK^{+}08] John T. Anderson, D. Van Holliday, Rudy Kloser, Dave G. Reid, and Yvan Simard. Acoustic seabed classification: current practice and future directions. ICES Journal of Marine Science, 65 (6):1004–1011, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/6/1004/601054.

Aprahamian:2007:AME

[AWW^{+}07] M. W. Aprahamian, A. M. Walker, B. Williams, A. Bark, and B. Knights. On the application of models of European eel (Anguilla anguilla) production and escapement to the development of eel management plans: the River Severn. ICES Journal of Marine Science, 64(7):1472–1482, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/7/1472/729051.

Berg:2003:CFC

[BA03] Erik Berg and Ole Thomas Albert. Cod in fjords and coastal waters of North Norway: distribution and variation in length and maturity at age. ICES Journal of Marine Science, 60 (4):787–797, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/4/787/692382.

Brierley:2004:STE

[BAB^{+}04] Andrew S. Brierley, Bjørn Eric Axelsen, David C. Boyer, Christopher P. Lynam, Carol A. Didcock, Helen J. Boyer, Conrad A. J. Sparks, Jennifer E. Purcell, and Mark J. Gibbons. Single-target echo detections of jellyfish. ICES Journal of Marine Science, 61(3):383–393, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/61/3/383/671936.

Baker:2008:RWP

[BABB08] Patrick Baker, James D. Austin, Brian W. Bowen, and Shirley M. Baker. Range-wide population structure and history of the northern quahog (Merceneria merceneria) inferred from mitochondrial DNA sequence data. ICES Journal of Marine Science, 65(2):155–163, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/2/155/736574.

Backer:2008:ISK

[Bac08] Hermanni Backer. Indicators and scientific knowledge in regional Baltic Sea environmental policy. ICES Journal of Marine Science, 65(8):1398–1401, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/8/1398/714665.

Bagge:2004:BGW

[Bag04] Ole Bagge. The biology of the greater weever (Trachinus draco) in the commercial fishery of the Kattegat. ICES Journal of Marine Science, 61(6):933–943, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/6/933/678152.

Baine:2002:NSR

[Bai02] Mark Baine. The North Sea rigs-to-reefs debate. ICES Journal of Marine Science, 59(S1):S277–S280, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S277/617964.

Baird:2009:AFV

[Bai09] Dan Baird. An assessment of the functional variability of selected coastal ecosystems in the context of local environmental changes. ICES Journal of Marine Science, 66(7):1520–1527, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/7/1520/656595.

Barrett:2002:FCS

[BANGC02] Robert T. Barrett, Tycho Anker-Nilssen, Geir W. Gabrielsen, and Gilles Chapdelaine. Food consumption by seabirds in Norwegian waters. ICES Journal of Marine Science, 59(1):43–57, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/1/43/650029.

Bode:2004:TVT

[BAO04] Antonio Bode and Maria Teresa Alvarez-Ossorio. Taxonomic versus trophic structure of mesozooplankton: a seasonal study of species succession and stable carbon and nitrogen isotopes in a coastal upwelling ecosystem. ICES Journal of Marine Science, 61(4):563–571, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/4/563/604390.

Bartsch:2005:IST

[Bar05] Joachim Bartsch. The influence of spatio-temporal egg production variability on the modelled survival of the early life history stages of mackerel (Scomber scombrus) in the eastern North Atlantic. ICES Journal of Marine Science, 62(6):1049–1060, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/6/1049/616028.

Beaulaton:2007:EMM

[BB07] Laurent Beaulaton and Cédric Briand. Effect of management measures on glass eel escapement. ICES Journal of Marine Science, 64(7):1402–1413, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/64/7/1402/726053.

Benoit-Bird:2009:ESL

[BB09] Kelly J. Benoit-Bird. The effects of scattering-layer composition, animal size, and numerical density on the frequency response of volume backscatter. ICES Journal of Marine Science, 66 (3):582–593, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/3/582/816452.

Benoit-Bird:2003:SDN

[BBA03] Kelly J. Benoit-Bird and Whitlow W. L. Au. Spatial dynamics of a nearshore, micronekton sound-scattering layer. ICES Journal of Marine Science, 60(4):899–913, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/4/899/694600.

Blanco-Bercial:2006:CMC

[BBÁMC06] Leocadio Blanco-Bercial, Florentina Álvarez-Marqués, and Jesús A. Cabal. Changes in the mesozooplankton community associated with the hydrography off the northwestern Iberian Peninsula. ICES Journal of Marine Science, 63(5):799–810, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/5/799/661565.

Bricknell:2006:DGF

[BBB06a] Ian R. Bricknell, James E. Bron, and Tim J. Bowden. Diseases of gadoid fish in cultivation: a review. ICES Journal of Marine Science, 63(2):253–266, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/2/253/638979.

Butcher:2006:MSW

[BBB06b] Paul A. Butcher, Matt K. Broadhurst, and Craig P. Brand. Mortality of sand whiting (Sillago ciliata) released by recreational anglers in an Australian estuary. ICES Journal of Marine Science, 63(3):567–571, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/63/3/567/723797.

Bjorge:2002:IBH

[BBBF02] Arne Bjørge, Trine Bekkby, Vegar Bakkestuen, and Erik Framstad. Interactions between harbour seals, Phoca vitulina, and fisheries in complex coastal waters explored by combined Geographic Information System (GIS) and energetics modelling. ICES Journal of Marine Science, 59(1):29–42, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/59/1/29/
650026.

Bertrand:2004:DVB

[BBC^{+}04] Arnaud Bertrand, Maria Angela Barbieri, Jose Córdova, Carola Hernández, Fabián Gómez, and Francisco Leiva. Diel vertical behaviour, predator–prey relationships, and occupation of space by jack mackerel (Trachurus murphyi) off Chile. ICES Journal of Marine Science, 61(7):1105–1112, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/7/1105/878772.

Bertrand:2005:LTP

[BBGA05] Sophie Bertrand, Julian M. Burgos, François Gerlotto, and Jaime Atiquipa. Lévy trajectories of Peruvian purse-seiners as an indicator of the spatial distribution of anchovy (Engraulis ringens). ICES Journal of Marine Science, 62(3):477–482, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/3/477/663434.

Burke:2008:OSA

[BBK08] Nóirín Burke, Deirdre Brophy, and Pauline A. King. Otolith shape analysis: its application for discriminating between stocks of Irish Sea and Celtic Sea herring (Clupea harengus) in the Irish Sea. ICES Journal of Marine Science, 65(9):1670–1675, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/9/1670/632202.

Beacham:2002:MSS

[BBM^{+}02] Terry D. Beacham, John Brattey, Kristina M. Miller, Khai D. Le, and Ruth E. Withler. Multiple stock structure of Atlantic cod (Gadus morhua) off Newfoundland and Labrador determined from genetic variation. ICES Journal of Marine Science, 59(4):650–665, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/4/650/676764.

Bridger:2001:SFD

[BBMS01] C. J. Bridger, R. K. Booth, R. S. McKinley, and D. A. Scruton. Site fidelity and dispersal patterns of domestic triploid steelhead trout (Oncorhynchus mykiss Walbaum) released to the wild. ICES Journal of Marine Science, 58(2):510–516, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/2/510/617863.

Brandon:2007:CCJ

[BBPW07] John R. Brandon, Jeffrey M. Breiwick, André E. Punt, and Paul R. Wade. Constructing a coherent joint prior while respecting biological realism: application to marine mammal stock assessments. ICES Journal of Marine Science, 64(6):1085– 1100, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/6/1085/615665.

Bayer:2008:RFM

[BBR08] Elizabeth Bayer, Richard A. Barnes, and Hubert L. Rees. The regulatory framework for marine dredging indicators and their operational efficiency within the UK: a possible model for other nations? ICES Journal of Marine Science, 65(8): 1402–1406, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/8/1402/714227.

Bakun:2009:RCA

[BBS09a] Andrew Bakun, Elizabeth A. Babcock, and Christine Santora. Regulating a complex adaptive system via its wasp-waist: grappling with ecosystem-based management of the New England herring fishery. ICES Journal of Marine Science, 66(8):1768– 1775, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/8/1768/673350.

Bourguignon:2009:MDI

[BBS^{+}09b] Sébastien Bourguignon, Laurent Berger, Carla Scalabrin, Ronan Fablet, and Valérie Mazauric. Methodological developments for improved bottom detection with the ME70 multibeam echosounder. ICES Journal of Marine Science, 66(6):1015–1022, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1015/694054.

Burke:2009:TTS

[BBSK09] Nóirín Burke, Deirdre Brophy, Pieter-Jan Schön, and Pauline A. King. Temporal trends in stock origin and abundance of juvenile herring (Clupea harengus) in the Irish Sea. ICES Journal of Marine Science, 66(8):1749–1753, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/8/1749/675071.

Barber:2007:ITD

[BC07] Julie S. Barber and J. Stanley Cobb. Injury in trapped Dungeness crabs (Cancer magister). ICES Journal of Marine Science, 64(3):464–472, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/64/3/464/817174.

Barrett:2006:SNP

[BCAN^{+}06] Robert T. Barrett, Gilles Chapdelaine, Tycho Anker-Nilssen, Anders Mosbech, William A. Montevecchi, James B. Reid, and Richard R. Veit. Seabird numbers and prey consumption in the North Atlantic. ICES Journal of Marine Science, 63(6): 1145–1158, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/6/1145/617302.

Barrett:2007:DSS

[BCAN^{+}07] Robert T. Barrett, Kees (C. J.) Camphuysen, Tycho Anker-Nilssen, John W. Chardine, Robert W. Furness, Stefan Garthe, Ommo Hüppop, Mardik F. Leopold, William A. Montevecchi, and Richard R. Veit. Diet studies of seabirds: a review and recommendations. ICES Journal of Marine Science, 64(9):1675– 1691, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/9/1675/784392.

Badalamenti:2002:ARC

[BCD^{+}02] F. Badalamenti, R. Chemello, G. D’anna, P. Henriquez Ramos, and S. Riggio. Are artificial reefs comparable to neighbouring natural rocky areas? A mollusc case study in the Gulf of Castellammare (NW Sicily). ICES Journal of Marine Science, 59(S1): S127–S131, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S127/617903.

Bode:2003:PFU

[BCL03] Antonio Bode, Pablo Carrera, and Santiago Lens. The pelagic foodweb in the upwelling ecosystem of Galicia (NW Spain) during spring: natural abundance of stable carbon and nitrogen isotopes. ICES Journal of Marine Science, 60(1):11–22, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/1/11/611423.

Barange:2005:SSO

[BCT05] Manuel Barange, Janet C. Coetzee, and Nandipha M. Twatwa. Strategies of space occupation by anchovy and sardine in the southern Benguela: the role of stock size and intra-species competition. ICES Journal of Marine Science, 62(4):645–654, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/4/645/673354.

Brophy:2002:TPA

[BD02] Deirdre Brophy and Bret S. Danilowicz. Tracing populations of Atlantic herring (Clupea harengus L.) in the Irish and Celtic Seas using otolith microstructure. ICES Journal of Marine Science, 59(6):1305–1313, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/6/1305/608196.

Brophy:2003:IPR

[BD03] Deirdre Brophy and Bret S. Danilowicz. The influence of pre-recruitment growth on subsequent growth and age at first spawning in Atlantic herring (Clupea harengus L.). ICES Journal of Marine Science, 60(5):1103–1113, ???? 2003. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/5/1103/766607.

Bjornsson:2004:QNF

[BD04] Björn Björnsson and Maria Álvaro Dongala Dombaxe. Quality of Nephrops as food for Atlantic cod (Gadus morhua L.) with possible implications for fisheries management. ICES Journal of Marine Science, 61(6):983–991, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/6/983/679150.

Bult:2007:EFS

[BD07] Tammo P. Bult and Willem Dekker. Experimental field study on the migratory behaviour of glass eels (Anguilla anguilla) at the interface of fresh and salt water. ICES Journal of Marine Science, 64(7):1396–1401, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/7/1396/727273.

Bez:2006:RFL

[BDD06] N. Bez, E. De Oliveira, and G. Duhamel. Repetitive fishing, local depletion, and fishing efficiencies in the Kerguelen Islands fisheries. ICES Journal of Marine Science, 63(3):532–542, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/3/532/722924.

Blanchard:2005:DCF

[BDJ^{+}05] Julia L. Blanchard, Nicholas K. Dulvy, Simon Jennings, James R. Ellis, John K. Pinnegar, Alex Tidd, and Laurence T. Kell. Do climate and fishing influence size-based indicators of Celtic Sea fish community structure? ICES Journal of Marine Science, 62(3):405–411, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/62/3/405/661410.

Bachelet:2000:SCM

[BdMAL00] Guy Bachelet, Xavier de Montaudouin, Isabelle Auby, and Pierre-Jean Labourg. Seasonal changes in macrophyte and macrozoobenthos assemblages in three coastal lagoons under varying degrees of eutrophication. ICES Journal of Marine Science, 57(5):1495–1506, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/57/5/1495/660968.

Bertrand:2004:IBF

[BDÑ04] Sophie Bertrand, Erich Díaz, and Miguel Ñiquen. Interactions between fish and fisher’s spatial distribution and behaviour: an empirical study of the anchovy (Engraulis ringens) fishery of Peru. ICES Journal of Marine Science, 61(7):1127–1136, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/7/1127/879273.

Breen:2004:SEH

[BDO^{+}04] Mike Breen, Jamie Dyson, Finbarr G. O’Neill, Emma Jones, and Michael Haigh. Swimming endurance of haddock (Melanogrammus aeglefinus L.) at prolonged and sustained swimming speeds, and its role in their capture by towed fishing gears. ICES Journal of Marine Science, 61(7):1071–1079, ???? 2004. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1071/878207.

Bertignac:2007:CBA

[BdP07] Michel Bertignac and Hélène de Pontual. Consequences of bias in age estimation on assessment of the northern stock of European hake (Merluccius merluccius) and on management advice. ICES Journal of Marine Science, 64(5):981–988, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/5/981/640661.

Brander:2001:MTP

[BDS01] K. M. Brander, R. R. Dickson, and J. G. Shepherd. Modelling the timing of plankton production and its effect on recruitment of cod (Gadus morhua). ICES Journal of Marine Science, 58 (5):962–966, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/5/962/730868.

Brooking:2006:STW

[BDTW06] Paul Brooking, Gino Doucette, Steve Tinker, and Frederick G. Whoriskey. Sonic tracking of wild cod, Gadus morhua, in an inshore region of the Bay of Fundy: a contribution to understanding the impact of cod farming for wild cod and endangered salmon populations. ICES Journal of Marine Science, 63 (7):1364–1371, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/7/1364/761915.

Beaugrand:2005:MPE

[Bea05] Grégory Beaugrand. Monitoring pelagic ecosystems using plankton indicators. ICES Journal of Marine Science, 62(3):333–338, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/3/333/658946.

Barbaro:2009:MSM

[BEB^{+}09] Alethea Barbaro, Baldvin Einarsson, Björn Birnir, Sven Sigurdsson, Hédinn Valdimarsson, Ólafur Karvel Pálsson, Sveinn Sveinbjörnsson, and orsteinn Sigurdsson. Modelling and simulations of the migration of pelagic fish. ICES Journal of Marine Science, 66(5):826–838, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/66/5/826/663038.

Benfey:2001:UST

[Ben01] Tillmann J. Benfey. Use of sterile triploid Atlantic salmon (Salmo salar L.) for aquaculture in New Brunswick, Canada. ICES Journal of Marine Science, 58(2):525–529, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/2/525/617865.

Bergeron:2000:ESW

[Ber00] Jean-Pierre Bergeron. Effect of strong winds on the nutritional condition of anchovy (Engraulis encrasicolus L.) larvae in the Bay of Biscay, Northeast Atlantic, as inferred from an early field application of the DNA/C index. ICES Journal of Marine Science, 57(2):249–255, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/2/249/620420.

Bergeron:2004:CYG

[Ber04] Jean-Pierre Bergeron. Contrasting years in the Gironde estuary (Bay of Biscay, NE Atlantic) springtime outflow and consequences for zooplankton pyruvate kinase activity and the nutritional condition of anchovy larvae: an early view. ICES Journal of Marine Science, 61(6):928–932, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/6/928/678084.

Bruce:2001:IMO

[BES^{+}01] B. D. Bruce, K. Evans, C. A. Sutton, J. W. Young, and D. M. Furlani. Influence of mesoscale oceanographic processes on larval distribution and stock structure in jackass morwong (Nemadactylus macropterus: Cheilodactylidae). ICES Journal of Marine Science, 58(5):1072–1080, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/5/1072/730831.

Bethke:2004:ENT

[Bet04] Eckhard Bethke. The evaluation of noise- and threshold-induced bias in the integration of single-fish echoes. ICES Journal of Marine Science, 61(3):405–415, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/3/405/672364.

Bjorn:2002:SLL

[BF02] Pål Arne Bjørn and Bengt Finstad. Salmon lice, Lepeophtheirus salmonis (Krøyer), infestation in sympatric populations of Arctic char, Salvelinus alpinus (L.), and sea trout, Salmo trutta (L.), in areas near and distant from salmon farms. ICES Journal of Marine Science, 59(1):131–139, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/1/131/649976.

Bonnet:2004:SCS

[BF04] D. Bonnet and C. Frid. Seven copepod species considered as indicators of water-mass influence and changes: results from a Northumberland coastal station. ICES Journal of Marine Science, 61(4):485–491, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/61/4/485/602964.

Bjorn:2007:DRC

[BFK^{+}07] P. A. Bjørn, B. Finstad, R. Kristoffersen, R. S. McKinley, and A. H. Rikardsen. Differences in risks and consequences of salmon louse, Lepeophtheirus salmonis (Krøyer), infestation on sympatric populations of Atlantic salmon, brown trout, and Arctic charr within northern fjords. ICES Journal of Marine Science, 64(2):386–393, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/2/386/2182381.

Ball:2000:LST

[BFM00] B. J. Ball, G. Fox, and B. W. Munday. Long- and short-term consequences of a Nephrops trawl fishery on the benthos and environment of the Irish Sea. ICES Journal of Marine Science, 57(5):1315–1320, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/5/1315/660923.

Burwen:2003:TBS

[BFMJ03] Debby L. Burwen, Steven J. Fleischman, James D. Miller, and Mark E. Jensen. Time-based signal characteristics as predictors of fish size and species for a side-looking hydroacoustic application in a river. ICES Journal of Marine Science, 60(3):662–668, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/3/662/660764.

Burton:2002:AKM

[BFSC02] W. H. Burton, J. S. Farrar, F. Steimle, and B. Conlin. Assessment of out-of-kind mitigation success of an artificial reef deployed in Delaware Bay, USA. ICES Journal of Marine Science, 59(S1):S106–S110, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/59/suppl/S106/617897.

Bundy:2005:BEC

[BFZ05] Alida Bundy, Paul Fanning, and Kees C. T. Zwanenburg. Balancing exploitation and conservation of the eastern Scotian Shelf ecosystem: application of a 4D ecosystem exploitation index. ICES Journal of Marine Science, 62(3):503–510, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/3/503/664229.

Bjordal:2004:OTE

[BG04] Åsmund Bjordal and François Gerlotto. Observation techniques and experimental design. ICES Journal of Marine Science, 61 (7):1233–1234, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/7/1233/882364.

Belpaire:2007:ECC

[BG07] C. Belpaire and G. Goemans. Eels: contaminant cocktails pinpointing environmental contamination. ICES Journal of Marine Science, 64(7):1423–1436, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/7/1423/728488.

Beare:2000:ALT

[BGAM00] D. J. Beare, A. Gislason, O. S. Astthorsson, and E. McKenzie. Assessing long-term changes in early summer zooplankton communities around Iceland. ICES Journal of Marine Science, 57 (6):1545–1561, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/6/1545/768538.

Brehmer:2006:EVU

[BGG^{+}06] P. Brehmer, J. Guillard, Y. Guennégan, J. L. Bigot, and B. Liorzou. Evidence of a variable “unsampled” pelagic fish biomass in shallow water (<20 m): the case of the Gulf of Lion. ICES Journal of Marine Science, 63(3):444–451, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/3/444/719818.

Belpaire:2008:PFE

[BGG^{+}08] C. Belpaire, G. Goemans, C. Geeraerts, P. Quataert, and K. Parmentier. Pollution fingerprints in eels as models for the chemical status of rivers. ICES Journal of Marine Science, 65(8): 1483–1491, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/8/1483/712835.

Bang:2008:RBC

[BGL08] A. Bang, P. Grønkjær, and B. Lorenzen. The relation between concentrations of ovarian trace elements and the body size of Atlantic cod Gadus morhua. ICES Journal of Marine Science, 65(7):1191–1197, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/7/1191/643589.

Brierley:2003:BME

[BGW03] Andrew S. Brierley, Stephen F. Gull, and Maged H. Wafy. A Bayesian maximum entropy reconstruction of stock distribution and inference of stock density from line-transect acousticsurvey data. ICES Journal of Marine Science, 60(3):446–452, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/3/446/657820.

Braccini:2005:SVF

[BGW05] J. Matías Braccini, Bronwyn M. Gillanders, and Terence I. Walker. Sources of variation in the feeding ecology of the piked spurdog (Squalus megalops): implications for inferring predator–prey interactions from overall dietary composition. ICES Journal of Marine Science, 62(6):1076–1094, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/6/1076/616640.

Burgos:2007:SAP

[BH07] Julian M. Burgos and John K. Horne. Sensitivity analysis and parameter selection for detecting aggregations in acoustic data. ICES Journal of Marine Science, 64(1):160–168, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/1/160/647013.

Burgos:2008:CCA

[BH08] Julian M. Burgos and John K. Horne. Characterization and classification of acoustically detected fish spatial distributions. ICES Journal of Marine Science, 65(7):1235–1247, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/7/1235/642470.

Barry:2008:BWR

[BHH^{+}08] Simon C. Barry, Keith R. Hayes, Chad L. Hewitt, Hanna L. Behrens, Egil Dragsund, and Siri M. Bakke. Ballast water risk assessment: principles, processes, and methods. ICES Journal of Marine Science, 65(2):121–131, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/2/121/735809.

Bradshaw:2004:SDF

[BHM^{+}04] Corey J. A. Bradshaw, Jane Higgins, Kelvin J. Michael, Simon J. Wotherspoon, and Mark A. Hindell. At-sea distribution of female southern elephant seals relative to variation in ocean surface properties. ICES Journal of Marine Science, 61 (6):1014–1027, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/6/1014/676039.

Barot:2005:MNA

[BHMD05] S. Barot, M. Heino, M. J. Morgan, and U. Dieckmann. Maturation of Newfoundland American plaice (Hippoglossoides platessoides): long-term trends in maturation reaction norms despite low fishing mortality? ICES Journal of Marine Science, 62 (1):56–64, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/1/56/669062.

Bradshaw:2002:OSS

[BHMS02] C. J. A. Bradshaw, M. A. Hindell, K. J. Michael, and M. D. Sumner. The optimal spatial scale for the analysis of elephant seal foraging as determined by geo-location in relation to sea surface temperatures. ICES Journal of Marine Science, 59(4):770–781, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/4/770/676795.

Bekkevold:2006:GIG

[BHN06] Dorte Bekkevold, Michael M. Hansen, and Einar E. Nielsen. Genetic impact of gadoid culture on wild fish populations: predictions, lessons from salmonids, and possibilities for minimizing adverse effects. ICES Journal of Marine Science, 63(2):198–208, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/2/198/637159.

Beverton:2004:GML

[BHØ^{+}04] Raymond J. H. Beverton, Arvid Hylen, Ole-Johan Østvedt, John Alvsvaag, and Terence C. Iles. Growth, maturation, and longevity of maturation cohorts of Norwegian spring-spawning herring. ICES Journal of Marine Science, 61(2):165–175, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/2/165/619854.

Bolle:2005:DTE

[BHR^{+}05] Loes J. Bolle, Ewan Hunter, Adriaan D. Rijnsdorp, Martin A. Pastoors, Julian D. Metcalfe, and John D. Reynolds. Do tagging experiments tell the truth? Using electronic tags to evaluate conventional tagging data. ICES Journal of Marine Science, 62(2):236–246, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/2/236/603589.

Bekkby:2008:MLS

[BI08] Trine Bekkby and Martin Isæus. Mapping large, shallow inlets and bays: modelling a Natura 2000 habitat with digital terrain and wave-exposure models. ICES Journal of Marine Science, 65(2):238–241, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/2/238/736559.

Bianchi:2000:IFS

[Bia00] G. Bianchi. Impact of fishing on size composition and diversity of demersal fish communities. ICES Journal of Marine Science, 57(3):558–571, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/558/635932.

Bernal:2008:UMM

[BIdL^{+}08] Miguel Bernal, Leire Ibaibarriaga, Ana Lago de Lanzós, Mike E. Lonergan, Carmen Hernández, Concha Franco, Inmaculada Rasines, Luis Valdés, and David L. Borchers. Using multinomial models to analyse data from Iberian sardine egg incubation experiments: a comparison with traditional techniques. ICES Journal of Marine Science, 65(1):51–59, January 2008. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/65/1/51/
611825.

Bertrand:2000:AELb

[BJ00a] A. Bertrand and E. Josse. Acoustic estimation of longline tuna abundance. ICES Journal of Marine Science, 57(5):1526, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/5/1526/660976.

Bertrand:2000:AELa

[BJ00b] Arnaud Bertrand and Erwan Josse. Acoustic estimation of longline tuna abundance. ICES Journal of Marine Science, 57(4): 919–926, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/4/919/647377.

Bertrand:2000:TTS

[BJ00c] Arnaud Bertrand and Erwan Josse. Tuna target-strength related to fish length and swimbladder volume. ICES Journal of Marine Science, 57(4):1143–1146, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/4/1143/647285.

Beamish:2006:EMS

[BJN^{+}06] Richard J. Beamish, Simon Jones, Chrys-Ellen Neville, Ruston Sweeting, Grace Karreman, Sonja Saksida, and Elysha Gordon. Exceptional marine survival of pink salmon that entered the marine environment in 2003 suggests that farmed Atlantic salmon and Pacific salmon can coexist successfully in a marine ecosystem on the Pacific coast of Canada. ICES Journal of Marine Science, 63(7):1326–1337, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/7/1326/759673.

Bjornsson:2002:EAF

[Bjö02] Björn Björnsson. Effects of anthropogenic feeding on the growth rate, nutritional status and migratory behaviour of free-ranging cod in an Icelandic fjord. ICES Journal of Marine Science, 59 (6):1248–1255, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/6/1248/608182.

Brophy:2007:LOG

[BK07] Deirdre Brophy and Pauline A. King. Larval otolith growth histories show evidence of stock structure in Northeast Atlantic blue whiting (Micromesistius poutassou). ICES Journal of Marine Science, 64(6):1136–1144, September 2007. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/6/1136/613621.

Bahou:2007:FCF

[BKN^{+}07] Laurent Bahou, Tidiani Koné, Valentin N’Douba, Kouassi J. N’Guessan, Essetchi P. Kouamélan, and Gooré B. Gouli. Food composition and feeding habits of little tunny (Euthynnus alletteratus) in continental shelf waters of Côte d’Ivoire (West Africa). ICES Journal of Marine Science, 64(5):1044–1052, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/5/1044/641726.

Bogaards:2009:BSB

[BKR09] Johannes A. Bogaards, Sarah B. M. Kraak, and Adriaan D. Rijnsdorp. Bayesian survey-based assessment of North Sea plaice (Pleuronectes platessa): extracting integrated signals from multiple surveys. ICES Journal of Marine Science, 66(4):665–679, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/4/665/604219.

Blaber:2000:EFS

[Bla00] S. Blaber. Effects of fishing on the structure and functioning of estuarine and nearshore ecosystems. ICES Journal of Marine Science, 57(3):590–602, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/3/590/635938.

Blanchard:2001:EFD

[Bla01] F. Blanchard. The effect of fishing on demersal fish community dynamics: an hypothesis. ICES Journal of Marine Science, 58 (3):711–718, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/3/711/810136.

Borja:2006:RBW

[BLMB06] Ángel Borja, Pedro Liria, Iñigo Muxika, and Juan Bald. Relationships between wave exposure and biomass of the goose barnacle (Pollicipes pollicipes, Gmelin, 1790) in the Gaztelugatxe Marine Reserve (Basque Country, northern Spain). ICES Journal of Marine Science, 63(4):626–636, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/4/626/692108.

Boyd:2005:EMS

[BLRC05] S. E. Boyd, D. S. Limpenny, H. L. Rees, and K. M. Cooper. The effects of marine sand and gravel extraction on the macrobenthos at a commercial dredging site (results 6 years post-dredging). ICES Journal of Marine Science, 62(2):145–162, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/2/145/602082.

Bergmann:2001:MAR

[BM01a] M. Bergmann and P. G. Moore. Mortality of Asterias rubens and Ophiura ophiura discarded in the Nephrops fishery of the Clyde Sea area, Scotland. ICES Journal of Marine Science, 58 (3):531–542, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/3/531/810055.

Bergmann:2001:SDC

[BM01b] M. Bergmann and P. G. Moore. Survival of decapod crustaceans discarded in the Nephrops fishery of the Clyde Sea area, Scotland. ICES Journal of Marine Science, 58(1):163–171, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/1/163/603510.

Basu:2002:EDB

[BM02] Atanu Basu and Shivani Malhotra. Error detection of bathymetry data by visualization using GIS. ICES Journal of Marine Science, 59(1):226–234, ???? 2002. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/1/226/650019.

Broadhurst:2009:SMC

[BM09] Matt K. Broadhurst and Russell B. Millar. Square-mesh codend circumference and selectivity. ICES Journal of Marine Science, 66(3):566–572, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/3/566/814996.

Bevacqua:2007:MOA

[BMC^{+}07] Daniele Bevacqua, Paco Melià, Alain J. Crivelli, Marino Gatto, and Giulio A. De Leo. Multi-objective assessment of conservation measures for the European eel (Anguilla anguilla): an application to the Camargue lagoons. ICES Journal of Marine Science, 64(7):1483–1490, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/7/1483/728692.

Buhl-Mortensen:2009:PBB

[BMDBM09] Pål Buhl-Mortensen, Margaret Dolan, and Lene Buhl-Mortensen. Prediction of benthic biotopes on a Norwegian offshore bank using a combination of multivariate analysis and GIS classification. ICES Journal of Marine Science, 66(9):2026–2032, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/9/2026/726975.

Blanchard:2008:PMS

[BMJ08] Julia L. Blanchard, David L. Maxwell, and Simon Jennings. Power of monitoring surveys to detect abundance trends in depleted populations: the effects of density-dependent habitat use, patchiness, and climate change. ICES Journal of Marine Science, 65(1):111–120, January 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/1/111/613983.

Brown:2005:MSH

[BML^{+}05] Craig J. Brown, Annika Mitchell, David S. Limpenny, Mike R. Robertson, Matthew Service, and Neil Golding. Mapping seabed habitats in the Firth of Lorn off the west coast of Scotland: evaluation and comparison of habitat maps produced using the acoustic ground-discrimination system, RoxAnn, and sidescan sonar. ICES Journal of Marine Science, 62(4):790–802, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/4/790/676525.

Brickman:2007:DPI

[BMLH07] David Brickman, Gudrun Marteinsdottir, Kai Logemann, and Ingo H. Harms. Drift probabilities for Icelandic cod larvae. ICES Journal of Marine Science, 64(1):49–59, January 2007. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/64/1/49/
649193.

Bowen:2003:SEP

[BMM03] W. D. Bowen, J. McMillan, and R. Mohn. Sustained exponential population growth of grey seals at Sable Island, Nova Scotia. ICES Journal of Marine Science, 60(6):1265–1274, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/6/1265/653081.

Barrento:2008:TLC

[BMP^{+}08] Sara Barrento, António Marques, Sónia Pedro, Paulo Vaz-Pires, and Maria Leonor Nunes. The trade of live crustaceans in Portugal: space for technological improvements. ICES Journal of Marine Science, 65(4):551–559, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/4/551/639240.

Broadhurst:2009:UDC

[BMU09] Matt K. Broadhurst, Russell B. Millar, and Sebastian S. Uhlmann. Using a double codend to reduce discard mortality. ICES Journal of Marine Science, 66(10):2077–2081, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/10/2077/678090.

Bagliniere:2005:ICR

[BMV05] Jean-Luc Baglinière, Frédéric Marchand, and Vincent Vauclin. Interannual changes in recruitment of the Atlantic salmon (Salmo salar) population in the River Oir (Lower Normandy, France): relationships with spawners and in-stream habitat.

ICES Journal of Marine Science, 62(4):695–707, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/4/695/674479.

Beare:2005:USD

[BNBR05] D. J. Beare, C. L. Needle, F. Burns, and D. G. Reid. Using survey data independently from commercial data in stock assessment: an example using haddock in ICES division VIa. ICES Journal of Marine Science, 62(5):996–1005, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/5/996/858616.

Burwen:2007:CNE

[BNF^{+}07] Debby L. Burwen, Patrick A. Nealson, Steven J. Fleischman, Timothy J. Mulligan, and John K. Horne. The complexity of narrowband echo envelopes as a function of fish sideaspect angle. ICES Journal of Marine Science, 64(5):1066–1074, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/5/1066/642614.

Brodziak:2005:DEF

[BO05] Jon Brodziak and Loretta O’Brien. Do environmental factors affect recruits per spawner anomalies of New England groundfish? ICES Journal of Marine Science, 62(7):1394–1407, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/7/1394/658865.

Bjornsson:2006:EWQ

[BÓ06] Björn Björnsson and Sólveig R. Ólafsdóttir. Effects of water quality and stocking density on growth performance of juvenile cod (Gadus morhua L.). ICES Journal of Marine Science, 63 (2):326–334, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/2/326/640846.

Birch:2008:SBH

[BO08] Gavin F. Birch and Marco A. Olmos. Sediment-bound heavy metals as indicators of human influence and biological risk in coastal water bodies. ICES Journal of Marine Science, 65(8):

1407–1413, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/8/1407/714053.

Bartolino:2008:BPJ

[BOC^{+}08] Valerio Bartolino, Alessandro Ottavi, Francesco Colloca, Gian Domenico Ardizzone, and Gunnar Stefánsson. Bathymetric preferences of juvenile European hake (Merluccius merluccius). ICES Journal of Marine Science, 65(6):963–969, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/6/963/604089.

Booth:2000:ISC

[Boo00] Anthony J. Booth. Incorporating the spatial component of fisheries data into stock assessment models. ICES Journal of Marine Science, 57(4):858–865, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/4/858/647352.

Botsford:2001:PIR

[Bot01] Louis W. Botsford. Physical influences on recruitment to California Current invertebrate populations on multiple scales. ICES Journal of Marine Science, 58(5):1081–1091, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/5/1081/730832.

Boxaspen:2006:RBG

[Box06] Karin Boxaspen. A review of the biology and genetics of sea lice. ICES Journal of Marine Science, 63(7):1304–1316, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/7/1304/758933.

Brooks:2007:GCS

[BP07] Elizabeth N. Brooks and Joseph E. Powers. Generalized compensation in stock-recruit functions: properties and implications for management. ICES Journal of Marine Science, 64 (3):413–424, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/3/413/813453.

Byers:2008:GAF

[BP08] James E. Byers and James M. Pringle. Going against the flow: how marine invasions spread and persist in the face of advection. ICES Journal of Marine Science, 65(5):723–724, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/5/723/712712.

Baumann:2003:REH

[BPD^{+}03] Hannes Baumann, Pierre Pepin, Fraser J. M. Davidson, Fran Mowbray, Dietrich Schnack, and John F. Dower. Reconstruction of environmental histories to investigate patterns of larval radiated shanny (Ulvaria subbifurcata) growth and selective survival in a large bay of Newfoundland. ICES Journal of Marine Science, 60(2):243–258, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/60/2/243/624855.

Babcock:2005:PUS

[BPM^{+}05] Elizabeth A. Babcock, Ellen K. Pikitch, Murdoch K. McAllister, Panayiota Apostolaki, and Christine Santora. A perspective on the use of spatialized indicators for ecosystem-based fishery management through spatial zoning. ICES Journal of Marine Science, 62(3):469–476, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/3/469/662960.

Bacon:2009:EAL

[BPM^{+}09] P. J. Bacon, S. C. F. Palmer, J. C. MacLean, G. W. Smith, B. D. M. Whyte, W. S. C. Gurney, and A. F. Youngson. Empirical analyses of the length, weight, and condition of adult Atlantic salmon on return to the Scottish coast between 1963 and 2006. ICES Journal of Marine Science, 66(5):844–859, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/5/844/663955.

Berger:2009:MRU

[BPT09] Laurent Berger, Cyrille Poncelet, and Verena M. Trenkel. A method for reducing uncertainty in estimates of fish-school frequency response using data from multifrequency and multibeam echosounders. ICES Journal of Marine Science, 66(6):

1155–1161, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1155/695853.

Buelens:2009:KMD

[BPWS09] Bart Buelens, Tim Pauly, Raymond Williams, and Arthur Sale. Kernel methods for the detection and classification of fish schools in single-beam and multibeam acoustic data. ICES Journal of Marine Science, 66(6):1130–1135, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1130/689053.

Balguerias:2000:OSB

[BQHG00] E. Balguerías, M. E. Quintero, and C. L. Hernández-González. The origin of the Saharan Bank cephalopod fishery. ICES Journal of Marine Science, 57(1):15–23, February 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/57/1/15/
641132.

Bez:2000:RSS

[BR00] Nicolas Bez and Jacques Rivoirard. On the role of sea surface temperature on the spatial distribution of early stages of mackerel using inertiograms. ICES Journal of Marine Science, 57 (2):383–392, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/2/383/620463.

Beare:2002:IST

[BR02] D. J. Beare and D. G. Reid. Investigating spatio-temporal change in spawning activity by Atlantic mackerel between 1977 and 1998 using generalized additive models. ICES Journal of Marine Science, 59(4):711–724, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/4/711/676775.

Blindheim:2004:WMF

[BR04] Johan Blindheim and Francisco Rey. Water-mass formation and distribution in the Nordic Seas during the 1990s. ICES Journal of Marine Science, 61(5):846–863, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/5/846/868837.

Barry:2008:USD

[BR08a] Jon Barry and Hubert L. Rees. Use of simulated data as a tool for testing the performance of diversity indices in response to an organic enrichment event. ICES Journal of Marine Science, 65 (8):1456–1461, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/8/1456/713365.

Butterworth:2008:SCA

[BR08b] Doug S. Butterworth and Rebecca A. Rademeyer. Statistical catch-at-age analysis vs. ADAPT-VPA: the case of Gulf of Maine cod. ICES Journal of Marine Science, 65(9):1717–1732, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/9/1717/632337.

Brander:2005:CRS

[Bra05] Keith M. Brander. Cod recruitment is strongly affected by climate when stock biomass is low. ICES Journal of Marine Science, 62(3):339–343, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/62/3/339/659193.

Brander:2007:RGC

[Bra07] Keith M. Brander. The role of growth changes in the decline and recovery of North Atlantic cod stocks since 1970. ICES Journal of Marine Science, 64(2):211–217, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/2/211/2182221.

Boswell:2009:SES

[BRC09] Kevin M. Boswell, Brian M. Roth, and James H. Cowan, Jr. Simulating the effects of side-aspect fish orientation on acoustic biomass estimates. ICES Journal of Marine Science, 66 (6):1398–1403, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1398/693618.

Bekkby:2008:SPM

[BRE^{+}08] Trine Bekkby, Eli Rinde, Lars Erikstad, Vegar Bakkestuen, Oddvar Longva, Ole Christensen, Martin Isæus, and Pål Erik Isachsen. Spatial probability modelling of eelgrass (Zostera marina) distribution on the west coast of Norway. ICES Journal of Marine Science, 65(7):1093–1101, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/7/1093/643728.

Bekkby:2009:SPD

[BREB09] Trine Bekkby, Eli Rinde, Lars Erikstad, and Vegar Bakkestuen. Spatial predictive distribution modelling of the kelp species Laminaria hyperborea. ICES Journal of Marine Science, 66(10): 2106–2115, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/10/2106/681328.

Buschmann:2006:RIS

[BRHG^{+}06] Alejandro H. Buschmann, Verónica A. Riquelme, María C. Hernández-González, Daniel Varela, Jaime E. Jiménez, Luis A. Henríquez, Pedro A. Vergara, Ricardo Guíñez, and Luis Filún. A review of the impacts of salmonid farming on marine coastal ecosystems in the southeast Pacific. ICES Journal of Marine Science, 63(7):1338–1345, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/7/1338/760174.

Briggs:2002:AFE

[Bri02] R. Briggs. The application of fecundity estimates to determine the spawning stock biomass of Irish Sea Nephrops norvegicus (L.) using the annual larval production method. ICES Journal of Marine Science, 59(1):109–119, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/1/109/649966.

Brown:2002:RSC

[Bro02a] E. Brown. Remote sensing of capelin and other biological features in the North Pacific using lidar and video technology. ICES Journal of Marine Science, 59(5):1120–1130, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/5/1120/675181.

Brown:2002:LHD

[Bro02b] Evelyn D. Brown. Life history, distribution, and size structure of Pacific capelin in Prince William Sound and the northern Gulf of Alaska. ICES Journal of Marine Science, 59(5):983–996, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/5/983/675223.

Bromley:2003:UMS

[Bro03] P. J. Bromley. The use of market sampling to generate maturity ogives and to investigate growth, sexual dimorphism and reproductive strategy in central and south-western North Sea sole (Solea solea L.). ICES Journal of Marine Science, 60(1):52–65, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/1/52/611445.

Beare:2002:STP

[BRP02] D. J. Beare, D. G. Reid, and P. Petitgas. Spatio-temporal patterns in herring (Clupea harengus L.) school abundance and size in the northwest North Sea: modelling space–time dependencies to allow examination of the impact of local school abundance on school size. ICES Journal of Marine Science, 59(3):469–479, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/3/469/610817.

Bozzano:2002:FDC

[BS02] A. Bozzano and F. Sardà. Fishery discard consumption rate and scavenging activity in the northwestern Mediterranean Sea. ICES Journal of Marine Science, 59(1):15–28, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/59/1/15/
649983.

Benoit:2003:ALD

[BS03] Hugues P. Benoít and Douglas P. Swain. Accounting for lengthand depth-dependent diel variation in catchability of fish and invertebrates in an annual bottom-trawl survey. ICES Journal of Marine Science, 60(6):1298–1317, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/6/1298/653532.

Bahamon:2009:FDP

[BSA09] Nixon Bahamon, Francesc Sardà, and Jacopo Aguzzi. Fuzzy diel patterns in catchability of deep-water species on the continental margin. ICES Journal of Marine Science, 66(10):2211–2218, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/10/2211/680052.

Beukers-Stewart:2003:PDP

[BSMB03] B. D. Beukers-Stewart, M. W. J. Mosley, and A. R. Brand. Population dynamics and predictions in the Isle of Man fishery for the great scallop, Pecten maximus L. ICES Journal of Marine Science, 60(2):224–242, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/2/224/624458.

Bjornsson:2001:OTG

[BSO01] Björn Björnsson, Agnar Steinarsson, and Matthías Oddgeirsson. Optimal temperature for growth and feed conversion of immature cod (Gadus morhua L.). ICES Journal of Marine Science, 58(1):29–38, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/1/29/603531.

Bahamon:2007:PBI

[BSS07] Nixon Bahamon, Francesc Sardà, and Petri Suuronen. Potential benefits from improved selectivity in the northwest Mediterranean multispecies trawl fishery. ICES Journal of Marine Science, 64(4):757–760, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/64/4/757/642796.

Battaglia:2006:EEE

[BTR06] André Battaglia, Verena M. Trenkel, and Marie-Joëlle Rochet. Estimating end effects in trawl catches. ICES Journal of Marine Science, 63(5):956–959, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/63/5/956/664608.

Bucklin:2000:PGV

[Buc00] A. Bucklin. Population genetic variation of Calanus finmarchicus in Icelandic waters: preliminary evidence of genetic differences between Atlantic and Arctic populations. ICES Journal of Marine Science, 57(6):1592–1604, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/6/1592/768556.

Bunce:2001:PCA

[Bun01] Ashley Bunce. Prey consumption of Australasian gannets (Morus serrator) breeding in Port Phillip Bay, southeast Australia, and potential overlap with commercial fisheries. ICES Journal of Marine Science, 58(4):904–915, ???? 2001. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/4/904/630227.

Butler:2001:SSE

[But01] E. Butler. A survey strategy and environmental monitoring network for an estuary supporting finfish cage culture. ICES Journal of Marine Science, 58(2):460–468, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/2/460/617850.

Butterworth:2007:WMP

[But07] Doug S. Butterworth. Why a management procedure approach? Some positives and negatives. ICES Journal of Marine Science, 64(4):613–617, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/4/613/639583.

Breteler:2007:APS

[BVB^{+}07] Jan Klein Breteler, Tim Vriese, Jost Borcherding, André Breukelaar, Lothar Jörgensen, Stefan Staas, Gerard de Laak, and Detlev Ingendahl. Assessment of population size and migration routes of silver eel in the River Rhine based on a 2-year combined mark-recapture and telemetry study. ICES Journal of Marine Science, 64(7):1450–1456, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/7/1450/728959.

Bulgakova:2001:WSS

[BVD01] Tatiana Bulgakova, Dmitri Vasilyev, and Niels Daan. Weighting and smoothing of stomach content data as input for MSVPA with particular reference to the Barents Sea. ICES Journal of Marine Science, 58(6):1208–1218, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/6/1208/641556.

Bishop:2008:SCR

[BVDS08] J. Bishop, W. N. Venables, C. M. Dichmont, and D. J. Sterling. Standardizing catch rates: is logbook information by itself enough? ICES Journal of Marine Science, 65(2):255–266, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/2/255/733342.

Borges:2008:WDP

[BvKvH^{+}08] Lisa Borges, Olvin A. van Keeken, Aloysius T. M. van Helmond, Bram Couperus, and Mark Dickey-Collas. What do pelagic freezer-trawlers discard? ICES Journal of Marine Science, 65 (4):605–611, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/4/605/640309.

Bergman:2000:MMB

[BvS00] Magda J. N. Bergman and Jan W. van Santbrink. Mortality in megafaunal benthic populations caused by trawl fisheries on the Dutch continental shelf in the North Sea in 1994. ICES Journal of Marine Science, 57(5):1321–1331, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/5/1321/660928.

Boswell:2008:SAT

[BW08] Kevin M. Boswell and Charles A. Wilson. Side-aspect targetstrength measurements of bay anchovy (Anchoa mitchilli) and Gulf menhaden (Brevoortia patronus) derived from ex situ experiments. ICES Journal of Marine Science, 65(6):1012– 1020, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/6/1012/601935.

Brodeur:2000:STV

[BWC00] Richard D. Brodeur, Matthew T. Wilson, and Lorenzo Ciannelli. Spatial and temporal variability in feeding and condition of age-0 walleye pollock (Theragra chalcogramma) in frontal regions of the Bering Sea. ICES Journal of Marine Science, 57(2):256–264, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/2/256/620422.

Boulcott:2007:RVM

[BWG^{+}07] Philip Boulcott, Peter J. Wright, Fiona M. Gibb, Henrik Jensen, and Iain M. Gibb. Regional variation in maturation of sandeels in the North Sea. ICES Journal of Marine Science, 64(2): 369–376, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/2/369/2182485.

Bark:2007:CST

[BWK07] A. Bark, B. Williams, and B. Knights. Current status and temporal trends in stocks of European eel in England and Wales. ICES Journal of Marine Science, 64(7):1368–1378, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/7/1368/727672.

Borges:2006:MDO

[BZRO06] Lisa Borges, Alain F. Zuur, Emer Rogan, and Rick Officer. Modelling discard ogives from Irish demersal fisheries. ICES Journal of Marine Science, 63(6):1086–1095, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/6/1086/616094.

Cardinale:2000:DWA

[CA00] M. Cardinale and F. Arrhenius. Decreasing weight-at-age of Atlantic herring (Clupea harengus) from the Baltic Sea between 1986 and 1996: a statistical analysis. ICES Journal of Marine Science, 57(4):882–893, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/4/882/647360.

Cripps:2002:ESE

[CA02] S. J. Cripps and J. P. Aabel. Environmental and socioeconomic impact assessment of Ekoreef, a multiple platform rigs-to-reefs development. ICES Journal of Marine Science, 59 (S1):S300–S308, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S300/617974.

Ceyhan:2007:AGR

[CAAJ07] Tevfik Ceyhan, Okan Akyol, Adnan Ayaz, and Francis Juanes. Age, growth, and reproductive season of bluefish (Pomatomus saltatrix) in the Marmara region, Turkey. ICES Journal of Marine Science, 64(3):531–536, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/3/531/818349.

Caddy:2000:MCB

[Cad00] J. F. Caddy. Marine catchment basin effects versus impacts of fisheries on semi-enclosed seas. ICES Journal of Marine Science, 57(3):628–640, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/628/635946.

Castillo:2005:PSM

[CAGV05] Ana G. F. Castillo, Paula Alvarez, and Eva Garcia-Vazquez. Population structure of Merluccius merluccius along the Iberian Peninsula coast. ICES Journal of Marine Science, 62(8): 1699–1704, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/8/1699/796293.

Callaway:2002:DCS

[Cal02] R. Callaway. Diversity and community structure of epibenthic invertebrates and fish in the North Sea. ICES Journal of Marine Science, 59(6):1199–1214, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/6/1199/608176.

Calbet:2008:TRM

[Cal08] Albert Calbet. The trophic roles of microzooplankton in marine systems. ICES Journal of Marine Science, 65(3):325–331, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/3/325/786057.

Campbell:2008:OIA

[Cam08] Marnie L. Campbell. Organism impact assessment: risk analysis for post-incursion management. ICES Journal of Marine Science, 65(5):795–804, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/5/795/714920.

Castro:2007:ECS

[Cas07] B. G. Castro. Element composition of sardine (Sardina pilchardus) otoliths along the Atlantic Coast of the Iberian Peninsula. ICES Journal of Marine Science, 64(3):512–518, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/3/512/816564.

Campbell:2009:ILB

[CAWD09] Neil Campbell, Lynda Allan, Adrian Weetman, and Helen Dobby. Investigating the link between Nephrops norvegicus burrow density and sediment composition in Scottish waters. ICES Journal of Marine Science, 66(9):2052–2059, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/9/2052/726193.

Coombs:2007:SMO

[CB07] Roger F. Coombs and Richard Barr. In situ measurements of orange roughy (Hoplostethus atlanticus) target strength. ICES Journal of Marine Science, 64(6):1220–1234, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/6/1220/614533.

Cadiou:2009:MAF

[CBBL09] Gwenaël Cadiou, Charles F. Boudouresque, Patrick Bonhomme, and Laurence Le Diréach. The management of artisanal fishing within the Marine Protected Area of the Port-Cros National Park (northwest Mediterranean Sea): a success story? ICES Journal of Marine Science, 66(1):41–49, January 2009. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/66/1/41/
633981.

Carlstrom:2002:FEU

[CBDB02] Julia Carlström, Per Berggren, Felicia Dinnétz, and Patrik Börjesson. A field experiment using acoustic alarms (pingers) to reduce harbour porpoise by-catch in bottom-set gillnets. ICES Journal of Marine Science, 59(4):816–824, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/4/816/676809.

Cordell:2008:ACM

[CBDS08] Jeffery R. Cordell, Stephen M. Bollens, Robyn Draheim, and Mark Sytsma. Asian copepods on the move: recent invasions in the Columbia–Snake River system, USA. ICES Journal of Marine Science, 65(5):753–758, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/5/753/712312.

Clausen:2007:AVO

[CBHM07] L. A. W. Clausen, D. Bekkevold, E. M. C. Hatfield, and H. Mosegaard. Application and validation of otolith microstructure as a stock identification method in mixed Atlantic herring (Clupea harengus) stocks in the North Sea and western Baltic. ICES Journal of Marine Science, 64(2):377–385, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/2/377/2182523.

Chittenden:2009:CRP

[CBM09] Cedar M. Chittenden, Richard J. Beamish, and R. Scott McKinley. A critical review of Pacific salmon marine research relating to climate. ICES Journal of Marine Science, 66(10): 2195–2204, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/10/2195/678223.

Cook:2006:ICM

[CBS^{+}06] E. J. Cook, K. D. Black, M. D. J. Sayer, C. J. Cromey, D. L. Angel, E. Spanier, A. Tsemel, T. Katz, N. Eden, I. Karakassis, M. Tsapakis, E. T. Apostolaki, and A. Malej. The influence of caged mariculture on the early development of sublittoral fouling communities: a pan-European study. ICES Journal of Marine Science, 63(4):637–649, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/4/637/692173.

Cury:2005:QEI

[CC05] Philippe M. Cury and Villy Christensen. Quantitative ecosystem indicators for fisheries management. ICES Journal of Marine Science, 62(3):307–310, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/62/3/307/658706.

Casini:2004:FPH

[CCA04] Michele Casini, Massimiliano Cardinale, and Fredrik Arrhenius. Feeding preferences of herring (Clupea harengus) and sprat (Sprattus sprattus) in the southern Baltic Sea. ICES Journal of Marine Science, 61(8):1267–1277, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/8/1267/629668.

Carrera:2006:CAL

[CCB^{+}06] P. Carrera, J. H. Churnside, G. Boyra, V. Marques, C. Scalabrin, and A. Uriarte. Comparison of airborne lidar with echosounders: a case study in the coastal Atlantic waters of southern Europe. ICES Journal of Marine Science, 63(9):1736–1750, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/9/1736/701078.

Comeau:2002:TAC

[CCC02] L. A. Comeau, S. E. Campana, and G. A. Chouinard. Timing of Atlantic cod (Gadus morhua L.) seasonal migrations in the southern Gulf of St Lawrence: interannual variability and proximate control. ICES Journal of Marine Science, 59(2):333–351, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/2/333/619646.

Casini:2005:TCR

[CCHV05] Michele Casini, Massimiliano Cardinale, Joakim Hjelm, and Francesca Vitale. Trends in cpue and related changes in spatial distribution of demersal fish species in the Kattegat and Skagerrak, eastern North Sea, between 1981 and 2003. ICES Journal of Marine Science, 62(4):671–682, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/4/671/673774.

Conti:2003:WBA

[CD03] Stéphane G. Conti and David A. Demer. Wide-bandwidth acoustical characterization of anchovy and sardine from reverberation measurements in an echoic tank. ICES Journal of Marine Science, 60(3):617–624, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/3/617/660099.

Conti:2006:IPS

[CD06] Stéphane G. Conti and David A. Demer. Improved parameterization of the SDWBA for estimating krill target strength. ICES Journal of Marine Science, 63(5):928–935, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/5/928/664115.

Cutter:2007:ASD

[CD07] George R. Cutter, Jr. and David A. Demer. Accounting for scattering directivity and fish behaviour in multibeam-echosounder surveys. ICES Journal of Marine Science, 64(9):1664–1674, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/9/1664/784270.

Churnside:2009:TSL

[CD09] James H. Churnside and Percy L. Donaghay. Thin scattering layers observed by airborne lidar. ICES Journal of Marine Science, 66(4):778–789, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/4/778/603126.

Conti:2005:BBS

[CDB05] Stéphane G. Conti, David A. Demer, and Andrew S. Brierley. Broad-bandwidth, sound scattering, and absorption from krill (Meganyctiphanes norvegica), mysids (Praunus flexuosus and Neomysis integer), and shrimp (Crangon crangon). ICES Journal of Marine Science, 62(5):956–965, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/5/956/858202.

Campbell:2009:IMU

[CDB09] N. Campbell, H. Dobby, and N. Bailey. Investigating and mitigating uncertainties in the assessment of Scottish Nephrops norvegicus populations using simulated underwater television data. ICES Journal of Marine Science, 66(4):646–655, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/4/646/604278.

Crosnier:2008:TOS

[CDBS08] L. Crosnier, M. Drévillon, S. Ramos Buarque, and F. Soulat. Three ocean state indices implemented in the Mercator–Ocean operational suite. ICES Journal of Marine Science, 65(8):1504– 1507, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/8/1504/713663.

Crawford:2007:TNC

[CDD^{+}07] Robert J. M. Crawford, Benedict L. Dundee, Bruce M. Dyer, Norbert T. W. Klages, Michael A. Meÿer, and Leshia Upfold. Trends in numbers of Cape gannets (Morus capensis), 1956/ 1957–2005/2006, with a consideration of the influence of food and other factors. ICES Journal of Marine Science, 64(1): 169–177, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/1/169/647775.

Cranfield:2005:BEE

[CDDM05] H. J. Cranfield, A. Dunn, I. J. Doonan, and K. P. Michael. Bonamia exitiosa epizootic in Ostrea chilensis from Foveaux Strait, southern New Zealand between 1986 and 1992. ICES Journal of Marine Science, 62(1):3–13, ???? 2005. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/1/3/667406.

Constable:2000:MFC

[CdlMA^{+}00] Andrew J. Constable, William K. de la Mare, David J. Agnew, Inigo Everson, and Denzil Miller. Managing fisheries to conserve the Antarctic marine ecosystem: practical implementation of the Convention on the Conservation of Antarctic Marine Living Resources (CCAMLR). ICES Journal of Marine Science, 57 (3):778–791, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/778/636012.

Churnside:2003:CLE

[CDM03] James H. Churnside, David A. Demer, and Behzad Mahmoudi. A comparison of lidar and echosounder measurements of fish schools in the Gulf of Mexico. ICES Journal of Marine Science, 60(1):147–154, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/1/147/611429.

Charrier:2006:IPG

[CDQL06] Grégory Charrier, Jean-Dominique Durand, Louis Quiniou, and Jean Laroche. An investigation of the population genetic structure of pollack (Pollachius pollachius) based on microsatellite markers. ICES Journal of Marine Science, 63(9):1705–1709, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/9/1705/699810.

Challier:2005:TAR

[CDR05] Laurence Challier, Matthew R. Dunn, and Jean-Paul Robin. Trends in age-at-recruitment and juvenile growth of cuttlefish, Sepia officinalis, from the English Channel. ICES Journal of Marine Science, 62(8):1671–1682, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/8/1671/794116.

Conti:2005:IMF

[CDSC05] Stéphane G. Conti, David A. Demer, Michael A. Soule, and Jean H. E. Conti. An improved multiple-frequency method for measuring in situ target strengths. ICES Journal of Marine Science, 62(8):1636–1646, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/8/1636/792372.

Castro:2003:IMN

[CEH03] M. Castro, P. Encarnação, and P. Henriques. Increment at molt for the Norway lobster (Nephrops norvegicus) from the south coast of Portugal. ICES Journal of Marine Science, 60 (5):1159–1164, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/5/1159/768692.

Collie:2000:PEI

[CEV00] Jeremy S. Collie, Galo A. Escanero, and Page C. Valentine. Photographic evaluation of the impacts of bottom fishing on benthic epifauna. ICES Journal of Marine Science, 57(4):987–1001, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/987/647405.

Carscadden:2002:TVC

[CF02] J. E. Carscadden and K. T. Frank. Temporal variability in the condition factors of Newfoundland capelin (Mallotus villosus) during the past two decades. ICES Journal of Marine Science, 59(5):950–958, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/5/950/675210.

Cadigan:2005:LID

[CF05] Noel G. Cadigan and Patrick J. Farrell. Local influence diagnostics for the retrospective problem in sequential population analysis. ICES Journal of Marine Science, 62(2):256–265, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/2/256/603899.

Corgos:2006:MGM

[CF06] Antonio Corgos and Juan Freire. Morphometric and gonad maturity in the spider crab Maja brachydactyla: a comparison of methods for estimating size at maturity in species with determinate growth. ICES Journal of Marine Science, 63(5):851–859, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/5/851/662699.

Carscadden:2000:EER

[CFL00] J. E. Carscadden, K. T. Frank, and W. C. Leggett. Evaluation of an environment–recruitment model for capelin (Mallotus villosus). ICES Journal of Marine Science, 57(2):412–418, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/2/412/620475.

Courbin:2007:HOM

[CFMdP07] Nicolas Courbin, Ronan Fablet, Capucine Mellon, and Hélène de Pontual. Are hake otolith macrostructures randomly deposited? Insights from an unsupervised statistical and quantitative approach applied to Mediterranean hake otoliths. ICES Journal of Marine Science, 64(6):1191–1201, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/6/1191/614419.

Clark:2003:LTP

[CFN03] Robin A. Clark, Chris L. J. Frid, and Kirsty R. Nicholas. Longterm, predation-based control of a central-west North Sea zooplankton community. ICES Journal of Marine Science, 60 (2):187–197, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/2/187/623642.

Chamberlain:2001:IBS

[CFR^{+}01] J. Chamberlain, T. F. Fernandes, P. Read, T. D. Nickell, and I. M. Davies. Impacts of biodeposits from suspended mussel (Mytilus edulis L.) culture on the surrounding surficial sediments. ICES Journal of Marine Science, 58(2):411–416, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/2/411/617839.

Cabranes:2008:GSO

[CFRM08] C. Cabranes, P. Fernandez-Rueda, and J. L. Martínez. Genetic structure of Octopus vulgaris around the Iberian Peninsula and Canary Islands as indicated by microsatellite DNA variation. ICES Journal of Marine Science, 65(1):12–16, January 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/1/12/613733.

Campbell:2007:ARE

[CG07] Marnie L. Campbell and Charmaine Gallagher. Assessing the relative effects of fishing on the New Zealand marine environment through risk analysis. ICES Journal of Marine Science, 64(2):256–270, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/2/256/2182455.

Christiansen:2004:GPF

[CGN^{+}04] Jørgen S. Christiansen, Asbjørn Gildberg, Kjell T. Nilssen, Charlotta Lindblom, and Tore Haug. The gastric properties of free-ranging harp (Pagophilus groenlandicus (Erxleben, 1777)) and hooded (Cystophora cristata (Erxleben, 1777)) seals. ICES Journal of Marine Science, 61(2):287–292, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/2/287/622215.

Collie:2009:EEF

[CGS09] Jeremy S. Collie, Dian J. Gifford, and John H. Steele. End-to-end foodweb control of fish production on Georges Bank. ICES Journal of Marine Science, 66(10):2223–2232, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/10/2223/679266.

Collie:2003:UAD

[CGV03] Jeremy S. Collie, Henrik Gislason, and Morten Vinther. Using AMOEBAs to display multispecies, multifleet fisheries advice. ICES Journal of Marine Science, 60(4):709–720, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/4/709/690679.

Campbell:2000:VCF

[CH00] R. W. Campbell and E. J. H. Head. Viability of Calanus finmarchicus eggs in situ: does the presence of intact phytoplankton reduce hatching success? ICES Journal of Marine Science, 57(6):1780–1785, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/6/1780/768664.

Cook:2005:IWC

[CH05] R. M. Cook and M. R. Heath. The implications of warming climate for the management of North Sea demersal fisheries. ICES Journal of Marine Science, 62(7):1322–1326, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/7/1322/656742.

Chambers:2006:PIC

[CH06] Michael D. Chambers and William H. Howell. Preliminary information on cod and haddock production in submerged cages off the coast of New Hampshire, USA. ICES Journal of Marine Science, 63(2):385–392, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/63/2/385/642047.

Cannaby:2009:ILF

[CH09] Heather Cannaby and Y. Sinan Hüsrevoğlu. The influence of low-frequency variability and long-term trends in North Atlantic sea surface temperature on Irish waters. ICES Journal of Marine Science, 66(7):1480–1489, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/7/1480/657991.

Chaput:2004:CUS

[Cha04] Gerald Chaput. Considerations for using spawner reference levels for managing single- and mixed-stock fisheries of Atlantic salmon. ICES Journal of Marine Science, 61(8):1379–1388, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/8/1379/632161.

Clemente:2009:ETR

[CHB09] Sabrina Clemente, José Carlos Hernández, and Alberto Brito. Evidence of the top–down role of predators in structuring sublittoral rocky-reef communities in a marine protected area and nearby areas of the Canary Islands. ICES Journal of Marine Science, 66(1):64–71, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/1/64/633137.

Christensen:2002:KPM

[Chr02] Jørgen Møller Christensen. Kristian Popp Madsen 28 April 1926–20 May 2001. ICES Journal of Marine Science, 59(1): 236–237, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/1/236/650023.

Cipriano:2009:AAI

[Cip09] Rocco C. Cipriano. Antibody against infectious salmon anaemia virus among feral Atlantic salmon (Salmo salar). ICES Journal of Marine Science, 66(5):865–870, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/5/865/664501.

Collins:2002:EIA

[CJM^{+}02] K. J. Collins, A. C. Jensen, J. J. Mallinson, V. Roenelle, and I. P. Smith. Environmental impact assessment of a scrap tyre artificial reef. ICES Journal of Marine Science, 59(S1):S243–S249, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL https://academic.oup.com/icesjms/
article/59/suppl/S243/617949.

Collins:2002:OFE

[CJS02] K. J. Collins, A. C. Jensen, and I. P. Smith. Oxygen fluxes of enclosed reef epibiota communities. ICES Journal of Marine Science, 59(S1):S94–S99, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/59/suppl/S94/618015.

Cooke:2003:AOA

[CKS03] K. Cooke, R. Kieser, and R. D. Stanley. Acoustic observation and assessment of fish in high-relief habitats. ICES Journal of Marine Science, 60(3):658–661, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/3/658/660610.

Claytor:2000:CRF

[Cla00] Ross R. Claytor. Conflict resolution in fisheries management using decision rules: an example using a mixed-stock Atlantic Canadian herring fishery. ICES Journal of Marine Science, 57 (4):1110–1127, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/4/1110/647278.

Caselle:2002:THF

[CLFS02] J. E. Caselle, M. S. Love, C. Fusaro, and D. Schroeder. Trash or habitat? Fish assemblages on offshore oilfield seafloor debris in the Santa Barbara Channel, California. ICES Journal of Marine Science, 59(S1):S258–S265, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S258/617955.

Chlaida:2009:EGC

[CLK^{+}09] M. Chlaida, V. Laurent, S. Kifani, T. Benazzou, H. Jaziri, and S. Planes. Evidence of a genetic cline for Sardina pilchardus along the Northwest African coast. ICES Journal of Marine Science, 66(2):264–271, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/2/264/594826.

Chang:2009:HCV

[CLL^{+}09] Shui-Kai Chang, Ta-Te Lin, Ghung-Hui Lin, Hsiang-Yun Chang, and Ching-Lu Hsieh. How to collect verifiable length data on tuna from photographs: an approach for sample vessels. ICES Journal of Marine Science, 66(5):907–915, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/5/907/664151.

Colonello:2007:RAA

[CLM07] Jorge H. Colonello, Luis O. Lucifora, and Ana M. Massa. Reproduction of the angular angel shark (Squatina guggenheim): geographic differences, reproductive cycle, and sexual dimorphism. ICES Journal of Marine Science, 64(1):131–140, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/1/131/646282.

Chaput:2005:PCA

[CLR^{+}05] G. Chaput, C. M. Legault, D. G. Reddin, F. Caron, and P. G. Amiro. Provision of catch advice taking account of nonstationarity in productivity of Atlantic salmon (Salmo salar L.) in the Northwest Atlantic. ICES Journal of Marine Science, 62(1):131–143, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/1/131/666998.

Crain:2000:DSS

[CM00] J. A. Crain and C. B. Miller. Detection of sex and sex ratio in Calanus finmarchicus early stage fifth copepodites. ICES Journal of Marine Science, 57(6):1773–1779, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/6/1773/768657.

Cabreira:2006:ASR

[CMC^{+}06a] A. G. Cabreira, A. Madirolas, G. Alvarez Colombo, E. M. Acha, and H. W. Mianzan. Acoustic study of the Río de la Plata estuarine front. ICES Journal of Marine Science, 63(9):1718–1725, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/9/1718/700601.

Carlsson:2006:GHA

[CMC^{+}06b] Jens Carlsson, Jan R. McDowell, Jeanette E. L. Carlsson, Droplaug Ólafsdóttir, and John E. Graves. Genetic heterogeneity of Atlantic bluefin tuna caught in the eastern North Atlantic Ocean south of Iceland. ICES Journal of Marine Science, 63 (6):1111–1117, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/6/1111/616553.

Carscadden:2002:TRA

[CMDN02] J. E. Carscadden, W. A. Montevecchi, G. K. Davoren, and B. S. Nakashima. Trophic relationships among capelin (Mallotus villosus) and seabirds in a changing ecosystem. ICES Journal of Marine Science, 59(5):1027–1033, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/5/1027/675131.

Cury:2005:VTE

[CMGS05] Philippe M. Cury, Christian Mullon, Serge M. Garcia, and Lynne J. Shannon. Viability theory for an ecosystem approach to fisheries. ICES Journal of Marine Science, 62(3):577–584, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/3/577/666438.

Christiansen:2005:DCM

[CMHN05] Jørgen S. Christiansen, Anne-Grethe Gamst Moen, Thomas H. Hansen, and Kjell T. Nilssen. Digestion of capelin, Mallotus villosus (Müller), herring, Clupea harengus L., and polar cod, Boreogadus saida (Lepechin), otoliths in a simulated seal stomach. ICES Journal of Marine Science, 62(1):86–92, ???? 2005. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/62/1/86/
669860.

Christensen:2009:SRF

[CMJ09] Asbjørn Christensen, Henrik Mosegaard, and Henrik Jensen. Spatially resolved fish population analysis for designing MPAs: influence on inside and neighbouring habitats. ICES Journal of Marine Science, 66(1):56–63, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/1/56/634296.

Clarke:2009:ELI

[CMK09] Shelley C. Clarke, Murdoch K. McAllister, and R. Craig Kirkpatrick. Estimating legal and illegal catches of Russian sockeye salmon from trade and market data. ICES Journal of Marine Science, 66(3):532–545, April 2009. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/66/3/532/816695.

Crawford:2001:VAE

[CMM01] C. M. Crawford, I. M. Mitchell, and C. K. A. Macleod. Video assessment of environmental impacts of salmon farms. ICES Journal of Marine Science, 58(2):445–452, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/2/445/617848.

Colombo:2003:ACG

[CMM03] Gustavo Alvarez Colombo, Hermes Mianzan, and Adrian Madirolas. Acoustic characterization of gelatinous plankton aggregations: four case studies from the Argentine continental shelf. ICES Journal of Marine Science, 60(3):650–657, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/3/650/660506.

Chakraborty:2007:ACS

[CMN^{+}07] Bishwajit Chakraborty, Vasudev Mahale, Gajanan Navelkar, B. Ramalingeswara Rao, R. G. Prabhudesai, Baban Ingole, and G. Janakiraman. Acoustic characterization of seafloor habitats on the western continental shelf of India. ICES Journal of Marine Science, 64(3):551–558, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/3/551/812734.

Coughlan:2006:TVI

[CMO^{+}06] James Coughlan, Philip McGinnity, Brian O’Farrell, Eileen Dillane, Ola Diserud, Elvira de Eyto, Killian Farrell, Ken Whelan, René J. M. Stet, and Thomas F. Cross. Temporal variation in an immune response gene (MHC I) in anadromous Salmo trutta in an Irish river before and during aquaculture activities. ICES Journal of Marine Science, 63(7):1248–1255, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/7/1248/756434.

Cotter:2007:ESP

[CMP07] A. J. R. Cotter, B. Mesnil, and G. J. Piet. Estimating stock parameters from trawl cpue-at-age series using year-class curves.

ICES Journal of Marine Science, 64(2):234–247, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/2/234/2182281.

Collins:2002:ASC

[Col02] M. Collins. Assessment of stone crab (Lithodidae) density on the South Georgia slope using baited video cameras. ICES Journal of Marine Science, 59(2):370–379, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/2/370/619656.

Cook:2004:EAS

[Coo04] R. M. Cook. Estimation of the age-specific rate of natural mortality for Shetland sandeels. ICES Journal of Marine Science, 61(2):159–164, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/2/159/619813.

Corfield:2000:EAS

[Cor00a] Jamie Corfield. The effects of acid sulphate run-off on a subtidal estuarine macrobenthic community in the Richmond River, NSW, Australia. ICES Journal of Marine Science, 57(5): 1517–1523, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/5/1517/660972.

Corten:2000:PAH

[Cor00b] A. Corten. A possible adaptation of herring feeding migrations to a change in timing of the Calanus finmarchicus season in the eastern North Sea. ICES Journal of Marine Science, 57(4): 1261–1270, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/4/1270/647318.

Cordue:2001:NIS

[Cor01] P. L. Cordue. A note on incorporating stochastic recruitment into deterministic age structured population models. ICES Journal of Marine Science, 58(4):794–798, ???? 2001. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/4/794/630203.

Cordue:2007:NNR

[Cor07] Patrick L. Cordue. A note on non-random error structure in trawl survey abundance indices. ICES Journal of Marine Science, 64(7):1333–1337, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/64/7/1333/729565.

Costa:2009:MVM

[Cos09] Ana Maria Costa. Macroscopic vs. microscopic identification of the maturity stages of female horse mackerel. ICES Journal of Marine Science, 66(3):509–516, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/3/509/812136.

Cotter:2001:INS

[Cot01] A. J. R. Cotter. Intercalibration of North Sea international bottom trawl surveys by fitting year-class curves. ICES Journal of Marine Science, 58(3):622–632, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/3/622/810099.

Chancollon:2006:FFE

[CPR06] Odile Chancollon, Claire Pusineri, and Vincent Ridoux. Food and feeding ecology of Northeast Atlantic swordfish (Xiphias gladius) off the Bay of Biscay. ICES Journal of Marine Science, 63(6):1075–1085, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/6/1075/615845.

Castillo:2004:SSG

[CR04] Jorge Castillo and Hugo Robotham. Spatial structure and geometry of schools of sardine (Sardinops sagax) in relation to abundance, fishing effort, and catch in northern Chile. ICES Journal of Marine Science, 61(7):1113–1119, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1113/878913.

Clarke:2008:BFU

[CRB08] Colleen S. L. Mercer Clarke, John C. Roff, and Shannon M. Bard. Back to the future: using landscape ecology to understand changing patterns of land use in Canada, and its effects on the sustainability of coastal ecosystems. ICES Journal of Marine Science, 65(8):1534–1539, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/8/1534/712599.

Cutter:2009:MTD

[CRC^{+}09] George R. Cutter, Josiah S. Renfree, Martin J. Cox, Andrew S. Brierley, and David A. Demer. Modelling three-dimensional directivity of sound scattering by Antarctic krill: progress towards biomass estimation using multibeam sonar. ICES Journal of Marine Science, 66(6):1245–1251, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1245/691358.

Catchpole:2008:EET

[CRIP08] Tom L. Catchpole, Andrew S. Revill, James Innes, and Sean Pascoe. Evaluating the efficacy of technical measures: a case study of selection device legislation in the UK Crangon crangon (brown shrimp) fishery. ICES Journal of Marine Science, 65 (2):267–275, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/2/267/737922.

Cuzin-Roudy:2004:LCS

[CRTS04] J. Cuzin-Roudy, G. A. Tarling, and J.-O. Strömberg. Life cycle strategies of northern krill (Meganyctiphanes norvegica) for regulating growth, moult, and reproductive activity in various environments: the case of fjordic populations. ICES Journal of Marine Science, 61(4):721–737, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/4/721/608235.

Certain:2008:DSD

[CRvCB08] G. Certain, V. Ridoux, O. van Canneyt, and V. Bretagnolle. Delphinid spatial distribution and abundance estimates over the shelf of the Bay of Biscay. ICES Journal of Marine Science, 65(4):656–666, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/4/656/639969.

Cherel:2001:CMV

[CRW^{+}01] Yves Cherel, Vincent Ridoux, Henri Weimerskirch, Torkild Tveraa, and Olivier Chastel. Capelin (Mallotus villosus) as an important food source for northern fulmars (Fulmarus glacialis) breeding at Bjørnøya (Bear Island), Barents Sea. ICES Journal of Marine Science, 58(1):355–361, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/1/355/603546.

Conners:2002:UAC

[CS02] M. Elizabeth Conners and Steven J. Schwager. The use of adaptive cluster sampling for hydroacoustic surveys. ICES Journal of Marine Science, 59(6):1314–1325, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/6/1314/608197.

Cadrin:2005:MVY

[CS05] Steven X. Cadrin and Vaughn M. Silva. Morphometric variation of yellowtail flounder. ICES Journal of Marine Science, 62(4):683–694, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/4/683/674104.

Crozier:2004:MAS

[CSC^{+}04] W. W. Crozier, P-J. Schön, G. Chaput, E. C. E. Potter, N.Ó Maoiléidigh, and J. C. MacLean. Managing Atlantic salmon (Salmo salar L.) in the mixed stock environment: challenges and considerations. ICES Journal of Marine Science, 61(8): 1344–1358, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/8/1344/631701.

Cutts:2006:DID

[CSdQB06] C. J. Cutts, J. Sawanboonchun, C. Mazorra de Quero, and J. G. Bell. Diet-induced differences in the essential fatty acid (EFA) compositions of larval Atlantic cod (Gadus morhua L.) with reference to possible effects of dietary EFAs on larval performance. ICES Journal of Marine Science, 63(2):302–310, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/2/302/640388.

Cochrane:2000:SSE

[CSH00] N. A. Cochrane, D. D. Sameoto, and A. W. Herman. Scotian Shelf euphausiid and silver hake population changes during 1984–1996 measured by multi-frequency acoustics. ICES Journal of Marine Science, 57(1):122–132, February 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/1/122/641128.

Charbonnel:2002:EIH

[CSR^{+}02] Eric Charbonnel, Christophe Serre, Sandrine Ruitton, Jean-Georges Harmelin, and Antony Jensen. Effects of increased habitat complexity on fish assemblages associated with large artificial reef units (French Mediterranean coast). ICES Journal of Marine Science, 59(S1):S208–S213, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S208/617939.

Coelho:2009:PEC

[CSVGTP09] Carlos Coelho, Raquel Silva, Fernando Veloso-Gomes, and Francisco Taveira-Pinto. Potential effects of climate change on northwest Portuguese coastal zones. ICES Journal of Marine Science, 66(7):1497–1507, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/7/1497/658590.

Chandrasekar:2006:DEO

[CSW06] Subhashini Chandrasekar, George A. Sorial, and James W. Weaver. Dispersant effectiveness on oil spills — impact of salinity. ICES Journal of Marine Science, 63(8):1418–1430, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/8/1418/712410.

Chereskin:2007:IDV

[CT07] T. K. Chereskin and G. A. Tarling. Interannual to diurnal variability in the near-surface scattering layer in Drake Passage. ICES Journal of Marine Science, 64(9):1617–1626, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/9/1617/779048.

Chou:2002:LTM

[CTF02] Wei-Rung Chou, Kwee Siong Tew, and Lee-Shing Fang. Longterm monitoring of the demersal fish community in a steelslag disposal area in the coastal waters of Kaohsiung, Taiwan.

ICES Journal of Marine Science, 59(S1):S238–S242, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S238/617948.

Cogan:2009:RMH

[CTLN09] Christopher B. Cogan, Brian J. Todd, Peter Lawton, and Thomas T. Noji. The role of marine habitat mapping in ecosystem-based management. ICES Journal of Marine Science, 66(9):2033–2042, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/9/2033/727270.

Cabreira:2009:ANN

[CTM09] Ariel G. Cabreira, Martín Tripode, and Adrián Madirolas. Artificial neural networks for fish-species identification. ICES Journal of Marine Science, 66(6):1119–1129, July 2009. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/1119/690251.

Churnside:2009:CDP

[CTW09] James H. Churnside, Eirik Tenningen, and James J. Wilson. Comparison of data-processing algorithms for the lidar detection of mackerel in the Norwegian Sea. ICES Journal of Marine Science, 66(6):1023–1028, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1023/690921.

Cury:2000:SPU

[Cur00] P. Cury. Small pelagics in upwelling systems: patterns of interaction and structural changes in “wasp-waist” ecosystems. ICES Journal of Marine Science, 57(3):603–618, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/3/603/635940.

Crawford:2007:ACC

[CUUD07] Robert J. M. Crawford, Les G. Underhill, Leshia Upfold, and Bruce M. Dyer. An altered carrying capacity of the Benguela upwelling ecosystem for African penguins (Spheniscus demersus). ICES Journal of Marine Science, 64(3):570–576, April 2007.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/3/570/816245.

Carscadden:2002:I

[CV02] J. E. Carscadden and H. Vilhjálmsson. Introduction. ICES Journal of Marine Science, 59(5):863–869, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/5/863/675194.

Coetzee:2008:FCM

[CvdLHF08] Janet C. Coetzee, Carl D. van der Lingen, Laurence Hutchings, and Tracey P. Fairweather. Has the fishery contributed to a major shift in the distribution of South African sardine? ICES Journal of Marine Science, 65(9):1676–1688, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/9/1676/632495.

Cunha:2008:ESD

[CVG08] M. Emília Cunha, Catarina Vendrell, and Patrícia Gonçalves. Experimental study of the dependence of embryonic development of Trachurus trachurus eggs on temperature. ICES Journal of Marine Science, 65(1):17–24, January 2008. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/65/1/17/
612089.

Chavanich:2009:MBS

[CVL^{+}09] Suchana Chavanich, Voranop Viyakarn, Thepsuda Loyjiw, Priyapat Pattaratamrong, and Anchalee Chankong. Mass bleaching of soft coral, Sarcophyton spp. in Thailand and the role of temperature and salinity stress. ICES Journal of Marine Science, 66(7):1515–1519, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/7/1515/657058.

Chu:2005:MSS

[CW05] D. Chu and P. H. Wiebe. Measurements of sound-speed and density contrasts of zooplankton in Antarctic waters. ICES Journal of Marine Science, 62(4):818–831, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/4/818/677444.

Carr:2006:EJF

[CW06] Jonathan W. Carr and Frederick G. Whoriskey. The escape of juvenile farmed Atlantic salmon from hatcheries into freshwater streams in New Brunswick, Canada. ICES Journal of Marine Science, 63(7):1263–1268, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/7/1263/756823.

Charles:2009:HDM

[CW09a] Anthony Charles and Lisette Wilson. Human dimensions of marine protected areas. ICES Journal of Marine Science, 66 (1):6–15, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/1/6/633456.

Chateau:2009:MPF

[CW09b] Olivier Chateau and Laurent Wantiez. Movement patterns of four coral reef fish species in a fragmented habitat in New Caledonia: implications for the design of marine protected area networks. ICES Journal of Marine Science, 66(1):50–55, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/1/50/632727.

Chu:2000:IMP

[CWC00] Dezhang Chu, Peter Wiebe, and Nancy Copley. Inference of material properties of zooplankton from acoustic and resistivity measurements. ICES Journal of Marine Science, 57(4): 1128–1142, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/4/1128/647282.

Chu:2003:MPN

[CWC^{+}03] Dezhang Chu, Peter H. Wiebe, Nancy J. Copley, Gareth L. Lawson, and Velmurugu Puvanendran. Material properties of North Atlantic cod eggs and early-stage larvae and their influence on acoustic scattering. ICES Journal of Marine Science, 60(3):508–515, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/3/508/658462.

Clarke:2002:CLF

[CWYM^{+}02] Shelley Clarke, Albert Leung Wai-Yin, Y. M. Mak, Robin Kennish, and Nigel Haggan. Consultation with local fishers on the Hong Kong artificial reefs initiative. ICES Journal of Marine Science, 59(S1):S171–S177, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S171/617920.

Chen:2007:INN

[CZC07] Xin Jun Chen, Xiao Hu Zhao, and Yong Chen. Influence of El Niño/La Niña on the western winter–spring cohort of neon flying squid (Ommastrephes bartramii) in the northwestern Pacific Ocean. ICES Journal of Marine Science, 64(6):1152– 1160, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/6/1152/615781.

R:2000:EST

[D00] F. M. E. N. A. R. D. Exploitation of small tunas by a purseseine fishery with fish aggregating devices and their feeding ecology in an eastern tropical Atlantic ecosystem. ICES Journal of Marine Science, 57(3):525–530, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/3/525/635927.

Daan:2001:E

[Daa01a] Niels Daan. Editorial. ICES Journal of Marine Science, 58(4):737, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/4/737/630191.

Daan:2001:F

[Daa01b] Niels Daan. Foreword. ICES Journal of Marine Science, 58(5):935, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/5/935/730855.

Daan:2003:EIJ

[Daa03] Niels Daan. Editing the ICES journal of marine science: a joy forever? ICES Journal of Marine Science, 60(6):1167–1168, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/6/1167/651053.

Daan:2005:AEM

[Daa05] Niels Daan. An afterthought: Ecosystem metrics and pressure indicators. ICES Journal of Marine Science, 62(3):612–613, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/3/612/667171.

Duarte:2009:SFC

[DAAD09] Rafael Duarte, Manuela Azevedo, and Manuel Afonso-Dias. Segmentation and fishery characteristics of the mixed-species multigear Portuguese fleet. ICES Journal of Marine Science, 66 (3):594–606, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/3/594/817215.

Dalley:2002:AFL

[DAd02] Edgar L. Dalley, John T. Anderson, and Brad deYoung. Atmospheric forcing, larval drift, and recruitment of capelin (Mallotus villosus). ICES Journal of Marine Science, 59(5):929–941, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/5/929/675208.

Dobby:2008:IQI

[DAH^{+}08] H. Dobby, L. Allan, M. Harding, C. H. Laurenson, and H. A. McLay. Improving the quality of information on Scottish anglerfish fisheries: making use of fishers’ data. ICES Journal of Marine Science, 65(7):1334–1345, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/7/1334/647453.

Davis:2007:SFE

[Dav07] M. W. Davis. Simulated fishing experiments for predicting delayed mortality rates using reflex impairment in restrained fish. ICES Journal of Marine Science, 64(8):1535–1542, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/8/1535/612639.

DeOliveira:2004:DRJ

[DB04] J. A. A. De Oliveira and D. S. Butterworth. Developing and refining a joint management procedure for the multispecies South African pelagic fishery. ICES Journal of Marine Science, 61 (8):1432–1442, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/8/1432/632739.

Dunstan:2008:MIM

[DB08] Piers K. Dunstan and Nicholas J. Bax. Management of an invasive marine species: defining and testing the effectiveness of ballast-water management options using management strategy evaluation. ICES Journal of Marine Science, 65(6):841– 850, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/6/841/602492.

Dalpadado:2001:DTA

[DBBM01] Padmini Dalpadado, Nina Borkner, Bjarte Bogstad, and Sigbjørn Mehl. Distribution of Themisto (Amphipoda) spp. in the Barents Sea and predator-prey interactions. ICES Journal of Marine Science, 58(4):876–895, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/4/876/630222.

Doonan:2003:SAS

[DBC03] I. J. Doonan, B. Bull, and R. F. Coombs. Star acoustic surveys of localized fish aggregations. ICES Journal of Marine Science, 60(1):132–146, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/1/132/2271507.

Doyle:2002:ELH

[DBDA^{+}02] Miriam J. Doyle, Morgan S. Busby, Janet T. Duffy-Anderson, Susan J. Picquelle, and Ann C. Matarese. Early life history of capelin (Mallotus villosus) in the northwest Gulf of Alaska: a historical perspective based on larval collections, October 1977– March 1979. ICES Journal of Marine Science, 59(5):997–1005, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/5/997/675224.

Duineveld:2007:EAC

[DBL07] Gerard C. A. Duineveld, Magda J. N. Bergman, and Marc S. S. Lavaleye. Effects of an area closed to fisheries on the composition of the benthic fauna in the southern North Sea. ICES Journal of Marine Science, 64(5):899–908, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/5/899/640415.

deBruyn:2009:AAS

[dBMS09] Paul A. de Bruyn, Coleen L. Moloney, and Michael H. Schleyer. Application of age-structured production models to assess oyster Striostrea margaritacea populations managed by rotational harvesting in KwaZulu-Natal, South Africa. ICES Journal of Marine Science, 66(2):408–419, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/2/408/594267.

deBoer:2002:CST

[dBP02] W. F. de Boer and H. H. T. Prins. The community structure of a tropical intertidal mudflat under human exploitation. ICES Journal of Marine Science, 59(6):1237–1247, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/6/1237/608180.

Dicken:2006:POT

[DBS06] M. L. Dicken, A. J. Booth, and M. J. Smale. Preliminary observations of tag shedding, tag reporting, tag wounds, and tag biofouling for raggedtooth sharks (Carcharias taurus) tagged off the east coast of South Africa. ICES Journal of Marine Science, 63(9):1640–1648, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/9/1640/698649.

Defeo:2001:EDA

[DC01] Omar Defeo and John F. Caddy. Evaluating a dynamic approach to yield-mortality models. ICES Journal of Marine Science, 58 (6):1253–1260, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1253/641565.

Demer:2003:RTV

[DC03a] David A. Demer and Stephane G. Conti. Reconciling theoretical versus empirical target strengths of krill: effects of phase variability on the distorted-wave Born approximation. ICES Journal of Marine Science, 60(2):429–434, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/2/429/627957.

Demer:2003:VSD

[DC03b] David A. Demer and Stéphane G. Conti. Validation of the stochastic distorted-wave Born approximation model with broad bandwidth total target strength measurements of Antarctic krill. ICES Journal of Marine Science, 60(3):625–635, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/3/625/660278.

Demer:2004:RTV

[DC04a] David A. Demer and Stéphane G. Conti. Reconciling theoretical versus empirical target strengths of krill: effects of phase variability on the distorted-wave Born approximation. ICES Journal of Marine Science, 61(1):157–158, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/1/157/701395.

Demer:2004:VSD

[DC04b] David A. Demer and Stéphane G. Conti. Validation of the stochastic distorted-wave Born approximation model with broad bandwidth total target strength measurements of Antarctic krill. ICES Journal of Marine Science, 61(1):155–156, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/1/155/701392.

DAmato:2005:PGS

[DC05a] M. E. D’Amato and G. R. Carvalho. Population genetic structure and history of the long-tailed hake, Macruronus magellanicus, in the SW Atlantic as revealed by mtDNA RFLP analysis. ICES Journal of Marine Science, 62(2):247–255, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/2/247/603627.

Demer:2005:NTS

[DC05b] David A. Demer and Stéphane G. Conti. New target-strength model indicates more krill in the Southern Ocean. ICES Journal of Marine Science, 62(1):25–32, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/1/25/667333.

daCosta:2003:UTB

[dCA03] Marcus R. da Costa and Francisco G. Araújo. Use of a tropical bay in southeastern Brazil by juvenile and subadult Micropogonias furnieri (Perciformes, Sciaenidae). ICES Journal of Marine Science, 60(2):268–277, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/2/268/625331.

Dickey-Collas:2009:LHD

[DCCS09] Mark Dickey-Collas, Maurice Clarke, and Aril Slotte. “Linking herring”: do we really understand plasticity? ICES Journal of Marine Science, 66(8):1649–1651, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/8/1649/674956.

Dawe:2000:EER

[DCD00] E. G. Dawe, E. B. Colbourne, and K. F. Drinkwater. Environmental effects on recruitment of short-finned squid (Illex illecebrosus). ICES Journal of Marine Science, 57(4):1002–1013, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/1002/647239.

Doonan:2003:ASS

[DCM03] Ian J. Doonan, Roger F. Coombs, and Sam McClatchie. The absorption of sound in seawater in relation to the estimation of deep-water fish biomass. ICES Journal of Marine Science, 60 (5):1047–1055, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/5/1047/764859.

Daponte:2004:SFR

[DCN^{+}04] M. C. Daponte, F. L. Capitanio, D. E. Nahabedian, M. D. Viñas, and R. M. Negri. Sagitta friderici Ritter–Záhony (Chaetognatha) from South Atlantic waters: abundance, population structure, and life cycle. ICES Journal of Marine Science, 61 (4):680–686, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/4/680/607459.

Dickey-Collas:2007:PWV

[DCPvK07] Mark Dickey-Collas, Martin A. Pastoors, and Olvin A. van Keeken. Precisely wrong or vaguely right: simulations of noisy discard data and trends in fishing effort being included in the stock assessment of North Sea plaice. ICES Journal of Marine Science, 64(9):1641–1649, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/9/1641/786481.

Demer:2009:SSM

[DCRB09] David A. Demer, George R. Cutter, Josiah S. Renfree, and John L. Butler. A statistical-spectral method for echo classification. ICES Journal of Marine Science, 66(6):1081–1090, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1081/691666.

Deysher:2002:DCA

[DDGJ02] Larry E. Deysher, T. A. Dean, R. S. Grove, and A. Jahn. Design considerations for an artificial reef to grow giant kelp (Macrocystis pyrifera) in Southern California. ICES Journal of Marine Science, 59(S1):S201–S207, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S201/617937.

DeLara:2007:MFB

[DDGR07] Michel De Lara, Luc Doyen, Thérèse Guilbaud, and Marie-Joëlle Rochet. Is a management framework based on spawningstock biomass indicators sustainable? A viability approach. ICES Journal of Marine Science, 64(4):761–767, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/4/761/641082.

Dransfeld:2005:EMS

[DDM^{+}05] Leonie Dransfeld, Oonagh Dwane, John Molloy, Sarah Gallagher, and Dave G. Reid. Estimation of mackerel (Scomber scombrus L., 1758) and horse mackerel (Trachurus trachurus L., 1758) daily egg production outside the standard ICES survey area. ICES Journal of Marine Science, 62(8):1705–1710, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/8/1705/796422.

Dinmore:2003:ILS

[DDR^{+}03] T. A. Dinmore, D. E. Duplisea, B. D. Rackham, D. L. Maxwell, and S. Jennings. Impact of a large-scale area closure on patterns of fishing disturbance and the consequences for benthic communities. ICES Journal of Marine Science, 60(2):371–380, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/2/371/627288.

DeBiasi:2004:IET

[De 04] Anna Maria De Biasi. Impact of experimental trawling on the benthic assemblage along the Tuscany coast (north Tyrrhenian Sea, Italy). ICES Journal of Marine Science, 61(8):1260–1266, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/8/1260/629478.

Degnbol:2005:IMC

[Deg05] Poul Degnbol. Indicators as a means of communicating knowledge. ICES Journal of Marine Science, 62(3):606–611, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/3/606/667041.

Dekker:2000:FGE

[Dek00a] Willem Dekker. The fractal geometry of the European eel stock. ICES Journal of Marine Science, 57(1):109–121, February 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/1/109/641103.

Dekker:2000:PAE

[Dek00b] Willem Dekker. A procrustean assessment of the European eel stock. ICES Journal of Marine Science, 57(4):938–947, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/938/647389.

Dekker:2004:WCD

[Dek04] Willem Dekker. What caused the decline of the Lake IJsselmeer eel stock after 1960? ICES Journal of Marine Science, 61 (3):394–404, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/3/394/672021.

Dempson:2001:EME

[Dem01] J. Dempson. Estimation of marine exploitation rates on Atlantic salmon (Salmo salar L.) stocks in Newfoundland, Canada. ICES Journal of Marine Science, 58(1):331–341, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/1/331/603540.

Dalpadado:2000:FFC

[DEMD00] P. Dalpadado, B. Ellertsen, W. Melle, and A. Dommasnes. Food and feeding conditions of Norwegian spring-spawning herring (Clupea harengus) through its feeding migrations. ICES Journal of Marine Science, 57(4):843–857, August 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/4/843/647348.

Denny:2008:DMR

[Den08] C. M. Denny. Development of a method to reduce the spread of the ascidian Didemnum vexillum with aquaculture transfers. ICES Journal of Marine Science, 65(5):805–810, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/5/805/713634.

Desprez:2000:PBI

[Des00] Michel Desprez. Physical and biological impact of marine aggregate extraction along the French coast of the Eastern English Channel: short- and long-term post-dredging restoration. ICES Journal of Marine Science, 57(5):1428–1438, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/5/1428/660948.

Daan:2000:RDE

[DF00] Niels Daan and Michael J. Fogarty. Recruitment dynamics of exploited marine populations: Physical–biological interactions. ICES Journal of Marine Science, 57(2):189–190, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/2/189/620403.

Durbin:2000:ADC

[DGC00a] E. G. Durbin, P. R. Garrahan, and M. C. Casas. Abundance and distribution of Calanus finmarchicus on the Georges Bank during 1995 and 1996. ICES Journal of Marine Science, 57(6): 1664–1685, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/6/1664/768606.

Durbin:2000:DDC

[DGC00b] E. G. Durbin, P. R. Garrahan, and M. C. Casas. Depth distribution of Calanus finmarchicus nauplii on the Georges Bank during 1995 and 1996. ICES Journal of Marine Science, 57(6): 1686–1693, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/6/1686/768614.

Dudas:2009:CRD

[DGK^{+}09] Sarah E. Dudas, Brian A. Grantham, Anthony R. Kirincich, Bruce A. Menge, Jane Lubchenco, and John A. Barth. Current reversals as determinants of intertidal recruitment on the central Oregon coast. ICES Journal of Marine Science, 66(2):396–407, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/2/396/594077.

Danovaro:2002:IAR

[DGMM02] R. Danovaro, C. Gambi, A. Mazzola, and S. Mirto. Influence of artificial reefs on the surrounding infauna: analysis of meiofauna. ICES Journal of Marine Science, 59(S1):S356–S362, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL https://academic.oup.com/icesjms/
article/59/suppl/S356/617997.

Doksaeter:2009:ESM

[DGO^{+}09] Lise Doksæter, Olav R. Godø, Erik Olsen, Leif Nøttestad, and Ruben Patel. Ecological studies of marine mammals using a seabed-mounted echosounder. ICES Journal of Marine Science, 66(6):1029–1036, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1029/697110.

Daan:2005:CNS

[DGPR05] Niels Daan, Henrik Gislason, John G. Pope, and Jake C. Rice. Changes in the North Sea fish community: evidence of indirect effects of fishing? ICES Journal of Marine Science, 62 (2):177–188, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/2/177/602590.

Didrikas:2004:STS

[DH04] Tomas Didrikas and Sture Hansson. In situ target strength of the Baltic Sea herring and sprat. ICES Journal of Marine Science, 61(3):378–382, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/61/3/378/671686.

DeRobertis:2007:PPT

[DH07] Alex De Robertis and Ian Higginbottom. A post-processing technique to estimate the signal-to-noise ratio and remove echosounder background noise. ICES Journal of Marine Science, 64(6):1282–1291, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/64/6/1282/616894.

Dell:2008:SBI

[DH08] James T. Dell and Alistair J. Hobday. School-based indicators of tuna population status. ICES Journal of Marine Science, 65(4):612–622, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/4/612/637047.

Didrikas:2009:ELI

[DH09] Tomas Didrikas and Sture Hansson. Effects of light intensity on activity and pelagic dispersion of fish: studies with a seabedmounted echosounder. ICES Journal of Marine Science, 66 (2):388–395, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/2/388/593972.

delHoyo:2004:DTE

[dHET04] Juan José García del Hoyo, David Castilla Espino, and Ramón Jiménez Toribio. Determination of technical efficiency of fisheries by stochastic frontier models: a case on the Gulf of Cádiz (Spain). ICES Journal of Marine Science, 61(3):416–421, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/3/416/672466.

Dippner:2001:ICV

[DHKV01] J. W. Dippner, J. Hänninen, H. Kuosa, and I. Vuorinen. The influence of climate variability on zooplankton abundance in the Northern Baltic Archipelago Sea (SW Finland). ICES Journal of Marine Science, 58(3):569–578, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/3/569/810076.

DeRobertis:2008:SSD

[DHWW08] Alex De Robertis, Vidar Hjellvik, Neal J. Williamson, and Christopher D. Wilson. Silent ships do not always encounter more fish: comparison of acoustic backscatter recorded by a noise-reduced and a conventional research vessel. ICES Journal of Marine Science, 65(4):623–635, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/4/623/636513.

Dagg:2008:HCF

[DiUVH08] Michael Dagg, Shin ichi Uye, Luis Valdés, and Roger Harris. Human and climate forcing of zooplankton populations: Introduction. ICES Journal of Marine Science, 65(3):277–278, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/3/277/788879.

Dahle:2006:GCB

[DJRO06] G. Dahle, K. E. Jørstad, H. E. Rusaas, and H. Otterå. Genetic characteristics of broodstock collected from four Norwegian coastal cod (Gadus morhua) populations. ICES Journal of Marine Science, 63(2):209–215, ???? 2006. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/2/209/637575.

Dale:2000:DPS

[DK00] T. Dale and S. Kaartvedt. Diel patterns in stage-specific vertical migration of Calanus finmarchicus in habitats with midnight sun. ICES Journal of Marine Science, 57(6):1800–1818, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/6/1800/768674.

Demer:2009:IPS

[DKMO09] David A. Demer, Rudy J. Kloser, David N. MacLennan, and Egil Ona. An introduction to the proceedings and a synthesis of the 2008 ICES Symposium on the Ecosystem Approach with Fisheries Acoustics and Complementary Technologies (SEAFACTS). ICES Journal of Marine Science, 66(6):961–965, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/961/697621.

Dutil:2003:WSC

[DLC03] Jean-Denis Dutil, Yvan Lambert, and Denis Chabot. Winter and spring changes in condition factor and energy reserves of wild cod compared with changes observed during fooddeprivation in the laboratory. ICES Journal of Marine Science, 60(4):780–786, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/4/780/692312.

Drinkwater:2005:I

[DLM^{+}05] Kenneth F. Drinkwater, Harald Loeng, Bernard A. Megrey, Nick Bailey, and Robin M. Cook. Introduction. ICES Journal of Marine Science, 62(7):1203–1204, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/7/1203/654597.

Moreno:2000:PDR

[dLMACC00] J. I. de Leiva Moreno, V. N. Agostini, J. F. Caddy, and F. Carocci. Is the pelagic-demersal ratio from fishery landings a useful proxy for nutrient availability? A preliminary data exploration for the semi-enclosed seas around Europe. ICES Journal of Marine Science, 57(4):1091–1102, August 2000.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/4/1091/647272.

deLestang:2006:IVM

[dLMS06] Simon de Lestang and Roy Melville-Smith. Interannual variation in the moult cycle and size at double breeding of mature female western rock lobster (Panulirus cygnus). ICES Journal of Marine Science, 63(9):1631–1639, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/9/1631/698509.

Denis:2002:STA

[DLR02] V. Denis, J. Lejeune, and J. P. Robin. Spatio-temporal analysis of commercial trawler data using general additive models: patterns of loliginid squid abundance in the north-east Atlantic. ICES Journal of Marine Science, 59(3):633–648, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/3/633/610849.

Diachok:2001:END

[DLS01] Orest Diachok, Bernard Liorzou, and Carla Scalabrin. Estimation of the number density of fish from resonance absorptivity and echo sounder data. ICES Journal of Marine Science, 58 (1):137–153, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/1/137/603505.

Drinkwater:2000:ERS

[DLT^{+}00] K. Drinkwater, S. Lochman, C. Taggart, K. Thompson, and K. Frank. Entrainment of redfish (Sebastes spp.) larvae off the Scotian Shelf. ICES Journal of Marine Science, 57(2):372–382, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/2/372/620460.

DiCapua:2004:PSC

[DM04] Iole Di Capua and Maria Grazia Mazzocchi. Population structure of the copepods Centropages typicus and Temora stylifera in different environmental conditions. ICES Journal of Marine Science, 61(4):632–644, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/61/4/632/606283.

Dunford:2006:PDS

[DM06] Adam J. Dunford and Gavin J. Macaulay. Progress in determining southern blue whiting (Micromesistius australis) target strength: results of swimbladder modelling. ICES Journal of Marine Science, 63(5):952–955, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/5/952/664482.

Daskalov:2007:IFA

[DM07a] Georgi M. Daskalov and Elchin V. Mamedov. Integrated fisheries assessment and possible causes for the collapse of anchovy kilka in the Caspian Sea. ICES Journal of Marine Science, 64 (3):503–511, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/3/503/814192.

Degnbol:2007:UPC

[DM07b] Poul Degnbol and Bonnie J. McCay. Unintended and perverse consequences of ignoring linkages in fisheries systems. ICES Journal of Marine Science, 64(4):793–797, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/4/793/641747.

Dommasnes:2004:HMC

[DMDE04] Are Dommasnes, Webjørn Melle, Padmini Dalpadado, and Bjørnar Ellertsen. Herring as a major consumer in the Norwegian Sea. ICES Journal of Marine Science, 61(5):739–751, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/5/739/863719.

Delaney:2007:IDD

[DMvD07] Alyne E. Delaney, H. Anne McLay, and Wim L. T. van Densen. Influences of discourse on decision-making in EU fisheries management: the case of North Sea cod (Gadus morhua). ICES Journal of Marine Science, 64(4):804–810, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/4/804/640773.

Duarte-Neto:2008:USO

[DNLSM08] Paulo Duarte-Neto, Rosângela Lessa, Borko Stosic, and Eric Morize. The use of sagittal otoliths in discriminating stocks of common dolphinfish (Coryphaena hippurus) off northeastern Brazil using multishape descriptors. ICES Journal of Marine Science, 65(7):1144–1152, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/7/1144/642740.

Dalen:2003:CAA

[DNP03] John Dalen, Kjell Nedreaas, and Ronald Pedersen. A comparative acoustic-abundance estimation of pelagic redfish (Sebastes mentella) from hull-mounted and deep-towed acoustic systems. ICES Journal of Marine Science, 60(3):472–479, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/3/472/658028.

Ditton:2002:DAR

[DOBT02] Robert B. Ditton, Hal R. Osburn, Troy L. Baker, and Carol E. Thailing. Demographics, attitudes, and reef management preferences of sport divers in offshore Texas waters. ICES Journal of Marine Science, 59(S1):S186–S191, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S186/617923.

Dolgov:2002:RCM

[Dol02] A. V. Dolgov. The role of capelin (Mallotus villosus) in the foodweb of the Barents Sea. ICES Journal of Marine Science, 59(5):1034–1045, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/5/1034/675134.

Dorn:2001:FBF

[Dor01] Martin W. Dorn. Fishing behavior of factory trawlers: a hierarchical model of information processing and decision-making. ICES Journal of Marine Science, 58(1):238–252, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/1/238/603525.

Dumas:2003:VDPa

[DP03a] J. Dumas and P. Prouzet. Variability of demographic parameters and population dynamics of Atlantic salmon Salmo salar L. in a south-west French river. ICES Journal of Marine Science, 60 (2):356–370, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/2/356/626995.

Dumas:2003:VDPb

[DP03b] J. Dumas and P. Prouzet. Variability of demographic parameters and population dynamics of Atlantic salmon (Salmo salar L.) in a southwest French river. ICES Journal of Marine Science, 60(5):1165, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/5/1165/768976.

dePontual:2003:PTE

[dPBB^{+}03] Hélène de Pontual, Michel Bertignac, André Battaglia, Gérard Bavouzet, Philippe Moguedet, and Anne-Laure Groison. A pilot tagging experiment on European hake (Merluccius merluccius): methodology and preliminary results. ICES Journal of Marine Science, 60(6):1318–1327, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/6/1318/653622.

dePontual:2006:EUE

[dPGPB06] Hélène de Pontual, Anne Laure Groison, Carmen Piñeiro, and Michel Bertignac. Evidence of underestimation of European hake growth in the Bay of Biscay, and its relationship with bias in the agreed method of age estimation. ICES Journal of Marine Science, 63(9):1674–1681, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/9/1674/699179.

dePuelles:2008:DCH

[dPM08] Maria Luz Fernández de Puelles and Juan Carlos Molinero. Decadal changes in hydrographic and ecological time-series in the Balearic Sea (western Mediterranean), identifying links between climate and zooplankton. ICES Journal of Marine Science, 65(3):311–317, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/3/311/786410.

Doray:2009:IEV

[DPN^{+}09] Mathieu Doray, Pierre Petitgas, Laetitia Nelson, Stéphanie Mahévas, Erwan Josse, and Lionel Reynal. The influence of the environment on the variability of monthly tuna biomass around a moored, fish-aggregating device. ICES Journal of Marine Science, 66(6):1410–1416, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1410/691133.

dePuelles:2004:HCZ

[dPVJM04] M. Luz Fernández de Puelles, Joaquín Valencia, Javier Jansá, and Ana Morillas. Hydrographical characteristics and zooplankton distribution in the Mallorca channel (Western Mediterranean): spring 2001. ICES Journal of Marine Science, 61 (4):654–666, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/4/654/606504.

dePuelles:2004:ZVC

[dPVV04] M. Luz Fernández de Puelles, Joaquín Valencia, and Laura Vicente. Zooplankton variability and climatic anomalies from 1994 to 2001 in the Balearic Sea (Western Mediterranean). ICES Journal of Marine Science, 61(4):492–500, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/4/492/603204.

Dekker:2007:TME

[DPW07] Willem Dekker, Mike Pawson, and Håkan Wickström. Is there more to eels than slime? An introduction to papers presented at the ICES Theme Session in September 2006. ICES Journal of Marine Science, 64(7):1366–1367, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/7/1366/728958.

deRobertis:2001:VAE

[dR01] Alex de Robertis. Validation of acoustic echo counting for studies of zooplankton behavior. ICES Journal of Marine Science, 58(3):543–561, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/3/543/810064.

Demer:2008:VET

[DR08a] David A. Demer and Josiah S. Renfree. Variations in echosounder–transducer performance with water temperature. ICES Journal of Marine Science, 65(6):1021–1035, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/6/1021/602098.

Duplisea:2008:PSR

[DR08b] Daniel E. Duplisea and Dominique Robert. Prerecruit survival and recruitment of northern Gulf of St Lawrence Atlantic cod. ICES Journal of Marine Science, 65(6):946–952, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/6/946/604211.

DeOliveira:2006:IUP

[DRDC06] J. A. A. De Oliveira, B. A. Roel, and M. Dickey-Collas. Investigating the use of proxies for fecundity to improve management advice for western horse mackerel Trachurus trachurus. ICES Journal of Marine Science, 63(1):25–35, ???? 2006. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/63/1/25/
624734.

Drinkwater:2005:RAC

[Dri05] Kenneth F. Drinkwater. The response of Atlantic cod (Gadus morhua) to future climate change. ICES Journal of Marine Science, 62(7):1327–1337, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/7/1327/656867.

Davies:2001:TEM

[DRRS01] I. M. Davies, G. K. Rodger, J. Redshaw, and R. M. Stagg. Targeted environmental monitoring for the effects of medicines used to treat sea-lice infestation on farmed fish. ICES Journal of Marine Science, 58(2):477–485, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/2/477/617853.

Dunn:2009:NGO

[DRSD09] M. R. Dunn, G. J. Rickard, P. J. H. Sutton, and I. J. Doonan. Nursery grounds of the orange roughy around New Zealand.

ICES Journal of Marine Science, 66(5):871–885, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/5/871/663853.

Darriba:2005:ESU

[DSG05] Susana Darriba, Fuencisla San Juan, and Alejandro Guerra. Energy storage and utilization in relation to the reproductive cycle in the razor clam Ensis arcuatus (Jeffreys, 1865). ICES Journal of Marine Science, 62(5):886–896, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/5/886/856445.

DeRobertis:2003:AOS

[DSJ03] Alex De Robertis, Chad Schell, and Jules S. Jaffe. Acoustic observations of the swimming behavior of the euphausiid Euphausia pacifica Hansen. ICES Journal of Marine Science, 60 (4):885–898, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/4/885/694370.

Diaz:2008:ADE

[DSV^{+}08] Paz Díaz, Juan Santos, Francisco Velasco, Alberto Serrano, and Nélida Pérez. Anglerfish discard estimates and patterns in Spanish Northeast Atlantic trawl fisheries. ICES Journal of Marine Science, 65(7):1350–1361, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/7/1350/646394.

Drabsch:2001:LIR

[DTC01] Sharon L. Drabsch, Jason E. Tanner, and Sean D. Connell. Limited infaunal response to experimental trawling in previously untrawled areas. ICES Journal of Marine Science, 58 (6):1261–1271, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1261/641569.

Dunn:2001:BEJ

[Dun01] M. R. Dunn. The biology and exploitation of John Dory, Zeus faber (Linnaeus, 1758) in the waters of England and Wales. ICES Journal of Marine Science, 58(1):96–105, ???? 2001. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/58/1/96/
603557.

Duplisea:2005:RGP

[Dup05] Daniel E. Duplisea. Running the gauntlet: the predation environment of small fish in the northern Gulf of St Lawrence, Canada. ICES Journal of Marine Science, 62(3):412–416, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/3/412/661526.

delValle:2007:IDF

[dVA07] Ikerne del Valle and Kepa Astorkiza. Institutional designs to face the dark side of total allowable catches. ICES Journal of Marine Science, 64(4):851–857, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/851/643173.

DeRobertis:2006:WPR

[DW06] Alex De Robertis and Christopher D. Wilson. Walleye pollock respond to trawling vessels. ICES Journal of Marine Science, 63(3):514–522, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/3/514/722520.

Dwyer:2003:ADV

[DWC03] Karen S. Dwyer, Stephen J. Walsh, and Steven E. Campana. Age determination, validation and growth of Grand Bank yellowtail flounder (Limanda ferruginea). ICES Journal of Marine Science, 60(5):1123–1138, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/5/1123/767244.

Desroy:2003:MRS

[DWDD03] N. Desroy, C. Warembourg, J. M. Dewarumez, and J. C. Dauvin. Macrobenthic resources of the shallow soft-bottom sediments in the eastern English Channel and southern North Sea. ICES Journal of Marine Science, 60(1):120–131, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/1/120/611426.

Egli:2004:UTR

[EB04] D. P. Egli and R. C. Babcock. Ultrasonic tracking reveals multiple behavioural modes of snapper (Pagrus auratus) in a temperate no-take marine reserve. ICES Journal of Marine Science, 61 (7):1137–1143, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/7/1137/880649.

Ebert:2006:RBC

[ECC06] David A. Ebert, Leonard J. V. Compagno, and Paul D. Cowley. Reproductive biology of catsharks (Chondrichthyes: Scyliorhinidae) off the west coast of southern Africa. ICES Journal of Marine Science, 63(6):1053–1065, ???? 2006. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/6/1053/614996.

Ebert:2008:ARB

[ECC08] David A. Ebert, Leonard J. V. Compagno, and Paul D. Cowley. Aspects of the reproductive biology of skates (Chondrichthyes: Rajiformes: Rajoidei) from southern Africa. ICES Journal of Marine Science, 65(1):81–102, January 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/1/81/612673.

Engelhard:2003:AMP

[EDG03] Georg H. Engelhard, Ulf Dieckmann, and Olav Rune Godø. Age at maturation predicted from routine scale measurements in Norwegian spring-spawning herring (Clupea harengus) using discriminant and neural network analyses. ICES Journal of Marine Science, 60(2):304–313, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/60/2/304/626008.

Ellingsen:2002:ACS

[EGB02] Kari E. Ellingsen, John S. Gray, and Erik Bjørnbom. Acoustic classification of seabed habitats using the QTC VIEW^{T M} system. ICES Journal of Marine Science, 59(4):825–835, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/4/825/676814.

Espeland:2007:HRE

[EGO^{+}07] Sigurd Heiberg Espeland, Ailin Fernløf Gundersen, Esben Moland Olsen, Halvor Knutsen, Jakob Gjøsæter, and Nils C. Stenseth. Home range and elevated egg densities within an inshore spawning ground of coastal cod. ICES Journal of Marine Science, 64(5):920–928, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/64/5/920/640344.

Escribano:2000:INN

[EH00] R. Escribano and P. Hidalgo. Influence of El Niño and La Niña on the population dynamics of Calanus chilensis in the Humboldt Current ecosystem of northern Chile. ICES Journal of Marine Science, 57(6):1867–1874, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/6/1867/768700.

Evans:2004:DSW

[EH04] Karen Evans and Mark A. Hindell. The diet of sperm whales (Physeter macrocephalus) in southern Australian waters. ICES Journal of Marine Science, 61(8):1313–1329, ???? 2004. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/8/1313/630486.

Ehrhold:2006:RMN

[EHG06] Axel Ehrhold, Dominique Hamon, and Brigitte Guillaumont. The REBENT monitoring network, a spatially integrated, acoustic approach to surveying nearshore macrobenthic habitats: application to the Bay of Concarneau (South Brittany, France). ICES Journal of Marine Science, 63(9):1604–1615, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/9/1604/697792.

Eayrs:2007:AJT

[EHL07] Steve Eayrs, Nguyen Phong Hai, and Janet Ley. Assessment of a juvenile and trash excluder device in a Vietnamese shrimp trawl fishery. ICES Journal of Marine Science, 64(8):1598– 1602, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/8/1598/613768.

Eigaard:2009:BAT

[Eig09] Ole Ritzau Eigaard. A bottom-up approach to technological development and its management implications in a commercial fishery. ICES Journal of Marine Science, 66(5):916–927, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/5/916/663733.

Erzini:2005:ATT

[EIS05] Karim Erzini, Cheikh A. O. Inejih, and Kim A. Stobberup. An application of two techniques for the analysis of short, multivariate non-stationary time-series of Mauritanian trawl survey data. ICES Journal of Marine Science, 62(3):353–359, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/3/353/659611.

Ewart:2001:LII

[EJR01] K. V. Ewart, S. C. Johnson, and N. W. Ross. Lectins of the innate immune system and their relevance to fish health. ICES Journal of Marine Science, 58(2):380–385, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/2/380/617828.

Ellis:2008:FNA

[EK08] Jim R. Ellis and Jennie Keable. Fecundity of Northeast Atlantic spurdog (Squalus acanthias). ICES Journal of Marine Science, 65(6):979–981, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/6/979/604161.

Eero:2007:EBC

[EKPT07] Margit Eero, Friedrich W. Köster, Maris Plikshs, and Fritz Thurow. Eastern Baltic cod (Gadus morhua callarias) stock dynamics: extending the analytical assessment back to the mid-1940s. ICES Journal of Marine Science, 64(6):1257– 1271, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/6/1257/616930.

Eleftheriou:2000:MBD

[Ele00] A. Eleftheriou. Marine benthos dynamics: Environmental and fisheries impacts: Introduction and overview. ICES Journal of Marine Science, 57(5):1299–1302, October 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/5/1299/660907.

Ehrhardt:2001:DDL

[ELR01] Nelson M. Ehrhardt, Christopher M. Legault, and Victor R. Restrepo. Density-dependent linkage between juveniles and recruitment for pink shrimp (Farfantepenaeus duorarum) in southern Florida. ICES Journal of Marine Science, 58(5):1100–1105, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/5/1100/730841.

Eastwood:2007:HAU

[EMA^{+}07] P. D. Eastwood, C. M. Mills, J. N. Aldridge, C. A. Houghton, and S. I. Rogers. Human activities in UK offshore waters: an assessment of direct, physical pressure on the seabed. ICES Journal of Marine Science, 64(3):453–463, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/3/453/815129.

Evans:2002:WMA

[EN02] Geoffrey T. Evans and Brian S. Nakashima. A weighted multiplicative analysis to estimate trends in year-class size of capelin. ICES Journal of Marine Science, 59(5):1116–1119, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/5/1116/675177.

Eno:2001:ECT

[Eno01] N. Eno. Effects of crustacean traps on benthic fauna. ICES Journal of Marine Science, 58(1):11–20, ???? 2001. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/58/1/11/
603498.

Edwards:2009:ICM

[ERBP09] Charles T. T. Edwards, Rébecca A. Rademeyer, Doug S. Butterworth, and Éva E. Plagányi. Investigating the consequences of marine protected areas for the South African deep-water hake (Merluccius paradoxus) resource. ICES Journal of Marine Science, 66(1):72–81, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/1/72/633887.

Exadactylos:2007:CAN

[ERGT07] Athanasios Exadactylos, Mark J. Rigby, Audrey J. Geffen, and John P. Thorpe. Conservation aspects of natural populations and captive-bred stocks of turbot (Scophthalmus maximus) and Dover sole (Solea solea) using estimates of genetic diversity. ICES Journal of Marine Science, 64(6):1173–1181, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/6/1173/2835575.

Ermolchev:2009:MRS

[Erm09] Viacheslav A. Ermolchev. Methods and results of in situ targetstrength measurements of Atlantic cod (Gadus morhua) during combined trawl-acoustic surveys. ICES Journal of Marine Science, 66(6):1225–1232, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/6/1225/690675.

Edwards:2001:LTR

[ERP01] Martin Edwards, Philip Reid, and Benjamin Planque. Longterm and regional variability of phytoplankton biomass in the Northeast Atlantic (1960–1995). ICES Journal of Marine Science, 58(1):39–49, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/58/1/39/603550.

Ehrenberg:2002:MEP

[ES02] John E. Ehrenberg and Tracey W. Steig. A method for estimating the “position accuracy” of acoustic fish tags. ICES Journal of Marine Science, 59(1):140–149, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/1/140/649980.

Ehrenberg:2003:ITS

[ES03] John E. Ehrenberg and Tracey W. Steig. Improved techniques for studying the temporal and spatial behavior of fish in a fixed location. ICES Journal of Marine Science, 60(3):700–706, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/3/700/661388.

Ehrenberg:2009:SRB

[ES09] John E. Ehrenberg and Tracey W. Steig. A study of the relationship between tag-signal characteristics and achievable performances in acoustic fish-tag studies. ICES Journal of Marine Science, 66(6):1278–1283, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1278/691419.

Esmaeili:2006:TEA

[Esm06] Abdoulkarim Esmaeili. Technical efficiency analysis for the Iranian fishery in the Persian Gulf. ICES Journal of Marine Science, 63(9):1759–1764, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/9/1759/701695.

Eggert:2007:PRO

[ET07] Håkan Eggert and Ragnar Tveterås. Potential rent and overcapacity in the Swedish Baltic Sea trawl fishery for cod (Gadus morhua). ICES Journal of Marine Science, 64(3):439–445, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/3/439/817008.

Everson:2007:IAE

[ETB07] Inigo Everson, Geraint A. Tarling, and Bo Bergström. Improving acoustic estimates of krill: experience from repeat sampling of northern krill (Meganyctiphanes norvegica) in Gullmarsfjord, Sweden. ICES Journal of Marine Science, 64(1):39–48, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/1/39/647706.

Ellis:2007:EFD

[EW07] Nick Ellis and You-Gan Wang. Effects of fish density distribution and effort distribution on catchability. ICES Journal of Marine Science, 64(1):178–191, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/1/178/648421.

Ermolchev:2003:REV

[EZ03] Viatcheslav A. Ermolchev and Michael L. Zaferman. Results of experiments on the video-acoustic estimation of fish target strength in situ. ICES Journal of Marine Science, 60 (3):544–547, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/3/544/658988.

Farina:2008:LWS

[FAL^{+}08] A. C. Fariña, M. Azevedo, J. Landa, R. Duarte, P. Sampedro, G. Costas, M. A. Torres, and L. Cañás. Lophius in the world: a synthesis on the common features and life strategies. ICES Journal of Marine Science, 65(7):1272–1280, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/7/1272/647165.

Falace:2002:EII

[FB02a] A. Falace and G. Bressan. Evaluation of the influence of inclination of substrate panels on seasonal changes in a macrophytobenthic community. ICES Journal of Marine Science, 59 (S1):S116–S121, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S116/617901.

Falace:2002:QQA

[FB02b] A. Falace and G. Bressan. A qualitative and quantitative analysis of the evolution of macroalgal vegetation on an artificial reef with anti-grazing nets (Loano-Ligurian Sea). ICES Journal of Marine Science, 59(S1):S150–S156, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S150/617912.

Fleischman:2003:MMS

[FB03] Steve J. Fleischman and Debby L. Burwen. Mixture models for the species apportionment of hydroacoustic data, with echoenvelope length as the discriminatory variable. ICES Journal of Marine Science, 60(3):592–598, ???? 2003. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/3/592/659652.

Faraj:2007:SCD

[FB07] Abdelmalek Faraj and Nicolas Bez. Spatial considerations for the dakhla stock of Octopus vulgaris: indicators, patterns, and fisheries interactions. ICES Journal of Marine Science, 64(9): 1820–1828, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/9/1820/788003.

Fiorentino:2008:CSS

[FBD^{+}08] F. Fiorentino, F. Badalamenti, G. D’Anna, G. Garofalo, P. Gianguzza, M. Gristina, C. Pipitone, P. Rizzo, and T. Fortibuoni. Changes in spawning-stock structure and recruitment pattern of red mullet, Mullus barbatus, after a trawl ban in the Gulf of Castellammare (central Mediterranean Sea). ICES Journal of Marine Science, 65(7):1175–1183, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/7/1175/645210.

Fassler:2009:BAE

[FBF09] Sascha M. M. Fässler, Andrew S. Brierley, and Paul G. Fernandes. A Bayesian approach to estimating target strength. ICES Journal of Marine Science, 66(6):1197–1204, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/1197/689763.

Figueiredo:2003:ORC

[FBMR^{+}03] I. Figueiredo, P. Bordalo-Machado, S. Reis, D. Sena-Carvalho, T. Blasdale, A. Newton, and L. S. Gordo. Observations on the reproductive cycle of the black scabbardfish (Aphanopus carbo Lowe, 1839) in the NE Atlantic. ICES Journal of Marine Science, 60(4):774–779, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/60/4/774/692220.

Friedland:2005:ERC

[FCM05] Kevin D. Friedland, Gerald Chaput, and Julian C. MacLean. The emerging role of climate in post-smolt growth of Atlantic salmon. ICES Journal of Marine Science, 62(7):1338–1349, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/7/1338/657365.

Farrell:2009:SGI

[FCM09] Edward D. Farrell, Maurice W. Clarke, and Stefano Mariani. A simple genetic identification method for Northeast Atlantic smoothhound sharks (Mustelus spp.). ICES Journal of Marine Science, 66(3):561–565, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/66/3/561/812519.

Freon:2005:SEI

[FDD^{+}05] Pierre Fréon, Laurent Drapeau, Jeremy H. M. David, Almudena Fernández Moreno, Rob W. Leslie, W. Herman Oosthuizen, Lynne J. Shannon, and Carl D. van der Lingen. Spatialized ecosystem indicators in the southern Benguela. ICES Journal of Marine Science, 62(3):459–468, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/3/459/662759.

Fernandes:2009:CTS

[Fer09] Paul G. Fernandes. Classification trees for species identification of fish-school echotraces. ICES Journal of Marine Science, 66 (6):1073–1080, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1073/691837.

Fredou:2006:UMS

[FFL06] Thierry Frédou, Beatrice P. Ferreira, and Yves Letourneur. A univariate and multivariate study of reef fisheries off northeastern Brazil. ICES Journal of Marine Science, 63(5):883–896, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/5/883/663451.

Fassler:2009:TSB

[FG09] Sascha M. M. Fässler and Natalia Gorska. On the target strength of Baltic clupeids. ICES Journal of Marine Science, 66(6):1184–1190, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1184/689176.

Fromentin:2000:BPE

[FGBS00] Jean-Marc Fromentin, Jakob Gjøsæter, Ottar N. Bjørnstad, and N. Chr. Stenseth. Biological processes and environmental factors regulating the dynamics of the Norwegian Skagerrak cod populations since 1919. ICES Journal of Marine Science, 57 (2):330–338, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/2/330/620444.

Fortunati:2002:TGT

[FGD02] L. Fortunati, G. Garofalo, and R. Demontis. TSDV: a GIS tool for inspecting trawl survey data. ICES Journal of Marine Science, 59(1):168–178, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/59/1/168/649996.

Fraser:2008:MSV

[FGFP08] Helen M. Fraser, Simon P. R. Greenstreet, Rob J. Fryer, and Gerjan J. Piet. Mapping spatial variation in demersal fish species diversity and composition in the North Sea: accounting for species- and size-related catchability in survey trawls. ICES Journal of Marine Science, 65(4):531–538, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/4/531/638964.

Fabi:2002:EFA

[FGLT02] G. Fabi, F. Grati, A. Lucchetti, and L. Trovarelli. Evolution of the fish assemblage around a gas platform in the northern Adriatic Sea. ICES Journal of Marine Science, 59(S1):S309–S315, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL https://academic.oup.com/icesjms/
article/59/suppl/S309/617976.

Fraser:2007:TAC

[FGP07] Helen M. Fraser, Simon P. R. Greenstreet, and Gerjan J. Piet. Taking account of catchability in groundfish survey trawls: implications for estimating demersal fish biomass. ICES Journal of Marine Science, 64(9):1800–1819, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/9/1800/779494.

Fraser:2009:SMC

[FGP09] Helen M. Fraser, S. P. R. Greenstreet, and Gerjan J. Piet. Selecting MPAs to conserve groundfish biodiversity: the consequences of failing to account for catchability in survey trawls. ICES Journal of Marine Science, 66(1):82–89, January 2009. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/66/1/82/
633287.

Fielding:2004:BVA

[FGR04] S. Fielding, G. Griffiths, and H. S. J. Roe. The biological validation of ADCP acoustic backscatter through direct comparison with net samples and model predictions based on acousticscattering models. ICES Journal of Marine Science, 61(2): 184–200, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/2/184/620162.

Friedland:2000:LBO

[FHDM00] Kevin D. Friedland, Lars P. Hansen, David A. Dunkley, and Julian C. MacLean. Linkage between ocean climate, post-smolt growth, and survival of Atlantic salmon (Salmo salar L.) in the North Sea area. ICES Journal of Marine Science, 57(2):419–429, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/2/419/620479.

Frid:2000:LTC

[FHHH00] C. L. J. Frid, K. G. Harwood, S. J. Hall, and J. A. Hall. Longterm changes in the benthic communities on North Sea fishing grounds. ICES Journal of Marine Science, 57(5):1303–1309, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/5/1303/660909.

Folkvord:2009:EPC

[FHJS09] Arild Folkvord, Hans Høie, Arne Johannessen, and Turid Solbakken. Effects of prey concentration, light regime, and parental origin on growth and survival of herring larvae under controlled experimental conditions. ICES Journal of Marine Science, 66 (8):1702–1709, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/8/1702/673212.

Fiksen:2000:ATD

[Fik00] Ø. Fiksen. The adaptive timing of diapause — a search for evolutionarily robust strategies in Calanus finmarchicus. ICES Journal of Marine Science, 57(6):1825–1833, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/6/1825/768681.

Fjalling:2005:EHS

[Fjä05] Arne Fjälling. The estimation of hidden seal-inflicted losses in the Baltic Sea set-trap salmon fisheries. ICES Journal of Marine Science, 62(8):1630–1635, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/8/1630/792145.

Ferro:2007:SSU

[FJK^{+}07] R. S. T. Ferro, E. G. Jones, R. J. Kynoch, R. J. Fryer, and B-E. Buckett. Separating species using a horizontal panel in the Scottish North Sea whitefish trawl fishery. ICES Journal of Marine Science, 64(8):1543–1550, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/8/1543/613073.

Fernandez-Jover:2008:SPD

[FJSJBS^{+}08] Damian Fernandez-Jover, Pablo Sanchez-Jerez, Just Tomás Bayle-Sempere, Carlos Valle, and Tim Dempster. Seasonal patterns and diets of wild fish assemblages associated with Mediterranean coastal fish farms. ICES Journal of Marine Science, 65 (7):1153–1160, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/7/1153/643026.

Fey:2006:PJB

[FL06] Dariusz P. Fey and Tomasz B. Linkowski. Predicting juvenile Baltic cod (Gadus morhua) age from body and otolith size measurements. ICES Journal of Marine Science, 63(6):1045–1052, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/6/1045/614860.

Fletcher:2005:AQR

[Fle05] W. J. Fletcher. The application of qualitative risk assessment methodology to prioritize issues for fisheries management.

ICES Journal of Marine Science, 62(8):1576–1587, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/8/1576/789703.

Ferno:2004:FBE

[FLH04] Anders Fernö, Svein Løkkeborg, and C. E. Hollingworth. Fish behaviour in exploited ecosystems: Preface. ICES Journal of Marine Science, 61(7):1029, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/61/7/1029/877518.

Fiske:2006:RBF

[FLH06] Peder Fiske, Roar A. Lund, and Lars P. Hansen. Relationships between the frequency of farmed Atlantic salmon, Salmo salar L., in wild salmon populations and fish farming activity in Norway, 1989–2004. ICES Journal of Marine Science, 63 (7):1182–1189, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/7/1182/753755.

Fablet:2009:CFS

[FLK^{+}09] Ronan Fablet, Riwal Lefort, Imen Karoui, Laurent Berger, Jacques Massé, Carla Scalabrin, and Jean-Marc Boucher. Classifying fish schools and estimating their species proportions in fishery-acoustic surveys. ICES Journal of Marine Science, 66 (6):1136–1142, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1136/695639.

Fabi:2002:EAR

[FLP^{+}02] G. Fabi, F. Luccarini, M. Panfili, C. Solustri, and A. Spagnolo. Effects of an artificial reef on the surrounding soft-bottom community (central Adriatic Sea). ICES Journal of Marine Science, 59(S1):S343–S349, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/59/suppl/S343/617991.

Fiori:2004:AIG

[FM04] Sandra M. Fiori and Enrique M. Morsán. Age and individual growth of Mesodesma mactroides (Bivalvia) in the southernmost range of its distribution. ICES Journal of Marine Science, 61 (8):1253–1259, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/8/1253/629347.

Fogarty:2001:RCH

[FMB01] Michael J. Fogarty, Ransom A. Myers, and Keith G. Bowen. Recruitment of cod and haddock in the North Atlantic: a comparative analysis. ICES Journal of Marine Science, 58(5):952–961, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/5/952/730866.

Frisk:2002:PDL

[FMF02] M. G. Frisk, T. J. Miller, and M. J. Fogarty. The population dynamics of little skate Leucoraja erinacea, winter skate Leucoraja ocellata, and barndoor skate Dipturus laevis: predicting exploitation limits using matrix analyses. ICES Journal of Marine Science, 59(3):576–586, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/3/576/610839.

Friedland:2009:RAS

[FMH^{+}09] Kevin D. Friedland, Julian C. MacLean, Lars P. Hansen, Arnaud J. Peyronnet, Lars Karlsson, David G. Reddin, Niall Ó Maoiléidigh, and Jennifer L. McCarthy. The recruitment of Atlantic salmon in Europe. ICES Journal of Marine Science, 66 (2):289–304, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/2/289/594931.

Friedland:2007:OCS

[FMK07] Kevin D. Friedland, Michael J. Miller, and Brian Knights. Oceanic changes in the Sargasso Sea and declines in recruitment of the European eel. ICES Journal of Marine Science, 64(3):519–530, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/3/519/817502.

Fryer:2000:USC

[FN00] R. J. Fryer and M. D. Nicholson. Using smoothers for comprehensive assessments of contaminant time series in marine biota. ICES Journal of Marine Science, 57(5):1525, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/5/1525/660975.

Fryer:2002:ACD

[FN02] R. J. Fryer and M. D. Nicholson. Assessing covariate-dependent contaminant time-series in the marine environment. ICES Journal of Marine Science, 59(1):1–14, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/1/1/649963.

Fogarty:2001:I

[Fog01] M. J. Fogarty. Introduction. ICES Journal of Marine Science, 58(5):936, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/5/936/730858.

Fox:2001:RTS

[Fox01] C. J. Fox. Recent trends in stock-recruitment of blackwater herring (Clupea harengus L.) in relation to larval production. ICES Journal of Marine Science, 58(4):750–762, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/4/750/630197.

Frie:2003:TAM

[FPKH03] Anne Kirstine Frie, Vladimir A. Potelov, Michael C. S. Kingsley, and Tore Haug. Trends in age-at-maturity and growth parameters of female Northeast Atlantic harp seals, Pagophilus groenlandicus (Erxleben, 1777). ICES Journal of Marine Science, 60 (5):1018–1032, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/5/1018/763939.

Frid:2006:EBM

[FPS06] Chris L. J. Frid, Odette A. L. Paramor, and Catherine L. Scott. Ecosystem-based management of fisheries: is science limiting? ICES Journal of Marine Science, 63(9):1567–1572, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/9/1567/696221.

Fu:2001:RSC

[FQS01] Caihong Fu, Terrance J. Quinn II, and Thomas C. Shirley. The role of sex change, growth and mortality in Pandalus population dynamics and management. ICES Journal of Marine Science, 58(3):607–621, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/3/607/810096.

Finlay:2004:RDD

[FR04] Kerri Finlay and John C. Roff. Radiotracer determination of the diet of calanoid copepod nauplii and copepodites in a temperate estuary. ICES Journal of Marine Science, 61(4):552–562, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/4/552/604308.

Fudge:2009:PAA

[FR09] Susan B. Fudge and George A. Rose. Passive- and activeacoustic properties of a spawning Atlantic cod (Gadus morhua) aggregation. ICES Journal of Marine Science, 66(6):1259–1263, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1259/694433.

Francis:2000:LE

[Fra00] R. I. C. Chris Francis. Letter to the Editors. ICES Journal of Marine Science, 57(1):185–187, February 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/1/185/641137.

Francis:2006:MSE

[Fra06] R. I. C. Chris Francis. Measuring the strength of environment– recruitment relationships: the importance of including predictor screening within cross-validations. ICES Journal of Marine Science, 63(4):594–599, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/63/4/594/691541.

Friedland:2003:OTC

[FRC03] Kevin D. Friedland, David G. Reddin, and Martin Castonguay. Ocean thermal conditions in the post-smolt nursery of North American Atlantic salmon. ICES Journal of Marine Science, 60(2):343–355, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/2/343/626629.

Friis-Rodel:2002:RCM

[FRK02] Elisabeth Friis-Rødel and Per Kanneworff. A review of capelin (Mallotus villosus) in Greenland waters. ICES Journal of Marine Science, 59(5):890–896, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/59/5/890/675200.

Frumkes:2002:SCR

[Fru02] D. R. Frumkes. The status of the California Rigs-to-Reefs Programme and the need to limit consumptive fishing activities. ICES Journal of Marine Science, 59(S1):S272–S276, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S272/617961.

Fabi:2002:ABD

[FS02] G. Fabi and A. Sala. An assessment of biomass and diel activity of fish at an artificial reef (Adriatic Sea) using a stationary hydroacoustic technique. ICES Journal of Marine Science, 59 (2):411–420, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/2/411/619665.

Fernandes:2003:AUV

[FSB^{+}03] Paul G. Fernandes, Pete Stevenson, Andrew S. Brierley, Frederick Armstrong, and E. John Simmonds. Autonomous underwater vehicles: future platforms for fisheries acoustics. ICES Journal of Marine Science, 60(3):684–691, ???? 2003. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/3/684/661076.

Fast:2009:DOC

[FSDB09] Mark D. Fast, Mark S. Sokolowski, Keith J. Dunton, and Paul R. Bowser. Dichelesthium oblongum (Copepoda: Dichelesthiidae) infestation in wild-caught Atlantic sturgeon, Acipenser oxyrinchus oxyrinchus. ICES Journal of Marine Science, 66 (10):2141–2147, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/10/2141/679476.

Flostrand:2009:MMP

[FSDC09] Linnea A. Flostrand, Jacob F. Schweigert, Kristen S. Daniel, and Jaclyn S. Cleary. Measuring and modelling Pacific herring spawning-site fidelity and dispersal using tag-recovery dispersal curves. ICES Journal of Marine Science, 66(8):1754– 1761, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/8/1754/675572.

Fabry:2008:IOA

[FSFO08] Victoria J. Fabry, Brad A. Seibel, Richard A. Feely, and James C. Orr. Impacts of ocean acidification on marine fauna and ecosystem processes. ICES Journal of Marine Science, 65 (3):414–432, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/3/414/789605.

Fulton:2005:WEI

[FSP05] Elizabeth A. Fulton, Anthony D. M. Smith, and André E. Punt. Which ecological indicators can robustly detect effects of fishing? ICES Journal of Marine Science, 62(3):540–551, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/3/540/665354.

Freitas:2003:ASC

[FSQ^{+}03] Rosa Freitas, Susana Silva, Victor Quintino, Ana Maria Rodrigues, Karl Rhynas, and William T. Collins. Acoustic seabed classification of marine habitats: studies in the western coastalshelf area of Portugal. ICES Journal of Marine Science, 60 (3):599–608, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/3/599/659720.

Frank:2000:EEW

[FSS00] Kenneth T. Frank, Nancy L. Shackell, and James E. Simon. An evaluation of the Emerald/Western Bank juvenile haddock closed area. ICES Journal of Marine Science, 57(4):1023–1034, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/1023/647245.

Floeter:2005:APS

[FT05] Jens Floeter and Axel Temming. Analysis of prey size preference of North Sea whiting, saithe, and grey gurnard. ICES Journal of Marine Science, 62(5):897–907, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/5/897/856686.

Fernandez-Torquemada:2008:DPO

[FTDVC^{+}08] Yolanda Fernández-Torquemada, Marta Díaz-Valdés, Francisco Colilla, Beatriz Luna, José Luis Sánchez-Lizaso, and Alfonso A. Ramos-Esplá. Descriptors from Posidonia oceanica (L.) Delile meadows in coastal waters of Valencia, Spain, in the context of the EU Water Framework Directive. ICES Journal of Marine Science, 65(8):1492–1497, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/8/1492/714280.

Furness:2002:MII

[Fur02] Robert W. Furness. Management implications of interactions between fisheries and sandeel-dependent seabirds and seals in the North Sea. ICES Journal of Marine Science, 59(2):261–269, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/2/261/619635.

Fjalling:2006:AHD

[FWW06] Arne Fjälling, Magnus Wahlberg, and Håkan Westerberg. Acoustic harassment devices reduce seal interaction in the Baltic salmon-trap, net fishery. ICES Journal of Marine Science, 63 (9):1751–1758, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/9/1751/701455.

Gislason:2000:WDO

[GA00] A. Gislason and O. S. Astthorsson. Winter distribution, ontogenetic migration, and rates of egg production of Calanus finmarchicus southwest of Iceland. ICES Journal of Marine Science, 57(6):1727–1739, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/6/1727/768631.

Goni:2005:AJN

[GA05] Nicolas Goñi and Haritz Arrizabalaga. Analysis of juvenile North Atlantic albacore (Thunnus alalunga) catch per unit effort by surface gears in relation to environmental variables. ICES Journal of Marine Science, 62(7):1475–1482, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/7/1475/660760.

Gaard:2000:SAD

[Gaa00] E. Gaard. Seasonal abundance and development of Calanus finmarchicus in relation to phytoplankton and hydrography on the Faroe Shelf. ICES Journal of Marine Science, 57(6):1605– 1611, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/6/1605/768560.

Guyonnet:2003:BEF

[GAA^{+}03] B. Guyonnet, C. Aliaume, J.-J. Albaret, C. Casellas, A. Zerbi, G. Lasserre, and T. Do Chi. Biology of Ethmalosa fimbriata (Bowdich) and fish diversity in the Ebrie Lagoon (Ivory Coast), a multipolluted environment. ICES Journal of Marine Science, 60(2):259–267, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/2/259/625151.

Goni:2004:RIE

[GAA^{+}04] Raquel Goñi, Sara Adlerstein, Federico Alvarez, Mariano García, Pilar Sánchez, Mario Sbrana, Francesc Maynou, and Claudio Viva. Recruitment indices of European hake, Merluccius merluccius (Linnaeus 1758), in the Northwest Mediterranean based on landings from bottom-trawl multispecies fisheries. ICES Journal of Marine Science, 61(5):760–773, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/5/760/864015.

Grandcourt:2006:FBS

[GAFA06] E. M. Grandcourt, T. Z. Al Abdessalaam, F. Francis, and A. T. Al Shamsi. Fisheries biology of a short-lived tropical species: Gerres longirostris (Lacépède, 1801) in the Arabian Gulf. ICES Journal of Marine Science, 63(3):452–459, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/3/452/720064.

Greenstreet:2006:VAS

[GAM^{+}06] Simon P. R. Greenstreet, Eric Armstrong, Henrik Mosegaard, Henrik Jensen, Iain M. Gibb, Helen M. Fraser, Beth E. Scott, Gayle J. Holland, and Jonathan Sharples. Variation in the abundance of sandeels Ammodytes marinus off southeast Scotland: an evaluation of area-closure fisheries management and stock abundance assessment methods. ICES Journal of Marine Science, 63(8):1530–1550, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/8/1530/714676.

Gislason:2000:LCC

[GAP^{+}00] A. Gislason, O. S. Astthorsson, H. Petursdottir, H. Gudfinnsson, and A. R. Bodvarsdottir. Life cycle of Calanus finmarchicus south of Iceland in relation to hydrography and chlorophyll a. ICES Journal of Marine Science, 57(6):1619–1627, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/6/1619/768573.

Gaspar:2002:IMS

[Gas02] M. Gaspar. Influence of mesh size and tooth spacing on the proportion of damaged organisms in the catches of the Portuguese clam dredge fishery. ICES Journal of Marine Science, 59(6): 1228–1236, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/6/1228/608179.

Gates:2000:IST

[Gat00] J. M. Gates. Input substitution in a trap fishery. ICES Journal of Marine Science, 57(1):89–108, February 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/57/1/89/
641153.

Goodsir:2008:USS

[GAW^{+}08] Freya Goodsir, Michael J. Armstrong, Peter R. Witthames, David L. Maxwell, and Clive J. Fox. The use of species-specific TaqMan probes for identifying early stage gadoid eggs following formaldehyde fixation. ICES Journal of Marine Science, 65(9): 1573–1577, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/9/1573/632418.

Gilbert:2006:GCT

[GAYR06] Antoine Gilbert, Serge Andréfouët, Laurent Yan, and Georges Remoissenet. The giant clam Tridacna maxima communities of three French Polynesia islands: comparison of their population sizes and structures at early stages of their exploitation. ICES Journal of Marine Science, 63(9):1573–1589, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/9/1573/696819.

Godoy:2002:FAE

[GAZ02] Eduardo A. S. Godoy, Tito C. M. Almeida, and Ilana R. Zalmon. Fish assemblages and environmental variables on an artificial reef north of Rio de Janeiro, Brazil. ICES Journal of Marine Science, 59(S1):S138–S143, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S138/617907.

Gerlotto:2006:WAI

[GBBG06] François Gerlotto, Sophie Bertrand, Nicolas Bez, and Mariano Gutierrez. Waves of agitation inside anchovy schools observed with multibeam sonar: a way to transmit information in response to predation. ICES Journal of Marine Science, 63(8): 1405–1417, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/8/1405/711825.

Gascuel:2005:TST

[GBC^{+}05] Didier Gascuel, Yves-Marie Bozec, Emmanuel Chassot, Audrey Colomb, and Martial Laurans. The trophic spectrum: theory and application as an ecosystem indicator. ICES Journal of Marine Science, 62(3):443–452, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/3/443/662465.

Gjosaeter:2002:AMB

[GBT02] Harald Gjøsæter, Bjarte Bogstad, and Sigurd Tjelmeland. Assessment methodology for Barents Sea capelin, Mallotus villosus (Müller). ICES Journal of Marine Science, 59(5):1086–1095, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/5/1086/675163.

Godoy:2002:CAB

[GC02] Eduardo A. S. Godoy and Ricardo Coutinho. Can artificial beds of plastic mimics compensate for seasonal absence of natural beds of Sargassum furcatum? ICES Journal of Marine Science, 59(S1):S111–S115, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/59/suppl/S111/617899.

Garcia:2005:EAF

[GC05a] Serge M. Garcia and Kevern L. Cochrane. Ecosystem approach to fisheries: a review of implementation guidelines. ICES Journal of Marine Science, 62(3):311–318, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/3/311/658728.

Gorska:2005:EIS

[GC05b] Natalia Gorska and Dezhang Chu. On the echo interference in sound backscattering by densely aggregated targets. ICES Journal of Marine Science, 62(4):771–778, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/4/771/676037.

Garcia:2007:FSL

[GC07] Serge M. Garcia and Anthony T. Charles. Fishery systems and linkages: from clockworks to soft watches. ICES Journal of Marine Science, 64(4):580–587, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/580/640385.

Gaspar:2009:CWI

[GCC^{+}09] M. B. Gaspar, S. Carvalho, R. Constantino, J. Tata-Regala, J. Cúrdia, and C. C. Monteiro. Can we infer dredge fishing effort from macrobenthic community structure? ICES Journal of Marine Science, 66(10):2121–2132, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/10/2121/681956.

Goncalves:2009:EBF

[GCM09] Patrícia Gonçalves, Ana Maria Costa, and Alberto G. Murta. Estimates of batch fecundity and spawning fraction for the southern stock of horse mackerel (Trachurus trachurus) in ICES division IXa. ICES Journal of Marine Science, 66(4):617–622, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/4/617/605084.

Gerlotto:2004:TDS

[GCS^{+}04] F. Gerlotto, J. Castillo, A. Saavedra, M. A. Barbieri, M. Espejo, and P. Cotel. Three-dimensional structure and avoidance behaviour of anchovy and common sardine schools in central southern Chile. ICES Journal of Marine Science, 61(7):1120–1126, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/7/1120/879087.

Gifford:2009:FDM

[GCS09] Dian J. Gifford, Jeremy S. Collie, and John H. Steele. Functional diversity in a marine fish community. ICES Journal of Marine Science, 66(5):791–796, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/5/791/663460.

Garcia:2005:ESF

[GD05] S. M. Garcia and J. I. De Leiva Moreno. Evolution of the state of fish stocks in the Northeast Atlantic within a precautionary framework, 1970–2003: a synoptic evaluation. ICES Journal of Marine Science, 62(8):1603–1605, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/8/1603/790439.

Gjosaeter:2002:GBS

[GDH02] Harald Gjøsæter, Padmini Dalpadado, and Arne Hassel. Growth of Barents Sea capelin (Mallotus villosus) in relation to zooplankton abundance. ICES Journal of Marine Science, 59 (5):959–967, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/5/959/675214.

Gaertner:2004:ANL

[GDL04] Daniel Gaertner and Michel Dreyfus-Leon. Analysis of nonlinear relationships between catch per unit effort and abundance in a tuna purse-seine fishery simulated with artificial neural networks. ICES Journal of Marine Science, 61(5):812–820, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/5/812/867451.

Geffen:2009:AHB

[Gef09] Audrey J. Geffen. Advances in herring biology: from simple to complex, coping with plasticity and adaptability. ICES Journal of Marine Science, 66(8):1688–1695, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/8/1688/672963.

Griffiths:2001:RNA

[GEM01] Gwyn Griffiths, Peter Enoch, and Nicholas W. Millard. On the radiated noise of the autosub autonomous underwater vehicle. ICES Journal of Marine Science, 58(6):1195–1200, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/6/1195/641551.

Groenewold:2000:EBS

[GF00] Stefan Groenewold and Mark Fonds. Effects on benthic scavengers of discards and damaged benthos produced by the beamtrawl fishery in the southern North Sea. ICES Journal of Marine Science, 57(5):1395–1406, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/5/1395/660942.

Gillis:2001:IEF

[GF01] D. M. Gillis and K. T. Frank. Influence of environment and fleet dynamics on catch rates of eastern Scotian Shelf cod through the early 1980s. ICES Journal of Marine Science, 58(1):61–69, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/1/61/603553.

Gorokhova:2004:PHC

[GFH04] Elena Gorokhova, Towe Fagerberg, and Sture Hansson. Predation by herring (Clupea harengus) and sprat (Sprattus sprattus) on Cercopagis pengoi in a western Baltic Sea bay. ICES Journal of Marine Science, 61(6):959–965, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/6/959/678475.

Graham:2007:FPG

[GFKM07] Norman Graham, Richard S. T. Ferro, William A. Karp, and Philip MacMullen. Fishing practice, gear design, and the ecosystem approach — three case studies demonstrating the effect of management strategy on gear selectivity and discards. ICES Journal of Marine Science, 64(4):744–750, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/4/744/643088.

Greenstreet:2009:UMA

[GFP09] Simon P. R. Greenstreet, Helen M. Fraser, and Gerjan J. Piet. Using MPAs to address regional-scale ecological objectives in the North Sea: modelling the effects of fishing effort displacement. ICES Journal of Marine Science, 66(1):90–100, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/1/90/635230.

Glass:2004:ISB

[GG04] Chris Glass and John Gunn. Influence of social behaviour and behavioural interactions in understanding temporal and spatial dynamics and their effect on availability and catchability. ICES Journal of Marine Science, 61(7):1236–1237, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1236/882374.

Giesecke:2008:RFS

[GG08] Ricardo Giesecke and Humberto E. González. Reproduction and feeding of Sagitta enflata in the Humboldt Current system off Chile. ICES Journal of Marine Science, 65(3):361–370, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/3/361/788381.

Green:2009:SSC

[GG09] Bridget S. Green and Caleb Gardner. Surviving a sea-change: survival of southern rock lobster (Jasus edwardsii) translocated to a site of fast growth. ICES Journal of Marine Science, 66 (4):656–664, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/4/656/603251.

Gilkinson:2005:IIR

[GGM^{+}05] Kent D. Gilkinson, Donald C. Gordon, Kevin G. MacIsaac, David L. McKeown, Ellen L. R. Kenchington, Cynthia Bourbonnais, and W. Peter Vass. Immediate impacts and recovery trajectories of macrofaunal communities following hydraulic clam dredging on banquereau, eastern Canada. ICES Journal of Marine Science, 62(5):925–947, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/5/925/856941.

Gestal:2007:AOP

[GGP07] Camino Gestal, Angel Guerra, and Santiago Pascual. Aggregata octopiana (Protista: Apicomplexa): a dangerous pathogen during commercial Octopus vulgaris ongrowing. ICES Journal of Marine Science, 64(9):1743–1748, December 2007. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/9/1743/786332.

Gonzalez:2004:CCT

[GGV^{+}04] H. E. González, R. Giesecke, C. A. Vargas, M. Pavez, J. Iriarte, P. Santibáñez, L. Castro, R. Escribano, and F. Pagès. Carbon cycling through the pelagic foodweb in the northern Humboldt Current off Chile (23^{o}S). ICES Journal of Marine Science, 61 (4):572–584, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/4/572/604594.

Gaard:2000:VAC

[GH00] E. Gaard and B. Hansen. Variations in the advection of Calanus finmarchicus onto the Faroe Shelf. ICES Journal of Marine Science, 57(6):1612–1618, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/6/1612/768567.

Gauthier:2004:PAD

[GH04] Stéphane Gauthier and John K. Horne. Potential acoustic discrimination within boreal fish assemblages. ICES Journal of Marine Science, 61(5):836–845, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/5/836/868591.

Gray:2007:ESN

[GH07] Tim S. Gray and Jenny Hatchard. Environmental stewardship as a new form of fisheries governance. ICES Journal of Marine Science, 64(4):786–792, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/64/4/786/638893.

Grandcourt:2008:RSA

[GHBR08] E. M. Grandcourt, T. Hecht, A. J. Booth, and J. Robinson. Retrospective stock assessment of the Emperor red snapper (Lutjanus sebae) on the Seychelles Bank between 1977 and 2006. ICES Journal of Marine Science, 65(6):889–898, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/6/889/601505.

Guethmundsson:2009:ASC

[GHC09] Kristinn Gudmundsson, Mike R. Heath, and Elizabeth D. Clarke. Average seasonal changes in chlorophyll a in Icelandic waters. ICES Journal of Marine Science, 66(10):2133–2140, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/10/2133/682550.

Graham:2009:TEA

[GHD^{+}09] Isla M. Graham, Robert N. Harris, Becks Denny, Dan Fowden, and Dave Pullan. Testing the effectiveness of an acoustic deterrent device for excluding seals from Atlantic salmon rivers in Scotland. ICES Journal of Marine Science, 66(5):860–864, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/5/860/664210.

Greenstreet:2009:MDF

[GHFA09] S. P. R. Greenstreet, G. J. Holland, T. W. K. Fraser, and V. J. Allen. Modelling demersal fishing effort based on landings and days absence from port, to generate indicators of “activity”. ICES Journal of Marine Science, 66(5):886–901, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/5/886/663115.

Godo:2004:BMS

[GHI^{+}04] Olav Rune Godø, Vidar Hjellvik, Svein A. Iversen, Aril Slotte, Eirik Tenningen, and Terje Torkelsen. Behaviour of mackerel schools during summer feeding migration in the Norwegian Sea, as observed from fishing vessel sonars. ICES Journal of Marine Science, 61(7):1093–1099, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/7/1093/878526.

Gurshin:2009:MAB

[GJH^{+}09] Christopher W. D. Gurshin, J. Michael Jech, W. Huntting Howell, Thomas C. Weber, and Larry A. Mayer. Measurements of acoustic backscatter and density of captive Atlantic cod with synchronized 300-kHz multibeam and 120-kHz splitbeam echosounders. ICES Journal of Marine Science, 66(6): 1303–1309, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1303/691468.

Gabbasov:2008:MMB

[GJL08] Mars B. Gabbasov, Nurbolat Zh. Jaichibekov, and Daniel V. Lebedev. A mathematical model of biological resource dynamics, using Caspian/Ural sturgeon as a case study. ICES Journal of Marine Science, 65(1):103–110, January 2008. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/1/103/613513.

Graham:2004:RTA

[GJR04] N. Graham, E. G. Jones, and D. G. Reid. Review of technological advances for the study of fish behaviour in relation to demersal fishing trawls. ICES Journal of Marine Science, 61 (7):1036–1043, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/7/1036/877540.

Griffiths:2009:TSR

[GKFM09] Shane P. Griffiths, Petra M. Kuhnert, Gary F. Fry, and Fiona J. Manson. Temporal and size-related variation in the diet, consumption rate, and daily ration of mackerel tuna (Euthynnus affinis) in neritic waters of eastern Australia. ICES Journal of Marine Science, 66(4):720–733, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/4/720/604888.

Gorska:2007:ABS

[GKO07] Natalia Gorska, Rolf J. Korneliussen, and Egil Ona. Acoustic backscatter by schools of adult Atlantic mackerel. ICES Journal of Marine Science, 64(6):1145–1151, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/6/1145/615444.

Gilbey:2005:NDM

[GKOV05] John Gilbey, David Knox, Martha O’Sullivan, and Eric Verspoor. Novel DNA markers for rapid, accurate, and costeffective discrimination of the continental origin of Atlantic salmon (Salmo salar L.). ICES Journal of Marine Science, 62 (8):1606–1616, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/8/1606/791241.

Garrison:2000:FES

[GL00] Lance P. Garrison and Jason S. Link. Fishing effects on spatial distribution and trophic guild structure of the fish community in the Georges Bank region. ICES Journal of Marine Science, 57(3):723–730, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/723/635964.

Guillard:2004:SSTa

[GLDB04a] J. Guillard, A. Lebourges-Dhaussy, and P. Brehmer. Simultaneous Sv and TS measurements on Young-of-the-Year (YOY) freshwater fish using three frequencies. ICES Journal of Marine Science, 61(2):267–273, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/61/2/267/621727.

Guillard:2004:SSTb

[GLDB04b] J. Guillard, A. Lebourges-Dhaussy, and P. Brehmer. Simultaneous Sv and TS measurements on Young-of-the-Year (YOY) freshwater fish using three frequencies. ICES Journal of Marine Science, 61(5):868–869, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/61/5/868/869066.

Ghertsos:2000:ISS

[GLDD00] Konstantinos Ghertsos, Christophe Luczak, Jean-Marie Dewarumez, and Jean-Claude Dauvin. Influence of spatial scales of observation on temporal change in diversity and trophic structure of fine-sand communities from the English Channel and the southern North Sea. ICES Journal of Marine Science, 57(5): 1481–1487, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/5/1481/660964.

Goudey:2001:MEE

[GLKPCP01] C. A. Goudey, G. Loverich, H. Kite-Powell, and B. A. Costa-Pierce. Mitigating the environmental effects of mariculture through single-point moorings (SPMs) and drifting cages. ICES Journal of Marine Science, 58(2):497–503, ???? 2001. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/2/497/617860.

Gordoa:2006:BCI

[GLR06] Ana Gordoa, Heinrich Lesch, and Silvia Rodergas. Bycatch: complementary information for understanding fish behaviour. Namibian Cape hake (M. capensis and M. paradoxus) as a case study. ICES Journal of Marine Science, 63(8):1513–1519, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/8/1513/714039.

Gaspar:2003:CDM

[GLS^{+}03] M. B. Gaspar, F. Leitão, M. N. Santos, L. Chícharo, M. D. Dias, A. Chícharo, and C. C. Monteiro. A comparison of direct macrofaunal mortality using three types of clam dredges. ICES Journal of Marine Science, 60(4):733–742, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/4/733/691398.

Guijarro:2006:SDS

[GM06] Beatriz Guijarro and Enric Massutí. Selectivity of diamond- and square-mesh codends in the deepwater crustacean trawl fishery off the Balearic Islands (western Mediterranean). ICES Journal of Marine Science, 63(1):52–67, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/1/52/625324.

Glass:2007:FFI

[GM07] Christopher W. Glass and Christopher A. Manning. Fishery forum on integrating fishing and ecosystem conservation: the way forward. ICES Journal of Marine Science, 64(8):1614– 1615, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/8/1614/614021.

Goetz:2006:UGG

[GMGN06] Frederick W. Goetz, Linda McCauley, Giles W. Goetz, and Birgitta Norberg. Using global genome approaches to address problems in cod mariculture. ICES Journal of Marine Science, 63 (2):393–399, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/2/393/642481.

Giannoulaki:2006:ECT

[GMKS06] Marianna Giannoulaki, Athanassios Machias, Constantin Koutsikopoulos, and Stylianos Somarakis. The effect of coastal topography on the spatial structure of anchovy and sardine. ICES Journal of Marine Science, 63(4):650–662, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/4/650/692418.

Gerritsen:2006:SMC

[GML06] Hans D. Gerritsen, David McGrath, and Colm Lordan. A simple method for comparing age–length keys reveals significant regional differences within a single stock of haddock (Melanogrammus aeglefinus). ICES Journal of Marine Science, 63(6): 1096–1100, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/6/1096/616149.

Garrido:2008:HMT

[GMM^{+}08] Susana Garrido, Alberto G. Murta, Ana Moreira, Maria João Ferreira, and Maria Manuel Angélico. Horse mackerel (Trachurus trachurus) stomach fullness off Portugal: index calibration and spatio-temporal variations in feeding intensity. ICES Journal of Marine Science, 65(9):1662–1669, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/9/1662/631710.

Gentleman:2008:MCD

[GNC08] W. C. Gentleman, A. B. Neuheimer, and R. G. Campbell. Modelling copepod development: current limitations and a new realistic approach. ICES Journal of Marine Science, 65(3):399–413, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/3/399/789374.

Gorska:2003:MAE

[GO03a] Natalia Gorska and Egil Ona. Modelling the acoustic effect of swimbladder compression in herring. ICES Journal of Marine Science, 60(3):548–554, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/60/3/548/659075.

Gorska:2003:MES

[GO03b] Natalia Gorska and Egil Ona. Modelling the effect of swimbladder compression on the acoustic backscattering from herring at normal or near-normal dorsal incidences. ICES Journal of Marine Science, 60(6):1381–1391, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/6/1381/654737.

Gordoa:2009:DAB

[GOA^{+}09] Ana Gordoa, Maria Pilar Olivar, Raquel Arevalo, Jordi Viñas, Balbina Molí, and Xenia Illas. Determination of Atlantic bluefin tuna (Thunnus thynnus) spawning time within a transport cage in the western Mediterranean. ICES Journal of Marine Science, 66(10):2205–2210, December 2009. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/66/10/2205/683204.

Gorska:2005:ABA

[GOK05] Natalia Gorska, Egil Ona, and Rolf Korneliussen. Acoustic backscattering by Atlantic mackerel as being representative of fish that lack a swimbladder. backscattering by individual fish. ICES Journal of Marine Science, 62(5):984–995, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/5/984/858560.

Gudmundsdottir:2007:EYC

[GOS07] Asta Gudmundsdottir, Gudmundur J. Oskarsson, and Sveinn Sveinbjörnsson. Estimating year-class strength of Icelandic summer-spawning herring on the basis of two survey methods. ICES Journal of Marine Science, 64(6):1182–1190, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/6/1182/614765.

Greene:2000:RCF

[GP00] C. H. Greene and A. J. Pershing. The response of Calanus finmarchicus populations to climate variability in the Northwest Atlantic: basin-scale forcing associated with the North Atlantic Oscillation. ICES Journal of Marine Science, 57(6): 1536–1544, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/6/1536/768532.

Godo:2009:DMS

[GPP09] Olav Rune Godø, Ruben Patel, and Geir Pedersen. Diel migration and swimbladder resonance of small fish: some implications for analyses of multifrequency echo data. ICES Journal of Marine Science, 66(6):1143–1148, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1143/694543.

Gislason:2008:CNS

[GPRD08] Henrik Gislason, John G. Pope, Jake C. Rice, and Niels Daan. Coexistence in North Sea fish communities: implications for growth and natural mortality. ICES Journal of Marine Science, 65(4):514–530, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/4/514/638453.

Grosjean:2004:EMI

[GPWG04] Philippe Grosjean, Marc Picheral, Caroline Warembourg, and Gabriel Gorsky. Enumeration, measurement, and identification of net zooplankton samples using the ZOOSCAN digital imaging system. ICES Journal of Marine Science, 61(4):518–525, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/4/518/603573.

Greve:2005:PNS

[GPZ^{+}05] Wulf Greve, Sabine Prinage, Heike Zidowitz, Jutta Nast, and Frank Reiners. On the phenology of North Sea ichthyoplankton. ICES Journal of Marine Science, 62(7):1216–1223, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/7/1216/654731.

Gonzalez-Quiros:2003:IDP

[GQCÁMI03] R. González-Quirós, J. Cabal, F. Álvarez-Marqués, and A. Isla. Ichthyoplankton distribution and plankton production related to the shelf break front at the Avilés canyon. ICES Journal of Marine Science, 60(2):198–210, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/2/198/623687.

Gauthier:2001:TSE

[GR01] Stéphane Gauthier and George A. Rose. Target strength of encaged Atlantic redfish (Sebastes spp.). ICES Journal of Marine Science, 58(3):562–568, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/58/3/562/810067.

Gauthier:2002:STS

[GR02] Stéphane Gauthier and George A. Rose. In situ target strength studies on Atlantic redfish (Sebastes spp.). ICES Journal of Marine Science, 59(4):805–815, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/4/805/676807.

Gauthier:2005:DVM

[GR05] Stéphane Gauthier and George A. Rose. Diel vertical migration and shoaling heterogeneity in Atlantic redfish: effects on acoustic and bottom-trawl surveys. ICES Journal of Marine Science, 62(1):75–85, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/1/75/669517.

Greenstreet:2006:IHN

[GR06] Simon P. R. Greenstreet and Stuart I. Rogers. Indicators of the health of the North Sea fish community: identifying reference levels for an ecosystem approach to management. ICES Journal of Marine Science, 63(4):573–593, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/4/573/690519.

Garcia:2006:ESD

[GRE06] Elena Guijarro Garcia, Stefán Áki Ragnarsson, and Hrafnkell Eiríksson. Effects of scallop dredging on macrobenthic communities in west Iceland. ICES Journal of Marine Science, 63 (3):434–443, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/3/434/719564.

Greenstreet:2008:BNS

[Gre08] Simon P. R. Greenstreet. Biodiversity of North Sea fish: why do the politicians care but marine scientists appear oblivious to this issue? ICES Journal of Marine Science, 65(8):1515– 1519, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/8/1515/712396.

Groger:2007:SRA

[GRMR07] Joachim P. Gröger, Rodney A. Rountree, Martin Missong, and Hans-Joachim Rätz. A stock rebuilding algorithm featuring risk assessment and an optimization strategy of single or multispecies fisheries. ICES Journal of Marine Science, 64(6):1101– 1115, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/6/1101/614574.

Grout:2006:IBS

[Gro06] Douglas E. Grout. Interactions between striped bass (Morone saxatilis) rebuilding programmes and the conservation of Atlantic salmon (Salmo salar) and other anadromous fish species in the USA. ICES Journal of Marine Science, 63(7):1346–1352, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/7/1346/760651.

Gislason:2000:EEF

[GS00] Henrik Gislason and Michael M. Sinclair. Ecosystem effects of fishing. ICES Journal of Marine Science, 57(3):466–467, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/3/466/635911.

Garthe:2003:UDO

[GS03] Stefan Garthe and Birgit Scherp. Utilization of discards and offal from commercial fisheries by seabirds in the Baltic Sea. ICES Journal of Marine Science, 60(5):980–989, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/5/980/770761.

Gomes:2001:SPG

[GSdFB01] Manuel C. Gomes, Ester Serrão, and Maria de Fátima Borges. Spatial patterns of groundfish assemblages on the continental shelf of Portugal. ICES Journal of Marine Science, 58 (3):633–647, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/3/633/810108.

Glover:2003:DSA

[GSN^{+}03] K. A. Glover, Ø. Skaala, F. Nilsen, R. Olsen, A. J. Teale, and J. B. Taggart. Differing susceptibility of anadromous brown trout (Salmo trutta L.) populations to salmon louse (Lepeophtheirus salmonis (Krøyer, 1837)) infection. ICES Journal of Marine Science, 60(5):1139–1148, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/5/1139/767852.

Glover:2008:GAI

[GSS08] K. A. Glover, O. T. Skilbrei, and Ø. Skaala. Genetic assignment identifies farm of origin for Atlantic salmon Salmo salar escapees in a Norwegian fjord. ICES Journal of Marine Science, 65(6): 912–920, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/6/912/600179.

Gislason:2000:SOI

[GSSO00] Henrik Gislason, Michael Sinclair, Keith Sainsbury, and Robert O’Boyle. Symposium overview: incorporating ecosystem objectives within fisheries management. ICES Journal of Marine Science, 57(3):468–475, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/3/468/635913.

Garatun-Tjeldsto:2006:FIJ

[GTOJA06] Oddvard Garatun-Tjeldstø, Håkon Otterå, Kåre Julshamn, and Erland Austreng. Food ingestion in juvenile cod estimated by inert lanthanide markers — effects of food particle size. ICES Journal of Marine Science, 63(2):311–319, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/2/311/640516.

Gudmundsson:2004:TSA

[Gud04] G. Gudmundsson. Time-series analysis of abundance indices of young fish. ICES Journal of Marine Science, 61(2):176–183, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/2/176/620104.

Gudmundsdottir:2002:PTA

[GV02] Asta Gudmundsdottir and Hjálmar Vilhjálmsson. Predicting total allowable catches for Icelandic capelin, 1978–2001. ICES Journal of Marine Science, 59(5):1105–1115, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/5/1105/675174.

Gosse:2004:VCA

[GW04] Karen R. Gosse and Joseph S. Wroblewski. Variant colourations of Atlantic cod (Gadus morhua) in Newfoundland and Labrador nearshore waters. ICES Journal of Marine Science, 61(5):752–759, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/5/752/863823.

Guttormsen:2009:SMC

[GW09] Michael A. Guttormsen and Christopher D. Wilson. In situ measurements of capelin (Mallotus villosus) target strength in the North Pacific Ocean. ICES Journal of Marine Science, 66 (2):258–263, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/2/258/594768.

Gaughan:2001:LBF

[GWF01] D. J. Gaughan, K. V. White, and W. J. Fletcher. The links between functionally distinct adult assemblages of Sardinops sagax: larval advection across management boundaries. ICES Journal of Marine Science, 58(3):597–606, ???? 2001. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/3/597/810091.

Galley:2006:CMO

[GWG06] Elizabeth A. Galley, Peter J. Wright, and Fiona M. Gibb. Combined methods of otolith shape analysis improve identification of spawning areas of Atlantic cod. ICES Journal of Marine Science, 63(9):1710–1717, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/9/1710/700248.

Goetz:2008:IBS

[GWSV08] Sabine Goetz, Matthias Wolff, Wolfgang Stotz, and Mario J. Villegas. Interactions between the South American sea lion (Otaria flavescens) and the artisanal fishery off coquimbo, northern Chile. ICES Journal of Marine Science, 65(9):1739–1746, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/9/1739/630972.

Glass:2007:FTC

[GWvM07] Christopher W. Glass, Stephen J. Walsh, and Bob van Marlen. Fishing technology in the 21st century: integrating fishing and ecosystem conservation. ICES Journal of Marine Science, 64(8):

1499–1502, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/8/1499/613997.

Grove:2002:ENN

[GZND02] Robert S. Grove, Karel Zabloudil, Tim Norall, and Lawrence Deysher. Effects of El Niño events on natural kelp beds and artificial reefs in southern California. ICES Journal of Marine Science, 59(S1):S330–S337, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S330/617987.

Ganzedo:2009:WDT

[GZS^{+}09] U. Ganzedo, E. Zorita, A. P. Solari, G. Chust, A. Santana del Pino, J. Polanco, and J. J. Castro. What drove tuna catches between 1525 and 1756 in southern Europe? ICES Journal of Marine Science, 66(7):1595–1604, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/7/1595/657212.

Huckstadt:2003:BSS

[HA03] L. A. Hückstädt and T. Antezana. Behaviour of the southern sea lion (Otaria flavescens) and consumption of the catch during purse-seining for jack mackerel (Trachurus symmetricus) off central Chile. ICES Journal of Marine Science, 60(5):1003–1011, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/5/1003/763517.

Hubley:2008:SBM

[HAG^{+}08] P. Bradford Hubley, Peter G. Amiro, A. Jamie F. Gibson, Gilles L. Lacroix, and Anna M. Redden. Survival and behaviour of migrating Atlantic salmon (Salmo salar L.) kelts in river, estuarine, and coastal habitat. ICES Journal of Marine Science, 65 (9):1626–1634, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/9/1626/628974.

Hall:2001:MPA

[Hal01] Stephen J. Hall. Marine protected areas: a picture paints a thousand words. ICES Journal of Marine Science, 58(4):738–739, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/4/738/630193.

Hamre:2006:NCG

[Ham06] Kristin Hamre. Nutrition in cod (Gadus morhua) larvae and juveniles. ICES Journal of Marine Science, 63(2):267–274, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/2/267/639707.

Hovde:2002:SSO

[HAN02] S. C. Hovde, O. T. Albert, and E. M. Nilssen. Spatial, seasonal and ontogenetic variation in diet of Northeast Arctic Greenland halibut (Reinhardtius hippoglossoides). ICES Journal of Marine Science, 59(2):421–437, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/2/421/619667.

Hansen:2006:MSF

[Han06] Lars P. Hansen. Migration and survival of farmed Atlantic salmon (Salmo salar L.) released from two Norwegian fish farms. ICES Journal of Marine Science, 63(7):1211–1217, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/7/1211/754971.

Harris:2000:FGR

[Har00] R. Harris. Feeding, growth, and reproduction in the genus Calanus. ICES Journal of Marine Science, 57(6):1708–1726, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/6/1708/768625.

Harbitz:2007:ESP

[Har07] Alf Harbitz. Estimation of shrimp (Pandalus borealis) carapace length by image analysis. ICES Journal of Marine Science, 64 (5):939–944, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/5/939/640689.

Hoolihan:2006:MDA

[HAvH06] John P. Hoolihan, Prem Anandh, and Lynne van Herwerden. Mitochondrial DNA analyses of narrow-barred Spanish mackerel (Scomberomorus commerson) suggest a single genetic stock in the ROPME sea area (Arabian Gulf, Gulf of Oman, and Arabian Sea). ICES Journal of Marine Science, 63(6):1066–1074, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/6/1066/615488.

He:2007:RCS

[HB07] Pingguo He and Vincent Balzano. Reducing the catch of small shrimps in the Gulf of Maine pink shrimp fishery with a sizesorting grid device. ICES Journal of Marine Science, 64(8): 1551–1557, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/8/1551/613021.

Higgason:2009:LSM

[HB09] Kelley D. Higgason and Maria Brown. Local solutions to manage the effects of global climate change on a marine ecosystem: a process guide for marine resource managers. ICES Journal of Marine Science, 66(7):1640–1646, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/7/1640/658770.

Haya:2001:EIC

[HBC01] K. Haya, L. E. Burridge, and B. D. Chang. Environmental impact of chemical wastes produced by the salmon aquaculture industry. ICES Journal of Marine Science, 58(2):492–496, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/2/492/617857.

Hanchet:2005:DEC

[HBD05] Stuart M. Hanchet, Ron G. Blackwell, and Alistair Dunn. Development and evaluation of catch per unit effort indices for southern blue whiting (Micromesistius australis) on the Campbell Island Rise, New Zealand. ICES Journal of Marine Science, 62(6):1131–1138, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/6/1131/617644.

Heath:2004:CEW

[HBG^{+}04] Michael R. Heath, Peter R. Boyle, Astthor Gislason, William S. C. Gurney, Stephen J. Hay, Erica J. H. Head, Steven Holmes, Anna Ingvarsdóttir, Sigrun H. Jónasdóttir, Pennie Lindeque, Raymond T. Pollard, Jens Rasmussen, Kelvin Richards, Katherine Richardson, Gary Smerdon, and Douglas Speirs. Comparative ecology of over-wintering Calanus finmarchicus in the northern North Atlantic, and implications for life-cycle patterns. ICES Journal of Marine Science, 61(4):698–708, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/4/698/608060.

Hinz:2006:HAP

[HBS^{+}06] Hilmar Hinz, Melanie Bergmann, Richard Shucksmith, Michel J. Kaiser, and Stuart I. Rogers. Habitat association of plaice, sole, and lemon sole in the English Channel. ICES Journal of Marine Science, 63(5):912–927, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/63/5/912/663966.

Host:2002:GDM

[HBST02] Gudmund Høst, Erlend Berg, Tore Schweder, and Sigurd Tjelmeland. A Gamma/Dirichlet model for estimating uncertainty in age-specific abundance of Norwegian spring-spawning herring. ICES Journal of Marine Science, 59(4):737–748, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/4/737/676783.

Hollowed:2009:FMF

[HBW^{+}09] Anne Babcock Hollowed, Nicholas A. Bond, Thomas K. Wilderbuer, William T. Stockhausen, Z. Teresa A’mar, Richard J. Beamish, James E. Overland, and Michael J. Schirripa. A framework for modelling fish and shellfish responses to future climate change. ICES Journal of Marine Science, 66(7):1584–1594, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/7/1584/657744.

Hart:2009:EBG

[HC09] Deborah R. Hart and Antonie S. Chute. Estimating von Bertalanffy growth parameters from growth increment data using a linear mixed-effects model, with an application to the sea scallop Placopecten magellanicus. ICES Journal of Marine Science, 66 (10):2165–2175, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/10/2165/679802.

Harvey:2003:EMF

[HCE^{+}03] Chris J. Harvey, Sean P. Cox, Timothy E. Essington, Sture Hansson, and James F. Kitchell. An ecosystem model of food web and fisheries interactions in the Baltic Sea. ICES Journal of Marine Science, 60(5):939–950, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/5/939/769206.

Holmes:2006:APF

[HCEM06] John A. Holmes, George M. W. Cronkite, Hermann J. Enzenhofer, and Timothy J. Mulligan. Accuracy and precision of fishcount data from a “dual-frequency identification sonar” (DID-SON) imaging system. ICES Journal of Marine Science, 63 (3):543–555, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/3/543/723100.

Hampel:2003:HVD

[HCV03] H. Hampel, A. Cattrijsse, and M. Vincx. Habitat value of a developing estuarine brackish marsh for fish and macrocrustaceans. ICES Journal of Marine Science, 60(2):278–289, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/2/278/625442.

Hoffmann:2000:ECA

[HD00] E. Hoffmann and P. Dolmer. Effect of closed areas on distribution of fish and epibenthos. ICES Journal of Marine Science, 57(5):1310–1314, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/5/1310/660910.

Heino:2002:ERN

[HDG02] Mikko Heino, Ulf Dieckmann, and Olav Rune Godø. Estimating reaction norms for age and size at maturation with reconstructed immature size distributions: a new technique illustrated by application to Northeast Arctic cod. ICES Journal of Marine Science, 59(3):562–575, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/59/3/562/610836.

Holliday:2009:HFA

[HDG^{+}09] D. V. Holliday, P. L. Donaghay, C. F. Greenlaw, J. M. Napp, and J. M. Sullivan. High-frequency acoustics and bio-optics in ecosystems research. ICES Journal of Marine Science, 66 (6):974–980, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/974/696674.

Heslenfeld:2008:OEQ

[HE08] Peter Heslenfeld and E. Lisette Enserink. OSPAR ecological quality objectives: the utility of health indicators for the North Sea. ICES Journal of Marine Science, 65(8):1392–1397, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/8/1392/714770.

Heath:2000:CAC

[Hea00a] M. Heath. Comparative analysis of Calanus finmarchicus demography at locations around the Northeast Atlantic. ICES Journal of Marine Science, 57(6):1562–1580, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/6/1562/768543.

Heath:2000:WDC

[Hea00b] M. Heath. Winter distribution of Calanus finmarchicus in the Northeast Atlantic. ICES Journal of Marine Science, 57(6): 1628–1635, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/6/1628/768583.

Heath:2005:CSFa

[Hea05a] Michael R. Heath. Changes in the structure and function of the North Sea fish foodweb, 1973–2000, and the impacts of fishing and climate. ICES Journal of Marine Science, 62(5):847–868, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/5/847/855875.

Heath:2005:CSFb

[Hea05b] Michael R. Heath. Changes in the structure and function of the North Sea fish foodweb, 1973–2000, and the impacts of fishing and climate. ICES Journal of Marine Science, 62(6):1202, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/6/1202/618575.

Heath:2005:RVT

[Hea05c] Michael R. Heath. Regional variability in the trophic requirements of shelf sea fisheries in the Northeast Atlantic, 1973–2000. ICES Journal of Marine Science, 62(7):1233–1244, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/7/1233/655115.

Heath:2007:CZE

[Hea07] Michael R. Heath. The consumption of zooplankton by early life stages of fish in the North Sea. ICES Journal of Marine Science, 64(9):1650–1663, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/9/1650/786613.

Herrera:2002:OFC

[HEGH02] R. Herrera, F. Espino, M. Garrido, and R. J. Haroun. Observations on fish colonization and predation on two artificial reefs in the Canary Islands. ICES Journal of Marine Science, 59 (S1):S69–S73, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S69/618008.

Hansen:2005:CEP

[HEGL05] Bogi Hansen, Sólvá K. Eliasen, Eilif Gaard, and Karin M. H. Larsen. Climatic effects on plankton and productivity on the Faroe Shelf. ICES Journal of Marine Science, 62(7):1224–1232, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/7/1224/654931.

Helle:2000:DCS

[Hel00] K. Helle. Distribution of the copepodite stages of Calanus finmarchicus from Lofoten to the Barents Sea in July 1989. ICES Journal of Marine Science, 57(6):1636–1644, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/6/1636/768590.

Helvey:2002:SCO

[Hel02] Mark Helvey. Are southern California oil and gas platforms essential fish habitat? ICES Journal of Marine Science, 59 (S1):S266–S271, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S266/617958.

Hendrickson:2004:PBN

[Hen04] Lisa C. Hendrickson. Population biology of northern shortfin squid (Illex illecebrosus) in the Northwest Atlantic Ocean and initial documentation of a spawning area. ICES Journal of Marine Science, 61(2):252–266, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/61/2/252/621327.

Herrero:2004:RSF

[Her04] Ines Herrero. Risk and strategy of fishers alternatively exploiting sea bream and tuna in the Gibraltar Strait from an efficiency perspective. ICES Journal of Marine Science, 61(2):211–217, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/2/211/620802.

Hoff:2008:MCH

[HF08a] A. Hoff and H. Frost. Modelling combined harvest and effort regulations: the case of the Dutch beam trawl fishery for plaice and sole in the North Sea. ICES Journal of Marine Science, 65 (6):822–831, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/6/822/600272.

Hopkins:2008:MOV

[HF08b] Grant A. Hopkins and Barrie M. Forrest. Management options for vessel hull fouling: an overview of risks posed by inwater cleaning. ICES Journal of Marine Science, 65(5):811–815, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/5/811/713497.

Hindar:2006:GEE

[HFMD06] Kjetil Hindar, Ian A. Fleming, Philip McGinnity, and Ola Diserud. Genetic and ecological effects of salmon farming on wild salmon: modelling from experimental results. ICES Journal of Marine Science, 63(7):1234–1247, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/7/1234/755889.

Hamazaki:2005:ABS

[HFWB05] Toshihide Hamazaki, Lowell Fair, Leslie Watson, and Elisabeth Brennan. Analyses of Bering Sea bottom-trawl surveys in Norton Sound: absence of regime shift effect on epifauna and demersal fish. ICES Journal of Marine Science, 62(8):1597–1602, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/8/1597/790193.

Heilskov:2001:EBF

[HH01] Anna C. Heilskov and Marianne Holmer. Effects of benthic fauna on organic matter mineralization in fish-farm sediments: importance of size and abundance. ICES Journal of Marine Science, 58(2):427–434, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/2/427/617844.

Hazen:2003:MEE

[HH03] Elliott L. Hazen and John K. Horne. A method for evaluating the effects of biological factors on fish target strength. ICES Journal of Marine Science, 60(3):555–562, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/3/555/659152.

Hazen:2004:CMM

[HH04] Elliott L. Hazen and John K. Horne. Comparing the modelled and measured target-strength variability of walleye pollock, Theragra chalcogramma. ICES Journal of Marine Science, 61(3):363–377, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/3/363/671467.

Harris:2009:CWA

[HHAB09] Peter T. Harris, Andrew D. Heap, Tara J. Anderson, and Brendan Brooke. Comment on: Williams et al. (2009) “Australia’s deep-water reserve network: implications of false homogeneity for classifying abiotic surrogates of biodiversity”.

ICES Journal of Marine Science, 66: 214–224. ICES Journal of Marine Science, 66(10):2082–2085, December 2009. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/

66/10/2082/682472. See [WBK^{+}09b, WBK09a].

Harms:2000:MNA

[HHB^{+}00] I. H. Harms, M. R. Heath, A. D. Bryant, J. O. Backhaus, and D. A. Hainbucher. Modelling the Northeast Atlantic circulation: implications for the spring invasion of shelf regions by Calanus finmarchicus. ICES Journal of Marine Science, 57(6): 1694–1707, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/6/1694/768621.

Hughes:2009:CST

[HHC^{+}09] Sarah L. Hughes, N. Penny Holliday, Eugene Colbourne, Vladimir Ozhigin, Hédinn Valdimarsson, Svein Østerhus, and Karen Wiltshire. Comparison of in situ time-series of temperature with gridded sea surface temperature datasets in the North Atlantic. ICES Journal of Marine Science, 66(7):1467–1479, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/7/1467/656378.

Holm:2000:STD

[HHH00] M. Holm, J. Chr. Holst, and L. P. Hansen. Spatial and temporal distribution of post-smolts of Atlantic salmon (Salmo salar L.) in the Norwegian Sea and adjacent areas. ICES Journal of Marine Science, 57(4):955–964, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/4/955/647395.

Haugland:2006:FAS

[HHHH06] Monika Haugland, Jens Christian Holst, Marianne Holm, and Lars Petter Hansen. Feeding of Atlantic salmon (Salmo salar L.) post-smolts in the Northeast Atlantic. ICES Journal of Marine Science, 63(8):1488–1500, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/8/1488/713660.

Hiddink:2006:PEA

[HHJK06] J. G. Hiddink, T. Hutton, S. Jennings, and M. J. Kaiser. Predicting the effects of area closures and fishing effort restrictions on the production, biomass, and species richness of benthic invertebrate communities. ICES Journal of Marine Science, 63 (5):822–830, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/5/822/662058.

Hofmann:2004:LMS

[HHKL04] Eileen E. Hofmann, A. G. Edward Haskell, John M. Klinck, and Cathy M. Lascara. Lagrangian modelling studies of Antarctic krill (Euphausia superba) swarm formation. ICES Journal of Marine Science, 61(4):617–631, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/4/617/606038.

Hinckley:2001:ISL

[HHMM01] Sarah Hinckley, Albert J. Hermann, Kathy L. Mier, and Bernard A. Megrey. Importance of spawning location and timing to successful transport to nursery areas: a simulation study of Gulf of Alaska walleye pollock. ICES Journal of Marine Science, 58(5):1042–1052, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/5/1042/730819.

Hermann:2001:ATC

[HHMN01] Albert J. Hermann, Sarah Hinckley, Bernard A. Megrey, and Jeffrey M. Napp. Applied and theoretical considerations for constructing spatially explicit individual-based models of marine larval fish that include multiple trophic levels. ICES Journal of Marine Science, 58(5):1030–1041, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/5/1030/730816.

Hjellvik:2008:CVA

[HHO08] Vidar Hjellvik, Nils Olav Handegard, and Egil Ona. Correcting for vessel avoidance in acoustic-abundance estimates for herring. ICES Journal of Marine Science, 65(6):1036–1045, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/6/1036/604262.

Hauton:2003:ESE

[HHSM03] C. Hauton, J. M. Hall-Spencer, and P. G. Moore. An experimental study of the ecological impacts of hydraulic bivalve dredging on maerl. ICES Journal of Marine Science, 60(2):381–392, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/2/381/627412.

Henderson:2008:IBP

[HHT08] M. J. Henderson, J. K. Horne, and R. H. Towler. The influence of beam position and swimming direction on fish target strength. ICES Journal of Marine Science, 65(2):226–237, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/2/226/734499.

Hollowed:2000:IPM

[HIL00] Anne Babcock Hollowed, James N. Ianelli, and Patricia A. Livingston. Including predation mortality in stock assessments: a case study for Gulf of Alaska walleye pollock. ICES Journal of Marine Science, 57(2):279–293, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/2/279/620429.

Hislop:2001:SEL

[His01] J. Hislop. A synthesis of the early life history of the anglerfish, Lophius piscatorius (Linnaeus, 1758) in northern British waters. ICES Journal of Marine Science, 58(1):70–86, ???? 2001. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/58/1/70/
603555.

Hansen:2003:OMW

[HJ03] Lars P. Hansen and Jan Arge Jacobsen. Origin and migration of wild and escaped farmed Atlantic salmon, Salmo salar L., in oceanic areas north of the Faroe Islands. ICES Journal of Marine Science, 60(1):110–119, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/1/110/611425.

Hannesson:2008:UTD

[HJB^{+}08] Sigurdur Hannesson, Audbjorg Jakobsdottir, James Begley, Lorna Taylor, and Gunnar Stefansson. On the use of tagging data in statistical multispecies multi-area models of marine populations. ICES Journal of Marine Science, 65(9):1762–1772, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/9/1762/629420.

Huse:2004:SST

[HJBG04] Geir Huse, Geir Odd Johansen, Bjarte Bogstad, and Harald Gjøsæter. Studying spatial and trophic interactions between capelin and cod using individual-based modelling. ICES Journal of Marine Science, 61(7):1201–1213, ???? 2004. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1201/881988.

Huse:2000:DVA

[HK00] Ingvar Huse and Rolf Korneliussen. Diel variation in acoustic density measurements of overwintering herring (Clupea harengus L.). ICES Journal of Marine Science, 57(4):903–910, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/903/647368.

Hvingel:2006:FMS

[HK06] Carsten Hvingel and Michael C. S. Kingsley. A framework to model shrimp (Pandalus borealis) stock dynamics and to quantify the risk associated with alternative management options, using Bayesian methods. ICES Journal of Marine Science, 63 (1):68–82, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/1/68/625497.

Hinrichsen:2009:IEB

[HKBK09] Hans-Harald Hinrichsen, Gerd Kraus, Uwe Böttcher, and Fritz Köster. Identifying eastern Baltic cod nursery grounds using hydrodynamic modelling: knowledge for the design of marine protected areas. ICES Journal of Marine Science, 66(1):101–108, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/1/101/635156.

Hassel:2004:ISS

[HKD^{+}04] Arne Hassel, Tor Knutsen, John Dalen, Kristian Skaar, Svein Løkkeborg, Ole Arve Misund, Øivind Østensen, Merete Fonn, and Eli Kyrkjebø Haugland. Influence of seismic shooting on the lesser sandeel (Ammodytes marinus). ICES Journal of Marine Science, 61(7):1165–1173, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/7/1165/881604.

Hayakawa:2001:SFM

[HKI01] Y. Hayakawa, M. Kobayashi, and M. Izawa. Sedimentation flux from mariculture of oyster (Crassostrea gigas) in Ofunato estuary, Japan. ICES Journal of Marine Science, 58(2):435–444, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/2/435/617846.

Hoolihan:2007:DSR

[HL07] John P. Hoolihan and Jiangang Luo. Determining summer residence status and vertical habitat use of sailfish (Istiophorus platypterus) in the Arabian Gulf. ICES Journal of Marine Science, 64(9):1791–1799, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/9/1791/783510.

Hoving:2009:FRB

[HL09] Hendrik Jan T. Hoving and Marek R. Lipiński. Female reproductive biology, and age of deep-sea squid Histioteuthis miranda from southern Africa. ICES Journal of Marine Science, 66(9): 1868–1872, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/9/1868/724104.

Halsband-Lenk:2004:LHS

[HLCG04] Claudia Halsband-Lenk, François Carlotti, and Wulf Greve. Life-history strategies of calanoid congeners under two different climate regimes: a comparison. ICES Journal of Marine Science, 61(4):709–720, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/61/4/709/608119.

He:2008:VME

[HLL^{+}08] Wenping He, Zhongjie Li, Jiashou Liu, Yuxuan Li, Brian R. Murphy, and Songguang Xie. Validation of a method of estimating age, modelling growth, and describing the age composition of Coilia mystus from the Yangtze Estuary, China. ICES Journal of Marine Science, 65(9):1655–1661, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/9/1655/630049.

Huse:2000:RST

[HLS00] Irene Huse, Svein Løkkeborg, and Aud Vold Soldal. Relative selectivity in trawl, longline and gillnet fisheries for cod and haddock. ICES Journal of Marine Science, 57(4):1271–1282, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/1271/647322.

Holmer:2001:ISE

[HLSW01] M. Holmer, P. Lassus, J. E. Stewart, and D. J. Wildish. ICES Symposium on Environmental Effects of Mariculture. ICES Journal of Marine Science, 58(2):363–368, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/2/363/617821.

Haugland:2004:ECS

[HM04a] Eli Kyrkjebø Haugland and Ole Arve Misund. Evidence for a clustered spatial distribution of clupeid fish schools in the Norwegian Sea and off the coast of southwest Africa. ICES Journal of Marine Science, 61(7):1088–1092, ???? 2004. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1088/878450.

Holley:2004:FSD

[HM04b] Jean-François Holley and Paul Marchal. Fishing strategy development under changing conditions: examples from the French offshore fleet fishing in the North Atlantic. ICES Journal of Marine Science, 61(8):1410–1431, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/8/1410/632553.

Hunt:2005:CBT

[HM05] George L. Hunt, Jr. and Bernard A. Megrey. Comparison of the biophysical and trophic characteristics of the Bering and Barents Seas. ICES Journal of Marine Science, 62(7):1245–1255, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/7/1245/655224.

Haring:2008:MFM

[HM08] Phil Haring and J-J. Maguire. The monkfish fishery and its management in the northeastern USA. ICES Journal of Marine Science, 65(7):1370–1379, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/7/1370/646844.

Hjellvik:2003:AEE

[HMAN03] Vidar Hjellvik, Kathrine Michalsen, Asgeir Aglen, and Odd Nakken. An attempt at estimating the effective fishing height of the bottom trawl using acoustic survey recordings. ICES Journal of Marine Science, 60(5):967–979, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/5/967/770555.

Hedger:2008:OIF

[HMD^{+}08] Richard D. Hedger, François Martin, Julian J. Dodson, Daniel Hatin, François Caron, and Fred G. Whoriskey. The optimized interpolation of fish positions and speeds in an array of fixed acoustic receivers. ICES Journal of Marine Science, 65(7): 1248–1259, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/7/1248/646012.

Hay:2009:SDP

[HMDS09] Douglas E. Hay, P. Bruce McCarter, Kristen S. Daniel, and Jacob F. Schweigert. Spatial diversity of Pacific herring (Clupea pallasi) spawning areas. ICES Journal of Marine Science, 66 (8):1662–1666, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/8/1662/674966.

Hirose:2009:ACT

[HMHI09] Miyuki Hirose, Tohru Mukai, Doojin Hwang, and Kohji Iida. The acoustic characteristics of three jellyfish species: Nemopilema nomurai, Cyanea nozakii, and Aurelia aurita. ICES Journal of Marine Science, 66(6):1233–1237, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/1233/696569.

Haapasaari:2007:MMF

[HMK^{+}07] P. Haapasaari, C. G. J. Michielsens, T. P. Karjalainen, K. Reinikainen, and S. Kuikka. Management measures and fishers’ commitment to sustainable exploitation: a case study of Atlantic salmon fisheries in the Baltic Sea. ICES Journal of Marine Science, 64(4):825–833, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/825/639245.

Hardman-Mountford:2008:OMS

[HMMB^{+}08] Nicholas J. Hardman-Mountford, Gerald Moore, Dorothee C. E. Bakker, Andrew J. Watson, Ute Schuster, Rosa Barciela, Adrian Hines, Gwenaelle Moncoiffé, Juan Brown, Stephen Dye, Jerry Blackford, Paul J. Somerfield, Jason Holt, David J. Hydes, and James Aiken. An operational monitoring system to provide indicators of CO_{2}-related variables in the ocean. ICES Journal of Marine Science, 65(8):1498–1503, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/8/1498/712704.

Hutton:2004:MFL

[HMPC04] Trevor Hutton, Simon Mardle, Sean Pascoe, and Robin A. Clark. Modelling fishing location choice within mixed fisheries: English North Sea beam trawlers in 2000 and 2001. ICES Journal of Marine Science, 61(8):1443–1452, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/8/1443/632932.

Hulson:2008:DCF

[HMQ^{+}08] Peter-John F. Hulson, Sara E. Miller, Terrance J. Quinn, Gary D. Marty, Steven D. Moffitt, and Frederick Funk. Data conflicts in fishery models: incorporating hydroacoustic data into the Prince William Sound Pacific herring assessment model. ICES Journal of Marine Science, 65(1):25–43, January 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/1/25/611634.

Hauge:2007:LTE

[HNK07] Kjellrun Hiis Hauge, Kåre Nolde Nielsen, and Knut Korsbrekke. Limits to transparency — exploring conceptual and operational aspects of the ICES framework for providing precautionary fisheries management advice. ICES Journal of Marine Science, 64 (4):738–743, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/4/738/642919.

Humborstad:2004:RBC

[HNLR04] Odd-Børre Humborstad, Leif Nøttestad, Svein Løkkeborg, and Hans Tore Rapp. RoxAnn bottom classification system, sidescan sonar and video-sledge: spatial resolution and their use in assessing trawling impacts. ICES Journal of Marine Science, 61(1):53–63, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/1/53/702027.

Hammond:2001:ABA

[HO01] T. R. Hammond and C. M. O’Brien. An application of the Bayesian approach to stock assessment model uncertainty. ICES Journal of Marine Science, 58(3):648–656, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/3/648/810111.

Horwood:2006:NSC

[HOD06] Joe Horwood, Carl O’Brien, and Chris Darby. North Sea cod recovery? ICES Journal of Marine Science, 63(6):961–968, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/6/961/617928.

Hoie:2004:TDF

[HOF04] Hans Høie, Erling Otterlei, and Arild Folkvord. Temperaturedependent fractionation of stable oxygen isotopes in otoliths of juvenile cod (Gadus morhua L.). ICES Journal of Marine Science, 61(2):243–251, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/61/2/243/621210.

Huveneers:2007:QDA

[HOGH07] Charlie Huveneers, Nicholas M. Otway, Susan E. Gibbs, and Robert G. Harcourt. Quantitative diet assessment of wobbegong sharks (genus Orectolobus) in New South Wales, Australia. ICES Journal of Marine Science, 64(6):1272–1281, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/6/1272/616342.

Hauge:2005:FMQ

[HOHS05] Kjellrun Hiis Hauge, Erik Olsen, Hilde Elise Heldal, and Hein Rune Skjoldal. A framework for making qualities of indicators transparent. ICES Journal of Marine Science, 62 (3):552–557, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/3/552/665630.

Hollingworth:2000:EEF

[Hol00a] Chuck Hollingworth. Ecosystem effects of fishing preface. ICES Journal of Marine Science, 57(3):465, June 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/3/465/635909.

Hollowed:2000:MMI

[Hol00b] A. Hollowed. Are multispecies models an improvement on single-species models for measuring fishing impacts on marine ecosystems? ICES Journal of Marine Science, 57(3):707–719, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/3/707/635962.

Hollingworth:2002:P

[Hol02] C. E. Hollingworth. Preface. ICES Journal of Marine Science, 59(5):861–862, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/5/861/675188.

Holland:2003:ISM

[Hol03] Daniel S. Holland. Integrating spatial management measures into traditional fishery management systems: the case of the Georges Bank multispecies groundfish fishery. ICES Journal of Marine Science, 60(5):915–929, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/5/915/768979.

Harbitz:2009:UAA

[HOP09] Alf Harbitz, Egil Ona, and Michael Pennington. The use of an adaptive acoustic-survey design to estimate the abundance of highly skewed fish populations. ICES Journal of Marine Science, 66(6):1349–1354, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1349/694019.

Horne:2003:IOP

[Hor03] John K. Horne. The influence of ontogeny, physiology, and behaviour on the target strength of walleye pollock (Theragra chalcogramma). ICES Journal of Marine Science, 60(5): 1063–1074, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/5/1063/765444.

Horbowy:2008:SPC

[Hor08] Jan Horbowy. Sensitivity of predicted cohort size and catches to errors in estimates of fishing mortality in the terminal year. ICES Journal of Marine Science, 65(7):1227–1234, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/7/1227/646314.

Hoydal:2007:VIB

[Hoy07] Kjartan Hoydal. Viewpoint: the interface between scientific advice and fisheries management. ICES Journal of Marine Science, 64(4):846–850, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/4/846/639923.

Horbowy:2001:MIB

[HP01] Jan Horbowy and Magdalena Podolska. Modelling infection of Baltic herring (Clupea harengus membras) by larval Anisakis simplex. ICES Journal of Marine Science, 58(1):321–330, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/1/321/603538.

Harbitz:2004:CSS

[HP04] Alf Harbitz and Michael Pennington. Comparison of shortest sailing distance through random and regular sampling points. ICES Journal of Marine Science, 61(1):140–147, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/1/140/700632.

Hammer:2005:RMP

[HP05] Astrid C. Hammer and Jonathan W. Pitchford. The role of mixotrophy in plankton bloom dynamics, and the consequences for productivity. ICES Journal of Marine Science, 62 (5):833–840, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/5/833/855604.

Hallfredsson:2007:MGM

[HP07] Elvar H. Hallfredsson and John G. Pope. Modelling the growth, mortality, and predation interactions of cod juveniles and capelin larvae in the Barents Sea using a novel protomoment population dynamics model. ICES Journal of Marine Science, 64(7):1313–1323, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/7/1313/729346.

Handegard:2009:ETB

[HPB09] Nils Olav Handegard, Geir Pedersen, and Ole Brix. Estimating tail-beat frequency using split-beam echosounders. ICES Journal of Marine Science, 66(6):1252–1258, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/1252/688797.

Hofmann:2004:MSI

[HPBK04] Eileen E. Hofmann, Eric N. Powell, Eleanor A. Bochenek, and John M. Klinck. A modelling study of the influence of environment and food supply on survival of Crassostrea gigas larvae. ICES Journal of Marine Science, 61(4):596–616, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/4/596/605613.

Hoffmann:2009:MPA

[HPR09] Erik Hoffmann and Angel Pérez-Ruzafa. Marine protected areas as a tool for fishery management and ecosystem conservation: an introduction. ICES Journal of Marine Science, 66(1):1–5, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/1/1/633884.

Huggett:2000:RBE

[HR00] J. A. Huggett and A. J. Richardson. A review of the biology and ecology of Calanus agulhensis off South Africa. ICES Journal of Marine Science, 57(6):1834–1849, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/6/1834/768685.

Holliday:2001:TCB

[HR01] N. Penny Holliday and Philip C. Reid. Is there a connection between high transport of water through the Rockall Trough and ecological changes in the North Sea? ICES Journal of Marine Science, 58(1):270–274, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/1/270/603528.

Hermes:2009:SSC

[HR09] Juliet C. Hermes and Chris J. C. Reason. The sensitivity of the Seychelles–Chagos thermocline ridge to large-scale wind anomalies. ICES Journal of Marine Science, 66(7):1455–1466, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/7/1455/658085.

Holbrook:2002:SEO

[HRB02] Sally J. Holbrook, Daniel C. Reed, and J. Scott Bull. Survival experiments with outplanted seedlings of surfgrass (Phyllospadix torreyi) to enhance establishment on artificial structures. ICES Journal of Marine Science, 59(S1):S350–S355, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL https://academic.oup.com/icesjms/
article/59/suppl/S350/617993.

Hygum:2000:RCC

[HRHC00] B. H. Hygum, C. Rey, B. W. Hansen, and F. Carlotti. Rearing cohorts of Calanus finmarchicus (Gunnerus) in mesocosms. ICES Journal of Marine Science, 57(6):1740–1751, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/6/1740/768638.

Heffernan:2004:UDS

[HRM04] O. Heffernan, D. Righton, and K. Michalsen. Use of data storage tags to quantify vertical movements of cod: effects on acoustic measures. ICES Journal of Marine Science, 61(7):1062–1070, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/7/1062/878004.

Hammond:2001:GPE

[HS01] T. R. Hammond and G. L. Swartzman. A general procedure for estimating the composition of fish school clusters using standard acoustic survey data. ICES Journal of Marine Science, 58 (6):1115–1132, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1115/641523.

Harris:2006:SGS

[HS06] Bradley P. Harris and Kevin D. E. Stokesbury. Shell growth of sea scallops (Placopecten magellanicus) in the southern and northern Great South Channel, USA. ICES Journal of Marine Science, 63(5):811–821, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/63/5/811/661668.

Hammill:2007:APA

[HS07] M. O. Hammill and G. B. Stenson. Application of the precautionary approach and conservation reference points to management of Atlantic seals. ICES Journal of Marine Science, 64(4):702–706, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/4/702/641570.

Hunter:2009:CEH

[HS09] W. R. Hunter and M. D. J. Sayer. The comparative effects of habitat complexity on faunal assemblages of northern temperate artificial and natural reefs. ICES Journal of Marine Science, 66(4):691–698, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/4/691/604376.

Hinrichsen:2001:TLD

[HSA^{+}01] H.-H. Hinrichsen, M. St. John, E. Aro, P. Grønkjær, and R. Voss. Testing the larval drift hypothesis in the Baltic Sea: retention versus dispersion caused by wind-driven circulation. ICES Journal of Marine Science, 58(5):973–984, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/5/973/730873.

Hutin:2005:ADS

[HSA05] Estelle Hutin, Yvan Simard, and Philippe Archambault. Acoustic detection of a scallop bed from a single-beam echosounder in the St. Lawrence. ICES Journal of Marine Science, 62 (5):966–983, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/5/966/858289.

Horne:2009:SUP

[HSA^{+}09] John K. Horne, Kouichi Sawada, Koki Abe, Richard B. Kreisberg, David H. Barbee, and Kazuhiro Sadayasu. Swimbladders under pressure: anatomical and acoustic responses by walleye pollock. ICES Journal of Marine Science, 66(6):1162–1168, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1162/694779.

Haug:2006:EHS

[HSCN06] Tore Haug, Garry B. Stenson, Peter J. Corkeron, and Kjell T. Nilssen. Estimation of harp seal (Pagophilus groenlandicus) pup production in the North Atlantic completed: results from surveys in the Greenland Sea in 2002. ICES Journal of Marine Science, 63(1):95–104, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/63/1/95/625803.

Hall:2006:DLC

[HSdLP06] Norman G. Hall, Kim D. Smith, Simon de Lestang, and Ian C. Potter. Does the largest chela of the males of three crab species undergo an allometric change that can be used to determine morphometric maturity? ICES Journal of Marine Science, 63 (1):140–150, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/1/140/623692.

Hall-Spencer:2000:SDP

[HSM00] J. M. Hall-Spencer and P. G. Moore. Scallop dredging has profound, long-term impacts on maerl habitats. ICES Journal of Marine Science, 57(5):1407–1415, October 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/5/1407/660943.

Hamilton:2009:PSH

[HSM09] K. M. Hamilton, P. W. Shaw, and D. Morritt. Prevalence and seasonality of Hematodinium (Alveolata: Syndinea) in a Scottish crustacean community. ICES Journal of Marine Science, 66(9):1837–1845, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/9/1837/722060.

Hinrichsen:2005:SPB

[HSPM05] H-H. Hinrichsen, J. O. Schmidt, C. Petereit, and C. Möllmann. Survival probability of Baltic larval cod in relation to spatial overlap patterns with their prey obtained from drift model studies. ICES Journal of Marine Science, 62(5):878–885, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/5/878/856286.

Hammond:2001:BEF

[HSR01] T. R. Hammond, G. L. Swartzman, and T. S. Richardson. Bayesian estimation of fish school cluster composition applied to a Bering Sea acoustic survey. ICES Journal of Marine Science, 58(6):1133–1149, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1133/641526.

Hastie:2005:EMP

[HSS^{+}05] Gordon D. Hastie, René J. Swift, George Slesser, Paul M. Thompson, and William R. Turrell. Environmental models for predicting oceanic dolphin habitat in the Northeast Atlantic.

ICES Journal of Marine Science, 62(4):760–770, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/4/760/675762.

Husebo:2007:GJN

[HSS07] Åse Husebø, Aril Slotte, and Erling Kåre Stenevik. Growth of juvenile Norwegian spring-spawning herring in relation to latitudinal and interannual differences in temperature and fish density in their coastal and fjord nursery areas. ICES Journal of Marine Science, 64(6):1161–1172, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/6/1161/613876.

Huse:2008:INI

[HSS08] Geir Huse, Are Salthaug, and Morten D. Skogen. Indications of a negative impact of herring on recruitment of Norway pout. ICES Journal of Marine Science, 65(6):906–911, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/6/906/603314.

Husebo:2009:EHT

[HSS^{+}09] Åse Husebø, Erling Kåre Stenevik, Aril Slotte, Petter Fossum, Are Salthaug, Frode Vikebø, Sondre Aanes, and Arild Folkvord. Effects of hatching time on year-class strength in Norwegian spring-spawning herring (Clupea harengus). ICES Journal of Marine Science, 66(8):1710–1717, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/8/1710/675413.

Hammond:2005:CCD

[HT05] T. R. Hammond and V. M. Trenkel. Censored catch data in fisheries stock assessment. ICES Journal of Marine Science, 62 (6):1118–1130, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/6/1118/617407.

Hill:2009:RFP

[HTA09] Simeon L. Hill, Philip N. Trathan, and David J. Agnew. The risk to fishery performance associated with spatially resolved management of Antarctic krill (Euphausia superba) harvesting. ICES Journal of Marine Science, 66(10):2148–2154, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/10/2148/678166.

Hindar:2004:CGV

[HTSB04] Kjetil Hindar, Jarle Tufto, Leif Magnus Sættem, and Torveig Balstad. Conservation of genetic variation in harvested salmon populations. ICES Journal of Marine Science, 61(8):1389–1397, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/8/1389/632252.

Harris:2004:DNL

[HU04] R. R. Harris and M. Ulmestrand. Discarding Norway lobster (Nephrops norvegicus L.) through low salinity layers — mortality and damage seen in simulation experiments. ICES Journal of Marine Science, 61(1):127–139, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/1/127/700187.

Huse:2004:PMS

[Hus04] Ingvar Huse. Physiological mechanisms of sensory organs and swimming performance. ICES Journal of Marine Science, 61 (7):1235, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/7/1235/882367.

Hansen:2006:IBA

[HW06] Lars P. Hansen and Malcolm L. Windsor. Interactions between aquaculture and wild Stocks of Atlantic salmon and other diadromous fish species: Science and management, challenges and solutions: an introduction by the conveners. ICES Journal of Marine Science, 63(7):1159–1161, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/7/1159/752717.

Handegard:2008:ATF

[HW08] Nils Olav Handegard and Kresimir Williams. Automated tracking of fish in trawls using the DIDSON (Dual frequency IDentification SONar). ICES Journal of Marine Science, 65(4):636–644, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/4/636/636791.

Holmes:2008:ESD

[HWF08] Steven J. Holmes, Peter J. Wright, and Robert J. Fryer. Evidence from survey data for regional variability in cod dynamics in the North Sea and West of Scotland. ICES Journal of Marine Science, 65(2):206–215, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/2/206/734744.

Isla:2004:MSF

[IA04] José Alejandro Isla and Ricardo Anadón. Mesozooplankton size-fractionated metabolism and feeding off NW Spain during autumn: effects of a poleward current. ICES Journal of Marine Science, 61(4):526–534, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/4/526/603653.

Iles:2000:CHS

[IB00] Terence C. Iles and R. J. H. Beverton. The concentration hypothesis: the statistical evidence. ICES Journal of Marine Science, 57(2):216–227, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/2/216/620410.

Ibaibarriaga:2008:TSB

[IFUR08] Leire Ibaibarriaga, Carmen Fernández, Andrés Uriarte, and Beatriz A. Roel. A two-stage biomass dynamic model for Bay of Biscay anchovy: a Bayesian approach. ICES Journal of Marine Science, 65(2):191–205, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/2/191/735664.

Jakupsstovu:2007:ERD

[íJCMR07] S. H. í Jákupsstovu, L. R. Cruz, J-J. Maguire, and J. Reinert. Effort regulation of the demersal fisheries at the Faroe Islands: a 10-year appraisal. ICES Journal of Marine Science, 64(4):730– 737, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/4/730/642865.

Jakupsstovu:2002:CFW

[íJR02] Stein Hjalti í Jákupsstovu and Jákup Reinert. Capelin in Faroese waters — a messenger of harsh times? ICES Journal of Marine Science, 59(5):884–889, ???? 2002. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/59/5/884/675199.

Irigoien:2000:EFD

[Iri00] X. Irigoien. The effect of food on the determination of sex ratio in Calanus spp.: evidence from experimental studies and field data. ICES Journal of Marine Science, 57(6):1752–1763, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/6/1752/768643.

Ingolfsson:2007:EMC

[ISHB07] Ólafur Arnar Ingólfsson, Aud Vold Soldal, Irene Huse, and Mike Breen. Escape mortality of cod, saithe, and haddock in a Barents Sea trawl fishery. ICES Journal of Marine Science, 64(9): 1836–1844, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/9/1836/784141.

Itoi:2007:IGP

[ISS^{+}07] Shiro Itoi, Takashi Saito, Mai Shimojo, Sayaka Washio, and Haruo Sugita. Identification of Girella punctata and G. leonina by PCR-RFLP analysis. ICES Journal of Marine Science, 64 (2):328–331, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/2/328/2182257.

Incze:2000:ASR

[IWP00] Lewis S. Incze, Richard A. Wahle, and Alvaro T. Palma. Advection and settlement rates in a benthic invertebrate: recruitment to first benthic stage in Homarus americanus. ICES Journal of Marine Science, 57(2):430–437, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/2/430/620485.

Jackson:2000:CLF

[JAC00] George D. Jackson, Ross A. Alford, and J. Howard Choat. Can length frequency analysis be used to determine squid growth? — An assessment of ELEFAN. ICES Journal of Marine Science, 57 (4):948–954, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/4/948/647391.

Jaffe:2006:UMA

[Jaf06] Jules S. Jaffe. Using multi-angle scattered sound to size fish swimbladders. ICES Journal of Marine Science, 63(8): 1397–1404, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/8/1397/711544.

Jaffe:2008:UMA

[Jaf08] Jules S. Jaffe. Using multiple-angle scattered sound to size fish swimbladders. ICES Journal of Marine Science, 65(6): 1092, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/6/1092/605571.

Jager:2002:ACD

[Jag02] Zwanette Jager. Across-channel distribution of flounder larvae (Platichthys flesus L.) in the eems–dollard estuary and its effects on larval transport estimates. ICES Journal of Marine Science, 59(6):1187–1198, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/6/1187/608175.

Josse:2000:SAT

[JB00] Erwan Josse and Arnaud Bertrand. In situ acoustic target strength measurements of tuna associated with a fish aggregating device. ICES Journal of Marine Science, 57(4):911–918, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/911/647373.

Jonsdottir:2006:OST

[JCM06] Ingibjörg G. Jónsdóttir, Steven E. Campana, and Gudrun Marteinsdottir. Otolith shape and temporal stability of spawning groups of Icelandic cod (Gadus morhua L.). ICES Journal of Marine Science, 63(8):1501–1512, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/8/1501/713939.

Jennings:2005:RPR

[JD05] Simon Jennings and Nicholas K. Dulvy. Reference points and reference directions for size-based indicators of community structure. ICES Journal of Marine Science, 62(3):397–404, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/3/397/661240.

Jadot:2006:APH

[JDA^{+}06] C. Jadot, A. Donnay, M. L. Acolas, Y. Cornet, and M. L. Bégout Anras. Activity patterns, home-range size, and habitat utilization of Sarpa salpa (Teleostei: Sparidae) in the Mediterranean Sea. ICES Journal of Marine Science, 63(1):128–139, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/1/128/623536.

Jonsdottir:2001:GDA

[JDN01] Ó. D. B. Jónsdóttir, A. K. Daníelsdóttir, and G. Nævdal. Genetic differentiation among Atlantic cod (Gadus morhua L.) in Icelandic waters: temporal stability. ICES Journal of Marine Science, 58(1):114–122, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/58/1/114/603499.

Jellyman:2007:SNZ

[Jel07] Don J. Jellyman. Status of New Zealand fresh-water eel stocks and management initiatives. ICES Journal of Marine Science, 64(7):1379–1386, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/64/7/1379/726306.

Jensen:2002:ARE

[Jen02a] Antony Jensen. Artificial reefs of Europe: perspective and future. ICES Journal of Marine Science, 59(S1):S3–S13, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL https://academic.oup.com/icesjms/
article/59/suppl/S3/617973.

Jensen:2002:SIC

[Jen02b] Antony C. Jensen. The Seventh International Conference on Artificial Reefs and Related Aquatic Habitats. Introduction. ICES Journal of Marine Science, 59(S1):S1–S2, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S1/617893.

Jennings:2009:RMP

[Jen09] Simon Jennings. The role of marine protected areas in environmental management. ICES Journal of Marine Science, 66 (1):16–21, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/1/16/632372.

Jech:2005:CTK

[JFCH05] J. Michael Jech, Kenneth G. Foote, Dezhang Chu, and Lawrence C. Hufnagle, Jr. Comparing two 38-kHz scientific echosounders. ICES Journal of Marine Science, 62(6): 1168–1179, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/6/1168/618060.

Johnsen:2007:DVA

[JG07] E. Johnsen and O. R. Godø. Diel variations in acoustic recordings of blue whiting (Micromesistius poutassou). ICES Journal of Marine Science, 64(6):1202–1209, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/6/1202/616153.

Joaquim:2008:RVS

[JGM^{+}08] Sandra Joaquim, Miguel B. Gaspar, Domitília Matias, Radhouan Ben-Hamadou, and William S. Arnold. Rebuilding viable spawner patches of the overfished Spisula solida (Mollusca: Bivalvia): a preliminary contribution to fishery sustainability. ICES Journal of Marine Science, 65(1):60–64, January 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/1/60/612148.

Joensen:2004:SSS

[JGN04] Hóraldur Joensen and Otto Grahl-Nielsen. Stock structure of Sebastes mentella in the North Atlantic revealed by chemometry of the fatty acid profile in heart tissue. ICES Journal of Marine Science, 61(1):113–126, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/1/113/699392.

Johansen:2009:UAT

[JGST09] Geir O. Johansen, Olav R. Godø, Morten D. Skogen, and Terje Torkelsen. Using acoustic technology to improve the modelling of the transportation and distribution of juvenile gadoids in the Barents Sea. ICES Journal of Marine Science, 66(6):1048–1054, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1048/692831.

Jacobsen:2001:FHW

[JH01a] Jan Arge Jacobsen and Lars Petter Hansen. Feeding habits of wild and escaped farmed Atlantic salmon, Salmo salar L., in the Northeast Atlantic. ICES Journal of Marine Science, 58 (4):916–933, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/4/916/630230.

Jech:2001:EST

[JH01b] J. Michael Jech and John K. Horne. Effects of in situ target spatial distributions on acoustic density estimates. ICES Journal of Marine Science, 58(1):123–136, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/1/123/603503.

Jiao:2009:HBA

[JHC09] Yan Jiao, Christopher Hayes, and Enric Cortés. Hierarchical Bayesian approach for population dynamics modelling of fish complexes without species-specific data. ICES Journal of Marine Science, 66(2):367–377, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/2/367/593660.

Jakob:2009:SDP

[JHKZ09] Eva Jakob, Reinhold Hanel, Sven Klimpel, and Karsten Zumholz. Salinity dependence of parasite infestation in the European eel Anguilla anguilla in northern Germany. ICES Journal of Marine Science, 66(2):358–366, March 2009. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/2/358/593513.

Jorgensen:2005:SSH

[JHL05] Hanne B. Hede Jørgensen, Michael Møller Hansen, and Volker Loeschcke. Spring-spawning herring (Clupea harengus L.) in the southwestern Baltic Sea: do they form genetically distinct spawning waves? ICES Journal of Marine Science, 62(6): 1065–1075, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/6/1065/616347.

Jouffre:2005:AIF

[JI05] Didier Jouffre and Cheikh A. Inejih. Assessing the impact of fisheries on demersal fish assemblages of the Mauritanian continental shelf, 1987–1999, using dominance curves. ICES Journal of Marine Science, 62(3):380–383, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/3/380/660597.

Jonsson:2006:CAS

[JJ06] Bror Jonsson and Nina Jonsson. Cultured Atlantic salmon in nature: a review of their ecology and interaction with wild fish. ICES Journal of Marine Science, 63(7):1162–1181, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/7/1162/752771.

Jorstad:2006:CGR

[JKSO06] K. E. Jørstad, Ø. Karlsen, T. Svåsand, and H. Otterå. Comparison of growth rate among different protein genotypes in Atlantic cod, Gadus morhua, under farmed conditions. ICES Journal of Marine Science, 63(2):235–245, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/2/235/638422.

Johnson:2008:CCF

[JLR^{+}08] Catherine L. Johnson, Andrew W. Leising, Jeffrey A. Runge, Erica J. H. Head, Pierre Pepin, Stéphane Plourde, and Edward G. Durbin. Characteristics of Calanus finmarchicus dormancy patterns in the Northwest Atlantic. ICES Journal of Marine Science, 65(3):339–350, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/3/339/782428.

Jorgensen:2002:RFV

[JLS02] Terje Jørgensen, Svein Løkkeborg, and Aud Vold Soldal. Residence of fish in the vicinity of a decommissioned oil platform in the North Sea. ICES Journal of Marine Science, 59(S1): S288–S293, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S288/617969.

Jonsdottir:2007:CDS

[JMC07] Ingibjörg G. Jónsdóttir, Gudrun Marteinsdottir, and Steven E. Campana. Contribution of different spawning components to the mixed stock fishery for cod in Icelandic waters. ICES Journal of Marine Science, 64(9):1749–1759, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/9/1749/786873.

Jurado-Molina:2005:TSS

[JMLG05] Jesús Jurado-Molina, Patricia A. Livingston, and Vincent F. Gallucci. Testing the stability of the suitability coefficients from an eastern Bering Sea multispecies virtual population analysis. ICES Journal of Marine Science, 62(5):915–924, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/5/915/856879.

Jackson:2008:DYC

[JMWJ08] George D. Jackson, Mark G. Meekan, Simon Wotherspoon, and Christine H. Jackson. Distributions of young cephalopods in the tropical waters of Western Australia over two consecutive summers. ICES Journal of Marine Science, 65(2):140–147, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/2/140/733818.

Johannessen:2009:TCN

[JNF^{+}09] Arne Johannessen, Leif Nøttestad, Anders Fernö, Lise Langård, and Georg Skaret. Two components of Northeast Atlantic herring within the same school during spawning: support for the existence of a metapopulation? ICES Journal of Marine Science, 66(8):1740–1748, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/66/8/1740/676584.

Jorgensen:2002:ATS

[JO02] Roar Jørgensen and Kjell Olsen. Acoustic target strength of capelin measured by single-target tracking in a controlled cage experiment. ICES Journal of Marine Science, 59(5):1081–1085, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/5/1081/675159.

Johansen:2002:TSV

[Joh02] Geir Odd Johansen. Temporal and spatial variation in predation on juvenile herring (Clupea harengus L.) by Northeast Arctic cod (Gadus morhua L.) in the Barents Sea in 1984–1997. ICES Journal of Marine Science, 59(2):270–292, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/2/270/619637.

Johnson:2008:EIT

[Joh08] David Johnson. Environmental indicators: their utility in meeting the OSPAR Convention’s regulatory needs. ICES Journal of Marine Science, 65(8):1387–1391, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/8/1387/714571.

Jorgensen:2003:ESS

[Jør03] Roar Jørgensen. The effects of swimbladder size, condition and gonads on the acoustic target strength of mature capelin. ICES Journal of Marine Science, 60(5):1056–1062, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/5/1056/765072.

Joyce:2002:ASC

[Joy02] W. Joyce. Analysis of stomach contents of the porbeagle shark (Lamna nasus Bonnaterre) in the northwest Atlantic. ICES Journal of Marine Science, 59(6):1263–1269, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/6/1263/608187.

Jackson:2003:DSS

[JP03] George D. Jackson and Gretta Pecl. The dynamics of the summer-spawning population of the loliginid squid Sepioteuthis australis in Tasmania, Australia — a conveyor belt of recruits. ICES Journal of Marine Science, 60(2):290–296, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/2/290/625794.

Johnsen:2009:SDF

[JPO09] Espen Johnsen, Ronald Pedersen, and Egil Ona. Size-dependent frequency response of sandeel schools. ICES Journal of Marine Science, 66(6):1100–1105, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1100/694183.

Janowicz:2001:MBI

[JR01] M. Janowicz and J. Ross. Monitoring for benthic impacts in the southwest New Brunswick salmon aquaculture industry. ICES Journal of Marine Science, 58(2):453–459, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/2/453/617849.

Jaworski:2006:FHD

[JR06] Andrzej Jaworski and Stefán Áki Ragnarsson. Feeding habits of demersal fish in Icelandic waters: a multivariate approach. ICES Journal of Marine Science, 63(9):1682–1694, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/9/1682/699283.

Jennings:2007:RGT

[JR07] Simon Jennings and Andrew S. Revill. The role of gear technologists in supporting an ecosystem approach to fisheries. ICES Journal of Marine Science, 64(8):1525–1534, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/8/1525/613187.

Johnson:2008:GRF

[JRCS08] A. K. Johnson, R. Anne Richards, Daniel W. Cullen, and Sandra J. Sutherland. Growth, reproduction, and feeding of large monkfish, Lophius americanus. ICES Journal of Marine Science, 65(7):1306–1315, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/65/7/1306/646950.

Junquera:2003:TSO

[JRM^{+}03] S. Junquera, E. Román, J. Morgan, M. Sainza, and G. Ramilo. Time scale of ovarian maturation in Greenland halibut (Reinhardtius hippoglossoides, Walbaum). ICES Journal of Marine Science, 60(4):767–773, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/60/4/767/692082.

Jiao:2006:VCY

[JRN06] Yan Jiao, Kevin Reid, and Tom Nudds. Variation in the catchability of yellow perch (Perca flavescens) in the fisheries of Lake Erie using a Bayesian error-in-variable approach. ICES Journal of Marine Science, 63(9):1695–1704, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/9/1695/699612.

Jounela:2006:IBG

[JSMK06] P. Jounela, P. Suuronen, R. B. Millar, and M-L. Koljonen. Interactions between grey seal (Halichoerus grypus), Atlantic salmon (Salmo salar), and harvest controls on the salmon fishery in the Gulf of Bothnia. ICES Journal of Marine Science, 63(5):936–945, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/5/936/664169.

Jaworski:2006:EAC

[JSR06] Andrzej Jaworski, Jon Solmundsson, and Stefan Aki Ragnarsson. The effect of area closures on the demersal fish community off the east coast of Iceland. ICES Journal of Marine Science, 63(5):897–911, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/5/897/663768.

Jonasson:2007:CFI

[JTE^{+}07] Jónas P. Jonasson, Gudrun Thorarinsdottir, Hrafnkell Eiriksson, Jon Solmundsson, and Gudrun Marteinsdottir. Collapse of the fishery for Iceland scallop (Chlamys islandica) in Breidafjordur, West Iceland. ICES Journal of Marine Science, 64(2):298–308, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/2/298/2182359.

Jonsson:2005:FAW

[JV05] Steingrímur Jónsson and Hédinn Valdimarsson. The flow of Atlantic water to the North Icelandic shelf and its relation to the drift of cod larvae. ICES Journal of Marine Science, 62(7): 1350–1359, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/7/1350/657580.

Johnson:2007:BOC

[JvD07] Teresa R. Johnson and Wim L. T. van Densen. Benefits and organization of cooperative research for fisheries management. ICES Journal of Marine Science, 64(4):834–840, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/4/834/640672.

Jansen:2007:JGF

[JWBP07] Henrice M. Jansen, Hendrik V. Winter, Maarten C. M. Bruijs, and Harry J. G. Polman. Just go with the flow? Route selection and mortality during downstream migration of silver eels in relation to river discharge. ICES Journal of Marine Science, 64(7): 1437–1443, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/7/1437/729131.

Jensen:2003:VDP

[JWM03] Henrik Jensen, Peter J. Wright, and Peter Munk. Vertical distribution of pre-settled sandeel (Ammodytes marinus) in the North Sea in relation to size and environmental variables. ICES Journal of Marine Science, 60(6):1342–1351, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/6/1342/654032.

Jirapunpipat:2009:BSB

[JYW09] Kanchana Jirapunpipat, Masashi Yokota, and Seiichi Watanabe. The benefits of species-based management of sympatric mud crabs migrating to a common fishing ground. ICES Journal of Marine Science, 66(3):470–477, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/3/470/812631.

Kidwai:2001:ADI

[KA01] Samina Kidwai and Shahid Amjad. Abundance and distribution of ichthyolarvae from upper pelagic waters of the northwestern Arabian Sea during different monsoon periods, 1992–1994. ICES Journal of Marine Science, 58(3):719–724, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/3/719/810139.

Kaartvedt:2000:LHC

[Kaa00] S. Kaartvedt. Life history of Calanus finmarchicus in the Norwegian Sea in relation to planktivorous fish. ICES Journal of Marine Science, 57(6):1819–1824, December 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/6/1819/768678.

Kalikhman:2001:PDF

[Kal01] I. Kalikhman. Patchy distribution fields: sampling distance unit and reconstruction adequacy. ICES Journal of Marine Science, 58(6):1184–1194, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1184/641545.

Kane:2007:ZAT

[Kan07] Joseph Kane. Zooplankton abundance trends on Georges Bank, 1977–2004. ICES Journal of Marine Science, 64(5):909–919, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/5/909/641810.

Karayev:2006:MCS

[Kar06] Robert A. Karayev. Modelling Caspian sturgeon population dynamics: a new paradigm and new technology. ICES Journal of Marine Science, 63(6):980–994, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/6/980/618403.

Kasatkina:2009:IUT

[Kas09] Svetlana M. Kasatkina. The influence of uncertainty in target strength on abundance indices based on acoustic surveys: examples of the Baltic Sea herring and sprat. ICES Journal of Marine Science, 66(6):1404–1409, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1404/693832.

Karlson:2007:ICB

[KASA07] Agnes M. L. Karlson, Gustaf Almqvist, Krzysztof E. Skóra, and Magnus Appelberg. Indications of competition between nonindigenous round goby and native flounder in the Baltic Sea. ICES Journal of Marine Science, 64(3):479–486, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/3/479/814725.

Katsukawa:2005:ECA

[Kat05] T. Katsukawa. Evaluation of current and alternative fisheries management scenarios based on spawning-per-recruit (SPR), revenue-per-recruit (RPR), and yield-per-recruit (YPR) diagrams. ICES Journal of Marine Science, 62(5):841–846, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/5/841/855778.

Kvamme:2007:EIL

[KB07] Cecilie Kvamme and Bjarte Bogstad. The effect of including length structure in yield-per-recruit estimates for northeast Arctic cod. ICES Journal of Marine Science, 64(2):357–368, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/2/357/2182328.

Kraak:2008:EMC

[KBDC^{+}08] S. B. M. Kraak, F. C. Buisman, M. Dickey-Collas, J. J. Poos, M. A. Pastoors, J. G. P. Smit, J. A. E. van Oostenbrugge, and N. Daan. The effect of management choices on the sustainability and economic performance of a mixed fishery: a simulation study. ICES Journal of Marine Science, 65(4):697–712, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/4/697/639883.

Kijewska:2009:MIE

[KBW09] Agnieszka Kijewska, Artur Burzyński, and Roman Wenne. Molecular identification of European flounder (Platichthys flesus) and its hybrids with European plaice (Pleuronectes platessa). ICES Journal of Marine Science, 66(5):902–906, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/5/902/664195.

Koueta:2000:BII

[KCBC00] N. Koueta, B. G. Castro, and E. Boucaud-Camou. Biochemical indices for instantaneous growth estimation in young cephalopod Sepia officinalis L. ICES Journal of Marine Science, 57(1):1–7, February 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/1/1/641102.

Kefalas:2003:ISD

[KCCM03] E. Kefalas, J. Castritsi-Catharios, and H. Miliou. The impacts of scallop dredging on sponge assemblages in the Gulf of Kalloni (Aegean Sea, northeastern Mediterranean). ICES Journal of Marine Science, 60(2):402–410, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/2/402/627638.

Kenny:2003:OSM

[KCD^{+}03] A. J. Kenny, I. Cato, M. Desprez, G. Fader, R. T. E. Schüttenhelm, and J. Side. An overview of seabed-mapping technologies in the context of marine habitat classification. ICES Journal of Marine Science, 60(2):411–418, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/2/411/627816.

Kell:2004:MMS

[KCL04] Laurence T. Kell, Walter W. Crozier, and Christopher M. Legault. Mixed and multi-stock fisheries. ICES Journal of Marine Science, 61(8):1330, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/61/8/1330/631058.

Kang:2009:EST

[KCL^{+}09] Donhyug Kang, Sungho Cho, Changwon Lee, Jung-Goo Myoung, and Jungyul Na. Ex situ target-strength measurements of Japanese anchovy (Engraulis japonicus) in the coastal Northwest Pacific. ICES Journal of Marine Science, 66(6):1219–1224, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1219/691397.

Kelly:2006:ISC

[KCR06] Ciaran J. Kelly, Edward A. Codling, and Emer Rogan. The Irish Sea cod recovery plan: some lessons learned. ICES Journal of Marine Science, 63(4):600–610, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/4/600/691607.

Kilada:2007:VAG

[KCR07] Raouf W. Kilada, Steven E. Campana, and Dale Roddick. Validated age, growth, and mortality estimates of the ocean quahog (Arctica islandica) in the western Atlantic. ICES Journal of Marine Science, 64(1):31–38, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/1/31/645669.

Kell:2009:LSE

[KDCH^{+}09] Laurence T. Kell, Mark Dickey-Collas, Niels T. Hintzen, Richard D. M. Nash, Graham M. Pilling, and Beatriz A. Roel. Lumpers or splitters? Evaluating recovery and management plans for metapopulations of herring. ICES Journal of Marine Science, 66 (8):1776–1783, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/8/1776/676197.

Korneliussen:2008:PCM

[KDO^{+}08] Rolf J. Korneliussen, Noel Diner, Egil Ona, Laurent Berger, and Paul G. Fernandes. Proposals for the collection of multifrequency acoustic data. ICES Journal of Marine Science, 65(6): 982–994, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/6/982/599852.

Kraak:2009:BSA

[KDP09] Sarah B. M. Kraak, Niels Daan, and Martin A. Pastoors. Biased stock assessment when using multiple, hardly overlapping, tuning series if fishing trends vary spatially. ICES Journal of Marine Science, 66(10):2272–2277, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/10/2272/679167.

Kjaersgaard:2008:EAM

[KF08] Jens Kjærsgaard and Hans Frost. Effort allocation and marine protected areas: is the North Sea plaice Box a management compromise? ICES Journal of Marine Science, 65(7):1203–1215, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/7/1203/642649.

Kang:2002:EAU

[KFM02] Myounghee Kang, Masahiko Furusawa, and Kazushi Miyashita. Effective and accurate use of difference in mean volume backscattering strength to identify fish and plankton. ICES Journal of Marine Science, 59(4):794–804, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/4/794/676804.

Kaiser:2007:ESA

[KGRW07] Michel J. Kaiser, Norman Graham, Craig S. Rose, and Peter H. Wiebe. Ecosystem-sensitive approaches to fishing: reconciling fisheries with conservation through improvements in fishing technology. ICES Journal of Marine Science, 64(8): 1610–1611, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/8/1610/614007.

Kvist:2001:SVA

[KGT01] T. Kvist, H. Gislason, and P. Thyregod. Sources of variation in the age composition of sandeel landings. ICES Journal of Marine Science, 58(4):842–851, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/4/842/630216.

Kang:2003:EST

[KH03a] Donhyug Kang and Doojin Hwang. Ex situ target strength of rockfish (Sebastes schlegeli) and red sea bream (Pagrus major) in the Northwest Pacific. ICES Journal of Marine Science, 60 (3):538–543, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/3/538/658886.

Kloser:2003:CUT

[KH03b] R. J. Kloser and J. K. Horne. Characterizing uncertainty in target-strength measurements of a deepwater fish: orange roughy (Hoplostethus atlanticus). ICES Journal of Marine Science, 60(3):516–523, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/60/3/516/658738.

Korneliussen:2009:CMS

[KHE^{+}09] Rolf J. Korneliussen, Yngve Heggelund, Inge K. Eliassen, Ola K. Øye, Tor Knutsen, and John Dalen. Combining multibeamsonar and multifrequency-echosounder data: examples of the analysis and imaging of large euphausiid schools. ICES Journal of Marine Science, 66(6):991–997, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/991/694325.

Korneliussen:2009:ASI

[KHEJ09] Rolf J. Korneliussen, Yngve Heggelund, Inge K. Eliassen, and Geir O. Johansen. Acoustic species identification of schooling fish. ICES Journal of Marine Science, 66(6):1111–1118, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1111/696107.

Krag:2009:VSF

[KHM09] Ludvig Ahm Krag, René Holst, and Niels Madsen. The vertical separation of fish in the aft end of a demersal trawl. ICES Journal of Marine Science, 66(4):772–777, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/4/772/603600.

Kenchington:2003:MMG

[KHN03] Ellen Kenchington, Mikko Heino, and Einar Eg Nielsen. Managing marine genetic diversity: time for action? ICES Journal of Marine Science, 60(6):1172–1176, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/6/1172/651070.

Kang:2006:ACW

[KHO06] Myounghee Kang, Satoshi Honda, and Tatsuki Oshima. Age characteristics of walleye pollock school echoes. ICES Journal of Marine Science, 63(8):1465–1476, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/8/1465/713370.

Ki:2008:RMJ

[KHS^{+}08] Jang-Seu Ki, Dae-Sik Hwang, Kyoungsoon Shin, Won Duk Yoon, Donghyun Lim, Young Shil Kang, Yoon Lee, and Jae-Seong Lee. Recent moon jelly (Aurelia sp.1) blooms in Korean coastal waters suggest global expansion: examples inferred from mitochondrial COI and nuclear ITS-5.8S rDNA sequences. ICES Journal of Marine Science, 65(3):443–452, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/3/443/786563.

Kaeriyama:2004:MCC

[KI04] Hideki Kaeriyama and Tsutomu Ikeda. Metabolism and chemical composition of mesopelagic ostracods in the western North Pacific Ocean. ICES Journal of Marine Science, 61(4):535–541, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/4/535/603941.

Koropitan:2009:IPP

[KIDY09] Alan F. Koropitan, Motoyoshi Ikeda, Ario Damar, and Yasuhiro Yamanaka. Influences of physical processes on the ecosystem of Jakarta Bay: a coupled physical–ecosystem model experiment. ICES Journal of Marine Science, 66(2):336–348, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/2/336/595705.

Kingsley:2002:FTI

[Kin02] Michael C. S. Kingsley. Food for thought ITQs and the economics of high-grading. ICES Journal of Marine Science, 59(4):649, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/4/649/676761.

Kastelle:2006:AVW

[KK06a] Craig R. Kastelle and Daniel K. Kimura. Age validation of walleye pollock (Theragra chalcogramma) from the Gulf of Alaska using the disequilibrium of pb-210 and ra-226. ICES Journal of Marine Science, 63(8):1520–1529, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/8/1520/714269.

Klevjer:2006:STS

[KK06b] Thor A. Klevjer and Stein Kaartvedt. In situ target strength and behaviour of northern krill (Meganyctiphanes norvegica). ICES Journal of Marine Science, 63(9):1726–1735, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/9/1726/700828.

Kingsley:2004:BRS

[KKC04] M. C. S. Kingsley, P. Kanneworff, and D. M. Carlsson. Buffered random sampling: a sequential inhibited spatial point process applied to sampling in a trawl survey for northern shrimp Pandalus borealis in West Greenland waters. ICES Journal of Marine Science, 61(1):12–24, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/1/12/699744.

Krafft:2006:GPP

[KKF^{+}06] Bjørn A. Krafft, Kit M. Kovacs, Anne Kirstine Frie, Tore Haug, and Christian Lydersen. Growth and population parameters of ringed seals (Pusa hispida) from Svalbard, Norway, 2002–2004. ICES Journal of Marine Science, 63(6):1136–1144, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/6/1136/617152.

Kim:2007:ENN

[KKS^{+}07] Heeyong Kim, Shingo Kimura, Akira Shinoda, Takashi Kitagawa, Yoshikazu Sasai, and Hideharu Sasaki. Effect of El Niño on migration and larval transport of the Japanese eel (Anguilla japonica). ICES Journal of Marine Science, 64(7):1387–1395, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/7/1387/726759.

Koster:2000:TCC

[KM00] Friedrich W. Köster and Christian Möllmann. Trophodynamic control by clupeid predators on recruitment success in Baltic cod? ICES Journal of Marine Science, 57(2):310–323, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/2/310/620438.

Kemp:2002:VFM

[KM02] Z. Kemp and G. Meaden. Visualization for fisheries management from a spatiotemporal perspective. ICES Journal of Marine Science, 59(1):190–202, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/59/1/190/650009.

Karlsson:2005:DHW

[KM05] Sten Karlsson and Jarle Mork. Deviation from Hardy–Weinberg equilibrium, and temporal instability in allele frequencies at microsatellite loci in a local population of Atlantic cod. ICES Journal of Marine Science, 62(8):1588–1596, ???? 2005. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/8/1588/789964.

Kell:2007:FOS

[KMG^{+}07] L. T. Kell, I. Mosqueira, P. Grosjean, J-M. Fromentin, D. Garcia, R. Hillary, E. Jardim, S. Mardle, M. A. Pastoors, J. J. Poos, F. Scott, and R. D. Scott. FLR: an open-source framework for the evaluation and development of management strategies. ICES Journal of Marine Science, 64(4):640–646, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/4/640/640024.

Koster:2005:BCR

[KMH^{+}05] Friedrich W. Köster, Christian Möllmann, Hans-Harald Hinrichsen, Kai Wieland, Jonna Tomkiewicz, Gerd Kraus, Rüdiger Voss, Andrei Makarchouk, Brian R. MacKenzie, Michael A. St. John, Dietrich Schnack, Norbert Rohlf, Tomasz Linkowski, and Jan E. Beyer. Baltic cod recruitment — the impact of climate variability on key processes. ICES Journal of Marine Science, 62 (7):1408–1425, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/7/1408/658985.

Kamenos:2004:SSD

[KMHS04] Nicholas A. Kamenos, P. Geoffrey Moore, and Jason M. Hall-Spencer. Small-scale distribution of juvenile gadoids in shallow inshore waters; what role does maerl play? ICES Journal of Marine Science, 61(3):422–429, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/3/422/672739.

Kang:2005:ITA

[KMI^{+}05] Donhyug Kang, Tohru Mukai, Kohji Iida, Doojin Hwang, and Jung-Goo Myoung. The influence of tilt angle on the acoustic target strength of the Japanese common squid (Todarodes pacificus). ICES Journal of Marine Science, 62(4):779–789, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/4/779/676303.

Kristiansen:2001:OST

[KMJH01] Tore S. Kristiansen, Kathrine Michalsen, Jan Arge Jacobsen, and Irene Huse. Optimal selection of temperature areas by juvenile cod (Gadus morhua L.) in the Barents Sea modelled by dynamic optimisation. ICES Journal of Marine Science, 58 (1):172–182, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/1/172/603514.

Kimoto:2007:MSR

[KMM07] A. Kimoto, T. Mouri, and T. Matsuishi. Modelling stock– recruitment relationships to examine stock management policies. ICES Journal of Marine Science, 64(5):870–877, July 2007.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/5/870/640713.

Korsbrekke:2001:SBA

[KMNP01] Knut Korsbrekke, Sigbjørn Mehl, Odd Nakken, and Michael Pennington. A survey-based assessment of the Northeast Arctic cod stock. ICES Journal of Marine Science, 58(4):763–769, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/4/763/630199.

Kobari:2008:SAE

[KMT08] Toru Kobari, Masatoshi Moku, and Kazutaka Takahashi. Seasonal appearance of expatriated boreal copepods in the Oyashio– Kuroshio mixed region. ICES Journal of Marine Science, 65 (3):469–476, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/3/469/784244.

Kiszka:2007:DER

[KMV^{+}07] Jeremy Kiszka, Kelly Macleod, Olivier Van Canneyt, Dylan Walker, and Vincent Ridoux. Distribution, encounter rates, and habitat characteristics of toothed cetaceans in the Bay of Biscay and adjacent waters from platform-of-opportunity data. ICES Journal of Marine Science, 64(5):1033–1043, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/5/1033/641915.

Karlsen:2006:EPE

[KNKT06] Ø. Karlsen, B. Norberg, O. S. Kjesbu, and G. L. Taranger. Effects of photoperiod and exercise on growth, liver size, and age at puberty in farmed Atlantic cod (Gadus morhua L.). ICES Journal of Marine Science, 63(2):355–364, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/2/355/641602.

Kimura:2000:BPM

[KNO00] Shingo Kimura, Hideaki Nakata, and Yuji Okazaki. Biological production in meso-scale eddies caused by frontal disturbances of the Kuroshio Extension. ICES Journal of Marine Science, 57 (1):133–142, February 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/1/133/641130.

Kawabe:2004:DMS

[KNS^{+}04] Ryo Kawabe, Yasuhiko Naito, Katsufumi Sato, Kazushi Miyashita, and Nariharu Yamashita. Direct measurement of the swimming speed, tailbeat, and body angle of Japanese flounder (Paralichthys olivaceus). ICES Journal of Marine Science, 61 (7):1080–1087, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/7/1080/878326.

Korsnes:2006:BGJ

[KNS^{+}06] Kjetil Korsnes, Ove Nicolaisen, Cecilie K. Skår, Audun H. Nerland, and Øivind Bergh. Bacteria in the gut of juvenile cod Gadus morhua fed live feed enriched with four different commercial diets. ICES Journal of Marine Science, 63(2):296–301, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/2/296/640303.

Knudsen:2009:LTE

[Knu09] Hans Petter Knudsen. Long-term evaluation of scientificechosounder performance. ICES Journal of Marine Science, 66 (6):1335–1340, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1335/690843.

Korneliussen:2002:OSP

[KO02] Rolf J. Korneliussen and Egil Ona. An operational system for processing and visualizing multi-frequency acoustic data. ICES Journal of Marine Science, 59(2):293–313, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/2/293/619638.

Korneliussen:2003:SEG

[KO03] Rolf J. Korneliussen and Egil Ona. Synthetic echograms generated from the relative frequency response. ICES Journal of Marine Science, 60(3):636–640, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/3/636/660334.

Koljonen:2006:ACP

[Kol06] Marja-Liisa Koljonen. Annual changes in the proportions of wild and hatchery Atlantic salmon (Salmo salar) caught in the Baltic Sea. ICES Journal of Marine Science, 63(7):1274–1285, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/7/1274/757261.

Korneliussen:2000:MRE

[Kor00] Rolf J. Korneliussen. Measurement and removal of echo integration noise. ICES Journal of Marine Science, 57(4):1204–1217, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/1204/647303.

Koslow:2000:CSD

[Kos00] J. Koslow. Continental slope and deep-sea fisheries: implications for a fragile ecosystem. ICES Journal of Marine Science, 57 (3):548–557, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/548/635930.

Koslow:2009:RAE

[Kos09] J. Anthony Koslow. The role of acoustics in ecosystem-based fishery management. ICES Journal of Marine Science, 66 (6):966–973, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/966/693235.

Koutsoubas:2000:MCS

[Kou00] D. Koutsoubas. Macrobenthic community structure and disturbance assessment in Gialova Lagoon, Ionian Sea. ICES Journal of Marine Science, 57(5):1472–1480, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/5/1472/660963.

Kraan:2007:DEC

[KPD^{+}07] Casper Kraan, Theunis Piersma, Anne Dekinga, Anita Koolhaas, and Jaap van der Meer. Dredging for edible cockles (Cerastoderma edule) on intertidal flats: short-term consequences of fisher patch-choice decisions for target and non-target benthic fauna. ICES Journal of Marine Science, 64(9):1735–1742, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/9/1735/785964.

Kraus:2009:MBE

[KPD^{+}09] Gerd Kraus, Dominique Pelletier, Julien Dubreuil, Christian Möllmann, Hans-Harald Hinrichsen, Francois Bastardie, Youen Vermard, and Stéphanie Mahévas. A model-based evaluation of marine protected areas: the example of eastern Baltic cod (Gadus morhua callarias L.). ICES Journal of Marine Science, 66(1):109–121, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/1/109/632831.

Kell:2005:EIM

[KPK^{+}05] L. T. Kell, G. M. Pilling, G. P. Kirkwood, M. Pastoors, B. Mesnil, K. Korsbrekke, P. Abaunza, R. Aps, A. Biseau, P. Kunzlik, C. Needle, B. A. Roel, and C. Ulrich-Rescan. An evaluation of the implicit management procedure used for some ICES roundfish stocks. ICES Journal of Marine Science, 62(4):750–759, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/4/750/675615.

Kell:2006:EMA

[KPK^{+}06] L. T. Kell, G. M. Pilling, G. P. Kirkwood, M. A. Pastoors, B. Mesnil, K. Korsbrekke, P. Abaunza, R. Aps, A. Biseau, P. Kunzlik, C. L. Needle, B. A. Roel, and C. Ulrich. An evaluation of multi-annual management strategies for ICES roundfish stocks. ICES Journal of Marine Science, 63(1):12–24, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/1/12/623288.

Kell:2005:ICC

[KPO05] Laurence T. Kell, Graham M. Pilling, and Carl M. O’Brien. Implications of climate change for the management of North Sea cod (Gadus morhua). ICES Journal of Marine Science, 62 (7):1483–1491, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/7/1483/660898.

Kell:2005:EMM

[KPS^{+}05] L. T. Kell, M. A. Pastoors, R. D. Scott, M. T. Smith, F. A. Van Beek, C. M. O’Brien, and G. M. Pilling. Evaluation of multiple management objectives for Northeast Atlantic flatfish stocks: sustainability vs. stability of yield. ICES Journal of Marine Science, 62(6):1104–1117, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/6/1104/617344.

Kieser:2005:DSN

[KRM05] Robert Kieser, Pall Reynisson, and Timothy J. Mulligan. Definition of signal-to-noise ratio and its critical role in split-beam measurements. ICES Journal of Marine Science, 62(1):123–130, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/1/123/666787.

Kloser:2009:AOM

[KRYL09] Rudy J. Kloser, Tim E. Ryan, Jock W. Young, and Mark E. Lewis. Acoustic observations of micronekton fish on the scale of an ocean basin: potential and challenges. ICES Journal of Marine Science, 66(6):998–1006, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/998/692393.

Kai:2008:EFM

[KS08] Mikihiko Kai and Kunio Shirakihara. Effectiveness of a feedback management procedure based on controlling the size of marine protected areas through catch per unit effort. ICES Journal of Marine Science, 65(7):1216–1226, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/7/1216/645727.

Kornilovs:2001:FZI

[KSD01] G. Kornilovs, L. Sidrevics, and J. W. Dippner. Fish and zooplankton interaction in the Central Baltic Sea. ICES Journal of Marine Science, 58(3):579–588, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/3/579/810081.

Karakassis:2000:ICF

[KTH^{+}00] I. Karakassis, M. Tsapakis, E. Hatziyanni, K.-N. Papadopoulou, and W. Plaiti. Impact of cage farming of fish on the seabed in three Mediterranean coastal areas. ICES Journal of Marine Science, 57(5):1462–1471, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/5/1462/660958.

Karaiskou:2005:DDA

[KTM^{+}05] Nikoletta Karaiskou, Alexander Triantafyllidis, Maritsa Margaroni, Dimitris Karatzas, and Costas Triantaphyllidis. A double DNA approach for identifying Macrorhamphosus scolopax (Pisces, Centriscidae). ICES Journal of Marine Science, 62(8): 1683–1690, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/8/1683/794512.

Kolstad:2006:BWS

[KTRG06] K. Kolstad, I. Thorland, T. Refstie, and B. Gjerde. Body weight, sexual maturity, and spinal deformity in strains and families of Atlantic cod (Gadus morhua) at two years of age at different locations along the Norwegian coast. ICES Journal of Marine Science, 63(2):246–252, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/63/2/246/638756.

Kress:2002:UCF

[KTS02] N. Kress, M. Tom, and E. Spanier. The use of coal fly ash in concrete for marine artificial reefs in the southeastern Mediterranean: compressive strength, sessile biota, and chemical composition. ICES Journal of Marine Science, 59(S1):S231–S237, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL https://academic.oup.com/icesjms/
article/59/suppl/S231/617946.

Kjesbu:2006:GMD

[KTT06] Olav Sigurd Kjesbu, Geir Lasse Taranger, and Edward A. Trippel. Gadoid mariculture: development and future challenges: Introduction. ICES Journal of Marine Science, 63(2):187–191, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/2/187/636912.

Kupschus:2004:TDR

[Kup04] Sven Kupschus. A temperature-dependent reproductive model for spotted seatrout (Cynoscion nebulosus) explaining spatiotemporal variations in reproduction and young-of-the-year recruitment in Florida estuaries. ICES Journal of Marine Science, 61(1):3–11, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/1/3/701456.

Katsanevakis:2006:SPD

[KV06] Stelios Katsanevakis and George Verriopoulos. Seasonal population dynamics of Octopus vulgaris in the eastern Mediterranean. ICES Journal of Marine Science, 63(1):151–160, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/1/151/623778.

Kingsley:2008:CBT

[KWBR08] Michael C. S. Kingsley, Kai Wieland, Bo Bergström, and Michael Rosing. Calibration of bottom trawls for northern shrimp. ICES Journal of Marine Science, 65(6):873– 881, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/6/873/602857.

Kennish:2002:SSL

[KWL^{+}02] Robin Kennish, Keith D. P. Wilson, John Lo, Shelley C. Clarke, and Steve Laister. Selecting sites for large-scale deployment of artificial reefs in Hong Kong: constraint mapping and prioritization techniques. ICES Journal of Marine Science, 59(S1): S164–S170, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S164/617916.

Kimura:2000:NCF

[KWZ00] Ryo Kimura, Yoshiro Watanabe, and Hiromu Zenitani. Nutritional condition of first-feeding larvae of Japanese sardine in the coastal and oceanic waters along the Kuroshio Current. ICES Journal of Marine Science, 57(2):240–248, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/2/240/620418.

Kalikhman:2003:DFA

[KYG03] I. Kalikhman, Y. Z. Yacobi, and M. Gophen. Distribution fields for aquatic ecosystem components: method of optimization of correlation zones. ICES Journal of Marine Science, 60(1):1–10, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/1/1/611418.

Livingston:2005:FEI

[LAB^{+}05] P. A. Livingston, K. Aydin, J. Boldt, J. Ianelli, and J. Jurado-Molina. A framework for ecosystem impacts assessment using an indicator approach. ICES Journal of Marine Science, 62 (3):592–597, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/3/592/666760.

Loland:2007:EDT

[LAO^{+}07] Anders Løland, Magne Aldrin, Egil Ona, Vidar Hjellvik, and Jens Christian Holst. Estimating and decomposing total uncertainty for survey-based abundance estimates of Norwegian spring-spawning herring. ICES Journal of Marine Science, 64 (7):1302–1312, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/64/7/1302/727524.

Law:2000:FSP

[Law00] Richard Law. Fishing, selection, and phenotypic evolution. ICES Journal of Marine Science, 57(3):659–668, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/3/659/635953.

Lawrence:2008:FFA

[Law08] Janice E. Lawrence. Furtive foes: algal viruses as potential invaders. ICES Journal of Marine Science, 65(5):716–722, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/5/716/713189.

Lawson:2001:SIP

[LBF01] Gareth L. Lawson, Manuel Barange, and Pierre Fréon. Species identification of pelagic fish schools on the South African continental shelf using acoustic descriptors and ancillary information.

ICES Journal of Marine Science, 58(1):275–287, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/1/275/603529.

Laffargue:2006:TPE

[LBL06] P. Laffargue, M-L. Bégout, and F. Lagardère. Testing the potential effects of shellfish farming on swimming activity and spatial distribution of sole (Solea solea) in a mesocosm. ICES Journal of Marine Science, 63(6):1014–1028, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/6/1014/614493.

Louanchi:2009:DCS

[LBN09] Ferial Louanchi, Meriem Boudjakdji, and Lamri Nacef. Decadal changes in surface carbon dioxide and related variables in the Mediterranean Sea as inferred from a coupled data-diagnostic model approach. ICES Journal of Marine Science, 66(7): 1538–1546, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/7/1538/657120.

Lang:2000:VGP

[LBNS00] Geoffrey M. Lang, Richard D. Brodeur, Jeffrey M. Napp, and Robert Schabetsberger. Variation in groundfish predation on juvenile walleye pollock relative to hydrographic structure near the Pribilof Islands, Alaska. ICES Journal of Marine Science, 57(2):265–271, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/2/265/620425.

Landry:2004:MPO

[LC04] Michael R. Landry and Albert Calbet. Microzooplankton production in the oceans. ICES Journal of Marine Science, 61 (4):501–507, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/4/501/603428.

Landaeta:2006:LDG

[LC06] Mauricio F. Landaeta and Leonardo R. Castro. Larval distribution and growth of the rockfish, Sebastes capensis (Sebastidae, Pisces), in the fjords of southern Chile. ICES Journal of Marine Science, 63(4):714–724, ???? 2006. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/4/714/694122.

LeQuesne:2009:MMS

[LC09a] W. J. F. Le Quesne and Edward A. Codling. Managing mobile species with MPAs: the effects of mobility, larval dispersal, and fishing mortality on closure size. ICES Journal of Marine Science, 66(1):122–131, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/1/122/634877.

Liu:2009:SIV

[LC09b] Guimei Liu and Fei Chai. Seasonal and interannual variability of primary and export production in the South China Sea: a three-dimensional physical–biogeochemical model study. ICES Journal of Marine Science, 66(2):420–431, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/2/420/595551.

Licandeo:2007:AGR

[LCC07] Roberto Licandeo, Francisco Cerna, and Renato Céspedes. Age, growth, and reproduction of the roughskin skate, Dipturus trachyderma, from the southeastern Pacific. ICES Journal of Marine Science, 64(1):141–148, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/1/141/647906.

Leon:2008:DMG

[LCC08] Roxana León, Leonardo R. Castro, and Mario Cáceres. Dispersal of Munida gregaria (Decapoda: Galatheidae) larvae in Patagonian channels of southern Chile. ICES Journal of Marine Science, 65(7):1131–1143, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/7/1131/643283.

Liu:2009:CSO

[LCC09] Yong Liu, Yong Chen, and Jiahua Cheng. A comparative study of optimization methods and conventional methods for sampling design in fishery-independent surveys. ICES Journal of Marine Science, 66(9):1873–1882, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/9/1873/722648.

Lennert-Cody:2008:EGC

[LCRS08] Cleridy E. Lennert-Cody, Jason J. Roberts, and Richard J. Stephenson. Effects of gear characteristics on the presence of bigeye tuna (Thunnus obesus) in the catches of the purse-seine fishery of the eastern Pacific Ocean. ICES Journal of Marine Science, 65(6):970–978, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/65/6/970/603434.

Lea:2003:FSL

[LD03a] Mary-Anne Lea and Laurent Dubroca. Fine-scale linkages between the diving behaviour of Antarctic fur seals and oceanographic features in the southern Indian Ocean. ICES Journal of Marine Science, 60(5):990–1002, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/5/990/771407.

Link:2003:THB

[LD03b] Jason S. Link and Chad Demarest. Trawl hangs, baby fish, and closed areas: a win–win scenario. ICES Journal of Marine Science, 60(5):930–938, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/60/5/930/769058.

Lindley:2005:VDC

[LD05] J. A. Lindley and S. Daykin. Variations in the distributions of Centropages chierchiae and Temora stylifera (Copepoda: Calanoida) in the north-eastern Atlantic Ocean and western European shelf waters. ICES Journal of Marine Science, 62 (5):869–877, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/5/869/856130.

Lebourges-Dhaussy:2009:ZSD

[LDCH^{+}09] Anne Lebourges-Dhaussy, Janet Coetzee, Larry Hutchings, Gildas Roudaut, and Cornelia Nieuwenhuys. Zooplankton spatial distribution along the South African coast studied by multifrequency acoustics, and its relationships with environmental parameters and anchovy distribution. ICES Journal of Marine Science, 66(6):1055–1062, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1055/696904.

Laurenson:2008:LBC

[LDM08] Chevonne H. Laurenson, Helen Dobby, and H. Anne McLay. The Lophius budegassa component of monkfish catches in Scottish waters. ICES Journal of Marine Science, 65(7):1346–1349, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/7/1346/644682.

Laurenson:2008:BFL

[LDML08] Chevonne H. Laurenson, Helen Dobby, H. Anne McLay, and Beth Leslie. Biological features of the Lophius piscatorius catch in Scottish waters. ICES Journal of Marine Science, 65(7): 1281–1290, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/7/1281/646155.

Liu:2008:MGC

[LDNS08] Hongbin Liu, Michael J. Dagg, Jeffrey M. Napp, and Riki Sato. Mesozooplankton grazing in the coastal Gulf of Alaska: Neocalanus spp. vs. other mesozooplankton. ICES Journal of Marine Science, 65(3):351–360, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/65/3/351/783181.

Landa:2008:GWA

[LDQ08] Jorge Landa, Rafael Duarte, and Iñaki Quincoces. Growth of white anglerfish (Lophius piscatorius) tagged in the Northeast Atlantic, and a review of age studies on anglerfish. ICES Journal of Marine Science, 65(1):72–80, January 2008. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/65/1/72/
613360.

Lloret:2002:ELS

[LdSSG02] Josep Lloret, Luis Gil de Sola, Arnauld Souplet, and René Galzin. Effects of large-scale habitat variability on condition of demersal exploited fish in the north-western Mediterranean. ICES Journal of Marine Science, 59(6):1215–1227, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/6/1215/608177.

LeQuesne:2009:FMG

[Le 09] Will J. F. Le Quesne. Are flawed MPAs any good or just a new way of making old mistakes? ICES Journal of Marine Science, 66(1):132–136, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/1/132/634780.

Laslett:2004:FGM

[LEP04] Geoff M. Laslett, J. Paige Eveson, and Tom Polacheck. Fitting growth models to length frequency data. ICES Journal of Marine Science, 61(2):218–230, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/2/218/620849.

Laurans:2009:MSA

[LFD^{+}09] Martial Laurans, Spyros Fifas, Sébastien Demaneche, Stéphane Brérette, and Olivier Debec. Modelling seasonal and annual variation in size at functional maturity in the European lobster (Homarus gammarus) from self-sampling data. ICES Journal of Marine Science, 66(9):1892–1898, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/9/1892/724978.

Lunneryd:2003:LMS

[LFW03] Sven Gunnar Lunneryd, Arne Fjälling, and Håkan Westerberg. A large-mesh salmon trap: a way of mitigating seal impact on a coastal fishery. ICES Journal of Marine Science, 60(6): 1194–1199, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/6/1194/652052.

Laran:2008:STP

[LG08] Sophie Laran and Alexandre Gannier. Spatial and temporal prediction of fin whale distribution in the northwestern Mediterranean Sea. ICES Journal of Marine Science, 65(7):1260–1269, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/7/1260/642360.

Large:2009:LAN

[LGH^{+}09] Philip A. Large, Norman G. Graham, Nils-Roar Hareide, Robert Misund, Dominic J. Rihan, Myles C. Mulligan, Peter J. Randall, David J. Peach, Philip H. McMullen, and Xavier Harlay. Lost and abandoned nets in deep-water gillnet fisheries in the Northeast Atlantic: retrieval exercises and outcomes. ICES Journal of Marine Science, 66(2):323–333, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/2/323/595693.

Lunde:2008:RTS

[LGR08] T. M. Lunde, O. R. Godø, and R. Rosland. Reliability of trawl surveys on cod in Norwegian fjords. ICES Journal of Marine Science, 65(6):937–945, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/65/6/937/602292.

Latour:2003:NTR

[LHHF03] Robert J. Latour, John M. Hoenig, Daniel A. Hepworth, and Stewart D. Frusher. A novel tag-recovery model with two size classes for estimating fishing and natural mortality, with implications for the southern rock lobster (Jasus edwardsii) in Tasmania, Australia. ICES Journal of Marine Science, 60(5): 1075–1085, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/5/1075/765950.

Laidre:2009:SSC

[LHHJ^{+}09] Kristin L. Laidre, Patrick J. Heagerty, Mads Peter Heide-Jørgensen, Lars Witting, and Malene Simon. Sexual segregation of common minke whales (Balaenoptera acutorostrata) in Greenland, and the influence of sea temperature on the sex ratio of catches. ICES Journal of Marine Science, 66(10):2253– 2266, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/10/2253/680315.

Lockyer:2001:ALR

[LHJJ^{+}01] C. Lockyer, M. P. Heide-Jørgensen, J. Jensen, C. C. Kinze, and T. Buus Sørensen. Age, length and reproductive parameters of harbour porpoises Phocoena phocoena (L.) from West Greenland. ICES Journal of Marine Science, 58(1):154–162, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/1/154/603507.

Laidre:2004:DOP

[LHJJT04] K. L. Laidre, M. P. Heide-Jørgensen, O. A. Jørgensen, and M. A. Treble. Deep-ocean predation by a high Arctic cetacean. ICES Journal of Marine Science, 61(3):430–440, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/3/430/673044.

Lokkeborg:2002:STV

[LHJS02] Svein Løkkeborg, Odd-Børre Humborstad, Terje Jørgensen, and Aud Vold Soldal. Spatio-temporal variations in gillnet catch rates in the vicinity of North Sea oil platforms. ICES Journal of Marine Science, 59(S1):S294–S299, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S294/617970.

LeGall:2000:DHB

[LHKGS00] A. C. Le Gall, D. J. Hydes, B. A. Kelly-Gerreyn, and D. J. Slinn. Development of a 2D horizontal biogeochemical model for the Irish Sea DYMONIS. ICES Journal of Marine Science, 57(4): 1050–1059, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/4/1050/647252.

Lindstrom:2002:PHC

[LHR02] Ulf Lindstrøm, Tore Haug, and Ingolf Røttingen. Predation on herring, Clupea harengus, by minke whales, Balaenoptera acutorostrata, in the Barents Sea. ICES Journal of Marine Science, 59(1):58–70, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/1/58/650031.

Link:2005:TEI

[Lin05] Jason S. Link. Translating ecosystem indicators into decision criteria. ICES Journal of Marine Science, 62(3):569–576, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/3/569/665783.

Litzow:2006:CRS

[Lit06] Michael A. Litzow. Climate regime shifts and community reorganization in the Gulf of Alaska: how do recent shifts compare with 1976/1977? ICES Journal of Marine Science, 63 (8):1386–1396, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/8/1386/711232.

Livingston:2000:MVP

[LJM00] Patricia A. Livingston and Jesus Jurado-Molina. A multispecies virtual population analysis of the eastern Bering Sea. ICES Journal of Marine Science, 57(2):294–299, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/2/294/620432.

Lee:2009:ADS

[LKK^{+}09] Gwang H. Lee, Han J. Kim, Dae C. Kim, Bo Y. Yi, Seong M. Nam, Boo K. Khim, and Moon S. Lim. The acoustic diversity of the seabed based on the similarity index computed from chirp seismic data. ICES Journal of Marine Science, 66(2):227–236, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/2/227/593105.

Lee:2008:MID

[LKL08] Ji Hoon Lee, Ludvig Karlsen, and Chun Woo Lee. A method for improving the dynamic simulation efficiency of underwater flexible structures by implementing non-active points in modelling. ICES Journal of Marine Science, 65(9):1552–1558, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/9/1552/628725.

Lan:2008:SRB

[LLC^{+}08] Yang-Chi Lan, Ming-An Lee, Wen-Yu Chen, Feng-Jen Hsieh, Jia-Yi Pan, Don-Chung Liu, and Wei-Cheng Su. Seasonal relationships between the copepod community and hydrographic conditions in the southern East China Sea. ICES Journal of Marine Science, 65(3):462–468, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/3/462/782894.

Lajus:2005:UHC

[LLD^{+}05] Dmitry L. Lajus, Julia A. Lajus, Zoya V. Dmitrieva, Alexei V. Kraikovski, and Daniel A. Alexandrov. The use of historical catch data to trace the influence of climate on fish populations: examples from the white and Barents Sea fisheries in the 17th and 18th centuries. ICES Journal of Marine Science, 62(7): 1426–1435, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/7/1426/659120.

Labansen:2007:SDR

[LLHK07] Aili L. Labansen, Christian Lydersen, Tore Haug, and Kit M. Kovacs. Spring diet of ringed seals (Phoca hispida) from northwestern Spitsbergen, Norway. ICES Journal of Marine Science, 64(6):1246–1256, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/64/6/1246/615021.

Lloret:2000:TSM

[LLS00] Josep Lloret, Jordi Lleonart, and Ignasi Solé. Time series modelling of landings in Northwest Mediterranean Sea. ICES Journal of Marine Science, 57(1):171–184, February 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/1/171/641135.

Lacoste:2001:ITS

[LMC^{+}01] Karine N. Lacoste, Jean Munro, Martin Castonguay, François J. Saucier, and Jacques A. Gagné. The influence of tidal streams on the pre-spawning movements of Atlantic herring, Clupea harengus L., in the st Lawrence estuary. ICES Journal of Marine Science, 58(6):1286–1298, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/6/1286/641576.

Lucifora:2002:REA

[LME02] Luis O. Lucifora, Roberto C. Menni, and Alicia H. Escalante. Reproductive ecology and abundance of the sand tiger shark, Carcharias taurus, from the southwestern Atlantic. ICES Journal of Marine Science, 59(3):553–561, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/3/553/610833.

Lucifora:2005:RSO

[LME05] Luis O. Lucifora, Roberto C. Menni, and Alicia H. Escalante. Reproduction and seasonal occurrence of the copper shark, Carcharhinus brachyurus, from north Patagonia, Argentina. ICES Journal of Marine Science, 62(1):107–115, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/1/107/665970.

LeGresley:2008:NIT

[LMM^{+}08] Murielle M. LeGresley, Jennifer L. Martin, Paul McCurdy, Bruce Thorpe, and Blythe D. Chang. Non-indigenous tunicate species in the Bay of Fundy, eastern Canada. ICES Journal of Marine Science, 65(5):770–774, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/65/5/770/712722.

Lok:2002:ART

[LMU^{+}02] Altan Lök, Cengiz Metin, Ali Ulaş, F. Ozan Düzbastilar, and Adnan Tokaç. Artificial reefs in Turkey. ICES Journal of Marine Science, 59(S1):S192–S195, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S192/617926.

Lavin:2007:EVN

[LMVdZ^{+}07] Alicia Lavín, Xabier Moreno-Ventas, Victoria Ortiz de Zárate, Pablo Abaunza, and José Manuel Cabanas. Environmental variability in the North Atlantic and Iberian waters and its influence on horse mackerel (Trachurus trachurus) and albacore (Thunnus alalunga) dynamics. ICES Journal of Marine Science, 64 (3):425–438, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/3/425/812225.

Lewy:2003:MSF

[LN03] P. Lewy and A. Nielsen. Modelling stochastic fish stock dynamics using Markov chain Monte Carlo. ICES Journal of Marine Science, 60(4):743–752, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/60/4/743/691687.

Lundgren:2008:MPS

[LN08] Bo Lundgren and J. Rasmus Nielsen. A method for the possible species discrimination of juvenile gadoids by broad-bandwidth backscattering spectra vs. angle of incidence. ICES Journal of Marine Science, 65(4):581–593, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/4/581/636909.

Lee:2005:SSS

[LND05] O. Lee, R. D. M. Nash, and B. S. Danilowicz. Small-scale spatiotemporal variability in ichthyoplankton and zooplankton distribution in relation to a tidal-mixing front in the Irish Sea. ICES Journal of Marine Science, 62(6):1021–1036, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/6/1021/615693.

Lambert:2009:MGP

[LNLS09] Gwladys Lambert, J. Rasmus Nielsen, Lena I. Larsen, and Henrik Sparholt. Maturity and growth population dynamics of Norway pout (Trisopterus esmarkii) in the North Sea, Skagerrak, and Kattegat. ICES Journal of Marine Science, 66(9): 1899–1914, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/9/1899/722216.

Lucena:2005:CDS

[LO05] Flávia Lucena and Carl M. O’Brien. The consequences of different scenarios in the management of the gillnet and purse-seine fisheries targeting Pomatomus saltatrix and Cynoscion guatucupa in southern Brazil: a bio-economic approach. ICES Journal of Marine Science, 62(2):201–213, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/2/201/603273.

Lopez:2006:IBM

[Lóp06] Bruno Díaz López. Interactions between Mediterranean bottlenose dolphins (Tursiops truncatus) and gillnets off Sardinia, Italy. ICES Journal of Marine Science, 63(5):946–951, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/5/946/664268.

Landa:2000:MLW

[LP00] J. Landa and C. Piñeiro. Megrim (Lepidorhombus whiffiagonis) growth in the North-eastern Atlantic based on backcalculation of otolith rings. ICES Journal of Marine Science, 57(4):1077–1090, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/4/1077/647266.

Litzow:2000:MTS

[LPA^{+}00] Michael A. Litzow, John F. Piatt, Alisa A. Abookire, Alexander K. Prichard, and Martin D. Robards. Monitoring temporal and spatial variability in sandeel (Ammodytes hexapterus) abundance with pigeon guillemot (Cepphus columba) diets. ICES Journal of Marine Science, 57(4):976–986, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/4/976/647401.

Lo:2008:EJA

[LPH^{+}08] Wen-Tseng Lo, Jennifer E. Purcell, Jia-Jang Hung, Huei-Meei Su, and Pei-Kai Hsu. Enhancement of jellyfish (Aurelia aurita) populations by extensive aquaculture rafts in a coastal lagoon in Taiwan. ICES Journal of Marine Science, 65(3):453–461, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/3/453/783857.

LePennec:2003:PLP

[LPL03] M. Le Pennec, A. Paugam, and G. Le Pennec. The pelagic life of the pectinid Pecten maximus — a review. ICES Journal of Marine Science, 60(2):211–223, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/2/211/623789.

Little:2009:DRA

[LPM^{+}09] L. Richard Little, André E. Punt, Bruce D. Mapstone, Gavin A. Begg, Barry Goldman, and Nick Ellis. Different responses to area closures and effort controls for sedentary and migratory harvested species in a multispecies coral reef linefishery. ICES Journal of Marine Science, 66(9):1931–1941, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/9/1931/724338.

Luna:2009:BIR

[LPSL09] Beatriz Luna, Carlos Valle Pérez, and Jose Luis Sánchez-Lizaso. Benthic impacts of recreational divers in a Mediterranean marine protected area. ICES Journal of Marine Science, 66 (3):517–523, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/3/517/817490.

lePape:2001:IVS

[lPV01] O. le Pape and J. Vigneau. The influence of vessel size and fishing strategy on the fishing effort for multispecies fisheries in northwestern France. ICES Journal of Marine Science, 58 (6):1232–1242, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1232/641560.

Lehtonen:2004:MSI

[LS04] Esa Lehtonen and Petri Suuronen. Mitigation of seal-induced damage in salmon and whitefish trapnet fisheries by modification of the fish bag. ICES Journal of Marine Science, 61(7): 1195–1200, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/7/1195/881968.

Lekve:2002:SRE

[LSGD02] Kyrre Lekve, Nils Chr. Stenseth, Jakob Gjøsæter, and Sylvain Dolédec. Species richness and environmental conditions of fish along the Norwegian Skagerrak coast. ICES Journal of Marine Science, 59(4):757–769, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/59/4/757/676792.

Lin:2009:ESS

[LSH^{+}09] Hsing-Juh Lin, Kwang-Tsao Shao, Hwey-Lian Hsieh, Wen-Tseng Lo, and Xiao-Xun Dai. The effects of system-scale removal of oyster-culture racks from Tapong Bay, southwestern Taiwan: model exploration and comparison with field observations. ICES Journal of Marine Science, 66(5):797–810, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/5/797/663519.

Leitao:2007:CAR

[LSM07] Francisco Leitão, Miguel N. Santos, and Carlos C. Monteiro. Contribution of artificial reefs to the diet of the white sea bream (Diplodus sargus). ICES Journal of Marine Science, 64 (3):473–478, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/3/473/818513.

Livingston:2000:FBE

[LT00] Patricia A. Livingston and Sigurd Tjelmeland. Fisheries in boreal ecosystems. ICES Journal of Marine Science, 57(3): 619–627, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/619/635942.

Link:2006:EIC

[LT06] P. Michael Link and Richard S. J. Tol. Economic impacts of changes in the population dynamics of fish on the fisheries of the Barents Sea. ICES Journal of Marine Science, 63(4):611–625, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/4/611/691788.

Lagardere:2000:FOM

[LTA00] F. Lagardère, K. Thibaudeau, and M. L. Bégout Anras. Feasibility of otolith markings in large juvenile turbot, Scophthalmus maximus, using immersion in alizarin-red S solutions. ICES Journal of Marine Science, 57(4):1175–1181, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/4/1175/647294.

Lorenzo:2009:STC

[LTI09] María N. Lorenzo, Juan J. Taboada, and Isabel Iglesias. Sensitivity of thermohaline circulation to decadal and multidecadal variability. ICES Journal of Marine Science, 66(7):1439–1447, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/7/1439/657910.

Lunneryd:2001:FPH

[Lun01] S. G. Lunneryd. Fish preference by the harbour seal (Phoca vitulina), with implications for the control of damage to fishing gear. ICES Journal of Marine Science, 58(4):824–829, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/4/824/630212.

Lindegarth:2000:ETD

[LVHU00] M. Lindegarth, D. Valentinsson, M. Hansson, and M. Ulmestrand. Effects of trawling disturbances on temporal and spatial structure of benthic soft-sediment assemblages in Gullmarsfjorden, Sweden. ICES Journal of Marine Science, 57(5):1369–1376, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/5/1369/660936.

Logerwell:2004:SDF

[LW04] Elizabeth A. Logerwell and Christopher D. Wilson. Species discrimination of fish using frequency-dependent acoustic backscatter. ICES Journal of Marine Science, 61(6):1004–1013, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/6/1004/675855.

Leporati:2009:ASS

[LZS09] Stephen C. Leporati, Philippe E. Ziegler, and Jayson M. Semmens. Assessing the stock status of holobenthic octopus fisheries: is catch per unit effort sufficient? ICES Journal of Marine Science, 66(3):478–487, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/3/478/813463.

Milligan:2009:ETP

[MAAN09] R. J. Milligan, A. Albalat, R. J. A. Atkinson, and D. M. Neil. The effects of trawling on the physical condition of the Norway lobster Nephrops norvegicus in relation to seasonal cycles in the Clyde Sea area. ICES Journal of Marine Science, 66(3):488–494, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/3/488/816719.

Mahe:2007:OSV

[MAB^{+}07] K. Mahe, R. Amara, T. Bryckaert, M. Kacher, and J. M. Brylinski. Ontogenetic and spatial variation in the diet of hake (Merluccius merluccius) in the Bay of Biscay and the Celtic Sea.

ICES Journal of Marine Science, 64(6):1210–1219, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/6/1210/615539.

Marchal:2007:ITC

[MAC^{+}07] Paul Marchal, Bo Andersen, B. Caillart, Ole Eigaard, Olivier Guyader, Holger Hovgaard, Ane Iriondo, Fanny Le Fur, Jacques Sacchi, and Marina Santurtún. Impact of technological creep on fishing effort and fishing mortality, for a selection of European fleets. ICES Journal of Marine Science, 64(1):192–209, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/1/192/648309.

MacCall:2009:DCA

[Mac09] Alec D. MacCall. Depletion-corrected average catch: a simple formula for estimating sustainable yields in data-poor situations. ICES Journal of Marine Science, 66(10):2267–2271, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/10/2267/682739.

Mamedov:2006:BAK

[Mam06] Elchin V. Mamedov. The biology and abundance of kilka (Clupeonella spp.) along the coast of Azerbaijan, Caspian Sea. ICES Journal of Marine Science, 63(9):1665–1673, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/9/1665/699036.

Martinez-Abrain:2002:DTW

[MAMO02] Alejandro Martínez-Abraín, Raquel Maestre, and Daniel Oro. Demersal trawling waste as a food source for Western Mediterranean seabirds during the summer. ICES Journal of Marine Science, 59(3):529–537, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/59/3/529/610826.

Mooney:2007:ASP

[MANT07] T. Aran Mooney, Whitlow W. L. Au, Paul E. Nachtigall, and Edward A. Trippel. Acoustic and stiffness properties of gillnets as they relate to small cetacean bycatch. ICES Journal of Marine Science, 64(7):1324–1332, October 2007. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/7/1324/729624.

Martins:2007:GVA

[Mar07] Maria Manuel Martins. Growth variability in Atlantic mackerel (Scomber scombrus) and Spanish mackerel (Scomber japonicus) off Portugal. ICES Journal of Marine Science, 64(9): 1785–1790, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/9/1785/788819.

Marchal:2008:CAM

[Mar08] Paul Marchal. A comparative analysis of métiers and catch profiles for some French demersal and pelagic fleets. ICES Journal of Marine Science, 65(4):674–686, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/4/674/639777.

Martinez:2006:GSP

[MASA06] Iciar Martinez, Michaela Aschan, Taran Skjerdal, and Salah M. Aljanabi. The genetic structure of Pandalus borealis in the Northeast Atlantic determined by RAPD analysis. ICES Journal of Marine Science, 63(5):840–850, ???? 2006. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/5/840/662516.

Macy:2001:SMS

[MB01] William K. Macy III and Jon K. T. Brodziak. Seasonal maturity and size at age of Loligo pealeii in waters of southern New England. ICES Journal of Marine Science, 58(4):852–864, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/4/852/630219.

Morgan:2005:ECR

[MB05] M. Joanne Morgan and John Brattey. Effect of changes in reproductive potential on perceived productivity of three Northwest Atlantic cod (Gadus morhua) stocks. ICES Journal of Marine Science, 62(1):65–74, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/62/1/65/669269.

Morris:2006:HSM

[MB06] Liz Morris and David Ball. Habitat suitability modelling of economically important fish species with commercial fisheries data. ICES Journal of Marine Science, 63(9):1590–1603, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/9/1590/696978.

Maiorano:2009:SCP

[MBC^{+}09] Luigi Maiorano, Valerio Bartolino, Francesco Colloca, Alvaro Abella, Andrea Belluscio, Paolo Carpentieri, Alessandro Criscoli, Giovanna Jona Lasinio, Alessandro Mannini, Fabio Pranovi, Bruno Reale, Giulio Relini, Claudio Viva, and Gian Domenico Ardizzone. Systematic conservation planning in the Mediterranean: a flexible tool for the identification of notake marine protected areas. ICES Journal of Marine Science, 66(1):137–146, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/1/137/632258.

Mohn:2002:OMF

[MBM02] Christian Mohn, Joachim Bartsch, and Jens Meincke. Observations of the mass and flow field at Porcupine Bank. ICES Journal of Marine Science, 59(2):380–392, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/2/380/619660.

Macbeth:2006:RST

[MBPW06] William G. Macbeth, Matt K. Broadhurst, Brian D. Paterson, and Michael E. L. Wooden. Reducing the short-term mortality of juvenile school prawns (Metapenaeus macleayi) discarded during trawling. ICES Journal of Marine Science, 63(5):831–839, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/5/831/662409.

Munch:2000:RDB

[MC00] Stephan B. Munch and David O. Conover. Recruitment dynamics of bluefish (Pomatomus saltatrix) from Cape Hatteras to Cape Cod, 1973–1995. ICES Journal of Marine Science, 57 (2):393–402, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/2/393/620468.

Mohn:2007:HCR

[MC07] R. K. Mohn and G. A. Chouinard. Harvest control rules for stocks displaying dynamic production regimes. ICES Journal of Marine Science, 64(4):693–697, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/693/641946.

McPhie:2009:BDA

[MC09] Romney P. McPhie and Steven E. Campana. Bomb dating and age determination of skates (family Rajidae) off the eastern coast of Canada. ICES Journal of Marine Science, 66(3):546–560, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/3/546/815192.

MacLennan:2004:EDF

[MCAS04] D. N. MacLennan, P. J. Copland, E. Armstrong, and E. J. Simmonds. Experiments on the discrimination of fish and seabed echoes. ICES Journal of Marine Science, 61(2):201–210, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/2/201/620601.

MacNeil:2009:SDR

[MCB09] M. Aaron MacNeil, John K. Carlson, and Lawrence R. Beerkircher. Shark depredation rates in pelagic longline fisheries: a case study from the Northwest Atlantic. ICES Journal of Marine Science, 66(4):708–719, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/4/708/602561.

Muino:2003:CSS

[MCI03] Ramón Muiño, Pablo Carrera, and Magdalena Iglesias. The characterization of sardine (Sardina pilchardus Walbaum) schools off the Spanish–Atlantic coast. ICES Journal of Marine Science, 60(6):1361–1372, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/6/1361/654532.

Melvin:2003:ECA

[MCL03] Gary D. Melvin, Norman A. Cochrane, and Yanchao Li. Extraction and comparison of acoustic backscatter from a calibrated multi- and single-beam sonar. ICES Journal of Marine Science, 60(3):669–677, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/3/669/660941.

Miller:2000:OSV

[MCM00] C. B. Miller, J. A. Crain, and C. A. Morgan. Oil storage variability in Calanus finmarchicus. ICES Journal of Marine Science, 57(6):1786–1799, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/6/1786/768668.

Moursund:2003:FAD

[MCP03] Russell A. Moursund, Thomas J. Carlson, and Rock D. Peters. A fisheries application of a dual-frequency identification sonar acoustic camera. ICES Journal of Marine Science, 60 (3):678–683, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/3/678/661054.

Mantovani:2006:ICE

[MCRF06] S. Mantovani, G. Castaldelli, R. Rossi, and E. A. Fano. The infaunal community in experimentally seeded low and high density Manila clam (Tapes philippinarum) beds in a Po River Delta lagoon (Italy). ICES Journal of Marine Science, 63(5):860–866, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/5/860/662817.

Mormede:2001:PPR

[MD01] Sophie Mormede and Ian M. Davies. Polychlorobiphenyl and pesticide residues in monkfish Lophius piscatorius and black scabbard Aphanopus carbo from the Rockall Trough. ICES Journal of Marine Science, 58(3):725–736, ???? 2001. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/3/725/810147.

Moschino:2003:ESD

[MDM03] V. Moschino, M. Deppieri, and M. G. Marin. Evaluation of shell damage to the clam Chamelea gallina captured by hydraulic dredging in the Northern Adriatic Sea. ICES Journal of Marine Science, 60(2):393–401, ???? 2003. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/2/393/627465.

Mandelman:2007:PSM

[MF07] John W. Mandelman and Marianne A. Farrington. The physiological status and mortality associated with otter-trawl capture, transport, and captivity of an exploited elasmobranch, Squalus acanthias. ICES Journal of Marine Science, 64(1):122–130, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/1/122/646027.

Morello:2007:UTF

[MFA07] Elisabetta B. Morello, Carlo Froglia, and R. J. A. Atkinson. Underwater television as a fishery-independent method for stock assessment of Norway lobster (Nephrops norvegicus) in the central Adriatic Sea (Italy). ICES Journal of Marine Science, 64 (6):1116–1123, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/64/6/1116/614219.

Mauna:2008:CFV

[MFB^{+}08] A. Cecilia Mauna, Bárbara C. Franco, Ana Baldoni, E. Marcelo Acha, Mario L. Lasta, and Oscar O. Iribarne. Cross-front variations in adult abundance and recruitment of Patagonian scallop (Zygochlamys patagonica) at the SW Atlantic shelf break front. ICES Journal of Marine Science, 65(7):1184–1190, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/7/1184/644139.

Maclennan:2002:CAD

[MFD02] David N. Maclennan, Paul G. Fernandes, and John Dalen. A consistent approach to definitions and symbols in fisheries acoustics. ICES Journal of Marine Science, 59(2):365–369, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/2/365/619652.

Matsushita:2004:RJF

[MFIO04] Yoshiki Matsushita, Kaoru Fujita, Naoya Ikegami, and Satoshi Ohata. Reaction of juvenile flounder to grid separators. ICES Journal of Marine Science, 61(7):1174–1178, ???? 2004. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1174/881710.

Morris:2002:BCR

[MG02] Corey J. Morris and John M. Green. Biological characteristics of a resident population of Atlantic cod (Gadus morhua L.) in southern Labrador. ICES Journal of Marine Science, 59 (4):666–678, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/4/666/676765.

Mabragana:2007:FEA

[MG07] Ezequiel Mabragaña and Diego A. Giberto. Feeding ecology and abundance of two sympatric skates, the shortfin sand skate Psammobatis normani McEachran, and the smallthorn sand skate P. rudis Günther (Chondrichthyes, Rajidae), in the southwest Atlantic. ICES Journal of Marine Science, 64(5): 1017–1027, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/5/1017/643282.

Madsen:2008:WEM

[MGH08] Matias L. Madsen, Eilif Gaard, and Benni W. Hansen. Waxester mobilization by female Calanus finmarchicus (Gunnerus) during spring ascendance and advection to the Faroe Shelf. ICES Journal of Marine Science, 65(7):1112–1121, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/7/1112/644042.

MacLennan:2003:AFA

[MGM03] David N. MacLennan, François Gerlotto, and Jacques Massé. Acoustics in fisheries and aquatic ecology: Part 1 introduction. ICES Journal of Marine Science, 60(3):435–436, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/3/435/657700.

Marteinsdottir:2000:SVH

[MGS00] Gudrun Marteinsdottir, Björn Gunnarsson, and Iain M. Suthers. Spatial variation in hatch date distributions and origin of pelagic juvenile cod in Icelandic waters. ICES Journal of Marine Science, 57(4):1182–1195, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/57/4/1182/647297.

Marteinsdottir:2001:E

[MGS01] Gudrun Marteinsdottir, Björn Gunnarsson, and Iain M. Suthers. Erratum. ICES Journal of Marine Science, 58(1):361, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/1/361/603547.

Marteinsdottir:2000:SVA

[MGTS00] Gudrun Marteinsdottir, Asta Gudmundsdottir, Vilhjalmur Thorsteinsson, and Gunnar Stefansson. Spatial variation in abundance, size composition and viable egg production of spawning cod (Gadus morhua L.) in Icelandic waters. ICES Journal of Marine Science, 57(4):824–830, August 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/4/824/647342.

McCollum:2006:MCF

[MGvH06] Artie McCollum, Jessica Geubtner, and Ione Hunt von Herbing. Metabolic cost of feeding in Atlantic cod (Gadus morhua) larvae using microcalorimetry. ICES Journal of Marine Science, 63 (2):335–339, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/2/335/641070.

Megrey:2001:ETF

[MH01] Bernard A. Megrey and Sarah Hinckley. Effect of turbulence on feeding of larval fishes: a sensitivity analysis using an individualbased model. ICES Journal of Marine Science, 58(5):1015–1029, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/5/1015/730812.

Megrey:2002:USV

[MHD02] Bernard A. Megrey, Sarah Hinckley, and Elizabeth L. Dobbins. Using scientific visualization tools to facilitate analysis of multi-dimensional data from a spatially explicit, biophysical, individual-based model of marine fish early life history.

ICES Journal of Marine Science, 59(1):203–215, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/1/203/650012.

Masden:2009:BMI

[MHF^{+}09] Elizabeth A. Masden, Daniel T. Haydon, Anthony D. Fox, Robert W. Furness, Rhys Bullman, and Mark Desholm. Barriers to movement: impacts of wind farms on migrating birds. ICES Journal of Marine Science, 66(4):746–753, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/4/746/603346.

Maunder:2006:IPU

[MHH06] Mark N. Maunder, Shelton J. Harley, and John Hampton. Including parameter uncertainty in forward projections of computationally intensive statistical population dynamic models. ICES Journal of Marine Science, 63(6):969–979, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/6/969/618044.

Moore:2009:SPD

[MHV09] Cordelia H. Moore, Euan S. Harvey, and Kimberly P. Van Niel. Spatial prediction of demersal fish distributions: enhancing our understanding of species–environment relationships. ICES Journal of Marine Science, 66(9):2068–2075, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/9/2068/727082.

Miller:2002:UEP

[Mil02] Margaret W. Miller. Using ecological processes to advance artificial reef goals. ICES Journal of Marine Science, 59(S1):S27–S31, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL https://academic.oup.com/icesjms/
article/59/suppl/S27/617960.

Miller:2008:CGD

[Mil08] Charles B. Miller. Copepod growth in detail: pattern similarity to decapod larvae. ICES Journal of Marine Science, 65 (3):332–338, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/3/332/783444.

Mistri:2002:PBC

[Mis02] Michele Mistri. Persistence of benthic communities: a case study from the Valli di Comacchio, a Northern Adriatic lagoonal ecosystem (Italy). ICES Journal of Marine Science, 59 (2):314–322, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/2/314/619640.

Miyashita:2003:DCA

[Miy03] Kazushi Miyashita. Diurnal changes in the acoustic-frequency characteristics of Japanese anchovy (Engraulis japonicus) postlarvae “shirasu” inferred from theoretical scattering models. ICES Journal of Marine Science, 60(3):532–537, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/3/532/658843.

Moloney:2005:CBH

[MJA^{+}05] Coleen L. Moloney, Astrid Jarre, Hugo Arancibia, Yves-Marie Bozec, Sergio Neira, Jean-Paul Roux, and Lynne J. Shannon. Comparing the Benguela and Humboldt marine upwelling ecosystems with indicators derived from inter-calibrated models. ICES Journal of Marine Science, 62(3):493–502, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/3/493/664131.

Michalsen:2008:FMC

[MJB08] Kathrine Michalsen, Edda Johannesen, and Bjarte Bogstad. Feeding of mature cod (Gadus morhua) on the spawning grounds in Lofoten. ICES Journal of Marine Science, 65(4):571–580, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/4/571/636015.

McDonald:2001:EWM

[MKB01] A. David McDonald, Terese H. Kendrick, and Paul A. Breen. Empirical weighting of multiple stock-abundance indices for parameter estimation and stock assessment in a multi-zone or multi-species fishery. ICES Journal of Marine Science, 58(1):

204–215, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/1/204/603518.

Mollmann:2005:CZP

[MKFK05] Christian Möllmann, Georgs Kornilovs, Marina Fetter, and Friedrich W. Köster. Climate, zooplankton, and pelagic fish growth in the central Baltic Sea. ICES Journal of Marine Science, 62(7):1270–1280, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/62/7/1270/656228.

Mecenero:2006:RFC

[MKR06] Silvia Mecenero, Stephen P. Kirkman, and Jean-Paul Roux. A refined fish consumption model for lactating Cape fur seals (Arctocephalus pusillus pusillus), based on scat analyses. ICES Journal of Marine Science, 63(8):1551–1566, ???? 2006. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/8/1551/715166.

Mantyniemi:2009:VIF

[MKR^{+}09] Samu Mäntyniemi, Sakari Kuikka, Mika Rahikainen, Laurence T. Kell, and Veijo Kaitala. The value of information in fisheries management: North Sea herring as an example. ICES Journal of Marine Science, 66(10):2278–2283, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/10/2278/682372.

Martin:2008:NPS

[ML08] Jennifer L. Martin and Murielle M. LeGresley. New phytoplankton species in the Bay of Fundy since 1995. ICES Journal of Marine Science, 65(5):759–764, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/5/759/712876.

Maggio:2009:MAR

[MLG^{+}09] Teresa Maggio, Sabrina Lo Brutto, Flavio Garoia, Fausto Tinti, and Marco Arculeo. Microsatellite analysis of red mullet Mullus barbatus (Perciformes, Mullidae) reveals the isolation of the Adriatic Basin in the Mediterranean Sea. ICES Journal of Marine Science, 66(9):1883–1891, October 2009. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/9/1883/722978.

Mumford:2009:IMM

[MLLK09] J. D. Mumford, A. W. Leach, P. Levontin, and L. T. Kell. Insurance mechanisms to mediate economic risks in marine fisheries. ICES Journal of Marine Science, 66(5):950–959, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/5/950/664078.

Mayer:2002:VPF

[MLM02] Larry Mayer, Yanchao Li, and Gary Melvin. 3D visualization for pelagic fisheries research and assessment. ICES Journal of Marine Science, 59(1):216–225, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/1/216/650015.

Megrey:2005:CAS

[MLM05] Bernard A. Megrey, Yong-Woo Lee, and S. Allen Macklin. Comparative analysis of statistical tools to identify recruitment– environment relationships and forecast recruitment strength. ICES Journal of Marine Science, 62(7):1256–1269, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/7/1256/655604.

Melvin:2002:CFV

[MLMC02] Gary Melvin, Yanchao Li, Larry Mayer, and Allan Clay. Commercial fishing vessels, automatic acoustic logging systems and 3D data visualization. ICES Journal of Marine Science, 59 (1):179–189, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/1/179/650003.

Manning:2001:MES

[MLNC01] J. P. Manning, R. G. Lough, C. E. Naimie, and J. H. Churchill. Modelling the effect of a slope-water intrusion on advection of fish larvae in May 1995 on Georges Bank. ICES Journal of Marine Science, 58(5):985–993, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/5/985/730876.

Moustahfid:2009:AEI

[MLOT09] H. Moustahfid, J. S. Link, W. J. Overholtz, and M. C. Tyrrell. The advantage of explicitly incorporating predation mortality into age-structured stock assessment models: an application for Atlantic mackerel. ICES Journal of Marine Science, 66 (3):445–454, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/3/445/812236.

Mqoqi:2007:ESA

[MLS07] Mandisile Mqoqi, Marek R. Lipiński, and Anne G. V. Salvanes. The ecology of Sepia australis (Cephalopoda: Sepiidae) along the south coast of South Africa. ICES Journal of Marine Science, 64(5):945–955, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/64/5/945/641427.

McGarvey:2001:INH

[MM01] R. McGarvey and J. M. Matthews. Incorporating numbers harvested in dynamic estimation of yearly recruitment: onshore wind in interannual variation of South Australian rock lobster (Jasus edwardsii). ICES Journal of Marine Science, 58 (5):1092–1099, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/5/1092/730838.

Megrey:2002:VSD

[MM02] Bernard A. Megrey and Erlend Moksness. Visualization of spatial data introduction. ICES Journal of Marine Science, 59(1):150, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/1/150/649987.

Massuti:2003:DADb

[MM03a] Enric Massutí and Joan Moranta. Demersal assemblages and depth distribution of elasmobrachs from the continental shelf and slope off the Balearic Islands (western Mediterranean). ICES Journal of Marine Science, 60(6):1398, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/6/1398/654962.

Massuti:2003:DADa

[MM03b] Enric Massutí and Joan Moranta. Demersal assemblages and depth distribution of elasmobranchs from the continental shelf and slope off the Balearic Islands (western Mediterranean). ICES Journal of Marine Science, 60(4):753–766, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/4/753/691866.

Mueter:2005:DPB

[MM05] Franz J. Mueter and Bernard A. Megrey. Distribution of population-based indicators across multiple taxa to assess the status of Gulf of Alaska and Bering Sea groundfish communities. ICES Journal of Marine Science, 62(3):344–352, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/3/344/659396.

Malik:2007:ISF

[MM07] Mashkoor A. Malik and Larry A. Mayer. Investigation of seabed fishing impacts on benthic structure using multi-beam sonar, sidescan sonar, and video. ICES Journal of Marine Science, 64 (5):1053–1065, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/5/1053/640915.

Merino:2009:BMT

[MMB09] Gorka Merino, Francesc Maynou, and Jean Boncoeur. Bioeconomic model for a three-zone Marine Protected Area: a case study of Medes Islands (northwest Mediterranean). ICES Journal of Marine Science, 66(1):147–154, January 2009. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/1/147/634750.

McClatchie:2003:RUL

[MMC03] S. McClatchie, G. J. Macaulay, and R. F. Coombs. A requiem for the use of 20 log_{10} length for acoustic target strength with special reference to deep-sea fishes. ICES Journal of Marine Science, 60(2):419–428, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/2/419/627863.

Mayfield:2008:ICR

[MMCD08] Stephen Mayfield, Richard McGarvey, Ian J. Carlson, and Cameron Dixon. Integrating commercial and research surveys to estimate the harvestable biomass, and establish a quota, for an “unexploited” abalone population. ICES Journal of Marine Science, 65(7):1122–1130, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/7/1122/645485.

McConnaughey:2000:ECT

[MMD00] R. A. McConnaughey, K. L. Mier, and C. B. Dew. An examination of chronic trawling effects on soft-bottom benthos of the eastern Bering Sea. ICES Journal of Marine Science, 57(5): 1377–1388, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/5/1377/660937.

McGarvey:2009:EMT

[MMF09] Richard McGarvey, Janet M. Matthews, and John E. Feenstra. Estimating mortality from times-at-large: testing accuracy and precision using simulated single tag-recovery data. ICES Journal of Marine Science, 66(3):573–581, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/3/573/811925.

Mollmann:2008:ECO

[MMKKJ08] Christian Möllmann, Bärbel Müller-Karulis, Georgs Kornilovs, and Michael A. St John. Effects of climate and overfishing on zooplankton dynamics and ecosystem structure: regime shifts, trophic cascade, and feedback loops in a simple ecosystem. ICES Journal of Marine Science, 65(3):302–310, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/3/302/784343.

McManus:2000:CRF

[MMKR^{+}00] John W. McManus, Lambert A. B. Meñez, Kathleen N. Kesner-Reyes, Sheila G. Vergara, and M. C. Ablan. Coral reef fishing and coral-algal phase shifts: implications for global reef status. ICES Journal of Marine Science, 57(3):572–578, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/3/572/635934.

Mazzola:2000:FFE

[MML^{+}00] A. Mazzola, S. Mirto, T. La Rosa, M. Fabiano, and R. Danovaro. Fish-farming effects on benthic community structure in coastal sediments: analysis of meiofaunal recovery. ICES Journal of Marine Science, 57(5):1454–1461, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/5/1454/660956.

Menendez:2009:FSI

[MML09] Melisa Menendez, Fernando J. Mendez, and Inigo J. Losada. Forecasting seasonal to interannual variability in extreme sea levels. ICES Journal of Marine Science, 66(7):1490–1496, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/7/1490/658258.

McClelland:2000:STD

[MMM00] G. McClelland, R. K. Misra, and D. J. Martell. Spatial and temporal distributions of larval sealworm (Pseudoterranova decipiens, Nematoda: Anisakinae), in Hippoglossoides platessoides (Pleuronectidae) in eastern Canada from 1980 to 1990. ICES Journal of Marine Science, 57(1):69–88, February 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/57/1/69/
641149.

Morita:2001:GCS

[MMS01] Shoko H. Morita, Kentaro Morita, and Hiroyuki Sakano. Growth of chum salmon (Oncorhynchus keta) correlated with sea-surface salinity in the North Pacific. ICES Journal of Marine Science, 58(6):1335–1339, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1335/641588.

Marin:2008:DMI

[MMV^{+}08] Valentina Marin, Mariapaola Moreno, Paolo Vassallo, Luigi Vezzulli, and Mauro Fabiano. Development of a multistep indicatorbased approach (MIBA) for the assessment of environmental quality of harbours. ICES Journal of Marine Science, 65(8): 1436–1441, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/8/1436/713109.

Morales-Nin:2002:IAR

[MN02] B. Morales-Nin. Intercalibration of age readings of deepwater black scabbardfish, Aphanopus carbo (Lowe, 1839). ICES Journal of Marine Science, 59(2):352–364, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/2/352/619648.

Mullers:2009:ILR

[MNCU09] Ralf H. E. Mullers, René A. Navarro, Robert J. M. Crawford, and Les G. Underhill. The importance of lipid-rich fish prey for Cape gannet chick growth: are fishery discards an alternative? ICES Journal of Marine Science, 66(10):2244–2252, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/10/2244/682935.

Motta:2007:RBB

[MNGB07] Fabio S. Motta, Rafael C. Namora, Otto B. F. Gadig, and F. M. S. Braga. Reproductive biology of the Brazilian sharpnose shark (Rhizoprionodon lalandii) from southeastern Brazil. ICES Journal of Marine Science, 64(9):1829–1835, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/9/1829/787746.

Marchal:2001:TCF

[MNHL01] Paul Marchal, J. Rasmus Nielsen, Holger Hovgård, and Hans Lassen. Time changes in fishing power in the Danish cod fisheries of the Baltic Sea. ICES Journal of Marine Science, 58 (1):298–310, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/1/298/603533.

Morales-Nin:2005:RFM

[MNMG^{+}05] Beatriz Morales-Nin, Joan Moranta, Cristina García, María Pilar Tugores, Antoni Maria Grau, Francisco Riera, and Margalida Cerdà. The recreational fishery off Majorca Island (western Mediterranean): some implications for coastal resource management. ICES Journal of Marine Science, 62(4):727–739, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/4/727/675256.

Murase:2009:AGA

[MNY^{+}09] Hiroto Murase, Hiroshi Nagashima, Shiroh Yonezaki, Ryuichi Matsukura, and Toshihide Kitakado. Application of a generalized additive model (GAM) to reveal relationships between environmental factors and distributions of pelagic fish and krill: a case study in Sendai Bay, Japan. ICES Journal of Marine Science, 66(6):1417–1424, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/6/1417/695432.

Moll:2000:ATD

[Mol00] Andreas Moll. Assessment of three-dimensional physicalbiological ECOHAM1 simulations by quantified validation for the North Sea with ICES and ERSEM data. ICES Journal of Marine Science, 57(4):1060–1068, August 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/4/1060/647257.

Moreno:2002:ESA

[Mor02] Isabel Moreno. Effects of substrate on the artificial reef fish assemblage in Santa Eulalia Bay (Ibiza, western Mediterranean). ICES Journal of Marine Science, 59(S1):S144–S149, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S144/617911.

Morgan:2004:RBF

[Mor04] M. Joanne Morgan. The relationship between fish condition and the probability of being mature in American plaice (Hippoglossoides platessoides). ICES Journal of Marine Science, 61(1):64–70, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/1/64/702342.

Mowbray:2002:CVD

[Mow02] Frances K. Mowbray. Changes in the vertical distribution of capelin (Mallotus villosus) off Newfoundland. ICES Journal of Marine Science, 59(5):942–949, ???? 2002. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/5/942/675209.

Maguire:2008:MAA

[MPD^{+}08] Jean-Jacques Maguire, Pilar Pereda, Rafael Duarte, Helen Dobby, and Manuela Azevedo. Monkfish/ anglerfish across the world; common problems and common solutions: an introduction to papers presented at the ICES Theme Session in September 2007. ICES Journal of Marine Science, 65(7):1270–1271, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/7/1270/647519.

McLeay:2009:DMR

[MPG^{+}09] L. J. McLeay, B. Page, S. D. Goldsworthy, T. M. Ward, D. C. Paton, M. Waterman, and M. D. Murray. Demographic and morphological responses to prey depletion in a crested tern (Sterna bergii) population: can fish mortality events highlight performance indicators for fisheries management? ICES Journal of Marine Science, 66(2):237–247, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/2/237/594396.

McLaughlin:2007:CTH

[MPJ07] Emma McLaughlin, Alex Portig, and Mark P. Johnson. Can traditional harvesting methods for cockles be accommodated in a special area of conservation? ICES Journal of Marine Science, 64(2):309–317, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/2/309/2182548.

Mathieson:2008:MAI

[MPN^{+}08] Arthur C. Mathieson, Judith R. Pederson, Christopher D. Neefus, Clinton J. Dawes, and Troy L. Bray. Multiple assessments of introduced seaweeds in the Northwest Atlantic. ICES Journal of Marine Science, 65(5):730–741, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/5/730/713661.

Mello:2005:SCW

[MR05] L. G. S. Mello and G. A. Rose. Seasonal cycles in weight and condition in Atlantic cod (Gadus morhua L.) in relation to fisheries. ICES Journal of Marine Science, 62(5):1006–1015, ???? 2005.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/5/1006/855204.

Mello:2009:ADZ

[MR09] L. G. S. Mello and G. A. Rose. The acoustic dead zone: theoretical vs. empirical estimates, and its effect on density measurements of semi-demersal fish. ICES Journal of Marine Science, 66(6):1364–1369, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1364/694638.

Murawski:2001:IDV

[MRT01] S. A. Murawski, P. J. Rago, and E. A. Trippel. Impacts of demographic variation in spawning characteristics on reference points for fishery management. ICES Journal of Marine Science, 58(5):1002–1014, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/5/1002/730807.

Mikkonen:2008:LMM

[MRV^{+}08] S. Mikkonen, M. Rahikainen, J. Virtanen, R. Lehtonen, S. Kuikka, and A. Ahvonen. A linear mixed model with temporal covariance structures in modelling catch per unit effort of Baltic herring. ICES Journal of Marine Science, 65(9):1645–1654, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/9/1645/629813.

Moran:2000:EOT

[MS00] M. J. Moran and P. C. Stephenson. Effects of otter trawling on macrobenthos and management of demersal scalefish fisheries on the continental shelf of north-western Australia. ICES Journal of Marine Science, 57(3):510–516, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/3/510/635922.

Maynou:2001:IEF

[MS01a] Francesc Maynou and Francesc Sardà. Influence of environmental factors on commercial trawl catches of Nephrops norvegicus (L.). ICES Journal of Marine Science, 58(6):1318–1325, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/6/1318/641582.

Munro:2001:MLN

[MS01b] Peter T. Munro and David A. Somerton. Maximum likelihood and non-parametric methods for estimating trawl footrope selectivity. ICES Journal of Marine Science, 58(1):220–229, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/1/220/603522.

Morton:2002:DOO

[MS02] Alexandra B. Morton and Helena K. Symonds. Displacement of Orcinus orca (L.) by high amplitude sound in British Columbia, Canada. ICES Journal of Marine Science, 59(1):71– 80, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/1/71/650034.

Moltschaniwskyj:2004:SSV

[MS04] Natalie A. Moltschaniwskyj and Michael A. Steer. Spatial and seasonal variation in reproductive characteristics and spawning of southern calamary (Sepioteuthis australis): spreading the mortality risk. ICES Journal of Marine Science, 61(6):921–927, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/6/921/678004.

Melvin:2007:DRF

[MS07] Gary D. Melvin and Robert L. Stephenson. The dynamics of a recovering fish stock: Georges Bank herring. ICES Journal of Marine Science, 64(1):69–82, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/1/69/648931.

McConnaughey:2009:SRB

[MS09] Robert A. McConnaughey and Stephen E. Syrjala. Statistical relationships between the distributions of groundfish and crabs in the eastern Bering Sea and processed returns from a single-beam echosounder. ICES Journal of Marine Science, 66(6):1425–1432, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1425/697625.

Mahevas:2004:QAV

[MSB04] S. Mahévas, Y. Sandon, and A. Biseau. Quantification of annual variations in fishing power due to vessel characteristics: an application to the bottom-trawlers of South-Brittany targeting anglerfish (Lophius budegassa and Lophius piscatorius). ICES Journal of Marine Science, 61(1):71–83, ???? 2004. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/61/1/71/
702447.

Maunder:2006:ICP

[MSF^{+}06] Mark N. Maunder, John R. Sibert, Alain Fonteneau, John Hampton, Pierre Kleiber, and Shelton J. Harley. Interpreting catch per unit effort data to assess the status of individual stocks and communities. ICES Journal of Marine Science, 63 (8):1373–1385, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/8/1373/710477.

Munoz:2009:SMS

[MSGC^{+}09] P. Durán Muñoz, M. Sayago-Gil, J. Cristobo, S. Parra, A. Serrano, V. Díaz del Rio, T. Patrocinio, M. Sacau, F. J. Murillo, D. Palomino, and L. M. Fernández-Salas. Seabed mapping for selecting cold-water coral protection areas on Hatton Bank, Northeast Atlantic. ICES Journal of Marine Science, 66(9): 2013–2025, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/9/2013/725884.

McAuley:2007:MEI

[MSH07] Rory B. McAuley, Colin A. Simpfendorfer, and Norm G. Hall. A method for evaluating the impacts of fishing mortality and stochastic influences on the demography of two long-lived shark stocks. ICES Journal of Marine Science, 64(9):1710–1722, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/9/1710/781282.

Miller:2007:OSS

[MSI07] Timothy J. Miller, John R. Skalski, and James N. Ianelli. Optimizing a stratified sampling design when faced with multiple objectives. ICES Journal of Marine Science, 64(1):97–109, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/1/97/648215.

Mustapha:2009:SMS

[MSIL09] M. A. Mustapha, S. Sei-Ichi, and T. Lihan. Satellite-measured seasonal variations in primary production in the scallop-farming region of the Okhotsk Sea. ICES Journal of Marine Science, 66 (7):1557–1569, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/7/1557/658895.

Morita:2006:RPS

[MSM^{+}06] Kentaro Morita, Toshihiko Saito, Yasuyuki Miyakoshi, Masa aki Fukuwaka, Toru Nagasawa, and Masahide Kaeriyama. A review of Pacific salmon hatchery programmes on Hokkaido Island, Japan. ICES Journal of Marine Science, 63(7):1353–1363, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/7/1353/760821.

Melvin:2009:ORS

[MSP09] Gary D. Melvin, Robert L. Stephenson, and Michael J. Power. Oscillating reproductive strategies of herring in the western Atlantic in response to changing environmental conditions. ICES Journal of Marine Science, 66(8):1784–1792, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/8/1784/675829.

Methven:2003:SPP

[MSR03] David A. Methven, D. C. Schneider, and George A. Rose. Spatial pattern and patchiness during ontogeny: post-settled Gadus morhua from coastal Newfoundland. ICES Journal of Marine Science, 60(1):38–51, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/60/1/38/611444.

McQuinn:2005:AIA

[MSS^{+}05] Ian H. McQuinn, Yvan Simard, Thomas W. F. Stroud, Jean-Louis Beaulieu, and Stephen J. Walsh. An adaptive, integrated “acoustic-trawl” survey design for Atlantic cod (Gadus morhua) with estimation of the acoustic and trawl dead zones. ICES Journal of Marine Science, 62(1):93–106, ???? 2005. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/62/1/93/
670207.

McAuley:2007:GMS

[MSW07] R. B. McAuley, C. A. Simpfendorfer, and I. W. Wright. Gillnet mesh selectivity of the sandbar shark (Carcharhinus plumbeus): implications for fisheries management. ICES Journal of Marine Science, 64(9):1702–1709, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/9/1702/778696.

Mills:2007:EHR

[MTJ^{+}07] Craig M. Mills, Sunny E. Townsend, Simon Jennings, Paul D. Eastwood, and Carla A. Houghton. Estimating high resolution trawl fishing effort from satellite-based vessel monitoring system data. ICES Journal of Marine Science, 64(2):248–255, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/2/248/2182309.

Muino:2003:CCS

[Mui03] R. Muiño. Consistency in the correlation of school parameters across years and stocks. ICES Journal of Marine Science, 60 (1):164–175, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/1/164/611434.

Marchal:2002:CTI

[MUK^{+}02] Paul Marchal, Clara Ulrich, Knut Korsbrekke, Martin Pastoors, and Brian Rackham. A comparison of three indices of fishing power on some demersal fisheries of the North Sea. ICES Journal of Marine Science, 59(3):604–623, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/3/604/610844.

Murawski:2000:DOE

[Mur00a] Steven A. Murawski. Definitions of overfishing from an ecosystem perspective. ICES Journal of Marine Science, 57(3): 649–658, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/649/635951.

Murta:2000:MVH

[Mur00b] Alberto G. Murta. Morphological variation of horse mackerel (Trachurus trachurus) in the Iberian and North African Atlantic: implications for stock identification. ICES Journal of Marine Science, 57(4):1240–1248, August 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/4/1240/647314.

Maes:2007:CGR

[MV07] Gregory E. Maes and Filip A. M. Volckaert. Challenges for genetic research in European eel management. ICES Journal of Marine Science, 64(7):1463–1471, October 2007. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/7/1463/727349.

Murta:2009:UEA

[MV09] Alberto G. Murta and Catarina Vendrell. Using the EM algorithm to age fish eggs. ICES Journal of Marine Science, 66 (4):607–616, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/4/607/602821.

Mackinson:2005:FRB

[MvdKN05] Steven Mackinson, Jeroen van der Kooij, and Suzanna Neville. The fuzzy relationship between trawl and acoustic surveys in the North Sea. ICES Journal of Marine Science, 62(8):1556–1575, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/8/1556/789331.

Moreno:2008:UMD

[MVM^{+}08] Mariapaola Moreno, Luigi Vezzulli, Valentina Marin, Paola Laconi, Giancarlo Albertelli, and Mauro Fabiano. The use of meiofauna diversity as an indicator of pollution in harbours. ICES Journal of Marine Science, 65(8):1428–1435, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/8/1428/712902.

Marrari:2004:SPM

[MVMH04] Marina Marrari, María Delia Viñas, Patricia Martos, and Daniel Hernández. Spatial patterns of mesozooplankton distribution in the Southwestern Atlantic Ocean (34^{o}–41^{o}S) during austral spring: relationship with the hydrographic conditions. ICES Journal of Marine Science, 61(4):667–679, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/4/667/606888.

McQuinn:2003:TAT

[MW03] Ian H. McQuinn and Paul D. Winger. Tilt angle and target strength: target tracking of Atlantic cod (Gadus morhua) during trawling. ICES Journal of Marine Science, 60(3):575–583, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/3/575/659462.

Murawski:2005:EDC

[MWF^{+}05] Steven A. Murawski, Susan E. Wigley, Michael J. Fogarty, Paul J. Rago, and David G. Mountain. Effort distribution and catch patterns adjacent to temperate MPAs. ICES Journal of Marine Science, 62(6):1150–1167, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/6/1150/617874.

Muxagata:2004:CTD

[MWS04] Erik Muxagata, John A. Williams, and Martin Sheader. Composition and temporal distribution of cirripede larvae in Southampton Water, England, with particular reference to the secondary production of Elminius modestus. ICES Journal of Marine Science, 61(4):585–595, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/61/4/585/604935.

Minegishi:2009:SIA

[MYAT09] Yuki Minegishi, Tatsuki Yoshinaga, Jun Aoyama, and Katsumi Tsukamoto. Species identification of Anguilla japonica by realtime PCR based on a sequence detection system: a practical application to eggs and larvae. ICES Journal of Marine Science, 66(9):1915–1918, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/9/1915/722792.

Myers:2001:SRG

[Mye01] Ransom A. Myers. Stock and recruitment: generalizations about maximum reproductive rate, density dependence, and variability using meta-analytic approaches. ICES Journal of Marine Science, 58(5):937–951, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/58/5/937/730862.

Niamaimandi:2008:RBG

[NAK^{+}08] Nassir Niamaimandi, Arshad Aziz, Daud Siti Khalijah, Saed Che Roos, and Bahram Kiabi. Reproductive biology of the green tiger prawn (Penaeus semisulcatus) in coastal waters of Bushehr, Persian Gulf. ICES Journal of Marine Science, 65(9): 1593–1599, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/9/1593/632042.

Naumenko:2002:DPC

[Nau02] E. A. Naumenko. The dynamics of prespawning capelin (Mallotus villosus socialis) off the West Kamchatka coast. ICES Journal of Marine Science, 59(5):1006–1010, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/5/1006/675125.

Nelson:2008:UPB

[NB08] Walter G. Nelson and Cheryl A. Brown. Use of probabilitybased sampling of water-quality indicators in supporting development of quality criteria. ICES Journal of Marine Science, 65 (8):1421–1427, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/8/1421/714720.

Nichol:2006:RBP

[NC06] D. G. Nichol and E. A. Chilton. Recuperation and behaviour of Pacific cod after barotrauma. ICES Journal of Marine Science, 63(1):83–94, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/1/83/625743.

Noblezada:2008:SDC

[NC08] Mary Mar P. Noblezada and Wilfredo L. Campos. Spatial distribution of chaetognaths off the northern Bicol Shelf, Philippines (Pacific coast). ICES Journal of Marine Science, 65(3):484–494, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/3/484/786729.

Nakhle:2007:CMS

[NCC^{+}07] K. Nakhlé, D. Cossa, D. Claisse, B. Beliaeff, and S. Simon. Cadmium and mercury in Seine Estuary flounders and mussels: the results of two decades of monitoring. ICES Journal of Marine Science, 64(5):929–938, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/5/929/640482.

Neilson:2003:DVD

[NCM^{+}03] John D. Neilson, Donald Clark, Gary D. Melvin, Peter Perley, and Chris Stevens. The diel-vertical distribution and characteristics of pre-spawning aggregations of pollock (Pollachius virens) as inferred from hydroacoustic observations: the implications for survey design. ICES Journal of Marine Science, 60(4):860–871, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/4/860/694044.

Niemela:2005:JSA

[NEJH05] E. Niemelä, J. Erkinaro, M. Julkunen, and E. Hassinen. Is juvenile salmon abundance related to subsequent and preceding catches? Perspectives from a long-term monitoring programme. ICES Journal of Marine Science, 62(8):1617–1629, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/8/1617/791507.

Nottestad:2002:HWI

[NFM^{+}02] Leif Nøttestad, Anders Fernö, Steve Mackinson, Tony Pitcher, and Ole Arve Misund. How whales influence herring school dynamics in a cold-front area of the Norwegian Sea. ICES Journal of Marine Science, 59(2):393–400, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/2/393/619662.

Nogueira:2004:CBS

[NGNB^{+}04] Enrique Nogueira, Gonzalo González-Nuevo, Antonio Bode, Manuel Varela, Xosé Anxelu G. Morán, and Luis Valdés. Comparison of biomass and size spectra derived from optical plankton counter data and net samples: application to the assessment of mesoplankton distribution along the northwest and North Iberian shelf. ICES Journal of Marine Science, 61(4):508–517, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/4/508/603472.

Nakamura:2009:SDV

[NH09] Takeshi Nakamura and Akira Hamano. Seasonal differences in the vertical distribution pattern of Japanese jack mackerel, Trachurus japonicus: changes according to age? ICES Journal of Marine Science, 66(6):1289–1295, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1289/695884.

Nesse:2009:MAS

[NHK09] Tonje Lexau Nesse, Halvor Hobæk, and Rolf J. Korneliussen. Measurements of acoustic-scattering spectra from the whole and parts of Atlantic mackerel. ICES Journal of Marine Science, 66 (6):1169–1175, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1169/693881.

Ndjaula:2009:ODC

[NHKJK09] Hilkka O. N. Ndjaula, Tom Hansen, Maria Krüger-Johnsen, and Olav Sigurd Kjesbu. Oocyte development in captive Atlantic horse mackerel Trachurus trachurus. ICES Journal of Marine Science, 66(4):623–630, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/4/623/603443.

Niehoff:2000:ESR

[Nie00] B. Niehoff. Effect of starvation on the reproductive potential of Calanus finmarchicus. ICES Journal of Marine Science, 57(6): 1764–1772, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/6/1764/768652.

Nishimori:2009:DET

[NIF^{+}09] Yasushi Nishimori, Kohji Iida, Masahiko Furusawa, Yong Tang, Kozo Tokuyama, Sanae Nagai, and Yoshihiro Nishiyama.

The development and evaluation of a three-dimensional, echointegration method for estimating fish-school abundance. ICES Journal of Marine Science, 66(6):1037–1042, July 2009. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/1037/691588.

Niwa:2007:RWD

[Niw07] Hiro-Sato Niwa. Random-walk dynamics of exploited fish populations. ICES Journal of Marine Science, 64(3):496–502, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/3/496/816032.

Nicholson:2004:TCI

[NJ04] Mike D. Nicholson and Simon Jennings. Testing candidate indicators to support ecosystem-based management: the power of monitoring surveys to detect temporal trends in fish community metrics. ICES Journal of Marine Science, 61(1):35–42, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/1/35/701738.

Neuenfeldt:2000:TCR

[NK00] Stefan Neuenfeldt and Friedrich W. Köster. Trophodynamic control on recruitment success in Baltic cod: the influence of cannibalism. ICES Journal of Marine Science, 57(2):300–309, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/2/300/620435.

Nakata:2000:IMS

[NKOK00] Hideaki Nakata, Shingo Kimura, Yuji Okazaki, and Akihide Kasai. Implications of meso-scale eddies caused by frontal disturbances of the Kuroshio Current for anchovy recruitment. ICES Journal of Marine Science, 57(1):143–152, February 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/1/143/641131.

Nunes:2008:EVD

[NM08] Paulo A. L. D. Nunes and Anil Markandya. Economic value of damage caused by marine bio-invasions: lessons from two European case studies. ICES Journal of Marine Science, 65 (5):775–780, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/5/775/714681.

North:2009:MSG

[NM09] Elizabeth W. North and Franz J. Mueter. Marine science with global vision: creating a place for early career scientists. An introduction to selected articles from the 2007 Early Career Scientists Conference. ICES Journal of Marine Science, 66 (2):334–335, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/2/334/595758.

Nishibori:2001:DFP

[NNT01] Naoyoshi Nishibori, Akinori Nishii, and Haruyoshi Takayama. Detection of free polyamine in coastal seawater using ion exchange chromatography. ICES Journal of Marine Science, 58 (6):1201–1207, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1201/641554.

Norro:2006:BDA

[NPPO06] Alain Norro, Georges Pichot, Virginie Pison, and José Ozer. A bi-dimensional approach to assessing the volumetric evolution of an exploited sandbank. ICES Journal of Marine Science, 63 (1):176–186, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/1/176/624596.

Neumann:2009:TVS

[NRR^{+}09] Hermann Neumann, Henning Reiss, Sebastian Rakers, Siegfried Ehrich, and Ingrid Kröncke. Temporal variability in southern North Sea epifauna communities after the cold winter of 1995/ 1996. ICES Journal of Marine Science, 66(10):2233–2243, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/10/2233/682161.

Noji:2009:SBI

[NRS09] Thomas Noji, Heye Rumohr, and Stephen J. Smith. Sediment– biota interactions and mapping marine habitats: an introduction. ICES Journal of Marine Science, 66(9):2012, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/9/2012/727265.

Neilson:2006:PPV

[NSP06] John D. Neilson, Wayne T. Stobo, and Peter Perley. Pollock (Pollachius virens) stock structure in the Canadian maritimes inferred from mark-recapture studies. ICES Journal of Marine Science, 63(4):749–765, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/63/4/749/694518.

Nakashima:2002:BSS

[NT02] Brian S. Nakashima and Christopher T. Taggart. Is beachspawning success for capelin, Mallotus villosus (Müller), a function of the beach? ICES Journal of Marine Science, 59(5): 897–908, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/5/897/675203.

Nero:2004:SAE

[NTJ04] Redwood W. Nero, Charles H. Thompson, and J. Michael Jech. In situ acoustic estimates of the swimbladder volume of Atlantic herring (Clupea harengus). ICES Journal of Marine Science, 61 (3):323–337, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/3/323/671089.

Nielsen:2007:SDJ

[NTSM07] Julie K. Nielsen, S. James Taggart, Thomas C. Shirley, and Jennifer Mondragon. Spatial distribution of juvenile and adult female Tanner crabs (Chionoecetes bairdi) in a glacial fjord ecosystem: implications for recruitment processes. ICES Journal of Marine Science, 64(9):1772–1784, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/9/1772/787151.

Nakashima:2002:CMV

[NW02] Brian S. Nakashima and John P. Wheeler. Capelin (Mallotus villosus) spawning behaviour in Newfoundland waters–the interaction between beach and demersal spawning. ICES Journal of Marine Science, 59(5):909–916, ???? 2002. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/5/909/675205.

Nissling:2002:RSR

[NWH02] Anders Nissling, Lars Westin, and Olle Hjerne. Reproductive success in relation to salinity for three flatfish species, dab (Limanda limanda), plaice (Pleuronectes platessa), and flounder (Pleuronectes flesus), in the brackish water Baltic Sea. ICES Journal of Marine Science, 59(1):93–108, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/59/1/93/
650039.

Otteraa:2006:DST

[OAJ06] Håkon Otterå, Ann-Lisbeth Agnalt, and Knut E. Jørstad. Differences in spawning time of captive Atlantic cod from four regions of Norway, kept under identical conditions. ICES Journal of Marine Science, 63(2):216–223, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/2/216/637794.

Orlov:2005:FRS

[OB05] Alexei M. Orlov and Igor A. Biryukov. First report of sablefish in spawning condition off the coast of Kamchatka and the Kuril Islands. ICES Journal of Marine Science, 62(5):1016–1020, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/5/1016/855461.

Orlova:2005:RBP

[OBD^{+}05] E. L. Orlova, V. D. Boitsov, A. V. Dolgov, G. B. Rudneva, and V. N. Nesterova. The relationship between plankton, capelin, and cod under different temperature conditions. ICES Journal of Marine Science, 62(7):1281–1292, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/7/1281/656411.

Orlova:2002:CDC

[OBNU02] Emma L. Orlova, Vladimir D. Boitsov, Valentina N. Nesterova, and Nikolai G. Ushakov. Composition and distribution of copepods, a major prey of capelin in the central Barents Sea, in moderate and warm years. ICES Journal of Marine Science, 59 (5):1053–1061, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/5/1053/675145.

OReilly:2006:DEA

[OCWV06] Patrick T. O’Reilly, Jonathan W. Carr, Frederick G. Whoriskey, and Eric Verspoor. Detection of European ancestry in escaped farmed Atlantic salmon, Salmo salar L., in the Magaguadavic River and Chamcook Stream, New Brunswick, Canada. ICES Journal of Marine Science, 63(7):1256–1262, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/7/1256/756724.

ODriscoll:2003:DSC

[O’D03] Richard L. O’Driscoll. Determining species composition in mixed-species marks: an example from the New Zealand hoki (Macruronus novaezelandiae) fishery. ICES Journal of Marine Science, 60(3):609–616, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/60/3/609/659931.

ODriscoll:2004:EUA

[O’D04] Richard L. O’Driscoll. Estimating uncertainty associated with acoustic surveys of spawning hoki (Macruronus novaezelandiae) in Cook Strait, New Zealand. ICES Journal of Marine Science, 61(1):84–97, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/1/84/702720.

Oeberst:2009:MDG

[ODCN09] Rainer Oeberst, Mark Dickey-Collas, and Richard D. M. Nash. Mean daily growth of herring larvae in relation to temperature over a range of 5–20^{o}C, based on weekly repeated cruises in the Greifswalder Bodden. ICES Journal of Marine Science, 66(8): 1696–1701, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/8/1696/677425.

Orlova:2005:EAB

[ODRN05] E. L. Orlova, A. V. Dolgov, G. B. Rudneva, and V. N. Nesterova. The effect of abiotic and biotic factors on the importance of macroplankton in the diet of Northeast Arctic cod in recent years. ICES Journal of Marine Science, 62(7):1463–1474, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/7/1463/660234.

Ostrowski:2009:RSS

[OdSBS09] Marek Ostrowski, José C. B. da Silva, and Bomba Bazik-Sangolay. The response of sound scatterers to El Niño- and La Niña-like oceanographic regimes in the southeastern Atlantic. ICES Journal of Marine Science, 66(6):1063–1072, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/1063/695091.

Ohman:2004:CSC

[OED^{+}04] M. D. Ohman, K. Eiane, E. G. Durbin, J. A. Runge, and H.-J. Hirche. A comparative study of Calanus finmarchicus mortality patterns at five localities in the North Atlantic. ICES Journal of Marine Science, 61(4):687–697, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/4/687/607875.

Otterlei:2002:E

[OFN02a] E. Otterlei, A. Folkvord, and G. Nyhammer. Erratum. ICES Journal of Marine Science, 59(4):850, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/4/850/676824.

Otterlei:2002:TDOa

[OFN02b] E. Otterlei, A. Folkvord, and G. Nyhammer. Temperature dependent otolith growth of larval and early juvenile Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 59 (2):401–410, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/2/401/619663.

Otterlei:2002:TDOb

[OFN02c] E. Otterlei, A. Folkvord, and G. Nyhammer. Temperature dependent otolith growth of larval and early juvenile Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 59 (4):851–860, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/4/851/676827.

Ozbilgin:2006:SVT

[ÖFR^{+}06] H. Özbilgin, R. S. T. Ferro, J. H. B. Robertson, G. Holtrop, and R. J. Kynoch. Seasonal variation in trawl codend selection of northern North Sea haddock. ICES Journal of Marine Science, 63(4):737–748, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/4/737/694351.

Ozbilgin:2004:RLM

[ÖG04] H. Özbilgin and C. W. Glass. Role of learning in mesh penetration behaviour of haddock (Melanogrammus aeglefinus). ICES Journal of Marine Science, 61(7):1190–1194, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1190/881893.

ODriscoll:2009:AEM

[OGD09] Richard L. O’Driscoll, Stéphane Gauthier, and Jennifer A. Devine. Acoustic estimates of mesopelagic fish: as clear as day and night? ICES Journal of Marine Science, 66(6):1310–1317, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1310/690469.

Orlov:2006:FBC

[OGL06] Alexander V. Orlov, Yuri V. Gerasimov, and Oleg M. Lapshin. The feeding behaviour of cultured and wild Atlantic salmon, Salmo salar L., in the Louvenga River, Kola Peninsula, Russia. ICES Journal of Marine Science, 63(7):1297–1303, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/7/1297/757758.

Olsen:2007:NEB

[OGR^{+}07] Erik Olsen, Harald Gjøsæter, Ingolf Røttingen, Are Dommasnes, Petter Fossum, and Per Sandberg. The Norwegian ecosystembased management plan for the Barents Sea. ICES Journal of Marine Science, 64(4):599–602, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/599/639604.

Oskarsson:2009:VSD

[ÓGS09] Gudmundur J. Óskarsson, A. Gudmundsdottir, and T. Sigurdsson. Variation in spatial distribution and migration of Icelandic summer-spawning herring. ICES Journal of Marine Science, 66 (8):1762–1767, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/8/1762/674571.

ODor:2000:DGL

[OH00] R. K. O’Dor and J. A. Hoar. Does geometry limit squid growth? ICES Journal of Marine Science, 57(1):8–14, February 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/1/8/641152.

ONeill:2007:PPM

[OH07] F. G. O’Neill and B. Herrmann. PRESEMO — a predictive model of codend selectivity — a tool for fishery managers. ICES Journal of Marine Science, 64(8):1558–1568, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/8/1558/613090.

Ojaveer:2005:ICC

[OK05] Evald Ojaveer and Margers Kalejs. The impact of climate change on the adaptation of marine fish in the Baltic Sea. ICES Journal of Marine Science, 62(7):1492–1500, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/7/1492/661024.

Oeberst:2009:WYC

[OKG^{+}09] Rainer Oeberst, Birgitt Klenz, Tomas Gröhsler, Mark Dickey-Collas, Richard D. M. Nash, and Christopher Zimmermann. When is year-class strength determined in western Baltic herring? ICES Journal of Marine Science, 66(8):1667– 1672, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/8/1667/675141.

Onsrud:2004:VDF

[OKRK04] M. S. R. Onsrud, S. Kaartvedt, A. Røstad, and T. A. Klevjer. Vertical distribution and feeding patterns in fish foraging on the krill Meganyctiphanes norvegica. ICES Journal of Marine Science, 61(8):1278–1290, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/8/1278/629825.

Ottersen:2000:CEG

[OL00] G. Ottersen and H. Loeng. Covariability in early growth and year-class strength of Barents Sea cod, haddock, and herring: the environmental link. ICES Journal of Marine Science, 57 (2):339–348, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/2/339/620446.

Overholtz:2007:CIM

[OL07] W. J. Overholtz and J. S. Link. Consumption impacts by marine mammals, fish, and seabirds on the Gulf of Maine–Georges Bank Atlantic herring (Clupea harengus) complex during the years 1977–2002. ICES Journal of Marine Science, 64(1):83–96, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/1/83/647386.

Ouellet:2001:CEC

[OLB01] Patrick Ouellet, Yvan Lambert, and Isabelle Bérubé. Cod egg characteristics and viability in relation to low temperature and maternal nutritional condition. ICES Journal of Marine Science, 58(3):672–686, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/58/3/672/810120.

Overholtz:2000:CIP

[OLS00] W. J. Overholtz, J. S. Link, and L. E. Suslowicz. Consumption of important pelagic fish and squid by predatory fish in the northeastern USA shelf ecosystem with some fishery comparisons. ICES Journal of Marine Science, 57(4):1147–1159, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/1147/647288.

ODriscoll:2005:UFP

[OM05] Richard L. O’Driscoll and Gavin J. Macaulay. Using fishprocessing time to carry out acoustic surveys from commercial vessels. ICES Journal of Marine Science, 62(2):295–305, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/2/295/604249.

Ona:2009:CMT

[OMA09] Egil Ona, Valerie Mazauric, and Lars Nonboe Andersen. Calibration methods for two scientific multibeam systems. ICES Journal of Marine Science, 66(6):1326–1334, July 2009. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/1326/696447.

OBrien-MacDonald:2006:GBD

[OMBP06] Kelly O’Brien-MacDonald, Joseph A. Brown, and Christopher C. Parrish. Growth, behaviour, and digestive enzyme activity in larval Atlantic cod (Gadus morhua) in relation to rotifer lipid. ICES Journal of Marine Science, 63(2):275–284, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/2/275/639890.

OMaoileidigh:2004:APF

[ÓMP^{+}04] N. Ó Maoiléidigh, P. McGinnity, E. Prévost, E. C. E. Potter, P. Gargan, W. W. Crozier, P. Mills, and W. Roche. Application of pre-fishery abundance modelling and Bayesian hierarchical stock and recruitment analysis to the provision of precautionary catch advice for Irish salmon (Salmo salar L.) fisheries. ICES Journal of Marine Science, 61(8):1370–1378, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/8/1370/632017.

Okumura:2003:ASA

[OMTS03] Tsuyoshi Okumura, Toshikazu Masuya, Yoshimi Takao, and Kouichi Sawada. Acoustic scattering by an arbitrarily shaped body: an application of the boundary-element method. ICES Journal of Marine Science, 60(3):563–570, ???? 2003. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/3/563/659318.

Ormseth:2009:CCL

[ON09] Olav A. Ormseth and Brenda L. Norcross. Causes and consequences of life-history variation in North American stocks of Pacific cod. ICES Journal of Marine Science, 66(2):349–357, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/2/349/593173.

Ona:2003:ETS

[Ona03] Egil Ona. An expanded target-strength relationship for herring. ICES Journal of Marine Science, 60(3):493–499, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/3/493/658268.

ODriscoll:2001:SAT

[OR01] Richard L. O’Driscoll and George A. Rose. In situ acoustic target strength of juvenile capelin. ICES Journal of Marine Science, 58(1):342–345, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/58/1/342/603542.

Oosthuizen:2009:BTS

[OR09] Anè Oosthuizen and Mike J. Roberts. Bottom temperature and in situ development of chokka squid eggs (Loligo vulgaris reynaudii) on mid-shelf spawning grounds, South Africa. ICES Journal of Marine Science, 66(9):1967–1971, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/9/1967/726852.

ODriscoll:2002:CCC

[ORA02] R. L. O’Driscoll, G. A. Rose, and J. T. Anderson. Counting capelin: a comparison of acoustic density and trawl catchability. ICES Journal of Marine Science, 59(5):1062–1071, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/5/1062/675150.

Orlowski:2000:DDA

[Orł00] A. Orłowski. Diel dynamics of acoustic measurements of Baltic fish. ICES Journal of Marine Science, 57(4):1196–1203, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/1196/647300.

Orlowski:2001:BPE

[Orl01] A. Orlowski. Behavioural and physical effect on acoustic measurements of Baltic fish within a diel cycle. ICES Journal of Marine Science, 58(6):1174–1183, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/6/1174/641541.

Orlowski:2003:AST

[Orl03] Andrzej Orlowski. Acoustic semi-tomography in studies of the structure and function of the marine ecosystem. ICES Journal of Marine Science, 60(6):1392–1397, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/6/1392/654880.

Orlowski:2005:EVA

[Orl05] Andrzej Orlowski. Experimental verification of the acoustic characteristics of the clupeoid diel cycle in the Baltic. ICES Journal of Marine Science, 62(6):1180–1190, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/6/1180/618344.

Otero:2009:RPF

[ORVP09] P. Otero, M. Ruiz-Villarreal, and Á. Peliz. River plume fronts off NW Iberia from satellite observations and model data. ICES Journal of Marine Science, 66(9):1853–1864, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/9/1853/722383.

Olyott:2006:STP

[OSB06] L. J. H. Olyott, W. H. H. Sauer, and A. J. Booth. Spatiotemporal patterns in maturation of the chokka squid (Loligo vulgaris reynaudii) off the coast of South Africa. ICES Journal of Marine Science, 63(9):1649–1664, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/9/1649/698877.

OBoyle:2005:IEB

[OSK^{+}05] R. O’Boyle, M. Sinclair, P. Keizer, K. Lee, D. Ricard, and P. Yeats. Indicators for ecosystem-based management on the Scotian Shelf: bridging the gap between theory and practice.

ICES Journal of Marine Science, 62(3):598–605, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/3/598/666896.

Opstad:2006:GSD

[OSLO06] I. Opstad, J. Suontama, E. Langmyhr, and R. E. Olsen. Growth, survival, and development of Atlantic cod (Gadus morhua L.) weaned onto diets containing various sources of marine protein. ICES Journal of Marine Science, 63(2):320–325, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/2/320/640643.

Ostrovsky:2009:AQF

[Ost09] Ilia Ostrovsky. The acoustic quantification of fish in the presence of methane bubbles in the stratified Lake Kinneret, Israel. ICES Journal of Marine Science, 66(6):1043–1047, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/1043/695289.

Osenberg:2002:QFE

[OSWL02] Craig W. Osenberg, Colette M. St. Mary, Jacqueline A. Wilson, and William J. Lindberg. A quantitative framework to evaluate the attraction–production controversy. ICES Journal of Marine Science, 59(S1):S214–S221, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S214/617942.

Oskarsson:2006:FVI

[ÓT06] Gudmundur J. Óskarsson and Christopher T. Taggart. Fecundity variation in Icelandic summer-spawning herring and implications for reproductive potential. ICES Journal of Marine Science, 63(3):493–503, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/63/3/493/721638.

Ozbilgin:2007:SVT

[ÖTTM07] Hüseyin Özbilgin, Zafer Tosunoğlu, Adnan Tokaç, and Gülnur Metin. Seasonal variation in the trawl codend selectivity of picarel (Spicare smaris). ICES Journal of Marine Science, 64(8):

1569–1572, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/8/1569/613741.

Orlova:2002:FSF

[OUNB02] E. L. Orlova, N. G. Ushakov, V. N. Nesterova, and V. D. Boitsov. Food supply and feeding of capelin (Mallotus villosus) of different size in the central latitudinal zone of the Barents Sea during intermediate and warm years. ICES Journal of Marine Science, 59(5):968–975, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/5/968/675217.

Oddone:2004:SMS

[OV04] María Cristina Oddone and Gonzalo Velasco. Size at maturity of the smallnose fanskate Sympterygia bonapartii (Müller & Henle, 1841) (Pisces, Elasmobranchii, Rajidae) in the SW Atlantic. ICES Journal of Marine Science, 61(2):293–296, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/2/293/622302.

Oddone:2005:RBA

[OV05] María Cristina Oddone and Carolus Maria Vooren. Reproductive biology of Atlantoraja cyclophora (Regan 1903) (Elasmobranchii: Rajidae) off southern Brazil. ICES Journal of Marine Science, 62(6):1095–1103, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/6/1095/616948.

Pascual:2006:HPI

[PAA06] S. Pascual, E. Abollo, and C. Azevedo. Host–parasite interaction of a muscle-infecting didymozoid in the Atlantic mackerel Scomber scombrus L. ICES Journal of Marine Science, 63 (1):169–175, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/1/169/624338.

Ponti:2002:DPA

[PAC02] Massimo Ponti, Marco Abbiati, and Victor Ugo Ceccherelli. Drilling platforms as artificial reefs: distribution of macrobenthic assemblages of the “paguro” wreck (northern Adriatic Sea). ICES Journal of Marine Science, 59(S1):S316–S323, ???? 2002.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S316/617980.

Perez:2005:GIH

[PÁMGV05] Juliana Perez, Paula Álvarez, Jose L. Martinez, and Eva Garcia-Vazquez. Genetic identification of hake and megrim eggs in formaldehyde-fixed plankton samples. ICES Journal of Marine Science, 62(5):908–914, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/62/5/908/856802.

Pascoe:2006:EFM

[Pas06] Sean Pascoe. Economics, fisheries, and the marine environment. ICES Journal of Marine Science, 63(1):1–3, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/1/1/622551.

Payne:2004:NFY

[Pay04] Andrew I. L. Payne. The next few years of the ICES Journal of Marine Science. ICES Journal of Marine Science, 61(1):1–2, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/1/1/699390.

Planque:2000:CFN

[PB00] B. Planque and S. D. Batten. Calanus finmarchicus in the North Atlantic: the year of Calanus in the context of interdecadal change. ICES Journal of Marine Science, 57(6):1528–1535, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/6/1528/768526.

Payne:2005:GPI

[PB05a] Andrew I. L. Payne and Christiane Barranguet. Growth and potential of the ICES Journal of Marine Science. ICES Journal of Marine Science, 62(1):1–2, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/1/1/665966.

Peltonen:2005:ATS

[PB05b] Heikki Peltonen and Helge Balk. The acoustic target strength of herring (Clupea harengus L.) in the northern Baltic Sea. ICES Journal of Marine Science, 62(4):803–808, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/4/803/676744.

Poulard:2005:ICC

[PB05c] Jean-Charles Poulard and Fabian Blanchard. The impact of climate change on the fish community structure of the eastern continental shelf of the Bay of Biscay. ICES Journal of Marine Science, 62(7):1436–1443, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/7/1436/659491.

Pederson:2008:FIC

[PB08a] Judith A. Pederson and April M. H. Blakeslee. Fifth international conference on marine bioinvasions: Introduction. ICES Journal of Marine Science, 65(5):713–715, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/5/713/715155.

Powers:2008:PRU

[PB08b] Joseph E. Powers and Elizabeth N. Brooks. Penalties and rewards for over- and underages of catch allocations. ICES Journal of Marine Science, 65(9):1541–1551, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/9/1541/629492.

Pitcher:2002:FBN

[PBH02] Tony J. Pitcher, Eny A. Buchary, and Trevor Hutton. Forecasting the benefits of no-take human-made reefs using spatial ecosystem simulation. ICES Journal of Marine Science, 59(S1):S17–S26, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S17/617918.

Peteiro:2006:GMG

[PBLFR06] Laura G. Peteiro, José M. F. Babarro, Uxío Labarta, and María José Fernández-Reiriz. Growth of Mytilus galloprovincialis after the Prestige oil spill. ICES Journal of Marine Science, 63(6):1005–1013, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/63/6/1005/614225.

Perry:2004:IGS

[PBM^{+}04] R. Ian Perry, Harold P. Batchelder, David L. Mackas, Sanae Chiba, Edward Durbin, Wulf Greve, and Hans M. Verheye. Identifying global synchronies in marine zooplankton populations: issues and opportunities. ICES Journal of Marine Science, 61 (4):445–456, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/4/445/602103.

Perez-Castaneda:2005:GMT

[PCD05] Roberto Pérez-Castañeda and Omar Defeo. Growth and mortality of transient shrimp populations (Farfantepenaeus spp.) in a coastal lagoon of Mexico: role of the environment and density-dependence. ICES Journal of Marine Science, 62(1):14– 24, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/1/14/667216.

Page:2008:OGP

[PCDM08] Henry M. Page, Carolynn S. Culver, Jenifer E. Dugan, and Brent Mardian. Oceanographic gradients and patterns in invertebrate assemblages on offshore oil platforms. ICES Journal of Marine Science, 65(6):851–861, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/6/851/600867.

Pascoe:2001:PVH

[PCM01] Sean Pascoe, Louisa Coglan, and Simon Mardle. Physical versus harvest-based measures of capacity: the case of the United Kingdom vessel capacity unit system. ICES Journal of Marine Science, 58(6):1243–1252, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/6/1243/641561.

Payne:2009:FSN

[PCM09] Mark R. Payne, Lotte Worsøe Clausen, and Henrik Mosegaard. Finding the signal in the noise: objective data-selection criteria improve the assessment of western Baltic springspawning herring. ICES Journal of Marine Science, 66(8):1673– 1680, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/8/1673/677024.

Poloczanska:2004:FVN

[PCRW04] Elvira S. Poloczanska, Robin M. Cook, Graeme D. Ruxton, and Peter J. Wright. Fishing vs. natural recruitment variation in sandeels as a cause of seabird breeding failure at Shetland: a modelling approach. ICES Journal of Marine Science, 61(5):788–797, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/5/788/866027.

Potter:2004:EFP

[PCS^{+}04] E. C. E. Potter, W. W. Crozier, P-J. Schön, M. D. Nicholson, D. L. Maxwell, E. Prévost, J. Erkinaro, G. Gudbergsson, L. Karlsson, L. P. Hansen, J. C. MacLean, N. Ó Maoiléidigh, and S. Prusov. Estimating and forecasting pre-fishery abundance of Atlantic salmon (Salmo salar L.) in the Northeast Atlantic for the management of mixed-stock fisheries. ICES Journal of Marine Science, 61(8):1359–1369, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/8/1359/631948.

Pepin:2007:OECb

[PCS^{+}07a] P. Pepin, K. A. Curtis, P. V. R. Snelgrove, B. de Young, and J. A. Heilbig. Optimal estimation of catch by the continuous underway fish egg sampler based on a model of the vertical distribution of American plaice (Hippoglossoides platessoides) eggs. ICES Journal of Marine Science, 64(5):1075, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/5/1075/643556.

Pepin:2007:OECa

[PCS^{+}07b] P. Pepin, K. A. Curtis, P. V. R. Snelgrove, B. de Young, and J. A. Helbig. Optimal estimation of catch by the continuous underway fish egg sampler based on a model of the vertical distribution of American plaice (Hippoglossoides platessoides) eggs. ICES Journal of Marine Science, 64(1):18–30, January 2007.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/1/18/648568.

Pauly:2000:PEP

[PCW00] Daniel Pauly, Villy Christensen, and Carl Walters. Ecopath, Ecosim, and Ecospace as tools for evaluating ecosystem impact of fisheries. ICES Journal of Marine Science, 57(3):697–706, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/3/697/635961.

Punt:2007:DMP

[PD07] André E. Punt and Greg P. Donovan. Developing management procedures that are robust to uncertainty: lessons from the International Whaling Commission. ICES Journal of Marine Science, 64(4):603–612, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/64/4/603/641398.

Pranovi:2004:MSI

[PDRG04] F. Pranovi, F. Da Ponte, S. Raicevich, and O. Giovanardi. A multidisciplinary study of the immediate effects of mechanical clam harvesting in the Venice lagoon. ICES Journal of Marine Science, 61(1):43–52, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/61/1/43/701859.

Pedchenko:2005:RIE

[Ped05] Andrey P. Pedchenko. The role of interannual environmental variations in the geographic range of spawning and feeding concentrations of redfish Sebastes mentella in the Irminger Sea. ICES Journal of Marine Science, 62(7):1501–1510, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/7/1501/661300.

Peltonen:2002:ADB

[Pel02] H. Peltonen. Age determination of Baltic herring from whole otoliths and from neutral red stained otolith cross sections. ICES Journal of Marine Science, 59(2):323–332, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/2/323/619643.

Penas:2007:FCP

[Pen07] Ernesto Penas. The fishery conservation policy of the European union after 2002: towards long-term sustainability. ICES Journal of Marine Science, 64(4):588–595, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/4/588/642862.

Pena:2008:STS

[Peñ08] Hector Peña. In situ target-strength measurements of Chilean jack mackerel (Trachurus symmetricus murphyi) collected with a scientific echosounder installed on a fishing vessel. ICES Journal of Marine Science, 65(4):594–604, May 2008. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/4/594/639571.

Petitgas:2001:RBS

[Pet01] P. Petitgas. On the relation between schools, clusters of schools, and abundance in pelagic fish stocks. ICES Journal of Marine Science, 58(6):1150–1160, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/6/1150/641529.

Petitgas:2003:MIC

[Pet03] Pierre Petitgas. A method for the identification and characterization of clusters of schools along the transect lines of fisheries-acoustic surveys. ICES Journal of Marine Science, 60 (4):872–884, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/4/872/694129.

Peterman:2004:PSS

[Pet04] Randall M. Peterman. Possible solutions to some challenges facing fisheries scientists and managers. ICES Journal of Marine Science, 61(8):1331–1343, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/8/1331/631060.

Pitois:2006:LTC

[PF06] Sophie G. Pitois and Clive J. Fox. Long-term changes in zooplankton biomass concentration and mean size over the Northwest European shelf inferred from continuous plankton recorder data. ICES Journal of Marine Science, 63(5):785–798, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/5/785/661324.

Pena:2008:MTS

[PF08a] Héctor Peña and Kenneth G. Foote. Modelling the target strength of Trachurus symmetricus murphyi based on highresolution swimbladder morphometry using an MRI scanner. ICES Journal of Marine Science, 65(9):1751–1761, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/9/1751/633074.

Pitois:2008:EMP

[PF08b] Sophie G. Pitois and Clive J. Fox. Empirically modelling the potential effects of changes in temperature and prey availability on the growth of cod larvae in UK shelf seas. ICES Journal of Marine Science, 65(9):1559–1572, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/9/1559/630620.

Pohle:2001:ARB

[PFF01] G. Pohle, B. Frost, and R. Findlay. Assessment of regional benthic impact of salmon mariculture within the Letang Inlet, Bay of Fundy. ICES Journal of Marine Science, 58(2):417–426, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/2/417/617841.

Pedersen:2009:NSF

[PFK^{+}09] Søren Anker Pedersen, Heino Fock, Jochen Krause, Christian Pusch, Anne L. Sell, Uwe Böttcher, Stuart I. Rogers, Mattias Sköld, Henrik Skov, Magdalena Podolska, Gerjan J. Piet, and Jake C. Rice. Natura 2000 sites and fisheries in German offshore waters. ICES Journal of Marine Science, 66(1):155–169, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/1/155/634370.

Peteiro:2008:GBR

[PFLFR08] Laura G. Peteiro, Ramón Filgueira, Uxío Labarta, and María José Fernández-Reiriz. Growth and biochemical responses of the offspring of mussels directly affected by the “Prestige” oil spill. ICES Journal of Marine Science, 65(4):509–513, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/4/509/635844.

Packard:2008:EFP

[PG08] Ted T. Packard and May Gómez. Exploring a first-principles-based model for zooplankton respiration. ICES Journal of Marine Science, 65(3):371–378, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/65/3/371/784567.

Potelov:2003:EPP

[PGB03] V. A. Potelov, A. P. Golikov, and V. A. Bondarev. Estimated pup production of harp seals Pagophilus groenlandicus in the White Sea, Russia, in 2000. ICES Journal of Marine Science, 60(5):1012–1017, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/5/1012/763806.

Pampoulie:2009:SGS

[PGD09] Christophe Pampoulie, David Gíslason, and Anna Kristin Daníelsdóttir. A “seascape genetic” snapshot of Sebastes marinus calls for further investigation across the North Atlantic. ICES Journal of Marine Science, 66(10):2219–2222, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/10/2219/681502.

Pascual:2005:RGI

[PGG05] S. Pascual, A. F. González, and A. Guerra. The recruitment of gill-infesting copepods as a categorical predictor of size-at-age data in squid populations. ICES Journal of Marine Science, 62(4):629–633, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/4/629/672971.

Pershing:2005:IVG

[PGJ^{+}05] Andrew J. Pershing, Charles H. Greene, Jack W. Jossi, Loretta O’Brien, Jon K. T. Brodziak, and Barbara A. Bailey. Interdecadal variability in the Gulf of Maine zooplankton community, with potential impacts on fish recruitment. ICES Journal of Marine Science, 62(7):1511–1523, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/7/1511/661385.

Pinho:2001:BAC

[PGM01] M. R. Pinho, J. M. Gonçalves, and H. R. Martins. Biology and abundance of Cancer bellianus (Decapoda, Brachyura) around the Azores. ICES Journal of Marine Science, 58(4):896–903, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/4/896/630225.

Petitgas:2009:CAC

[PGMB09] Pierre Petitgas, Anne Goarant, Jacques Massé, and Paul Bourriau. Combining acoustic and CUFES data for the quality control of fish-stock survey estimates. ICES Journal of Marine Science, 66(6):1384–1390, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/6/1384/689698.

Podolska:2003:IBH

[PH03] Magdalena Podolska and Jan Horbowy. Infection of Baltic herring (Clupea harengus membras) with Anisakis simplex larvae, 1992–1999: a statistical analysis using generalized linear models. ICES Journal of Marine Science, 60(1):85–93, ???? 2003. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/60/1/85/
611460.

Polovina:2005:EID

[PH05] Jeffrey J. Polovina and Evan A. Howell. Ecosystem indicators derived from satellite remotely sensed oceanographic data for the North Pacific. ICES Journal of Marine Science, 62(3):319–327, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/3/319/658873.

Payne:2009:RCE

[PHDC^{+}09] Mark R. Payne, Emma M. C. Hatfield, Mark Dickey-Collas, Tone Falkenhaug, Alejandro Gallego, Joachim Gröger, Priscilla Licandro, Marcos Llope, Peter Munk, Christine Röckmann, Jörn O. Schmidt, and Richard D. M. Nash. Recruitment in a changing environment: the 2000s North Sea herring recruitment failure. ICES Journal of Marine Science, 66(2):272–277, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/2/272/595050.

Patel:2004:BHC

[PHG04] Ruben Patel, Nils Olav Handegard, and Olav Rune Godø. Behaviour of herring (Clupea harengus L.) towards an approaching autonomous underwater vehicle. ICES Journal of Marine Science, 61(7):1044–1049, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/7/1044/877719.

Pedersen:2009:LAT

[PHO09] Geir Pedersen, Nils Olav Handegard, and Egil Ona. Lateralaspect, target-strength measurements of in situ herring (Clupea harengus). ICES Journal of Marine Science, 66(6):1191–1196, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1191/696175.

Piet:2000:QEI

[Pie00] G. Piet. A quantitative evaluation of the impact of beam trawling on benthic fauna in the southern North Sea. ICES Journal of Marine Science, 57(5):1332–1339, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/5/1332/660931.

Piet:2002:UEI

[Pie02] G. J. Piet. Using external information and GAMs to improve catch-at-age indices for North Sea plaice and sole. ICES Journal of Marine Science, 59(3):624–632, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/3/624/610847.

Pielou:2008:PLI

[Pie08] E. C. Pielou. Plankton, from the last ice age to the year 3007. ICES Journal of Marine Science, 65(3):296–301, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/3/296/785576.

Piet:2005:RPF

[PJ05] G. J. Piet and S. Jennings. Response of potential fish community indicators to fishing. ICES Journal of Marine Science, 62 (2):214–225, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/2/214/603422.

Pederson:2008:GAS

[PJ08] Hugh G. Pederson and Craig R. Johnson. Growth and age structure of sea urchins (Heliocidaris erythrogramma) in complex barrens and native macroalgal beds in eastern Tasmania. ICES Journal of Marine Science, 65(1):1–11, January 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/1/1/612269.

Piet:2008:EPI

[PJR08] Gerjan J. Piet, Henrice M. Jansen, and Marie-Joëlle Rochet. Evaluating potential indicators for an ecosystem approach to fishery management in European waters. ICES Journal of Marine Science, 65(8):1449–1455, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/8/1449/713894.

Payne:2007:PIJ

[PK07] Andrew I. L. Payne and Cathy Kennedy. Publishing the ICES Journal of Marine Science: facing fresh challenges in a competitive environment. ICES Journal of Marine Science, 64(1):1–2, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/1/1/647378.

Pedersen:2009:RFR

[PK09] Geir Pedersen and Rolf J. Korneliussen. The relative frequency response derived from individually separated targets of northeast Arctic cod (Gadus morhua), saithe (Pollachius virens), and Norway pout (Trisopterus esmarkii). ICES Journal of Marine Science, 66(6):1149–1154, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1149/692294.

Pilling:2008:CEB

[PKH^{+}08] Graham M. Pilling, Laurence T. Kell, Trevor Hutton, Peter J. Bromley, Alex N. Tidd, and Loes J. Bolle. Can economic and biological management objectives be achieved by the use of MSY-based reference points? A North Sea plaice (Pleuronectes platessa) and sole (Solea solea) case study. ICES Journal of Marine Science, 65(6):1069–1080, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/6/1069/601340.

Pawson:2007:SSB

[PKP07] Mike G. Pawson, Sven Kupschus, and Graham D. Pickett. The status of sea bass (Dicentrarchus labrax) stocks around England and Wales, derived using a separable catch-at-age model, and implications for fisheries management. ICES Journal of Marine Science, 64(2):346–356, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/2/346/2182413.

Probst:2006:PEE

[PKRT06] W. Nikolaus Probst, Gerd Kraus, Rick M. Rideout, and Edward A. Trippel. Parental effects on early life history traits of haddock Melanogrammus aeglefinus. ICES Journal of Marine Science, 63(2):224–234, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/63/2/224/638234.

Pope:2001:RIP

[PLJ01] John Pope, Philip Large, and Tore Jakobsen. Revisiting the influences of parent stock, temperature, and predation on the recruitment of the Northeast Arctic cod stock, 1930–1990. ICES Journal of Marine Science, 58(5):967–972, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/5/967/730871.

Peltonen:2007:PFA

[PLP^{+}07] Heikki Peltonen, Miska Luoto, Jari-Pekka Pääkkönen, Miina Karjalainen, Antti Tuomaala, Jukka Pönni, and Markku Viitasalo. Pelagic fish abundance in relation to regional environmental variation in the Gulf of Finland, northern Baltic Sea. ICES Journal of Marine Science, 64(3):487–495, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/3/487/812900.

Peralba:2004:VSD

[PM04] Àurea Peralba and Maria Grazia Mazzocchi. Vertical and seasonal distribution of eight Clausocalanus species (Copepoda: Calanoida) in oligotrophic waters. ICES Journal of Marine Science, 61(4):645–653, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/61/4/645/606387.

Pecl:2006:LHS

[PM06] Gretta T. Pecl and Natalie A. Moltschaniwskyj. Life history of a short-lived squid (Sepioteuthis australis): resource allocation as a function of size, growth, maturation, and hatching season. ICES Journal of Marine Science, 63(6):995–1004, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/6/995/619446.

Pavlov:2000:OMC

[PMB00] Dimitri A. Pavlov, Erlend Moksness, and Vladimir A. Burmenski. Otolith microstructure characteristics in White Sea spring-spawning herring (Clupea pallasi marisalbi Berg) larvae. ICES Journal of Marine Science, 57(4):1069–1076, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/4/1069/647261.

Petitgas:2003:SVS

[PMB^{+}03a] Pierre Petitgas, Jacques Massé, Pierre Beillois, Emilie Lebarbier, and Arnaud Le Cann. Sampling variance of species identification in fisheries acoustic surveys based on automated procedures associating acoustic images and trawl hauls. ICES Journal of Marine Science, 60(3):437–445, ???? 2003. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/3/437/657703.

Poder:2003:RIO

[PMB^{+}03b] Tõnis Põder, Serge Y. Maestrini, Maija Balode, Urmas Lips, Christian Béchemin, Andris Andrushaitis, and Ingrida Purina. The role of inorganic and organic nutrients on the development of phytoplankton along a transect from the Daugava River mouth to the Open Baltic, in spring and summer 1999. ICES Journal of Marine Science, 60(4):827–835, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/4/827/693327.

Parsons:2008:SIM

[PMB^{+}08] Matt Parsons, Ian Mitchell, Adam Butler, Norman Ratcliffe, Morten Frederiksen, Simon Foster, and James B. Reid. Seabirds as indicators of the marine environment. ICES Journal of Marine Science, 65(8):1520–1526, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/8/1520/714606.

Pope:2000:GIF

[PMD^{+}00] John G. Pope, Dave S. Macdonald, Niels Daan, John D. Reynolds, and Simon Jennings. Gauging the impact of fishing mortality on non-target species. ICES Journal of Marine Science, 57(3):689–696, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/3/689/635959.

Parsons:2009:LIM

[PMM^{+}09] Miles J. Parsons, Robert D. McCauley, Michael C. Mackie, Paulus Siwabessy, and Alec J. Duncan. Localization of individual mulloway (Argyrosomus japonicus) within a spawning aggregation and their behaviour throughout a diel spawning period. ICES Journal of Marine Science, 66(6):1007–1014, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/1007/690513.

Petrakis:2001:DND

[PMN01] G. Petrakis, D. N. MacLennan, and A. W. Newton. Day– night and depth effects on catch rates during trawl surveys in the North Sea. ICES Journal of Marine Science, 58(1):50–60, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/1/50/603552.

Podolska:2006:AAH

[PN06] Magdalena Podolska and Dorota Napierska. Acetylcholinesterase activity in hosts (herring Clupea harengus) and parasites (Anisakis simplex larvae) from the southern Baltic. ICES Journal of Marine Science, 63(1):161–168, ???? 2006. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/1/161/623964.

Patel:2009:MHD

[PO09] Ruben Patel and Egil Ona. Measuring herring densities with one real and several phantom research vessels. ICES Journal of Marine Science, 66(6):1264–1269, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1264/696844.

Powles:2000:APE

[Pow00] H. Powles. Assessing and protecting endangered marine species. ICES Journal of Marine Science, 57(3):669–676, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/3/669/635955.

Peterson:2008:ICR

[PP08] Jay O. Peterson and William T. Peterson. Influence of the Columbia River plume (USA) on the vertical and horizontal distribution of mesozooplankton over the Washington and Oregon shelf. ICES Journal of Marine Science, 65(3):477–483, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/3/477/785383.

Petitgas:2003:CCR

[PPB03] Pierre Petitgas, Jean Charles Poulard, and Alain Biseau. Comparing commercial and research survey catch per unit of effort: megrim in the Celtic Sea. ICES Journal of Marine Science, 60(1):66–76, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/1/66/611450.

Prevost:2003:SBR

[PPC^{+}03] Etienne Prévost, Eric Parent, Walter Crozier, Ian Davidson, Jacques Dumas, Gudni Gudbergsson, Kjetil Hindar, Phil McGinnity, Julian MacLean, and Leif M. Sættem. Setting biological reference points for Atlantic salmon stocks: transfer of information from data-rich to sparse-data situations by Bayesian hierarchical modelling. ICES Journal of Marine Science, 60 (6):1177–1193, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/6/1177/651308.

Plourde:2009:LTS

[PPH09] Stéphane Plourde, Pierre Pepin, and Erica J. H. Head. Longterm seasonal and spatial patterns in mortality and survival of Calanus finmarchicus across the Atlantic zone monitoring programme region, Northwest Atlantic. ICES Journal of Marine Science, 66(9):1942–1958, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/66/9/1942/725089.

Pitcher:2000:IET

[PPHB00] C. R. Pitcher, I. R. Poiner, B. J. Hill, and C. Y. Burridge. Implications of the effects of trawling on sessile megazoobenthos on a tropical shelf in northeastern Australia. ICES Journal of Marine Science, 57(5):1359–1368, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/5/1359/660934.

Pitcher:2001:E

[PPHB01] C. R. Pitcher, I. R. Poiner, B. J. Hill, and C. Y. Burridge. Erratum. ICES Journal of Marine Science, 58(1):49, ???? 2001. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/58/1/49/
603551.

Park:2006:EER

[PPK^{+}06] Heum Gi Park, Velmurugu Puvanendran, Anne Kellett, Christopher C. Parrish, and Joseph A. Brown. Effect of enriched rotifers on growth, survival, and composition of larval Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 63 (2):285–295, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/2/285/640037.

Pastoors:2007:VMS

[PPKM07] Martin A. Pastoors, Jan Jaap Poos, Sarah B. M. Kraak, and Marcel A. M. Machiels. Validating management simulation models and implications for communicating results to stakeholders. ICES Journal of Marine Science, 64(4):818–824, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/4/818/642713.

Pawson:2007:MFI

[PPL^{+}07] M. G. Pawson, G. D. Pickett, J. Leballeur, M. Brown, and M. Fritsch. Migrations, fishery interactions, and management units of sea bass (Dicentrarchus labrax) in Northwest Europe. ICES Journal of Marine Science, 64(2):332–345, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/2/332/2182514.

Popper:2004:RCF

[PPMH04] Arthur N. Popper, Dennis T. T. Plachta, David A. Mann, and Dennis Higgs. Response of clupeid fish to ultrasound: a review. ICES Journal of Marine Science, 61(7):1057–1061, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1057/877888.

Pedersen:2009:IST

[PPTS09] O. P. Pedersen, T. Pedersen, K. S. Tande, and D. Slagstad. Integrating spatial and temporal mortality from herring on capelin larvae: a study in the Barents Sea. ICES Journal of Marine Science, 66(10):2183–2194, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/10/2183/680435.

Preuss:2009:CMS

[PPW^{+}09] Bastien Preuss, Dominique Pelletier, Laurent Wantiez, Yves Letourneur, Sébastien Sarramégna, Michel Kulbicki, René Galzin, and Jocelyne Ferraris. Considering multiple-species attributes to understand better the effects of successive changes in protection status on a coral reef fish assemblage. ICES Journal of Marine Science, 66(1):170–179, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/1/170/634976.

Piet:2007:PPI

[PQRG07] G. J. Piet, F. J. Quirijns, L. Robinson, and S. P. R. Greenstreet. Potential pressure indicators for fishing, and their data requirements. ICES Journal of Marine Science, 64(1):110–121, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/1/110/646890.

Pinn:2001:FAE

[PR01] Eunice H. Pinn and M. R. Robertson. Further analysis of the effect of bioturbation by Nephrops norvegicus (L.) on the acoustic return of the RoxAnn^{T M} seabed discrimination system. ICES Journal of Marine Science, 58(1):216–219, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/1/216/603520.

Pinn:2003:ETS

[PR03] Eunice H. Pinn and M. R. Robertson. Effect of track spacing and data interpolation on the interpretation of benthic community distributions derived from RoxAnn^{T M} acoustic surveys. ICES Journal of Marine Science, 60(6):1288–1297, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/6/1288/653433.

Piet:2004:PPR

[PR04] G. J. Piet and J. C. Rice. Performance of precautionary reference points in providing management advice on North Sea fish stocks. ICES Journal of Marine Science, 61(8):1305–1312, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/8/1305/630446.

Patterson:2007:CSL

[PR07] Kenneth Patterson and Martine Résimont. Change and stability in landings: the responses of fisheries to scientific advice and TACs. ICES Journal of Marine Science, 64(4):714–717, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/4/714/641564.

Plaganyi:2007:MMP

[PRB^{+}07] Éva E. Plagányi, Rebecca A. Rademeyer, Doug S. Butterworth, Carryn L. Cunningham, and Susan J. Johnston. Making management procedures operational — innovations implemented in South Africa. ICES Journal of Marine Science, 64(4):626–632, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/4/626/641982.

Pope:2006:MEM

[PRD^{+}06] John G. Pope, Jake C. Rice, Niels Daan, Simon Jennings, and Henrik Gislason. Modelling an exploited marine fish community with 15 parameters — results from a simple size-based model. ICES Journal of Marine Science, 63(6):1029–1044, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/6/1029/614699.

Pranovi:2000:RTN

[PRF^{+}00] F. Pranovi, S. Raicevich, G. Franceschini, M. G. Farrace, and O. Giovanardi. Rapido trawling in the northern Adriatic Sea: effects on benthic communities in an experimental area. ICES Journal of Marine Science, 57(3):517–524, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/3/517/635924.

Pastoors:2000:EPC

[PRvB00] M. A. Pastoors, A. D. Rijnsdorp, and F. A. van Beek. Effects of a partially closed area in the North Sea (“plaice box”) on stock development of plaice. ICES Journal of Marine Science, 57(4): 1014–1022, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/4/1014/647242.

Pineiro:2003:AEG

[PS03] C. Piñeiro and M. Saínza. Age estimation, growth and maturity of the European hake (Merluccius merluccius (Linnaeus, 1758)) from Iberian Atlantic waters. ICES Journal of Marine Science, 60(5):1086–1102, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/5/1086/766268.

Prokopchuk:2006:DHM

[PS06] Irina Prokopchuk and Evgeniy Sentyabov. Diets of herring, mackerel, and blue whiting in the Norwegian Sea in relation to Calanus finmarchicus distribution and temperature conditions. ICES Journal of Marine Science, 63(1):117–127, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/1/117/623146.

Palmer:2009:WPD

[PS09] Ryan M. Palmer and Jen D. Snowball. The willingness to pay for dusky kob (Argyrosomus japonicus) restocking: using recreational linefishing licence fees to fund stock enhancement in South Africa. ICES Journal of Marine Science, 66(5):839–843, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/5/839/663346.

Pondella:2002:FPT

[PSC02] Daniel J. Pondella II, John S. Stephens, Jr., and Matthew T. Craig. Fish production of a temperate artificial reef based on the density of embiotocids (Teleostei: Perciformes). ICES Journal of Marine Science, 59(S1):S88–S93, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/59/suppl/S88/618013.

Platt:2007:BOF

[PSFY07] Trevor Platt, Shubha Sathyendranath, and César Fuentes-Yaco. Biological oceanography and fisheries management: perspective after 10 years. ICES Journal of Marine Science, 64(5):863–869, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/5/863/642505.

Parker-Stetter:2009:IMH

[PSHL09] Sandra L. Parker-Stetter, John K. Horne, and Mariko M. Langness. The influence of midwater hypoxia on nekton vertical migration. ICES Journal of Marine Science, 66(6):1296–1302, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1296/689387.

Park:2004:IAP

[PSO^{+}04] Wongyu Park, Molly Sturdevant, Joseph Orsi, Alex Wertheimer, Emily Fergusson, William Heard, and Thomas Shirley. Interannual abundance patterns of copepods during an ENSO event in Icy Strait, southeastern Alaska. ICES Journal of Marine Science, 61(4):464–477, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/61/4/464/602451.

Pampoulie:2008:GDA

[PSSD08] Christophe Pampoulie, Pétur Steingrund, Magnus Ö. Stefánsson, and Anna K. Daníelsdóttir. Genetic divergence among East Icelandic and Faroese populations of Atlantic cod provides evidence for historical imprints at neutral and non-neutral markers. ICES Journal of Marine Science, 65(1):65–71, January 2008. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/65/1/65/
613915.

Petihakis:2007:STF

[PST^{+}07] G. Petihakis, C. J. Smith, G. Triantafyllou, G. Sourlantzis, K-N. Papadopoulou, A. Pollani, and G. Korres. Scenario testing of fisheries management strategies using a high resolution ERSEM– POM ecosystem model. ICES Journal of Marine Science, 64(9): 1627–1640, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/9/1627/788451.

Pedersen:2000:PCB

[PTTS00] O-P. Pedersen, K. S. Tande, A. Timonin, and T. Semenova. A possible connection between hydrography and the distribution of Calanus finmarchicus on the Norwegian midshelf in 1997. ICES Journal of Marine Science, 57(6):1645–1655, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/6/1645/768596.

Pusineri:2005:FFE

[PVH^{+}05] C. Pusineri, Y. Vasseur, S. Hassani, L. Meynier, J. Spitz, and V. Ridoux. Food and feeding ecology of juvenile albacore, Thunnus alalunga, off the Bay of Biscay: a case study. ICES Journal of Marine Science, 62(1):116–122, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/1/116/666626.

Piet:2009:MDI

[PvHG09] G. J. Piet, R. van Hal, and S. P. R. Greenstreet. Modelling the direct impact of bottom trawling on the North Sea fish community to derive estimates of fishing mortality for non-target fish species. ICES Journal of Marine Science, 66(9):1985–1998, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/9/1985/723772.

Peltonen:2004:SFP

[PVLP04] Heikki Peltonen, Mika Vinni, Antti Lappalainen, and Jukka Pönni. Spatial feeding patterns of herring (Clupea harengus L.), sprat (Sprattus sprattus L.), and the three-spined stickleback (Gasterosteus aculeatus L.) in the Gulf of Finland, Baltic Sea. ICES Journal of Marine Science, 61(6):966–971, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/6/966/678771.

Pedersen:2005:EFC

[PZTE05] O. P. Pedersen, M. Zhou, K. S. Tande, and A. Edvardsen. Eddy formation on the coast of North Norway — evidenced by synoptic sampling. ICES Journal of Marine Science, 62(4):615–628, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/4/615/672889.

Quinta:2004:MDP

[QGdS04] Ricardo Quinta, Laurentina Gomes, and Ana Teia dos Santos. A mitochondrial DNA PCR–RFLP marker for population studies of the black scabbardfish (Aphanopus carbo). ICES Journal of Marine Science, 61(5):864–867, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/5/864/868984.

Reid:2007:ACS

[RAB^{+}07] D. G. Reid, V. J. Allen, D. J. Bova, E. G. Jones, R. J. Kynoch, K. J. Peach, P. G. Fernandes, and W. R. Turrell. Anglerfish catchability for swept-area abundance estimates in a new survey trawl. ICES Journal of Marine Science, 64(8):1503– 1511, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/8/1503/613521.

Radtke:2003:EEE

[Rad03] Krzysztof Radtke. Evaluation of the exploitation of Eastern Baltic cod (Gadus morhua callarias L.) stock in 1976–1997. ICES Journal of Marine Science, 60(5):1114–1122, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/5/1114/766882.

Rikardsen:2006:SMF

[RAKS06] A. H. Rikardsen, P-A. Amundsen, R. Knudsen, and S. Sandring. Seasonal marine feeding and body condition of sea trout (Salmo trutta) at its northern distribution. ICES Journal of Marine Science, 63(3):466–475, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/3/466/720862.

Reid:2007:CSM

[RAR^{+}07] David G. Reid, John Annala, Shale Rosen, Mike Pol, Steve X. Cadrin, and Stephen J. Walsh. Current status of mobile and static sampling gears used in resource surveys. ICES Journal of Marine Science, 64(8):1607–1609, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/8/1607/614006.

Rothschild:2009:VRB

[RASS09] Brian J. Rothschild, Charles F. Adams, Christopher L. Sarro, and Kevin D. E. Stokesbury. Variability in the relationship between sea scallop shell height and meat weight. ICES Journal of Marine Science, 66(9):1972–1977, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/9/1972/726292.

Rayns:2007:AGH

[Ray07] Nick Rayns. The Australian government’s harvest strategy policy. ICES Journal of Marine Science, 64(4):596–598, May 2007.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/4/596/641228.

Reid:2000:IFP

[RBBB00] P. C. Reid, E. J. V. Battle, S. D. Batten, and K. M. Brander. Impacts of fisheries on plankton community structure. ICES Journal of Marine Science, 57(3):495–502, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/3/495/635917.

Ratz:2007:SMM

[RBD^{+}07] Hans-Joachim Rätz, Eckhard Bethke, Hendrik Dörner, Doug Beare, and Joachim Gröger. Sustainable management of mixed demersal fisheries in the North Sea through fleet-based management — a proposal from a biological perspective. ICES Journal of Marine Science, 64(4):652–660, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/652/641141.

Rotherham:2008:DBT

[RBGJ08] Douglas Rotherham, Matt K. Broadhurst, Charles A. Gray, and Daniel D. Johnson. Developing a beam trawl for sampling estuarine fish and crustaceans: assessment of a codend cover and effects of different sizes of mesh in the body and codend. ICES Journal of Marine Science, 65(4):687–696, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/4/687/639428.

Rice:2007:FMS

[RC07] Jake C. Rice and Paul L. Connolly. Fisheries management strategies: an introduction by the conveners. ICES Journal of Marine Science, 64(4):577–579, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/577/643237.

Reid:2005:AEM

[RCBM05] Keith Reid, John P. Croxall, Dirk R. Briggs, and Eugene J. Murphy. Antarctic ecosystem monitoring: quantifying the response of ecosystem indicators to variability in Antarctic krill. ICES Journal of Marine Science, 62(3):366–373, ???? 2005.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/3/366/660016.

Rothschild:2005:MFS

[RCL05] Brian J. Rothschild, Changsheng Chen, and R. Greg Lough. Managing fish stocks under climate uncertainty. ICES Journal of Marine Science, 62(7):1531–1541, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/7/1531/661658.

Reiss:2008:VBA

[RCLD08] Christian S. Reiss, Anthony M. Cossio, Valerie Loeb, and David A. Demer. Variations in the biomass of Antarctic krill (Euphausia superba) around the South Shetland Islands, 1996–2006. ICES Journal of Marine Science, 65(4):497–508, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/4/497/637755.

Rueda:2001:SAI

[RD01] Mario Rueda and Omar Defeo. Survey abundance indices in a tropical estuarine lagoon and their management implications: a spatially-explicit approach. ICES Journal of Marine Science, 58 (6):1219–1231, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1219/641559.

Rueda:2003:BMA

[RD03] Mario Rueda and Omar Defeo. A bioeconomic multispecies analysis of an estuarine small-scale fishery: spatial structure of biovalue. ICES Journal of Marine Science, 60(4):721–732, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/4/721/691067.

Roel:2007:HCR

[RD07] Beatriz A. Roel and José A. A. De Oliveira. Harvest control rules for the western horse mackerel (Trachurus trachurus) stock given paucity of fishery-independent data. ICES Journal of Marine Science, 64(4):661–670, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/661/640975.

Roel:2009:TSB

[RDB09] Beatriz A. Roel, José A. A. De Oliveira, and Steven Beggs. A two-stage biomass model for Irish Sea herring allowing for additional variance in the recruitment index caused by mixing of stocks. ICES Journal of Marine Science, 66(8):1808– 1813, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/8/1808/675299.

Rijnsdorp:2006:PFM

[RDD06] Adriaan D. Rijnsdorp, Niels Daan, and Willem Dekker. Partial fishing mortality per fishing trip: a useful indicator of effective fishing effort in mixed demersal fisheries. ICES Journal of Marine Science, 63(3):556–566, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/3/556/723358.

Ramster:2003:SVJ

[RDF^{+}03] John Ramster, Niels Daan, Chris Frid, Bernard Megrey, Andrew Payne, and Pierre Pepin. Sixty volumes of the Journal du Conseil /ICES Journal of Marine Science. ICES Journal of Marine Science, 60(6):1169–1171, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/6/1169/651056.

Rijnsdorp:2000:EFP

[RDHP00] A. D. Rijnsdorp, W. Dol, M. Hoyer, and M. A. Pastoors. Effects of fishing power and competitive interactions among vessels on the effort allocation on the trip level of the Dutch beam trawl fleet. ICES Journal of Marine Science, 57(4):927–937, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/927/647383.

Rogers:2000:CDF

[RE00] S. I. Rogers and J. R. Ellis. Changes in the demersal fish assemblages of British coastal waters during the 20th century. ICES Journal of Marine Science, 57(4):866–881, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/4/866/647356.

Reeves:2003:SSI

[Ree03] Stuart A. Reeves. A simulation study of the implications of age-reading errors for stock assessment and management advice. ICES Journal of Marine Science, 60(2):314–328, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/2/314/626253.

Ravier:2001:LTF

[RF01] Christelle Ravier and Jean-Marc Fromentin. Long-term fluctuations in the eastern Atlantic and Mediterranean bluefin tuna population. ICES Journal of Marine Science, 58(6):1299–1317, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/6/1299/641577.

Rudels:2002:EGC

[RFM^{+}02] Bert Rudels, Eberhard Fahrbach, Jens Meincke, Gereon Budéus, and Patrick Eriksson. The East Greenland Current and its contribution to the Denmark Strait overflow. ICES Journal of Marine Science, 59(6):1133–1154, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/6/1133/608168.

Ruckert:2002:EHM

[RFT02] Chris Rückert, Jens Floeter, and Axel Temming. An estimate of horse mackerel biomass in the North Sea, 1991–1997. ICES Journal of Marine Science, 59(1):120–130, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/1/120/649971.

Rosenberg:2007:FRM

[RG07] Andrew A. Rosenberg and Christopher W. Glass. Fishers’ responses to management measures and their socio-economic effects. ICES Journal of Marine Science, 64(8):1612–1613, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/64/8/1612/614018.

Rocha:2004:RMB

[RGG^{+}04] Francisco Rocha, Joaquín Gracia, Ángel F. González, Carlos M. Jardón, and Ángel Guerra. Reliability of a model based on a short fishery statistics survey: application to the Northeast Atlantic monkfish fishery. ICES Journal of Marine Science, 61(1):25–34, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/1/25/701397.

Rowe:2004:DPF

[RHBR04] Sherrylynn Rowe, Jeffrey A. Hutchings, Dorte Bekkevold, and Ana Rakitin. Depensation, probability of fertilization, and the mating system of Atlantic cod (Gadus morhua L.). ICES Journal of Marine Science, 61(7):1144–1150, ???? 2004. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1144/880763.

Renfree:2009:SSS

[RHD09] Josiah S. Renfree, Sean A. Hayes, and David A. Demer. Sound-scattering spectra of steelhead (Oncorhynchus mykiss), Coho (O. kisutch), and Chinook (O. tshawytscha) salmonids. ICES Journal of Marine Science, 66(6):1091–1099, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/1091/692249.

Rees:2008:EIU

[RHH^{+}08] Hubert L. Rees, Jeffrey L. Hyland, Ketil Hylland, Colleen S. L. Mercer Clarke, John C. Roff, and Suzanne Ware. Environmental indicators: utility in meeting regulatory needs. an overview. ICES Journal of Marine Science, 65(8):1381– 1386, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/8/1381/714382.

Rhodes:2008:EAA

[Rho08] Christopher J. Rhodes. Excess acoustic absorption attributable to the biological modification of seawater viscosity. ICES Journal of Marine Science, 65(9):1747–1750, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/9/1747/632155.

Rice:2000:EFI

[Ric00a] Jake C. Rice. Evaluating fishery impacts using metrics of community structure. ICES Journal of Marine Science, 57 (3):682–688, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/682/635958.

Richardson:2000:IEF

[Ric00b] Katherine Richardson. Integrating environment and fisheries management objectives in the ICES area: reflections of a past ACME chair. ICES Journal of Marine Science, 57(3):766–770, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/3/766/636009.

Richardson:2008:HWZ

[Ric08] Anthony J. Richardson. In hot water: zooplankton and climate change. ICES Journal of Marine Science, 65(3):279–295, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/3/279/787309.

Rice:2009:GTS

[Ric09] Jake C. Rice. A generalization of the three-stage model for advice using the precautionary approach in fisheries, to apply broadly to ecosystem properties and pressures. ICES Journal of Marine Science, 66(3):433–444, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/3/433/818029.

Rumohr:2000:ITF

[RK00] Heye Rumohr and Thomas Kujawski. The impact of trawl fishery on the epifauna of the southern North Sea. ICES Journal of Marine Science, 57(5):1389–1394, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/5/1389/660940.

Reiss:2004:SVE

[RK04] Henning Reiss and Ingrid Kröncke. Seasonal variability of epibenthic communities in different areas of the southern North Sea. ICES Journal of Marine Science, 61(6):882–905, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/6/882/677440.

Reiss:2006:ECE

[RKE06] Henning Reiss, Ingrid Kröncke, and Siegfried Ehrich. Estimating the catching efficiency of a 2-m beam trawl for sampling epifauna by removal experiments. ICES Journal of Marine Science, 63 (8):1453–1464, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/8/1453/713096.

Rostad:2006:FAV

[RKKM06] Anders Røstad, Stein Kaartvedt, Thor A. Klevjer, and Webjørn Melle. Fish are attracted to vessels. ICES Journal of Marine Science, 63(8):1431–1437, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/8/1431/712848.

Ryan:2009:MVV

[RKM09] Tim E. Ryan, Rudy J. Kloser, and Gavin J. Macaulay. Measurement and visual verification of fish target strength using an acoustic-optical system attached to a trawlnet. ICES Journal of Marine Science, 66(6):1238–1244, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1238/696363.

Rahikainen:2003:MEE

[RKP03] Mika Rahikainen, Sakari Kuikka, and Raimo Parmanne. Modelling the effect of ecosystem change on spawning per recruit of Baltic herring. ICES Journal of Marine Science, 60(1):94–109, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/1/94/611461.

Ratz:2005:LTV

[RL05] Hans-Joachim Rätz and Josep Lloret. Long-term variability of growth and recruitment of cod (Gadus morhua) off Greenland. ICES Journal of Marine Science, 62(7):1310–1321, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/7/1310/656621.

Rice:2007:WCR

[RL07] Jake C. Rice and Èmilie Legacè. When control rules collide: a comparison of fisheries management reference points and IUCN criteria for assessing risk of extinction. ICES Journal of Marine Science, 64(4):718–722, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/64/4/718/639981.

Reusser:2008:PIW

[RL08] Deborah A. Reusser and Henry Lee II. Predictions for an invaded world: a strategy to predict the distribution of native and non-indigenous species at multiple scales. ICES Journal of Marine Science, 65(5):742–745, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/5/742/712821.

Ruocco:2006:RBA

[RLdAW06] Natalia L. Ruocco, Luis O. Lucifora, Juan M. Díaz de Astarloa, and Otto Wöhler. Reproductive biology and abundance of the white-dotted skate, Bathyraja albomaculata, in the Southwest Atlantic. ICES Journal of Marine Science, 63(1):105–116, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/1/105/622604.

Ross:2001:UUP

[RLF01] B. P. Ross, J. Lien, and R. W. Furness. Use of underwater playback to reduce the impact of eiders on mussel farms. ICES Journal of Marine Science, 58(2):517–524, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/2/517/617864.

Rajasilta:2001:OWB

[RLH01] M. Rajasilta, P. Laine, and J. Hänninen. Ovarian weight of the Baltic herring (Clupea harengus membras) in relation to spawning time in the Archipelago Sea, northern Baltic. ICES Journal of Marine Science, 58(1):106–113, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/1/106/603495.

Reid:2001:RBH

[RM01] D. G. Reid and C. D. Maravelias. Relationships between herring school distribution and seabed substrate derived from RoxAnn. ICES Journal of Marine Science, 58(6):1161–1173, ???? 2001.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/6/1161/641537.

Rodriguez-Marin:2003:SBT

[RMAO^{+}03] E. Rodríguez-Marín, H. Arrizabalaga, M. Ortiz, C. Rodríguez-Cabello, G. Moreno, and L. T. Kell. Standardization of bluefin tuna, Thunnus thynnus, catch per unit effort in the baitboat fishery of the Bay of Biscay (Eastern Atlantic). ICES Journal of Marine Science, 60(6):1216–1231, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/6/1216/652519.

Russo:2009:PMH

[RMB^{+}09] Tommaso Russo, Stefano Mariani, Paolo Baldi, Antonio Parisi, Giuseppe Magnifico, Lotte Worsøe Clausen, and Stefano Cataudella. Progress in modelling herring populations: an individual-based model of growth. ICES Journal of Marine Science, 66(8):1718–1725, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/8/1718/677543.

Ragonese:2005:RRI

[RMDB05] Sergio Ragonese, Gaia Morizzo, Alberto De Santi, and Marco L. Bianchini. Rapid-response indicators of changes in resource state based on Mediterranean bottom-trawl surveys. ICES Journal of Marine Science, 62(3):511–515, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/3/511/664412.

Rhodes:2001:HAM

[RMKT01] L. L. Rhodes, A. L. Mackenzie, H. F. Kaspar, and K. E. Todd. Harmful algae and mariculture in New Zealand. ICES Journal of Marine Science, 58(2):398–403, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/2/398/617834.

Rideout:2006:VFS

[RML06] R. M. Rideout, M. J. Morgan, and G. R. Lilly. Variation in the frequency of skipped spawning in Atlantic cod (Gadus morhua) off Newfoundland and Labrador. ICES Journal of Marine Science, 63(6):1101–1110, ???? 2006. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/6/1101/616363.

Ryan:2005:GDB

[RMM05] Anthony W. Ryan, Valeria Mattiangeli, and Jarle Mork. Genetic differentiation of blue whiting (Micromesistius poutassou Risso) populations at the extremes of the species range and at the Hebrides–Porcupine Bank spawning grounds. ICES Journal of Marine Science, 62(5):948–955, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/5/948/857368.

Reiss:2005:CHV

[RNK05] Henning Reiss, Hermann Neumann, and Ingrid Kröncke. Chelaheight vs. body-weight relationships for North Sea hermit crabs (Paguridae). ICES Journal of Marine Science, 62(4):723–726, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/4/723/675122.

Richards:2008:PBM

[RNS08] R. Anne Richards, Paul C. Nitschke, and Katherine A. Sosebee. Population biology of monkfish Lophius americanus. ICES Journal of Marine Science, 65(7):1291–1305, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/7/1291/645826.

Roth:2008:APC

[RNWS08] Jennifer E. Roth, Nadav Nur, Pete Warzybok, and William J. Sydeman. Annual prey consumption of a dominant seabird, the common murre, in the California current system. ICES Journal of Marine Science, 65(6):1046–1056, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/6/1046/603636.

Rose:2002:CGC

[RO02] G. A. Rose and R. L. O’Driscoll. Capelin are good for cod: can the northern stock rebuild without them? ICES Journal of Marine Science, 59(5):1018–1026, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/5/1018/675129.

Rodionov:2005:ASR

[RO05] Sergei Rodionov and James E. Overland. Application of a sequential regime shift detection method to the Bering Sea ecosystem. ICES Journal of Marine Science, 62(3):328–332, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/3/328/658905.

Roel:2004:MOB

[ROB04] Beatriz A. Roel, Carl M. O’Brien, and Marinelle Basson. Management options for the Blackwater herring, a local springspawning stock in the Thames Estuary. ICES Journal of Marine Science, 61(3):297–307, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/61/3/297/670918.

Roberts:2005:CSL

[Rob05] Michael J. Roberts. Chokka squid (Loligo vulgaris reynaudii) abundance linked to changes in South Africa’s Agulhas Bank ecosystem during spawning and the early life cycle. ICES Journal of Marine Science, 62(1):33–55, ???? 2005. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/62/1/33/
667900.

Robinson:2008:MLS

[Rob08] Martin Robinson. Minimum landing size for Northeast Atlantic stocks of deep-water red crab, Chaceon affinis (Milne Edwards and Bouvier, 1894). ICES Journal of Marine Science, 65 (2):148–154, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/2/148/734408.

Rochet:2000:CAL

[Roc00a] Marie-Joëlle Rochet. A comparative approach to life-history strategies and tactics among four orders of teleost fish. ICES Journal of Marine Science, 57(2):228–239, April 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/2/228/620413.

Rochet:2000:DCS

[Roc00b] Marie-Joëlle Rochet. Does the concept of spawning per recruit make sense? ICES Journal of Marine Science, 57(4):1160–1174, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/1160/647290.

Rose:2003:MCN

[Ros03] George A. Rose. Monitoring coastal northern cod: towards an optimal survey of Smith Sound, Newfoundland. ICES Journal of Marine Science, 60(3):453–462, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/3/453/657946.

Rose:2005:CMV

[Ros05a] G. A. Rose. Capelin (Mallotus villosus) distribution and climate: a sea “canary” for marine ecosystem change. ICES Journal of Marine Science, 62(7):1524–1530, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/7/1524/661510.

Rose:2005:DRN

[Ros05b] G. A. Rose. On distributional responses of North Atlantic fish to climate change. ICES Journal of Marine Science, 62(7): 1360–1374, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/7/1360/657622.

Rose:2009:VTS

[Ros09] George A. Rose. Variations in the target strength of Atlantic cod during vertical migration. ICES Journal of Marine Science, 66(6):1205–1211, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1205/693702.

Rothschild:2000:FSR

[Rot00] Brian J. Rothschild. “Fish stocks and recruitment”: the past thirty years. ICES Journal of Marine Science, 57(2):191–201, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/2/191/620406.

Reeves:2007:ESB

[RP07] Stuart A. Reeves and Martin A. Pastoors. Evaluating the science behind the management advice for North Sea cod. ICES Journal of Marine Science, 64(4):671–678, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/671/639046.

Rademeyer:2007:TTD

[RPB07] Rebecca A. Rademeyer, Éva E. Plagányi, and Doug S. Butterworth. Tips and tricks in designing management procedures. ICES Journal of Marine Science, 64(4):618–625, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/4/618/642589.

Rochet:2008:ETE

[RPB^{+}08] Marie-Joëlle Rochet, Magali Prigent, Jacques A. Bertrand, André Carpentier, Franck Coppin, Jean-Paul Delpech, Guy Fontenelle, Eric Foucher, Kelig Mahé, Emilie Rostiaux, and Verena M. Trenkel. Ecosystem trends: evidence for agreement between fishers’ perceptions and scientific information. ICES Journal of Marine Science, 65(6):1057–1068, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/6/1057/601172.

Rudstam:2003:AST

[RPE^{+}03] L. G. Rudstam, S. L. Parker, D. W. Einhouse, L. D. Witzel, D. M. Warner, J. L. Stritzel, D. L. Parrish, and P. J. Sullivan. Application of in situ target-strength estimations in lakes: examples from rainbow-smelt surveys in Lakes Erie and Champlain. ICES Journal of Marine Science, 60(3):500–507, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/3/500/658346.

Rijnsdorp:2009:REC

[RPE^{+}09] Adriaan D. Rijnsdorp, Myron A. Peck, Georg H. Engelhard, Christian Möllmann, and John K. Pinnegar. Resolving the effect of climate change on fish populations. ICES Journal of Marine Science, 66(7):1570–1583, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/7/1570/657377.

Romakkaniemi:2003:DWA

[RPK^{+}03] A. Romakkaniemi, I. Perä, L. Karlsson, E. Jutila, U. Carlsson, and T. Pakarinen. Development of wild Atlantic salmon stocks in the rivers of the northern Baltic Sea in response to management measures. ICES Journal of Marine Science, 60(2):329–342, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/2/329/626466.

Royer:2002:SAE

[RPR02] J. Royer, P. Périès, and J. P. Robin. Stock assessments of English Channel loliginid squids: updated depletion methods and new analytical methods. ICES Journal of Marine Science, 59 (3):445–457, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/3/445/610813.

Rudstam:2009:TSO

[RPSSW09] Lars G. Rudstam, Sandra L. Parker-Stetter, Patrick J. Sullivan, and David M. Warner. Towards a standard operating procedure for fishery acoustic surveys in the Laurentian Great Lakes, North America. ICES Journal of Marine Science, 66(6):1391–1397, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/6/1391/690339.

Rochet:2002:ADF

[RPT02] Marie-Joëlle Rochet, Isabelle Péronnet, and Verena M. Trenkel. An analysis of discards from the French trawler fleet in the Celtic Sea. ICES Journal of Marine Science, 59(3):538–552, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/3/538/610830.

Robichaud:2002:AER

[RR02a] David Robichaud and George A. Rose. Assessing evacuation rates and spawning abundance of marine fishes using coupled telemetric and acoustic surveys. ICES Journal of Marine Science, 59(2):254–260, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/59/2/254/619633.

Robichaud:2002:RCT

[RR02b] David Robichaud and George A. Rose. The return of cod transplanted from a spawning ground in southern Newfoundland. ICES Journal of Marine Science, 59(6):1285–1293, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/6/1285/608192.

Rice:2005:FSS

[RR05a] Jake C. Rice and Marie-Joëlle Rochet. A framework for selecting a suite of indicators for fisheries management. ICES Journal of Marine Science, 62(3):516–527, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/3/516/664653.

Rochet:2005:DEC

[RR05b] Marie-Joëlle Rochet and Jake C. Rice. Do explicit criteria help in selecting indicators for ecosystem-based fisheries management? ICES Journal of Marine Science, 62(3):528–539, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/3/528/665311.

Robichaud:2006:DDD

[RR06] David Robichaud and George A. Rose. Density-dependent distribution of demersal juvenile Atlantic cod (Gadus morhua) in Placentia Bay, Newfoundland. ICES Journal of Marine Science, 63(4):766–774, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/4/766/694554.

Rice:2007:DRI

[RR07] Jake C. Rice and Denis Rivard. The dual role of indicators in optimal fisheries management strategies. ICES Journal of Marine Science, 64(4):775–778, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/775/641264.

Rochet:2009:SBM

[RR09] Marie-Joëlle Rochet and Jake C. Rice. Simulation-based management strategy evaluation: ignorance disguised as mathematics? ICES Journal of Marine Science, 66(4):754–762, May 2009.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/4/754/602618.

Rutherford:2003:ESC

[RRC03] Edward S. Rutherford, Kenneth A. Rose, and James H. Cowan, Jr. Evaluation of the Shepherd and Cushing (1980) model of density-dependent survival: a case study using striped bass (Morone saxatilis) larvae in the Potomac River, Maryland, USA. ICES Journal of Marine Science, 60(6):1275–1287, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/6/1275/653331.

Relini:2000:RFG

[RRT00] G. Relini, M. Relini, and G. Torchia. The role of fishing gear in the spreading of allochthonous species: the case of Caulerpa taxifolia in the Ligurian Sea. ICES Journal of Marine Science, 57(5):1421–1427, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/5/1421/660946.

Relini:2002:TRB

[RRTdA02] Giulio Relini, Marco Relini, Giovanni Torchia, and Giulia de Angelis. Trophic relationships between fishes and an artificial reef. ICES Journal of Marine Science, 59(S1):S36–S42, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S36/617999.

Relini:2002:TYC

[RRTP02] Giulio Relini, Marco Relini, Giovanni Torchia, and Giovanni Palandri. Ten years of censuses of fish fauna on the Loano artificial reef. ICES Journal of Marine Science, 59(S1):S132–S137, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL https://academic.oup.com/icesjms/
article/59/suppl/S132/617905.

Ryan:2008:DAS

[RRY08] Scott A. Ryan, John C. Roff, and Philip A. Yeats. Development and application of seasonal indices of coastal-zone eutrophication. ICES Journal of Marine Science, 65(8):1469–1474, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/8/1469/713637.

Ragnarsson:2003:SDO

[RS03] Stefán Áki Ragnarsson and Sigmar Arnar Steingrímsson. Spatial distribution of otter trawl effort in Icelandic waters: comparison of measures of effort and implications for benthic community effects of trawling activities. ICES Journal of Marine Science, 60(6):1200–1215, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/6/1200/652072.

Rahikainen:2004:CGV

[RS04] Mika Rahikainen and Robert L. Stephenson. Consequences of growth variation in northern Baltic herring for assessment and management. ICES Journal of Marine Science, 61(3):338–350, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/3/338/671185.

Rikardsen:2006:DSS

[RS06a] A. H. Rikardsen and S. Sandring. Diet and size-selective feeding by escaped hatchery rainbow trout Oncorhynchus mykiss (Walbaum). ICES Journal of Marine Science, 63(3):460–465, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/3/460/720686.

Rosenlund:2006:WSP

[RS06b] Grethe Rosenlund and Magnus Skretting. Worldwide status and perspective on gadoid culture. ICES Journal of Marine Science, 63(2):194–197, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/2/194/636976.

Rocklin:2009:CCC

[RSC^{+}09] Delphine Rocklin, Marie-Catherine Santoni, Jean-Michel Culioli, Jean-Antoine Tomasini, Dominique Pelletier, and David Mouillot. Changes in the catch composition of artisanal fisheries attributable to dolphin depredation in a Mediterranean marine reserve. ICES Journal of Marine Science, 66(4):699–707, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/4/699/604061.

Rossi-Santos:2008:ODH

[RSNB^{+}08] Marcos R. Rossi-Santos, Elitieri S. Neto, Clarêncio G. Baracho, Sérgio R. Cipolotti, Enrico Marcovaldi, and Marcia H. Engel. Occurrence and distribution of humpback whales (Megaptera novaeangliae) on the north coast of the State of Bahia, Brazil, 2000–2006. ICES Journal of Marine Science, 65(4):667–673, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/4/667/637918.

Rottingen:2003:EAL

[RT03] Ingolf Røttingen and Sigurd Tjelmeland. Evaluation of the absolute levels of acoustic estimates of the 1983 year class of Norwegian spring-spawning herring. ICES Journal of Marine Science, 60(3):480–485, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/3/480/658068.

Rochet:2005:CIT

[RTB^{+}05] Marie-Joëlle Rochet, Verena Trenkel, Robert Bellail, Franck Coppin, Olivier Le Pape, Jean-Claude Mahé, Jocelyne Morin, Jean-Charles Poulard, Ivan Schlaich, Arnauld Souplet, Yves Vérin, and Jacques Bertrand. Combining indicator trends to assess ongoing changes in exploited fish communities: diagnostic of communities off the coasts of France. ICES Journal of Marine Science, 62(8):1647–1664, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/8/1647/793215.

Rabalais:2009:GCE

[RTDJ09] Nancy N. Rabalais, R. Eugene Turner, Robert J. Díaz, and Dubravko Justić. Global change and eutrophication of coastal waters. ICES Journal of Marine Science, 66(7):1528–1537, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/7/1528/656749.

Roa-Ureta:2007:STS

[RUA07] Rubén Roa-Ureta and Alexander I. Arkhipkin. Short-term stock assessment of Loligo gahi at the Falkland Islands: sequential use of stochastic biomass projection and stock depletion models. ICES Journal of Marine Science, 64(1):3–17, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/1/3/648750.

Rotherham:2007:SDS

[RUCG07] D. Rotherham, A. J. Underwood, M. G. Chapman, and C. A. Gray. A strategy for developing scientific sampling tools for fishery-independent surveys of estuarine fish in New South Wales, Australia. ICES Journal of Marine Science, 64(8):1512– 1516, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/8/1512/612804.

Roa-Ureta:2007:BES

[RUN07] Rubén Roa-Ureta and Edwin Niklitschek. Biomass estimation from surveys with likelihood-based geostatistics. ICES Journal of Marine Science, 64(9):1723–1734, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/9/1723/783804.

Rijnsdorp:2000:CIA

[RvMBV00] A. D. Rijnsdorp, P. L. van Mourik Broekman, and E. G. Visser. Competitive interactions among beam trawlers exploiting local patches of flatfish in the North Sea. ICES Journal of Marine Science, 57(4):894–902, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/4/894/647364.

Rivoirard:2001:CED

[RW01] Jacques Rivoirard and Kai Wieland. Correcting for the effect of daylight in abundance estimation of juvenile haddock (Melanogrammus aeglefinus) in the North Sea: an application of kriging with external drift. ICES Journal of Marine Science, 58 (6):1272–1285, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/6/1272/641572.

Ryer:2004:LEB

[Rye04] Clifford H. Ryer. Laboratory evidence for behavioural impairment of fish escaping trawls: a review. ICES Journal of Marine Science, 61(7):1157–1164, ???? 2004. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/7/1157/881030.

Smolyar:2003:QDB

[SA03] I. Smolyar and N. Adrov. The quantitative definition of the Barents Sea Atlantic water: mapping of the annual climatic cycle and interannual variability. ICES Journal of Marine Science, 60(4):836–845, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/4/836/693651.

Santiago:2005:IGS

[SA05] J. Santiago and H. Arrizabalaga. An integrated growth study for North Atlantic albacore (Thunnus alalunga Bonn. 1788). ICES Journal of Marine Science, 62(4):740–749, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/4/740/675421.

Sanvicente-Anorve:2007:RBF

[SAAFCA07] L. Sanvicente-Añorve, M. A. Alatorre, C. Flores-Coto, and C. Alba. Relationships between fish larvae and siphonophores in the water column: effect of wind-induced turbulence and thermocline depth. ICES Journal of Marine Science, 64(5):878–888, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/5/878/640762.

Sabates:2004:DVD

[Sab04] A. Sabatés. Diel vertical distribution of fish larvae during the winter-mixing period in the Northwestern Mediterranean. ICES Journal of Marine Science, 61(8):1243–1252, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/8/1243/629260.

Sato:2009:SOS

[SAM09] Noriyosi Sato, Satoshi Awata, and Hiroyuki Munehara. Seasonal occurrence and sexual maturation of Japanese pygmy squid (Idiosepius paradoxus) at the northern limits of their distribution. ICES Journal of Marine Science, 66(5):811–815, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/5/811/665303.

Stensholt:2002:VDD

[SAMS02] Boonchai K. Stensholt, Asgeir Aglen, Sigbjørn Mehl, and Eivind Stensholt. Vertical density distributions of fish: a balance between environmental and physiological limitation. ICES Journal of Marine Science, 59(4):679–710, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/4/679/676771.

Skaret:2005:BSH

[SAN^{+}05] Georg Skaret, Bjørn Erik Axelsen, Leif Nøttestad, Anders Fernö, and Arne Johannessen. The behaviour of spawning herring in relation to a survey vessel. ICES Journal of Marine Science, 62 (6):1061–1064, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/6/1061/616263.

Simonini:2004:OES

[SAPP04] R. Simonini, I. Ansaloni, A. M. Bonvicini Pagliai, and D. Prevedelli. Organic enrichment and structure of the macrozoobenthic community in the northern Adriatic Sea in an area facing adige and po mouths. ICES Journal of Marine Science, 61(6):871–881, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/6/871/677041.

Sarafanov:2009:ENA

[Sar09] Artem Sarafanov. On the effect of the North Atlantic Oscillation on temperature and salinity of the subpolar North Atlantic intermediate and deep waters. ICES Journal of Marine Science, 66(7):1448–1454, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/7/1448/658172.

Shiganova:2000:EGP

[SB00a] Tamara A. Shiganova and Yulia V. Bulgakova. Effects of gelatinous plankton on Black Sea and Sea of Azov fish and their food resources. ICES Journal of Marine Science, 57(3):641–648, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/3/641/635948.

Skjaeraasen:2000:DFE

[SB00b] J. E. Skjæraasen and O. A. Bergstad. Distribution and feeding ecology of Raja radiata in the northeastern North Sea and Skagerrak (Norwegian Deep). ICES Journal of Marine Science, 57(4):1249–1260, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/4/1249/647316.

Skjaeraasen:2001:NDL

[SB01] J. E. Skjæraasen and O. A. Bergstad. Notes on the distribution and length composition of Raja lintea, R. fyllae, R. hyperborea and Bathyraja spinicauda (Pisces: Rajidae) in the deep northeastern North Sea and on the slope of the eastern Norwegian Sea. ICES Journal of Marine Science, 58(1):21–28, ???? 2001. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/58/1/21/
603519.

Svedang:2003:STA

[SB03] Henrik Svedäng and Gwenola Bardon. Spatial and temporal aspects of the decline in cod (Gadus morhua L.) abundance in the Kattegat and eastern Skagerrak. ICES Journal of Marine Science, 60(1):32–37, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/60/1/32/611439.

Smolyar:2004:DMF

[SB04] Igor V. Smolyar and Timothy G. Bromage. Discrete model of fish scale incremental pattern: a formalization of the 2D anisotropic structure. ICES Journal of Marine Science, 61 (6):992–1003, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/6/992/679363.

Salvanes:2006:NUB

[SB06] Anne Gro Vea Salvanes and Victoria Braithwaite. The need to understand the behaviour of fish reared for mariculture or restocking. ICES Journal of Marine Science, 63(2):345–354, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/2/345/641309.

Skreslet:2005:SEU

[SBB^{+}05] Stig Skreslet, Angel Borja, Luca Bugliaro, Georg Hansen, Ralf Meerkötter, Ketil Olsen, and Jean Verdebout. Some effects of ultraviolet radiation and climate on the reproduction of Calanus finmarchicus (Copepoda) and year class formation in arcto-Norwegian cod (Gadus morhua). ICES Journal of Marine Science, 62(7):1293–1300, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/7/1293/656497.

Schabetsberger:2000:DVM

[SBC^{+}00] R. Schabetsberger, R. D. Brodeur, L. Ciannelli, J. M. Napp, and G. L. Swartzman. Diel vertical migration and interaction of zooplankton and juvenile walleye pollock (Theragra chalcogramma) at a frontal region near the Pribilof Islands, Bering Sea. ICES Journal of Marine Science, 57(4):1283–1295, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/1283/647325.

Schwach:2007:PKF

[SBC^{+}07] Vera Schwach, Denis Bailly, Anne-Sofie Christensen, Alyne E. Delaney, Poul Degnbol, Wim L. T. van Densen, Petter Holm, H. Anne McLay, Kåre Nolde Nielsen, Martin A. Pastoors, Stuart A. Reeves, and Douglas C. Wilson. Policy and knowledge in fisheries management: a policy brief. ICES Journal of Marine Science, 64(4):798–803, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/798/641022.

Sweeting:2009:SBS

[SBD^{+}09] Christopher J. Sweeting, Fabio Badalamenti, Giovanni D’Anna, Carlo Pipitone, and N. V. C. Polunin. Steeper biomass spectra of demersal fish communities after trawler exclusion in Sicily. ICES Journal of Marine Science, 66(1):195–202, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/1/195/634936.

Stevens:2000:EFS

[SBDW00] J. D. Stevens, R. Bonfil, N. K. Dulvy, and P. A. Walker. The effects of fishing on sharks, rays, and chimaeras (chondrichthyans), and the implications for marine ecosystems. ICES Journal of Marine Science, 57(3):476–494, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/3/476/635915.

Saillant:2006:GIS

[SBG06] Eric Saillant, S. Coleen Bradfield, and John R. Gold. Genetic impacts of shrimp trawling on red snapper (Lutjanus campechanus) in the northern Gulf of Mexico. ICES Journal of Marine Science, 63(4):705–713, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/4/705/693875.

Sparholt:2007:MAS

[SBL07] Henrik Sparholt, Mette Bertelsen, and Hans Lassen. A metaanalysis of the status of ICES fish stocks during the past half century. ICES Journal of Marine Science, 64(4):707–713, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/4/707/641682.

Smith:2007:IQE

[SBP07] C. J. Smith, A. C. Banks, and K-N. Papadopoulou. Improving the quantitative estimation of trawling impacts from sidescansonar and underwater-video imagery. ICES Journal of Marine Science, 64(9):1692–1701, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/9/1692/788942.

Skogen:2007:IVN

[SBR07] Morten D. Skogen, W. Paul Budgell, and Francisco Rey. Interannual variability in Nordic seas primary production. ICES Journal of Marine Science, 64(5):889–898, July 2007. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/5/889/641227.

Smith:2009:ICF

[SBT^{+}09] Stephen J. Smith, Jerry Black, Brian J. Todd, Vladimir E. Kostylev, and Mark J. Lundy. The impact of commercial fishing on the determination of habitat associations for sea scallops (Placopecten magellanicus, Gmelin). ICES Journal of Marine Science, 66(9):2043–2051, October 2009. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/9/2043/726717.

Stanton:2000:RRM

[SC00] Timothy K. Stanton and Dezhang Chu. Review and recommendations for the modelling of acoustic scattering by fluid-like elongated zooplankton: euphausiids and copepods. ICES Journal of Marine Science, 57(4):793–807, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/4/793/647331.

Sirovic:2009:RST

[ŠCBD09] Ana Širović, George R. Cutter, John L. Butler, and David A. Demer. Rockfish sounds and their potential use for population monitoring in the Southern California Bight. ICES Journal of Marine Science, 66(6):981–990, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/981/692029.

Saura:2006:GVR

[SCCM06] María Saura, Pablo Caballero, Armando Caballero, and Paloma Morán. Genetic variation in restored Atlantic salmon (Salmo salar L.) populations in the Ulla and Lérez rivers, Galicia, Spain. ICES Journal of Marine Science, 63(7):1290–1296, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/7/1290/757660.

Shephard:2007:ESC

[SCHR07] Samuel Shephard, Paul Connolly, Nils-Roar Hareide, and Emer Rogan. Establishing stakeholder connections for management of the Irish orange roughy fishery. ICES Journal of Marine Science, 64(4):841–845, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/4/841/638752.

Shannon:2000:MEF

[SCJ00] Lynne J. Shannon, Philippe M. Cury, and Astrid Jarre. Modelling effects of fishing in the Southern Benguela ecosystem. ICES Journal of Marine Science, 57(3):720–722, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/3/720/635963.

Schultz:2000:EAD

[SCLG00] Eric T. Schultz, Robert K. Cowen, Kamazima M. M. Lwiza, and Angela M. Gospodarek. Explaining advection: do larval bay anchovy (Anchoa mitchilli) show selective tidal-stream transport? ICES Journal of Marine Science, 57(2):360–371, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/2/360/620453.

StJohn:2001:DPM

[SCLK01] M. A. St. John, C. Clemmesen, T. Lund, and T. Köster. Diatom production in the marine environment: implications for larval fish growth and condition. ICES Journal of Marine Science, 58 (5):1106–1113, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/5/1106/730846.

Somerfield:2008:AFD

[SCWD08] Paul J. Somerfield, K. Robert Clarke, Richard M. Warwick, and Nick K. Dulvy. Average functional distinctness as a measure of the composition of assemblages. ICES Journal of Marine Science, 65(8):1462–1468, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/8/1462/713228.

Segers:2007:PSN

[SDCR07] F. H. I. D. Segers, M. Dickey-Collas, and A. D. Rijnsdorp. Prey selection by North Sea herring (Clupea harengus), with special reference to fish eggs. ICES Journal of Marine Science, 64 (1):60–68, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/1/60/645806.

Santos:2001:SHM

[SdFBG01] A. Miguel P. Santos, Maria de Fátima Borges, and Steve Groom. Sardine and horse mackerel recruitment and upwelling off Portugal. ICES Journal of Marine Science, 58(3):589–596, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/3/589/810087.

Soualili:2008:ASP

[SDG^{+}08] Dina Soualili, Philippe Dubois, Pol Gosselin, Philippe Pernet, and Monique Guillou. Assessment of seawater pollution by heavy metals in the neighbourhood of Algiers: use of the sea urchin, Paracentrotus lividus, as a bioindicator. ICES Journal of Marine Science, 65(2):132–139, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/2/132/733408.

Scenna:2006:TEPa

[SdlRdA06a] L. B. Scenna, S. B. García de la Rosa, and J. M. Díaz de Astarloa. Trophic ecology of the Patagonian skate, Bathyraja macloviana, on the Argentine continental shelf. ICES Journal of Marine Science, 63(5):867–874, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/5/867/663048.

Scenna:2006:TEPb

[SdlRdA06b] L. B. Scenna, S. B. García de la Rosa, and J. M. Díaz de Astarloa. Trophic ecology of the Patagonian skate, Bathyraja macloviana, on the Argentine continental shelf. ICES Journal of Marine Science, 63(9):1765, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/9/1765/701755.

Saglam:2009:REI

[SDÖ09] Hacer Sağlam, Ertuğ Düzgüneş, and Hamdi Öğüt. Reproductive ecology of the invasive whelk Rapana venosa Valenciennes, 1846, in the southeastern Black Sea (Gastropoda: Muricidae). ICES Journal of Marine Science, 66(9):1865–1867, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/9/1865/726645.

Sanchez:2000:IOT

[SDRK00] P. Sanchez, M. Demestre, M. Ramon, and M. J. Kaiser. The impact of otter trawling on mud communities in the northwestern Mediterranean. ICES Journal of Marine Science, 57(5): 1352–1358, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/5/1352/660933.

Shen:2009:SPE

[SDWQ09] Haixue Shen, Martin W. Dorn, Vidar Wespestad, and Terrance J. Quinn. Schooling pattern of eastern Bering Sea walleye pollock and its effect on fishing behaviour. ICES Journal of Marine Science, 66(6):1284–1288, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1284/692357.

Stram:2009:FMR

[SE09] Diana L. Stram and Diana C. K. Evans. Fishery management responses to climate change in the North Pacific. ICES Journal of Marine Science, 66(7):1633–1639, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/7/1633/658823.

Seaman:2002:UTO

[Sea02] William Seaman, Jr. Unifying trends and opportunities in global artificial reef research, including evaluation. ICES Journal of Marine Science, 59(S1):S14–S16, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/59/suppl/S14/617908.

Secor:2000:LRC

[Sec00a] D. H. Secor. Longevity and resilience of Chesapeake Bay striped bass. ICES Journal of Marine Science, 57(4):808–815, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/808/647335.

Secor:2000:SNT

[Sec00b] David H. Secor. Spawning in the nick of time? Effect of adult demographics on spawning behaviour and recruitment in Chesapeake Bay striped bass. ICES Journal of Marine Science, 57 (2):403–411, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/2/403/620472.

Secor:2002:ILC

[Sec02] D. Secor. Inter-laboratory comparison of Atlantic and Mediterranean bluefin tuna otolith microconstituents. ICES Journal of Marine Science, 59(6):1294–1304, ???? 2002. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/6/1294/608194.

Smith:2009:DBP

[SEOR09] Robert J. Smith, Paul D. Eastwood, Yoshitaka Ota, and Stuart I. Rogers. Developing best practice for using marxan to locate marine protected areas in European waters. ICES Journal of Marine Science, 66(1):188–194, January 2009. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/1/188/634672.

Seytre:2009:CRM

[SF09] Catherine Seytre and Patrice Francour. The Cap Roux MPA (Saint–Raphaël, French Mediterranean): changes in fish assemblages within four years of protection. ICES Journal of Marine Science, 66(1):180–187, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/1/180/634642.

Smith:2007:STS

[SFH^{+}07] A. D. M. Smith, E. J. Fulton, A. J. Hobday, D. C. Smith, and P. Shoulder. Scientific tools to support the practical implementation of ecosystem-based fisheries management. ICES Journal of Marine Science, 64(4):633–639, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/633/641888.

Steimle:2002:BMP

[SFKC02] F. Steimle, K. Foster, R. Kropp, and B. Conlin. Benthic macrofauna productivity enhancement by an artificial reef in Delaware Bay, USA. ICES Journal of Marine Science, 59(S1):S100–S105, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL https://academic.oup.com/icesjms/
article/59/suppl/S100/617895.

Swain:2001:DSA

[SFM01] D. P. Swain, K. T. Frank, and G. Maillet. Delineating stocks of Atlantic cod (Gadus morhua) in the Gulf of St Lawrence and Cabot Strait areas using vertebral number. ICES Journal of Marine Science, 58(1):253–269, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/1/253/603526.

Skaug:2007:HPA

[SFØ07] Hans J. Skaug, Lennart Frimannslund, and Nils I. Øien. Historical population assessment of Barents Sea harp seals (Pagophilus groenlandicus). ICES Journal of Marine Science, 64(7):1356– 1365, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/7/1356/727990.

Sanchez:2000:HMS

[SG00] Francisco Sánchez and Julio Gil. Hydrographic mesoscale structures and poleward current as a determinant of hake (Merluccius merluccius) recruitment in southern Bay of Biscay. ICES Journal of Marine Science, 57(1):152–170, February 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/1/152/641134.

Steingrund:2005:RBP

[SG05] Petur Steingrund and Eilif Gaard. Relationship between phytoplankton production and cod production on the Faroe Shelf. ICES Journal of Marine Science, 62(2):163–176, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/2/163/602260.

Sumaila:2000:AEE

[SGAC00] Ussif Rashid Sumaila, Sylvie Guénette, Jackie Alder, and Ratana Chuenpagdee. Addressing ecosystem effects of fishing using marine protected areas. ICES Journal of Marine Science, 57(3):752–760, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/752/636007.

Simmonds:2009:ODA

[SGC^{+}09] E. John Simmonds, Mariano Gutiérrez, Andres Chipollini, Francois Gerlotto, Mathieu Woillez, and Arnaud Bertrand. Optimizing the design of acoustic surveys of Peruvian anchoveta. ICES Journal of Marine Science, 66(6):1341–1348, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/1341/696051.

Schirripa:2009:TDM

[SGM09] Michael J. Schirripa, C. Phillip Goodyear, and Richard M. Methot, Jr. Testing different methods of incorporating climate data into the assessment of US West Coast sablefish. ICES Journal of Marine Science, 66(7):1605–1613, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/7/1605/656457.

Sousa:2009:ANE

[SGMMGB09] Arturo Sousa, Pablo García-Murillo, Julia Morales, and Leoncio García-Barrón. Anthropogenic and natural effects on the coastal lagoons in the southwest of Spain (Doñana National Park). ICES Journal of Marine Science, 66(7):1508–1514, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/7/1508/658285.

Swan:2006:OCA

[SGMN^{+}06] S. C. Swan, A. J. Geffen, B. Morales-Nin, J. D. M. Gordon, T. Shimmield, T. Sawyer, and E. Massutí. Otolith chemistry: an aid to stock separation of Helicolenus dactylopterus (bluemouth) and Merluccius merluccius (European hake) in the Northeast Atlantic and Mediterranean. ICES Journal of Marine Science, 63(3):504–513, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/3/504/722042.

Sanz:2008:SGE

[SGMV^{+}08] Núria Sanz, José-Luis García-Marín, Jordi Viñas, Marina Roldán, and Carles Pla. Spawning groups of European anchovy: population structure and management implications. ICES Journal of Marine Science, 65(9):1635–1644, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/9/1635/628901.

Sherman:2002:ARD

[SGS02] R. L. Sherman, D. S. Gilliam, and R. E. Spieler. Artificial reef design: void space, complexity, and attractants. ICES Journal of Marine Science, 59(S1):S196–S200, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S196/617933.

Stransky:2005:ADG

[SGS^{+}05] Christoph Stransky, Sif Gudmundsdóttir, Thorsteinn Sigurdsson, Svend Lemvig, Kjell Nedreaas, and Fran Saborido-Rey. Age determination and growth of Atlantic redfish (Sebastes marinus and S. mentella): bias and precision of age readers and otolith preparation methods. ICES Journal of Marine Science, 62(4):655–670, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/4/655/673428.

Smith:2008:RCA

[SGY08] Melissa D. Smith, Jonathan H. Grabowski, and Philip O. Yund. The role of closed areas in rebuilding monkfish populations in the Gulf of Maine. ICES Journal of Marine Science, 65(7): 1326–1333, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/7/1326/646852.

Schnute:2006:RPM

[SH06] Jon T. Schnute and Rowan Haigh. Reference points and management strategies: lessons from quantum mechanics. ICES Journal of Marine Science, 63(1):4–11, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/1/4/625085.

Schnute:2007:CAC

[SH07] Jon T. Schnute and Rowan Haigh. Compositional analysis of catch curve data, with an application to Sebastes maliger. ICES Journal of Marine Science, 64(2):218–233, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/2/218/2182266.

Skilbrei:2009:VME

[SHAH09] Ove T. Skilbrei, Jens Christian Holst, Lars Asplin, and Marianne Holm. Vertical movements of “escaped” farmed Atlantic salmon (Salmo salar L.) — a simulation study in a western Norwegian fjord. ICES Journal of Marine Science, 66(2):278–288, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/2/278/595287.

Smith:2004:PBE

[SHdLP04] Kim D. Smith, Norman G. Hall, Simon de Lestang, and Ian C. Potter. Potential bias in estimates of the size of maturity of crabs derived from trap samples. ICES Journal of Marine Science, 61 (6):906–912, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/6/906/677685.

Shelton:2005:DRC

[She05] Peter A. Shelton. Did over-reliance on commercial catch rate data precipitate the collapse of northern cod? ICES Journal of Marine Science, 62(6):1139–1149, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/6/1139/617793.

Shelton:2007:WRS

[She07] Peter A. Shelton. The weakening role of science in the management of groundfish off the east coast of Canada. ICES Journal of Marine Science, 64(4):723–729, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/723/639866.

Smith:2001:SDA

[SHS01] P. E. Smith, J. K. Horne, and D. C. Schneider. Spatial dynamics of anchovy, sardine, and hake pre-recruit stages in the California current. ICES Journal of Marine Science, 58(5):1063–1071, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/5/1063/730827.

St-Hilaire:2001:SBU

[SHSKR01] S. St-Hilaire, C. Stephen, M. Kent, and C. Ribble. Sentinels in the bay: using farm fish to monitor for pathogens in the environment. ICES Journal of Marine Science, 58(2):369–373, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/2/369/617825.

Schweigert:2009:RFU

[SHT^{+}09] Jacob F. Schweigert, Douglas E. Hay, Thomas W. Therriault, Matthew Thompson, and Carl W. Haegele. Recruitment forecasting using indices of young-of-the-year Pacific herring (Clupea pallasi) abundance in the Strait of Georgia, (BC). ICES Journal of Marine Science, 66(8):1681–1687, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/8/1681/676357.

Silva:2003:MVA

[Sil03] Alexandra Silva. Morphometric variation among sardine (Sardina pilchardus) populations from the northeastern Atlantic and the western Mediterranean. ICES Journal of Marine Science, 60 (6):1352–1360, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/6/1352/654334.

Simmonds:2003:WAT

[Sim03] E. J. Simmonds. Weighting of acoustic- and trawl-survey indices for the assessment of North Sea herring. ICES Journal of Marine Science, 60(3):463–471, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/3/463/657966.

Simmonds:2007:CTP

[Sim07a] E. John Simmonds. Comparison of two periods of North Sea herring stock management: success, failure, and monetary value. ICES Journal of Marine Science, 64(4):686–692, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/4/686/642385.

Simon:2007:AGC

[Sim07b] Janek Simon. Age, growth, and condition of European eel (Anguilla anguilla) from six lakes in the River Havel system (Germany). ICES Journal of Marine Science, 64(7):1414–1422, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/7/1414/727065.

Simmonds:2009:EQN

[Sim09] E. John Simmonds. Evaluation of the quality of the North Sea herring assessment. ICES Journal of Marine Science, 66(8): 1814–1822, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/8/1814/673783.

Sinclair:2001:NMC

[Sin01] A. F. Sinclair. Natural mortality of cod (Gadus morhua) in the Southern Gulf of st Lawrence. ICES Journal of Marine Science, 58(1):1–10, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/58/1/1/603492.

Sinclair:2009:HIH

[Sin09] Mike Sinclair. Herring and ICES: a historical sketch of a few ideas and their linkages. ICES Journal of Marine Science, 66 (8):1652–1661, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/8/1652/673842.

Stobberup:2005:ASS

[SIT^{+}05] Kim A. Stobberup, Cheikh A. O. Inejih, Sory Traoré, Carlos Monteiro, Patrícia Amorim, and Karim Erzini. Analysis of size spectra off northwest Africa: a useful indicator in tropical areas? ICES Journal of Marine Science, 62(3):424–429, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/3/424/661801.

Smith:2008:DCA

[SJ08] I. Philip Smith and Antony C. Jensen. Dynamics of closed areas in Norway lobster fisheries. ICES Journal of Marine Science, 65 (9):1600–1609, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/9/1600/631744.

Sanchez-Jerez:2002:EAR

[SJGRRRE02] P. Sánchez-Jerez, B. M. Gillanders, S. Rodríguez-Ruiz, and A. A. Ramos-Esplá. Effect of an artificial reef in Posidonia meadows on fish assemblage and diet of Diplodus annularis. ICES Journal of Marine Science, 59(S1):S59–S68, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S59/618004.

Saloniemi:2004:SRW

[SJKN^{+}04] I. Saloniemi, E. Jokikokko, I. Kallio-Nyberg, E. Jutila, and P. Pasanen. Survival of reared and wild Atlantic salmon smolts: size matters more in bad years. ICES Journal of Marine Science, 61(5):782–787, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/5/782/865891.

Stobutzki:2003:CFB

[SJM03] Ilona Stobutzki, Peter Jones, and Margaret Miller. A comparison of fish bycatch communities between areas open and closed to prawn trawling in an Australian tropical fishery. ICES Journal of Marine Science, 60(5):951–966, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/5/951/769561.

Shirakihara:2004:EMR

[SK04] Kunio Shirakihara and Shuichi Kitada. Estimating migration rates from two tag–release/one recovery experiments. ICES Journal of Marine Science, 61(5):821–828, ???? 2004. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/5/821/867726.

Simmonds:2007:MIO

[SK07] E. John Simmonds and Stephen Keltz. Management implications and options for a stock with unstable or uncertain dynamics: west of Scotland herring. ICES Journal of Marine Science, 64(4):679–685, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/4/679/642436.

Skagen:2007:MSR

[Ska07] Dankert W. Skagen. Management strategies for reducing variation in annual yield: when can they work? ICES Journal of Marine Science, 64(4):698–701, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/4/698/641076.

Stanley:2000:EWR

[SKC^{+}00] R. D. Stanley, R. Kieser, K. Cooke, A. M. Surry, and B. Mose. Estimation of a widow rockfish (Sebastes entomelas) shoal off British Columbia, Canada as a joint exercise between stock assessment staff and the fishing industry. ICES Journal of Marine Science, 57(4):1035–1049, August 2000. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/4/1035/647248.

Secor:2009:CEP

[SKC09] David H. Secor, Lisa A. Kerr, and Steven X. Cadrin. Connectivity effects on productivity, stability, and persistence in a herring metapopulation model. ICES Journal of Marine Science, 66(8): 1726–1732, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/8/1726/675636.

Stanley:2002:TDV

[SKH02] R. D. Stanley, R. Kieser, and M. Hajirakar. Three-dimensional visualization of a widow rockfish (Sebastes entomelas) shoal over interpolated bathymetry. ICES Journal of Marine Science, 59 (1):151–155, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/1/151/649989.

Sigurethsson:2006:FPR

[SKR^{+}06] Th. Sigurdsson, K. Kristinsson, H-J. Rätz, K. H. Nedreaas, S. P. Melnikov, and J. Reinert. The fishery for pelagic redfish (Sebastes mentella) in the Irminger Sea and adjacent waters. ICES Journal of Marine Science, 63(4):725–736, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/4/725/694308.

Sakurai:2000:CIS

[SKS^{+}00] Y. Sakurai, H. Kiyofuji, S. Saitoh, T. Goto, and Y. Hiyama. Changes in inferred spawning areas of Todarodes pacificus (Cephalopoda: Ommastrephidae) due to changing environmental conditions. ICES Journal of Marine Science, 57(1):24–30, February 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/1/24/641140.

Seitz:2001:VTB

[SL01] R. D. Seitz and R. N. Lipcius. Variation in top-down and bottom-up control of marine bivalves at differing spatial scales. ICES Journal of Marine Science, 58(3):689–699, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/3/689/810129.

Sheridan:2004:YIT

[SL04] Cecelia C. Sheridan and Michael R. Landry. A 9-year increasing trend in mesozooplankton biomass at the Hawaii Ocean Timeseries Station ALOHA. ICES Journal of Marine Science, 61 (4):457–463, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/4/457/602294.

Sosa-Lopez:2005:EIB

[SLMCRM05] Atahualpa Sosa-López, David Mouillot, Thang Do Chi, and Julia Ramos-Miranda. Ecological indicators based on fish biomass distribution along trophic levels: an application to the terminos coastal lagoon, Mexico. ICES Journal of Marine Science, 62(3):453–458, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/3/453/662697.

Sparholt:2002:NPN

[SLN02a] Henrik Sparholt, Lena I. Larsen, and J. Rasmus Nielsen. Nonpredation natural mortality of Norway pout (Trisopterus esmarkii) in the North Sea. ICES Journal of Marine Science, 59 (6):1276–1284, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/6/1276/608191.

Sparholt:2002:VMI

[SLN02b] Henrik Sparholt, Lena I. Larsen, and J. Rasmus Nielsen. Verification of multispecies interactions in the North Sea by trawl survey data on Norway pout (Trisopterus esmarkii). ICES Journal of Marine Science, 59(6):1270–1275, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/6/1270/608188.

Schipper:2009:RAE

[SLvdB^{+}09] Cor A. Schipper, Joost Lahr, Paul J. van den Brink, Steve G. George, Peter-Diedrich Hansen, Helena C. da Silva de Assis, Ron van der Oost, John E. Thain, Dave Livingstone, Carys Mitchelmore, Frederik-Jan van Schooten, Freek Ariese, Albertinka J. Murk, Guy C. M. Grinwis, Hans Klamer, Belinda J. Kater, Jaap F. Postma, Bert van der Werf, and A. Dick Vethaak. A retrospective analysis to explore the applicability of fish biomarkers and sediment bioassays along contaminated salinity transects. ICES Journal of Marine Science, 66(10):2089–2105, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/10/2089/680732.

Stratoudakis:2002:SSD

[SM02] Yorgos Stratoudakis and Ana Marçalo. Sardine slipping during purse-seining off northern Portugal. ICES Journal of Marine Science, 59(6):1256–1262, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/6/1256/608185.

Scalabrin:2009:HMF

[SMB09] Carla Scalabrin, Christian Marfia, and Jean Boucher. How much fish is hidden in the surface and bottom acoustic blind zones? ICES Journal of Marine Science, 66(6):1355–1363, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/6/1355/697379.

Smith:2001:TYR

[SMEK01] Aaron C. Smith, Jennifer L. Martin, James M. Ehrman, and Irena Kaczmarska. Ten-year record of Thalassiosira nordenskioeldii population dynamics: comparison of aquaculture and non-aquaculture sites in the quoddy region. ICES Journal of Marine Science, 58(2):391–397, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/2/391/617831.

Santos:2002:DVF

[SMG02] M. N. Santos, C. C. Monteiro, and M. B. Gaspar. Diurnal variations in the fish assemblage at an artificial reef. ICES Journal of Marine Science, 59(S1):S32–S35, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/59/suppl/S32/617982.

Saunders:2009:USM

[SMH09] T. Saunders, S. Mayfield, and A. Hogg. Using a simple morphometric marker to identify spatial units for abalone fishery management. ICES Journal of Marine Science, 66(2):305–314, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/2/305/595072.

Schnute:2007:DTE

[SMI07] Jon T. Schnute, Mark N. Maunder, and James N. Ianelli. Designing tools to evaluate fishery management strategies: can the scientific community deliver? ICES Journal of Marine Science, 64(6):1077–1084, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/6/1077/616074.

Skjaeraasen:2008:ESS

[SMK08] Jon Egil Skjæraasen, Justin J. Meager, and Ørjan Karlsen. The expression of secondary sexual characteristics in recruit- and repeat-spawning farmed and wild Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 65(9):1710–1716, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/9/1710/630396.

Sutherland:2001:CSP

[SML01] T. F. Sutherland, A. J. Martin, and C. D. Levings. Characterization of suspended particulate matter surrounding a salmonid net-pen in the Broughton Archipelago, British Columbia. ICES Journal of Marine Science, 58(2):404–410, ???? 2001. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/2/404/617837.

Stephenson:2009:PIC

[SMP09] Robert L. Stephenson, Gary D. Melvin, and Michael J. Power. Population integrity and connectivity in Northwest Atlantic herring: a review of assumptions and evidence. ICES Journal of Marine Science, 66(8):1733–1739, September 2009. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/8/1733/677110.

Sundby:2008:SSS

[SN08] Svein Sundby and Odd Nakken. Spatial shifts in spawning habitats of arcto-Norwegian cod related to multidecadal climate oscillations and climate change. ICES Journal of Marine Science, 65(6):953–962, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/6/953/604353.

Stentiford:2001:RHI

[SNA01] G. D. Stentiford, D. M. Neil, and R. J. A. Atkinson. The relationship of Hematodinium infection prevalence in a Scottish Nephrops norvegicus population to season, moulting and sex. ICES Journal of Marine Science, 58(4):814–823, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/4/814/630208.

Swartzman:2002:SPP

[SNB^{+}02] Gordon Swartzman, Jeffrey Napp, Richard Brodeur, Andreas Winter, and Lorenzo Ciannelli. Spatial patterns of pollock and zooplankton distribution in the Pribilof Islands, Alaska nursery area and their relationship to pollock recruitment. ICES Journal of Marine Science, 59(6):1167–1186, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/6/1167/608171.

Santos:2005:SSL

[SNM05] Aires J. P. Santos, João Nogueira, and Helder Martins. Survival of sardine larvae off the Atlantic Portuguese coast: a preliminary numerical study. ICES Journal of Marine Science, 62 (4):634–644, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/4/634/673080.

Sequeira:2009:AGB

[SNV^{+}09] Vera Sequeira, Ana Neves, Ana Rita Vieira, Ivone Figueiredo, and Leonel S. Gordo. Age and growth of bluemouth, Helicolenus dactylopterus, from the Portuguese continental slope. ICES Journal of Marine Science, 66(3):524–531, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/3/524/816243.

Sala:2007:EMQ

[SOB^{+}07] A. Sala, F. G. O’Neill, G. Buglioni, A. Lucchetti, V. Palumbo, and R. J. Fryer. Experimental method for quantifying resistance to the opening of netting panels. ICES Journal of Marine Science, 64(8):1573–1578, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/8/1573/613649.

Somerton:2004:DPC

[Som04] David A. Somerton. Do Pacific cod (Gadus macrocephalus) and walleye pollock (Theregra chalcogramma) lack a herding response to the doors, bridles, and mudclouds of survey trawls? ICES Journal of Marine Science, 61(7):1186–1189, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1186/881860.

Skreslet:2000:SSH

[SOMT00] S. Skreslet, K. Olsen, Å. Mohus, and K. S. Tande. Stage-specific habitats of Calanus finmarchicus and Calanus helgolandicus in a stratified northern Norwegian fjord. ICES Journal of Marine Science, 57(6):1656–1663, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/6/1656/768603.

Stephens:2002:LPM

[SP02] John Stephens, Jr. and Daniel Pondella, II. Larval productivity of a mature artificial reef: the ichthyoplankton of King Harbor, California, 1974–1997. ICES Journal of Marine Science, 59 (S1):S51–S58, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S51/618002.

Smith:2003:BDS

[SP03a] C. J. Smith and K.-N. Papadopoulou. Burrow density and stock size fluctuations of Nephrops norvegicus in a semi-enclosed bay. ICES Journal of Marine Science, 60(4):798–805, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/4/798/692660.

Sondre:2003:EAC

[SP03b] Aanes Sondre and Michael Pennington. On estimating the age composition of the commercial catch of Northeast Arctic cod from a sample of clusters. ICES Journal of Marine Science, 60 (2):297–303, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/2/297/625881.

Simonit:2005:IEI

[SP05] Silvio Simonit and Charles Perrings. Indirect economic indicators in bio-economic fishery models: agricultural price indicators and fish stocks in Lake Victoria. ICES Journal of Marine Science, 62(3):483–492, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/62/3/483/663558.

Shertzer:2007:DFM

[SP07a] Kyle W. Shertzer and Michael H. Prager. Delay in fishery management: diminished yield, longer rebuilding, and increased probability of stock collapse. ICES Journal of Marine Science, 64(1):149–159, January 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/64/1/149/646621.

Staal:2007:QAA

[SP07b] Johan Stål and Leif Pihl. Quantitative assessment of the area of shallow habitat for fish on the Swedish west coast. ICES Journal of Marine Science, 64(3):446–452, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/3/446/816706.

Smith:2000:IOT

[SPD00] C. J. Smith, K. N. Papadopoulou, and S. Diliberto. Impact of otter trawling on an eastern Mediterranean commercial trawl fishing ground. ICES Journal of Marine Science, 57(5):1340–1351, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/5/1340/660932.

Serra-Pereira:2008:DDD

[SPFF^{+}08] B. Serra-Pereira, I. Figueiredo, I. Farias, T. Moura, and L. S. Gordo. Description of dermal denticles from the caudal region of Raja clavata and their use for the estimation of age and growth. ICES Journal of Marine Science, 65(9):1701–1709, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/9/1701/631313.

Somarakis:2004:DEP

[SPG^{+}04] Stylianos Somarakis, Isabel Palomera, Alberto Garcia, Luis Quintanilla, Constantin Koutsikopoulos, Andrés Uriarte, and Lorenzo Motos. Daily egg production of anchovy in European waters. ICES Journal of Marine Science, 61(6):944–958, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/6/944/678267.

Sarda:2000:CDS

[SPGT00] R. Sardá, S. Pinedo, A. Gremare, and S. Taboada. Changes in the dynamics of shallow sandy-bottom assemblages due to sand extraction in the Catalan Western Mediterranean Sea. ICES Journal of Marine Science, 57(5):1446–1453, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/5/1446/660952.

Solmundsson:2005:FMI

[SPK05] Jon Solmundsson, Jonbjorn Palsson, and Hjalti Karlsson. Fidelity of mature Icelandic plaice (Pleuronectes platessa) to spawning and feeding grounds. ICES Journal of Marine Science, 62(2):189–200, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/2/189/603047.

Sainsbury:2000:DOM

[SPS00a] Keith J. Sainsbury, André E. Punt, and Anthony D. M. Smith. Design of operational management strategies for achieving fishery ecosystem objectives. ICES Journal of Marine Science, 57 (3):731–741, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/731/635993.

Swain:2000:EWT

[SPS00b] D. P. Swain, G. A. Poirier, and A. F. Sinclair. Effect of water temperature on catchability of Atlantic cod (Gadus morhua) to the bottom-trawl survey in the southern Gulf of St Lawrence. ICES Journal of Marine Science, 57(1):56–68, February 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/1/56/641147.

Storr-Paulsen:2004:SSA

[SPWHR04] Marie Storr-Paulsen, Kai Wieland, Holger Hovgård, and Hans-Joachim Rätz. Stock structure of Atlantic cod (Gadus morhua) in West Greenland waters: implications of transport and migration. ICES Journal of Marine Science, 61(6):972–982, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/6/972/679025.

Sellanes:2008:MCS

[SQN08] J. Sellanes, E. Quiroga, and C. Neira. Megafauna community structure and trophic relationships at the recently discovered Concepción Methane Seep Area, Chile, ≈ 36^{o}S. ICES Journal of Marine Science, 65(7):1102–1111, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/7/1102/644368.

Sanchez:2003:SSC

[SR03] Ricardo F. Sánchez and Paulo Relvas. Spring–summer climatological circulation in the upper layer in the region of Cape St. Vincent, Southwest Portugal. ICES Journal of Marine Science, 60(6):1232–1250, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/6/1232/652814.

Saborido-Rey:2004:AGR

[SRGC04] Fran Saborido-Rey, Dolores Garabana, and Santiago Cerviño. Age and growth of redfish (Sebastes marinus, S. mentella, and S. fasciatus) on the Flemish Cap (Northwest Atlantic). ICES Journal of Marine Science, 61(2):231–242, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/2/231/621064.

Stanton:2003:IFO

[SRJ03] Timothy K. Stanton, D. Benjamin Reeder, and J. Michael Jech. Inferring fish orientation from broadband-acoustic echoes. ICES Journal of Marine Science, 60(3):524–531, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/3/524/658787.

Shin:2005:USB

[SRJ^{+}05] Yunne-Jai Shin, Marie-Joëlle Rochet, Simon Jennings, John G. Field, and Henrik Gislason. Using size-based indicators to evaluate the ecosystem effects of fishing. ICES Journal of Marine Science, 62(3):384–396, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/62/3/384/660710.

Sabatini:2000:DPP

[SRM00] M. E. Sabatini, F. C. Ramírez, and P. Martos. Distribution pattern and population structure of Calanus australis Brodsky, 1959 over the southern Patagonian Shelf off Argentina in summer. ICES Journal of Marine Science, 57(6):1856–1866, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/6/1856/768695.

Stelzenmuller:2008:STP

[SRM08] Vanessa Stelzenmüller, Stuart I. Rogers, and Craig M. Mills. Spatio-temporal patterns of fishing pressure on UK marine landscapes, and their implications for spatial planning and management. ICES Journal of Marine Science, 65(6):1081– 1091, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/6/1081/603159.

Stares:2007:DPC

[SRMB07] J. C. Stares, R. M. Rideout, M. J. Morgan, and J. Brattey. Did population collapse influence individual fecundity of Northwest Atlantic cod? ICES Journal of Marine Science, 64(7): 1338–1347, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/7/1338/728716.

Saborido-Rey:2000:GVS

[SRN00] Fran Saborido-Rey and Kjell H. Nedreaas. Geographic variation of Sebastes mentella in the Northeast Arctic derived from a morphometric approach. ICES Journal of Marine Science, 57 (4):965–975, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/4/965/647398.

Spares:2007:IMD

[SRS^{+}07] Aaron D. Spares, Jeffery M. Reader, Michael J. W. Stokesbury, Tom McDermott, Lubomir Zikovsky, Trevor S. Avery, and Michael J. Dadswell. Inferring marine distribution of Canadian and Irish Atlantic salmon (Salmo salar L.) in the North Atlantic from tissue concentrations of bio-accumulated caesium 137. ICES Journal of Marine Science, 64(2):394–404, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/2/394/2182595.

Stefansson:2009:DPD

[SRS^{+}09] Magnús Örn Stefánsson, Jákup Reinert, orsteinn Sigurdsson, Kristján Kristinsson, Kjell Nedreaas, and Christophe Pampoulie. Depth as a potential driver of genetic structure of Sebastes mentella across the North Atlantic Ocean. ICES Journal of Marine Science, 66(4):680–690, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/4/680/604659.

Soiland:2000:VTD

[SS00] H. Søiland and M. D. Skogen. Validation of a three-dimensional biophysical model using nutrient observations in the North Sea. ICES Journal of Marine Science, 57(4):816–823, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/4/816/647339.

Suuronen:2007:RTM

[SS07] Petri Suuronen and Francesc Sardà. The role of technical measures in European fisheries management and how to make them work better. ICES Journal of Marine Science, 64(4):751–756, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/4/751/642494.

Simard:2009:DCA

[SS09] Yvan Simard and Marc Sourisseau. Diel changes in acoustic and catch estimates of krill biomass. ICES Journal of Marine Science, 66(6):1318–1325, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1318/691722.

Stenevik:2008:BVD

[SSA08] Erling Kåre Stenevik, Svein Sundby, and Ann Lisbeth Agnalt. Buoyancy and vertical distribution of Norwegian coastal cod (Gadus morhua) eggs from different areas along the coast. ICES Journal of Marine Science, 65(7):1198–1202, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/7/1198/644714.

Silva:2006:TGV

[SSC^{+}06] A. Silva, M. B. Santos, B. Caneco, G. Pestana, C. Porteiro, P. Carrera, and Y. Stratoudakis. Temporal and geographic variability of sardine maturity at length in the northeastern Atlantic and the western Mediterranean. ICES Journal of Marine Science, 63(4):663–676, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/63/4/663/692666.

Slotte:2007:SMR

[SSI07] Aril Slotte, Dankert Skagen, and Svein A. Iversen. Size of mackerel in research vessel trawls and commercial purse-seine catches: implications for acoustic estimation of biomass. ICES Journal of Marine Science, 64(5):989–994, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/5/989/642461.

Soldal:2002:RRN

[SSJL02] Aud Vold Soldal, Ingvald Svellingen, Terje Jørgensen, and Svein Løkkeborg. Rigs-to-reefs in the North Sea: hydroacoustic quantification of fish in the vicinity of a “semi-cold” platform. ICES Journal of Marine Science, 59(S1):S281–S287, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S281/617966.

Siira:2006:SWH

[SSKE06] Antti Siira, Petri Suuronen, Petri Kreivi, and Jaakko Erkinaro. Size of wild and hatchery-reared Atlantic salmon populations in the northern Baltic Sea estimated by a stratified mark-recapture method. ICES Journal of Marine Science, 63(8):1477–1487, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/8/1477/713486.

Silva:2009:GVS

[SSU^{+}09] A. Silva, D. W. Skagen, A. Uriarte, J. Massé, M. B. Santos, V. Marques, P. Carrera, P. Beillois, G. Pestana, C. Porteiro, and Y. Stratoudakis. Geographic variability of sardine dynamics in the Iberian Biscay region. ICES Journal of Marine Science, 66(3):495–508, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/3/495/814587.

Sawada:2009:TSL

[STA^{+}09] Kouichi Sawada, Hideyuki Takahashi, Koki Abe, Taro Ichii, Kazutoshi Watanabe, and Yoshimi Takao. Target-strength, length, and tilt-angle measurements of Pacific saury (Cololabis saira) and Japanese anchovy (Engraulis japonicus) using an acoustic-optical system. ICES Journal of Marine Science, 66 (6):1212–1218, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1212/692544.

Saetre:2002:FAR

[STAN02] R. Sætre, R. Toresen, and T. Anker-Nilssen. Factors affecting the recruitment variability of the Norwegian spring-spawning herring (Clupea harengus L.). ICES Journal of Marine Science, 59(4):725–736, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/4/725/676780.

Stensholt:2001:CMP

[Ste01a] Boonchai K. Stensholt. Cod migration patterns in relation to temperature: analysis of storage tag data. ICES Journal of Marine Science, 58(4):770–793, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/4/770/630200.

Stephen:2001:RRA

[Ste01b] C. Stephen. Role of risk assessment in fish health policy and management. ICES Journal of Marine Science, 58(2):374–379, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/2/374/617826.

Stenson:2002:TEI

[Ste02] G. Stenson. Is there evidence of increased pup production in northwest Atlantic harp seals, Pagophilus groenlandicus? ICES Journal of Marine Science, 59(1):81–92, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/59/1/81/
650036.

Stentiford:2008:DEE

[Ste08] Grant D. Stentiford. Diseases of the European edible crab (Cancer pagurus): a review. ICES Journal of Marine Science, 65(9): 1578–1592, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/9/1578/629544.

Sigurdsson:2006:STD

[STG06] Thorsteinn Sigurdsson, Vilhjalmur Thorsteinsson, and Leifur Gústafsson. In situ tagging of deep-sea redfish: application of an underwater, fish-tagging system. ICES Journal of Marine Science, 63(3):523–531, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/63/3/523/722766.

Suuronen:2007:FAR

[STJ^{+}07] Petri Suuronen, Vesa Tschernij, Pekka Jounela, Daniel Valentinsson, and P-O. Larsson. Factors affecting rule compliance with mesh size regulations in the Baltic cod trawl fishery. ICES Journal of Marine Science, 64(8):1603–1606, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/8/1603/613776.

Salas:2008:ATI

[STM^{+}08] Fuensanta Salas, Heliana Teixeira, Concepción Marcos, João Carlos Marques, and Angel Pérez-Ruzafa. Applicability of the trophic index TRIX in two transitional ecosystems: the Mar Menor lagoon (Spain) and the Mondego estuary (Portugal). ICES Journal of Marine Science, 65(8):1442–1448, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/8/1442/713704.

Stamatis:2006:AVN

[STMM06] Costas Stamatis, Alexander Triantafyllidis, Katerina A. Moutou, and Zissis Mamuris. Allozymic variation in Northeast Atlantic and Mediterranean populations of Norway lobster, Nephrops norvegicus. ICES Journal of Marine Science, 63(5):875–882, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/5/875/663315.

Stransky:2005:GVG

[Str05] Christoph Stransky. Geographic variation of golden redfish (Sebastes marinus) and deep-sea redfish (S. mentella) in the North Atlantic based on otolith shape analysis. ICES Journal of Marine Science, 62(8):1691–1698, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/8/1691/795877.

Sumaila:2008:FPI

[STW^{+}08] Ussif Rashid Sumaila, Louise Teh, Reg Watson, Peter Tyedmers, and Daniel Pauly. Fuel price increase, subsidies, overcapacity, and resource sustainability. ICES Journal of Marine Science, 65(6):832–840, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/6/832/602588.

Svedang:2003:IDF

[Sve03] Henrik Svedäng. The inshore demersal fish community on the Swedish Skagerrak coast: regulation by recruitment from offshore sources. ICES Journal of Marine Science, 60(1):23–31, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/1/23/611436.

Svendsen:2008:RDP

[SVRF08] Tore C. Svendsen, Katrin Vorkamp, Bent Rønsholdt, and Jens-Ole Frier. Retrospective determination of primary feeding areas of Atlantic salmon (Salmo salar) using fingerprinting of chlorinated organic contaminants. ICES Journal of Marine Science, 65(6):921–929, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/6/921/602683.

Sayer:2002:PLS

[SW02] M. D. J. Sayer and T. A. Wilding. Planning, licensing, and stakeholder consultation in an artificial reef development: the Loch Linnhe reef, a case study. ICES Journal of Marine Science, 59(S1):S178–S185, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/59/suppl/S178/617921.

Simpson:2006:ICI

[SW06a] Anne W. Simpson and Les Watling. An investigation of the cumulative impacts of shrimp trawling on mud-bottom fishing grounds in the Gulf of Maine: effects on habitat and macrofaunal community structure. ICES Journal of Marine Science, 63 (9):1616–1630, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/9/1616/698100.

Skilbrei:2006:SGS

[SW06b] Ove T. Skilbrei and Vidar Wennevik. Survival and growth of sea-ranched Atlantic salmon, Salmo salar L., treated against sea lice before release. ICES Journal of Marine Science, 63 (7):1317–1325, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/7/1317/759509.

Skilbrei:2006:UCS

[SW06c] Ove T. Skilbrei and Vidar Wennevik. The use of catch statistics to monitor the abundance of escaped farmed Atlantic salmon and rainbow trout in the sea. ICES Journal of Marine Science, 63(7):1190–1200, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/7/1190/754029.

Skaala:2006:ETG

[SWG06] Øystein Skaala, Vidar Wennevik, and Kevin A. Glover. Evidence of temporal genetic change in wild Atlantic salmon, Salmo salar L., populations affected by farm escapees. ICES Journal of Marine Science, 63(7):1224–1233, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/7/1224/755604.

Symes:2007:FMI

[Sym07] David Symes. Fisheries management and institutional reform: a European perspective. ICES Journal of Marine Science, 64 (4):779–785, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/4/779/639669.

Syrjala:2000:CUD

[Syr00] Stephen E. Syrjala. Critique on the use of the delta distribution for the analysis of trawl survey data. ICES Journal of Marine Science, 57(4):831–842, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/4/831/647345.

Stemmann:2008:GZF

[SYR^{+}08] Lars Stemmann, Marsh Youngbluth, Kevin Robert, Aino Hosia, Marc Picheral, Harriet Paterson, Frederic Ibanez, Lionel Guidi, Fabien Lombard, and Gabriel Gorsky. Global zoogeography of fragile macrozooplankton in the upper 100–1000 m inferred from the underwater video profiler. ICES Journal of Marine Science, 65(3):433–442, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/3/433/785872.

Siddeek:2007:EPM

[SZ07] M. S. M. Siddeek and J. Zheng. Evaluating the parameters of a MSY control rule for the Bristol Bay, Alaska, stock of red king crabs. ICES Journal of Marine Science, 64(5):995–1005, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/5/995/642353.

Triantafillos:2005:GEN

[TA05] Lianos Triantafillos and Mark Adams. Genetic evidence that the northern calamary, Sepioteuthis lessoniana, is a species complex in Australian waters. ICES Journal of Marine Science, 62 (8):1665–1670, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/8/1665/793881.

Tuck:2000:PBN

[TAC00] I. D. Tuck, R. J. A. Atkinson, and C. J. Chapman. Population biology of the Norway lobster, Nephrops norvegicus (L.) in the Firth of Clyde, Scotland II: fecundity and size at onset of sexual maturity. ICES Journal of Marine Science, 57(4):1227–1239, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/4/1227/647310.

Taranger:2006:CLD

[TAHK06] G. L. Taranger, L. Aardal, T. Hansen, and O. S. Kjesbu. Continuous light delays sexual maturation and increases growth of Atlantic cod (Gadus morhua L.) in sea cages. ICES Journal of Marine Science, 63(2):365–375, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/2/365/641875.

Tallack:2007:ERS

[Tal07] S. M. L. Tallack. Escape ring selectivity, bycatch, and discard survivability in the New England fishery for deep-water red crab, Chaceon quinquedens. ICES Journal of Marine Science, 64(8): 1579–1586, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/8/1579/613597.

Tande:2000:PVC

[Tan00] K. Tande. Patterns in the variations of copepod spring and summer abundance in the northeastern Norwegian Sea and the Barents Sea in cold and warm years during the 1980s and 1990s. ICES Journal of Marine Science, 57(6):1581–1591, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/6/1581/768549.

Tasker:2000:IFM

[Tas00] M. Tasker. The impacts of fishing on marine birds. ICES Journal of Marine Science, 57(3):531–547, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/3/531/635929.

Turpin:2002:PPH

[TB02] Robert K. Turpin and Stephen A. Bortone. Pre- and posthurricane assessment of artificial reefs: evidence for potential use as refugia in a fishery management strategy. ICES Journal of Marine Science, 59(S1):S74–S82, ???? 2002. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/59/suppl/S74/618009.

Tsou:2001:PMR

[TC01] T. S. Tsou and J. S. Collie. Predation-mediated recruitment in the Georges Bank fish community. ICES Journal of Marine Science, 58(5):994–1001, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/5/994/730878.

Tunez:2008:RFA

[TCC08] Juan I. Túnez, Humberto L. Cappozzo, and Marcelo H. Cassini. Regional factors associated with the distribution of South American fur seals along the Atlantic coast of South America. ICES Journal of Marine Science, 65(9):1733–1738, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/9/1733/631508.

Tett:2008:UPC

[TCM^{+}08] P. Tett, C. Carreira, D. K. Mills, S. van Leeuwen, J. Foden, E. Bresnan, and R. J. Gowen. Use of a phytoplankton community index to assess the health of coastal waters. ICES Journal of Marine Science, 65(8):1475–1482, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/8/1475/714779.

Tudela:2005:DOR

[TCP05] Sergi Tudela, Marta Coll, and Isabel Palomera. Developing an operational reference framework for fisheries management on the basis of a two-dimensional index of ecosystem impact. ICES Journal of Marine Science, 62(3):585–591, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/3/585/666660.

Tian:2009:DSS

[TCS^{+}09] Rucheng C. Tian, Changsheng Chen, Kevin D. E. Stokesbury, Brian J. Rothschild, Geoffrey W. Cowles, Qichun Xu, Song Hu, Bradley P. Harris, and Michael C. Marino. Dispersal and settlement of sea scallop larvae spawned in the fishery closed areas on Georges Bank. ICES Journal of Marine Science, 66(10):

2155–2164, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/10/2155/679059.

Tenningen:2006:LTS

[TCSW06] Eirik Tenningen, James H. Churnside, Aril Slotte, and James J. Wilson. Lidar target-strength measurements on Northeast Atlantic mackerel (Scomber scombrus). ICES Journal of Marine Science, 63(4):677–682, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/63/4/677/693027.

Tzeng:2009:MMH

[TCTC09] Chih-Hsiang Tzeng, Chih-Shin Chen, Pei-Ciao Tang, and Tai-Sheng Chiu. Microsatellite and mitochondrial haplotype differentiation in blue mackerel (Scomber australasicus) from the western North Pacific. ICES Journal of Marine Science, 66 (5):816–825, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/5/816/664298.

Tegner:2000:EEF

[TD00] M. J. Tegner and P. K. Dayton. Ecosystem effects of fishing in kelp forest communities. ICES Journal of Marine Science, 57(3):579–589, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/3/579/635936.

Tyrrell:2008:PFM

[TDE08] Megan C. Tyrrell, Michele Dionne, and Jessica A. Edgerly. Physical factors mediate effects of grazing by a non-indigenous snail species on saltmarsh cordgrass (Spartina alterniflora) in New England marshes. ICES Journal of Marine Science, 65(5):746– 752, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/5/746/712585.

Tereshchenko:2002:DPF

[Ter02] Elena S. Tereshchenko. The dynamics of population fecundity in Barents Sea capelin. ICES Journal of Marine Science, 59 (5):976–982, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/5/976/675219.

Thorley:2005:CBA

[TES^{+}05] J. L. Thorley, D. M. R. Eatherley, A. B. Stephen, I. Simpson, J. C. MacLean, and A. F. Youngson. Congruence between automatic fish counter data and rod catches of Atlantic salmon (Salmo salar) in Scottish rivers. ICES Journal of Marine Science, 62(4):808–817, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/62/4/808/677058.

Tamura:2002:GSC

[TF02] Tsutomu Tamura and Yoshihiro Fujise. Geographical and seasonal changes of the prey species of minke whale in the Northwestern Pacific. ICES Journal of Marine Science, 59(3):516–528, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/3/516/610825.

Thiele:2004:MFB

[TF04] Wilfried Thiele and Anders Fernö. Modelling of fish behaviour. ICES Journal of Marine Science, 61(7):1240–1241, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1240/883504.

Trygonis:2009:OSA

[TGS09] Vasilis Trygonis, Stratis Georgakarakos, and E. John Simmonds. An operational system for automatic school identification on multibeam sonar echoes. ICES Journal of Marine Science, 66 (5):935–949, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/5/935/664949.

Torgersen:2005:VRC

[TH05] Thomas Torgersen and Geir Huse. Variability in retention of Calanus finmarchicus in the Nordic Seas. ICES Journal of Marine Science, 62(7):1301–1309, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/7/1301/656583.

Therriault:2008:PPD

[TH08a] Thomas W. Therriault and Leif-Matthias Herborg. Predicting the potential distribution of the vase tunicate Ciona intestinalis in Canadian waters: informing a risk assessment. ICES Journal of Marine Science, 65(5):788–794, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/5/788/714324.

Therriault:2008:QBR

[TH08b] Thomas W. Therriault and Leif-Matthias Herborg. A qualitative biological risk assessment for vase tunicate Ciona intestinalis in Canadian waters: using expert knowledge. ICES Journal of Marine Science, 65(5):781–787, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/5/781/714476.

Triantafillos:2004:AIS

[TJAS04] L. Triantafillos, G. D. Jackson, M. Adams, and B. L. McGrath Steer. An allozyme investigation of the stock structure of arrow squid Nototodarus gouldi (Cephalopoda: Ommastrephidae) from Australia. ICES Journal of Marine Science, 61(5):829–835, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/5/829/867880.

Tjelmeland:2002:MUA

[Tje02] Sigurd Tjelmeland. A model for the uncertainty around the yearly trawl-acoustic estimate of biomass of Barents Sea capelin, Mallotus villosus (Müller). ICES Journal of Marine Science, 59 (5):1072–1080, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/5/1072/675153.

Totland:2009:QRS

[TJG^{+}09] Atle Totland, Geir O. Johansen, Olav R. Godø, Egil Ona, and Terje Torkelsen. Quantifying and reducing the surface blind zone and the seabed dead zone using new technology. ICES Journal of Marine Science, 66(6):1370–1376, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1370/691043.

Torgersen:2001:SSB

[TK01] T. Torgersen and S. Kaartvedt. In situ swimming behaviour of individual mesopelagic fish studied by split-beam echo target tracking. ICES Journal of Marine Science, 58(1):346–354, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/1/346/603545.

Tolonen:2003:PFN

[TK03] A. Tolonen and E. Karlsbakk. The parasite fauna of the Norwegian spring spawning herring (Clupea harengus L.). ICES Journal of Marine Science, 60(1):77–84, ???? 2003. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/60/1/77/
611453.

Tjelmeland:2005:EEA

[TL05] Sigurd Tjelmeland and Ulf Lindstrøm. An ecosystem element added to the assessment of Norwegian spring-spawning herring: implementing predation by minke whales. ICES Journal of Marine Science, 62(2):285–294, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/62/2/285/604194.

Trenkel:2004:DVT

[TLM04] V. M. Trenkel, P. Lorance, and S. Mahévas. Do visual transects provide true population density estimates for deepwater fish? ICES Journal of Marine Science, 61(7):1050–1056, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1050/877824.

Tyrrell:2008:EEP

[TLMO08] Megan C. Tyrrell, Jason S. Link, Hassan Moustahfid, and William J. Overholtz. Evaluating the effect of predation mortality on forage species population dynamics in the Northeast US continental shelf ecosystem using multispecies virtual population analysis. ICES Journal of Marine Science, 65(9):1689– 1700, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/9/1689/632822.

Tande:2000:PDC

[TM00] K. S. Tande and C. B. Miller. Population dynamics of Calanus in the North Atlantic: Results from the trans-Atlantic study of Calanus finmarchicus. ICES Journal of Marine Science, 57 (6):1527, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/6/1527/768520.

Temming:2002:EMT

[TM02] A. Temming and N. Mergardt. Estimating the mean time between meals in the field from stomach content data and gastric evacuation functions of whiting (Merlangius merlangus L.) feeding on sandeel (Ammodytes marinus Raitt). ICES Journal of Marine Science, 59(4):782–793, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/4/782/676798.

Tallack:2009:DRE

[TM09] Shelly M. L. Tallack and John W. Mandelman. Do rare-earth metals deter spiny dogfish? A feasibility study on the use of electropositive “mischmetal” to reduce the bycatch of Squalus acanthias by hook gear in the Gulf of Maine. ICES Journal of Marine Science, 66(2):315–322, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/2/315/595345.

Trenkel:2008:NFM

[TMB08] Verena M. Trenkel, Valérie Mazauric, and Laurent Berger. The new fisheries multibeam echosounder ME70: description and expected contribution to fisheries research. ICES Journal of Marine Science, 65(4):645–655, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/4/645/640328.

Tsagarakis:2008:STT

[TMG^{+}08] K. Tsagarakis, A. Machias, M. Giannoulaki, S. Somarakis, and I. Karakassis. Seasonal and temporal trends in metrics of fish community for otter-trawl discards in a Mediterranean ecosystem. ICES Journal of Marine Science, 65(4):539–550, May 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/4/539/636371.

Takagi:2004:MMF

[TMI^{+}04] Tsutomu Takagi, Yutaka Moritomi, Jyun Iwata, Hiroshi Nakamine, and Nobuo Sannomiya. Mathematical model of fish schooling behaviour in a set-net. ICES Journal of Marine Science, 61(7):1214–1223, ???? 2004. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/7/1214/882212.

Tang:2009:ATD

[TNF09] Yong Tang, Yasushi Nishimori, and Masahiko Furusawa. The average three-dimensional target strength of fish by spheroid model for sonar surveys. ICES Journal of Marine Science, 66 (6):1176–1183, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1176/692693.

Trenkel:2004:DSP

[TPRR04] Verena M. Trenkel, John K. Pinnegar, Marie-Joëlle Rochet, and Brian D. Rackham. Different surveys provide similar pictures of trends in a marine fish community but not of individual fish populations. ICES Journal of Marine Science, 61(3):351–362, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/3/351/671355.

Todd:2009:DEA

[TPT^{+}09] Victoria L. G. Todd, William D. Pearse, Nick C. Tregenza, Paul A. Lepper, and Ian B. Todd. Diel echolocation activity of harbour porpoises (Phocoena phocoena) around North Sea offshore gas installations. ICES Journal of Marine Science, 66 (4):734–745, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/4/734/603648.

Tjelmeland:2009:OHC

[TR09] Sigurd Tjelmeland and Ingolf Røttingen. Objectives and harvest control rules in the management of the fishery of Norwegian spring-spawning herring. ICES Journal of Marine Science, 66 (8):1793–1799, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/8/1793/677529.

Triantafyllou:2000:TVB

[Tri00] G. Triantafyllou. Temporal variations in benthic communities and their response to physicochemical forcing: a numerical approach. ICES Journal of Marine Science, 57(5):1507–1516, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/5/1507/660970.

Trenkel:2007:MBP

[TRM07] Verena M. Trenkel, Marie-Joëlle Rochet, and Benoît Mesnil. From model-based prescriptive advice to indicator-based interactive advice. ICES Journal of Marine Science, 64(4):768–774, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/4/768/639621.

Trenkel:2005:DEC

[TS05] Verena M. Trenkel and Hans J. Skaug. Disentangling the effects of capture efficiency and population abundance on catch data using random effects models. ICES Journal of Marine Science, 62(8):1543–1555, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/8/1543/789126.

Treasurer:2006:GSD

[TSH^{+}06] James W. Treasurer, Harald Sveier, Warren Harvey, Roddy Allen, Christopher J. Cutts, Carlos Mazorra de Quero, and Leslie Ford. Growth, survival, diet, and on-growing husbandry of haddock Melanogrammus aeglefinus in tanks and netpens. ICES Journal of Marine Science, 63(2):376–384, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/2/376/641967.

Tuya:2000:AEM

[TSK00] Fernando C. Tuya, Mark L. Soboil, and Janine Kido. An assessment of the effectiveness of Marine Protected Areas in the San Juan Islands, Washington, USA. ICES Journal of Marine Science, 57(4):1218–1226, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/4/1218/647307.

Tichy:2003:NLE

[TSK03] F. E. Tichy, H. Solli, and H. Klaveness. Non-linear effects in a 200-kHz sound beam and the consequences for target-strength measurement. ICES Journal of Marine Science, 60(3):571–574, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/3/571/659428.

Takagi:2007:EIG

[TSK07] Tsutomu Takagi, Takashi Shimizu, and Holger Korte. Evaluating the impact of gillnet ghost fishing using a computational analysis of the geometry of fishing gear. ICES Journal of Marine Science, 64(8):1517–1524, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/8/1517/612904.

Tremblay:2009:ALH

[TST^{+}09] M. John Tremblay, Stephen J. Smith, Brian J. Todd, Pierre M. Clement, and David L. McKeown. Associations of lobsters (Homarus americanus) off southwestern Nova Scotia with bottom type from images and geophysical maps. ICES Journal of Marine Science, 66(9):2060–2067, October 2009. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/9/2060/726399.

Thorne:2008:HEV

[TT08] Richard E. Thorne and Gary L. Thomas. Herring and the “Exxon Valdez” oil spill: an investigation into historical data conflicts. ICES Journal of Marine Science, 65(1):44–50, January 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/1/44/613524.

Turan:2004:SIM

[Tur04] Cemal Turan. Stock identification of Mediterranean horse mackerel (Trachurus mediterraneus) using morphometric and meristic characters. ICES Journal of Marine Science, 61(5):774–781, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/5/774/865729.

Thain:2008:CME

[TVH08] John E. Thain, A. Dick Vethaak, and Ketil Hylland. Contaminants in marine ecosystems: developing an integrated indicator framework using biological-effect techniques. ICES Journal of Marine Science, 65(8):1508–1514, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/8/1508/713411.

Tzeng:2004:PGS

[TYH04] Tzong-Der Tzeng, Shean-Yeh Yeh, and Cho-Fat Hui. Population genetic structure of the kuruma prawn (Penaeus japonicus) in East Asia inferred from mitochondrial DNA sequences. ICES Journal of Marine Science, 61(6):913–920, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/6/913/677789.

Tollefsen:2003:EDS

[TZ03] Cristina D. S. Tollefsen and Len Zedel. Evaluation of a Doppler sonar system for fisheries applications. ICES Journal of Marine Science, 60(3):692–699, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/60/3/692/661328.

Ulrich:2004:DFF

[UA04] Clara Ulrich and Bo Sølgaard Andersen. Dynamics of fisheries, and the flexibility of vessel activity in Denmark between 1989 and 2001. ICES Journal of Marine Science, 61(3):308–322, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/3/308/670985.

Ulrich:2007:TFB

[UASN07] Clara Ulrich, Bo Sølgaard Andersen, Per J. Sparre, and J. Rasmus Nielsen. TEMAS: fleet-based bio-economic simulation software to evaluate management strategies accounting for fleet behaviour. ICES Journal of Marine Science, 64(4):647–651, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/4/647/642189.

Uhlmann:2009:MBL

[UBP^{+}09] Sebastian S. Uhlmann, Matt K. Broadhurst, Brian D. Paterson, David G. Mayer, Paul Butcher, and Craig P. Brand. Mortality and blood loss by blue swimmer crabs (Portunus pelagicus) after simulated capture and discarding from gillnets. ICES Journal of Marine Science, 66(3):455–461, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/3/455/813084.

Underhill:2005:IHE

[UC05] L. G. Underhill and R. J. M. Crawford. Indexing the health of the environment for breeding seabirds in the Benguela ecosystem. ICES Journal of Marine Science, 62(3):360–365, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/3/360/659661.

Ulmestrand:2001:GNL

[UE01] Mats Ulmestrand and Håkan Eggert. Growth of Norway lobster, Nephrops norvegicus (Linnaeus 1758), in the Skagerrak, estimated from tagging experiments and length frequency data. ICES Journal of Marine Science, 58(6):1326–1334, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/6/1326/641586.

Uusitalo:2005:EAS

[UKR05] Laura Uusitalo, Sakari Kuikka, and Atso Romakkaniemi. Estimation of Atlantic salmon smolt carrying capacity of rivers using expert knowledge. ICES Journal of Marine Science, 62 (4):708–722, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/4/708/674638.

Ungfors:2009:LSG

[UMSA09] Anette Ungfors, Niall J. McKeown, Paul W. Shaw, and Carl André. Lack of spatial genetic variation in the edible crab (Cancer pagurus) in the Kattegat–Skagerrak area. ICES Journal of Marine Science, 66(3):462–469, April 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/3/462/813187.

Ungfors:2007:SME

[Ung07] Anette Ungfors. Sexual maturity of the edible crab (Cancer pagurus) in the Skagerrak and the Kattegat, based on reproductive and morphometric characters. ICES Journal of Marine Science, 64(2):318–327, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/2/318/2182562.

Uzars:2000:CGM

[UP00] D. Uzars and M. Plikshs. Cod (Gadus morhua L.) cannibalism in the Central Baltic: interannual variability and influence of recruit abundance and distribution. ICES Journal of Marine Science, 57(2):324–329, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/2/324/620441.

Ushakov:2002:BSC

[UP02] Nikolai G. Ushakov and Dmitry V. Prozorkevich. The Barents Sea capelin — a review of trophic interrelations and fisheries. ICES Journal of Marine Science, 59(5):1046–1052, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/5/1046/675139.

Urquhart:2008:PAD

[UPK^{+}08] K. Urquhart, C. C. Pert, R. Kilburn, R. J. Fryer, and I. R. Bricknell. Prevalence, abundance, and distribution of Lepeoptheirus salmonis (Krøyer, 1837) and Caligus elongatus (Nordmann, 1832) on wild sea trout Salmo trutta L. ICES Journal of Marine Science, 65(2):171–173, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/2/171/734348.

Urban-Rich:2001:SEF

[UR01] Juanita Urban-Rich. Seston effects on faecal pellet carbon concentrations from a mixed community of copepods in Balsfjord, Norway, and the Antarctic Polar Front. ICES Journal of Marine Science, 58(3):700–710, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/58/3/700/810134.

Urban-Rich:2004:EFC

[URMS04] Juanita Urban-Rich, James T. McCarty, and Mark Shailer. Effects of food concentration and diet on chromophoric dissolved organic matter accumulation and fluorescent composition during grazing experiments with the copepod Calanus finmarchicus. ICES Journal of Marine Science, 61(4):542–551, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/4/542/604202.

Uye:2000:WDC

[Uye00] S. Uye. Why does Calanus sinicus prosper in the shelf ecosystem of the Northwest Pacific Ocean? ICES Journal of Marine Science, 57(6):1850–1855, December 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/6/1850/768691.

Vella:2009:EPI

[VBF09] Prassede Vella, Robert E. Bowen, and Anamarija Frankic. An evolving protocol to identify key stakeholder-influenced indicators of coastal change: the case of marine protected areas. ICES Journal of Marine Science, 66(1):203–213, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/1/203/633754.

Vause:2007:FFF

[VBSB07] B. J. Vause, B. D. Beukers-Stewart, and A. R. Brand. Fluctuations and forecasts in the fishery for queen scallops (Aequipecten opercularis) around the Isle of Man. ICES Journal of Marine Science, 64(6):1124–1135, September 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/6/1124/614795.

Vilhjalmsson:2002:ASC

[VC02] Hjálmar Vilhjálmsson and James E. Carscadden. Assessment surveys for capelin in the Iceland–East Greenland–Jan Mayen area, 1978–2001. ICES Journal of Marine Science, 59(5): 1096–1104, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/5/1096/675167.

Vaz:2007:EEC

[VCC07] S. Vaz, A. Carpentier, and F. Coppin. Eastern English Channel fish assemblages: measuring the structuring effect of habitats on distinct sub-communities. ICES Journal of Marine Science, 64 (2):271–287, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/2/271/2182425.

vanDamme:2009:RNS

[vDBF^{+}09] Cindy J. G. van Damme, Loes J. Bolle, Clive J. Fox, Petter Fossum, Gerd Kraus, Peter Munk, Norbert Rohlf, Peter R. Witthames, and Mark Dickey-Collas. A reanalysis of North Sea plaice spawning-stock biomass using the annual egg production method. ICES Journal of Marine Science, 66(9):1999–2011, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/9/1999/725470.

vanDalfsen:2000:DRM

[vDEM^{+}00] J. A. van Dalfsen, K. Essink, H. Toxvig Madsen, J. Birklund, J. Romero, and M. Manzanera. Differential response of macrozoobenthos to marine sand extraction in the North Sea and the Western Mediterranean. ICES Journal of Marine Science, 57(5): 1439–1445, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/5/1439/660950.

vanderKooij:2007:LUP

[vdKRS^{+}07] Jeroen van der Kooij, David Righton, Espen Strand, Kathrine Michalsen, Vilhjalmur Thorsteinsson, Henrik Svedäng, Francis C. Neat, and Stefan Neuenfeldt. Life under pressure: insights from electronic data-storage tags into cod swimbladder function. ICES Journal of Marine Science, 64(7):1293–1301, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/7/1293/728121.

vanDensen:2007:ICM

[vDM07] Wim L. T. van Densen and Bonnie J. McCay. Improving communication from managers to fishers in Europe and the US. ICES Journal of Marine Science, 64(4):811–817, May 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/4/811/641368.

vanderMeer:2000:PDT

[vdMBD00] Jaap van der Meer, Jan J. Beukema, and Rob Dekker. Population dynamics of two marine polychaetes: the relative role of density dependence, predation, and winter conditions. ICES Journal of Marine Science, 57(5):1488–1494, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/5/1488/660967.

vanderVeer:2000:RFS

[vdVBMR00] Henk W. van der Veer, Rüdiger Berghahn, John M. Miller, and Adriaan D. Rijnsdorp. Recruitment in flatfish, with special emphasis on North Atlantic species: Progress made by the Flatfish Symposia. ICES Journal of Marine Science, 57(2):202–215, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/57/2/202/620407.

Vecchione:2000:IAT

[Vec00] M. Vecchione. Importance of assessing taxonomic adequacy in determining fishing effects on marine biodiversity. ICES Journal of Marine Science, 57(3):677–681, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/3/677/635956.

Velikanov:2002:SDR

[Vel02] A. Ya. Velikanov. Spatial differences in reproduction of capelin (Mallotus villosus socialis) in the coastal waters of Sakhalin. ICES Journal of Marine Science, 59(5):1011–1017, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/5/1011/675127.

Venables:2009:SSF

[VEP^{+}09] William N. Venables, Nick Ellis, André E. Punt, Catherine M. Dichmont, and Roy A. Deng. A simulation strategy for fleet dynamics in Australia’s northern prawn fishery: effort allocation at two scales. ICES Journal of Marine Science, 66(4):631–645, May 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/4/631/604796.

Vainikka:2009:TTD

[VGBH09] Anssi Vainikka, Anna Gårdmark, Barbara Bland, and Joakim Hjelm. Two- and three-dimensional maturation reaction norms for the eastern Baltic cod, Gadus morhua. ICES Journal of Marine Science, 66(2):248–257, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/2/248/594640.

Verissimo:2003:RBE

[VGF03] Ana Veríssimo, Leonel Gordo, and Ivone Figueiredo. Reproductive biology and embryonic development of Centroscymnus coelolepis in Portuguese mainland waters. ICES Journal of Marine Science, 60(6):1335–1341, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/6/1335/653948.

Valle:2008:NVV

[VH08] Sonia R. Valle and Sharon Z. Herzka. Natural variability in δ ^{18}O values of otoliths of young Pacific sardine captured in Mexican waters indicates subpopulation mixing within the first year of life. ICES Journal of Marine Science, 65(2):174–190, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/2/174/735046.

Vabo:2004:SSB

[VHF^{+}04] Rune Vabø, Geir Huse, Anders Fernö, Terje Jørgensen, Svein Løkkeborg, and Georg Skaret. Simulating search behaviour of fish towards bait. ICES Journal of Marine Science, 61(7): 1224–1232, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/7/1224/882244.

Valdes:2004:RZG

[VHI^{+}04] Luis Valdés, Roger Harris, Tsutomu Ikeda, Skip McKinnell, and William T. Peterson. The role of zooplankton in global ecosystem dynamics: Comparative studies from the world oceans. ICES Journal of Marine Science, 61(4):441–444, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/4/441/602093.

Vilhjalmsson:2002:CMV

[Vil02] Hjálmar Vilhjálmsson. Capelin (Mallotus villosus) in the Iceland–East Greenland–Jan Mayen ecosystem. ICES Journal of Marine Science, 59(5):870–883, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/5/870/675196.

Vinther:2001:AHM

[Vin01] Morten Vinther. Ad hoc multispecies VPA tuning applied for the Baltic and North Sea fish stocks. ICES Journal of Marine Science, 58(1):311–320, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/58/1/311/603536.

Velasco:2008:DAG

[VLBB08] Francisco Velasco, Jorge Landa, Joaquín Barrado, and Marian Blanco. Distribution, abundance, and growth of anglerfish (Lophius piscatorius) on the Porcupine Bank (west of Ireland). ICES Journal of Marine Science, 65(7):1316–1325, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/7/1316/646734.

Villablanca:2007:HDG

[VLJM^{+}07] Roberto Villablanca, Guillermo Luna-Jorquera, Victor H. Marín, Stefan Garthe, and Alejandro Simeone. How does a generalist seabird species use its marine habitat? The case of the kelp gull in a coastal upwelling area of the Humboldt Current. ICES Journal of Marine Science, 64(7):1348–1355, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/7/1348/728236.

Vigneau:2007:DSO

[VM07] Joël Vigneau and Stéphanie Mahévas. Detecting sampling outliers and sampling heterogeneity when catch-at-length is estimated using the ratio estimator. ICES Journal of Marine Science, 64(5):1028–1032, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/64/5/1028/643177.

Vantrepotte:2009:TVY

[VM09] V. Vantrepotte and F. Mélin. Temporal variability of 10-year global SeaWiFS time-series of phytoplankton chlorophyll a concentration. ICES Journal of Marine Science, 66(7):1547–1556, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/7/1547/658441.

Vorberg:2000:ESF

[Vor00] R. Vorberg. Effects of shrimp fisheries on reefs of Sabellaria spinulosa (Polychaeta). ICES Journal of Marine Science, 57(5): 1416–1420, October 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/5/1416/660944.

Valdes:2009:OCO

[VPC^{+}09] Luis Valdés, William Peterson, John Church, Keith Brander, and Marta Marcos. Our changing oceans: conclusions of the first International Symposium on the Effects of climate change on the world’s oceans. ICES Journal of Marine Science, 66(7): 1435–1438, August 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/7/1435/658820.

Vinther:2004:SSA

[VRP04] Morten Vinther, Stuart A. Reeves, and Kenneth R. Patterson. From single-species advice to mixed-species management: taking the next step. ICES Journal of Marine Science, 61(8): 1398–1409, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/61/8/1398/632503.

Vikebo:2005:CET

[VSÅF05] Frode Vikebø, Svein Sundby, Bjørn Ådlandsvik, and Øyvind Fiksen. The combined effect of transport and temperature on distribution and growth of larvae and pelagic juveniles of arcto-Norwegian cod. ICES Journal of Marine Science, 62(7): 1375–1386, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/7/1375/658647.

Vitale:2006:HAI

[VSC06] F. Vitale, H. Svedäng, and M. Cardinale. Histological analysis invalidates macroscopically determined maturity ogives of the Kattegat cod (Gadus morhua) and suggests new proxies for estimating maturity status of individual fish. ICES Journal of Marine Science, 63(3):485–492, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/3/485/721544.

Voss:2007:VDB

[VSS07] Rüdiger Voss, Jörn O. Schmidt, and Dietrich Schnack. Vertical distribution of Baltic sprat larvae: changes in patterns of diel migration? ICES Journal of Marine Science, 64(5):956–962, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/5/956/642119.

Vuorinen:2002:PPP

[Vuo02] P. Vuorinen. PCDD, PCDF, PCB and thiamine in Baltic herring (Clupea harengus L.) and sprat [Sprattus sprattus (L.)] as a background to the M74 syndrome of Baltic salmon (Salmo salar L.). ICES Journal of Marine Science, 59(3):480–496, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/3/480/610819.

Walline:2007:GSE

[Wal07] Paul D. Walline. Geostatistical simulations of eastern Bering Sea walleye pollock spatial distributions, to estimate sampling precision. ICES Journal of Marine Science, 64(3):559–569, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/3/559/813271.

Ward:2001:EMM

[War01a] T. Ward. Effects of the 1995 and 1998 mass mortality events on the spawning biomass of sardine, Sardinops sagax, in South Australian waters. ICES Journal of Marine Science, 58(4):865–875, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/4/865/630220.

Warren:2001:SMA

[War01b] J. Warren. In situ measurements of acoustic target strengths of gas-bearing siphonophores. ICES Journal of Marine Science, 58(4):740–749, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/4/740/630195.

Watkins:2002:VAT

[WB02] J. L. Watkins and A. S. Brierley. Verification of the acoustic techniques used to identify Antarctic krill. ICES Journal of Marine Science, 59(6):1326–1336, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/6/1326/608200.

Walsh:2004:FBE

[WB04] Stephen J. Walsh and Åsmund Bjordal. Fish behaviour in exploited ecosystems: Introduction. ICES Journal of Marine Science, 61(7):1030–1035, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/7/1030/877523.

Witting:2005:AGW

[WB05] Lars Witting and Erik W. Born. An assessment of Greenland walrus populations. ICES Journal of Marine Science, 62 (2):266–284, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/2/266/603938.

Walker:2006:MIE

[WBC^{+}06] Alan M. Walker, Malcolm C. M. Beveridge, Walter Crozier, Niall Ó Maoiléidigh, and Nigel Milner. Monitoring the incidence of escaped farmed Atlantic salmon, Salmo salar L., in rivers and fisheries of the United Kingdom and Ireland: current progress and recommendations for future programmes. ICES Journal of Marine Science, 63(7):1201–1210, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/7/1201/754103.

Wahle:2008:NAD

[WBC^{+}08] Richard A. Wahle, Charlene E. Bergeron, Antonie S. Chute, Larry D. Jacobson, and Yong Chen. The Northwest Atlantic deep-sea red crab (Chaceon quinquedens) population before and after the onset of harvesting. ICES Journal of Marine Science, 65(6):862–872, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/65/6/862/600399.

Whoriskey:2006:MSS

[WBD^{+}06] Frederick G. Whoriskey, Paul Brooking, Gino Doucette, Stephen Tinker, and Jonathan W. Carr. Movements and survival of sonically tagged farmed Atlantic salmon released in Cobscook Bay, Maine, USA. ICES Journal of Marine Science, 63(7):1218–1223, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/7/1218/755290.

Williams:2009:RCW

[WBK09a] Alan Williams, Nicholas J. Bax, and Rudy J. Kloser. Remarks on “Comment on: Williams et al. (2009) Australia’s deep-water reserve network: implications of false homogeneity for classifying abiotic surrogates of biodiversity, ICES Journal of Marine Science, 66: 214–224” by Peter T. Harris, Andrew D. Heap, Tara J. Anderson, and Brendan Brooke. ICES Journal of Marine Science, 66(10):2086–2088, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/

66/10/2086/683355. See [WBK^{+}09b, HHAB09].

Williams:2009:ADW

[WBK^{+}09b] Alan Williams, Nicholas J. Bax, Rudy J. Kloser, Franziska Althaus, Bruce Barker, and Gordon Keith. Australia’s deep-water reserve network: implications of false homogeneity for classifying abiotic surrogates of biodiversity. ICES Journal of Marine Science, 66(1):214–224, January 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/1/214/634078. See

comment [HHAB09] and remarks [WBK09a].

Wagler:2009:FWS

[WBV09] H. Wagler, R. Berghahn, and R. Vorberg. The fishery for whiteweed, Sertularia cupressina (Cnidaria, Hydrozoa), in the Wadden Sea, Germany: history and anthropogenic effects. ICES Journal of Marine Science, 66(10):2116–2120, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/10/2116/681694.

Whoriskey:2001:RTA

[WC01] F. G. Whoriskey and J. W. Carr. Returns of transplanted adult, escaped, cultured Atlantic salmon to the Magaguadavic River, New Brunswick. ICES Journal of Marine Science, 58 (2):504–509, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/2/504/617862.

Walters:2005:PEI

[WCMK05] Carl J. Walters, Villy Christensen, Steven J. Martell, and James F. Kitchell. Possible ecosystem impacts of applying MSY policies from single-species assessment. ICES Journal of Marine Science, 62(3):558–568, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/62/3/558/665660.

Wells:2008:HUE

[WCP08] R. J. David Wells, James H. Cowan, Jr., and William F. Patterson III. Habitat use and the effect of shrimp trawling on fish and invertebrate communities over the northern Gulf of Mexico continental shelf. ICES Journal of Marine Science, 65(9): 1610–1619, December 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/9/1610/630202.

West:2009:POM

[WDRP09] Christopher D. West, Calvin Dytham, David Righton, and Jonathan W. Pitchford. Preventing overexploitation of migratory fish stocks: the efficacy of marine protected areas in a stochastic environment. ICES Journal of Marine Science, 66 (9):1919–1930, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/66/9/1919/722834.

Weinberg:2005:BSD

[Wei05] James R. Weinberg. Bathymetric shift in the distribution of Atlantic surfclams: response to warmer ocean temperature. ICES Journal of Marine Science, 62(7):1444–1453, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
62/7/1444/659651.

Wespestad:2000:RBC

[WFIM00] Vidar G. Wespestad, Lowell W. Fritz, W. James Ingraham, and Bernard A. Megrey. On relationships between cannibalism, climate variability, physical transport, and recruitment success of Bering Sea walleye pollock (Theragra chalcogramma). ICES Journal of Marine Science, 57(2):272–278, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/2/272/620427.

Weichler:2004:SDH

[WGLJM04] Tanja Weichler, Stefan Garthe, Guillermo Luna-Jorquera, and Julio Moraga. Seabird distribution on the Humboldt Current in northern Chile in relation to hydrography, productivity, and fisheries. ICES Journal of Marine Science, 61(1):148–154, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/1/148/700850.

Walsh:2004:FBR

[WGM04] Stephen J. Walsh, Olav Rune Godø, and Kathrine Michalsen. Fish behaviour relevant to fish catchability. ICES Journal of Marine Science, 61(7):1238–1239, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/61/7/1238/883499.

Was:2008:EPS

[WGMM08] Anna Was, Elizabeth Gosling, Karen McCrann, and Jarle Mork. Evidence for population structuring of blue whiting (Micromesistius poutassou) in the Northeast Atlantic. ICES Journal of Marine Science, 65(2):216–225, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/2/216/733944.

Williams:2008:DCA

[WHA08] Tom Williams, Kristin Helle, and Michaela Aschan. The distribution of chondrichthyans along the northern coast of Norway. ICES Journal of Marine Science, 65(7):1161–1174, October 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/7/1161/644798.

Walter:2007:CEA

[WHG07] John F. Walter III, John M. Hoenig, and Todd Gedamke. Correcting for effective area fished in fishery-dependent depletion estimates of abundance and capture efficiency. ICES Journal of Marine Science, 64(9):1760–1771, December 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/9/1760/782346.

Wildish:2001:CEM

[WHP01] D. J. Wildish, B. T. Hargrave, and G. Pohle. Cost-effective monitoring of organic enrichment resulting from salmon mariculture. ICES Journal of Marine Science, 58(2):469–476, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/2/469/617851.

Wiggert:2008:MSD

[WHP08] Jerry D. Wiggert, Eileen E. Hofmann, and Gustav-Adolf Paffenhöfer. A modelling study of developmental stage and environmental variability effects on copepod foraging. ICES Journal of Marine Science, 65(3):379–398, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/65/3/379/784056.

Wieland:2005:CRG

[Wie05] Kai Wieland. Changes in recruitment, growth, and stock size of northern shrimp (Pandalus borealis) at West Greenland: temperature and density-dependent effects at released predation pressure. ICES Journal of Marine Science, 62(7):1454–1462, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/7/1454/659990.

Winter:2007:SEM

[WJB07] Hendrik V. Winter, Henrice M. Jansen, and André W. Breukelaar. Silver eel mortality during downstream migration in the River Meuse, from a population perspective. ICES Journal of Marine Science, 64(7):1444–1449, October 2007. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/7/1444/728881.

Wieland:2000:CTS

[WJTH00] Kai Wieland, Astrid Jarre-Teichmann, and Katarzyna Horbowa. Changes in the timing of spawning of Baltic cod: possible causes and implications for recruitment. ICES Journal of Marine Science, 57(2):452–464, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/2/452/620493.

Waldron:2001:AVH

[WK01] Miranda E. Waldron and Michael Kerstan. Age validation in horse mackerel (Trachurus trachurus) otoliths. ICES Journal of Marine Science, 58(4):806–813, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/58/4/806/630207.

Wilson:2002:RHK

[WLK02] Keith D. P. Wilson, Albert W. Y. Leung, and Robin Kennish. Restoration of Hong Kong fisheries through deployment of artificial reefs in marine protected areas. ICES Journal of Marine Science, 59(S1):S157–S163, ???? 2002. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S157/617914.

Walmsley:2007:MSA

[WLS07a] Sarah A. Walmsley, Rob W. Leslie, and Warwick H. H. Sauer. Managing South Africa’s trawl bycatch. ICES Journal of Marine Science, 64(2):405–412, March 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/2/405/2182239.

Westerberg:2007:SEM

[WLS07b] Håkan Westerberg, Ingvar Lagenfelt, and Henrik Svedäng. Silver eel migration behaviour in the Baltic. ICES Journal of Marine Science, 64(7):1457–1462, October 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/7/1457/726626.

Willis:2001:MHR

[WM01] T. J. Willis and R. B. Millar. Modified hooks reduce incidental mortality of snapper (Pagrus auratus: Sparidae) in the New Zealand commercial longline fishery. ICES Journal of Marine Science, 58(4):830–841, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/58/4/830/630215.

Walsh:2004:ONB

[WM04] Stephen J. Walsh and M. Joanne Morgan. Observations of natural behaviour of yellowtail flounder derived from data storage tags. ICES Journal of Marine Science, 61(7):1151–1156, ???? 2004. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/61/7/1151/880972.

Wilhelmsson:2006:IOW

[WMÖ06] Dan Wilhelmsson, Torleif Malm, and Marcus C. Öhman. The influence of offshore windpower on demersal fish. ICES Journal of Marine Science, 63(5):775–784, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/5/775/661096.

Wanzenbock:2003:QAH

[WMS^{+}03] Josef Wanzenböck, Thomas Mehner, Michael Schulz, Hubert Gassner, and Ian J. Winfield. Quality assurance of hydroacoustic surveys: the repeatability of fish-abundance and biomass estimates in lakes within and between hydroacoustic systems. ICES Journal of Marine Science, 60(3):486–492, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/3/486/658138.

Wang:2003:STP

[WPB^{+}03] Jianjun Wang, Graham J. Pierce, Peter R. Boyle, Vincent Denis, Jean-Paul Robin, and Jose M. Bellido. Spatial and temporal patterns of cuttlefish (Sepia officinalis) abundance and environmental influences — a case study using trawl fishery data in French Atlantic coastal, English Channel, and adjacent waters. ICES Journal of Marine Science, 60(5):1149–1158, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/5/1149/768290.

Witherell:2000:EBA

[WPF00] David Witherell, Clarence Pautzke, and David Fluharty. An ecosystem-based approach for Alaska groundfish fisheries. ICES Journal of Marine Science, 57(3):771–777, June 2000. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
57/3/771/636011.

Weber:2009:CAO

[WPJ09] Thomas C. Weber, Héctor Peña, and J. Michael Jech. Consecutive acoustic observations of an Atlantic herring school in the Northwest Atlantic. ICES Journal of Marine Science, 66 (6):1270–1277, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/6/1270/694116.

Wheeler:2009:TCM

[WPM^{+}09] J. P. Wheeler, C. F. Purchase, P. D. M. Macdonald, R. Fill, L. Jacks, H. Wang, and C. Ye. Temporal changes in maturation, mean length-at-age, and condition of spring-spawning Atlantic herring (Clupea harengus) in Newfoundland waters. ICES Journal of Marine Science, 66(8):1800–1807, September 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/8/1800/674659.

Woillez:2007:ICS

[WPR^{+}07] Mathieu Woillez, Jean-Charles Poulard, Jacques Rivoirard, Pierre Petitgas, and Nicolas Bez. Indices for capturing spatial patterns and their evolution in time, with application to European hake (Merluccius merluccius) in the Bay of Biscay. ICES Journal of Marine Science, 64(3):537–550, April 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
64/3/537/818031.

Woillez:2009:EUA

[WRF09] Mathieu Woillez, Jacques Rivoirard, and Paul G. Fernandes. Evaluating the uncertainty of abundance estimates from acoustic surveys using geostatistical simulations. ICES Journal of Marine Science, 66(6):1377–1383, July 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/6/1377/697513.

Wilding:2002:EAR

[WS02a] Thomas A. Wilding and Martin D. J. Sayer. Evaluating artificial reef performance: approaches to pre- and post-deployment research. ICES Journal of Marine Science, 59(S1):S222–S230, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL https://academic.oup.com/icesjms/
article/59/suppl/S222/617945.

Wilding:2002:PCP

[WS02b] Thomas A. Wilding and Martin D. J. Sayer. The physical and chemical performance of artificial reef blocks made using quarry by-products. ICES Journal of Marine Science, 59(S1): S250–S257, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S250/617952.

Winter:2006:ICD

[WS06] Andreas G. Winter and Gordon L. Swartzman. Interannual changes in distribution of age-0 walleye pollock near the Pribilof Islands, Alaska, with reference to the prediction of pollock year-class strength. ICES Journal of Marine Science, 63 (6):1118–1135, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/63/6/1118/616641.

Watson:2006:CPT

[WSC^{+}06] J. T. Watson, S. Sales, G. Cumming, S. D. Fitzsimmons, J. Walden, G. Arthur, S. Saravanan, and L. A. McEvoy. Comparison of performance of two size groups of farmed cod (Gadus morhua L.) juveniles following transfer to sea cages. ICES Journal of Marine Science, 63(2):340–345, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
63/2/340/641195.

Shyue:2002:ITC

[wScY02] Shiahn wern Shyue and Kuang che Yang. Investigating terrain changes around artificial reefs by using a multi-beam echosounder. ICES Journal of Marine Science, 59(S1):S338– S342, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/59/suppl/S338/617989.

Workman:2002:HPS

[WSFH02] Ian Workman, Arvind Shah, Dan Foster, and Bret Hataway. Habitat preferences and site fidelity of juvenile red snapper (Lutjanus campechanus). ICES Journal of Marine Science, 59 (S1):S43–S50, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/59/suppl/S43/618001.

Wilding:2003:FAP

[WSP03] Thomas A. Wilding, Martin D. J. Sayer, and Paul G. Provost. Factors affecting the performance of the acoustic ground discrimination system RoxAnn^{T M}. ICES Journal of Marine Science, 60(6):1373–1380, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/60/6/1373/654649.

Warren:2003:IBP

[WSWS03] Joseph D. Warren, Timothy K. Stanton, Peter H. Wiebe, and Harvey E. Seim. Inference of biological and physical parameters in an internal wave using multiple-frequency, acoustic-scattering data. ICES Journal of Marine Science, 60(5):1033–1046, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/5/1033/764580.

Wasmund:2003:PTB

[WU03] N. Wasmund and S. Uhlig. Phytoplankton trends in the Baltic Sea. ICES Journal of Marine Science, 60(2):177–186, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/2/177/623558.

Woll:2006:SVA

[WvdMF06] Astrid K. Woll, Gro I. van der Meeren, and Inge Fossen. Spatial variation in abundance and catch composition of Cancer pagurus in Norwegian waters: biological reasoning and implications for assessment. ICES Journal of Marine Science, 63(3):421–433, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/3/421/719284.

Wlodarczyk:2001:MDV

[WW01] E. Wlodarczyk and R. Wenne. Mitochondrial DNA variation in sea trout from coastal rivers in the southern Baltic region. ICES Journal of Marine Science, 58(1):230–237, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
58/1/230/603523.

Winger:2007:FEM

[WW07] Paul D. Winger and Philip J. Walsh. The feasibility of escape mechanisms in conical snow crab traps. ICES Journal of Marine Science, 64(8):1587–1591, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/64/8/1587/613808.

Wright:2002:VFA

[WWGG02] P. J. Wright, D. A. Woodroffe, F. M. Gibb, and J. D. M. Gordon. Verification of first annulus formation in the illicia and otoliths of white anglerfish, Lophius piscatorius using otolith microstructure. ICES Journal of Marine Science, 59(3):587–593, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/3/587/610841.

Winger:2004:STH

[WWHB04] Paul D. Winger, Stephen J. Walsh, Pingguo He, and Joseph A. Brown. Simulating trawl herding in flatfish: the role of fish length in behaviour and swimming characteristics. ICES Journal of Marine Science, 61(7):1179–1185, ???? 2004. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
61/7/1179/881770.

Whomersley:2008:BID

[WWR^{+}08] Paul Whomersley, Suzanne Ware, Hubert L. Rees, Claire Mason, Thi Bolam, Mark Huxham, and Helen Bates. Biological indicators of disturbance at a dredged-material disposal site in Liverpool Bay, UK: an assessment using time-series data. ICES Journal of Marine Science, 65(8):1414–1420, November 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/8/1414/713928.

Woodd-Walker:2003:ISO

[WWWB03] Rachel S. Woodd-Walker, Jonathan L. Watkins, and Andrew S. Brierley. Identification of Southern Ocean acoustic targets using aggregation backscatter and shape characteristics. ICES Journal of Marine Science, 60(3):641–649, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/3/641/660369.

Wang:2009:EDS

[WYM09] You-Gan Wang, Yimin Ye, and David A. Milton. Efficient designs for sampling and subsampling in fisheries research based on ranked sets. ICES Journal of Marine Science, 66(5):928–934, June 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/66/5/928/664248.

Wiafe:2008:ICC

[WYMF08] George Wiafe, Hawa B. Yaqub, Martin A. Mensah, and Christopher L. J. Frid. Impact of climate change on long-term zooplankton biomass in the upwelling region of the Gulf of Guinea. ICES Journal of Marine Science, 65(3):318–324, April 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/3/318/788889.

Xia:2005:ESC

[XZW05] Junhong Xia, Jinsong Zheng, and Ding Wang. Ex situ conservation status of an endangered Yangtze finless porpoise population (Neophocaena phocaenoides asiaeorientalis) as measured from microsatellites and mtDNA diversity. ICES Journal of Marine Science, 62(8):1711–1716, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/8/1711/796465.

Ye:2003:SVG

[YBF^{+}03] Y. Ye, J. M. Bishop, N. Fetta, E. Abdulqader, J. Al-Mohammadi, A. H. Alsaffar, and S. Almatar. Spatial variation in growth of the green tiger prawn (Penaeus semisulcatus) along the coastal waters of Kuwait, eastern Saudi Arabia, Bahrain, and Qatar. ICES Journal of Marine Science, 60(4):806–817, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/4/806/692816.

Yu:2007:RFE

[YCCH07] Congda Yu, Zhihai Chen, Lianyuan Chen, and Pingguo He. The rise and fall of electrical beam trawling for shrimp in the East China Sea: technology, fishery, and conservation implications. ICES Journal of Marine Science, 64(8):1592–1597, November 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/64/8/1592/613921.

Ye:2000:RRS

[Ye00] Yimin Ye. Is recruitment related to spawning stock in penaeid shrimp fisheries? ICES Journal of Marine Science, 57(4): 1103–1109, August 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/4/1103/647276.

Yemane:2005:EEF

[YFL05] Dawit Yemane, John G. Field, and Rob W. Leslie. Exploring the effects of fishing on fish assemblages using Abundance Biomass Comparison (ABC) curves. ICES Journal of Marine Science, 62(3):374–379, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/3/374/660355.

Yamada:2008:WEG

[YG08] Sylvia Behrens Yamada and Graham E. Gillespie. Will the European green crab (Carcinus maenas) persist in the Pacific Northwest? ICES Journal of Marine Science, 65(5):725–729, July 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/65/5/725/712074.

Yaragina:2000:TII

[YM00] N. A. Yaragina and C. T. Marshall. Trophic influences on interannual and seasonal variation in the liver condition index of Northeast Arctic cod (Gadus morhua). ICES Journal of Marine Science, 57(1):42–55, February 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/57/1/42/641144.

Yeung:2008:UAB

[YM08] Cynthia Yeung and Robert A. McConnaughey. Using acoustic backscatter from a sidescan sonar to explain fish and invertebrate distributions: a case study in Bristol Bay, Alaska. ICES Journal of Marine Science, 65(2):242–254, March 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
65/2/242/737090.

Youngson:2002:RCT

[YMF02] A. F. Youngson, J. C. MacLean, and R. J. Fryer. Rod catch trends for early-running MSW salmon in Scottish rivers (1952– 1997): divergence among stock components. ICES Journal of Marine Science, 59(4):836–849, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/59/4/836/676820.

Yndestad:2001:ENI

[Ynd01] H. Yndestad. Earth nutation influence on Northeast Arctic cod management. ICES Journal of Marine Science, 58(4):799–805, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/4/799/630205.

Yndestad:2003:CLT

[Ynd03] Harald Yndestad. The code of the long-term biomass cycles in the Barents Sea. ICES Journal of Marine Science, 60(6): 1251–1264, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/60/6/1251/653039.

Yndestad:2006:ILN

[Ynd06] Harald Yndestad. The influence of the lunar nodal cycle on Arctic climate. ICES Journal of Marine Science, 63(3):401–420, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/3/401/718604.

Y:2005:TIE

[YNX^{+}05] P. C. U. R. Y., L. S. H. A. N. N. O. N., J. R. O. U. X., G. D. A. S. K. A. L. O. V., A. J. A. R. R. E., C. M. O. L. O. N. E. Y., and D. P. A. U. L. Y. Trophodynamic indicators for an ecosystem approach to fisheries. ICES Journal of Marine Science, 62 (3):430–442, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/62/3/430/661904.

Yndestad:2002:SDB

[YS02] Harald Yndestad and Anne Stene. System dynamics of the Barents Sea capelin. ICES Journal of Marine Science, 59(6): 1155–1166, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/59/6/1155/608170.

Yemane:2009:EEM

[YSF09] Dawit Yemane, Yunne-Jai Shin, and John G. Field. Exploring the effect of marine protected areas on the dynamics of fish communities in the southern Benguela: an individual-based modelling approach. ICES Journal of Marine Science, 66(2): 378–387, March 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/66/2/378/593802.

Yasuma:2003:TSM

[YSO^{+}03] Hiroki Yasuma, Kouichi Sawada, Tatsuki Ohshima, Kazushi Miyashita, and Ichiro Aoki. Target strength of mesopelagic lanternfishes (family Myctophidae) based on swimbladder morphology. ICES Journal of Marine Science, 60(3):584–591, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/3/584/659568.

Yasuma:2006:TSL

[YTS^{+}06] Hiroki Yasuma, Yoshimi Takao, Kouichi Sawada, Kazushi Miyashita, and Ichiro Aoki. Target strength of the lanternfish, Stenobrachius leucopsarus (family Myctophidae), a fish without an airbladder, measured in the Bering Sea. ICES Journal of Marine Science, 63(4):683–692, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/4/683/693140.

Yoneda:2005:ETF

[YW05] Michio Yoneda and Peter J. Wright. Effect of temperature and food availability on reproductive investment of first-time spawning male Atlantic cod, Gadus morhua. ICES Journal of Marine Science, 62(7):1387–1393, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/62/7/1387/658752.

Yoneda:2002:AGJ

[YYY^{+}02] Michio Yoneda, Shunji Yamasaki, Keisuke Yamamoto, Hiroshi Horikawa, and Michiya Matsuyama. Age and growth of John Dory, Zeus faber (Linnaeus, 1758), in the East China Sea. ICES Journal of Marine Science, 59(4):749–756, ???? 2002.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
59/4/749/676786.

Zerbi:2001:GJT

[ZAJ01] A. Zerbi, C. Aliaume, and J.-C. Joyeux. Growth of juvenile tarpon in Puerto Rican Estuaries. ICES Journal of Marine Science, 58(1):87–95, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/58/1/87/603556.

Zhang:2006:PGS

[ZCH06] Junbin Zhang, Zeping Cai, and Liangmin Huang. Population genetic structure of crimson snapper Lutjanus erythropterus in East Asia, revealed by analysis of the mitochondrial control region. ICES Journal of Marine Science, 63(4):693–704, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/63/4/693/693353.

Zedel:2009:EFW

[ZCR09] Len Zedel and Francis-Yan Cyr-Racine. Extracting fish and water velocity from Doppler profiler data. ICES Journal of Marine Science, 66(9):1846–1852, October 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/66/9/1846/725323.

Zuur:2001:MSS

[ZFFT01] G. Zuur, R. J. Fryer, R. S. T. Ferro, and T. Tokai. Modelling the size selectivities of a trawl codend and an associated square mesh panel. ICES Journal of Marine Science, 58(3):657–671, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/58/3/657/810116.

Zhao:2006:STS

[Zha06] Xianyong Zhao. In situ target-strength measurement of young hairtail (Trichiurus haumela) in the Yellow Sea. ICES Journal of Marine Science, 63(1):46–51, ???? 2006. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/63/1/46/625216.

Zitko:2001:ACM

[Zit01] V. Zitko. Analytical chemistry in monitoring the effects of aquaculture: one laboratory’s perspective. ICES Journal of Marine Science, 58(2):486–491, ???? 2001. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/58/2/486/617855.

Zheng:2000:RPA

[ZK00] Jie Zheng and Gordon H. Kruse. Recruitment patterns of Alaskan crabs in relation to decadal shifts in climate and physical oceanography. ICES Journal of Marine Science, 57(2): 438–451, April 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/57/2/438/620488.

Zedel:2003:ADC

[ZKP03] Len Zedel, Tor Knutsen, and Ranjan Patro. Acoustic Doppler current profiler observations of herring movement. ICES Journal of Marine Science, 60(4):846–859, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/60/4/846/693863.

Zwolinski:2007:DVV

[ZMM^{+}07] Juan Zwolinski, Alexandre Morais, Vitor Marques, Yorgos Stratoudakis, and Paul G. Fernandes. Diel variation in the vertical distribution and schooling behaviour of sardine (Sardina pilchardus) off Portugal. ICES Journal of Marine Science, 64 (5):963–972, July 2007. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/64/5/963/642736.

Zalmon:2002:ERA

[ZNGF02] Ilana R. Zalmon, Ronaldo Novelli, Marcelo P. Gomes, and Vicente V. Faria. Experimental results of an artificial reef programme on the Brazilian coast north of Rio de Janeiro. ICES Journal of Marine Science, 59(S1):S83–S87, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL https://academic.oup.com/icesjms/article/
59/suppl/S83/618011.

Zhao:2003:ECM

[ZO03] Xianyong Zhao and Egil Ona. Estimation and compensation models for the shadowing effect in dense fish aggregations.

ICES Journal of Marine Science, 60(1):155–163, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
60/1/155/611430.

Zarraonaindia:2009:MVE

[ZPI^{+}09] Iratxe Zarraonaindia, Miguel Angel Pardo, Mikel Iriondo, Carmen Manzano, and Andone Estonba. Microsatellite variability in European anchovy (Engraulis encrasicolus) calls for further investigation of its genetic structure and biogeography. ICES Journal of Marine Science, 66(10):2176–2182, December 2009. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/
66/10/2176/679546.

Zedel:2005:FBO

[ZPK05] Len Zedel, Ranjan Patro, and Tor Knutsen. Fish behaviour and orientation-dependent backscatter in acoustic Doppler profiler data. ICES Journal of Marine Science, 62(6):1191–1201, ???? 2005. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/62/6/1191/618475.

Zheng:2002:DNA

[ZPRJ02] X. Zheng, G. J. Pierce, D. G. Reid, and I. T. Jolliffe. Does the North Atlantic current affect spatial distribution of whiting? Testing environmental hypotheses using statistical and GIS techniques. ICES Journal of Marine Science, 59(2):239–253, ???? 2002. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/59/2/239/619632.

Zwanenburg:2000:EFD

[Zwa00] K. C. T. Zwanenburg. The effects of fishing on demersal fish communities of the Scotian Shelf. ICES Journal of Marine Science, 57(3):503–509, June 2000. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/57/3/503/635920.

Zhao:2008:DDT

[ZWD08] Xianyong Zhao, Yong Wang, and Fangqun Dai. Depthdependent target strength of anchovy (Engraulis japonicus) measured in situ. ICES Journal of Marine Science, 65(6):

882–888, September 2008. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/65/6/882/600071.

Zimmermann:2003:IIP

[ZWW^{+}03] Mark Zimmermann, Mark E. Wilkins, Kenneth L. Weinberg, Robert R. Lauth, and Franklin R. Shaw. Influence of improved performance monitoring on the consistency of a bottom trawl survey. ICES Journal of Marine Science, 60(4):818–826, ???? 2003. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/60/4/818/693150.